A Complete Bibliography of Publications in Annals of the Institute of Statistical Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

26 April 2024 Version 1.35

Title word cross-reference

(0, θ) [969]. (k, t) [673, 651]. 0 [1407, 1426]. 1 [1407, 1426]. 1/p [460]. 2 [1840, 31]. 2^{7} [777]. 2^{l} [837]. 2^{m} [1453, 862, 1018, 1361, 1708, 400, 837, 597]. 2^{m_{1}+m_{2}} [1376]. 2^{r} [1361]. 2l + 1 [1453, 862, 1018]. 2 × 2 [1301]. 2 × 2 × 2 [1150, 1150]. 3 [48, 2752]. 3^{n} [838]. 3 × 3 [1030]. [θ, θ + 1) [1112, 1027]. 0[Hebrew letter tsade] [557]. _{1} [722]. _{2}F_{1} [791]. _{V C} [463]. a [1541]. α [2287, 2388, 2189, 2280, 2576, 2903]. AR(p) [2933, 993]. β [746]. C [3223, 854, 216, 2537, 3285, 1150, 402]. c + 1 [216]. C_{p} [1945]. χ^{2} [101, 782, 468]. D [2137, 1713, 2512, 2349, 457, 2167]. D_{n}

[1167]. δ [3198]. E [2537, 1347, 1489, 1432]. ϵ ; [465, 550, 1171, 2445]. ϵ_{n} [2482]. EX(p) [993]. ; F ; [1958, 1940, 2596, 1644, 2086, 468, 3195, 449, ; 748, 1323, 308, 451, 978, 1670, 1715, 2072]. ; f^{−2} [225]. F_{1} [722]. \frac{1}{2} ∑ \frac{∂^{2}}{∂x^{i2}} + ∑ b^{i}(x)\frac{∂}{∂x^{i}} ; [228]. g [3033]. Γ [127]. H [866]. H_{M} [463]. ; H_{MV C} [463]. I [2032, 2369, 516]. ∞ [2107]. ; J [2528]. K [1261, 608, 583, 3007, 897, 1287, ; 1314, 1525, 1841, 1907, 1958, 2210, 581, 2036, ; 1370, 3242, 1262, 1327, 2643, 868, 1397, 1427, ; 1833, 1983, 365, 2434, 2361, 2780, 248, 3286, ; 1951, 525, 1326, 1884, 3195, 1027, 1740, 1494, ; 509, 720, 1780, 2228, 176, 664, 1097, 983, ; 2067, 1068, 1188]. k(≥ 2) [1912]. k^{n} [604]. ; k ≥ 3 [1068]. k ≥ v [1489]. L ; [1245, 1626, 866, 1817, 1570, 2361]. L^{1} ; [3423]. L^{2} [3119]. L^{p} [787]. L_{0} [3136]. L_{1}

[1575, 1910, 1750, 2006, 2192, 2364, 3289, 3288]. L_{2} [2394, 2521, 2192]. L_{mvc} [1031]. L_{p} [1192, 1245, 3181, 3181]. L_{q} ≃ α · ρ^{β}/(1 − ρ) [929]. L_{V C} [1110]. Λ [2770]. LR [3029]. M [2515, 1684, 3247, 2083, 2621, 2622, 1626, 1604, 2425, 3351, 2544, 2586, 2836, 3329, 2691, 1746, 1979, 2297, 3208, 3336, 1517, 1597, 2323, 2481, 1293, 3100, 2779, 1838, 449, 748, 2771]. m^{2} [974]. R [3177]. R^{+} [3317]. M [2673]. MM [2758]. μ ± kδ [159]. μ ± kσ [132]. MV [1347, 1432]. N [852, 1398, 281, 1958, 2643, 2328, 2481, 1983, 605, 1070, 3195, 1102, 1740, 764, 1479, 597, 1652, 2228, 2597, 2985]. n ≤ 10 [1167]. N ≤ 42 [777]. Ω = O(n) [666, 999]. ω^{2} [2004]. orthomin(k) [1633]. P [2166, 3056, 1203, 1258, 3168, 1927, 798, 3291]. P (X < Y) [486, 901, 2139]. p∗ [2159]. ψ [1385, 3046]. q [2069]. R [300, 2530, 1951, 29, 36]. R^{d} [1701]. R^{m} [1824]. r × c [757]. S [2811, 2758]. S^{M} [711, 624]. S_{1} [765]. S_{1}S_{2}^{−1} [939, 1054]. S_{2}^{−1} [765]. S_{h}S_{e}^{−1} [940]. σ [899]. T [547, 1238, 487, 2493, 1239, 1348, 1141, 1324]. T^{∗} [1269]. T^{2} [1064, 121]. T_{0}^{2} [807, 578, 677, 1107, 1063]. T_{3} [374, 621]. T_{m} [778]. θ [1152]. θ(1 − p) [875]. θ(p) [875]. U [2948, 659, 2552, 3424, 2460, 366, 2915, 2536, 1465, 2828, 787, 3091, 683]. u(x)C(θ) exp(−x/θ) [1727]. V [777, 128, 2915, 2536, 548, 1708, 2685]. φ [1086]. X [901, 745]. X^{2} [2155]. x′V^{−1}x [1042]. Y [901]. Z [300, 2116, 101, 1590]. Z_{+} [2126, 2658]. || || [111].

-ary [605, 1070, 1102, 764]. -class [868, 720, 664, 1097, 621]. -component [1833]. -connections [1541, 2576]. -consistency [1192]. -convergence [1245, 787]. -correct [1385]. -cube [852]. -decomposability [2673]. -dependence [1838, 2771]. -Dimensional [2069, 1840, 31, 48, 2597, 281]. -Dissimilarity [978, 2072]. -distance [1750]. -distributed [127]. -distribution [1644]. -distributions [468]. -divergence [1670, 1715, 3223]. -entropy [465, 550]. -equivalence [1171, 1152]. -Estimates [2297, 2323, 1245, 2621, 2622, 1979]. -estimation [1684, 3247, 2544, 2836]. -estimator [1604, 3351, 1597]. -estimators [2515, 1945, 1626, 2425, 2586, 3329, 2691, 1746, 3208, 3336, 1517]. -extended [1027]. -family [1727]. -Function [2528]. -Functional [2083]. -goodness-of-fit [3119]. -grouped [1293]. -HNBUE [1326]. -in-a-row [1427]. -Markov [2445]. -means [3007]. -method [547]. -minimisers [2482]. -Mixing [2189, 2280]. -Monotonicity [2287]. -neighborhood [3198]. -norm [3181, 1575, 3181]. -optimal [1086, 3285, 2512, 2349, 2137, 1713, 2537, 1347]. -optimality [1432]. -order [1203, 1261, 1262, 1327]. -orthogonal [2685]. -Out-Of- [2228, 1958, 2643, 1983, 3195, 1740]. -overlapping [2361]. -parameter [1188]. -paths [2811]. -point [983, 3423]. -prior [3033]. -Projection [2032, 2369]. -records [2434]. -regression [2758]. -risk [1750]. -sample [854, 897, 3242, 509, 402]. -sets [673, 651]. -splines [3056]. -squares [1323]. -stable [2903]. -state [1068]. -static [748]. -Statistic [2004, 2948, 2596]. -statistics [659, 2552, 807, 3424, 2460, 2915, 1465, 449, 2828, 1817, 1570, 787, 121, 457, 3091]. -Sufficient [2166]. -tensors [2752]. -test [487, 308, 451]. -tests [1940]. -th [1258, 2530, 3286, 516, 3100]. -transforms [1590]. -type [2493, 778]. -unimodality [2388]. -valued [2126]. -values [3168, 1927, 3291]. -way [1398]. -weak [3046].

1 [314, 327, 1205]. 1/p [1335]. 1987/88 [1665].

2 [249]. 2020 [3353, 3356, 3354]. 2nd [1646]. 2th [1646].

3/2th [1646].

43-3-01 [1676].

A. [841]. abelian [836, 1008, 2269]. aberration [2745, 3339]. ABIC [1649]. ABS [1675]. abscissa [18]. Absence [2579]. Absolute [2061, 31, 48, 1723, 1653, 3405]. absolutely [1081, 1082, 2460, 853]. absorbing [814]. Abstract [2269]. Accelerated [3073, 1894, 3076, 3067]. Acceptance [180, 2706]. according [1678]. accumulated [603]. accumulation [116]. Accuracy [2049, 2391, 3054, 1813, 2621, 2622, 3235, 3358, 2953]. Accurate [2615, 1212, 2683]. Acknowledgement [1810, 1886, 2001, 2277, 2410, 2461, 1010, 640]. Acknowledgment [1757]. action [118, 789, 967]. Actions [2012]. Active [2311]. acts [3251]. actuarial [2725]. adaptation [3036, 2649]. adapted [2479]. Adaptive [3366, 884, 2346, 2146, 1881, 2544, 2284, 2292, 3153, 3145, 2792, 2380, 1147, 1011, 2957, 3171, 3170, 3281, 3237, 2913, 3332, 2863, 3429]. adaptive-to-model [3171, 3170]. Addendum [711]. addition [3190]. Additional [3124, 1133, 1198, 1500]. Additive [2776, 3066, 3379, 3212, 2784, 3325, 3257, 494, 2488, 3116, 3277, 351, 3189, 1230, 2638, 932, 2692, 3026]. Additivity [2283, 2693, 3011]. adequacy [1547, 2921]. adequate [1114, 1195]. adjacent [2454]. Adjusted [2080, 2399, 2990, 3145, 3058, 1983, 2765, 941, 2913]. Adjustment [1158]. Adjustments [2816]. Admissibility [1623, 1807, 1681, 1241, 585, 2097, 1896, 2108, 408, 2472, 1625, 1674, 1162]. Admissible [1572, 2631, 1224, 1209]. admitting [398]. Advection [2161]. Affine [2240, 2655]. Affinity [2186, 1081, 1082, 1166, 978, 718, 441, 641, 818]. after [1271, 1771, 3281, 3380, 1088]. aftershock [2718]. Against [2019, 1131, 1440, 1939, 2965, 665, 681, 1179, 713, 1431]. age [1493, 1946]. age-period-cohort [1493]. aggregates [1464]. aging [1992, 1925]. agricultural [3434, 41]. AIC [966, 2514, 2563, 2031, 3032, 985, 3185]. aided [3064]. AISM [1676]. Akaike [1267, 3355, 1268, 3353, 3356, 3354, 2710, 1162]. Aki [1439]. Alexander [2931]. algebra [2912]. algebraic [2742, 2749]. Algorithm [2175, 1709, 2440, 2568, 2405, 2570, 3047, 3103]. Algorithms [2706, 1507, 2714, 2874, 693, 729]. Alias [1018]. Aliasing [2026, 776, 3138]. aligned [1291, 1686]. Allocation [2172, 302, 1101, 396, 3359, 480, 851, 1168, 1252, 827]. allocations [3205, 867]. almost [1655, 1128]. alternating [2874]. Alternative [2266, 1295, 1681, 1531, 3191, 1927, 1064, 1273, 1342, 1071, 1431]. Alternatives [2019, 2272, 1468, 3119, 2510, 717, 1939, 665, 681, 762, 819, 874, 1179, 1392, 1606, 1567]. alternatives-I [665]. alternatives-II [681]. ambiguous [1224]. among [1543, 1632, 759, 697, 1573]. amount [1582]. ample [2743]. amplitude [2448]. analogue [633, 741]. Analogues [2288, 1558, 1273, 1313]. analyses [2615, 510, 326, 1377]. Analysis [2079, 2558, 861, 2195, 2424, 2216, 2217, 1675, 2099, 2226, 2787, 2290, 2229, 507, 1401, 2207, 2214, 985, 2300, 2063, 1443, 2297, 2768, 2180, 653, 172, 195, 817, 966, 2677, 343, 1015, 1134, 1450, 244, 1267, 1268, 972, 333, 3238, 2428, 2840, 2712, 1417, 2750, 2447, 3410, 1485, 1198, 1631, 3065, 636, 926, 1491, 3387, 2888, 2841, 3304, 1666, 1354, 2570, 438, 166, 375, 598, 1531, 3333, 2846, 2760, 744, 843, 3290, 3195, 2970, 55, 1459, 260, 546, 3433, 307, 308, 2920, 885, 4, 191, 264, 1437, 422, 2755, 273, 1770, 163, 998, 443, 927, 722, 791, 3392, 3297]. analysis [1418, 1313, 2930, 1536, 991, 1415, 1979, 1474, 1588, 2958, 1792, 2794, 2927, 2844, 3067, 2703]. analytic [1255]. analytical [1]. ANCOVA [3188].

Anderson [971]. Angular [2253, 2652]. angular-linear [2652]. animal [1044]. Annals [249]. Announcement [1663]. anomalous [3309]. ANOVA [2596, 362, 3081, 2559, 2637, 794]. Antedependent [2121]. anti [3103]. anti-ridge [3103]. Antipersistent [2298]. Antithetic [3074]. Any [2276, 3267, 1701, 923]. appearances [3099]. applicability [2395]. Application [3434, 1491, 2157, 1169, 2056, 1195, 2059, 1474, 2269, 810, 1245, 172, 195, 817, 272, 2730, 688, 475, 835, 1134, 1265, 2407, 3074, 2385, 419, 2351, 692, 1386, 3257, 1799, 254, 258, 1649, 1320, 1521, 1787, 1175, 3362, 1730, 469, 1185, 2908, 3246, 1766, 3318, 1388, 264, 134, 1242, 1954, 1734, 1908, 3369, 1148, 182, 3435, 2927]. Applications [2203, 1055, 2257, 2170, 2143, 2273, 2121, 2105, 2178, 2068, 1036, 1081, 133, 224, 2926, 2739, 2618, 3119, 1635, 3284, 1992, 2869, 1917, 2976, 892, 1358, 3353, 2421, 2428, 578, 1722, 1964, 3157, 3146, 2621, 2622, 1612, 2369, 829, 1661, 2460, 3354, 577, 113, 982, 317, 864, 2780, 1877, 619, 616, 1269, 2744, 2979, 2915, 2839, 2812, 441, 641, 2919, 931, 744, 843, 1811, 1905, 836, 1866, 795, 4, 1580, 444, 482, 3355, 2716, 1132, 1552, 1798, 2469, 3245, 2810]. applied [107, 401, 1170, 1419]. Approach [2325, 2579, 2197, 2290, 2271, 2264, 2241, 2292, 2218, 2543, 2859, 803, 2948, 2994, 1600, 2651, 2849, 3280, 3230, 500, 2851, 2756, 2512, 2458, 3229, 3302, 2451, 3337, 3114, 1306, 1149, 2951, 2608, 774, 867, 1545, 824, 826, 1214, 3147, 1014, 1218, 1296, 1624, 1961, 3432, 1186, 3099, 1941, 3120, 3121, 2764, 1049, 1310, 3116, 1527, 2475, 1294, 3183, 2639, 2464, 1323, 2820, 1502, 191, 1580, 849, 1678, 443, 927, 2940, 1418, 2716, 3322, 2827, 1891, 2749, 1464, 2961, 3331, 3153]. approaches [2377, 2927]. Approximate [2979, 1863, 1764, 3064, 2159, 2142, 1751, 1769, 750, 1298, 1365, 998, 132]. Approximated [1514, 2995].

Approximately [896, 419, 278]. Approximating [1538, 3424, 1942]. Approximation [2307, 2134, 2133, 2169, 2334, 2318, 80, 1507, 1888, 1144, 3000, 856, 1673, 209, 108, 1237, 844, 635, 2605, 3093, 732, 929, 959, 1223, 1638, 356, 610, 3136, 702, 754]. Approximations [2524, 2170, 1730, 950, 869, 1343, 2594, 1140, 2565, 1509, 2621, 2622, 1840, 2888, 2841, 1175, 1108, 3168, 701, 1259, 976, 1588, 1301]. arbitrarily [1022]. arbitrary [2391, 2901]. ARC [478]. arch [2506, 2584]. Archimedean [2593]. arcsinhed [3016]. Area [1932, 2230, 2818]. Area-interaction [1932]. areas [899, 132, 159]. argument [2939, 892, 722, 791, 1798]. arguments [1358, 1055, 3347]. ARIMA [1330]. arising [500, 1408, 2743]. arithmetic [3365]. ARMA [1888, 2890, 3049, 1929, 2478, 1533]. armed [2945]. array [106]. arrays [723, 2799, 3377, 840, 1324, 1694, 837]. arrival [1664]. arrivals [1758]. articles [775]. ARX [1906]. ary [605, 1070, 1102, 764]. ASN [1256]. ASP [2477]. Aspect [2032]. Aspects [2132, 3124, 2968, 94]. assay [378]. assays [960]. assembling [136]. Assessing [2099, 1743, 2030, 1752, 1389]. Assessment [2208, 2439]. assigned [598]. assisted [3208]. associate [1455, 1121, 796, 979, 279, 449, 748, 669]. Associated [2122, 2327, 2010, 2155, 1909, 2980, 2358, 2419, 2684, 1217, 3167, 2839, 1980, 1901, 1160, 1161, 1406, 2685]. Association [435, 1069, 979, 1308, 733, 598, 1251, 2681, 3128, 576, 2797]. assuming [1003, 1443]. Assumption [2070]. assumptions [3359, 3002]. astronomy [3157]. Asymmetric [2116, 2299, 2340, 3382, 1919, 1632, 2489, 3019, 2474]. asymmetrical [1308, 584]. asymmetry [3097]. Asymptotic [1131, 1156, 1142, 1106, 1575, 1016, 1174, 1719, 2739, 1115, 3160, 2248, 3319, 2589, 3315, 2539, 2778, 939, 972, 970, 2700, 1481, 2165, 3335, 3374, 1306, 2887, 2007, 2150, 667, 763, 808, 834, 1264, 2186, 1092, 2014, 2390, 1429, 1977, 792, 2596, 1107, 1305, 2630, 1499, 317, 498, 731, 1874, 940, 1355, 301, 1626, 2305, 2311, 2847, 410, 45, 3150, 3202, 953, 1196, 1480, 2312, 866, 3167, 2425, 2041, 2785, 2246, 2783, 2082, 2045, 2586, 3204, 762, 819, 1277, 2444, 2304, 2270, 2708, 2534, 2663, 2357, 1551, 828, 3237, 2916, 822, 1817, 938, 1717, 2445, 1442, 1546, 1247]. Asymptotic [2022, 2313, 883, 2940, 666, 722, 791, 1391, 1050, 1307, 2093, 543, 3390, 1798, 3336, 1286, 1597, 2247, 880, 2705, 882, 1017, 1402, 1482, 1646, 3052, 2391, 2955, 1910, 868, 905, 647, 765, 918, 419, 2428, 974, 1523, 846, 2951, 952, 1412, 1929, 3182, 488, 670, 695, 677, 721, 761, 951, 971, 1456, 385, 615, 678, 865, 1236, 767, 1787, 1861, 1604, 1229, 749, 3233, 999, 1378, 1374, 1905, 988, 353, 748, 527, 930, 1259, 1172, 2691, 379, 1629, 2729, 320, 1449, 2843, 1392, 716, 1162, 742, 1777, 1800, 995, 1930, 409, 540, 592, 1000, 1210]. Asymptotically [2234, 1987, 2203, 1317, 574, 701, 2880, 700, 1058, 1606, 906, 225, 2354, 738, 739, 608, 583, 2828, 2767]. Asymptotics [1950, 2963, 3360, 2303, 2112, 2452, 2201, 2225, 2426, 2600, 1574, 2504, 3352, 2407, 2458, 3205, 2888, 2841, 2999, 2915, 2758, 2363]. Asymtotic [807]. asynchronously [3022]. atomic [2636]. atomic-bomb [2636]. attachment [2650, 3367]. attaining [2372]. attitude [631]. attitudes [227, 222]. Attraction [2051, 593]. attributable [1150]. Auflösung [393]. Author [3356]. auto [867]. auto-correlations [867]. autocorrelation [1131, 955, 993]. autocorrelations [1990]. autocovariance [262]. Automatic [3407, 2371]. Autoregression [2064, 2895, 1395, 1098]. Autoregressive [643, 674, 2160, 2041, 2058, 2605, 2296, 882, 538, 556, 817, 1203, 1776, 996, 884, 1987, 2950, 2944, 3216, 2840, 2479, 1264, 1616, 1360, 957, 956, 1874, 2689, 3397, 2470, 3235, 3358, 1372, 1560, 3057, 1157, 2403, 989, 1474, 3411, 1849, 3369, 1286, 3426]. Auxiliary [2543, 2005, 1177, 1206, 3206]. available [1500]. Average [1735, 817, 1776, 2745, 1630, 1545, 3295, 3150, 3368, 1157, 162, 1263, 1067, 3426]. averaged [558]. averages [1122]. averaging [3186, 2380, 3221, 3220, 975, 3303, 3400, 3420]. axioms [124].

B [2398, 3125]. B-spline [2398]. B-splines [3125]. Back [2335]. backcross [25]. backfitted [2692]. backfitting [3066, 3379]. background [3261]. Bahadur [3365, 2060, 1967, 1559, 1039]. balance [823, 1250]. Balanced [837, 1592, 1534, 777, 605, 723, 3424, 812, 840, 1070, 1085, 1084, 1120, 279, 284, 507, 1376, 1453, 449, 1102, 764, 1324, 862, 1018, 400]. balancing [56]. balking [434]. ball [2808]. ball-bin [2808]. balls [1462]. Ban [459]. band [2848]. Bandit [2182, 2945]. bands [1383, 3328, 1581, 2441, 3211, 3117, 2368, 3042]. Bandwidth [2106, 2821, 2413, 1977, 2626, 3244, 2824, 3429]. bandwidths [1736]. Bartlett [1948, 3269, 298, 762, 1989, 1552]. barycentres [2507]. Based [2561, 2560, 2166, 2236, 2174, 2184, 2308, 2237, 2192, 2526, 2285, 2170, 2233, 2292, 2173, 2315, 2225, 2025, 2557, 2324, 2705, 802, 1154, 1350, 1575, 1016, 649, 2990, 3193, 3382, 1394, 1672, 2914, 659, 3174, 595, 3360, 3319, 3188, 2796, 2373, 1029, 3092, 1547, 1725, 684, 2451, 3114, 3031, 3413, 1759, 1177, 1503, 2666, 2568, 2541, 2608, 3205, 2380, 2454, 695, 697, 3267, 2582, 948, 925, 3147, 767, 248, 1522, 3210, 3432, 770, 3371, 2789, 2574, 866, 3199, 1183, 2495, 2583, 3008, 3368, 3197, 3078, 3223, 900, 1152, 103, 111, 131, 2360, 1469, 1465, 3033, 3324, 462]. based [942, 630, 3289, 3288, 236, 3246, 1073, 1766, 2670, 2989, 518, 1020, 2899, 3433, 1598, 3154, 2893, 1169, 1985, 379, 2682, 2509, 1686, 1780, 83, 176, 2940, 285, 456, 436, 471, 539, 613, 3297, 2406, 3407, 2720, 609, 572, 614, 2437, 3299, 687, 3343, 2954, 3030]. baseline [2651, 3064, 2554]. bases [3403, 3107, 2751, 3018]. Basic [2169, 154, 980]. basis [2676, 2624, 2727, 2699, 49, 1248, 1548, 2908, 2412]. Basu [2079]. bathtub [2811]. Bayes [607, 536, 919, 2339, 1258, 1164, 1650, 500, 2869, 3352, 3251, 523, 1783, 638, 2057, 2404, 3102, 3050, 1471, 2115, 1926, 2811, 2120, 483, 1014, 1624, 1962, 1996, 2140, 908, 395, 1322, 1255, 1702, 703, 1191, 2048, 1527, 1743, 889, 432, 3290, 909, 2284, 1057, 872, 849, 1303, 1727, 2313, 1572, 784, 1028, 1050, 1113, 804, 1528, 1893, 324, 1862, 573, 544, 1338, 1514]. Bayes-type [3290]. Bayesian [3226, 3225, 1420, 3227, 2590, 2079, 924, 966, 2677, 2651, 1742, 2352, 1635, 2558, 1450, 1610, 785, 1051, 3315, 2144, 1725, 2840, 454, 1417, 1416, 2216, 2447, 2217, 2666, 2274, 499, 786, 2380, 633, 774, 867, 2850, 1649, 1218, 1296, 1639, 3256, 1186, 2574, 2713, 1805, 2207, 2957, 1843, 2285, 1049, 1310, 1859, 1906, 1537, 2004, 1863, 3005, 3068, 3204, 704, 2969, 3033, 1993, 1377, 2970, 1827, 1609, 2208, 2947, 2438, 1244, 2920, 1329, 1502, 1206, 1580, 2300, 2721, 1678, 466, 582, 2860, 2637, 1221, 1844, 2638, 741, 891, 2716, 3224, 2958, 1792, 2863, 1946, 1711]. be [1583]. Bechhofer [1052]. become [2435]. beflanking [724]. Behavior [2130, 2231, 1142, 1115, 557, 589, 632, 3319, 3335, 3377, 2738, 1604, 1229, 889, 919, 1817, 1717, 1464, 1597]. behaviour [3133, 2365, 1852, 988, 880]. Behrens [2948, 28, 2069]. belief [1266]. Benford [1280]. bent [3085]. Berkson [3361]. Bernoulli [2000, 466]. Bernstein [982, 2881, 2795]. Berry [2054, 3217, 1908]. Berry-Esséen [2054]. Bessel [531, 2187]. Best [2018, 253, 2120, 2230, 3061, 2097, 647, 526, 2826, 1077, 2893, 1727]. best-subset [2893]. bestness [526]. Beta [2182, 2318, 835, 1134, 1612, 1229, 2908, 1233, 2417, 1850, 2508]. better [1261, 1262, 1327, 1325]. Between [2550, 2049, 2059, 1081, 1082, 1182, 591, 1790, 1115, 1778, 3166, 2500, 1509, 3314, 1964, 2186, 467, 773, 731, 147, 1492, 1626, 1701, 1037, 1410, 2719, 91, 20, 1045, 1023, 958, 1185, 3281, 2670, 1057, 2276, 1963, 863, 83, 56, 543, 3106]. Betweeness [1295]. Betweenness [980, 1007]. betweennesses [1007, 1295]. Bézier [2375]. Bhattacharyya [1352, 2372]. bi [3351]. bi-level [3351]. Bias [2952, 2193, 2722, 915, 3273, 3206, 2418, 2200, 473, 66, 2765, 1744, 1560, 3333, 2717, 720, 1800, 1286, 2902]. bias-adjusted [2765]. Bias-corrected [2952, 3273]. Biased [2518, 735, 2441, 2451, 3065, 485, 631, 3283, 3034, 2639, 3010, 1978, 2728]. biasedness [713]. BIB [1454]. BIC [1669]. biclusters [3412]. big [3366]. bilateral [448]. bilinear [1753, 3]. bin [2808, 3407]. Binary [2136, 2036, 2799, 2230, 2302, 1314, 2980, 2628, 1738, 1218, 1248, 2409, 1821, 1329, 1420, 1406, 2747, 2797]. Binomial [2210, 2524, 2008, 805, 847, 2565, 1004, 316, 2485, 2361, 230, 248, 2748, 418, 255, 288, 534, 1183, 3406, 950, 3172, 174, 1366, 3194, 2606, 849, 320, 848, 134, 176, 1572, 1074, 267, 1850, 2852]. biological [1474]. biomechanical [3435]. biophysics [385, 323]. birth [1561]. birth-and-death [1561]. birthday [1730]. bisection [1406]. Bivariate [2079, 2104, 1111, 2125, 2248, 309, 2593, 2245, 2780, 2123, 2619, 2192, 2306, 217, 1902, 2344, 973, 2978, 268, 1836, 2627, 1408, 3279, 3031, 2439, 1731, 2875, 3377, 755, 1461, 1617, 574, 1478, 2636, 942, 1211, 452, 2987, 2403, 391, 752, 1213, 328, 1019, 2509]. black [110]. Blackwell [873]. Blind [2336]. Block [2079, 2195, 1361, 1592, 861, 1534, 1362, 605, 1738, 812, 1489, 1038, 1070, 1085, 1084, 1120, 279, 284, 1432, 3024, 260, 449, 822, 863, 1324, 561, 1606]. blocked [3059]. blocking [1116]. blocks [604, 1432, 547]. blockwise [2424, 3369]. Blomqvist [2508]. blue [1077]. blurring [2998]. BLUS [926]. bodies [2654]. bomb [2636]. Bonferroni [1652]. Bonferroni-type [1652]. Boolean [3319]. Boosting [2724, 2655, 2818]. Bootstrap [1654, 2003, 1630, 2303, 2027, 2413, 3155, 1628, 2264, 2164, 2374, 1598, 2737, 1629, 3396, 2034, 2391, 1950, 2356, 2458, 2481, 2480, 2621, 2622, 1975, 1784, 2722, 2915, 1993, 2639, 2989, 1682, 2721, 3380, 1892]. Bootstrapped [1383]. Bootstrapping [2136, 2950, 3184, 2031, 3053, 1785, 1501, 1862, 1384]. Bootstraps [2004]. both [1101, 113]. bottles [3138]. Bouman [385]. Bound [2305, 1352, 1448, 1166, 419, 908, 112, 283, 624, 711, 1242, 367, 2372, 2474]. boundaries [2654]. Boundary [2407, 2093, 2200, 3095, 351, 3274, 2981]. Bounded [2370, 2074, 93, 321, 235, 2653, 2669, 3172, 3207, 2751, 104, 165, 1849, 324]. bounded-risk [3172]. Boundedness [565]. Bounding [2065, 2006]. Bounds [2221, 2150, 2521, 1673, 205, 1063, 2050, 261, 2149, 2022, 2932, 1992, 2421, 2424, 1412, 1750, 486, 901, 1968, 1120, 418, 271, 376, 616, 1653, 3096, 534, 2764, 931, 2642, 1811, 1211, 2357, 320, 1449, 214, 3018, 3217]. Box [856, 1771, 1995]. branching [3124, 502, 726, 3015, 1475]. Breakdown [2231, 2370]. breaking [3335]. Bregman [2568, 2885]. Bridge [2309, 3095, 2407, 3316, 3199]. bridge-type [3316]. Broken [3332]. Brownian [3095, 2352, 2407, 2385, 2514, 2563, 1759, 433, 2995, 199, 2309]. Broyden [1607]. Broyden-like [1607]. Buckley [2604]. Build [961]. Build-up [961]. Bumps [2138]. Burr [380]. burst [1758]. busy [414].

C. [841]. Calculation [1030, 533]. Calibration [2924, 2241, 1411, 1374, 2383, 2362, 1946]. calibrations [3195]. Cancellation [482]. cancer [1792]. Canonical [2450, 2099, 2209, 1873, 2428, 1198, 1183, 1109, 3310]. Cantelli [429, 397]. capacity [1204]. capital [331, 116]. caps [1508, 2435]. capture [245, 2958]. capture-recapture [245, 2958]. CAR [2920]. carbon [110]. Carlo [2079, 107, 115, 3074, 2301, 326, 2220, 2380, 2949, 2968, 199, 2972, 1609, 2218]. CARMA [2976, 3150, 2222]. carried [3287]. Case [2037, 2335, 2139, 2157, 1402, 2891, 1751, 3242, 1367, 2456, 3319, 3268, 974, 3252, 1808, 1782, 3065, 689, 634, 676, 1107, 208, 1014, 292, 703, 1762, 1822, 828, 1054, 663, 1727, 134, 657, 707, 728, 747, 105, 104, 1407, 1426, 995, 827, 2604, 1188, 1000, 1514]. case-control [3065]. case-III [747]. Cases [2014, 2093, 1156, 1352, 1646, 1187, 749, 2605, 954, 1050, 1895]. Categorical [2316, 800, 3277, 1443]. categories [690, 958, 1022]. Cauchy [3365, 3216, 2174, 2475, 167]. causal [3414, 1185]. cause [2855, 3067]. cause-of-death [2855]. caused [1292]. Cautionary [2191]. cdf [1175, 2856]. cell [1321, 1504, 639, 3018]. cells [670]. cement [1491]. Censored [2533, 2234, 2073, 2582, 2264, 2280, 2249, 2106, 2232, 2555, 2034, 2265, 2250, 2247, 1998, 459, 3048, 2964, 2842, 3301, 3152, 2373, 243, 2829, 3165, 122, 2906, 2894, 2734, 2785, 3034, 3414, 749, 1283, 2374, 2393, 3154, 1928, 2987, 895, 2682, 2830, 830, 2804, 2803, 2940, 3332, 612, 2467, 1718, 2632, 1862, 3320, 2838, 2633, 2810]. Censoring [2165, 2312, 2285, 2251, 2702, 2618, 2695, 3242, 3236, 2416, 2451, 1997, 1244, 2735, 3077, 3343]. Censorship [2033, 1581, 2641]. center [1106]. centered [1945]. Central [2489, 105, 3240, 101, 3031, 2394, 676, 516, 2904, 679, 140, 546, 799, 104]. centrality [298]. centre [3254]. Century [338, 352]. Certain [2210, 913, 2132, 294, 372, 2074, 123, 149, 126, 1804, 1578, 918, 1956, 846, 1736, 1100, 636, 67, 1393, 385, 53, 940, 645, 1070, 278, 665, 681, 269, 2, 22, 432, 1986, 1173, 622, 405, 587, 176, 474, 265]. Chain [2079, 2307, 2211, 2067, 2090, 2923, 2949, 1941, 2619, 626, 640, 1920]. Chains [2175, 2010, 2490, 814, 319, 2912, 2430, 162, 1087, 1068]. challenges [3353, 3354, 3355]. chance [1065]. Change [2203, 2053, 2585, 2134, 2253, 2526, 1925, 3032, 2314, 3158, 3163, 3357, 2807, 3017, 3252, 2494, 631, 1801, 1962, 3362, 3129, 2384, 2495, 3057, 3430, 2880, 1785, 3433, 870, 3113, 2418, 1431, 3085]. Change-Point [2203, 2134, 2253, 3032, 3163, 3252, 1801, 3362, 1785, 3433, 1431]. Change-Points [2053, 2807]. change-set [3129]. change-Supplement [631]. changepoint [2677, 3293, 1178]. Changes [1266, 2021, 1158, 1639]. channels [2380]. character [115, 221, 824, 826, 542]. characterisation [2433]. Characteristic [2325, 2188, 2174, 2305, 484, 555, 588, 765, 376, 478, 516, 853, 943, 2475, 91, 518, 1054, 1995, 372, 706, 1215, 354, 899]. Characteristics [2074, 1667, 1665, 1596, 611]. Characterization [2039, 2385, 2081, 2431, 832, 2151, 2342, 1522, 1070, 1085, 2333, 1435, 718, 1436, 1463, 2024, 1138, 293, 981, 672, 1600, 1386, 1270, 1117, 1126, 2454, 1360, 210, 274, 280, 164, 1454, 57, 78, 1569, 1183, 1422, 1811, 2432, 1986, 1885, 563, 1901, 328, 256, 1047]. Characterizations [1216, 893, 2386, 3347, 1311, 1853, 1985, 1295, 3284, 1722, 879, 1956, 1429, 1217, 1542, 1653, 1691, 1125, 1922, 1060, 1137, 2412]. Characterizing [1433, 917, 3348, 369, 714, 1478]. Charier [1771]. chart [1089]. charts [2494]. Chebyshev [348, 1652]. Chebyshev-type [1652]. check [3025]. Checking [2204, 2282, 2921, 3003, 2639]. chemical [1250]. Chernoff [1080]. chi [61, 773, 2979, 950, 1867, 168, 998, 1391]. chi-square [61, 773, 2979, 950, 1867, 168, 998, 1391]. chirp [2444]. choice [3388, 3238, 1881, 1628, 626, 640, 1387, 891, 124]. Cholesky [3218]. Choosing [2053]. Circle [2349, 3095, 2999, 944, 3244]. circuits [3401]. Circular [2281, 2939, 388, 580, 2484, 3027, 668, 3024]. circulas [3027]. city [26]. Class [2039, 2552, 2539, 2593, 2314, 2093, 2250, 1594, 810, 2511, 2777, 1455, 1672, 532, 868, 2355, 1163, 1526, 1135, 2712, 2883, 1675, 738, 739, 829, 488, 2429, 821, 1493, 654, 1837, 3027, 278, 2574, 1317, 3181, 1537, 2918, 1469, 1436, 920, 890, 968, 2395, 3098, 2828, 1817, 308, 451, 1128, 720, 3132, 669, 379, 402, 793, 621, 2682, 372, 664, 830, 1097, 1446, 2876, 59, 1209, 2824, 1066, 458, 1643, 1567, 2833, 1951]. class-convergences [59]. Classes [2179, 2922, 1121, 796, 979, 279, 1914, 395, 449, 748, 33]. Classical [846, 2757, 2604]. Classification [2551, 2549, 464, 21, 523, 607, 2866, 759, 1599, 135, 150, 346, 1014, 1398, 300, 908, 131, 958, 872, 1252, 784, 3436]. classified [1556, 985]. classifier [2990, 3193]. classifiers [3371]. Classifying [2080, 2676]. clinical [1474]. close [3405]. Closed [1948]. closeness [1981]. Closer [1555, 1259]. Cluster [2121, 3230, 1267, 1268, 3178, 2876, 3392, 2977]. Clustered [2136, 3363, 3201, 2797]. Clustering [1359, 3289, 3288, 3132, 1693, 1797, 3007]. clusters [3373, 3377]. coarse [2796]. code [1399]. Coefficient [2110, 2139, 2151, 2100, 2194, 3160, 1814, 2934, 2500, 2668, 2886, 2952, 1540, 3283, 227, 733, 428, 478, 3300, 942, 820, 1712, 734, 1284, 1879, 707, 728, 747, 656, 3026, 3144, 3200, 3400, 3083, 2954, 3111, 3153, 3234]. Coefficients [2238, 2046, 2163, 2541, 198, 125, 2964, 2603, 467, 316, 1806, 117, 1683, 3061, 1372, 2988, 644, 90, 955, 993, 1371, 2457, 342, 2759, 1952]. cohort [1493, 1377, 2604]. cointegrated [2496]. cointegration [2469].

Collapsibility [3128, 2597]. collectives [19]. collector [2926, 3126]. collision [1281]. colloidal [84]. colors [1796]. column [1660, 1620, 650, 3399]. column-wise [3399]. Combination [2146, 596, 1148]. Combinations [2039, 1292, 1298, 931, 1867]. combinatorial [918]. combinatrial [140]. Combined [2096]. Combining [2846, 447, 1927, 1301]. come [314, 327, 1205]. come-and-stay [314, 327]. Comedians [2044]. coming [1568]. comment [608]. comments [1326]. Common [2209, 2139, 2069, 2344, 1658, 1299, 1402, 445, 852, 1002, 1340, 1769, 1354, 1202, 1689, 1555, 1683, 1191, 1733, 1466, 1583]. communication [162, 197]. communication-pattern [162]. Compact [2269, 2869, 1306, 1008]. compactly [3403]. comparaisons [222]. Comparative [3419, 824, 826, 143]. compare [2731]. Comparing [2158, 3423, 3278, 1370, 3370, 2906, 1398, 1593, 157, 811, 93]. Comparison [2343, 333, 813, 2494, 986, 2271, 2636, 1073, 1150, 2056, 771, 466, 582, 2040, 1091, 1402, 3052, 688, 1121, 345, 453, 481, 3213, 2453, 716, 512, 1391]. Comparisons [2500, 2179, 2082, 1573, 1103, 54, 1992, 1029, 2986, 2815, 1584, 2660, 686, 432, 222, 518, 1054, 527, 547, 804]. Competing [3067, 3370, 2787, 3147, 2838]. Competitors [1430]. Complete [1914, 868, 1590, 1078, 920, 890, 968, 3148, 1446, 1046, 1001, 1068]. complete-data [3148]. Completeness [1240, 714, 1232]. completion [3393, 2742]. complex [892, 3037, 667, 696, 634, 695, 676, 860, 516, 3393, 393, 599, 635, 1054, 461, 760, 682, 2752, 741]. complexes [836]. complexity [1382, 3239]. complier [3414]. Component [1380, 2069, 887, 309, 1833, 1769, 1471, 1996, 2570, 943, 328, 1475, 3378, 698]. Components [2111, 2063, 1197, 2761, 2501, 1667, 1159, 3249, 1624, 389, 909, 3209, 1353, 932, 573]. Componentwise [1912, 2500]. Composite [1549, 3085, 3110, 3364, 1681, 991, 3299]. composition [1158]. Compound [2524, 2565, 2307, 2196, 426, 3000, 2530, 2356, 500, 1622, 772, 970, 495, 392, 496, 2386, 2684, 230, 969, 1027, 1112, 1173, 1866, 789, 407]. compounded [1461]. compression [1169]. Compromise [2581]. computation [172, 1012, 108, 1548, 181]. Computational [2968, 94]. computed [1746]. computer [3180, 3156]. Computing [2289, 1570, 1312]. concave [2385, 3185]. concavity [1615, 1511]. Concentration [477, 707, 728, 747]. Concentration-curve [477]. concepts [1963, 1138, 1019]. Concerned [2322]. concerning [852, 1197, 1706, 667, 696, 516, 448, 744, 843, 942, 599, 566, 197, 265]. Concomitant [2326, 3205, 1951, 456, 436]. concomitant-based [3205]. concomitants [1148]. concordance [2609]. Condition [2280, 2952, 3260, 3155, 374, 1602, 1633]. Conditional [2150, 1540, 2189, 2815, 3082, 2201, 1827, 2329, 2681, 3375, 2282, 3245, 3087, 1216, 2385, 3229, 1344, 2840, 2664, 1956, 2725, 1177, 2887, 3019, 3260, 3012, 3383, 3359, 2454, 2460, 2763, 1522, 3169, 3210, 3231, 1269, 1401, 3179, 1322, 3057, 3430, 1853, 3368, 2785, 2873, 3135, 3142, 1762, 1822, 1512, 1463, 2374, 1818, 1985, 328, 2412, 2735, 2675, 3322, 2749, 1902, 2469, 1539, 1687, 1828, 3091, 2728]. conditionally [807]. Conditionals [2123, 3018]. conditioning [533]. Conditions [2519, 2164, 1003, 1266, 286, 1354, 1656, 435, 1432, 937, 2473, 353, 1885, 116]. Cone [2155]. cones [1899, 2369, 3192, 1943]. Confidence [2033, 2560, 2116, 1945, 2539, 1581, 2441, 2016, 2249, 2047, 2540, 2062, 3145, 3054, 1383, 125, 2615, 489, 3328, 1705, 1813, 1411, 192, 740, 486, 901, 3211, 1673, 1355, 271, 376, 616, 1339, 2907, 3117, 1465, 1980, 1211, 2368, 3042, 1835, 1659, 422, 2641, 2721, 1544, 214, 881, 267, 2848, 1212, 1513, 2474, 1514]. configuration [2727]. configurations [797, 876, 2743, 622]. conflicting [2922]. confounded [810]. congestion [415, 265]. conjecture [1971, 852, 1335]. conjectured [1767]. conjugate [1312]. connected [1738, 1084, 2435, 1854]. Connectedness [1308, 1534]. Connections [2550, 2757, 1541, 2576]. Conover [509]. Consecutive [2103, 2036, 1740, 2067, 1841, 1907, 1958, 3187, 2799, 1088]. consecutive- [1958]. consequent [1363]. Conservatism [2116]. Conserved [2153]. consideration [291, 771, 159]. considerations [508, 891]. consistencies [1445]. Consistency [2165, 1354, 2345, 2105, 2688, 2263, 1036, 1192, 2739, 2869, 2998, 3431, 1149, 3359, 3084, 3221, 3220, 1355, 2949, 2371, 3033, 1818, 1395, 2463, 3007, 699, 2459, 2502]. Consistent [2354, 1484, 3293, 2625, 2232, 2535, 2247, 2595, 1672, 772, 2460, 1227, 424, 3090, 383, 612]. constancy [3210, 3070]. constant [2886, 3182, 1256, 188, 1431]. Constellation [977]. constituents [1316]. Constrained [2665, 2809, 2856, 1274, 386, 2471, 3312, 2938, 2610, 2426]. constraint [2671, 3341]. constraints [53, 1284, 2503]. constructed [3117]. Constructing [2047, 1671, 3251, 3114]. Construction [1102, 1315, 2978, 2385, 1344, 1135, 605, 633, 774, 1393, 1660, 840, 1250, 1332, 776, 1323, 778, 764, 1324, 1936, 1549]. constructions [1309, 1455, 1617]. constructive [3173]. Contacts [11, 17, 30, 44, 58, 74, 79, 89, 97, 120, 146, 170, 177, 186, 194, 211, 212, 220, 223, 233, 242, 251, 275, 305, 329, 358, 394, 430, 470, 514, 569, 629, 675, 730, 783, 825, 877, 922, 964, 1013, 1059, 1105, 1151, 1207, 1253, 1297, 1349, 1400, 1457, 1473, 1490, 1505, 1519, 1535, 1550, 1565, 1579, 1595, 1608, 1621, 1634, 1648, 1662, 1677, 1692, 1710, 1724, 1739, 1756, 1773, 1791, 1809, 1825, 1842, 1857, 1871, 1887, 1904, 1918, 1931, 1944, 1959, 1974, 1988, 2002, 2015, 2029, 2042, 2055]. Contacts [2066, 2078, 2091, 2102, 2114, 2129, 2141, 2154, 2168, 2183, 2198, 2213, 2227, 2243, 2260, 2278, 2294, 2310, 2330, 2350, 2366, 2382, 2397, 2411, 2423, 2436, 2446, 2462, 2476, 2487, 2498, 2513, 2525, 2538, 2553, 2566, 2577, 2588, 2601, 2611, 2623, 2635, 2647, 2659, 2672, 2686, 2696, 2709, 2723, 2732, 2740, 2753, 2762, 2769, 2781, 2793, 2802, 2813, 2823, 2834, 2845, 2857, 2868, 2878, 2889, 2898, 2909, 2917, 2928, 2937, 2946, 2956, 2966, 2975, 2983, 2991, 3001]. containing [1285]. contaminated [2413, 1935, 1342, 1392]. contamination [1768]. contaminations [1772]. content [852]. Context [2175]. contiguity [1367]. Contiguous [1567]. contingency [1468, 2624, 2858, 240, 249, 598, 2743, 1150, 757, 2751, 3018, 512, 2597]. continuation [3375]. continued [1700]. Continuity [2545, 210, 1006]. continuous [1081, 1082, 1878, 3284, 1035, 1888, 2950, 3143, 772, 1731, 3338, 814, 494, 1316, 2367, 210, 280, 3250, 3256, 2789, 853, 1056, 2919, 3117, 199, 3334, 2971, 354, 946, 1313, 93, 132, 159, 3136, 1557, 1749, 1336]. continuous-discrete [3256]. continuous-time [2950, 3143, 3338, 814, 2919]. continuum [1796]. contoured [1488]. Contrast [2095, 2720, 1821, 3006]. Contrast-based [2720]. contrasts [1705, 1493]. contributions [616, 1716]. Control [156, 190, 643, 674, 1086, 3065, 1491, 2494, 1011, 1458, 2453, 967]. Controlled [360, 156]. Controlling [2047]. controls [1593]. conventional [1301]. Convergence [2065, 268, 3243, 2016, 726, 3063, 1465, 1771, 155, 788, 1245, 2941, 1755, 1607, 2998, 1078, 2629, 1368, 1510, 1290, 231, 286, 564, 864, 1656, 1794, 3371, 1331, 2482, 703, 1869, 295, 1561, 1173, 936, 207, 1302, 1388, 511, 1129, 787, 3127, 1727, 789, 975, 657, 33, 105, 1001, 3159, 18, 1893, 1812, 3103, 1633]. convergences [59]. Converting [2892]. Convex [2039, 2196, 2124, 2028, 1899, 1969, 229, 2654, 1191, 2359, 1889, 373, 2742, 1113]. convexe [229]. convexity [1615]. Convolution [805, 1521]. convolutional [3371]. convolutions [3398]. Conway [3182]. coordinate [927]. coordinate-free [927]. Copas [3333]. Copas-like [3333]. Copula [3299, 3198, 2978, 2789, 3402, 3030]. copula-based [3030]. Copulas [2593, 3248, 2875, 3027]. Core [2261, 1562]. Corpuscle [2076, 2259]. correct [1385, 928, 759, 150]. correctability [1989]. Corrected [2389, 2952, 2396, 3273]. correcting [1586, 1795]. Correction [674, 1267, 673, 888, 3221, 1456, 51, 920, 3171, 3289, 1058, 3385, 609, 962, 2200, 3269, 1006, 3333, 1552]. Corrections [996, 1238, 1755, 1261, 1327, 607, 606, 1807, 497, 1808, 481, 948, 536, 957, 249, 1328, 1348, 919, 842, 843, 1045, 1420, 588, 429, 455, 456, 610, 1419, 1933, 1009]. Corrective [2012]. Correlated [2121, 68, 106, 3280, 144, 1713, 1917, 1428, 2944, 1141, 2663, 1413, 1879, 3385, 3384, 2674]. Correlation [2238, 467, 2099, 2496, 1198, 859, 1806, 428, 478, 1196, 3061, 1109, 32, 1495, 942, 2708, 820, 1873, 1775, 656, 3298, 560, 3310]. Correlations [2531, 2450, 1752, 2962, 867, 1304, 1618, 1435, 1176]. correlogram [303, 625, 661]. corresponding [228]. Corrosion [2202]. cost [331, 1382, 112, 285]. costs [3030]. costwise [561, 623]. Count [2046, 2777, 2715, 3137]. countable [319, 1087]. countably [3325]. Counterexamples [1669]. Counting [2138, 259, 1577, 3147]. counts [1530, 2887, 3362]. Coupon [2926, 3126]. covariables [3350, 3025]. Covariance [1012, 2519, 2942, 941, 2080, 2208, 2580, 586, 3166, 1029, 2765, 3337, 1357, 2761, 1616, 1976, 1720, 792, 859, 821, 971, 2630, 2850, 3094, 3305, 3002, 1732, 756, 953, 1603, 3063, 3024, 2478, 2943, 680, 819, 842, 911, 958, 1463, 2708, 1898, 452, 2491, 655, 3331, 2902, 1913, 2929, 1695]. covariances [517, 2656, 301, 413, 975, 1176]. Covariate [2128, 2859, 3145, 3328, 3301, 3262, 2870, 2856, 1631, 2945, 2649, 2662]. covariate-adjusted [3145]. Covariates [2556, 3287, 2825, 333, 2835, 2815, 3256, 2959, 2661, 3068, 1375, 3189, 3219, 3390, 2921]. covariation [3022]. covariogram [2730]. cover [1508]. Coverage [2157, 1351, 3117, 1980]. coverages [1469]. covering [2799]. Cox [3262, 3410, 3349, 2488, 3005, 1771, 3101, 1995, 2820, 3372]. Craig [1943]. Cramér [1106, 2221, 1003, 348, 937, 1928, 237, 409, 1567]. Cramér-Rao [1003]. Cramér-type [937, 1567]. Cramér-von [1928, 237]. Criteria [2197, 2347, 2120, 1374, 2746, 1029, 3422, 667, 2389, 856, 27, 564, 865, 2398, 756, 2722, 579, 775, 1365, 680, 819, 842, 911, 1020, 2893, 4]. Criterion [1268, 2242, 1267, 3028, 974, 3247, 1856, 2791, 2772, 634, 526, 902, 951, 1456, 948, 925, 315, 3005, 644, 1031, 1110, 422, 3148, 561, 623, 105, 104, 2720, 109, 1812, 3020, 3014]. critical [1424]. critique [1301]. Cross [2963, 143, 3255, 1386, 2389, 824, 826, 2943, 2708, 985, 1252, 2613]. cross-entropy [1386]. Cross-national [143]. Cross-over [2963]. cross-societal [824, 826]. cross-validation [2389, 2613]. crossed [872]. crossing [183, 1520]. crossings [2407]. crushed [110]. cube [852]. Cumulant [2516]. cumulants [533]. cumulative [200, 973, 3370, 337, 511]. cup [3126]. cure [2883, 2640, 2554]. cured [2776]. current [2414, 2544, 2776]. curvature [2576]. Curve [2148, 2208, 2314, 2300, 2555, 905, 960, 3270, 1616, 2389, 1218, 1732, 401, 2375, 2818, 1337, 1898, 1244, 2899, 1466, 477, 1470, 1897, 3436]. Curved [1949, 1208, 1270]. Curves [2117, 3423, 1078, 3314, 899]. cuscore [2494]. cusp [3130, 3268]. cusp-type [3130]. cusps [1482]. customers [552]. cusum [2494, 2384, 3213, 3214, 2418, 2526]. Cutting [239]. cycle [156]. cyclic [2690, 2897, 623]. cyclically [2354]. cylinder [852]. cylindrically [1293].

d.f. [542]. dam [1204]. dans [229]. Dantzig [3349]. Data [2136, 1676, 2533, 2234, 2046, 2281, 2073, 2216, 2049, 2253, 2582, 2230, 3397, 3432, 2087, 2080, 2264, 2583, 2280, 2285, 2202, 2233, 2161, 2534, 2348, 2314, 2249, 2121, 2106, 2580, 2232, 2251, 2555, 2034, 2265, 2297, 2093, 2250, 2028, 2543, 2247, 2068, 653, 213, 800, 2676, 2990, 3193, 21, 3048, 3363, 2777, 3160, 3054, 3158, 3366, 3201, 2822, 2615, 1868, 2964, 2842, 3328, 2960, 2796, 2936, 3301, 2416, 3357, 510, 637, 3307, 2441, 2829, 3364, 3021, 3165, 3262, 2447, 1976, 738, 739, 3065, 2894, 3196, 23, 39, 345, 453, 481, 3040, 2734, 2648, 2942, 2414, 3147, 3053, 1665, 1196]. data [1618, 866, 1255, 3081, 3008, 3406, 3394, 2873, 3135, 3197, 2846, 2924, 3034, 3142, 2544, 2776, 3414, 3404, 646, 2969, 2817, 1283, 2636, 2639, 3010, 3194, 2663, 2374, 2438, 3318, 3072, 2694, 1928, 2987, 3309, 1169, 985, 2737, 1213, 1733, 2805, 2682, 2755, 2830, 3148, 830, 2804, 2803, 2940, 2876, 3332, 3392, 2381, 2930, 2715, 1470, 2467, 1718, 3407, 1443, 1170, 1419, 2827, 1474, 977, 1089, 237, 321, 2353, 2632, 2662, 2855, 3299, 3409, 3412, 2674, 1862, 1978, 3025, 3331, 2797, 2838, 2927, 3047, 3091, 2768, 3343, 2810, 2844, 3234, 2929, 3067, 3137, 2613, 2728, 1860]. data-analytic [1255]. data-based [3407]. Data-driven [3397]. data-mining [2648]. data-taper [1860]. datasets [3345]. de-biased [3283]. Dealing [2378]. death [1561, 2855]. decay [2861]. Decentered [1868]. Decision [2245, 82, 103, 131, 3055, 331, 244, 1385, 454, 136, 53, 1624, 408, 34, 51, 47, 111, 350, 1294, 920, 890, 968, 1222, 466, 582, 1223, 443, 161, 407, 416, 455, 109]. Decision-theoretic [3055, 1294]. decisions [1164, 3226, 3225, 2716, 3224, 3227]. decoding [2713]. decomposability [1240, 2673]. decomposable [2658]. decompose [1667]. Decomposition [853, 943, 2069, 1220, 1229, 2789, 298, 3218, 630, 2708, 706, 354]. Decompositions [2748]. Decompounding [2756, 2456]. Deconvolution [2336, 2329, 2200, 2269, 2792]. Deconvolving [1935]. deep [3431, 2839]. defects [239]. deficiency [1908]. defined [1016]. defining [1214]. Definite [2155, 1030, 431, 515, 535, 3347, 1141]. Definiteness [2302]. Definition [2306, 1188]. definitive [3059]. Degenerate [2915, 2828]. degradation [3195]. degree [2745, 1454, 356, 1330]. degrees [244, 4, 3112]. delay [959]. delayed [416, 455]. delta [1130, 1627, 994, 2459]. delta-sequences [994]. Demand [2130]. demands [1563]. Demonstration [2012, 2167, 2980, 3109]. denoising [2440]. Densities [2293, 2261, 2023, 2105, 2159, 500, 1855, 2673, 1601, 1481, 3278, 1736, 1748, 3207, 2816, 2993, 1775, 906]. Density [2234, 1249, 2355, 2190, 2542, 2199, 2006, 2280, 2246, 2774, 2201, 2329, 994, 2106, 2224, 2885, 3258, 1036, 1130, 1192, 1181, 1154, 1482, 80, 225, 1937, 1909, 2914, 3177, 2869, 3087, 2842, 411, 3319, 2490, 1078, 1194, 1344, 2992, 1319, 2413, 3157, 2602, 2725, 3317, 1092, 1193, 2394, 575, 1808, 1782, 3050, 1831, 1093, 1977, 1935, 2680, 1296, 645, 1492, 2448, 1571, 1748, 1865, 1656, 1041, 3231, 1401, 2495, 2910, 3404, 3204, 2626, 894, 1201, 2947, 1867, 1598, 1395, 1129, 2641, 141, 1994, 1099, 1001, 1602, 3244, 142, 1066, 1978, 2985, 2981, 3241, 2613, 1336, 1860]. density-based [2495]. Density-ratio [2885]. denumerable [814]. Departures [2116, 2020, 271]. Dependence [2012, 2257, 2246, 2164, 2275, 2059, 2427, 1851, 2679, 3357, 3238, 1731, 3383, 3046, 1409, 1524, 1838, 2867, 2771, 1965, 1138, 1019, 1602, 3039]. Dependent [2104, 2037, 2122, 2256, 2591, 2332, 2530, 2184, 2199, 2598, 2312, 2270, 1036, 1080, 1130, 1181, 1999, 3160, 2354, 3315, 2837, 2351, 3374, 3076, 2835, 3395, 901, 1935, 2900, 1787, 311, 866, 1372, 2785, 2873, 3135, 2839, 1884, 2817, 2415, 1244, 1867, 3132, 2737, 2735, 1353, 521, 2426, 725, 3435, 2768]. depending [414]. depends [524]. Depletion [2052, 1367]. Depth [3031, 2960, 2866]. Depth-based [3031]. Derivation [1984, 2503, 1977, 677, 1256]. derivative [1319, 1306, 1968, 2985, 1336]. derivatives [1755, 3360, 1194, 1290, 1779, 3129]. derived [25, 1640, 548, 1381, 1361]. descent [3431, 1661]. Design [2137, 2579, 2070, 2197, 2148, 2242, 2034, 276, 490, 2440, 2479, 2819, 3285, 3017, 1744, 24, 26, 136, 1038, 279, 284, 507, 41, 260, 449, 748, 3361, 1518, 1936, 1361, 891, 3020, 2604]. design-adapted [2479]. designing [1917]. Designs [2550, 2549, 2203, 1086, 2196, 2349, 2537, 2149, 810, 1940, 2746, 1592, 2790, 3145, 823, 1309, 1455, 2745, 861, 2665, 1713, 1534, 2512, 1632, 1121, 777, 1362, 450, 497, 1856, 2603, 508, 549, 605, 3114, 1973, 1738, 839, 812, 1184, 1293, 1347, 1398, 1489, 1593, 1660, 3339, 796, 1070, 1085, 1084, 1120, 1250, 1308, 1454, 336, 349, 3156, 708, 1620, 650, 668, 374, 435, 1376, 1453, 2425, 1432, 2882, 2931, 1472, 811, 2453, 3178, 1102, 1815, 836, 822, 622, 795, 308, 451, 669, 778, 764, 863, 1324, 862, 1018, 1708, 1549, 561, 623, 809, 946, 1252]. designs [1879, 1606, 3059, 837, 3020]. designs-Some [508]. desired [415]. detected [2418]. Detecting [3265, 3337, 2438, 967, 2282, 2494, 870]. Detection [2585, 2253, 1488, 2021, 1665, 2380, 2448, 1962, 1639, 1186, 2979, 2719, 1898, 1169, 3113, 3375, 3369, 2961]. Determinant [1917, 123, 761]. Determination [236, 2100, 118, 1660, 198, 373, 1812]. determined [235]. Determining [3310, 1398]. deterministic [2973]. developments [1731, 1666]. Deviation [2061, 335, 2095, 1182, 102, 1145, 1799, 1193, 1599, 1300, 2806, 384, 1653, 3405, 283, 83, 383, 2794]. deviations [2515, 1723, 370, 3265, 1101, 3051, 2901, 1041, 2933, 1567]. device [923]. Diagnosing [2003]. diagnosis [95]. Diagnostic [2282]. Diagnostics [2209, 2128, 1680, 2705, 3056, 1846]. Diagonal [2523, 852, 2564]. diagonals [2378]. diagram [3407]. dichotomies [335]. dichotomous [1394, 1100]. difference [1790, 3162, 1115, 1428, 773, 1281, 2670, 1102, 1835, 622, 1733, 83, 3123]. differences [3337]. differencing [1330]. Different [2080, 2329, 2695, 868, 2887, 93]. differentiability [1656, 937]. Differentiable [2027]. Differential [1208, 1160, 1161, 1679, 1991, 892, 578, 3267, 1146, 1793, 3396, 182, 2794]. differentials [1706]. Diffusion [2046, 2526, 2161, 228, 2313, 3425, 2630, 550, 2854, 2910, 173, 3290, 2704, 2880, 2691, 2862, 2505, 3113, 2720, 1517, 1597]. diffusions [2782]. digamma [1048, 1062]. digital [1169]. dilation [2510]. Dimension [2303, 2337, 2083, 3208, 3193, 2842, 2996, 2673, 3267, 671, 3273]. Dimensional [2060, 2069, 2028, 2948, 2420, 3330, 2676, 2990, 2960, 3212, 3229, 3243, 1840, 697, 3040, 3391, 3094, 3305, 2979, 2959, 3008, 3063, 2924, 3142, 3005, 3351, 31, 48, 1441, 3209, 3312, 3309, 2938, 281, 707, 728, 747, 63, 2597, 977, 3200, 3421, 713, 3372, 3310, 3241]. dimensionality [234, 3270, 2625, 2427]. dimensions [1184]. Direct [2335, 693, 2649, 1594, 2943]. Directed [2103, 3367, 1923]. Direction [2097, 1605, 1704]. Directional [2867, 1448, 1868, 3081]. directions [831]. Dirichlet [1143, 2773, 2497, 427, 2169]. disaster [343]. disc [2455]. discipline [1613]. discontinuities [906]. discontinuous [1736]. discordant [100]. Discovering [3228]. discrepancy [1152]. Discrete [1314, 2578, 2442, 479, 703, 2157, 2017, 161, 407, 2094, 2028, 1021, 537, 1287, 1525, 1721, 2865, 2388, 3360, 2456, 1216, 1397, 970, 419, 691, 879, 3000, 392, 1014, 502, 3053, 3250, 568, 3169, 384, 3256, 292, 1542, 3179, 2783, 3277, 1811, 2782, 1986, 2754, 207, 1124, 506, 1442, 134, 2412, 1698, 1353, 2720, 321, 1528, 2632]. discretely [2862, 2843, 2903]. Discretized [1017]. Discriminant [2195, 2080, 307, 972, 1631, 636, 580, 792, 758, 941, 663, 2457, 357, 710]. discrimination [2676, 2665, 1964, 1767, 254, 258, 1153, 1187, 181]. discriminatory [333]. discs [1046]. Discussion [1636, 3226, 3354, 3417, 3416, 3225, 3139, 3355, 3227, 3140, 3356]. disease [2452]. diseases [3222, 1088]. disguised [774]. Disparity [2085, 1878]. Dispersion [2080, 2071, 290, 1684, 1803, 1910, 376, 318, 163, 912, 3047]. dissimilarity [697, 978, 2072]. Distance [3193, 2084, 2006, 350, 2990, 684, 685, 1072, 3381, 1969, 1750, 1701, 1083, 642, 770, 1037, 64, 103, 111, 131, 1343, 1965, 789, 60]. Distance-based [3193, 2990]. Distances [2276, 2719, 1548, 187]. distinct [765, 3335]. distinguishability [801]. distortion [2911]. Distributed [2328, 127, 68, 1543, 916, 3364, 1423, 3394, 3345, 828, 1479, 281]. Distribution [2079, 2037, 2188, 2248, 2135, 2328, 2837, 2018, 2043, 594, 684, 2150, 2185, 2186, 2134, 2011, 486, 636, 634, 2174, 580, 2596, 2607, 2308, 2016, 2124, 2151, 2342, 2245, 1083, 3321, 2237, 531, 2547, 2155, 2285, 2719, 1109, 2085, 2276, 2121, 2296, 2118, 2022, 2250, 2169, 2334, 2187, 2318, 3240, 200, 803, 981, 1005, 1094, 1356, 127, 195, 1076, 3365, 1142, 1106, 1350, 1958, 1203, 1111, 735, 459, 2573, 973, 1984, 2408, 2612, 2730, 388, 519, 1982, 2671, 3119, 475, 369, 714, 102, 2955, 1278, 1134, 370, 532, 1145, 1855, 2634, 1216, 2627, 765, 1647, 1590, 2373, 772]. distribution [807, 691, 216, 1408, 1725, 1568, 974, 1055, 1411, 1722, 3037, 3232, 3413, 1306, 2887, 1852, 1612, 1799, 1412, 1976, 3182, 1075, 1117, 696, 2452, 42, 893, 901, 1126, 3211, 2431, 774, 206, 1975, 792, 3387, 3327, 1444, 639, 670, 773, 412, 431, 515, 535, 695, 677, 676, 694, 761, 951, 971, 1107, 1456, 1605, 1227, 743, 769, 1254, 1540, 390, 392, 496, 2386, 3347, 1499, 135, 346, 1014, 2497, 1883, 1839, 571, 1874, 230, 147, 1832, 301, 1063, 198, 2455, 833, 855, 1043, 366, 568, 850, 1962, 2687, 2506, 1229, 255, 288, 428, 756, 898, 1217, 1461, 1942, 953]. distribution [1480, 1644, 2000, 3164, 652, 1569, 1880, 2881, 3008, 269, 1410, 1049, 3090, 1139, 1527, 579, 644, 679, 1031, 1699, 2475, 1298, 2360, 3117, 1541, 1337, 1345, 91, 931, 2817, 110, 31, 48, 999, 1079, 1095, 1241, 1333, 3324, 2432, 2880, 1045, 1023, 236, 353, 449, 748, 1211, 2771, 1124, 361, 933, 294, 398, 820, 2374, 2357, 440, 635, 1054, 1020, 1110, 1180, 452, 1972, 822, 930, 2726, 1570, 1125, 1167, 1985, 328, 403, 1546, 2417, 320, 2509, 256, 293, 1047, 1040, 1060, 1137, 1215, 1364, 1449, 1544, 1737, 355, 1062, 896, 134, 1221, 881, 341, 2804, 2803, 2940, 666]. distribution [682, 383, 356, 657, 707, 728, 747, 987, 1009, 380, 539, 63, 954, 1307, 513, 10, 16, 1119, 2852, 1704, 1849, 540, 1902, 1155, 1232, 687, 713, 1103, 1671, 1557, 1749, 2728, 1830]. Distribution-free [486, 3321, 803, 475, 714, 901, 135, 346]. Distributional [2088]. Distributions [2244, 2339, 1841, 2103, 2210, 2256, 2523, 2036, 2317, 2589, 2255, 2343, 2235, 2020, 2089, 2517, 2332, 2530, 2014, 2131, 2598, 2172, 2257, 292, 2123, 2010, 2171, 2333, 2086, 2182, 2170, 3259, 1337, 2254, 2329, 2152, 2594, 29, 2097, 682, 712, 2067, 2286, 2038, 3106, 2072, 2939, 200, 1998, 1081, 1082, 1182, 1200, 805, 1299, 1448, 1574, 1970, 2891, 3052, 1219, 1287, 1314, 1525, 1907, 3075, 3187, 2511, 3296, 1776, 2922, 3398, 2877, 3382, 68, 106, 2564, 2618, 1261, 1262, 1327, 1878, 101, 268, 3284, 835, 500, 1836, 404, 1894, 2442, 2658, 1804, 847, 1216, 918, 1397, 2808]. distributions [939, 972, 522, 606, 970, 2700, 35, 1163, 1481, 917, 495, 879, 3325, 426, 1386, 1416, 1964, 122, 2906, 1731, 814, 3102, 1510, 1921, 667, 763, 808, 834, 952, 1359, 2486, 3051, 1487, 856, 759, 152, 1429, 1599, 1033, 978, 3275, 755, 3327, 1488, 695, 721, 1305, 1316, 2501, 860, 479, 832, 1360, 150, 210, 280, 317, 498, 865, 1649, 2361, 2419, 2466, 2484, 2684, 164, 940, 806, 1311, 1522, 3027, 3250, 427, 619, 384, 255, 278, 516, 616, 1326, 1202, 1237, 1108, 1542, 1653, 2546, 1941, 1914, 1805, 3179, 2744, 1317, 1569, 2425, 1183, 2904, 1322, 1943, 3063, 3406]. distributions [3181, 567, 1289, 1869, 2783, 2933, 1585, 2669, 3089, 775, 1276, 889, 919, 1152, 1298, 1365, 441, 618, 1789, 503, 1697, 869, 1864, 1436, 680, 762, 819, 842, 432, 744, 843, 1478, 2867, 468, 1691, 969, 1027, 1112, 1173, 207, 3056, 1366, 1901, 3016, 1494, 546, 599, 1591, 1866, 3209, 1570, 1955, 1259, 914, 1343, 167, 1172, 1963, 1854, 1138, 1315, 1922, 2430, 1434, 328, 1019, 415, 593, 1392, 168, 1048, 1123, 1511, 1652, 841, 121, 2786, 742, 1844, 789, 1074, 1475, 306, 1452, 611, 33, 59, 76, 2412, 1734, 267, 2372, 155, 818, 1823, 1353, 521, 2503, 1234]. distributions [1902, 259, 1212, 1513, 3245, 1091]. Disturbance [664, 868, 1915]. Disturbances [2177, 912, 880]. Divergence [2316, 2532, 2233, 2914, 2892, 2568, 2772, 3050, 3051, 2495, 2910, 3223, 3404, 1259, 3148, 2885, 1670, 1715, 2503]. Divergences [2081, 3199, 2395]. diversity [1790]. divided [83]. Divisibility [2212, 1232]. divisible [450, 497, 508, 853, 943, 372, 1549, 354]. division [921]. Dixie [3126]. DLE [1017]. Does [2716, 2435]. Domain [2051, 3269, 1934, 1667, 1889, 593]. domains [3207]. Dominance [1097, 3378]. dominate [2852]. dominating [2473]. Doob [353]. Dose [1088]. Dose-response [1088]. doses [1103, 1091]. Double [1834, 1032, 1402, 2837, 1551, 720, 1097]. double-indexed [1551]. doubled [1985]. doubly [1107, 1043, 3120, 3121, 1928, 403, 2755, 2804, 2803, 2940]. down [2614]. Dr. [2710]. drawing [3190]. drift [1281, 2506, 433]. Driven [2281, 2222, 3422, 3397, 3396, 2903]. dropouts [3072, 3306]. drug [1088]. drug-induced [1088]. Dual [2757, 2228]. Duality [2032, 2931, 845, 962]. due [1506, 117, 353, 827]. duration [3057, 2997, 1088]. Dyadic [2292, 1536]. Dynamic [2585, 2236, 417, 2994, 3131, 3107, 2494, 3313, 1576, 883, 3137]. Dynamical [2293, 2153]. dynamics [3132, 2715].

E-optimality [1620]. Earthquake

[2077, 2718, 1676, 1170, 1419]. earthquakes [1649, 2839, 1693]. east [344]. EBIC [3040]. EBLUP [2826]. econometrics [1358, 1447]. economic [400]. Edgeworth [1208, 2356, 226, 1877, 1786, 1838, 2801, 2393, 954, 2767]. editors [3282]. Effect [213, 2542, 867, 1604, 298, 262, 2651, 68, 838, 1845, 604, 1651, 1228, 3368, 2719, 260, 548, 624, 711, 451, 584, 3138]. Effective [2543, 1444]. Effects [2579, 2026, 2082, 1378, 2320, 106, 3373, 1632, 3263, 2965, 871, 3395, 689, 3387, 1214, 608, 583, 745, 1685, 3029, 2661, 3414, 872, 794, 1812, 3105, 3421, 1913, 3420]. Efficiencies [1203, 488, 1389, 1176, 1559]. Efficiency [1878, 2007, 1738, 2832, 2304, 2085, 2149, 1131, 882, 1646, 1971, 3365, 2655, 326, 3036, 1270, 1429, 1967, 750, 1582, 2815, 816, 410, 2872, 915, 2764, 289, 596, 1247, 620, 116, 1188]. Efficient [2533, 2341, 3081, 3125, 3324, 3334, 752, 165, 2692, 3161, 3366, 3271, 1987, 2796, 2996, 2835, 3176, 2882, 574, 701, 3064, 1606, 1849, 906, 3421, 2954]. EIC [2031]. eigenvalue [172, 3193, 1464].

Eigenvalues [2086, 1467, 1644, 1495, 2417]. eigenvectors [1495]. Einstein [2046]. EISE [2475]. Eisenhart [279]. elections [1026]. Elementary [77, 440, 545, 1671]. elements [697]. Elfving [1856]. elicitability [3266]. elimination [1385, 1347, 3089, 1472]. ellipsoid [1351]. elliptic [1541]. elliptical [3088, 2486, 1425, 1995, 1734, 2383]. elliptically [1488]. embeddable [2619]. Emission [2163]. empiric [657]. Empirical [2033, 2963, 2096, 2539, 523, 607, 226, 2608, 3359, 2174, 2308, 2120, 2230, 1962, 1996, 3023, 3333, 2873, 1527, 2475, 2284, 2899, 2249, 2987, 2641, 1303, 1727, 2930, 2250, 2392, 2788, 2557, 2633, 1654, 1142, 1350, 3162, 1672, 3071, 3230, 1813, 3269, 684, 1144, 2886, 1471, 3082, 483, 536, 1014, 3002, 1624, 2506, 2654, 2789, 1322, 3063, 1702, 703, 2492, 2924, 3090, 1743, 2360, 1469, 3183, 3289, 3288, 909, 2697, 1835, 1995, 2820, 1678, 1528, 3306, 1893, 3077, 3123, 2399, 1514]. empty [670, 3122]. endpoints [3206]. energy [3146]. enflanking [724]. Engen [2485]. enhancement [3171, 3170]. enough [3238]. ensembles [3344]. Entropy [989, 1483, 2267, 419, 1386, 3051, 1763, 2738, 1571, 465, 550, 1180]. Entry [2126]. enumeration [2361]. Envelope [2292, 1949]. environment [3410]. environmental [1492]. epidemic [2718, 191, 264]. epidemic-type [2718]. Equal [2518, 2902, 639, 365, 1289, 851, 1210]. Equality [2124, 3166, 1029, 1004, 1720, 517, 859, 3094, 1806, 525, 1542, 665, 681, 680, 819, 842]. Equally [2137, 2262, 1843]. Equally-Weighted [2137]. equation [2994, 2511, 1061, 578, 3267, 3415, 3418, 1281, 1447, 3086, 3417, 3416, 173, 886, 484, 555, 588, 1060, 3396, 1301, 2927, 3103]. Equations [2092, 2178, 107, 1991, 1948, 2890, 868, 1607, 892, 3049, 1189, 67, 2923, 1276, 393, 3189, 1793, 1413, 620, 664, 912, 976, 1097, 1313, 1552, 3299, 182, 2794, 3047, 3153]. equicorrelated [539]. Equilibrium [2052, 554]. equiprobable [1589]. equireplicated [1085]. Equitable [899]. Equivalence [3281, 3374, 2390, 317, 498, 865, 1869, 2812, 1171, 1152, 1365, 1057, 1452]. Equivariance [2097, 3014]. Equivariant [2240, 2231, 3061, 2471, 872, 573]. Ergodic [148, 3360, 3422, 3035, 528, 3290, 2704, 3113, 2720, 3396]. ergodicity [2915]. Erlang [601, 259]. Erlangian [1563]. Errata [160, 232, 86, 87, 88, 96, 73, 145, 169, 241, 250, 130, 184, 138, 185, 119, 43, 219, 202]. Erratum [2888, 3045, 3121, 2804]. erroneously [216]. Error [2150, 1412, 2205, 2521, 2058, 2583, 2522, 1641, 1449, 2022, 2030, 2047, 2108, 2520, 2177, 2180, 2181, 3241, 2859, 213, 1203, 123, 3301, 2992, 180, 2791, 1675, 2394, 1444, 113, 3176, 2904, 844, 2559, 2913, 1060, 1737, 899, 1127, 2562, 1263, 2353, 1679, 1695, 1336]. error-areas [899]. error-in-response [2353]. Error-in-Variables [2205, 2181]. Errors [2298, 2589, 2320, 3258, 1394, 1619, 2700, 3216, 3374, 2602, 2952, 3395, 3084, 485, 631, 2900, 2698, 934, 117, 1201, 2415, 2663, 3361, 1879, 2383, 2674, 3025, 3435, 2595, 1765, 2911, 3083, 3426]. errors-in-covariables [3025]. Errors-in-Variables [2320, 2952, 3084, 2698, 3361, 2595]. Escort [3051]. esophagus [1792]. espace [229]. Esseen [3217, 1908, 2054]. estate [1919]. estimable [1493]. estimate [445, 446, 996, 647, 3431, 1264, 1092, 1226, 460, 1044, 625, 957, 956, 46, 1997, 1283, 1655, 828, 257, 895, 83, 3322]. Estimated [2186, 2315, 1106, 3379, 2475, 1695]. Estimates [2144, 243, 2521, 2279, 2289, 2345, 361, 2105, 2225, 2297, 2263, 2323, 2718, 802, 1192, 1245, 735, 459, 884, 491, 868, 2456, 905, 50, 1078, 423, 1783, 1526, 1101, 1072, 2424, 2621, 2622, 1359, 1193, 1334, 3220, 301, 1877, 1861, 45, 2371, 112, 315, 1537, 3404, 41, 704, 1821, 1042, 1827, 1598, 938, 1965, 1445, 100, 883, 1176, 471, 612, 505, 610, 1001, 1698, 1979, 2502, 204, 2426, 1155, 2985, 1338, 3221]. Estimating [2295, 2655, 3022, 2046, 2851, 1910, 2223, 2343, 2807, 2798, 2057, 2616, 1645, 2516, 1093, 2690, 1748, 2293, 3231, 1603, 1762, 3108, 2669, 1053, 2296, 2241, 1466, 1800, 2218, 2142, 2935, 1687, 2178, 1860, 1174, 359, 1923, 310, 1774, 1340, 35, 334, 1122, 2603, 1189, 1736, 1934, 1316, 1499, 1649, 731, 1498, 1492, 850, 278, 1644, 2881, 1960, 2667, 1056, 3189, 1381, 620, 371, 1819, 2997, 513, 3299, 1091, 3331, 2927, 3047, 3153, 3420]. Estimation [1658, 2238, 1299, 1525, 2561, 2325, 2533, 2234, 49, 2221, 2107, 2341, 1802, 411, 2627, 3028, 2190, 2301, 2936, 332, 2944, 3307, 216, 2992, 2235, 2479, 2784, 1530, 1403, 1759, 960, 2517, 3260, 1688, 575, 2117, 2115, 2220, 3292, 1831, 2139, 2119, 2501, 2582, 2734, 948, 925, 2199, 2189, 2488, 2084, 2342, 2245, 1225, 1158, 1571, 2982, 311, 3118, 2340, 1689, 2006, 2904, 1880, 2532, 2086, 2268, 2280, 3041, 3414, 1585, 2045, 2202, 2774, 646, 90, 2064, 503, 2008, 3376, 894, 1136, 2559, 2085, 2100, 2284, 2266, 1995, 2106, 663, 2017, 403, 2105, 2417, 2118, 2862, 2224, 1727, 1284, 1775]. Estimation [2491, 2580, 2876, 2232, 611, 1664, 1947, 1850, 2034, 2265, 1583, 2025, 2098, 2353, 60, 1895, 1330, 687, 2176, 2344, 3253, 3030, 2320, 2267, 2181, 2831, 2269, 299, 1668, 3255, 3258, 1036, 1081, 1082, 1130, 1181, 1154, 1574, 1356, 1556, 276, 277, 377, 473, 556, 591, 1104, 1958, 1380, 1937, 1755, 1577, 2739, 586, 1909, 2821, 3054, 1650, 3366, 1878, 2352, 3180, 2645, 3177, 2683, 1684, 2869, 3388, 3174, 1829, 1484, 532, 1894, 3087, 785, 2842, 847, 2490, 2679, 1249, 2355, 1405, 1515, 1194, 2996, 3130, 2792, 2481, 2416, 3152, 437, 3216, 637, 1833, 1601, 419, 3092, 3264, 1319]. estimation [1274, 3316, 2451, 2413, 3247, 3110, 2829, 2664, 1467, 3017, 2379, 1416, 1726, 2602, 2725, 3036, 3165, 1379, 1177, 3338, 1848, 2973, 3270, 2809, 2856, 2925, 2666, 1799, 2369, 3317, 1290, 1243, 3050, 3293, 3012, 3383, 3196, 1926, 1779, 1763, 1975, 1977, 1516, 755, 3327, 386, 2400, 252, 2680, 2763, 3283, 1582, 2815, 2850, 303, 661, 1190, 1586, 1218, 1296, 1627, 1832, 913, 1611, 2448, 1951, 660, 1865, 1801, 3295, 2405, 1996, 1657, 263, 3276, 1248, 1665, 3149, 3231, 2766, 3235, 3358, 1981, 1497, 3408, 1683, 1834, 2879, 1820, 1317, 1255, 3203, 3008, 3086, 3368, 3394, 3024, 703, 1863, 2640]. estimation [3078, 3125, 3345, 1527, 2544, 2988, 3062, 3405, 1764, 1558, 2472, 64, 790, 1925, 2836, 701, 1864, 3172, 3207, 1079, 2642, 1157, 2972, 2625, 630, 969, 1027, 1112, 1173, 909, 1562, 236, 3246, 1025, 1032, 398, 3334, 1244, 1866, 3101, 3072, 1712, 289, 705, 994, 1285, 753, 520, 506, 3055, 752, 1580, 1629, 2729, 849, 312, 780, 734, 1303, 848, 1335, 238, 2505, 664, 912, 1242, 1446, 1994, 1879, 2804, 2803, 2649, 2885, 3006, 3219, 789, 342, 975, 1176, 162, 245, 539, 613, 3297, 741, 2977, 2997, 1694, 891, 1424, 2540, 1127, 1353, 2383, 3004, 3244, 2733, 1113, 2953, 1735, 1716]. estimation [1170, 1419, 1912, 609, 572, 614, 648, 655, 656, 142, 737, 1528, 2657, 2662, 2855, 3136, 3409, 1369, 1893, 1978, 1966, 3399, 2736, 1103, 1148, 1614, 1557, 3323, 2595, 1765, 2617, 3085, 2929, 3426, 2703, 3241, 2613, 2728, 1946]. estimations [3218, 1095]. Estimative [2159]. Estimator [2061, 2018, 2534, 2266, 2177, 2600, 1076, 3365, 2955, 1051, 3212, 772, 1357, 3184, 846, 1799, 1750, 2394, 1767, 2452, 1429, 3425, 2630, 1227, 769, 3340, 390, 1404, 1271, 731, 767, 1784, 3002, 1604, 816, 1354, 2847, 410, 2375, 2872, 424, 2995, 1500, 1447, 3061, 1560, 1685, 2495, 2910, 2785, 3135, 3060, 3351, 2801, 2471, 2473, 2626, 1697, 1224, 1333, 1766, 1625, 1674, 1057, 1077, 2987, 585, 1546, 621, 320, 3127, 1737, 341, 2940, 383, 1028, 1800, 1050, 1718, 3217, 3200, 1849, 1862, 2418, 1908, 1286, 1828, 1597, 1749, 2903, 1188, 1210]. Estimators [2166, 2146, 2343, 522, 606, 2179, 2165, 2095, 2231, 2526, 2583, 2246, 2082, 2201, 2233, 2164, 2852, 2247, 2324, 2590, 882, 1017, 1156, 1352, 1448, 1402, 1482, 1646, 1203, 2515, 2730, 1945, 1002, 2354, 1145, 2483, 3360, 1804, 3319, 970, 1623, 1807, 2765, 813, 768, 3424, 122, 1306, 2404, 1976, 1382, 2390, 3205, 3084, 1545, 1360, 3046, 571, 1318, 1498, 1626, 1656, 1339, 3176, 424, 1196, 915, 866, 1555, 1153, 1187, 2482, 2425, 1685, 3261, 1191, 2812, 1743, 2586, 736, 3204, 1337, 3233, 2468, 1241, 3290, 955, 2444, 894, 988, 3281, 2771, 2758, 933, 3098, 1818, 2357, 3329]. estimators [373, 405, 1128, 1024, 1395, 720, 1717, 1129, 872, 2691, 1573, 1733, 1381, 1746, 621, 2457, 2830, 1911, 1936, 1572, 1097, 1775, 2554, 2804, 2803, 1483, 1734, 1930, 2759, 1602, 1209, 2724, 2824, 788, 2459, 1715, 1066, 3208, 904, 442, 592, 3336, 1643, 906, 1908, 1282, 1517, 2604, 573, 1336, 3429, 1000]. Euclidean [2428]. Eulerian [1647, 1492]. evaluation [123, 183, 3419, 385, 844]. evaluations [199]. even [933]. event [658, 1741, 2416, 176]. events [2668, 2974, 3099]. everywhere [1655]. Evidence [2716, 1301]. Ewens [2321, 2984]. Exact [2523, 2564, 2671, 2407, 2373, 3021, 3381, 1921, 2332, 2894, 2607, 2670, 2368, 2050, 1020, 2276, 794, 1713, 1754, 3387, 196, 3169, 478, 756, 652, 3089, 579, 679, 775, 1031, 1256, 1298, 1110, 1867, 1972, 1570, 1167, 457, 2824, 1210]. examination [226]. example [1900, 2725, 424, 381]. Examples [2336, 1190, 1170, 1419]. Exceedance [2088, 2887]. exceedances [1794, 379]. Exceeding [2173]. exchange [2381]. exchangeable [2865, 3080]. exciting [3146, 3261, 1025]. exhibiting [1563]. Existence [1021, 2144, 2336, 1283, 373, 3395, 424, 374, 435, 1821, 622, 886, 946, 904]. Expansion [2150, 2596, 2270, 1016, 2796, 807, 1799, 2951, 952, 3182, 761, 951, 971, 1456, 953, 2801, 999, 1771, 2393, 2445, 2691, 2729, 1449, 666, 954, 3336, 1000, 1210]. Expansions [1723, 2014, 2022, 2286, 2113, 1719, 1208, 2699, 2356, 939, 1481, 667, 808, 1412, 1305, 940, 1877, 1787, 1786, 1480, 1838, 1697, 1980, 762, 1378, 2708, 1876, 2443, 1798, 2767]. Expectation [3047, 1261, 1262, 1327, 1325, 685, 2454, 1269, 3008, 1849, 3039]. expectation-based [3008]. Expectation-robust [3047]. Expectations [2041, 2050, 2315, 1433, 1481, 1956, 1512, 1985, 3291]. Expected [2112, 3075, 1706, 832, 1876]. expected/observed [1876]. expectile [3342, 3286, 3191]. expectiles [1916]. Experiment [2214]. experimental [2745, 1744, 1506, 385, 2931, 1518, 1252]. Experiments [2197, 2040, 810, 2614, 3180, 1728, 1100, 654, 248, 3156, 745, 1422, 527, 822, 597, 771, 176]. explanatory [3281, 1329, 1420]. Explicit [2771, 975, 3285]. explosive [788, 1132, 1263]. Exponent [2023, 3118]. Exponential [2079, 2188, 2569, 2319, 2235, 2537, 2186, 2016, 2237, 2171, 2285, 2266, 2864, 2071, 1998, 981, 1005, 1094, 1021, 1402, 3052, 1208, 2573, 2671, 2695, 1272, 1538, 3201, 369, 1836, 1829, 1855, 1433, 2627, 2373, 1408, 992, 495, 1417, 1386, 1416, 1270, 1379, 122, 1510, 1243, 3192, 1126, 1516, 1254, 2386, 1584, 2631, 1832, 1043, 3169, 418, 1996, 1657, 801, 1217, 534, 1949, 1322, 1880, 269, 1049, 1585, 790, 1337, 1864, 20, 1224, 1478, 1042, 1986, 236, 3093, 1125, 732, 415, 1047, 1040, 1060, 1137, 270, 789, 3392, 1912, 1528, 3411, 1830]. exponential-type [1386]. exponentiality [1672]. exponentially [2463]. exposure [1496]. expression [1038]. expressions [1754, 833]. Extended [3005, 2795, 2485, 3086, 1027, 1891]. Extending [2939, 2572, 1055, 1135]. Extension [1051, 2031, 954, 2932, 935, 1003, 1159, 1320, 2872, 1822, 364, 461]. Extensions [1276, 2989, 1778, 70, 1410, 760, 707, 794]. extentions [728, 747]. extra [1708]. extraction [2919]. extraneous [342]. Extrapolating [2578]. Extrapolation [2196, 336, 349]. extrema [235]. extremal [339, 1312]. Extreme [2529, 2252, 3080, 2045, 2118, 268, 1403, 1368, 1938, 3260, 1260, 1237, 869, 1277, 217, 187, 1452]. Extreme-Value [2252]. Extremes [2121, 3398, 368, 1955]. eye [2636].

Factor [2140, 2207, 3352, 1632, 2750, 2888, 2841, 1354, 3313, 1927, 1437, 1708]. Factorial [2550, 2549, 1375, 810, 2746, 2790, 777, 708, 1376, 1453, 900, 1472, 3064, 836, 624, 711, 597, 862, 1018, 1361, 1708, 400, 3007, 837, 3020]. factorials [838, 1632, 604, 584]. factorization [973]. Factors [2048, 2313, 636, 193, 946, 995]. fading [2380]. faible [229]. Failure [1564, 2048, 2228, 2093, 1999, 1909, 1983, 1417, 1729, 3076, 2811, 3414, 2632, 3073, 3067]. Failures [2036, 2167, 2255, 1841, 1907, 3058, 2419]. Falls [2185]. familial [1196, 1618]. Families [2186, 725, 2323, 2978, 1433, 992, 1379, 1243, 3192, 1033, 1584, 3169, 1996, 1949, 1986, 1165, 1528]. Family [2244, 2339, 2316, 2124, 2071, 2932, 1448, 1970, 2891, 3052, 1208, 1272, 1855, 1270, 2738, 2806, 801, 3223, 790, 1224, 468, 969, 1027, 1112, 1173, 3016, 3093, 732, 1123, 1727, 789, 3392, 2372, 1830]. Farlie [1410]. Fast [3323, 2692]. Faster [3089]. Fay [2826]. feasibility [1665]. feasible [3347]. feature [3212, 3040, 3142, 3005, 3312, 3380, 3372, 3343]. Featured [2648]. Fechner [798]. feedback [504, 1576]. Fenchel [2032]. few [775, 737]. Fibers [3018]. Fibonacci [2780]. field [2770, 1787, 1665, 2979, 2752, 1597]. Fields [2279, 2592, 2026, 3427, 2629, 2942, 864, 2943, 936, 1388]. Fieller [70]. filling [3114]. filter [490, 1760, 1872, 2968]. Filtering [2215, 2218, 2951, 2971]. Filters [2220, 3304, 3246]. Finding [2715]. finely [1006]. Finite [2325, 2057, 2258, 529, 2026, 1933, 2059, 2557, 1477, 1482, 1356, 1115, 1610, 2589, 2778, 522, 606, 2602, 3338, 3211, 1316, 2501, 2923, 1014, 3023, 915, 530, 1146, 3329, 3093, 1395, 621, 3392, 356, 505, 610, 2463, 3151, 1209]. finite-sample [1395]. finitely [3325, 306]. Finiteness [2750]. First [2599, 2126, 2036, 2041, 2067, 299, 1841, 1907, 2612, 35, 1264, 3260, 1560, 403, 1205]. first-passage [2612]. Fisher [2326, 2948, 1566, 28, 1870, 2575, 2434, 1220, 1521, 3115, 2069, 809, 2038, 3103, 2810]. Fisher-scoring [3103]. Fit [2110, 2174, 2308, 2252, 2271, 2237, 2315, 1575, 3119, 595, 782, 691, 3238, 974, 2070, 3037, 3413, 2894, 3107, 3275, 1589, 158, 770, 1389, 389, 598, 1805, 3181, 2475, 2360, 2817, 2704, 2782, 1259, 1928, 2406, 3412, 409, 458, 1559]. fits [2477]. Fitting [538, 2298, 2578, 2043, 2581, 2147, 1098, 2098, 556, 643, 674, 996, 113, 957, 956, 934, 288]. five [2612]. Fixed [2256, 3145, 3284, 532, 1705, 2016, 3358, 2010, 2047, 2034, 2062, 2624, 3054, 3119, 3373, 2440, 3126, 526, 1355, 311, 3235, 2425, 1980, 3361, 156]. fixed-accuracy [3054]. fixed-cycle [156]. Fixed-Width [2016, 2062, 3145, 1705, 1355, 1980]. flanking [632, 724, 961]. flat [2683]. flat-top [2683]. Flats [2161]. flattening [2842]. Flexible [3382, 3279, 3156, 2522, 2850]. floating [1322]. fluctuations [139, 1690]. fluid [1758]. Fold [3129]. Fold-up [3129]. folding [1045, 1023]. follow [1733]. follow-up [1733]. footrule [2508]. forcing [2653]. Forecasting [2401, 2919, 3226, 2653, 3225, 85, 2864, 3224, 3227, 1679]. forecasts [3419, 1859, 1906, 2716]. foreign [2381]. foreshocks [1676]. forest [347, 98]. form [1035, 2765, 2701, 2496, 412, 694, 721, 761, 1699]. Forms [2790, 2040, 1238, 2730, 1948, 3037, 2486, 1260, 1480, 1239, 1348, 1435, 1753, 1141, 3, 1972, 1074]. Formula [2321, 6, 12, 577, 385, 1872, 929, 959, 2984]. Formulae [2011, 1948]. formulas [1917, 763, 834, 368, 2445, 742, 791]. formulation [1504, 315]. Forward [3391, 3287]. four [2790, 1054, 1020, 403, 963]. four-word [2790]. Fourier [2065, 2665, 1334, 3038, 7, 13]. fourth [1971]. FPE [1169, 1387]. fraction [1700]. Fractional [2313, 2746, 777, 3124, 2995, 1376, 1453, 776, 836, 862, 1018, 1361, 837, 3020]. fractionally [2605]. fractions [3064, 400]. fragility [2887]. fragmentary [1428]. Fragments [6, 12]. Frailty [2312, 3201, 2883, 2986, 3318]. framework [3352, 3175, 1524, 3351, 1642, 2885, 3399]. fraud [3369]. Fréchet [3133]. free [803, 475, 714, 486, 901, 135, 346, 3321, 3142, 2880, 2650, 927, 3421, 3343]. Freedman [2717]. freedom [244, 4, 3112]. French [229, 222, 304, 338, 352]. frequencies [2808, 2603, 512]. Frequency [1667, 277, 377, 3427, 3269, 3307, 35, 1225, 3150, 2667, 2381]. Frequentist [3186, 1808, 1782, 2721, 3262]. fresh [3138]. Full [2820, 1282]. Fully [2217]. Function [2551, 2963, 2018, 2190, 2057, 2591, 2516, 2174, 2305, 2237, 2080, 2534, 2334, 2528, 3255, 80, 262, 277, 377, 1350, 1016, 1755, 1577, 1325, 3360, 1216, 2837, 2416, 2668, 1581, 3264, 1319, 1725, 684, 2701, 3232, 2439, 1852, 2616, 1290, 1688, 3383, 2452, 3211, 779, 2454, 580, 792, 3327, 1968, 1227, 3267, 1499, 77, 731, 1355, 1627, 524, 758, 3295, 7, 13, 2506, 1083, 478, 3099, 953, 531, 1714, 1569, 3197, 1191, 2475, 999, 894, 1201, 1625, 398, 2374, 440, 1867, 1995, 746, 663, 601, 2830, 1737, 141, 1445, 722, 791, 356, 657, 1800, 1001, 357, 142, 2503, 2855, 710, 1849]. function [1862, 1902, 1155, 2474, 2981, 1557, 1749]. Functional [2699, 2480, 2083, 1060, 2224, 2511, 3158, 2822, 1061, 3265, 2512, 3188, 2792, 3337, 3341, 726, 1447, 3203, 3197, 3041, 2970, 1371, 484, 555, 588, 322, 687]. Functionals [2127, 2027, 2250, 1142, 3110, 1306, 2489, 494, 1831, 1571, 3348, 351, 2445, 1629, 1307, 2437, 2985, 3346]. Functions [2295, 2037, 2549, 2158, 2223, 2343, 2119, 2022, 2094, 2098, 2142, 200, 1245, 1142, 973, 2676, 330, 1984, 68, 2790, 1923, 1145, 1754, 1078, 892, 772, 691, 2798, 594, 684, 3247, 2761, 3370, 3175, 1334, 1779, 1178, 1661, 721, 1293, 1498, 147, 1532, 645, 427, 3295, 2570, 3129, 1041, 1401, 408, 941, 2881, 269, 3090, 2943, 1762, 853, 943, 34, 51, 3117, 91, 931, 3172, 811, 545, 599, 1591, 1129, 1381, 2641, 372, 484, 555, 588, 706, 354, 1230, 1775, 2554, 63, 1819, 2997, 2759, 1715, 1798, 1513, 3298, 3331, 2633, 3429, 1711, 1860]. Fundamental [2092, 1552, 591, 19, 111, 92, 602]. Further [578, 1731, 616, 1269, 291, 881, 742, 1074, 159, 70, 508, 78]. fused [3293]. future [1231]. fuzziness [1506, 1728]. Fuzzy [1952].

G [2107, 686, 841, 314, 1205].

G-peakedness [686]. gain [827]. Galton [1373, 502, 766, 1331]. Gambler [2779]. gambling [2620]. game [2653, 1065, 2864, 2707]. game-theoretic [2707]. games [2779]. Gamma [2317, 2343, 2190, 2020, 2666, 2151, 2342, 2312, 2594, 2504, 575, 297, 427, 1183, 2667, 1204, 1699, 1345, 1854, 3132, 1823, 1671, 42]. gamma-distributions [1854]. Gang [2325]. gap [190]. gappy [2760]. GARCH [3279, 3043, 3210, 3213]. GARCH-type [3043]. Gauss [3374, 984, 364, 461, 760]. Gaussian [1009, 1658, 417, 1012, 1090, 1111, 1564, 2699, 1950, 2221, 1450, 3352, 2962, 2700, 2761, 3422, 2217, 3293, 696, 3377, 1926, 2387, 695, 2308, 1540, 496, 303, 2570, 465, 1872, 2215, 2766, 3308, 2051, 943, 2667, 3311, 2064, 2288, 3273, 361, 3334, 3318, 2024, 492, 989, 2224, 682, 987, 1533, 2219, 2703]. Gaussian-sum [1872]. Gaussians [2742]. GEE [2636]. Gene [3410, 3106]. Gene-environment [3410]. General [1707, 2550, 3174, 2539, 2519, 2766, 1256, 2164, 2314, 2266, 2891, 3054, 557, 589, 632, 1363, 2490, 905, 1632, 2883, 2379, 2616, 3260, 2772, 856, 3276, 3432, 1997, 3213, 3430, 1541, 1822, 1377, 1766, 1936, 1446, 1307, 3159, 2759, 1424, 1567, 3346, 3343]. Generalised [2303, 3163, 288, 2812, 3132]. generalities [980]. generality [2814]. Generalization [2593, 897, 2612, 1774, 1485, 1788, 1075, 978, 683, 375, 1269, 1518, 1162, 987, 1009, 109]. generalizations [178, 1339, 463]. Generalized [2210, 2287, 1940, 2558, 2486, 152, 2279, 2327, 2361, 2128, 2333, 2661, 2051, 2584, 1789, 2697, 3318, 2024, 1928, 167, 2522, 1048, 2241, 2735, 2997, 3083, 2859, 1776, 2800, 3398, 3198, 3048, 2871, 1982, 3366, 3201, 28, 1647, 2421, 685, 578, 1726, 3175, 3124, 206, 677, 1107, 3045, 3044, 2466, 235, 1492, 3164, 1603, 3008, 2004, 3062, 679, 542, 3324, 1283, 3281, 1366, 294, 1302, 3072, 3190, 506, 720, 1659, 1745, 1387, 628, 1123, 163, 187, 666, 1483, 457, 2610, 1050, 995, 2362, 3306, 1232]. generalizing [2511]. generate [910, 2716]. generated [1889]. Generating [2516, 2094, 2837, 779, 3267, 3099, 1233, 281, 945]. Generation [3401, 1722, 990]. generator [137, 291]. generators [270]. genetic [1281]. Geometric [2212, 2171, 2023, 1568, 3335, 1651, 1612, 2386, 850, 1702, 1478, 1876, 513, 1891, 2617]. Geometric-Stable [2023]. Geometrical [1697, 1679]. Geometry [2155, 1670, 1208, 2770, 1408, 2571, 2757, 2742]. German [393]. Ghosh [1971]. gi [414, 727, 327, 1205]. GI/ [727]. GI/G/1 [1205]. GI/M/1 [327]. gi/m/2 [414]. Gibbs [2973, 3179, 3080, 2261]. Gibbs-type [3080]. Gibbsian [1596, 1562]. Gini [1337]. given [692, 1315, 3018, 267, 1704, 99]. Gleichungen [393]. Glivenko [429, 397]. Global [3304, 1313, 2636]. GLRT [2494]. GLSE [1175]. GMANOVA [808, 1412]. Gnanadesikan [1197]. Golay [1399]. Golf [2214]. Gompertz [2882]. good [3095, 3254]. Goodness [2110, 2070, 3413, 3107, 2174, 2308, 2252, 2271, 2237, 598, 2704, 2782, 2315, 1575, 3119, 782, 3238, 974, 3037, 2894, 3275, 1589, 770, 1389, 389, 3181, 2475, 2817, 1259, 1928, 2406, 3412, 1559]. Goodness-of-Fit [2174, 2308, 2252, 2271, 2237, 2315, 2070, 3413, 1575, 782, 974, 3037, 2894, 3275, 1589, 3181, 2475, 2817, 1259, 2406, 3412]. Govindarajulu [692]. Gower [974]. Grabarnik [3408]. grade [1698]. Gradient [2805, 195, 3431, 2853, 3029, 2432, 2457, 1312]. Gradual [3433]. Gram [1771]. Graph [2744, 2908, 2747]. graphical [3293, 3311, 977]. Graphics [2337]. graphite [84]. graphs [2908]. Graver [2908]. Grid [2026]. Gröbner [2727]. Group [450, 497, 508, 3237, 3385, 3384, 2614, 2727, 149, 3395, 3283, 530, 690, 3120, 3121, 1770, 1549]. Group-divisible [450]. group-testing [690]. group-wise [2727]. Grouped [2233, 738, 739, 1293, 278, 3068, 237]. grouping [3312]. groupings [106]. Groups [2269, 1769, 135, 22, 1008, 3194, 836, 83, 3173]. growing [3390]. Growth [2148, 2208, 2017, 2300, 905, 1616, 2389, 493, 554, 1732, 401, 1898, 2650, 1678, 1466, 1470, 1897, 3436]. Grubbs [1883]. guaranteed [2953]. Gumbel [1410].

H [463]. Haar [3297, 544]. habitat [1774]. Hájek [732]. Haldane [1320]. half [1747, 1982, 1093, 2471]. half-interval [2471]. half-sample [1747]. handling [333, 3409]. Hard [2261]. harmonic [1261, 1262, 1327, 1325, 2955]. Hastings [2706]. Hausdorff [671]. Having [2143, 377, 1014, 1489, 1660, 1308, 1432, 3376]. Hawkes [1025, 2861, 3323]. Hazard [2057, 3165, 2280, 1537, 2273, 2265, 2855, 2821, 3058, 3036, 1779, 3135, 2988, 3233, 2432, 1785, 2641, 1364, 2554, 3026, 1282, 1431, 2633]. Hazards [2165, 2702, 3048, 3287, 3212, 3152, 2870, 3349, 1688, 2640, 3189, 3372]. health [3054]. heaviness [1103]. Heavy [2075, 1758, 3357, 2900, 3043, 2726]. heavy-tailed [2900, 3043]. height [2652, 1678]. Help [11, 17, 30, 44, 58, 74, 79, 89, 97, 120, 146, 170, 177, 186, 194, 211, 212, 220, 223, 233, 242, 251, 275, 305, 329, 358, 394, 430, 470, 514, 569, 629, 675, 730, 783, 825, 877, 922, 964, 1013, 1059, 1105, 1151, 1207, 1253, 1297, 1349, 1400, 1457, 1473, 1490, 1505, 1519, 1535, 1550, 1565, 1579, 1595, 1608, 1621, 1634, 1648, 1662, 1677, 1692, 1710, 1724, 1739, 1756, 1773, 1791, 1809, 1825, 1842, 1857, 1871, 1887, 1904, 1918, 1931, 1944, 1959, 1974, 1988, 2002, 2015, 2029, 2042, 2055]. Help [2066, 2078, 2091, 2102, 2114, 2129, 2141, 2154, 2168, 2183, 2198, 2213, 2227, 2243, 2260, 2278, 2294, 2310, 2330, 2350, 2366, 2382, 2397, 2411, 2423, 2436, 2446, 2462, 2476, 2487, 2498, 2513, 2525, 2538, 2553, 2566, 2577, 2588, 2601, 2611, 2623, 2635, 2647, 2659, 2672, 2686, 2696, 2709, 2723, 2732, 2740, 2753, 2762, 2769, 2781, 2793, 2802, 2813, 2823, 2834, 2845, 2857, 2868, 2878, 2889, 2898, 2909, 2917, 2928, 2937, 2946, 2956, 2966, 2975, 2983, 2991, 3001, 478]. hemisphere [1665]. heredity [632, 3341]. Hermite [3177, 1334, 1789, 933]. Herriot [2826]. Heterogeneity [2341, 487, 1347, 1665, 1472, 3055]. heterogeneous [2618, 2737]. Heteroscedastic [2774, 2282, 1915, 1414, 1393, 3176, 745, 3430, 2595].

Heteroscedasticity

[2242, 1940, 2879, 3214, 2469].

heteroskedasticity [2765]. heuristic [290]. Hidden [2325, 2972, 2460, 3308, 3386]. Hierarchical [2558, 2148, 3024, 2438, 561, 623, 1252, 2638, 3006, 2715]. High [2083, 3312, 2028, 2948, 2676, 2990, 3193, 3212, 3229, 3307, 2370, 3040, 3391, 3094, 3305, 3150, 2959, 3008, 2924, 3005, 3351, 1435, 2667, 2969, 3273, 3209, 3309, 2938, 325, 2381, 3200, 3421, 3372, 3310, 3241]. high-dimension [3193, 3273]. High-dimensional [3312, 2948, 2676, 2990, 3229, 3040, 3305, 2959, 3008, 2924, 3005, 3351, 3209, 3309, 2938, 3421, 3372, 3310, 3241]. high-frequency [3307, 3150, 2667]. high-resolution [2969]. Higher [1574, 2645, 2683, 933, 1930, 2067, 2090, 810, 406, 1999, 1754, 1582, 2546, 1714, 1378, 2430]. Higher-Order [2067, 2090, 2683, 1582, 2546, 1714]. highest [1808, 1782]. Hilbert [3280, 3409]. Hilbertian [3249]. Hirotugu [2710]. historical [1486, 1693]. Hitting [1622, 199]. HKM [3254]. HNBUE [2510, 1440, 1326]. Hodges [767, 895, 3215]. Hoeffding [140]. Hoel [1518]. homeostasis [1363]. Homogeneity [2347, 2520, 902, 248, 1531, 1927, 3194, 387, 1179, 176, 2562]. homogeneous [3248, 1008, 1889, 815, 2831]. homoscedastic [2660]. homotopy [3375]. honor [2710]. Hooke [887]. Hotelling [807, 578, 677, 1107, 1064, 121]. hulls [1889]. Hybrid [2718, 3169, 2373, 1417, 3367]. Hyperbolic [2584]. hypercube [1518]. hypercube-generalization [1518]. hypercubic [1069, 979]. hypergeometric [892, 913, 1183, 618, 1048, 1123, 722, 791, 1798]. hypergraphs [3107]. hyperparameter [1076]. hypersphere [1234]. Hypotheses [2205, 2156, 2140, 2048, 2072, 1081, 1619, 2914, 1503, 1837, 254, 258, 745, 644, 81, 2697, 518, 599, 1442, 1770, 896, 474, 991, 907]. Hypothesis [3305, 1202, 857, 454, 1651, 667, 763, 834, 926, 365, 1681, 449, 748, 3428, 742, 2827, 3173, 3253]. hypothesis-decision [454]. hypothetical [389].

i.i.d [1274]. I.I.D. [1645, 930, 954, 2286, 2890, 1701]. ideals [2748, 3311]. identical [1707, 909]. Identically [2328, 1543, 916, 828, 1479]. Identifiability [1477, 3131, 3415, 3418, 3417, 3416, 1922, 1200, 3109, 3023, 2786]. identifiable [3348]. Identification [1714, 2977, 417, 504, 537, 591, 590, 2854, 3300, 1458, 963]. Identifying [3314, 3158]. identities [1707]. identity [692, 382, 316, 602]. IFR [532]. Ignorance [1076]. Ignored [2242]. II [1187, 588, 614, 655, 1998, 2891, 2702, 589, 1007, 2373, 739, 3275, 826, 2084, 258, 279, 349, 1997, 3099, 681, 352, 294, 3154, 1161, 264, 293, 555, 2047, 728, 2259, 1426, 804]. II. [1537]. IID [1182]. III [656, 632, 1295, 301, 3099, 747]. illustrations [3054]. illustrative [2974]. Image [2592, 1635, 3371, 1826]. imbedding [2923]. immigration [1373]. imperfect [3231]. implementation [1414, 1872]. implication [295, 2449]. implications [2790]. implies [1971]. Import [2130]. importance [3149, 2649, 2935]. Improper [3325, 1783, 1437]. Improved [2591, 2687, 1339, 1108, 1981, 2268, 2907, 1864, 1544, 3306, 3292, 1911, 1936, 512, 1483]. Improvements [2133, 2406, 1659]. Improving [2346, 2471, 3103]. imputation [3262, 3256, 2437]. inaccuracy [746]. Inadmissibility [1500, 341, 1382]. Inagaki [732]. incense [2651]. incidence [1363, 3370]. incidental [2798]. including [1392]. income [3104]. incomparability [564]. Incomplete [2578, 2087, 2348, 809, 359, 1592, 1390, 3424, 812, 279, 284, 507, 449, 822, 863, 1324, 2827, 3091]. incompletely [3386]. inconsistency [1733]. Incorporating [1486, 2661]. incorrectly [1444]. increasing [2996, 1236, 2359, 540]. Increments [2321, 815]. indefinite [1972]. Independence [2529, 2347, 2151, 2727, 1836, 1590, 1390, 488, 3082, 3347, 317, 615, 2470, 2789, 947, 1328, 2, 103, 680, 842, 3, 757, 1020, 3209, 2681, 4, 543, 2749]. Independent [2321, 805, 1543, 1707, 1134, 807, 773, 128, 209, 1, 873, 1345, 828, 447, 1479, 596, 587, 176, 1989, 1823, 932, 543, 815]. Index [2035, 2045, 3258, 473, 183, 2955, 3174, 3152, 2808, 2887, 2628, 2689, 3043, 3210, 3300, 3086, 3078, 3009, 3218, 3041, 1337, 3171, 3170, 3098, 2726, 3219, 3026, 2911, 3173, 3253, 3111, 2833, 2703]. indexed [1551]. Indicator [2549, 2790]. Indices [290, 2731]. indifference [1149, 1202]. Indirect [2439, 3266, 118]. individual [135, 227, 518, 1678, 187]. Individuals [2148]. induced [1969, 1088]. Inequalities [2244, 2339, 1396, 1706, 2156, 348, 959, 1364, 1230, 1870, 1265, 2962, 3272, 296, 616, 3239, 900, 22, 174, 1652, 2864, 2794]. Inequality [2317, 386, 38, 3088, 1003, 1051, 797, 2725, 858, 888, 178, 1521, 876, 1038, 284, 1512]. infectious [3222]. Inference [3058, 2514, 1373, 3373, 3163, 2075, 3232, 2035, 2541, 2239, 3387, 2230, 3362, 2918, 3183, 942, 2322, 2074, 3434, 2671, 1742, 2643, 2563, 3379, 1428, 2373, 2480, 2761, 3381, 3157, 3076, 2952, 2884, 2965, 3155, 3272, 2811, 3169, 3043, 1034, 1401, 3199, 1843, 3197, 2988, 3041, 3223, 1762, 8, 14, 2969, 1993, 3324, 2754, 1766, 3122, 2989, 3016, 2971, 3361, 1641, 597, 1686, 2860, 2637, 1221, 3380, 3375, 2437, 3306, 3409, 1671, 3346, 2633]. Inferences [1618, 2509, 3137, 2870, 2835, 3104, 2899, 3154, 2820, 1711]. Inferential [1805, 2980]. inferring [758]. Infinite [2212, 1232, 1115, 2960, 2679, 3243, 2680, 2470, 529, 1053, 2415]. infinite-dimensional [3243]. Infinitely [2234, 853, 943, 372, 354]. infinitesimal [2390]. infinity [319]. inflated [3078]. inflation [1488]. Influence [2099, 2209, 3056, 2208, 2297, 2180, 2370, 1532, 2703]. Information [2326, 2244, 2339, 1268, 1729, 2236, 1582, 319, 2302, 2204, 2232, 2251, 2038, 2543, 2005, 1970, 2891, 1219, 2994, 1566, 2922, 2978, 1870, 2892, 1267, 2842, 3422, 2517, 1133, 1198, 2575, 1506, 1767, 2757, 3272, 485, 631, 948, 925, 1540, 2434, 479, 210, 231, 254, 258, 274, 280, 286, 2398, 1220, 1521, 1486, 2722, 650, 3115, 3005, 889, 919, 1152, 1365, 1541, 1642, 1224, 630, 3093, 2893, 1206, 2862, 3148, 809, 342, 456, 436, 154, 2720, 1891, 1812, 2810, 3073, 1553]. information-geometric [1891]. Information-theoretic [2543, 1642]. Informative [2507, 2416]. informatively [3414]. inhomogeneity [1459]. Inhomogeneous [2145, 3386, 2504, 3069, 3261, 3101]. INID [3154]. initial [1604]. Innovation [2296, 3269]. Innovations [2075, 2944, 3043, 3210, 2426]. Input [2239, 377, 108, 117, 166]. input-output [108, 117, 166]. inputs [1204]. inspection [180, 899, 203]. instability [3313]. instance [3389]. Instrument [3302]. Instrumental [2859, 2505, 2688]. instruments [2879]. insurance [1866]. intake [1088]. integer [3279, 3276, 3213]. integer-valued [3279, 3276, 3213]. Integers [2547]. Integral [3275, 833, 2594, 2315, 7, 13, 1700, 1276, 1256, 886, 18]. integrals [1548]. Integrated [1181, 2237, 2583, 1737, 2932, 1504, 1126, 3060, 2605, 2985, 1067, 1336]. integration [2488]. integro [182]. integro-differential [182]. intensities [1614]. Intensity [2119, 1577, 3092, 2690, 3261, 2468, 2754, 55, 1185, 3297, 2831, 3429]. inter [650]. inter-column [650]. inter-row [650]. interaction [1932, 1632, 3410, 1214, 781, 1861, 1053, 1136, 1562, 1686]. interactions [3395, 745, 2979, 1708, 3138]. interactive [3040]. interarrival [314, 327]. intercept [3285]. interchangeability [1587]. interchangeable [1019]. interclass [1196, 1618]. interdeparture [947, 1328]. Interest [2218]. Interior [2093]. Intermediate [2271, 1523, 1777]. internal [95]. international [222]. internationales [222]. interpoint [1701]. interpolation [336, 349, 1285, 215]. interpolator [2364]. Interpretation [1786]. intersection [156]. intertemporal [116]. Interval [2116, 2539, 2165, 2139, 1424, 2247, 2062, 3145, 3054, 3266, 263, 315, 3414, 2471, 1827, 969, 1112, 849, 238, 1772, 1096, 267]. Interval-Censored [2247, 3414]. Intervals [2033, 2560, 2016, 875, 2249, 2157, 2047, 272, 2615, 1705, 740, 1355, 619, 2907, 947, 1328, 1465, 1980, 1835, 1659, 2641, 168, 881, 3006, 1896, 2540, 1212, 1513, 1514]. interviewing [66]. intra [821]. intra-class [821]. Intrablock [308]. Intraclass [2238, 240, 249, 859, 1806, 1304, 1618]. Intrinsic [2999, 2048, 2313]. Intrinsically [3035]. Invariance [2108, 1613, 2730, 103, 1529, 1696]. Invariant [2018, 1055, 2293, 1114, 2624, 1358, 1135, 1488, 1806, 472, 529, 530, 1820, 1195, 1557, 1749, 544]. inventory [1563]. Inverse [2308, 2288, 2024, 2017, 2269, 1658, 1111, 1564, 2612, 1950, 1450, 3066, 3264, 1926, 1540, 806, 2667, 900, 2625, 361, 3318, 693, 729, 1572, 3399]. inverse-Gaussian [2667, 3318]. inverses [628, 2610]. invertible [2605]. Investigation [551, 1185, 95, 794]. Investigations [489, 147]. investment [2716]. involved [645]. involves [1382, 526]. involving [1723, 144, 382, 316, 3267, 527, 822, 506, 1329, 1420, 1823]. irradiation [118]. Irrational [2548]. irregular [1694]. Islands [41]. Isoseparation [2860]. Isotonic [903, 2677, 2369, 3348]. Issue [2215, 2710]. issues [3131, 1266, 3055]. item [3376]. Iterated [2201, 1955, 173, 2864]. iterates [726]. iteration [975]. Iterative [2032, 2341, 729]. IV [1376].

J. [1971]. Jackknife [3162, 1685, 3077, 2390, 3090, 1495, 1746, 3123]. Jackknifing [1745]. Jacobi [618]. Jaeckel [1147]. James [2473, 2604]. James-type [2604]. Japan [1266, 1730, 304, 193]. Japanese [221, 824, 826, 1665, 338, 352, 963]. Japon [304]. japonais [338, 352]. Jensen [1512]. jittered [1947]. Jodges [750]. Joint [1907, 2036, 3212, 2185, 2011, 3377, 2131, 2419, 2484, 2598, 2546, 2207, 2067, 2469, 1235, 102, 939, 1852, 1883, 1839, 678, 1236, 1063, 749, 844, 1124, 1690, 540, 2929]. judgment [2809, 2925, 3122]. Jump [2117, 2581, 2694, 2961, 3146, 3425, 3304, 3167]. Jump-Preserving [2581, 2694]. jumps [3022]. justify [1673].

K. [128]. Kac [594, 684, 1144, 712]. Kac-statistics [712]. Kalman [1760]. Kaplan [3184, 2375, 2272]. Kernel [2551, 1909, 2490, 1977, 3046, 2345, 2280, 2246, 2164, 1285, 3079, 2718, 1755, 3280, 3243, 1601, 2413, 2489, 3317, 1290, 1865, 3295, 1656, 2847, 2570, 2371, 2812, 2801, 3204, 2626, 3233, 3273, 894, 1655, 1395, 1129, 1737, 1445, 1638, 1994, 1602, 3244, 2824, 2502, 2692, 3208, 3409, 2378, 1908, 2613, 1336, 3429]. kernel-assisted [3208]. Kernel-type [1909, 894, 1737]. Kernels [2190, 2299, 2007, 3019, 1736, 1978, 3091]. Khatri [841]. Kiefer [1052, 2060]. kind [271]. kinds [5]. Kinsaku [171]. KL [2503]. KL-divergence [2503]. known [1802, 1540, 3061, 542, 958, 283, 734, 566, 1096, 899, 1514]. Kolmogorov [897, 2407, 751, 2018, 691, 175, 465, 1037, 1167, 457, 321, 2038, 2318]. komplexen [393]. kriging [1638]. Kronecker [1472]. Kullback [2892, 2874, 2236, 3051, 231, 280, 286, 2532, 2232, 2852, 1828]. Kurtosis [2288, 734].

Lack [2306, 1126, 1118, 1040]. LAD [2807]. lag [276, 2683]. lag-windows [2683]. Lagrange [2905]. LAN [1938]. landscape [3434]. Langevin [1605, 2097, 1704]. language [26]. Laplace [1672, 3413, 2308, 2396, 2340, 2547, 1700, 2360, 18, 109]. Large [2515, 1272, 3262, 1147, 2806, 2095, 2933, 976, 1698, 189, 1234, 510, 1799, 3051, 2942, 3002, 2901, 428, 1041, 2653, 3063, 462, 3194, 1817, 2378, 1567, 2794]. Large-sample [976]. larger [698]. largest [446, 581, 1258, 516, 1108, 546, 635, 682, 543]. Lasso [2957, 2736, 2913, 3283, 3281, 3237]. Lasso-type [2736]. last [3287, 1205]. last-come-first-served [1205]. Latent [2712, 439, 1415, 939, 972, 952, 431, 515, 535, 676, 971, 743, 860, 3415, 3418, 940, 953, 1108, 665, 681, 3417, 3416, 869, 999, 546, 682, 1443]. latent-variable [3415, 3418, 3417, 3416]. Later [2104, 2184, 1999, 1920]. Lattice [2279, 1323, 557, 2979, 2969, 548]. Law [2048, 1742, 1363, 1530, 2666, 2653, 1435, 173, 20, 1280, 563, 627, 2864]. Laws [2561, 2051, 2132, 2201, 2023, 2628, 1885, 1690, 1381, 2814]. Layout [2578, 2450, 1214, 1927, 1273]. layouts [2477, 871]. LBI [1836]. LCFS [1613]. learned [3431]. Learning [2311, 2377, 523, 607]. Least [2061, 2346, 2096, 3405, 2176, 2385, 1321, 1741, 970, 2452, 3425, 390, 1354, 2872, 1248, 112, 1560, 2812, 2444, 1388, 1546, 3127, 912, 2759, 3385, 3384, 1415, 788, 2657, 1286, 1633, 2903]. least-squares [3127, 2759]. leave [2993]. leave-one-out [2993]. Lebesgue [2645]. Lecture [3353, 3356, 3354, 3355]. Leffler [3296, 2943, 1591]. Left [2165, 805, 3301, 3165, 1488, 2873, 3135, 2682, 2735, 2467]. left-truncated [805, 3301, 2873, 3135, 2682]. Lehmann [750, 767, 1153, 1187, 895, 1342]. Leibler [2892, 2236, 3051, 231, 280, 286, 2532, 2232, 2852, 1828]. lemma [633, 92]. Length [2210, 2256, 2175, 2010, 2090, 2441, 2451, 3034, 1884, 2639, 162, 267, 2728]. length-biased [2451, 3034, 2639, 2728]. Lengths [2011, 2598, 2408, 2484]. lens [2636]. Leontief [108, 117, 166]. level [2746, 2550, 2549, 2790, 3051, 553, 965, 1673, 708, 3351, 1781, 899, 93]. levels [165]. Levenberg [1313]. Lévy [1174, 3271, 2222, 3422, 3396]. Lévy-Driven [2222]. Life [2057, 2236, 2314, 309, 370, 2365, 1417, 1964, 1266, 158, 175, 1311, 1326, 1869, 1310, 1963, 266, 285]. life-testing [309]. Life-Time [2057, 1964]. Lifetime [1958, 1383, 2835, 3327, 3034, 1963, 2927, 3030]. lifetimes [1311, 1563]. light [1758]. Lightly [2051]. Lights [156]. like [1607, 3333]. Likelihood [2033, 2158, 1370, 2096, 2319, 1939, 2541, 2347, 1837, 2279, 2031, 2340, 2058, 2532, 2064, 2270, 1459, 1562, 2249, 2191, 2250, 2437, 2557, 2556, 2324, 1017, 935, 1090, 3365, 2994, 1016, 3162, 330, 2800, 3071, 3230, 1829, 1855, 905, 647, 1813, 3269, 3302, 2373, 3216, 2874, 1274, 2829, 846, 1799, 1133, 1264, 1976, 2369, 856, 634, 1661, 386, 951, 1456, 2832, 1589, 2815, 3082, 731, 864, 1318, 3304, 524, 1861, 1801, 1341, 410, 756, 1997, 2995, 2764, 1843, 3333, 2873, 2492, 2924, 3090, 2640, 3125, 1764, 644, 1256, 1337, 1697, 1822, 3183, 1422, 680, 842, 3324, 3290, 1283, 2499, 3289, 3288]. likelihood [936, 988, 1136, 353, 2697, 1025, 1876, 2443, 2816, 1818, 828, 2971, 2993, 2916, 3209, 1835, 2899, 492, 705, 1057, 2820, 2987, 2363, 541, 585, 2449, 2641, 320, 1273, 1391, 1989, 2930, 1930, 2724, 2733, 3185, 1715, 1891, 204, 2392, 2503, 3306, 1966, 1282, 1301, 1421, 3077, 3123, 2794, 2788, 2633, 2399, 2703]. Likelihood-based [2541, 2437, 3324, 2899]. likelihood-ratio [634]. Likelihoods [2295, 2142, 1923]. Limit [2127, 2321, 2628, 1300, 1701, 1997, 2152, 1263, 3119, 1578, 1143, 1344, 2489, 1852, 2394, 502, 766, 323, 319, 873, 140, 2357, 1674, 587, 799, 76, 104, 10, 16, 932, 1132, 3322]. limit-theorem [319]. limited [2994].

Limiting [1496, 2523, 3133, 1278, 2365, 2808, 2332, 2738, 3209, 1553, 3240, 805, 225, 2564, 268, 1921, 1969, 1260, 639, 1360, 571, 642, 567, 306, 1464]. Limits [2051, 125, 3325, 198]. LIML [2879]. line [3184, 1093, 113, 1961, 3085]. Linear [1094, 2053, 2338, 2343, 2303, 2299, 2075, 2428, 992, 2070, 2516, 2541, 2581, 2239, 2115, 2199, 2021, 2156, 2305, 2128, 2080, 2085, 2522, 2149, 2040, 2025, 2109, 2177, 1148, 2180, 2320, 1594, 2859, 1021, 107, 417, 504, 537, 459, 2800, 2821, 68, 1619, 2822, 491, 2354, 1610, 1802, 1061, 2934, 2964, 489, 868, 647, 1249, 3373, 1405, 1813, 1257, 916, 3278, 1726, 3175, 2952, 3341, 667, 763, 834, 1133, 2791, 1675, 2369, 1929, 3124, 3107, 926, 1977, 773, 3040, 3228, 2690, 1667, 2652, 2680, 2763, 3415, 3418, 2484, 46, 1398, 758, 1626, 311, 427, 3321, 1447, 2904, 3203, 2661, 269]. linear [2364, 3009, 2988, 1764, 644, 3417, 1298, 3068, 90, 3416, 2468, 931, 3033, 3183, 1277, 1882, 3195, 958, 1185, 3281, 2639, 2438, 1551, 1371, 1867, 1302, 2393, 3072, 506, 1573, 2363, 663, 1745, 621, 3127, 1936, 1770, 1303, 461, 760, 1284, 1242, 1446, 742, 2638, 789, 179, 2735, 2759, 1424, 3385, 3384, 2383, 3375, 357, 1415, 2675, 788, 932, 1263, 1716, 2362, 2392, 2353, 2662, 3200, 3306, 3411, 710, 1704, 1893, 3303, 3336, 3105, 2921, 3025, 3421, 1301, 3077, 2788, 1765, 2911, 3083, 3253, 2954, 3153, 1695, 2703, 3241, 1711, 1860, 1435]. linear-fractional [3124]. linear-loss [789]. Linear/ [1435]. Linearity [2587, 144, 2965]. Linearization [1991]. linearized [647]. linearly [464]. lines [25, 980, 1007, 1295, 1622]. Linex [2206, 1625]. literacy [24]. Local [3342, 2573, 2335, 2821, 2234, 2298, 2822, 1991, 2589, 2299, 2537, 2007, 1929, 2581, 2099, 2607, 2949, 2654, 2345, 2532, 2147, 2045, 1905, 2164, 1682, 2843, 1342, 2292, 1162, 2098, 2703, 1143, 1390, 2773, 1651, 2884, 3221, 3220, 502, 766, 2853, 3029, 873, 2572, 701, 762, 819, 2499, 2663, 2724, 3331]. Local-to-finite [2589]. localement [229]. localization [3268, 3252, 3308]. Locally [558, 1487, 608, 583, 2284, 2269, 3265, 229, 1454]. locally-convex [229]. Located [2239]. Locating [1201]. Location [2179, 2517, 2231, 2080, 2284, 2038, 2323, 1154, 1356, 3330, 659, 1346, 1484, 1804, 1428, 467, 218, 1516, 1488, 2738, 2806, 2734, 1865, 2982, 2574, 1685, 1585, 2471, 3117, 736, 574, 701, 3214, 2993, 476, 1172, 895, 752, 287, 402, 1734]. location-parameter [701]. location-scale [2734, 2574, 3117, 3214, 1172]. Locations [2080]. loci [2979]. locus [2999, 2896]. Loewner [482, 444]. Log [2295, 2031, 2058, 1511, 2675, 1016, 2369, 3107, 1256, 1095, 492, 2457]. Log-concavity [1511]. log-likelihood [1016]. log-likelihood-ratio [492]. Log-linear [2675, 2369, 3107]. log-normal [1095]. log-odds [2457]. Logarithm [2201, 173, 2864]. Logarithmic [2335, 101, 1319, 1599]. Logistic [2118, 2676, 1982, 1709, 1075, 1117, 1451, 1073, 1150, 3016]. Logrank [2073]. Long [2298, 2585, 2199, 2130, 2583, 2400, 2426, 1088]. Long-Memory [2298]. Long-Range [2199, 2400, 2426]. longest [3100]. longitudinal [2846, 3218, 1073, 3072, 2930, 3299, 3306, 2844, 3234, 2929, 3137]. look [1651, 3138]. looking [1089]. loop [1781]. Lorenz [1337, 2899, 2555]. Loss [1970, 2891, 2018, 2266, 2267, 1919, 2404, 1382, 1506, 1178, 1603, 1191, 3172, 1625, 373, 1483, 789, 1127, 1113, 3411, 2474]. Low [490, 3104, 3239, 3273]. low-complexity [3239]. low-income [3104]. low-sample-size [3273]. Lower [3096, 1242, 1448, 419, 2603, 2424, 908, 2642, 624, 711, 367]. LR [1913]. lymphatic [1792].

M [2679, 3197, 327, 3409, 414, 2107, 727, 314].

M-based [3197]. M-estimation [2679, 3409]. M. [684]. M/ [727]. M/G/ [2107]. M/G/1 [314]. M4 [2617]. MA [949]. machine [3273, 129]. macromolecules [1798]. Mad [2044]. Maesono [2304]. magnitude [1649]. Mahalanobis [685]. Main [604, 838, 871, 1599, 548, 624, 711, 584]. maintenance [1740]. majorant [2385]. Making [2711, 923]. manifold [2569, 2896, 2576]. manifolds [2869]. manipulation [1786]. Mann [2358, 2906, 261]. MANOCOVA [793]. MANOVA [1058, 519, 3263, 940, 401, 362, 1686, 700, 1606]. manufacturing [967]. Many [1762, 2879, 551, 306, 63]. many-dimensional [63]. map [1572]. margin [2478]. Marginal [3234, 1776, 1468, 939, 3278, 2666, 2656, 2488, 3082, 1401, 1569, 1531, 2817, 328, 3421, 2797]. marginals [2624, 1315, 2509, 3018]. margins [2887]. Marked [2527, 2528, 2825, 3069, 3035, 1905, 1459]. market [3307, 2715]. markets [3146]. Markoff [45, 182]. Markov [640, 2079, 1999, 2104, 1016, 2127, 2037, 2122, 2256, 2624, 2727, 2012, 3366, 782, 2175, 2490, 2837, 2307, 3374, 984, 814, 2332, 2530, 3107, 2211, 2460, 2084, 2923, 2279, 2598, 2949, 319, 2592, 2546, 1941, 2010, 2619, 2839, 3259, 1884, 503, 2972, 2270, 936, 626, 2912, 1598, 705, 994, 1057, 2751, 520, 938, 2445, 2430, 3386, 3018, 364, 461, 760, 306, 371, 2412, 1087, 1920, 2067, 2090, 2861, 1068]. Markov-Dependent [2332]. Markov-modulated [2861]. Markov-sequences [1598]. Markovian [817, 2607, 757, 2381]. marks [2839]. Marquardt [1313]. Marshall [1436]. Martingale [2147, 1350, 2464, 2940]. masked [2838]. masking [2838]. Mass [1994, 1014]. massive [3345]. Matching [2157, 2351, 52, 2885, 2902]. material [239]. Mathematical [1038, 348]. mathematico [23, 39]. mathematico-statistical [23, 39]. mathematics [557, 589, 632]. matings [1292]. matric [1238, 1239, 1348]. matric- [1238, 1239, 1348]. Matrices [2148, 2155, 2080, 2268, 892, 1029, 2370, 1135, 1720, 3192, 1030, 792, 859, 365, 3094, 1393, 376, 438, 516, 756, 3063, 869, 680, 819, 842, 958, 545, 546, 3433, 628, 444, 482, 2491, 2610, 3004, 2378]. Matrix [2519, 263, 2086, 3393, 2302, 2580, 2267, 1012, 3296, 586, 2655, 1684, 868, 1358, 2765, 1055, 1357, 2761, 952, 439, 431, 515, 535, 676, 971, 860, 2850, 3002, 953, 1108, 1644, 652, 298, 1603, 3061, 1820, 1109, 1541, 911, 999, 1495, 1898, 440, 635, 3093, 2417, 693, 729, 238, 809, 912, 2742, 666, 722, 791, 1312, 3412, 1798, 3399, 3047]. Matusita [1082, 1166, 642, 770, 1037, 718, 818]. max [224]. maxima [2449, 880]. Maximal [1145, 673, 651, 1874, 2752]. maximization [2818, 989]. Maximizing [2993, 2874, 1661]. Maximum [1829, 3216, 2829, 768, 2220, 1801, 1341, 2340, 2995, 2640, 2202, 2064, 1025, 1903, 2009, 705, 2105, 2076, 2259, 442, 592, 2556, 2324, 1140, 1182, 1646, 935, 3365, 2994, 905, 647, 1274, 846, 1799, 1264, 1976, 2369, 386, 431, 535, 731, 1318, 410, 1997, 2979, 1764, 1337, 1697, 1283, 988, 1136, 1818, 828, 1180, 930, 1057, 2987, 938, 585, 320, 1954, 1930, 2733, 204, 904, 1282]. Maxwell [3182]. may [1583]. MCLE [3052, 1272]. MCMC [2974]. Mean [2325, 2203, 2057, 2151, 2087, 2583, 1596, 2249, 2097, 2047, 2108, 2177, 2334, 2344, 1695, 1336, 1181, 1658, 2948, 445, 446, 581, 735, 1723, 2695, 3158, 102, 2655, 370, 1002, 1894, 1754, 2365, 3319, 1515, 2998, 1340, 637, 35, 2421, 2856, 696, 1147, 1975, 2494, 1545, 1605, 3094, 3347, 2896, 231, 280, 286, 1404, 1271, 1877, 384, 2687, 1339, 1202, 1653, 1689, 1555, 1820, 3197, 1191, 937, 551, 2472, 790, 646, 1064, 1980, 3172, 680, 842, 1079, 1095, 1241, 1333, 958, 1441, 1625, 753, 1717, 959, 870, 1247, 1544, 83, 1096, 341, 383, 356, 371, 471, 539, 142, 737, 2437]. mean [1704, 2474, 3331, 2902, 2929, 1338]. mean-covariance [2902, 2929]. mean-shift [2998]. meaning [2896]. meaningful [3263]. Means [2013, 2062, 2181, 1556, 3365, 106, 3133, 49, 3028, 50, 1705, 2458, 768, 1645, 517, 965, 2999, 1657, 1486, 3035, 1558, 1864, 448, 3289, 3288, 1211, 1835, 83, 582, 464, 975, 471, 613, 3007, 1583, 3047, 1210]. measurable [565, 529, 530]. Measure [2338, 1081, 1082, 1166, 1964, 1092, 1193, 1599, 697, 210, 254, 258, 274, 1199, 642, 770, 1981, 853, 718, 1815, 818, 544]. Measurement [2522, 2180, 2859, 3301, 983, 2911, 3083]. Measurements [2774, 2194, 180, 3062, 1443, 3436]. Measures [1346, 2049, 2892, 2798, 564, 1611, 1037, 529, 530, 1183, 2495, 1152, 528, 2438, 732, 2721, 2609, 1165, 3128, 576, 1103, 1553]. Measuring [2651, 3097]. mechanism [727, 1244, 3343]. mechanisms [3367]. Median [3092, 2483, 1803, 1910, 1117, 1300, 1320, 282, 1119, 165, 204, 215, 2736, 2633]. Median-based [3092]. Medians [2060, 2763]. medical [3030]. medicine [95]. Meier [3184, 2375, 2272]. Meijer [3317]. Meixner [1183]. Mellin [3317]. membership [1698]. Memorial [3353, 3356, 3354, 3355]. Memory [2298, 2583, 2306, 1126, 1118, 1040]. Merging [2707]. merits [3215]. mesh [878]. meta [3387, 3333]. meta-analysis [3387, 3333]. metallurgy [1954]. Method [2065, 2057, 2591, 2140, 2194, 1654, 107, 172, 195, 537, 935, 1577, 1991, 2807, 1319, 1414, 3413, 1379, 856, 1661, 1093, 345, 481, 1044, 2396, 303, 1393, 178, 850, 1175, 278, 108, 1217, 2818, 1, 3116, 1960, 64, 5, 199, 2636, 393, 630, 2499, 1609, 2368, 2620, 878, 1233, 85, 547, 1746, 281, 664, 1071, 245, 1312, 907, 3396, 2733, 1415, 3185, 2953, 977, 1089, 60, 2788, 2399, 2613]. Methode [393]. methodologies [2739]. Methodology [334, 1557, 3054]. Methods [2079, 2215, 2008, 2204, 1594, 1036, 1181, 1017, 1747, 2980, 3074, 1607, 3301, 510, 333, 2379, 3275, 986, 1649, 3038, 2949, 157, 2636, 3334, 693, 729, 780, 512, 477, 1313, 995, 945, 2797, 182, 3103, 1633]. metric [2869, 3115, 2395]. metrics [2433, 3106]. metrization [59]. Metropolis [2396, 2706]. Microbiology [157]. microstructure [3307]. midrange [1451]. Midzuno [38, 688]. Migration [963]. milieu [338, 352]. min [224]. min-max [224]. Minimal [2624, 2745, 279, 234, 3359, 942, 2412]. Minimax [2234, 1899, 3036, 2007, 1848, 472, 315, 630, 1734, 1978, 2474, 3020, 1749, 1856, 1357, 3102, 1487, 1228, 1271, 57, 78, 1191, 1387, 896]. Minimaxity [1404, 1333, 1674, 3004, 1390, 2472]. minimisers [2482]. minimize [112, 1042]. minimizers [2503]. Minimizing [2038, 3051, 1599, 2495]. Minimum [1878, 2084, 2095, 2910, 2233, 2085, 398, 1965, 1715, 2177, 2113, 966, 1219, 2978, 847, 1072, 1386, 1750, 1767, 1545, 743, 3339, 1701, 1714, 1317, 2471, 64, 1821, 1201, 1096, 3006, 60]. mining [2648]. Minke [1665]. Mirakyan [3327]. misclassification [2990, 1394, 1166, 758, 663]. Mises [1106, 2004, 831, 1928, 2509, 237, 409]. mismeasured [2661]. missed [1098]. Missing [2195, 2556, 1668, 1104, 2402, 3230, 2934, 3328, 2936, 3262, 3003, 2815, 660, 3256, 3350, 2364, 2663, 3189, 1285, 3148, 3219, 2930, 1470, 2437, 2662, 2855, 3208, 3409, 3303, 2921, 3047, 3067]. misspecification [3315]. misspecified [3155, 1573, 3396]. Misspecifying [2542]. Mittag [3296, 2943, 1591]. Mittag-Leffler [3296, 2943, 1591]. Mixed [2579, 2111, 2302, 2247, 861, 3373, 3256, 3029, 2661, 1885, 2843, 2638, 3105, 3137]. mixed-effects [3373, 3029, 2661, 3105]. mixed-models [861]. Mixing [1140, 2189, 2280, 500, 3326, 3196, 1227, 2681, 2445, 3367]. Mixture [2558, 1645, 2257, 2171, 3406, 2191, 2109, 2093, 2557, 1564, 3109, 522, 606, 2568, 1926, 1316, 2787, 3045, 3044, 3023, 2572, 811, 3056, 2449, 3151, 2503]. Mixtures [2319, 2150, 2013, 2170, 2022, 2786, 1200, 1477, 1240, 2573, 1855, 1833, 2501, 1843, 1288, 468, 3093, 1922, 1768, 1449, 2463, 3378, 3134]. ML [2405, 1447]. MLE [3052, 2695, 1272, 1412, 1584, 2475, 820, 2097, 2463, 2852, 1539, 1687, 1828]. MLEs [1021]. mobilité [304]. mobility [143, 304, 193]. Möbius [2455, 2789]. Modal [168, 3266, 2954]. Mode [2201, 332, 3266, 1726, 3046, 1501]. Model [2012, 2579, 3271, 2175, 2053, 2101, 2043, 2165, 2542, 2591, 2148, 2239, 2209, 3003, 2289, 2111, 2229, 2312, 3300, 2207, 3142, 2717, 2202, 2584, 2161, 2008, 2972, 2291, 2208, 2913, 2522, 2017, 2118, 2300, 2030, 2056, 1761, 245, 2025, 2159, 2297, 3303, 3400, 2557, 3420, 1668, 2718, 115, 504, 556, 643, 674, 1090, 1564, 3095, 996, 2746, 1953, 3048, 2671, 1619, 3201, 3058, 1987, 1802, 1367, 489, 3319, 3287, 905, 3315, 1915, 1774, 3188, 1405, 1813, 3212, 2668, 2992, 3238, 1547, 3279, 2829, 2761, 1726, 3076, 2870, 3410, 1133, 1412, 1616, 2791, 3349, 2835, 3186, 1688, 2452, 3292]. model [689, 2380, 3084, 1926, 2986, 1779, 821, 3228, 2787, 1493, 2367, 2652, 2485, 2815, 957, 956, 934, 2466, 1026, 1281, 1492, 2826, 3038, 1951, 855, 1626, 3397, 1732, 1354, 279, 1175, 401, 108, 166, 2766, 3358, 1941, 1447, 3061, 1372, 2425, 1685, 3203, 3086, 3333, 1702, 3350, 2640, 3078, 3005, 1795, 3009, 2776, 2988, 3062, 3277, 3108, 3068, 1693, 744, 843, 1821, 1157, 2880, 3195, 94, 3032, 1905, 2754, 3171, 3170, 449, 748, 2908, 2639, 2559, 1073, 1766, 1898, 1371, 2393, 3072, 509, 2820, 1172, 506, 1169, 872, 191, 264, 2737, 1413, 3222, 2678, 1957, 2650, 3127, 3148, 273, 1770, 1303, 1230, 1284]. model [1168, 1446, 1162, 2554, 2491, 2876, 3006, 197, 285, 1664, 741, 1099, 2759, 1698, 1424, 1850, 1470, 1718, 521, 788, 1132, 1263, 1716, 794, 3217, 1893, 2674, 1952, 3215, 3320, 3369, 2958, 1796, 1282, 1687, 1913, 3435, 2838, 3372, 3343, 2954, 3073, 3111, 3153, 2929, 3067, 1695, 3241, 1897, 3436]. Model-free [3142, 3343]. model-robustness [2746]. Modeling [2274, 2314, 2219, 3434, 2865, 2651, 2699, 3131, 997, 2642, 2970, 2499, 2805, 2675, 2469, 3103]. Modelling [2922, 2848, 2401, 1474, 3378]. Models [2136, 2221, 2558, 2203, 2223, 2144, 2110, 2070, 2349, 2537, 2316, 2217, 2035, 2541, 2205, 2215, 2128, 2147, 2536, 2064, 2288, 2302, 2077, 2085, 2296, 2149, 2191, 2313, 2063, 2520, 2093, 2177, 2071, 2282, 2028, 2194, 2180, 2320, 2181, 2859, 3434, 3258, 1021, 538, 924, 2994, 3330, 2800, 3193, 2871, 3109, 3160, 1370, 3366, 1538, 1878, 2352, 3180, 1814, 861, 1992, 309, 2354, 2665, 3174, 1713, 3074, 1829, 3088, 1630, 2934, 2964, 2714, 3373, 3379, 1143, 2936, 3229, 2996, 2500, 2808, 2944, 3216, 1526, 2351, 2840, 2479, 2603, 2883, 2480, 3017, 2379, 2750, 2496, 1759, 2775, 2529]. models [3338, 2628, 2952, 3257, 3341, 2666, 2389, 2568, 2369, 3186, 3293, 3155, 3395, 3107, 2429, 580, 859, 2460, 3040, 3228, 2734, 2832, 2488, 3283, 3391, 3045, 3044, 2698, 2631, 3415, 3418, 2398, 1398, 913, 2689, 311, 2405, 3256, 3023, 3176, 3276, 3313, 117, 3231, 2574, 3300, 2384, 2764, 3057, 3213, 3430, 3029, 2882, 3261, 2661, 3024, 2492, 1310, 1859, 1906, 2846, 1617, 3125, 3116, 3218, 2478, 3041, 3404, 2572, 2918, 3259, 1558, 3417, 3311, 1925, 3416, 704, 2969, 3033, 1377, 3344, 3274, 3170, 1185, 3214, 2639, 3010, 3246, 3334, 2438, 3318, 2971, 3189, 1712, 434, 2606, 2920, 1573, 2449, 3386, 1745, 1936, 1758]. models [2637, 461, 760, 1466, 2747, 2638, 2876, 3219, 3151, 2735, 2715, 2997, 1694, 1476, 3385, 3384, 3396, 265, 1415, 2562, 3128, 2362, 2392, 2353, 2692, 2662, 3026, 3144, 3200, 3306, 3390, 3411, 1846, 3303, 1812, 3105, 3161, 2921, 3025, 1286, 1103, 1091, 2961, 3077, 2788, 3091, 2595, 3400, 1765, 2911, 3083, 3173, 3253, 3030, 3234, 3137, 3426, 2833, 2703, 3171]. modern [1491]. Modes [1434, 2228]. modification [2879]. modifications [577, 1255, 1911]. Modified [1365, 3214, 1445, 2324, 3028, 684, 2568, 1661, 2888, 2841, 806, 3218, 1139, 727, 1057, 1074, 314, 327]. modifying [2902]. modulated [2839, 2861]. modulation [2448]. Moment [2558, 2252, 2342, 2170, 1957, 2118, 3127, 1600, 2955, 868, 2451, 3155, 3267, 301, 1611, 1869, 933]. Moment-Based [2170]. Moments [1292, 2125, 1351, 910, 2058, 2051, 2322, 2113, 3240, 1998, 1543, 1778, 1982, 101, 1754, 423, 35, 3338, 3045, 3044, 147, 235, 1043, 850, 438, 478, 2744, 1853, 542, 31, 32, 48, 933, 545, 1434, 403, 2786, 325, 1235, 1210]. Momentum [2951]. Momentum-space [2951]. Monitor [2228]. Monotone [2081, 857, 2664, 1976, 2433, 1702, 542, 1422, 1364]. monotonic [1925]. Monotonicity [2287, 1507, 2322, 599, 541, 3075, 709, 2695, 645, 560]. Monte [2079, 107, 115, 3074, 2301, 326, 2220, 2380, 2949, 2968, 199, 2972, 1609, 2218]. Moore [693, 729]. Morgenstern [2627, 1410]. Morisita [1774, 1492]. morphisms [1696]. Mortality [2341]. most [658, 1504, 1741, 1246, 608, 583, 1703, 700, 1058]. Motion [2514, 2214, 2385, 2563, 433, 2995]. motions [2903]. moving [817, 1776, 1630, 3295, 3150, 3432, 1157, 1263, 1067, 2617, 3426]. MRE [610]. MSE [1732]. Multi [2578, 1267, 1268, 3364, 1177, 2530, 2131, 1187, 1045, 1023, 476, 617, 2063, 2450, 1566, 2777, 2477, 244, 633, 1667, 2419, 2684, 1801, 2979, 287, 1342, 3018, 1223, 977, 2935, 1338]. Multi-auxiliary [1177]. multi-Bayes [1338]. multi-components [1667]. multi-decision [244]. multi-dimensional [2979, 977]. Multi-folding [1045, 1023].

Multi-parameter [1187]. multi-path [1801]. Multi-round [3364]. Multi-sample [1267, 1268, 476, 617, 2777, 287, 1342]. multi-stage [1223]. Multi-State [2131, 2530, 2419, 2684]. multi-terminal [1566]. Multi-Unit [2063]. multi-variate [633]. Multi-way [2578, 2450, 2477, 3018, 2935]. Multicanonical [2974]. Multicategorical [2216]. multiclass [2990]. multicollinearity [271, 1573]. Multidimensional [72, 345, 481, 462, 104, 2825, 453, 986, 105, 357, 710]. Multinomial [1709, 1143, 658, 359, 1321, 1504, 639, 1589, 833, 1527, 950, 1073, 1366, 355, 1221]. multinomials [2406]. multinormal [3037, 753, 1338]. Multiparameter [2338, 2186, 1353, 1380, 1433, 1808, 1782, 1527, 1822, 1494, 1050]. Multiparametric [1189]. Multiperiod [1859, 1906]. Multiple [2578, 2307, 192, 2120, 2660, 2140, 2100, 1898, 2056, 2693, 377, 2655, 3315, 3294, 972, 3316, 3403, 3262, 952, 3293, 1355, 1624, 3256, 428, 3167, 1469, 920, 890, 968, 1222, 1873, 3190, 547, 1936, 1303, 466, 582, 995, 804, 1893, 1643, 1301]. multiple-design [1936]. multiples [1256]. Multiplications [2287]. Multiplicative [2531, 2754, 3253]. Multiplicity [1475]. multiplier [2905]. multiply [1997]. Multiresolution [3297]. multiresponse [561]. Multiscale [3157, 3357, 2721, 3380]. multisectoral [554]. Multistate [2598, 3363, 2837]. multistep [1987]. Multivariate [3296, 1554, 2842, 3427, 2240, 1414, 2150, 2257, 2086, 2147, 2254, 2348, 2594, 2118, 1062, 2742, 2580, 505, 2609, 2715, 2508, 2282, 2267, 897, 1192, 421, 302, 3382, 3180, 363, 1015, 857, 835, 1134, 491, 1803, 2890, 411, 3229, 1078, 1340, 1055, 1411, 3157, 2621, 2622, 667, 763, 834, 1133, 2389, 1976, 1929, 696, 3226, 926, 2429, 821, 1488, 2888, 2841, 2387, 365, 412, 694, 721, 761, 1425, 2501, 3254, 1014, 1393, 1404, 2466, 1199, 1320, 1532, 806, 913, 3002, 1571, 311, 2660, 2405, 2687, 1794, 3150, 438, 516, 1339, 3118, 2789, 3070, 3100, 1618, 1644, 375, 1689]. multivariate [3203, 3181, 1191, 2933, 775, 1699, 1789, 1548, 1541, 3225, 1576, 2432, 2867, 1691, 3376, 1374, 2771, 820, 546, 518, 599, 1020, 3042, 1371, 3433, 1995, 1854, 1315, 1413, 600, 2805, 1768, 1544, 1936, 841, 187, 214, 340, 364, 420, 464, 461, 715, 760, 809, 927, 722, 742, 791, 682, 2491, 975, 1176, 1452, 380, 539, 1664, 610, 741, 1470, 1353, 2383, 1415, 1089, 3224, 3227, 1286, 3323, 1426]. multivariete [1197]. multiway [1472]. Müntz [982]. mutual [1045, 1023]. mutually [1323]. MV [1489]. MV-optimality [1489]. myopic [331].

national [221, 143]. natural

[3052, 343, 1272]. Near [2334, 3216, 2654]. Nearly [2160, 2706, 574, 1210]. nearness [3028]. necessarily [1336]. Necessary [286, 374, 2473, 1885, 435, 2646, 1633]. NEF [1894]. Negative [2008, 2050, 355, 805, 847, 2656, 2485, 2361, 205, 833, 2982, 1183, 3172, 415, 320, 1994]. Negative-Binomial [2008]. neighborhood [3198]. Neighbourhoods [2038]. Nested [2579, 2140, 2118, 1940, 1989]. Network [2521, 3021, 3431, 3107, 3390, 3369, 3323]. networks [3292, 3371, 2650, 3367]. Neumann [115, 2796, 94]. Neural [2521, 3431, 3292, 3371, 2713]. neuroblastoma [3318]. neurodynamics [1460]. never [324]. news [3254]. Newton [1746, 1313]. Neyman [2817, 449, 748, 2697, 92, 2977]. No [249, 1845, 3285, 2424, 1768]. no-signal [2424]. node [3390]. Noise [2336, 2591, 3095, 3271, 2890, 3307, 2766, 3308, 3290, 3101]. noisy [2694]. Non [854, 2339, 2504, 2221, 487, 2107, 2341, 2328, 2596, 516, 2215, 424, 2051, 680, 842, 2743, 2291, 2050, 2262, 546, 2527, 2219, 2177, 3429, 1200, 2948, 1156, 1646, 1970, 2891, 417, 2800, 2746, 519, 1543, 1707, 101, 2615, 489, 2700, 807, 1274, 3247, 3403, 1645, 2401, 667, 808, 834, 2454, 676, 2738, 2806, 1837, 208, 205, 645, 2982, 2570, 756, 997, 298, 1805, 1714, 3104, 3035, 679, 1925, 819, 671, 909, 748, 3334, 1110, 2605, 622, 548, 584, 1479, 287, 379, 2843, 742, 954, 991, 3004, 3185, 932, 182, 2831, 1830, 1860]. non-additive [932]. non-adjacent [2454]. non-Bayesian [1805]. Non-central [516, 546, 101, 676, 679]. non-centrality [298]. non-compactly [3403]. non-concave [3185]. Non-Equally [2262]. non-ergodic [3035]. Non-existence [424, 622]. Non-Gaussian [2215, 2700, 2570, 3334]. non-homogeneous [2831]. non-i.i.d [1274]. non-I.I.D. [1645]. non-identical [1707, 909]. Non-Identically [2328, 1543, 1479]. non-identifiability [1200]. non-independent [1707]. non-invertible [2605]. Non-Iterative [2341]. non-linear [489, 932, 1860]. non-minimum [1714]. non-monotonic [1925]. Non-Negative [2050, 205, 2982]. non-normal [2615, 671]. Non-normality [487, 2596, 2948]. Non-null [680, 842, 519, 807, 667, 808, 834, 756, 819, 748, 1110, 742]. non-oblique [1837]. non-orthogonal [548, 584]. Non-parametric [854, 2107, 3429, 208, 3104, 287, 379, 991]. non-quadratic [2800]. non-recurrent [2843]. Non-Regular [2339, 1156, 1646, 1970, 2891, 2746, 2738, 2806, 645, 954]. non-restricted [3004]. Non-simple [2527]. non-smooth [3247]. Non-Spherical [2177]. Non-standard [2504]. Non-Stationary [2291, 2401, 997, 182]. non-steep [1830]. Non-very [2743]. Nonasymptotic [2521]. nonbinomial [1164]. noncentral [1107, 366, 2744, 999, 1972, 2417, 666]. noncentrality [2417]. Nonconservative [2142]. nondifferentiable [1607, 2874]. Nonexistence [1042, 2074, 1068]. noniformative [1221]. nonignorable [3302, 3376, 3306]. Noninformative [2300].

Nonlinear

[2676, 2552, 2223, 2197, 2895, 2035, 2215, 2583, 2147, 2499, 2133, 2296, 2219, 2109, 2297, 2071, 2176, 1668, 2355, 1607, 2807, 2479, 3415, 3418, 2872, 3213, 3029, 1435, 3417, 3416, 3246, 886, 1313, 2930, 2426, 2657, 3435, 1711]. nonminimaxity [1404]. Nonnegative [2155, 1278]. nonnormal [1203]. nonnormality [2708]. Nonparametric [3255, 1081, 1082, 1154, 3363, 2579, 2643, 1015, 2341, 3087, 3379, 1428, 637, 2451, 2283, 3263, 3403, 3257, 2274, 3326, 689, 2380, 1779, 1968, 2582, 2290, 2199, 2189, 2414, 2680, 2763, 2900, 1627, 1587, 3295, 3149, 2006, 2905, 2264, 2536, 2754, 2534, 2204, 2415, 1244, 1866, 1179, 520, 597, 2755, 2830, 1847, 2034, 2265, 142, 204, 215, 737, 1513, 3025, 2200, 2728, 2693, 1130, 1438, 1824, 1755, 1577, 2777, 3177, 3280, 2869, 2756, 3328, 2679, 3188, 1845, 2700, 1319, 2441, 2664, 2701, 2886, 3011, 3374, 2602, 3036, 2856, 2884, 1290, 738, 739, 467, 871, 2757, 2734, 2398, 2384, 1537]. nonparametric [3168, 1993, 1439, 2970, 2670, 2947, 3042, 2987, 1395, 1641, 1502, 422, 463, 547, 596, 1445, 2804, 2803, 2940, 612, 165, 189, 3217, 3134, 2961, 3426]. Nonparametrics [2105]. nonrandomly [3394]. Nonregular [2014, 1352, 1370, 969, 1027, 1112]. nonresponse [3302, 3376]. nonstandard [2605, 1895]. Nonstationarity [2589]. Nonstationary [2160, 3362, 3235, 2384, 1418, 1816]. nonsynchronously [2630]. norm [1575, 3337, 3181, 798, 2685]. Normal [330, 2043, 2020, 2150, 2139, 2245, 2237, 2087, 2051, 2254, 2121, 2047, 1761, 2040, 2747, 2580, 2109, 2113, 2344, 2267, 200, 1140, 1556, 421, 581, 735, 459, 973, 106, 1394, 857, 1450, 2354, 404, 2615, 3028, 1515, 1340, 1344, 423, 1833, 35, 495, 3037, 667, 1976, 1767, 2429, 2431, 633, 755, 1545, 412, 694, 721, 761, 769, 1673, 3347, 209, 1404, 1271, 3002, 198, 311, 2687, 376, 516, 616, 1339, 1202, 1480, 1689, 1820, 1700, 1191, 2472, 1435, 1753, 844, 1548, 3033, 31, 48, 1079, 1095, 671, 1241, 1333, 448, 744, 843, 942, 2432, 1045, 1023, 1441, 3, 1211, 2771]. normal [3056, 283, 563, 1625, 820, 599, 2828, 452, 1972, 3209, 1717, 1138, 990, 1213, 2363, 600, 663, 1768, 403, 256, 293, 1449, 1544, 1911, 141, 291, 841, 83, 187, 340, 582, 881, 464, 715, 341, 1168, 899, 2491, 3006, 975, 1176, 356, 1583, 1716, 2767, 2474, 1830]. normalised [3132]. Normality [2116, 2165, 2246, 2045, 2586, 2948, 672, 487, 918, 1416, 1523, 1969, 1929, 1092, 1977, 2596, 670, 6, 12, 2630, 3267, 678, 1236, 1199, 1532, 1861, 2847, 866, 1389, 749, 1905, 1551, 828, 1302, 2843, 1777, 132, 159]. Normalized [2286]. Normalizing [1425, 1304, 1533, 3182, 1771, 2406, 1930]. normally [68]. Norms [2192]. Note [406, 272, 586, 28, 101, 717, 2018, 721, 153, 453, 136, 128, 743, 113, 150, 661, 164, 796, 840, 876, 1250, 2536, 2, 47, 91, 32, 2009, 874, 391, 2191, 862, 163, 92, 129, 179, 983, 380, 65, 181, 124, 710, 1130, 709, 21, 1953, 887, 519, 1384, 1440, 404, 1894, 691, 692, 549, 1072, 984, 1523, 1193, 856, 316, 901, 494, 673, 670, 1227, 982, 209, 274, 248, 1787, 1043, 816, 263, 376, 428, 401, 478, 1237, 1644, 534, 567, 1289, 469, 1691, 1441, 1703, 283, 509, 1024, 863, 2363, 320, 282, 841, 238, 356, 613, 1819]. note [1119, 815, 1431, 2576]. Notes [502, 2667, 214]. Notice [201]. Notion [2287, 961, 441, 1421]. Nuisance [2178, 3258, 1870, 1651, 1762, 1703, 2443]. Null [2589, 1590, 2596, 2140, 519, 807, 667, 808, 834, 856, 756, 680, 819, 842, 449, 748, 1110, 742]. null-distribution [449]. null-hypothesis [449, 748]. Number [2053, 2332, 2111, 3099, 2090, 2408, 1650, 414, 2837, 3373, 3335, 1722, 1921, 670, 1316, 2501, 1236, 137, 979, 255, 889, 919, 2972, 3194, 930, 624, 711, 379, 1387, 2984, 2752, 56, 325, 3390, 3412, 324, 540, 2935, 1235, 3310]. Numbers [2210, 2256, 2036, 2255, 2598, 2067, 2286, 1841, 1907, 382, 639, 2484, 1492, 2653, 990, 1233, 1511]. Numerical [693, 729, 512, 1991, 108, 393, 134, 181]. numerische [393].

Objective [2920, 2958, 924, 1609, 456, 436]. objects [1423]. oblique [1837]. Observation [2135, 299, 21, 3319, 3287, 1738, 112, 2754, 1201, 702, 754]. Observations [2195, 2270, 2173, 1668, 1556, 1104, 823, 3280, 144, 3230, 1713, 1917, 1274, 396, 1506, 1935, 660, 278, 525, 2785, 942, 2782, 936, 3329, 1285, 1098, 1168, 416, 455, 2720, 189, 543, 324, 906]. observe [324]. observed [3271, 3022, 2829, 2630, 311, 2918, 1876, 3386, 2862, 2843, 2381, 2903]. Observing [2167]. obstructive [1421]. obtained [996, 2656, 345, 453, 481, 957, 956, 255]. OC [577, 1256]. occasions [1279]. Occupancy [1462, 639, 670]. occurrence [2718, 1496, 2668, 1231, 1730, 2620]. occurrence/exposure [1496]. Occurrences [2077, 2276, 2090]. occurring [1795]. Odds [2289, 2229, 2986, 2815, 2636, 2457]. Oil [2130]. oki [1676]. old [3138]. Olkin [1436]. omnibus [3037]. One [2135, 2019, 2596, 2526, 2268, 2945, 2409, 2023, 2157, 1746, 1770, 299, 1448, 133, 3365, 3330, 857, 1855, 1632, 2673, 691, 1939, 135, 198, 1604, 1454, 530, 1531, 1685, 3172, 1224, 969, 1112, 2368, 626, 640, 2993, 447, 548, 585, 422, 1708, 566, 2637, 1168, 1065, 975, 2814, 794, 1212, 1812, 2935, 713]. One- [1770]. One-armed [2945]. one-chance [1065]. one-dimensional [3330, 713]. one-directional [1448]. one-parameter [1855, 530, 1224]. one-restrictional [548]. One-Sample [2268, 857, 3172, 447]. One-Sided [2019, 2157, 2409, 691, 1939, 198, 1531, 2368]. one-stage [969, 1112]. One-step [2526, 1746, 3365, 1604]. One-way [2596, 1685, 2637, 1168, 794, 1812]. ones [2656]. online [3246]. only [1014, 1256]. onset [2452]. onto [2369]. opacity [2636]. open [108, 117, 166]. operating [899]. operation [262]. Operational [2177]. operations [19]. Operator [2051, 1667, 3399]. Operator-Stable [2051]. operators [3166, 3337, 3327, 1160, 1161]. opinion [338, 352]. opinions [2707]. Optimal [2693, 331, 2221, 2203, 2512, 777, 1601, 396, 1856, 2196, 2349, 2603, 2819, 2602, 2007, 2870, 2115, 2396, 1656, 2172, 1376, 2882, 2182, 3060, 1332, 1712, 1768, 2228, 1168, 1252, 1022, 2733, 1890, 2746, 2137, 2665, 3388, 1713, 1086, 2537, 3285, 1973, 738, 739, 1977, 136, 252, 1159, 1347, 1398, 1593, 1660, 3348, 949, 2879, 1322, 3277, 1458, 373, 623, 809, 129, 285, 371, 995, 2997, 2827]. Optimality [2166, 2514, 2563, 2197, 559, 3059, 2590, 1156, 1482, 489, 1362, 1489, 1620, 1432, 3033, 2697, 561, 1050]. optimally [1503]. Optimisation [2468, 2482]. optimization [3110]. Optimizing [1892, 3281]. Optimum [1736, 336, 349, 1518, 562, 719, 780, 851, 115, 195, 302, 1720, 867, 1806, 1250, 1447, 630, 236, 626, 640, 391, 967, 56, 456, 436, 657]. Oracle [2725, 3239, 3215]. Order [2326, 2210, 2061, 2125, 2248, 2235, 2197, 2530, 2274, 2041, 2058, 2143, 2273, 2322, 2453, 2204, 2050, 2009, 2275, 134, 2315, 2108, 2059, 1816, 2067, 2090, 2113, 2705, 1998, 1402, 1482, 1574, 1646, 1971, 3052, 406, 1287, 1314, 1525, 1999, 1203, 709, 2739, 1543, 1707, 1778, 1982, 2849, 369, 1261, 1262, 1327, 2645, 2510, 2683, 3388, 3265, 1086, 1397, 423, 1983, 2421, 1973, 1270, 1503, 1368, 1523, 1938, 1852, 1264, 3260, 246, 839, 2454, 1582, 832, 1184, 982, 1839, 615, 2361, 2780, 1043, 568, 3015, 1229, 292, 2470, 2872, 949, 1237, 2546, 1714, 1447, 1560, 2359, 1435, 1298, 1980]. order [1277, 1882, 3013, 1378, 2587, 2559, 1903, 1815, 626, 640, 1494, 527, 3154, 2691, 1479, 2430, 541, 793, 989, 1230, 1777, 325, 611, 610, 1099, 1930, 1947, 1476, 2657, 1890, 1148]. Ordered [2560, 2343, 2169, 272, 595, 3028, 1396, 779, 3347, 2497, 427, 619, 1657, 269, 1927, 874, 1179, 1343, 1273, 539, 1022, 1443, 1912]. Ordering [2124, 2254, 3071, 3242, 2386, 227, 1463, 1781, 444, 482, 471, 613, 1116]. orderings [1851, 1731, 1409, 1963, 560]. Orders [2273]. ordinal [2636, 1415]. Ornstein [3163, 2843, 2903]. orthogeodesic [1829]. Orthogonal [878, 188, 1036, 1181, 2761, 1092, 1226, 1093, 3281, 1901, 548, 1323, 584, 2685]. orthogonality [1375]. Orthogonally [1820, 1488]. oscillation [262]. Other [2095, 68, 1571, 368, 270, 1474]. Outcome [3368]. outcomes [1623, 1807, 3376]. outlier [2448, 1186, 1898, 3375]. Outliers [2053, 1488, 1453, 2438, 737]. outline [9, 15]. outlying [189]. output [108, 117, 166]. outputs [117]. over-parametrized [3431]. Overdispersion [2089]. overlapping [2361]. overtime [151].

p [3340, 1335]. p-values [3340]. p.b.i.b. [669]. P.P.S. [688]. Packing [2261, 247, 1026, 1399, 1046]. pair [1701, 1775, 3298]. paired [54, 2433, 345, 453, 481, 527]. pairs [3310]. Pairwise [1592, 812, 1861, 3289, 3288, 3059, 3402]. Pairwise-balanced [1592]. Palm [2916]. panel [2994, 2777, 3404]. paper [1326, 608, 51, 3140]. papers [841]. paradox [2717]. Parallel [2972, 3204, 1913]. Parameter [2179, 2379, 2517, 2016, 1586, 2526, 3057, 2268, 3062, 1157, 2284, 3258, 1094, 1299, 1402, 3052, 537, 3054, 1272, 1370, 1855, 1804, 2627, 1246, 3130, 1725, 1530, 1651, 2973, 575, 517, 1516, 2400, 1254, 2631, 1832, 524, 2982, 850, 45, 530, 1187, 1880, 2495, 2933, 1762, 2471, 2669, 790, 1053, 3117, 701, 1224, 2880, 236, 3246, 2697, 2993, 895, 1442, 2729, 849, 2417, 320, 896, 760, 883, 789, 1734, 267, 513, 3004, 2733, 1155, 1212, 1188]. Parameterizing [3045, 3044]. Parameters [2317, 2343, 2075, 2162, 2342, 2340, 2207, 2586, 2178, 1154, 1356, 1525, 1174, 459, 234, 1870, 2851, 489, 868, 1405, 1428, 522, 606, 772, 2798, 638, 122, 1243, 902, 2398, 301, 645, 278, 525, 3358, 1834, 1880, 1628, 1585, 2475, 1056, 574, 704, 432, 2444, 1703, 3281, 2559, 2443, 361, 933, 476, 752, 403, 2862, 312, 1221, 700, 1058, 1466, 2463, 1850, 1127, 788, 1912, 3390, 1695, 1711]. Parametric [2533, 3408, 2092, 1615, 1924, 2074, 2556, 854, 2107, 3307, 2701, 1503, 2870, 3003, 2832, 2367, 208, 3321, 2764, 3261, 2492, 3104, 3117, 3274, 3171, 3170, 2606, 287, 379, 2860, 2759, 991, 1952, 3137, 3429]. Parametrization [2108, 3064]. parametrized [3431]. Pareto [3198, 917, 3324, 707, 728, 747, 1902]. parsimony [1669]. part [1028, 826, 254, 258]. Partial [2221, 892, 2349, 560, 2557, 2911, 595, 2962, 2996, 1856, 3377, 3267, 2506, 3203, 3009, 449, 748, 174, 593, 2392, 2353, 2662, 2985, 2921, 3025, 1282, 182, 3253]. Partially [2934, 2035, 3026, 2600, 2964, 1534, 2829, 723, 3175, 2952, 3228, 840, 1120, 279, 1376, 2661, 2988, 2918, 449, 1102, 3072, 1018, 400, 3200, 3411, 3083, 2954, 2703]. Particle [2220, 3246, 2379, 2973, 1053, 2971]. Particle-based [3246]. particles [1666]. particular [974, 271]. Partition [2301, 2838]. Partition-Weighted [2301]. partitioning [2858]. Partitions [1165, 779, 3080, 2984]. partly [3264, 3112]. Parzen [894]. Pascal [426]. pass [490]. passage [2612]. patchily [1423]. path [1801]. paths [1622, 1759, 2811, 1248]. Pattern [2327, 2580, 2420, 3187, 523, 607, 921, 153, 1214, 3089, 3259, 162, 197]. patterned [975, 1176]. Patterns [2331, 2530, 2094, 2528, 2865, 2837, 1840, 2419, 2684, 1136, 1459, 1562, 2678, 3245, 1792, 2617]. Paulson [633]. PBIB [279, 810, 1309, 1455, 1121, 796, 1308, 374, 435, 748, 622, 795, 308, 451, 778, 623]. pdf [1175]. peakedness [686]. Pearson [2932, 2545, 128, 301, 2123, 2333, 2697, 1901, 92, 383]. pediatrics [85]. Penalized [2800, 2346, 3086, 3191, 2874, 3116, 2544, 2733, 3185, 2703]. penalties [2800, 3316]. Penalty [2346, 2578, 1577, 2874, 1319, 3136, 2736]. Penrose [693, 729]. Penultimate [1260]. percent [2902]. percentage [1031, 142]. Percentile [2348, 1383, 1311]. percentiles [875]. Performance [2074, 1743, 2449]. period [246, 1493]. periodic [2766, 1614, 880]. Periodograms [2065]. Permissible [2981]. permit [1708]. permitting [3357]. permutation [1921, 1529, 1551, 3320]. permutations [560]. persistent [2879]. perspective [2775]. Perturbations [1087]. perturbed [886]. phase [437, 813, 867, 1714]. phenomena [39]. phenomenon [265]. Phylogenetic [2158, 3133, 2719]. piece [3382]. Piecewise [2225, 2973, 3375]. piecewise-linear [3375]. piggyback [2480]. Pillai [1788]. Pitman [3028, 1981, 1555, 1269, 736, 2984]. Pitman-type [736]. places [745]. plan [1361]. planar [1775, 1954]. plane [1346, 3268, 3252]. Plans [2285, 838, 604, 548, 624, 711, 584, 899, 203]. plant [1491]. plants [1458]. plot [861]. Plug [2626]. Plug-in [2626]. plus [3290]. Plutonium [2161]. Point [2705, 2145, 2203, 2134, 2253, 2139, 2140, 2261, 2077, 2314, 2527, 2113, 2528, 1932, 2644, 3423, 3284, 3419, 785, 2385, 2825, 1405, 1515, 3163, 2874, 3092, 3069, 3017, 3252, 1359, 3377, 1516, 23, 39, 1586, 1801, 1794, 3362, 3408, 3261, 1795, 3035, 2918, 1596, 1789, 1925, 2468, 3172, 1889, 1079, 3032, 1905, 988, 1136, 1185, 1459, 1562, 1025, 933, 1785, 2916, 3433, 3222, 3386, 2678, 1775, 983, 3297, 2418, 1431, 3085, 2831]. Point-Process [2077]. Points [2053, 2585, 2117, 2645, 3357, 2807, 1403, 1014, 1701, 2979, 1031, 281, 187, 945, 142]. Pointwise [2590, 3159, 2263]. Poisson [662, 1776, 2145, 2356, 2456, 1622, 50, 3268, 2307, 216, 3279, 1509, 2089, 638, 3000, 3252, 2895, 1645, 1840, 3182, 152, 42, 206, 2119, 2690, 2897, 1787, 2111, 1486, 1461, 1853, 2839, 3078, 853, 469, 2302, 20, 1889, 1366, 339, 2309, 2650, 239, 354, 1074, 267, 1127, 1614]. polaron [71]. Poles [3298]. policy [331, 2422]. pólya [1796, 2327, 806, 2523, 2564, 3143, 2466]. Polychotomous [2144]. polygon [3427]. polygonal [1961, 2592]. Polynomial [2298, 2196, 2007, 2582, 2345, 2147, 2098, 2794, 3342, 2512, 1547, 1856, 3285, 252, 2619, 618, 2663, 2961, 2595]. polynomials [1358, 1055, 1135, 1030, 2780, 652, 2881, 1901, 1160, 1161, 188]. polyspectral [2683]. polytopal [1251]. polytope [2745]. pooled [1299]. pooling [359, 965, 646]. Population [2325, 2057, 2120, 2258, 2024, 2059, 2344, 2600, 581, 1258, 302, 1164, 857, 2778, 1515, 972, 952, 3211, 526, 1425, 1044, 1665, 3008, 269, 3108, 942, 3, 283, 1032, 820, 1933, 753, 391, 1213, 621, 83, 340, 464, 3392, 245, 471, 539, 613, 1209, 609, 572, 614, 648, 655, 204, 687, 1210]. Populations [2347, 2050, 2009, 1658, 421, 581, 1790, 1370, 3166, 1610, 419, 667, 297, 253, 365, 902, 568, 2660, 1657, 875, 516, 525, 2359, 1548, 542, 448, 2670, 464, 715, 356, 477, 505, 610, 1912, 93, 132, 159, 3346]. portfolio [2716]. Positive [1851, 1409, 2302, 3054, 2394, 1030, 1093, 431, 515, 535, 1524, 2472, 1141, 114, 1717, 1019, 1028, 713, 3039]. positively [725]. Positron [2163]. Possible [1539, 3315, 1064]. possibly [2483, 2737, 1353, 93, 737]. Post [2049, 1035, 2809, 2925, 1266, 3122]. Post-Data [2049]. post-stratification [1035, 2809, 2925]. post-war [1266]. Posterior [2135, 2112, 1726, 2014, 2134, 2013, 2483, 1433, 2700, 1481, 1510, 1808, 1782, 2933, 889, 919, 744, 843, 1772].

Posteriori [2220, 1416]. potentially [2656]. potentials [1359, 1136, 1562]. Powell [2847]. Power [556, 2539, 1029, 2316, 524, 2048, 518, 213, 473, 591, 1742, 2914, 3294, 594, 917, 1530, 1133, 3050, 645, 3286, 1317, 1569, 2910, 2853, 3029, 1139, 3404, 1458, 294, 599, 1054, 1110, 1259, 1922, 1392, 2786, 1391, 409]. power-series [1922]. Powerful [2156, 3158, 1246, 608, 583, 1703, 700, 1058]. powers [1651, 1342]. PPS [437]. Practical [613, 37, 424, 132]. Pragmatic [1437]. Precision [2047, 2267, 532, 1483, 2491, 3399, 827]. predictands [585]. Predicting [2897, 2351]. Prediction [2791, 2162, 2230, 2300, 2984, 2076, 2259, 2362, 2206, 1094, 538, 1987, 1610, 785, 1919, 39, 2942, 2652, 472, 1341, 1732, 2687, 2574, 1049, 1277, 1678, 1761, 1114, 1195, 1896, 65, 1067, 1765, 3014, 1695]. predictions [3397]. Predictive [2159, 1090, 1805, 1993, 2816]. predictor [591, 590, 1115, 585, 1638, 3385, 3384]. predictors [559, 3249, 3277]. Preface [2567, 2376, 2648, 2710, 2741, 2967]. prefectural [963]. prefectures [963]. preference [1774, 124]. preferential [2650, 3367]. Prejudicing [2193]. Preliminary [2582, 49, 126, 510, 1382, 1767, 553, 948, 925, 965, 1254, 1404, 1271, 1241, 1333]. preliminary-test [1382]. Presence [2041, 2005, 3242, 1870, 2690, 1202, 1453, 1375, 1576, 1703, 2443, 2827, 2437]. present [208]. presentations [2744]. preserve [1567]. Preserving [2581, 2694, 1772, 2902]. prewhitening [262]. Price [2130, 1919, 166]. Principal [2069, 2761, 1769, 1159, 3249, 389]. principle [800, 1386, 2575, 3051, 1529, 1180, 989, 2427]. Prior [2135, 2182, 2204, 1076, 1219, 1450, 1145, 2772, 479, 2850, 3033, 630, 914, 1221, 456, 436, 2795, 2981]. priority [1010].

Priors [2157, 2300, 2109, 2863, 1433, 1783, 1725, 2351, 2574, 1826, 1310, 3239, 3132, 1772]. Probabilistic [3015, 2132]. Probabilities [237, 1468, 2612, 359, 388, 3325, 639, 196, 833, 418, 1248, 3167, 534, 2979, 1730, 950, 469, 3117, 1980, 174, 663, 621, 719, 1781, 457, 99]. Probability [2037, 344, 2190, 2351, 2293, 2030, 2315, 2169, 1646, 195, 709, 1984, 1164, 3284, 1166, 2892, 1344, 1623, 1807, 768, 1386, 759, 1599, 1033, 150, 210, 280, 317, 385, 498, 564, 865, 1296, 205, 235, 758, 112, 1520, 1152, 1365, 528, 82, 40, 207, 2395, 257, 313, 1955, 732, 938, 506, 1129, 1580, 141, 1772, 1096, 1572, 367, 76, 1312, 3159, 1165, 2707, 321, 904, 442, 592]. probable [658, 1321, 1504, 1741]. probit [273]. Problem [2255, 2134, 2076, 2259, 2206, 897, 2948, 935, 1142, 234, 475, 28, 3163, 2792, 928, 3264, 638, 1969, 42, 526, 1178, 52, 53, 781, 3348, 1801, 3129, 1730, 47, 131, 3172, 1027, 1463, 909, 476, 1717, 391, 752, 422, 1223, 420, 407, 975, 1176, 181, 804, 576, 1633]. Problems [2104, 2122, 2039, 2203, 2184, 2211, 2327, 2548, 2333, 2268, 2182, 2074, 2094, 2187, 172, 1721, 1958, 1999, 2420, 2926, 1600, 3280, 659, 857, 835, 425, 2796, 3110, 3285, 3126, 2884, 2608, 696, 786, 926, 670, 1190, 317, 913, 568, 1624, 3362, 401, 908, 2482, 2904, 703, 103, 98, 1294, 920, 890, 968, 1222, 744, 843, 94, 1020, 1502, 848, 2786, 789, 161, 416, 455, 1920, 109, 3389, 1088]. Procedure [2731, 2143, 966, 581, 649, 1258, 1752, 244, 1003, 1427, 523, 607, 454, 297, 3040, 1254, 483, 536, 288, 949, 997, 1702, 1136, 1609, 283, 734, 371, 3113, 2418, 3372]. Procedures [2032, 526, 2120, 2087, 2074, 2535, 126, 659, 3315, 1515, 1741, 3294, 1257, 1163, 192, 740, 871, 786, 553, 1033, 948, 925, 57, 78, 1389, 574, 958, 1150, 3101, 289, 1780, 239, 998, 716, 715, 285, 699, 324, 1301, 1671, 544]. Process [2126, 2160, 2075, 2199, 2041, 2058, 2048, 2077, 2291, 2224, 2705, 882, 1654, 133, 148, 190, 225, 417, 1012, 1016, 1174, 1577, 1115, 2504, 1742, 1992, 3352, 3419, 1622, 1051, 2998, 3143, 2773, 3092, 3335, 1530, 2401, 2884, 1264, 2608, 3124, 2690, 2897, 3147, 1360, 303, 1586, 1874, 323, 3015, 2506, 45, 2789, 1331, 501, 1560, 1853, 3261, 2945, 2839, 1795, 2933, 2918, 503, 199, 228, 2880, 1185, 492, 339, 2403, 885, 3222, 601, 989, 1797, 1638, 1475, 1418, 1099, 2467, 3113, 932, 2861, 1849, 259, 1330, 702, 182, 1597, 2810]. Processes [2127, 2145, 2321, 2523, 2222, 2552, 2516, 2189, 2084, 2526, 2261, 2023, 2309, 2527, 2176, 2005, 653, 817, 1203, 662, 1776, 2356, 1932, 2644, 3363, 2564, 884, 3360, 782, 1888, 2950, 2976, 1249, 2355, 1373, 1086, 2825, 3069, 1481, 1144, 2886, 3374, 3278, 3146, 2973, 2666, 1929, 3326, 3377, 3425, 1977, 2460, 2630, 2680, 2763, 502, 726, 766, 1787, 1861, 3397, 1794, 465, 550, 319, 3150, 2654, 2470, 1248, 3235, 3408, 1949, 2854, 3308, 2910, 3405, 3035, 1596, 2919, 1789, 1561, 1008, 351, 2468, 1889, 993, 2704, 1905, 988, 1025, 2916, 3101, 492, 705, 994, 1057, 520, 938, 2445]. processes [2691, 1442, 3386, 1615, 2862, 2843, 466, 582, 2505, 1775, 306, 371, 603, 3297, 2977, 1533, 1536, 2795, 1087, 2720, 815, 1067, 754, 1875, 1517, 1614, 3323, 2617, 2903, 2831]. processing [3271]. produces [225]. Product [2149, 2105, 2062, 1998, 127, 1134, 531, 1674, 1024, 1775]. production [41]. productive [2711]. products [973, 2651, 2656]. Professor [3140]. Profile [2295, 2492, 1923, 1015, 2832, 2764, 2816, 1913]. profiled [3153]. Profiles [2063]. prognosis [85]. programming [348, 179]. Progressive [2618, 1998, 2702]. progressively [1318, 1498, 3154]. projected [390, 155]. Projection [2032, 1554, 3403, 2866, 1379, 2369, 3425]. projections [494]. projectors [2685]. promotion [2651]. Proof [2335, 2065, 173, 1363, 1989]. proofs [77]. propagation [1675]. Proper [1191, 3251, 2571, 2981]. Properization [3251]. Properties [2644, 2188, 724, 2552, 2338, 2539, 2257, 2231, 2143, 911, 2534, 2133, 2157, 2554, 2093, 2088, 845, 962, 2247, 2068, 2705, 802, 1181, 1017, 1575, 2877, 1496, 1272, 905, 647, 2778, 1358, 970, 2700, 1568, 846, 1264, 829, 893, 779, 1147, 769, 252, 2386, 1159, 806, 1656, 3015, 45, 642, 3202, 1196, 1269, 3029, 2785, 2783, 641, 3204, 3233, 1222, 1889, 2444, 1827, 1374, 294, 1655, 2663, 3016, 1696, 3329, 1395, 938, 1442, 379, 883, 1452, 2610, 1947, 1698, 1052, 818, 3244, 204, 321, 458, 1232, 2902, 1553]. Property [384, 2023, 2306, 148, 709, 2730, 1118, 571, 1447, 501, 492, 339, 2445, 1040, 1215, 3215]. proportion [690]. Proportional [2702, 2165, 2986, 2289, 2229, 3048, 3287, 2870, 3349, 1688, 792, 2815, 2640, 827, 1282, 3372]. proportionate [40, 99, 2902]. proportions [1004, 1101, 1881, 3104, 3194]. proximal [2874]. Pseudo [2279, 3179, 3212, 3302, 1861, 2905, 3329, 1077, 990]. pseudo-best [1077]. Pseudo-Gibbs [3179]. pseudo-Lagrange [2905]. Pseudo-Likelihood [2279, 3302]. pseudo-observations [3329]. pseudo-random [990]. pseudo-score [3212]. Pseudolikelihood [2136]. psychological [654]. public [338, 352]. publication [3333]. publique [338, 352]. pure [1576, 1963]. Purely [3172, 425]. purposes [1744]. pursuit [1554]. PVF [1894].

Quadratic [1238, 423, 1239, 1348, 2266, 2040, 2800, 2730, 3022, 1507, 2792, 3037, 2616, 2486, 1092, 1193, 412, 694, 721, 761, 1480, 3060, 937, 1141, 3, 1972]. quadratic-approximation [1507]. Qualitative [2962, 23, 39]. quality [1266, 899]. Quantal [2144, 2251, 637, 378]. quantification [3292, 23, 39, 72, 345, 453, 481, 697, 62]. quantifying [85]. Quantile [2189, 3147, 3034, 3009, 3218, 2266, 2467, 2005, 3111, 1258, 1384, 3388, 3236, 1246, 1581, 2784, 3011, 3364, 2965, 1688, 2900, 3391, 1784, 1748, 3149, 3430, 2785, 2873, 3191, 3116, 2801, 3117, 2670, 1580, 3299, 1908, 1895, 3421, 3400, 3085, 3073, 3153, 3234]. Quantiles [2545, 2560, 2240, 1916, 310, 2778, 2481, 3260, 678, 1236, 3053, 1748, 3231, 2783, 2907, 844, 236, 93, 540, 1155, 3123, 2399, 1039]. Quantitative [824, 826, 3389, 2169, 2805, 1443]. Quantities [2153, 749]. Quantlets [2226]. quantum [2575, 385, 323]. quantum-biophysics [385]. Quasi [2295, 1923, 814, 3290, 1891, 2142, 3365, 3114, 3422, 3304, 1186, 3125, 2828, 2724, 1715, 1966, 2794]. quasi-arithmetic [3365]. quasi-information [3422]. Quasi-likelihood [3290, 1891, 3304, 3125, 2724, 1715, 1966, 2794]. Quasi-Likelihoods [2142]. Quasi-Profile [2295]. Quasi-stationary [814]. quasi-Sudoku-based [3114]. quaternion [3063]. Quelques [222]. Quenouille [1990]. Quenouille-type [1990]. Queue [2107, 414, 727, 1205]. Queueing [148, 434, 603, 551, 947, 1328, 704, 959, 552, 314, 327, 521, 845, 962, 1116]. queues [602, 1613]. quota [1984]. Quotas [2037]. quotient [531].

radial [2676, 2570]. radioligand [960]. railroad [183]. railway [343]. rainfall [1960]. Randić [2628]. Random [2033, 2331, 2579, 2039, 2328, 2053, 2110, 2070, 2163, 2239, 689, 2858, 2279, 137, 1281, 2258, 2026, 2051, 2291, 1406, 2251, 2094, 2286, 2113, 2838, 2320, 3434, 1182, 1721, 1984, 75, 272, 302, 2612, 2730, 1292, 1543, 1814, 1278, 1134, 2756, 2934, 3236, 2456, 3427, 2770, 3373, 2962, 1581, 1408, 2428, 2629, 1722, 3017, 3037, 1351, 2628, 1921, 3126, 1728, 910, 3003, 3387, 431, 515, 535, 493, 136, 247, 2942, 1499, 209, 864, 205, 235, 1225, 1026, 1399, 1158, 1787, 2901, 619, 3295, 1701, 3099, 3100, 2995, 1, 1685, 3350, 2943, 1508, 2435, 873, 3080, 950, 1171, 1345, 114, 686]. random [1045, 1023, 1463, 936, 3, 2908, 1124, 930, 3190, 1712, 886, 1388, 1955, 914, 1343, 506, 885, 872, 990, 1233, 1479, 391, 1690, 1206, 587, 270, 291, 1797, 2984, 896, 1284, 1168, 360, 3006, 1954, 954, 1407, 1426, 1046, 2930, 945, 609, 572, 614, 648, 655, 656, 794, 2662, 2855, 3208, 3367, 3303, 1812, 2921, 1235, 702, 754, 687, 827, 725, 698, 1913, 1597]. random-effects [3387, 1685, 1913]. randomization [1034, 260]. randomizations [3401]. Randomized [2614, 312, 3340, 260, 449, 748, 2453, 547, 561, 923, 1606, 1338]. randomizing [923]. Randomly [2534, 2964, 2458, 1718]. randomness [1131]. Range [2199, 1772, 2400, 248, 524, 368, 337, 511, 141, 176, 2426]. ranges [282, 1777]. Rank [1257, 1526, 2054, 2305, 1273, 2173, 1476, 802, 1131, 1751, 709, 1719, 1937, 1900, 363, 717, 3174, 829, 488, 218, 3345, 1430, 527, 822, 1302, 541, 318, 402, 793, 2682, 1291, 1342, 1392, 700, 830, 1058, 2752, 1606, 1718, 1812, 560]. rank-based [3174, 2682]. rank-sum [318, 402]. Ranked [2560, 2172, 2059, 2323, 2324, 2627, 768, 3232, 2856, 2608, 1880, 2670, 3178, 1933, 1924]. Ranked-Set [2323, 2856]. ranking [928, 638, 786, 297, 715, 891, 699, 3389]. rankings [2926]. ranks [649, 1686, 1780]. Rao [2221, 348, 366, 683, 1442, 1003]. Raphson [1313]. rare [2974, 379]. Rasch [3108]. Rate [2016, 2280, 2273, 2265, 2941, 1564, 2990, 1909, 2821, 1496, 414, 2416, 2668, 2883, 3036, 3165, 2811, 3371, 3135, 1537, 1561, 1173, 1785, 1302, 511, 1733, 1364, 3103]. Rates [1368, 703, 909, 2152, 789, 2218, 2718, 1245, 3058, 3243, 1983, 1510, 1153, 1187, 1465, 1965, 2862, 1445, 1727, 2824, 1893, 1812]. rating [3406]. Ratio [2341, 2319, 2542, 2347, 2245, 2058, 2270, 2249, 2191, 3255, 709, 1370, 125, 2851, 1855, 1802, 50, 807, 813, 1939, 3370, 1133, 634, 951, 1456, 1589, 1837, 864, 524, 756, 915, 1834, 1585, 644, 1256, 82, 3233, 1822, 1422, 680, 842, 2636, 936, 353, 1211, 1766, 1867, 3209, 1835, 405, 492, 1128, 1024, 541, 1273, 780, 163, 1071, 2885, 1391, 367, 1989]. Ratios [2541, 2706, 1679]. ray [3157]. Rayleigh [2020, 2360]. Real [1919, 980, 1007, 1295, 498, 865, 2752, 3298]. realization [3397]. realized [3202]. Reasoning [2214]. recapture [245, 2958]. reciprocal [575]. recombination [2719]. reconstructed [915]. reconstruction [2694, 3323]. Record [1005, 1094, 1568, 2454]. records [1231, 2434]. Recovery [650, 1065]. Recurrence [1998, 1543, 1479]. Recurrent [2084, 1948, 2416, 2668, 2843]. Recursions [2011]. Recursive [2923, 2586, 2729, 1602, 2098, 2263, 3222]. Recursively [2456]. reduced [1611, 1116]. Reducing [2193, 2842]. Reduction [2337, 979, 796, 2722, 915, 920, 890, 968, 1873, 3206, 1734, 3208, 2427, 2902]. redundancy [1631]. reference [1527]. refined [1398]. reflected [3254]. regard [94]. region [1484, 489, 526, 633, 774, 784, 1195]. Regions [955, 993, 1813, 1411, 1808, 1782, 253, 472, 3117, 462, 422, 2721, 340]. Regression [2061, 2578, 2203, 2146, 2303, 2337, 3264, 2110, 2283, 2349, 2537, 1503, 2216, 2007, 2163, 2591, 2541, 3065, 2581, 3196, 2521, 755, 2582, 263, 2345, 2264, 2147, 2536, 2534, 2085, 2133, 2262, 2522, 3392, 2225, 2034, 2265, 2098, 2297, 2263, 2177, 2535, 2556, 2693, 3342, 802, 854, 2702, 1755, 2800, 2699, 1953, 1950, 3160, 3423, 3280, 125, 144, 1814, 2822, 2665, 3174, 1713, 3066, 1709, 2615, 1802, 3236, 489, 1216, 3328, 2679, 1915, 2512, 1405, 1813, 1845, 3229, 1078, 1257, 3152, 2700, 3243, 2807, 2440, 1547, 2441, 2370, 2351, 2829, 1856, 2603, 2664, 2701, 2784, 2886, 3011]. regression [3314, 3285, 3374, 3364, 3017, 3403, 2602, 3262, 1177, 2389, 2791, 2835, 1750, 2965, 1290, 3326, 893, 1226, 1334, 3084, 3221, 3220, 1444, 365, 1968, 2734, 2414, 2900, 3391, 3147, 2398, 46, 1355, 1627, 3286, 1951, 3249, 3348, 855, 311, 1175, 1961, 3256, 3321, 2371, 2766, 1683, 3061, 1569, 2425, 2904, 2959, 2764, 2882, 3368, 3191, 3078, 3345, 3009, 1764, 1558, 644, 90, 3033, 3183, 3274, 2625, 894, 2499, 3281, 3010, 2758, 2947, 1655, 3334, 2415, 2663, 3042, 3361, 2393, 1518, 1712, 2893, 3309, 1395, 1965, 1641, 1329, 1420, 1413, 2363, 36, 585, 600, 2938, 1247, 1387, 3127, 780, 1303]. regression [1445, 238, 1761, 620, 700, 976, 1058, 1879, 3219, 3332, 342, 505, 610, 741, 2930, 3151, 2759, 1847, 3385, 3384, 3375, 2502, 1680, 2426, 3217, 3299, 1846, 1893, 2674, 1902, 1966, 1232, 2736, 3421, 2961, 2797, 3435, 3400, 1765, 3085, 2633, 2844, 2954, 3111, 3234, 2929, 3426, 3079, 1711]. regression-based [3368, 3299]. regression-type [1965]. Regressional [2024]. Regressions [2465, 3003, 816, 3134]. regret [1271]. Regular [2339, 2081, 1156, 1646, 1970, 2891, 2746, 2460, 2738, 2806, 645, 3118, 374, 435, 954]. regularity [1003, 353]. regularization [3281]. Regularized [3316, 3435, 2676, 2699, 2499, 3289, 3288, 2938, 3127, 3409, 3336]. regularizers [3112]. regularly [1914, 1098]. reinforced [2795]. reinforcement [1367]. reinforcement-depletion [1367]. Rejection [2167, 910, 1233, 189]. Rejoinder [1637, 3418, 3141, 3356]. Related [2548, 2215, 2333, 2201, 2153, 2023, 2187, 3075, 3187, 2922, 1543, 1778, 244, 1163, 846, 2884, 3192, 1228, 2763, 2466, 323, 1229, 263, 3202, 1152, 1789, 193, 1696, 1591, 1324, 1922, 238, 1844, 1679, 576, 1597].

Relatedness [1007]. Relating [1916, 1411, 174, 36]. Relation [147, 591, 1037, 91, 20, 1366, 3018, 1116]. relational [3412]. Relations [2092, 1998, 1543, 1707, 731, 1626, 1185, 863, 1479, 1613]. Relationship [1492, 1509, 1447, 1371]. Relationships [1778, 1982, 2049, 1410, 1963]. Relative [1790, 1389, 1176, 2428, 488, 2738, 1269, 2626, 1150, 289, 373]. release [2422, 2884]. relevant [3337]. Reliability [2101, 3195, 2017, 653, 139, 2352, 1802, 3188, 334, 1759, 631, 2642, 1580, 1844]. remained [799]. Remark [2065, 231, 564, 46, 166, 57, 78, 8, 14, 229]. Remarks [1200, 766, 530, 408, 399, 775, 600, 2637]. Remarque [229]. reneging [434]. Renewal [2127, 653, 1992, 782, 1563, 1853, 873, 503, 306, 1875, 2810]. renewals [325]. Renyi [196, 2738, 567]. Rényi-type [567]. repairable [1578]. repartition [229, 229]. Repeated [2194, 823, 2798, 3062, 2438, 156, 1443, 3436]. repetitions [3187]. Replacement [2258, 688, 313]. Replenishment [2052]. Replenishment-Depletion [2052]. Replicate [2774, 669]. replicated [597]. replicates [776]. Reply [3227]. Report [3282, 41]. reported [216]. Representation [2060, 7, 13, 2171, 817, 767, 3150, 732, 544, 1039]. Representations [2061, 2796, 2594, 2606]. representing [977]. reproducing [3280, 3409]. required [1393]. requiring [425]. Resampling [2402, 2493, 2893, 2391, 1588]. Resampling-based [2893]. Rescaled [2309]. research [824, 826, 84, 385]. Residual [2057, 2236, 2365, 1964, 1311, 3034, 3214, 2947, 1963, 1849, 1421]. Residuals [1897, 2644, 926, 2506]. resistant [1592, 1454]. resolution [1971, 777, 1376, 1453, 2969, 548, 862, 1018, 1708]. Resolutions [2550]. resolvent [433].

Respect [2108, 1856, 809]. Response [2144, 485, 631, 2251, 277, 377, 3145, 637, 960, 3003, 385, 1218, 2764, 1821, 1374, 1073, 1150, 1815, 1580, 623, 923, 2353, 3208, 1088]. response-adaptive [3145]. response-selective [2764]. responses [2934, 3309, 995, 2930, 1415, 2437, 3303]. responses-multiple [995]. restoration [1635, 1826]. Restricted [2343, 2322, 1468, 3212, 2952, 1033, 1681, 1587, 3060, 790, 2559, 2453, 312, 464, 1606, 3004]. restriction [1566, 1936]. restrictional [548]. Restrictions [2317, 2517, 517, 311]. result [1197, 495, 1149, 3102, 1228, 1518]. Results [2539, 2095, 887, 835, 2976, 594, 684, 3262, 1788, 2454, 855, 3167, 368, 1146, 2642, 622, 566, 881, 2597]. retracted [1173]. return [246]. Reversed [2331, 2273]. reversibility [3402]. reversible [2912]. review [2974]. Revisited [2046, 1592, 2485, 3291]. rhinoceros [1946]. rho [3070]. ridge [3281, 3332, 3103]. ridit [3188]. Right [2073, 2264, 1998, 3048, 1982, 3301, 3165, 2906, 3034, 2987, 2641, 2682, 3332, 2735, 2467, 2632, 3320, 3077]. Right-Censored [2073, 3048, 3301, 3034, 2682, 3332, 3320]. right-truncated [1982]. Rises [2185]. Risk [2339, 2025, 558, 1380, 183, 1496, 3110, 1750, 1545, 908, 2776, 2471, 3172, 1222, 1150, 1717, 1242, 1483, 2852, 1828]. risks [3370, 2787, 3147, 1073, 2838, 3067]. RKHS [3116]. Robust [3258, 2533, 1684, 491, 1833, 3050, 3012, 3383, 3276, 3313, 1275, 1683, 1826, 3203, 3394, 3350, 3345, 3223, 3404, 736, 3072, 3309, 1770, 3151, 3411, 1517, 2954, 2929, 802, 3271, 2866, 1306, 2698, 3064, 3171, 3170, 3056, 596, 2716, 3375, 3161, 1614, 3047]. robust-efficiency [596]. Robustness [2145, 2558, 2179, 1084, 784, 1279, 2746, 2618, 1878, 2955, 1967, 1453, 1642, 1020, 2860, 3389]. ROC [3162, 2818]. Rocky [2161]. role [1034]. roles [1994]. Root [2041, 2058, 2985, 3216, 431, 535, 743, 652, 635, 682]. Roots [2160, 765, 939, 972, 952, 439, 515, 676, 860, 940, 376, 516, 953, 1108, 665, 681, 3405, 869, 999, 440, 545, 546, 518, 1054]. rose [3407]. rotatability [1815]. rotatable [450, 497, 508, 1973, 839, 1184, 1293]. rotational [1234]. Rotations [2548]. round [3364]. rounded [2768]. row [1427, 1660, 1620, 650]. row-column [1660]. Roy [1197]. ruin [781, 2779]. Rule [2551, 658, 82, 111, 131, 1572, 1162, 1052, 804]. Rules [2210, 2109, 331, 3251, 2571, 1322, 103, 350, 1529, 432, 1114]. Run [2599, 2130, 3075, 2408]. run-related [3075]. Runs [2103, 2104, 2210, 2122, 2256, 388, 2036, 2307, 2332, 2011, 2131, 2598, 1941, 2010, 2090, 1907, 1999, 3031, 2484, 3099, 3100, 2546, 1884, 2430, 944, 2067, 3245, 1235]. Ryukyu [41].

sacrifice [2452]. Saddlepoint [2621, 2622, 2133, 1588, 2334, 2989]. saddlepoint-based [2989]. safe [1103, 1091]. safety [3054]. Sakamoto [1943]. sales [2651, 2715]. same [249, 3397, 468, 328]. same-realization [3397]. Sample [2545, 2151, 2268, 2518, 854, 3240, 803, 897, 2948, 262, 1142, 1751, 445, 446, 581, 2730, 1384, 1747, 2777, 1272, 3242, 857, 487, 1754, 1267, 1268, 2778, 2481, 35, 216, 1526, 1101, 3335, 2413, 1759, 3262, 1799, 3126, 1969, 1117, 2608, 218, 1147, 2888, 2841, 639, 412, 721, 761, 69, 1673, 3347, 2896, 208, 2497, 678, 1236, 767, 1784, 3053, 428, 1248, 1034, 1322, 1880, 3063, 269, 2783, 368, 749, 844, 1469, 3172, 1146, 462, 1283, 3273, 630, 3, 236, 2670, 283, 3329, 452, 447, 476, 1835, 617, 480, 509, 1955, 1395, 752, 287, 402, 403, 1342, 83, 187]. sample [976, 1071, 3392, 56, 471, 1698, 1718, 609, 572, 614, 165, 189, 737, 2632, 1234, 540, 3320, 687, 1039]. Sampled [2026, 213, 3150]. sampled-data [213]. sampler [3179]. samplers [2973, 2972]. Samples [2560, 2059, 2324, 200, 1299, 459, 61, 359, 1164, 1015, 1428, 2373, 1390, 243, 3232, 122, 1645, 2608, 694, 1500, 103, 744, 843, 942, 2670, 3122, 3178, 895, 1780, 141, 2984, 1777, 356, 539, 613, 612]. Sampling [139, 2321, 2135, 2552, 326, 1122, 24, 26, 2172, 2258, 2026, 2285, 2017, 2241, 2047, 1879, 2059, 2219, 2323, 1279, 658, 2624, 37, 75, 123, 302, 688, 3230, 3074, 425, 2627, 437, 813, 2451, 2856, 3126, 499, 3211, 633, 867, 153, 2974, 1225, 1158, 3149, 2425, 2764, 3393, 2667, 98, 1056, 9, 15, 40, 3178, 1933, 257, 289, 313, 480, 914, 878, 391, 1206, 621, 320, 2984, 719, 466, 582, 1924, 899, 360, 203, 56, 371, 2412, 648, 655, 656, 2935, 99, 827]. SAR [2969]. satisfying [484, 555, 588]. saturated [711]. Savage [1080]. scalability [3226, 3225, 3224, 3227]. scalar [1015]. Scale [2343, 2150, 2086, 2586, 2022, 1154, 3330, 659, 2851, 1246, 510, 1530, 575, 467, 1488, 2734, 1254, 1865, 2982, 2574, 1644, 1834, 1288, 1585, 3117, 432, 236, 3214, 2559, 3056, 476, 617, 1179, 1172, 402, 2650, 1449, 1392, 789, 1734, 2463, 3004, 1415]. scale-alternatives [1392]. scale-free [2650]. scaled [1594, 1952]. scaling [986, 995, 1022]. scan [3167, 2464]. scans [1941]. Scatter [2083, 2370, 3424]. scheduling [151]. scheme [688, 1069, 2361, 1011, 1251, 2327, 2017]. schemes [2695, 425, 3126, 979, 1308, 3169, 435, 1056, 776]. Schweder [3340]. science [2648]. Score [2335, 2607, 1470, 2557, 1751, 2796, 3212, 1967, 3029, 1697, 1680]. Scores [2207, 598, 626, 640, 830]. scoring [3251, 2571, 3103]. Scott [841, 2977]. screening [3212, 3082, 3142, 1294, 3059, 3343]. SDE [3422]. Search [1632, 3339, 3302, 949, 2409, 1708, 946]. searching [1632]. Second [1402, 1482, 3052, 2061, 2197, 2143, 2587, 2108, 2657, 2113, 2705, 2739, 3265, 1086, 1270, 3338, 839, 3015, 2872, 1980, 1815, 440, 1947].

Second-Order [2143, 3052, 2587, 2657, 2705, 2739, 3265, 1086, 2872, 1980, 1947]. Section [2290, 2648, 482]. sections [1954]. seemingly [816, 1175, 3061, 620, 976]. Segmentation [2292, 3357, 3432, 2592]. Segrè [2727]. seismic [2447, 1664]. Seismograms [1676]. selected [1503, 1298, 1735]. Selecting [2120, 581, 658, 1258, 1504, 2389, 2791]. Selection [421, 662, 2871, 2346, 2175, 2591, 1100, 2398, 2106, 1387, 2030, 2056, 2025, 2159, 649, 2730, 2821, 68, 475, 3271, 1987, 659, 3352, 1385, 1321, 1427, 1741, 3229, 928, 2712, 2413, 1706, 1149, 3257, 3341, 3349, 2772, 871, 786, 297, 1033, 1444, 3040, 1254, 3391, 3038, 3397, 2766, 3120, 3300, 3333, 3350, 3345, 2717, 3351, 3009, 1548, 3068, 3033, 1146, 3032, 1374, 1332, 3010, 2913, 3312, 2737, 1247, 989, 3148, 716, 1761, 1162, 360, 2876, 3151, 1099, 3380, 891, 3407, 1716, 3136, 3144, 3306, 3411, 2674, 3105, 3215, 2736, 3372, 2844, 2954, 3111, 2833, 3429, 3121]. selection-biased [3010]. selections [2727]. Selective [3380, 3381, 2764, 3375]. selector [3349, 2626]. selectors [3244]. Self [3146, 2247, 1240, 2658, 1586, 3261, 1795, 1025]. Self-Consistent [2247]. self-correcting [1586, 1795]. self-decomposability [1240]. Self-exciting [3146, 3261, 1025]. Semi [2658, 2764, 3274, 1835, 1291, 3366, 2442, 2634, 2832, 3147, 3261, 1310, 3386, 593, 2685, 1952, 3137]. Semi-aligned [1291]. semi-competing [3147]. Semi-empirical [1835]. semi-Markov [3366, 3386]. semi-norm [2685]. semi-or-used [1310]. Semi-parametric [2764, 3274, 2832, 3261, 1952, 3137]. Semi-self-decomposable [2658]. semi-Sibuya [2634]. semi-stable [2442, 593]. semicontinuous [3346]. semidefinite [2742]. semimartingale [3022, 3322]. Semiparametric [3201, 1814, 3236, 3301, 3247, 3076, 2216, 2835, 2400, 3256, 3176, 2478, 3219, 3378, 3134, 2797, 3346, 2557, 2556, 3048, 2936, 2992, 1526, 2883, 2480, 2787, 2492, 3078, 3062, 2554, 2876, 2930, 3161, 3400, 2844, 2954]. Semistable [2132, 2511]. semivarying [3144]. Sen [688]. sense [1555, 1153, 1187]. Sensitivity [1164, 2135, 2338, 2472, 1979, 2297, 1882]. Sensors [2239]. separability [2825]. separate [1401]. Separation [759, 1858, 983]. Sequence [2036, 2332, 2530, 2220, 2131, 2184, 2598, 2094, 2718, 1130, 1314, 1721, 2865, 2837, 571, 2419, 2684, 2546, 2000, 789, 1235]. sequence-compound [789]. Sequences [2331, 2765, 2607, 2021, 1080, 1909, 2483, 2358, 2484, 1627, 2901, 3099, 3100, 424, 1598, 994, 2768]. Sequential [1405, 1163, 1983, 2235, 2016, 1832, 3235, 1558, 2202, 790, 2224, 324, 2062, 658, 2849, 3054, 2643, 3074, 425, 1894, 1003, 2422, 1515, 3229, 3294, 740, 1471, 2380, 3040, 1355, 1624, 1996, 801, 2968, 3089, 1864, 1980, 3172, 1079, 1095, 2972, 1905, 753, 2678, 1406, 1770, 848, 1335, 715, 1844, 416, 455, 266, 2935, 3372]. sequentially [1983]. Serial [2054, 2305, 2496, 2789]. Series [2223, 2585, 2316, 2226, 2290, 2225, 2282, 1668, 1036, 1181, 2402, 491, 3265, 2796, 3269, 226, 3337, 2496, 3270, 2401, 1092, 3226, 1226, 1334, 1491, 1093, 1667, 2652, 3362, 3276, 3432, 997, 2789, 3070, 1317, 1569, 2384, 3430, 2779, 2364, 1139, 900, 1056, 90, 2760, 3225, 955, 1042, 3214, 294, 1873, 1817, 3433, 795, 1285, 1395, 1922, 1708, 443, 2786, 1664, 2715, 1947, 1476, 1816, 1847, 1132, 1263, 1464, 3224, 3227, 2961]. served [1205]. server [947, 1328, 1116, 1613]. servers [414, 551, 1890]. service [414, 727, 415, 603, 521, 1613]. Set [2560, 2172, 2276, 2059, 2323, 2324, 2627, 3232, 2856, 2608, 1236, 1593, 311, 3129, 1339, 598, 565, 1880, 1222, 969, 1027, 1112, 1173, 2670, 3178, 1933, 622, 1924, 543]. set-compound [969, 1027, 1112]. set-valued [3129]. sets [1543, 1778, 1945, 673, 651, 2942, 671, 1102, 1544, 566]. setting [2948, 3163, 2424, 129]. settings [2871, 3273]. Settlement [1335]. setup [3392]. Several [2238, 2317, 2011, 2347, 2082, 2097, 2072, 2181, 1658, 1015, 3166, 1135, 978, 902, 2660, 1542, 2359, 441, 552, 611, 818, 1823, 1127, 1513]. Shannon [1566, 210, 274]. Shape [110, 3058, 2367, 2896, 2491]. shape-adjusted [3058]. shapes [2904]. Sharp [3239, 2022]. sharpness [296]. shift [2998, 2738, 2806, 2649]. shifted [2612, 3406]. shifting [1310, 1994]. shifts [3314]. Shirakawa [26]. shock [1992, 2405]. Shocks [2130]. Short [2298]. Short-Memory [2298]. shortness [315]. shot [3101]. shot-noise [3101]. Shrinkage [2346, 1804, 2655, 2424, 1545, 948, 925, 2574, 1834, 1863, 1643]. Sibuya [2634, 3164]. Sided [2019, 2156, 2157, 1574, 1900, 1029, 691, 1939, 198, 1531, 2409, 2368, 1770, 381, 2518, 2113]. siécle [338, 352]. sightings [1423]. Sign [2193, 2192, 3312, 215]. sign-constrained [3312]. Signal [1694, 3271, 2424, 2919, 1169]. signals [1650, 3130, 1858, 2444]. Signed [2305, 1900, 363, 1430, 1780]. significance [49, 553, 965, 1064, 447, 100, 93, 165]. signs [1235]. Simes [3088]. similar [321]. Simple [2335, 2073, 2070, 122, 2530, 2087, 1301, 1953, 1496, 2671, 2851, 2701, 3114, 136, 1586, 2653, 914, 2527, 1781]. Simple-Tree [2073]. simplified [3379, 303, 625]. Simulated [2261, 2971]. Simulation [1793, 3231, 1734]. Simultaneous [2238, 1356, 1556, 3328, 1467, 1243, 3211, 271, 665, 681, 3361, 2097, 1127, 2736, 2181, 107, 868, 2500, 740, 67, 616, 375, 362, 1447, 3197, 2368, 413, 422, 793, 881, 664, 912, 1097, 1735]. sine [478]. Single [2035, 1998, 3258, 3388, 3174, 244, 3319, 3152, 1738, 1354, 3300, 3086, 3078, 3009, 3218, 3041, 947, 1328, 3171, 3170, 899, 3219, 699, 3026, 1116, 1613, 2911, 3173, 3253, 3111, 2833, 2703]. Single-Index [2035, 3258, 3152, 3300, 3086, 3078, 3009, 3218, 3041, 3171, 3170, 3219, 3026, 3173, 3253, 3111, 2833, 2703]. single-stage [699]. singly [35, 1830]. singular [404, 3021, 2708, 624, 711]. singularity [1795]. sinh [3016]. sinh-arcsinhed [3016]. Situations [2089, 1566, 2605]. six [1029]. Size [2325, 2076, 2259, 2600, 2391, 2808, 3126, 1044, 1673, 1322, 1432, 1508, 110, 3273, 480, 1955, 118, 56, 1954, 245]. Sizes [2518, 1101, 1660, 3080, 3108, 40]. Skew [2547, 2877, 2486, 2429, 2431, 1820, 477]. skew-elliptical [2486]. skew-normal [2429, 2431]. skew-symmetric [2877, 1820]. Skewness [2288, 1199, 3202, 1541]. Slash [3398, 3250]. SLEX [2291]. sliced [3114, 3156, 2625]. slippage [1488, 1275, 420]. slope [3183]. Small [2230, 2632, 3271, 868, 2750, 2888, 2841, 3308, 2782, 512, 2463, 165]. small-disturbance [868]. Small-sample [2632]. smallest [1108, 652, 543]. Smirnov [897, 751, 2018, 691, 175, 898, 237, 321, 2318]. Smirnov-test [897]. Smirnov-type [751]. Smoluchowski [2046]. Smooth [2234, 3066, 2281, 3017, 2117, 3327, 2225, 3379, 3247, 1688, 2390, 2817, 3112, 2502]. Smoothed [2027, 3168, 3123, 3364, 1784, 1819, 1892]. smoother [1872, 2968]. Smoothers [2299]. Smoothing [2714, 2416, 2581, 1589, 2194, 1845, 1547, 2819, 2631, 1649, 2398, 1760, 2375, 1872, 1826, 2812, 2638, 3159, 2692, 3331, 2863]. Sobel [1052]. Sobolev [3177]. Social [149, 26, 1266, 77, 62, 143, 222, 193, 304]. sociale [304]. sociales [222]. societal [824, 826]. Sociometric [153]. soft [1562]. soft-core [1562]. Software [2101, 2422]. Soil [2161]. sojourn [2906]. Solution [234, 107, 28, 1991, 638, 393, 886, 1957, 1338]. Solutions [2336, 2890, 3285, 3049, 67, 3348, 395, 622, 1437, 722, 975]. solving [1313, 182]. Some [810, 1350, 735, 1776, 2746, 2877, 2980, 823, 838, 1309, 1455, 475, 70, 659, 857, 835, 1754, 647, 1143, 1358, 510, 2179, 3353, 1568, 1411, 1722, 3285, 508, 1612, 2884, 871, 779, 2454, 2519, 3272, 3354, 860, 769, 252, 577, 1184, 865, 2271, 147, 1320, 1347, 1593, 855, 568, 619, 1120, 1454, 642, 1326, 112, 1331, 2087, 2155, 2041, 1432, 1255, 2646, 2182, 1617, 900, 1171, 98, 641, 114, 199, 931, 1222, 1889, 1146, 744, 843, 2642, 2008, 55, 193, 174, 2133, 1371, 622, 795, 548, 1172, 1213, 600, 1019, 1911, 848, 444, 482, 2313, 566, 2637, 400, 715]. Some [1009, 3355, 1452, 612, 2412, 995, 891, 1052, 818, 265, 2597, 1132, 1716, 1170, 1419, 2286, 1369, 458, 2068, 802, 133, 224, 1575, 3075, 1600, 1778, 1440, 861, 425, 1145, 3294, 2428, 1731, 1631, 740, 829, 2608, 2390, 553, 839, 1305, 27, 113, 1159, 317, 53, 77, 384, 465, 616, 1620, 2425, 567, 3142, 775, 441, 1056, 1294, 618, 704, 542, 911, 1241, 599, 387, 1125, 1963, 1854, 1324, 1690, 1573, 379, 463, 1060, 771, 342, 457, 266, 76, 1533, 1353, 521, 3128, 204, 1798, 3138, 1895, 687, 1553, 222]. sometimes [216]. Sooner [1999, 2104, 2184, 1920]. SOUPS [1310]. source [3268, 3252, 3308]. southern [1665]. Space [2221, 2101, 2874, 2216, 2217, 3261, 2077, 2219, 2063, 2718, 2651, 1363, 3133, 3280, 3074, 2714, 2490, 3114, 2379, 3338, 2951, 517, 2656, 921, 229, 2631, 1649, 1961, 3005, 1693, 630, 3246, 760, 3409]. space-filling [3114]. Space-Time [2077, 3261, 2718, 1693]. Spaced [2262]. Spaces [2169, 2869, 2960, 3243, 2395]. spacing [252]. Spacings [3013, 2105, 981, 595, 2465, 3096, 1049, 236, 1559]. spanning [1854]. Sparse [2996, 2483, 3212, 3229, 2858, 3120, 3121, 3116, 3259, 3344, 3194, 3329, 1733, 2406, 3399, 3241, 3237]. sparsity [3212, 3002, 3399]. sparsity-restricted [3212]. Spatial [2060, 2160, 2064, 2644, 1635, 3092, 3049, 1359, 2942, 1760, 3408, 2425, 1561, 1136, 1459, 1638, 2977, 3144, 3411, 1695, 3429]. spatial-temporal [3408]. Spatially [3248, 2239, 2969]. spatio [2478, 3331]. spatio-temporal [2478, 3331]. spatiotemporal [3323]. Spearman [829, 1429, 3070, 3433, 2508]. Special [2290, 2215, 2710, 1527, 94]. species [3106]. specific [2869]. specification [1177, 2850]. Specified [2210, 2598, 2090, 1444, 150, 2484]. spectra [276, 406, 473]. Spectral [2448, 885, 2224, 2292, 225, 996, 957, 956, 1877, 3008, 3063, 853, 1099, 932, 1066, 3402, 1860]. Spectrum [660, 213, 556, 591, 1714, 354]. Speed [1755, 1290, 3103]. sphere [1508, 885, 945, 1369]. spheres [247, 281, 1954]. Spherical [2026, 2177, 1683, 2646, 3181, 2435]. Spherically [2517, 1804, 3102, 1113]. sphericity [762, 1413]. Sphering [2068]. spike [2713]. spiked [3193]. spin [1858]. Spjøtvoll [3340]. Spline [3078, 3200, 1845, 2996, 1547, 2398, 2692, 2848, 2961, 2863]. spline-backfitted [2692]. Spline-based [3078]. splines [2500, 2819, 3125, 3056, 2638, 3159]. split [861, 1069]. split-plot [861]. Spots [347]. spread [903]. spreading [1792]. spring [823]. SPRT [1256]. square [1181, 61, 773, 2979, 950, 1867, 1323, 168, 998, 1391, 788]. Squared [2583, 2108, 2177, 2962, 1590, 1737, 789, 1127, 3411, 1695, 1336]. squared-distance [789]. Squares [2346, 2096, 2176, 970, 2452, 3425, 390, 1354, 2872, 2465, 1248, 1560, 2384, 2812, 2444, 1323, 1388, 1546, 3127, 912, 2759, 3385, 3384, 1415, 2657, 1286, 1633, 2903]. squeeze [1233]. Stability [2212, 1125, 2275, 2313, 2381, 2896, 554, 1060, 1137]. stabilization [413]. stabilizing [1425, 1304]. Stable [2561, 2171, 2051, 2023, 2176, 2442, 3383, 3295, 1691, 1381, 593, 706, 2903]. Stage [2143, 299, 1719, 1937, 2739, 3229, 3357, 1516, 1500, 1465, 969, 1112, 3178, 1032, 2438, 466, 582, 1223, 912, 3392, 699, 1116]. stage-wise [2438]. Staircase [2580]. standard [2504, 102, 1101, 1398, 283, 2913, 83]. Standardized [2341, 586]. star [2491]. star-shape [2491]. Start [2012, 2167, 2980, 3109]. Start-Up [2012, 2167, 2980, 3109]. State [2221, 2101, 2217, 2332, 2131, 2211, 2010, 2219, 2063, 2067, 2090, 1012, 2651, 2745, 3074, 2714, 2490, 2379, 3338, 2530, 2419, 2684, 1961, 2546, 2839, 3246, 1920, 521, 1068, 71]. state-dependent [2839, 521]. state-space [2651, 3074, 2714, 2379, 1961, 3246]. states [319, 3089, 2972, 306]. Static [2973, 748]. Static-parameter [2973]. Stationarity [2589, 3265, 2905]. Stationary [2552, 2584, 2291, 2224, 2225, 1012, 1115, 1278, 3360, 2890, 3265, 2490, 3427, 2355, 3278, 3049, 2761, 3338, 814, 2401, 3377, 303, 2506, 45, 2470, 997, 955, 1042, 988, 2916, 371, 1312, 1099, 1536, 1068, 182]. Statistic [2326, 2545, 149, 2319, 2004, 2069, 2318, 1970, 2891, 1106, 519, 1292, 3388, 1969, 858, 888, 2454, 3387, 2596, 971, 1589, 767, 366, 683, 3167, 3408, 3013, 398, 1167, 261, 1777, 1114, 1195, 560, 2948]. Statistical [653, 3434, 590, 2865, 3110, 2761, 2035, 2239, 2548, 2311, 3199, 2092, 1458, 2322, 3122, 1696, 2214, 3154, 1686, 1780, 1536, 1823, 2662, 3409, 2535, 2072, 1245, 172, 190, 277, 377, 537, 2711, 2926, 2914, 126, 425, 2377, 2569, 510, 3238, 3353, 3381, 1306, 2952, 3272, 23, 39, 1044, 3354, 2648, 84, 247, 934, 2780, 1034, 3223, 8, 14, 34, 51, 81, 2606, 95, 1629, 3355, 3297, 1307, 907, 265, 1132, 1170, 1419, 1552, 1846, 2576, 3014]. Statistics [2125, 2248, 2166, 2539, 2235, 2054, 2305, 2526, 2536, 2273, 2304, 2587, 2050, 2009, 2275, 2315, 2059, 2088, 2705, 1998, 1350, 3330, 1723, 1543, 1707, 1778, 1982, 2849, 3119, 475, 369, 714, 1440, 1635, 659, 1061, 2552, 807, 423, 1983, 2421, 3069, 594, 916, 992, 3424, 1503, 1368, 1523, 1938, 1852, 763, 808, 834, 829, 2390, 246, 2460, 695, 1305, 1425, 67, 3347, 832, 982, 1883, 1839, 615, 940, 2493, 1063, 300, 1043, 568, 279, 292, 1237, 3202, 1269, 2915, 567, 3089, 3259, 1276, 950, 1298, 2, 542, 1465, 1495, 1277, 1882, 448, 449, 2395, 1903, 2507]. statistics [1551, 2828, 2464, 1371, 822, 1302, 1817, 1570, 3154, 1259, 4, 1479, 787, 399, 217, 121, 134, 1230, 742, 457, 712, 798, 611, 2406, 1533, 1476, 1816, 321, 1680, 1068, 1567, 1148, 3091]. staturated [624]. status [2414, 2544, 2776, 62]. stay [314, 327]. steep [1830]. steepest [1661]. Stein [935, 1076, 3413, 1228, 1255, 2473, 1659, 1028, 1828, 1000, 1210]. Stein-type [1659]. Step [2084, 2194, 3365, 2671, 2849, 1604, 2526, 3101, 1746, 3331]. step-stress [2671, 2849]. Stepwise [998, 2861]. stereological [1611, 1666]. stick [3335]. stick-breaking [3335]. Stieltjes [7, 13, 18]. Stigmatized [2325]. Stirling [382, 1511]. Stochastic [2695, 1992, 1915, 2274, 2153, 2161, 2254, 1460, 182, 817, 1115, 3071, 3242, 370, 1991, 523, 607, 1481, 2666, 2951, 858, 888, 2986, 2386, 323, 1281, 1497, 1949, 2425, 295, 1008, 1576, 1185, 2971, 886, 1793, 1442, 1615, 1284, 883, 1065, 197, 1312, 1694, 3396, 932, 1263, 3378, 702, 754, 2794]. Stochastically [698]. Stopped [2127]. Stopping [2210, 1529, 1016, 1851, 1095, 1052]. Storage [1204]. Stoyan [3408]. straight [1622, 113]. strata [1158, 3122, 56]. Strategies [2182]. strategy [2653]. Stratification [27, 1035, 2809, 2925, 143, 562, 719, 780, 851, 456, 436, 827]. stratifications [657]. Stratified [302, 914, 1233, 648, 655, 656, 827, 75, 2829, 3211, 867, 1158, 3122, 480, 391, 1206, 471, 613, 609, 572, 614, 687]. Stratum [2181, 1101]. Strauss [1787]. strength [2352, 3131, 1759, 2775, 1324, 837]. stress [2671, 2849, 2352, 3131, 1759, 2775, 2884]. stress-strength [2352, 3131, 1759, 2775]. Strictly [2890, 3049, 3278]. Strong [1036, 2869, 3238, 1144, 3084, 2901, 1794, 2371, 2482, 1288, 2105, 2463, 3007, 2263, 2502, 2679, 3341, 3326, 3196, 3221, 3220, 2653, 1171, 627]. strongly [3193, 2900, 3385, 3384]. Structural [2207, 2994, 1953, 1802, 3415, 3418, 1639, 3313, 3417, 3416, 1371, 2715, 788, 3103]. structural-equation [3415, 3418, 3417, 3416]. Structure [2218, 803, 1684, 1616, 1976, 3226, 821, 3228, 3225, 1876, 2381, 2715, 1704, 3224, 3227, 1913]. structures [3305, 1732, 3024, 477, 2675, 1679]. Student [2493, 1180, 163]. Studentizations [2082]. Studentized [1975, 1967, 1839, 1784, 287]. studies [1073, 1020, 1110, 2632]. study [1279, 75, 2428, 1631, 221, 824, 826, 247, 1665, 3213, 1310, 143]. studying [1908]. sub [2995, 1777, 56]. sub-fractional [2995]. sub-samples [1777]. sub-sampling [56]. subbases [2751]. Subclasses [628]. subcritical [1373]. subexponential [2922]. subfamily [1379]. subgroup [2776]. subhypotheses [1257]. subject [180, 113, 2503]. subjected [311]. submanifolds [1552]. submatrices [365]. Submodels [2069]. subordinators [2667]. subpattern [2620]. subpopulation [1934, 1332]. Subsample [1747, 3289, 3288]. Subset [1321, 1033, 2893, 716, 1735]. Subsets [2297, 1503, 760]. subspaces [1769]. Success [2210, 2599, 2256, 2036, 2003, 2332, 2011, 1884, 2067, 2090, 1907, 1999, 1164, 2484]. Success-Runs [2210, 2036, 2067, 1907]. Successes [2036, 2255, 2332, 2067, 1841, 1907, 2419]. Succession [2037, 1984]. successions [1921]. successive [1279, 195, 688, 3058]. Sudoku [3114]. Suetuna [570]. Sufficiency [1728, 1512, 499, 1529, 771, 1165, 592]. Sufficient [234, 2166, 2337, 279, 1432, 398, 399, 1068, 1633]. Sum [2328, 717, 1590, 1969, 2656, 209, 2506, 875, 3432, 1872, 469, 40, 1345, 114, 174, 318, 402, 1407, 1426]. Summary [3069]. Summations [2286]. Sums [2065, 2051, 1140, 1182, 1278, 595, 2756, 2456, 1396, 3000, 3377, 128, 1787, 2465, 337, 1885, 930, 511, 1690, 587, 954]. Sun [1089]. sup [3337, 2192]. sup-norm [3337]. supercritical [1331]. superefficiency [3215]. superimposed [1373]. superiority [1539]. Superpopulation [2181]. superposed [2977]. Supplement [631]. supplements [1009]. support [1356, 2602, 3273]. supported [3403]. sure [3082]. surface [1815, 2694, 707, 728, 747, 945]. surfaces [3162]. surname [1730]. Surrogate [3343, 3180, 3292, 3206]. Surrogate-variable-based [3343]. Survey [3129, 2241, 2543, 499, 24, 26, 347, 1423, 1665, 3393, 98, 5, 41]. survey-sampling [499]. surveys [1034, 5, 1071]. Survival [2533, 2229, 2250, 3201, 1325, 3301, 2439, 2452, 1244, 2830, 3332, 991, 3128, 1862, 3343]. survival-sacrifice [2452]. survivors [2636]. Suzuki [692]. sweeping [351]. sweeping-out [351]. Swing [2214]. Switching [2217]. symbol [723]. Symbolic [1570]. Symmetric [2116, 2336, 2517, 2172, 2009, 2323, 803, 2877, 1804, 1632, 1143, 2673, 3000, 3102, 1030, 431, 515, 535, 3295, 1603, 1820, 1943, 3181, 2669, 440, 545, 1768, 444, 482, 1113, 804, 2903]. symmetrical [1487, 875, 708, 374, 435, 622, 204]. symmetrically [200]. Symmetries [2153, 1234]. Symmetrized [1751, 3424]. Symmetry [2019, 2288, 2304, 1916, 1106, 1438, 1824, 2871, 1554, 751, 3031, 1939, 3019, 2875, 3254, 1683, 2646, 1439, 3097, 1995, 713]. synchronized [1858]. system [377, 417, 2352, 868, 1729, 1759, 1542, 1447, 551, 1053, 40, 3195, 1901, 552, 967, 314, 327, 521, 1116]. systematic [37, 67]. Systems [2103, 2293, 2153, 1594, 504, 537, 1958, 2643, 309, 1578, 1983, 1675, 1266, 1563, 947, 1328, 9, 15, 1576, 1740, 313, 959, 415, 845, 962]. Szasz [3327, 982].

T [3434]. T-tessellations [3434]. table [1468, 388, 2858, 598, 1377]. Tables [200, 69, 141, 176, 2624, 240, 249, 2743, 1150, 757, 2751, 3018, 512, 2597, 2935, 1301]. tactical [797, 876, 622]. Tail [2075, 2529, 2231, 2955, 2842, 3260, 3383, 3359, 833, 2689, 3043, 3118, 3210, 2979, 2731, 3098, 2726, 1963, 1452, 903, 1103]. tail-first-order [3260]. tailed [2900, 3043, 2726]. tails [3357, 2387, 1091, 3079]. Taiwan [2447]. Takano [171]. takes [1256]. Tandem [148, 959, 845, 962, 1116, 1890]. tangent [3246]. taper [1860]. tapering [2942]. target [118, 3318]. Tchebycheff [178, 296]. Tchebycheff-type [296]. TE [553]. teams [2620]. technical [117]. technique [948, 925, 360, 923]. Techniques [2056, 2219, 2402, 1845, 2722, 1993]. Tempering [2261]. temporal [3408, 2478, 3331]. tending [1874]. tensor [1541]. Tensors [2443, 1640, 2752]. Term [2058, 2218, 1350, 799, 1060, 188]. terminal [1566]. terms [1166, 1607, 2852, 544]. territories [921]. tessellations [3434]. Test [2281, 2283, 2070, 158, 3210, 2087, 2058, 2495, 2526, 3430, 2030, 713, 2557, 897, 1350, 1937, 1900, 54, 61, 139, 388, 3158, 3242, 717, 1440, 487, 370, 2407, 782, 1246, 1390, 691, 1547, 2701, 2358, 3037, 1969, 1133, 1198, 1769, 1382, 1720, 218, 3387, 902, 1305, 6, 12, 3254, 69, 948, 925, 965, 1254, 2896, 208, 1404, 1271, 175, 940, 1199, 1398, 1593, 524, 1587, 2470, 3313, 770, 375, 2384, 3057, 3213, 2853, 1702, 3090, 1430, 3277, 775, 1276, 2731, 1822, 1439, 762, 819, 911, 1241, 1333, 1576, 2704, 2782, 2880, 3171, 3170, 3214, 2670, 283, 1020, 874, 1371, 3209, 308]. test [451, 1502, 1442, 2682, 1273, 998, 381, 285, 991, 237, 3412, 3320, 3421, 1913, 3173, 3253, 3402]. testability [926]. Testing [3198, 2579, 1619, 3071, 1325, 2914, 2160, 2203, 3166, 2825, 3188, 1845, 1004, 2020, 2886, 3011, 2496, 2236, 2529, 3019, 2965, 2205, 517, 859, 3267, 2689, 2111, 2140, 3070, 2959, 2048, 2359, 1927, 1064, 2304, 3010, 3194, 2313, 2520, 2562, 2071, 3039, 2072, 2693, 803, 1081, 2948, 1106, 2727, 2849, 2510, 857, 309, 3315, 1590, 3294, 1257, 1651, 696, 3275, 365, 971, 3094, 254, 258, 248, 1532, 525, 3321, 1486, 1202, 690, 2905, 1531, 644, 81, 1441, 449, 748, 2908, 2697, 3428, 3097, 1442, 2721, 896, 176, 512, 474, 907, 2827, 1148, 1431, 2595]. Tests [800, 1468, 2012, 2167, 144, 2019, 2317, 2203, 595, 2589, 2073, 2110, 2316, 2875, 3395, 2519, 2174, 2308, 2252, 2156, 2271, 2237, 2789, 2192, 2041, 2272, 2360, 2348, 1413, 2191, 2173, 2518, 2315, 1704, 802, 854, 1131, 1438, 1575, 1824, 1203, 1751, 1719, 1940, 3363, 2777, 2980, 1370, 1672, 49, 126, 363, 1015, 1836, 50, 751, 1029, 3021, 1417, 1899, 3031, 1939, 3413, 1503, 2906, 738, 739, 467, 3051, 1487, 488, 871, 2894, 553, 1605, 1967, 1968, 1837, 1806, 2698, 196, 3305, 1681, 645, 875, 1275, 1914, 362, 665, 681, 608, 583, 745, 3029, 2646, 3181]. tests [1310, 3168, 2475, 1469, 2817, 448, 1703, 757, 518, 599, 1054, 387, 447, 527, 617, 1179, 1928, 1213, 29, 36, 2449, 287, 318, 379, 402, 422, 463, 596, 793, 870, 1768, 1291, 1342, 1392, 1770, 896, 700, 830, 1058, 1391, 266, 1989, 1823, 1470, 1606, 93, 165, 215, 409, 458, 3161, 903, 1559]. th [1258, 2530, 3286, 516, 3100]. Their [2257, 2705, 810, 475, 3284, 1992, 892, 1358, 3370, 2486, 3205, 1779, 147, 645, 806, 616, 1153, 1187, 1269, 1949, 2744, 2979, 2082, 1730, 3233, 955, 1366, 1020, 795, 4, 2449, 1445, 2610, 995, 1132, 3245, 182]. Theorem [2054, 1080, 224, 632, 70, 1578, 1509, 984, 2489, 2394, 502, 766, 323, 319, 1990, 1943, 873, 889, 919, 111, 140, 207, 732, 587, 799, 364, 461, 760, 429, 397, 932, 1908]. Theorems [2127, 2026, 714, 1143, 1300, 1499, 982, 77, 1701, 1997, 1331, 873, 114, 1283, 1125, 76, 154, 788, 1132, 1263, 3322]. Theoretic [2245, 1294, 1642, 3055, 443, 2707, 1812, 2543]. Theoretical [3244, 191]. theories [995]. Theory [2248, 2558, 2046, 2311, 2312, 2170, 264, 2178, 1174, 3160, 557, 589, 632, 1145, 2377, 1061, 2976, 3315, 1427, 1055, 1411, 1729, 3036, 1306, 1260, 335, 1491, 2806, 67, 254, 258, 855, 319, 1, 8, 14, 34, 51, 9, 15, 3064, 1463, 3237, 2916, 527, 466, 582, 602, 65, 154, 1930, 1307, 3390, 1234, 1369, 99]. theory-aided [3064]. there [857]. thermal [1458]. thesis [798]. thinking [2711]. Third [1971, 2691, 1973, 1184, 1447]. Third-order [2691, 1184]. thoracic [1792]. those [435]. Three [2549, 723, 1632, 35, 763, 808, 834, 697, 69, 1184, 516, 690, 869, 546, 669, 2721, 1708, 912, 742, 3331]. three-associate [669]. three-factor [1632, 1708]. three-replicate [669]. three-sample [69]. three-stage [912]. three-step [3331]. Three-word [2549]. Threshold [2228, 1888, 2944, 2965, 3186, 3358, 2854, 1508, 1960]. thresholding [3297]. ties [208, 2613]. Tilted [2767]. Time [2104, 2210, 2331, 2599, 2126, 2122, 1363, 2255, 2223, 2585, 2057, 2236, 2216, 2316, 2530, 2162, 2184, 2211, 2226, 2290, 2327, 2010, 2077, 2314, 2225, 2094, 2282, 1668, 2718, 537, 1721, 1999, 2420, 3187, 2402, 2612, 2671, 491, 2354, 3360, 1888, 2950, 3265, 3143, 3269, 3337, 2761, 2496, 1729, 1964, 3076, 2906, 3338, 814, 3270, 2401, 2884, 2835, 3226, 1226, 910, 2656, 1491, 1667, 2652, 502, 1649, 2684, 3362, 3276, 3432, 997, 2789, 3070, 2000, 1372, 2384, 3430, 3261, 2364, 3414, 551, 1056, 2919, 90, 2760, 1693, 3225, 955, 1042, 2782, 2754, 3214, 1073, 3334, 1873, 1817, 3433, 1285]. time [1395, 959, 2430, 415, 443, 967, 314, 327, 1664, 2735, 2715, 1947, 1476, 1816, 1847, 1920, 521, 1132, 1263, 2848, 1464, 3224, 3227, 3245, 2961, 3402, 3073, 3067]. time-dependent [2354, 2835, 1372, 2735]. time-orthogonal [2761]. Time-Varying [2162, 1226]. times [1016, 1851, 1417, 2906, 1095, 1664, 2632]. timing [213]. timing-error [213]. Tippett [368]. together [112]. Tolerance [2560, 340, 253, 526, 633, 774, 136, 198, 462, 2540]. Tomography [2163]. top [2683]. topological [3106]. topologies [2719]. Toric [2606, 3311]. total [2408, 1960, 1209]. totals [1122, 1934]. trace [561, 623]. traces [438]. Traffic [156, 1758]. trains [2713]. transfer [1714]. Transform [2308, 1477, 1672, 3413, 3275, 2360, 587, 322]. Transformation [2128, 2939, 2941, 195, 3193, 2992, 1725, 2883, 427, 565, 529, 3274, 1045, 1023, 1771, 1995, 1718, 3077, 2788]. transformations [2842, 2870, 1425, 1304, 530, 1378, 2406, 1533, 1930]. Transformed [2315, 2942, 661, 428, 32]. transforms [1590]. transition [2612, 2906, 1248, 503, 1598]. Treating [1744]. treatment [290, 1398, 1461, 745, 3368, 3414, 2453, 1437, 265, 3420]. treatments [1398, 1593, 608, 583]. Tree [2073, 2292, 2628, 1406, 3106]. Tree-Based [2292]. Trees [2103, 2158, 3133, 2409, 1854]. Trend [1104, 2041, 717, 491, 2407, 1929, 2690, 1486, 2848]. Trials [2104, 2037, 2122, 2256, 2332, 2530, 2131, 2598, 1999, 2865, 25, 2837, 2419, 2684, 255, 2546, 2000, 1884]. triangular [3143, 3377, 525]. tridiagonal [3192]. trigamma [1048]. Trigonometric [2349, 2603]. Trimmed [1357, 2051, 2458, 1147, 1558]. trimming [1881]. Trinomial [690, 2612]. Triple [1516, 2047]. Truncated [1741, 2020, 2534, 2106, 2555, 200, 1998, 805, 3052, 1982, 1272, 500, 2842, 2456, 847, 3301, 35, 3165, 773, 1318, 1498, 301, 1043, 850, 2873, 3135, 361, 2374, 1985, 403, 2682, 2755, 513, 2467, 2953, 1830]. Truncation [2165, 378, 2048, 3076, 1779, 1499, 451, 391, 2735]. Tukey [887, 2338, 1882, 874]. tuned [1006]. Tuning [2175, 1628, 2733, 3421]. tuning-free [3421]. Twentieth [338, 352]. twice [2830]. Two [299, 2060, 2550, 2549, 2579, 1265, 2343, 437, 3357, 2332, 926, 2211, 697, 2156, 934, 2010, 2268, 2143, 3178, 3101, 912, 2518, 2067, 2090, 2113, 2194, 2062, 1081, 2948, 1094, 1574, 127, 1142, 2420, 1751, 445, 446, 581, 1719, 1937, 1900, 3162, 1468, 2746, 2739, 3382, 1543, 1778, 2790, 1455, 1370, 1635, 487, 309, 2851, 3028, 1632, 3229, 1121, 807, 1029, 333, 1004, 1526, 1055, 813, 3314, 3278, 1964, 1969, 667, 1769, 2608, 1720, 759, 488, 871, 3395, 867, 218, 2888, 2841, 986, 1214, 2896, 135, 208, 346, 564, 767, 1347, 279, 1657, 438, 2470, 1872, 1202, 2546]. two [708, 389, 507, 650, 1500, 745, 1880, 1853, 1410, 1585, 157, 1256, 103, 5, 1927, 1465, 819, 958, 1462, 1441, 2670, 1124, 545, 518, 1867, 1835, 313, 1329, 1420, 752, 585, 1273, 1342, 1686, 1708, 1770, 83, 466, 582, 623, 381, 789, 3392, 1475, 707, 728, 747, 1718, 1920, 1583, 93, 3378, 3320, 1116, 1890, 3123, 3346, 3173]. two-action [789]. two-and [1632]. two-associate [1455]. two-component [1475, 3378]. two-dimensional [2420, 1441, 707, 728, 747]. two-factor [2888, 2841, 1927]. two-filter [1872]. Two-level [2550, 2549, 2746, 2790]. two-parameter [1370, 1880]. Two-phase [437, 813]. two-piece [3382]. two-response [623]. Two-Sample [2268, 2948, 1142, 1751, 445, 446, 581, 1526, 1969, 2608, 218, 208, 767, 2670, 1342, 1718, 3320]. Two-Sided [2156, 2518, 2113, 1574, 1900, 1029, 1770, 381]. Two-Stage [2143, 3357, 3178, 912, 1719, 1937, 2739, 3229, 1500, 1465, 3392, 1116]. Two-State [2332, 2211, 2010, 2067, 2090, 2546, 1920]. Two-Step [2194, 3101]. Two-way [2579, 1468, 871, 3395, 1214, 346, 1347, 507, 650, 1273, 1686]. Tyler [2083]. Type [2033, 1997, 2123, 2047, 2718, 1574, 148, 1106, 1909, 1002, 2627, 751, 3130, 2808, 3316, 1386, 1503, 1092, 3012, 128, 1293, 296, 301, 2493, 3043, 2619, 1990, 374, 3080, 1204, 950, 736, 618, 1436, 3290, 1042, 894, 746, 1024, 3154, 1965, 778, 1659, 1737, 1652, 2804, 2803, 383, 954, 132, 2736, 2794, 2604, 3079, 1998, 2702, 2373, 2451, 3099, 2285, 567, 937, 1567]. Type-II [1998, 2702]. Types [2449, 986, 348, 435, 1462, 552].

Uhlenbeck [3163, 2843, 2903]. ultra [3212, 3391, 3200]. ultra-high [3391, 3200]. ultra-high-dimensional [3212]. ultrahigh [3142]. ultrahigh-dimensional [3142]. Umbrella [2272]. UMPS [3428]. UMVUE [1584]. unavailable [1738]. Unbalanced [2113, 1940, 1213]. Unbiased [2295, 1783, 460, 1153, 1187, 1335, 2518, 1352, 1448, 1900, 847, 647, 2404, 559, 1681, 424, 1317, 2468, 398, 405, 1128, 312, 848, 1242, 381, 471, 609, 572, 614]. Unbiasedness [2206]. unbounded [3207, 1889, 2993, 2864]. Uncatchable [2600]. Uncertainty [2575, 2092, 3226, 3292, 3225, 3224, 3227]. uncombined [1500]. unconditional [1539]. Undamped [262]. Underdispersion [2089]. underlying [1034]. undirected [2908]. undominated [1165]. Unequal [2080, 247, 3094, 1660, 525, 1432, 1730, 257, 313, 1780]. Unequally [2121]. unidimensional [227]. unification [1956]. Unified [2932, 2325, 2197, 3207, 1461, 1941, 3116, 1463, 927, 2885, 3399]. Uniform [2244, 2328, 2013, 1236, 749, 1152, 2270, 2758, 2824, 1299, 3071, 1078, 419, 1163, 2421, 1722, 1368, 1523, 2965, 564, 678, 865, 1237, 2646, 3181, 1298, 1365, 1045, 1023, 820, 1570, 2403, 1215, 2463, 945, 1001, 3367]. Uniformity [2281, 2236, 571, 3339]. Uniformly [2156, 2698, 1703, 281]. unilateral [448]. Unimodal [2105, 3258, 2673, 3102, 164, 1434]. Unimodality [2421, 2388, 879, 1288, 1811, 1511, 1772]. union [2435]. unionintersection [800]. Unique [973]. uniqueness [495, 2999, 3115, 2786]. Unit [2160, 2041, 2058, 2063, 3216]. Units [2167, 690, 1252]. unity [3405]. Univariate [2561, 1658, 3382, 3284, 1833, 3250, 1389, 2636, 663, 657, 1232, 1407]. Universal [2345, 2459, 2263, 3431, 3221, 3220, 3322, 2502, 2749]. Universally [2460]. universe [915]. Unknown [476, 3258, 3269, 2765, 2496, 792, 3254, 1254, 2982, 2687, 45, 690, 1191, 1585, 1201, 870, 1708, 896, 341, 1734]. unobserved [1623, 1807]. Unordered [2497]. unrelated [816, 1175, 3061, 620, 976]. Unrestricted [2343]. unsaturated [548]. unscented [2941]. unstable [1359]. Unstructured [2208, 1898]. Until [2210, 2036, 2067, 1841, 1907]. untruncated [1996]. upon [2167, 517, 524]. Upper [1750, 1166, 112, 3096, 1811, 283]. Urakawa [1676]. Urakawa-oki [1676]. Urn [2327, 2052, 1367, 1143, 2466, 3190, 1957, 2795, 1796, 1941]. Usage [2543, 3340, 1398]. Use [2079, 2536, 2214, 417, 473, 504, 924, 1090, 126, 633, 774, 1178, 1760, 282, 342, 923]. used [1261, 1262, 1327, 1325, 1310]. useful [636, 2642, 722, 791]. uses [764]. Using [2190, 2299, 2057, 2220, 2342, 2140, 2170, 2048, 2300, 2292, 2219, 2159, 3436, 2718, 1130, 1477, 1937, 2402, 2922, 3054, 1267, 1268, 2842, 2481, 637, 2370, 1725, 1101, 2479, 1530, 2480, 2379, 1416, 2866, 3370, 3019, 2973, 2809, 2856, 2925, 3050, 2581, 893, 3107, 3387, 3327, 661, 2398, 1532, 3149, 3231, 3408, 652, 1714, 1880, 2881, 1310, 3125, 1558, 5, 199, 2636, 2432, 2499, 3056, 1102, 3401, 2971, 994, 1206, 1746, 466, 582, 3219, 3392, 1176, 456, 436, 3206, 245, 2997, 1694, 3375, 2675, 237, 3136, 3409, 2736, 1103, 1860]. Uspensky [1509]. usual [341]. utilization [734, 163]. Uusipaikka [2368].

vague [2517]. Valid [3291]. validation [2389, 2353, 3025, 2613]. Validity [2204, 132, 1808, 1782]. Value [2325, 2529, 2252, 2045, 2118, 268, 1403, 1260, 2454, 2708, 187, 1452, 1424]. valued [2126, 3279, 3129, 3276, 3213]. values [1005, 1094, 3075, 2402, 216, 1568, 3335, 3340, 832, 311, 2364, 3168, 1927, 3291, 142, 132, 159, 737, 99]. vanishing [3311]. Variable [2175, 3341, 1865, 3136, 3144, 2844, 2833, 2859, 1012, 3352, 2351, 2712, 3257, 3349, 3391, 3415, 3418, 3120, 3121, 3345, 3351, 3009, 3417, 3416, 3033, 270, 291, 2505, 456, 436, 3151, 1443, 3306, 3411, 3105, 2736, 3343, 2954, 3111]. Variables [2328, 2205, 2121, 2094, 2520, 2286, 2320, 2181, 1036, 1130, 1181, 805, 127, 1721, 1984, 68, 106, 1394, 1543, 1707, 1619, 1278, 1134, 1428, 2428, 180, 3000, 2952, 1631, 2791, 488, 910, 3084, 1444, 113, 2698, 661, 209, 205, 235, 1951, 2901, 1480, 2619, 1, 3142, 873, 1753, 950, 1171, 1141, 1345, 114, 1045, 1023, 1374, 1463, 3281, 1124, 3361, 1867, 930, 1343, 506, 1329, 1420, 1479, 1690, 587, 1247, 1387, 456, 436, 3206, 954, 1407, 1426, 3385, 3384, 2562, 2688, 725, 2595]. Variance [2341, 2589, 2481, 2150, 2925, 1293, 1423, 2082, 811, 2009, 2241, 2063, 2932, 1448, 887, 1384, 1592, 847, 1051, 2679, 3269, 35, 1197, 2664, 2701, 1744, 3175, 1848, 1485, 1767, 2390, 2888, 2841, 1425, 812, 769, 2680, 625, 1085, 2687, 2722, 1304, 1653, 375, 1603, 1317, 1531, 1685, 2384, 1191, 679, 2760, 2836, 744, 843, 2642, 1811, 2970, 2559, 1032, 1903, 398, 2415, 257, 1717, 3055, 261, 4, 1659, 422, 870, 601, 320, 1911, 734, 100, 163, 1096, 664, 341, 2804, 2803, 666, 1483, 367, 610, 1734, 3322, 609, 572, 614, 648, 737, 1679, 2902, 573, 1514]. variance-balanced [1592, 1085]. Variances [2245, 1352, 106, 517, 301, 3096, 387, 413, 1629, 1780, 1775, 104]. Variants [2177]. variate [1238, 633, 860, 1239, 1348, 2417]. Variates [2209, 773, 128, 311, 427, 376, 531]. Variation [2139, 2151, 2886, 1540, 3118]. variations [1649, 2409]. variety [2862]. variogram [2730]. various [3172]. Varying [2964, 2668, 2162, 2071, 2194, 2169, 3160, 2934, 2500, 2952, 1226, 3283, 255, 1914, 3300, 2988, 2969, 2947, 621, 719, 3026, 3200, 3400, 3083, 2954, 3153, 3234]. Varying-Coefficient [2194, 2668, 3300, 3153]. Vector [2087, 1752, 2962, 1340, 971, 1689, 2943, 1697, 1064, 3273, 1625, 753, 1717, 698, 1000]. Vectors [2051, 2097, 1140, 1182, 2730, 980, 1007, 1295, 3037, 439, 3094, 686, 680, 842, 958, 1972, 3209, 1734, 1312, 2767, 2985, 3310]. Velden [385]. velocity [1694]. Verbal [2214]. Veronese [2727]. Versatile [3152]. Version [2338, 2584]. Versions [2057, 2389, 2508]. Versus [2109, 3177, 3325, 2453]. vertex [2385]. very [2743]. Via [2013, 2676, 2699, 3071, 1554, 3251, 533, 3379, 1845, 3424, 2424, 2811, 3347, 2405, 1853, 1310, 3345, 3218, 3117, 3032, 1885, 2368, 2913, 1995, 2721, 3148, 2786, 3380, 2795, 2848, 1902, 3399, 2935, 2831]. view [3423, 23, 39]. VII [2123]. Vinberg [3344]. vingtiéme [338, 352]. violations [1054, 1020]. virtually [2981]. vision [385, 323]. visual [385]. Vol [249]. volatility [2666, 3304, 3210, 3060, 2971]. volume [2396]. volume-corrected [2396]. volumes [3162]. Votaw [579].

Waiting

[1721, 2104, 2331, 2420, 2599, 3187, 2122, 2255, 2530, 2184, 2211, 2327, 2010, 2430, 2094, 1999, 2684, 2000, 551, 1920, 3245]. Wald [2804, 577, 3029, 8, 14, 2803, 1470]. Wald-type [2804, 2803]. Walk [2113, 2612, 493, 2995]. wandering [565]. war [1266]. Waring [1232]. warped [3326]. Wasserstein [3330, 3381]. Watson [1373, 502, 766, 1331]. wave [2652]. wave-height [2652]. Wavelet [2561, 3270, 2583, 2760, 2262, 2284, 2225, 2590, 2792, 2424, 2401, 2394, 3084, 2836, 3297]. Wavelet-Based [2561, 2583, 3297]. wavelets [2479, 3326, 2831]. waves [1664]. way [2450, 1468, 2579, 2477, 2578, 871, 3395, 2596, 1214, 346, 1347, 1398, 507, 650, 1685, 1273, 1686, 3018, 2637, 1168, 794, 1812, 2935]. weak [2629, 229, 3046, 864, 1869, 936, 1388, 1965, 1602, 3399]. weakly [1115, 1787, 565, 2653]. Web [2226]. Weibull [1574, 1564, 3413, 3012, 1617, 1844, 2520, 2562, 1875, 3079]. Weibull-type [3012, 3079]. weighing [823, 549, 1250, 863]. weight [1874, 1633]. Weighted [2137, 2073, 2301, 2089, 3205, 2272, 3189, 2249, 2759, 1916, 2705, 61, 3119, 2512, 1122, 3000, 2452, 1599, 875, 1843, 3135, 2812, 3035, 318, 402, 1546, 1415, 2688, 2503, 3217, 1952, 2633]. weighted-symmetry [1916]. Weights [2155]. welfare [77]. well [542, 1210]. well-known [542]. whales [1665]. wheat [25]. whether [1325]. which [225, 1900, 1398, 112, 2471, 1042, 468, 1708, 381, 1583, 324, 1567]. While [2047]. Whitney [2358, 2906, 261]. Whittle [3338, 1877]. Whitworth [944]. whole [3184]. Wicksell [2076, 2259]. wide [106]. Width [2016, 2062, 3145, 1705, 1355, 1980, 3407]. Wiener [210, 274, 501, 65]. Wigner [3344]. Wilcoxon [1900, 717, 767, 1430, 2587, 381, 2518]. Wilks [888, 519, 2770, 634, 858, 1031, 1276, 1110, 463]. window [276, 2810]. windows [2683]. wine [3138]. winning [2779]. wise [2727, 2438, 3399]. Wishart [952, 893, 3192, 439, 774, 3275, 676, 860, 1942, 1108, 652, 298, 2744, 1943, 2086, 999, 3344, 635, 666]. Within [2332, 2148]. Within-Individuals [2148]. Without [2193, 2258, 688, 1003, 2807, 1656, 943, 2357, 313, 188, 923]. word [2549, 2790]. Words [2276]. Work [151]. worse [2636]. worse-eye [2636]. worth [1735]. Wrong [2043]. Wu [3140].

X [3157]. X-ray [3157]. XII [249].

Yamamoto [385]. Yates [1002]. years [1266]. Yield [25, 468]. Yokes [1640].

Zelliner [816]. zero [133, 3078, 1520, 566, 2814, 2935]. zero-crossing [1520]. zeroinflated [3078]. zero-one [133, 566, 2814, 2935]. Zipf [2808]. Zipf-type [2808]. zonal [1055, 1135, 1030, 652, 1160, 1161]. zone [1149, 1202]. zur [393]. Zyoiti [570].

References

Kunisawa:1949:AMT

[1] K. Kunisawa. On an analytical method in the theory of independent random variables. Annals of the Institute of Statistical Mathematics, 1(1):1–77, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF03029242.

Matusita:1949:NIC

[2] Kameo Matusita. Note on the independence of certain statistics. Annals of the Institute of Statistical Mathematics, 1(1):79–82, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                              //link.springer.com/article/10.
                                              1007/BF03029243.

Ogawa:1949:IBQ

[3] Junjiro Ogawa. On the independence of bilinear and quadratic forms of a random sample from a normal population. Annals of the Institute of Statistical Mathematics, 1(1):83–108, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF03029244.

Sakamoto:1949:CID

[4] Heihachi Sakamoto. On the criteria of the independence and the degrees of freedom of statistics and their applications to the analysis of variance. Annals of the Institute of Statistical Mathematics, 1(1):109–122, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03029245.

Midzuno:1949:SMU

[5] Hiroshi Midzuno. A survey method using two kinds of surveys. Annals of the Institute of Statistical Mathematics, 1(1):123–124, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF03029246.

Hayashi:1949:FNTa

[6] Chikio Hayashi. Fragments of a new test formula of normality. Annals of the Institute of Statistical Mathematics, 1(1):125–130, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03029247.

Kawata:1949:RFFa

[7] Tatsuo Kawata. Representation of a function by the Fourier–Stieltjes integral. Annals of the Institute of Statistical Mathematics, 1(1):131–139, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03029248.

Matusita:1949:RWTa

[8] Kameo Matusita. A remark to the Wald’s theory of statistical inference. Annals of the Institute of Statistical Mathematics, 1(1):141–148, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03029249.

Midzuno:1949:OTSa

[9] Hiroshi Midzuno. An outline of the theory of sampling systems. Annals of the Institute of Statistical Mathematics, 1(1):149–156, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03029250.

Ugaheri:1949:LDa

[10] Tadashi Ugaheri. On a limit distribution. Annals of the Institute of Statistical Mathematics, 1(1):157–160, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03029251.

Anonymous:1949:HCa

[11] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 1(1):??, March 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayashi:1949:FNTb

[12] Chikio Hayashi. Fragments of a new test formula of normality. Annals of the Institute of Statistical Mathematics, 1(2):125–130, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02974852.

Kawata:1949:RFFb

[13] Tatsuo Kawata. Representation of a function by the Fourier–Stieltjes integral. Annals of the Institute of Statistical Mathematics, 1(2):131– 139, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02974853.

Matusita:1949:RWTb

[14] Kameo Matusita. A remark to the Wald’s theory of statistical inference. Annals of the Institute of Statistical Mathematics, 1(2):141–148, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02974854.

Midzuno:1949:OTSb

[15] Hiroshi Midzuno. An outline of the theory of sampling systems. Annals of the Institute of Statistical Mathematics, 1(2):149–156, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02974855.

Ugaheri:1949:LDb

[16] Tadashi Ugaheri. On a limit distribution. Annals of the Institute of Statistical Mathematics, 1(2): 157–160, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02974856.

Anonymous:1949:HCb

[17] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 1(2):??, September 1949. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Ugaheri:1950:ACL

[18] Tadashi Ugaheri. On the abscissa of convergence of Laplace–Stieltjes integral. Annals of the Institute of Statistical Mathematics, 2(1):1–3, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919495.

Matusita:1950:FOC

[19] Kameo Matusita. On the fundamental operations of collectives. Annals of the Institute of Statistical Mathematics, 2(1):5–11, December 1950– 1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919496.

Nabeya:1950:RBE

[20] Seiji Nabeya. On a relation between exponential law and Poisson’s law. Annals of the Institute of Statistical Mathematics, 2(1):13–16, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919497.

Aoyama:1950:NCO

[21] Hirojiro Aoyama. A note on the classification of observation data. Annals of the Institute of Statistical Mathematics, 2(1):17–19, December 1950– 1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919498.

Midzuno:1950:CGI

[22] Hiroshi Midzuno. On certain groups of inequalities. Annals of the Institute of Statistical Mathematics, 2(1): 21–33, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02919499.

Hayashi:1950:QQD

[23] Chikio Hayashi. On the quantification of qualitative data from the mathematico-statistical point of view. Annals of the Institute of Statistical Mathematics, 2(1):35–47, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02919500.

Hayashi:1950:SDL

[24] Chikio Hayashi, Fumiyuki Maruyama, Masatsugu D. Ishida, Setsuko Takakura, Masako Taguma, and Michiyo Suzuki. Sampling design in literacy survey. Annals of the Institute of Statistical Mathematics, 2(1):49–59, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02919501.

Baker:1950:YTB

[25] G. A. Baker and F. N. Briggs. Yield trials with backcross derived lines of wheat. Annals of the Institute of Statistical Mathematics, 2(1):61– 67, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02919502.

Hayashi:1950:SDS

[26] Chikio Hayashi. Sampling design in the social survey of language at the city of Shirakawa. Annals of the Institute of Statistical Mathematics, 2(1):69–75, December 1950– 1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919503.

Hayashi:1950:SCS

[27] Chikio Hayashi, Fumiyuki Maruyama, and Masatsugu D. Ishida. On some criteria for stratification. Annals of the Institute of Statistical Mathematics, 2(1):77–86, December 1950– 1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919504.

Bennett:1950:NSG

[28] B. M. Bennett. Note on a solution of the generalized Behrens–Fisher problem. Annals of the Institute of Statistical Mathematics, 2(1):87–90, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02919505.

Sato:1950:DT

[29] Ryoichiro Sato. “r” distributions and “r” tests. Annals of the Institute of Statistical Mathematics, 2(1):91– 124, December 1950–1951. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02919506.

Anonymous:1950:HC

[30] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 2(1):??, December 1950–1951.

CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nabeya:1951:AMD

[31] Seiji Nabeya. Absolute moments in 2-dimensional normal distribution. Annals of the Institute of Statistical Mathematics, 3(1):1, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02949770.

Nabeya:1951:NMT

[32] Seiji Nabeya. Note on the moments of the transformed correlation. Annals of the Institute of Statistical Mathematics, 3(1):2–6, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949771.

Takano:1951:CCD

[33] Kinsaku Takano. On the convergence of classes of distributions. Annals of the Institute of Statistical Mathematics, 3(1):7–15, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949772.

Matusita:1951:TSD

[34] Kameo Matusita. On the theory of statistical decision functions. Annals of the Institute of Statistical Mathematics, 3(1):17–35, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

1007/BF02949773. See correction [51].

Cohen:1951:EMV

[35] A. C. Cohen. On estimating the mean and variance of singly truncated normal frequency distributions from the first three sample moments. Annals of the Institute of Statistical Mathematics, 3(1):37–44, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949774.

Sato:1951:TRR

[36] Ryoichiro Sato. The r tests relating to the regression. Annals of the Institute of Statistical Mathematics, 3(1):45–56, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949775.

Aoyama:1951:PSS

[37] Hirojiro Aoyama. On practical systematic sampling. Annals of the Institute of Statistical Mathematics, 3 (1):57–63, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02949776.

Aoyama:1951:MI

[38] Hirojiro Aoyama. On Midzuno’s inequality. Annals of the Institute of Statistical Mathematics, 3(1):65–67, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949777.

Hayashi:1951:PPQ

[39] Chikio Hayashi. On the prediction of phenomena from qualitative data and the quantification of qualitative data from the mathematico-statistical point of view. Annals of the Institute of Statistical Mathematics, 3(1):69–98, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949778.

Midzuno:1951:SSP

[40] Hiroshi Midzuno. On the sampling system with probability proportionate to sum of sizes. Annals of the Institute of Statistical Mathematics, 3(1):99–107, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949779.

Midzuno:1951:RSD

[41] Hiroshi Midzuno. Report of the survey design for agricultural production estimates in the Ryukyu Islands. Annals of the Institute of Statistical Mathematics, 3(1):109–121, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949780.

Goodman:1951:PGD

[42] Leo A. Goodman. On the Poisson– Gamma distribution problem. Annals of the Institute of Statistical Mathematics, 3(1):123–125, ???? 1951– 1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949781.

Sato:1951:E

[43] R. Sato. Errata. Annals of the Institute of Statistical Mathematics, 3(1): 127–128, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02949782.

Anonymous:1951:HC

[44] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 3(1):??, ???? 1951–1952. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kazami:1952:APE

[45] Akiko Kazami. Asymptotic properties of the estimates of an unknown parameter in stationary Markoff process. Annals of the Institute of Statistical Mathematics, 4(1):1–6, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949783.

Isida:1952:RLR

[46] Masatugu D. Isida. A remark on the linear regression estimate. Annals of the Institute of Statistical Mathematics, 4(1):7–9, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949784.

Matusita:1952:NDP

[47] Kameo Matusita and Hirotugu Akaike. Note on the decision problem. Annals of the Institute of Statistical Mathematics, 4(1):11–14, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949785.

Nabeya:1952:AMD

[48] Seiji Nabeya. Absolute moments in 3-dimensional normal distribution. Annals of the Institute of Statistical Mathematics, 4(1):15–30, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949786.

Bennett:1952:EMB

[49] B. M. Bennett. Estimation of means on the basis of preliminary tests of significance. Annals of the Institute of Statistical Mathematics, 4(1):31–43, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949787.

Chapman:1952:TER

[50] Douglas G. Chapman. On tests and estimates for the ratio of Poisson means. Annals of the Institute of Statistical Mathematics, 4(1):45–49, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949788.

Matusita:1952:CPT

[51] Kameo Matusita. Correction to the paper “On the theory of statistical decision functions”. Annals of the Institute of Statistical Mathematics, 4 (1):51–53, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02949789. See [34].

Hayashi:1952:MP

[52] Chikio Hayashi and Hirotugu Akaike. On a matching problem. Annals of the Institute of Statistical Mathematics, 4(1):55–64, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949790.

Inada:1952:CDP

[53] Ken iti Inada. On a certain decision problem under some constraints. Annals of the Institute of Statistical Mathematics, 4(1):65–82, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949791.

Aoyama:1952:TPC

[54] Hirojiro Aoyama. On a test in paired comparisons. Annals of the Institute of Statistical Mathematics, 4(1): 83–87, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02949792.

Nisihira:1952:SAI

[55] Sigeki Nisihira. Some analysis on the intensity. Annals of the Institute of Statistical Mathematics, 4(1):89–94, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949793.

Taga:1952:OBB

[56] Yasushi Taga. On optimum balancing between sample size and number of strata in sub-sampling. Annals of the Institute of Statistical Mathematics, 4(1):95–102, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949794.

Konijn:1952:RCM

[57] H. S. Konijn. A remark on the characterization of minimax procedures. Annals of the Institute of Statistical Mathematics, 4(1):103–105, ???? 1952– 1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949795.

Anonymous:1952:HC

[58] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 4(1):??, ???? 1952–1953. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takano:1953:MCC

[59] Kinsaku Takano. A metrization of class-convergences of distributions. Annals of the Institute of Statistical Mathematics, 5(1):1–7, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949796.

Wolfowitz:1953:EMD

[60] J. Wolfowitz. Estimation by the minimum distance method. Annals of the Institute of Statistical Mathematics, 5(1):9–23, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949797.

Aoyama:1953:CST

[61] Hirojiro Aoyama. On the chi-square test for weighted samples. Annals of the Institute of Statistical Mathematics, 5(1):25–28, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949798.

Nisihira:1953:QSS

[62] Sigelri Nisihira. A quantification of social status. Annals of the Institute of Statistical Mathematics, 5(1): 29–40, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02949799.

Takano:1953:MDD

[63] Kinsakn Takano. On the manydimensional distribution functions. Annals of the Institute of Statistical Mathematics, 5(1):41–58, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949800.

Matusita:1953:EMD

[64] Kameo Matusita. On the estimation by the minimum distance method. Annals of the Institute of Statistical Mathematics, 5(1):59–65, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949801.

Takano:1953:NWP

[65] Kinsaku Takano. Note on Wiener’s prediction theory. Annals of the Institute of Statistical Mathematics, 5 (1):67–72, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02949802.

Aoyama:1953:IB

[66] Hirojiro Aoyama. On the interviewing bias. Annals of the Institute of Statistical Mathematics, 5(1):73–76, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949803.

Higuchi:1953:SCS

[67] Isao Higuchi. On the solutions of certain simultaneous equations in the theory of systematic statistics. Annals of the Institute of Statistical Mathematics, 5(1):77–90, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949804.

Baker:1953:ESL

[68] G. A. Baker. The effect of selection on linear functions of normally distributed correlated variables on the distributions of other linear functions. Annals of the Institute of Statistical Mathematics, 5(1):91–95, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02949805.

Hiraga:1953:TTS

[69] Yoshihiko Hiraga, Hidenori Morimura, and Hisao Watanabe. Tables for threesample test. Annals of the Institute of Statistical Mathematics, 5(1):

97–102, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02949806.

Bennett:1953:SFE

[70] B. M. Bennett. Some further extensions of Fieller’s theorem. Annals of the Institute of Statistical Mathematics, 5(1):103–106, ???? 1953– 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949807.

Yokota:1953:PS

[71] Toshio Yokota. On the polaron state. Annals of the Institute of Statistical Mathematics, 5(1):107–119, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02949808.

Hayashi:1953:MQ

[72] Chikio Hayashi. Multidimensional quantification. Annals of the Institute of Statistical Mathematics, 5(1): 121–143, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02949809.

Hayashi:1953:E

[73] Chikio Hayashi and H. S. Konijn. Errata. Annals of the Institute of Statistical Mathematics, 5(1):144, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02949810.

Anonymous:1953:HC

[74] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 5(1):??, ???? 1953–1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aoyama:1954:SSR

[75] Hirojiro Aoyama. A study of the stratified random sampling. Annals of the Institute of Statistical Mathematics, 6 (1):1–36, December 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02960514.

Takano:1954:SLT

[76] Kinsaku Takano. On some limit theorems of probability distributions. Annals of the Institute of Statistical Mathematics, 6(1):37–113, December 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02960515.

Inada:1954:EPS

[77] Ken ichi Inada. Elementary proofs of some theorems about the social welfare function. Annals of the Institute of Statistical Mathematics, 6(1):115–122, December 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02960516.

Konijn:1954:FRC

[78] H. S. Konijn. A further remark on the characterization of minimax procedures. Annals of the Institute of Statistical Mathematics, 6(1):123, December 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02960517.

Anonymous:1954:HCa

[79] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 6(1):??, December 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akaike:1954:ADF

[80] Hirotugu Akaike. An approximation to the density function. Annals of the Institute of Statistical Mathematics, 6(2):127–132, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02900741.

Matusita:1954:TSH

[81] Kameo Matusita, Yukio Suzuki, and Hirosi Hudimoto. On testing statistical hypotheses. Annals of the Institute of Statistical Mathematics, 6(2):133–141, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02900742.

Matusita:1954:DRP

[82] Kameo Matusita. Decision rule by probability ratio. Annals of the Institute of Statistical Mathematics, 6 (2):143–151, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02900743.

Siotani:1954:ESD

[83] Minoru Siotani. An estimate of standard deviation of normal population based on the difference between means of two groups divided by sample mean. Annals of the Institute of Statistical Mathematics, 6(2):153–160, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02900744.

Higuti:1954:SRC

[84] Isao Higuti. A statistical research on colloidal graphite. I. Annals of the Institute of Statistical Mathematics, 6(2):161–172, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02900745.

Sakino:1954:FPP

[85] Sigeki Sakino and Goro Kono. On the forecasting of prognosis in pediatrics by a quantifying method. Annals of the Institute of Statistical Mathematics, 6(2):173–178, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02900746.

Aoyama:1954:Ea

[86] Hirojiro Aoyama. Errata. Annals of the Institute of Statistical Mathematics, 6(2):179, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02900747.

Aoyama:1954:Eb

[87] Hirojiro Aoyama. Errata. Annals of the Institute of Statistical Mathematics, 6(2):179, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02900748.

Aoyama:1954:Ec

[88] Hirojiro Aoyama. Errata. Annals of the Institute of Statistical Mathematics, 6(2):179–180, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02900749.

Anonymous:1954:HCb

[89] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 6(2):??, June 1954. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mine:1954:ELR

[90] Akiko Mine. Estimation of linear regression coefficients in time series. Annals of the Institute of Statistical Mathematics, 6(3):181–189, ???? 1954– 1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02905912.

Motoo:1954:NRB

[91] Minoru Motoo. Note on a relation between the distribution functions and characteristic functions. Annals of the Institute of Statistical Mathematics, 6(3):191–195, ???? 1954– 1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02905913.

Suzuki:1954:NNP

[92] Yukio Suzuki. Note on the Neyman– Pearson’s fundamental lemma. Annals of the Institute of Statistical Mathematics, 6(3):197–211, ???? 1954– 1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02905914.

Walsh:1954:BSL

[93] John E. Walsh. Bounded significance level tests for comparing quantiles of two possibly different continuous populations. Annals of the Institute of Statistical Mathematics, 6(3): 213–222, ???? 1954–1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02905915.

Nikaido:1954:NAN

[94] Hukukane Nikaidô. New aspects of von Neumann’s model with special regard to computational problems. Annals of the Institute of Statistical Mathematics, 6(3):223–230, ???? 1954– 1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02905916.

Sakino:1954:SID

[95] Sigeki Sakino and Umeji Hirata. On the statistical investigation of diagnosis in the internal medicine. Annals of the Institute of Statistical Mathematics, 6(3):231–235, ???? 1954– 1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02905917.

Bennett:1954:E

[96] B. M. Bennett. Errata. Annals of the Institute of Statistical Mathematics, 6(3):237, ???? 1954–1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02905918.

Anonymous:1954:HCc

[97] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 6(3):??, ???? 1954–1955. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matusita:1955:SPS

[98] Kameo Matusita, Chikio Hayashi, Masatugu Isida, Hirobumi Uzawa, Hirosi Hudimoto, Hirotugu Akaike, and Tosio Uematu. Some problems of sampling in the forest survey. Annals of the Institute of Statistical Mathematics, 7(1):1–23, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02951442.

Yamamoto:1955:TSP

[99] Sumiyasu Yamamoto. On the theory of sampling with probabilities proportionate to given values. Annals of the Institute of Statistical Mathematics, 7(1):25–38, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02951443.

Siotani:1955:SDV

[100] Minoru Siotani. The significance of the discordant variance estimates. Annals of the Institute of Statistical Mathematics, 7(1):39–55, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951444.

Bennett:1955:NML

[101] B. M. Bennett. Note on the moments of the logarithmic non-central χ^{2} and z distributions. Annals of the Institute of Statistical Mathematics, 7(1):57–61, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951445.

Bennett:1955:JDM

[102] B. M. Bennett. On the joint distribution of the mean and standard deviation. Annals of the Institute of Statistical Mathematics, 7(1):63–66, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951446.

Matusita:1955:DRB

[103] Kameo Matusita and Hirotugu Akaike. Decision rules, based on the distance, for the problems of independence, invariance and two samples. Annals of the Institute of Statistical Mathematics, 7(1):67–80, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951447.

Takano:1955:MCL

[104] Kinsaku Takano. Multidimensional central limit criterion in the case of bounded variances. Annals of the Institute of Statistical Mathematics, 7(1):81–93, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951448.

Takano:1955:CCC

[105] Kinsaku Takano. Central convergence criterion in the multidimensional case. Annals of the Institute of Statistical Mathematics, 7(1):95–102, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951449.

Baker:1955:EWG

[106] G. A. Baker. The effects of wide groupings on the distributions of array means and variances for correlated normal variables. Annals of the Institute of Statistical Mathematics, 7(1):103–106, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951450.

Akaike:1955:MCM

[107] Hirotugu Akaike. Monte Carlo method applied to the solution of simultaneous linear equations. Annals of the Institute of Statistical Mathematics, 7(1):107–113, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951451.

Kimura:1955:AMN

[108] Hitosi Kimura. An approximation method in numerical computation of the Leontief’s open input-output model. Annals of the Institute of Statistical Mathematics, 7(1):115–122, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951452.

Uzawa:1955:GLC

[109] Hirofumi Uzawa. A generalization of Laplace criterion for decision problems. Annals of the Institute of Statistical Mathematics, 7(1):123–129, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951453.

Muta:1955:SSD

[110] Akinori Muta and Isao Higuti. Shape and size distribution of carbon black when it is crushed. Annals of the Institute of Statistical Mathematics, 7(1):131–135, ???? 1955– 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951454.

Matusita:1955:FTD

[111] Kameo Matusita and Minoru Motoo. On the fundamental theorem for the decision rule based on distance || ||. Annals of the Institute of Statistical Mathematics, 7(1):137–142, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951455.

Konijn:1955:SEW

[112] H. S. Konijn. Some estimates which minimize the least upper bound of a probability together with the cost of observation. Annals of the Institute of Statistical Mathematics, 7(1): 143–158, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02951456.

Hudimoto:1955:NFS

[113] Hirosi Hudimoto. Note on fitting a straight line when both variables are subject to error and some applications. Annals of the Institute of Statistical Mathematics, 7(1):159–167, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02951457.

Motoo:1955:STS

[114] Minoru Motoo. Some theorems on the sum of positive random variables. Annals of the Institute of Statistical Mathematics, 7(1):169–181, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02951458.

Akaike:1955:OCN

[115] Hirotugu Akaike. On optimum character of von Neumann’s Monte Carlo model. Annals of the Institute of Statistical Mathematics, 7(1):183–193, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02951459.

Uzawa:1955:IEC

[116] Hirofumi Uzawa. On intertemporal efficiency conditions of capital accumulation (i). Annals of the Institute of Statistical Mathematics, 7(1): 195–204, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02951460.

Kimura:1955:EOD

[117] Hitosi Kimura. On the errors of outputs due to errors of technical coefficients in Leontief’s open input-output models. Annals of the Institute of Statistical Mathematics, 7(1):205–213, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02951461.

Sakino:1955:DTS

[118] Sigeki Sakino. Determination of the target size by the indirect action of irradiation. Annals of the Institute of Statistical Mathematics, 7(1): 215–220, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02951462.

Sakino:1955:E

[119] Sigeki Sakino, Umeji Hirata, Kameo Matusita, and Hirotugu Akaike. Errata. Annals of the Institute of Statistical Mathematics, 7(1):221, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02951463.

Anonymous:1955:HC

[120] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 7(1):??, ???? 1955–1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Siotani:1956:DHS

[121] Minoru Siotani. On the distributions of the Hotelling’s T^{2}-statistics. Annals of the Institute of Statistical Mathematics, 8(1):1–14, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863561.

Epstein:1956:SEP

[122] Benjamin Epstein. Simple estimators of the parameters of exponential distributions when samples are censored. Annals of the Institute of Statistical Mathematics, 8(1):15–26, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863562.

Aoyama:1956:ESE

[123] Hirojiro Aoyama. On the evaluation of the sampling error of a certain determinant. Annals of the Institute of Statistical Mathematics, 8(1):27–33, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863563.

Uzawa:1956:NPA

[124] Hirofumi Uzawa. Note on preference and axioms of choice. Annals of the Institute of Statistical Mathematics, 8(1):35–40, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863564.

Bennett:1956:CLR

[125] B. M. Bennett. On confidence limits for the ratio of regression coefficients. Annals of the Institute of Statistical Mathematics, 8(1):41–43, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863565.

Bennett:1956:UPT

[126] B. M. Bennett. On the use of preliminary tests in certain statistical procedures. Annals of the Institute of Statistical Mathematics, 8(1):45–52, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863566.

Akaike:1956:DPT

[127] Hirotugu Akaike. On the distribution of the product of two Γ-distributed variables. Annals of the Institute of Statistical Mathematics, 8(1):53–54, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02863567.

Higuti:1956:NSI

[128] Isao Higuti. Note on the sums of the independent variates of K. Pearson’s type V. Annals of the Institute of Statistical Mathematics, 8(1):55–59, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863568.

Suzuki:1956:NOM

[129] Yukio Suzuki. Note on optimal machine setting. Annals of the Institute of Statistical Mathematics, 8(1):61–64, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863569.

Kimura:1956:E

[130] Hitosi Kimura. Errata. Annals of the Institute of Statistical Mathematics, 8(1):65, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02863570.

Matusita:1956:DRB

[131] Kameo Matusita. Decision rule, based on the distance, for the classification problem. Annals of the Institute of Statistical Mathematics, 8(1):67–77, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863571.

Walsh:1956:VAN

[132] John E. Walsh. Validity of approximate normality values for μ ± kσ areas of practical type continuous populations. Annals of the Institute of Statistical Mathematics, 8(1):79–86, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863572.

Akaike:1956:ZOP

[133] Hirotugu Akaike. On a zero-one process and some of its applications. Annals of the Institute of Statistical Mathematics, 8(1):87–94, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863573.

Siotani:1956:OSD

[134] Minoru Siotani. Order statistics for discrete case with a numerical application to the binomial distribution. Annals of the Institute of Statistical Mathematics, 8(1):95–104, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863574.

Hudimnoto:1956:DFC

[135] Hirosi Hudimnoto. On the distributionfree classification of an individual into one of two groups. Annals of the Institute of Statistical Mathematics, 8 (1):105–112, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02863575.

Higuti:1956:NDO

[136] Isao Higuti. Note on the decision of optimal tolerance in the design of a simple random assembling. Annals of the Institute of Statistical Mathematics, 8(1):113–118, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863576.

Isida:1956:RNG

[137] Masatugu Isida and Hiroji Ikeda. Random number generator. Annals of the Institute of Statistical Mathematics, 8(1):119–126, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863577.

Matusita:1956:E

[138] Kameo Matusita, Hitosi Kimura, H. S. Konijn, and Isao Higuti. Errata. Annals of the Institute of Statistical Mathematics, 8(1):127, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02863578.

Aoyama:1956:SFT

[139] Hirojiro Aoyama. Sampling fluctuations of the test reliability. Annals of the Institute of Statistical Mathematics, 8(1):129–143, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863579.

Motoo:1956:HCC

[140] Minoru Motoo. On the Hoeffding’s combinatrial central limit theorem. Annals of the Institute of Statistical Mathematics, 8(1):145–154, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02863580.

Sibuya:1956:TPD

[141] Masaaki Sibuya and Hideo Toda. Tables of the probability density function of range in normal samples. Annals of the Institute of Statistical Mathematics, 8(1):155–165, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863581.

Walsh:1956:NME

[142] John E. Walsh. Nonparametric mean estimation of percentage points and density function values. Annals of the Institute of Statistical Mathematics, 8 (1):167–180, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02863582.

Nisihira:1956:CNC

[143] Sigeki Nisihira. Cross-national comparative study on social stratification and social mobility. Annals of the Institute of Statistical Mathematics, 8(1): 181–191, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02863583.

Bennett:1956:TLR

[144] B. M. Bennett. Tests for linearity of regression involving correlated observations. Annals of the Institute of Statistical Mathematics, 8(1):193–195, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02863584.

Higuti:1956:E

[145] Isao Higuti. Errata. Annals of the Institute of Statistical Mathematics, 8(1):195, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02863585.

Anonymous:1956:HC

[146] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 8(1):??, December 1956. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Isii:1957:SIR

[147] Keiiti Isii. Some investigations of the relation between distribution functions and their moments. Annals of the Institute of Statistical Mathematics, 9 (1):1–11, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02892487.

Akaike:1957:EPT

[148] Hirotugu Akaike. On ergodic property of a tandem type queueing process. Annals of the Institute of Statistical Mathematics, 9(1):13–21, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892488.

Aoyama:1957:CSS

[149] Hirojiro Aoyama. On a certain statistic in a social group. Annals of the Institute of Statistical Mathematics, 9(1):23–30, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892489.

Hudimoto:1957:NPC

[150] Hirosi Hudimoto. A note on the probability of the correct classification when the distributions are not specified. Annals of the Institute of Statistical Mathematics, 9(1):31–36, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892490.

Suzuki:1957:SOW

[151] Yukio Suzuki. On scheduling of overtime work. Annals of the Institute of Statistical Mathematics, 9 (1):37–42, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892491.

Gold:1957:GPD

[152] Louis Gold. Generalized Poisson distributions. Annals of the Institute of Statistical Mathematics, 9(1):43–47, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892492.

Hayashi:1957:NSS

[153] Chikio Hayashi. Note on sampling from a sociometric pattern. Annals of the Institute of Statistical Mathematics, 9(1):49–52, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892493.

Takano:1957:BTI

[154] Kinsaku Takano. On the basic theorems of information theory. Annals of the Institute of Statistical Mathematics, 9(1):53–77, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892494.

Teicher:1957:CPD

[155] Henry Teicher. ”on the convergence of projected distributions”. Annals of the Institute of Statistical Mathematics, 9(1):79–86, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892495.

Uematu:1957:TCI

[156] Tosio Uematu. On the traffic control at an intersection controlled by the repeated fixed-cycle traffic lights. Annals of the Institute of Statistical Mathematics, 9(1):87–107, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892496.

Maly:1957:CTM

[157] Vladimír Malý. The comparing of two methods in microbiology. Annals of the Institute of Statistical Mathematics, 9(1):109–115, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892497.

Ishii:1957:TFL

[158] Goro Ishii. Test of fit in life test. Annals of the Institute of Statistical Mathematics, 9(1):117–125, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892498.

Walsh:1957:FCN

[159] John E. Walsh. Further consideration of normality values for μ ± kδ areas of continuous populations. Annals of the Institute of Statistical Mathematics, 9(1):127–129, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892499.

Akaike:1957:E

[160] Hirotugu Akaike. Errata. Annals of the Institute of Statistical Mathematics, 9(1):130, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02892500.

Suzuki:1957:DDP

[161] Yukio Suzuki. Discrete decision problems. Annals of the Institute of Statistical Mathematics, 9(1):131–148, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892501.

Taga:1957:EAL

[162] Yasushi Taga and Tatsuzo Suzuki. On the estimation of average length of chains in the communication-pattern.

Annals of the Institute of Statistical Mathematics, 9(1):149–156, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892502.

Siotani:1957:NUG

[163] Minoru Siotani. Note on the utilization of the generalized Student ratio in the analysis of variance or dispersion. Annals of the Institute of Statistical Mathematics, 9(1):157–171, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892503.

Isii:1957:NCU

[164] Keiiti Isii. Note on a characterization of unimodal distributions. Annals of the Institute of Statistical Mathematics, 9(1):173–184, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892504.

Walsh:1957:ESS

[165] John E. Walsh. Efficient small sample nonparametric median tests with bounded significance levels. Annals of the Institute of Statistical Mathematics, 9(1):185–199, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892505.

Kimura:1957:RPA

[166] Hitisi Kimura. A remark on price analysis in Leontief’s open input-output model. Annals of the Institute of Statistical Mathematics, 9(1):201–213, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892506.

Rider:1957:GCD

[167] Paul R. Rider. Generalized Cauchy distributions. Annals of the Institute of Statistical Mathematics, 9(1): 215–223, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892507.

Sibuya:1957:MIC

[168] Masaaki Sibuya. Modal intervals for chi-square distributions. Annals of the Institute of Statistical Mathematics, 9(1):225–236, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892508.

Ishii:1957:E

[169] Goro Ishii. Errata. Annals of the Institute of Statistical Mathematics, 9(1):236, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02892509.

Anonymous:1957:HC

[170] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 9(1):??, December 1957. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matusita:1959:KT

[171] Kameo Matusita. Kinsaku Takano 1915–1958. Annals of the Institute of Statistical Mathematics, 10(1):i–ii, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02883982.

Akaike:1959:CME

[172] Hirotugu Akaike. On a computation method for eigenvalue problems and its application to statistical analysis. Annals of the Institute of Statistical Mathematics, 10(1):1–20, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02883983.

Motoo:1959:PLI

[173] Minoru Motoo. Proof of the law of iterated logarithm through diffusion equation. Annals of the Institute of Statistical Mathematics, 10(1):21–28, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02883984.

Okamoto:1959:SIR

[174] Masashi Okamoto. Some inequalities relating to the partial sum of binomial probabilities. Annals of the Institute of Statistical Mathematics, 10(1):29–35, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02883985.

Ishii:1959:KST

[175] Goro Ishii. Kolmogorov–Smirnov test in life test. Annals of the Institute of Statistical Mathematics, 10(1):37–46, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883986.

Siotani:1959:TTH

[176] Minoru Siotani and Masaru Ozawa. Tables for testing the homogeneity of k independent binomial experiments on a certain event based on the range. Annals of the Institute of Statistical Mathematics, 10(1):47–63, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883987.

Anonymous:1959:HCa

[177] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 10(1):??, March 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Isii:1959:MGT

[178] Keiiti Isii. On a method for generalizations of Tchebycheff’s inequality. Annals of the Institute of Statistical Mathematics, 10(2):65–88, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883863.

Suzuki:1959:NLP

[179] Yukio Suzuki. Note on linear programming. Annals of the Institute of Statistical Mathematics, 10(2):89–105, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883864.

David:1959:AIV

[180] Herbert T. David, Edward A. Fay, and John E. Walsh. Acceptance inspection by variables when the measurements are subject to error. Annals of the Institute of Statistical Mathematics, 10(2):107–129, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883865.

Uematu:1959:NNC

[181] Tosio Uematu. Note on the numerical computation in the discrimination problem. Annals of the Institute of Statistical Mathematics, 10(2):131–135, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883866.

Yokota:1959:SMS

[182] Toshio Yokota. Stochastic methods of solving partial integro-differential equations and their application to non-stationary Markoff process: I. Annals of the Institute of Statistical Mathematics, 10(2):137–161, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883867.

Aoyama:1959:ERI

[183] Hirojiro Aoyama. On the evaluation of the risk index of the railroad crossing. Annals of the Institute of Statistical Mathematics, 10(2):163–180, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883868.

Kimura:1959:E

[184] Hitisi Kimura. Errata. Annals of the Institute of Statistical Mathematics, 10(2):181, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02883869.

Okamoto:1959:E

[185] Masashi Okamoto. Errata. Annals of the Institute of Statistical Mathematics, 10(2):181, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02883870.

Anonymous:1959:HCb

[186] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 10(2):??, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Siotani:1959:EVG

[187] Minoru Siotani. The extreme value of the generalized distances of the individual points in the multivariate normal sample. Annals of the Institute of Statistical Mathematics, 10(3): 183–208, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883941.

Sibuya:1959:OPC

[188] Masaaki Sibuya and Toshiro Haga. Orthogonal polynomials without constant term. Annals of the Institute of Statistical Mathematics, 10(3): 209–222, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883942.

Walsh:1959:LSN

[189] John E. Walsh. Large sample nonparametric rejection of outlying observations. Annals of the Institute of Statistical Mathematics, 10(3): 223–232, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883943.

Akaike:1959:SCG

[190] Hirotugu Akaike. On the statistical control of the gap process. Annals of the Institute of Statistical Mathematics, 10(3):233–259, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02883944.

Sakino:1959:AEM

[191] Sigeki Sakino and Chikio Hayashi. On the analysis of epidemic model I (theoretical approach). Annals of the Institute of Statistical Mathematics, 10 (3):261–275, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883945.

Dwass:1959:MCP

[192] Meyer Dwass. Multiple confidence procedures. Annals of the Institute of Statistical Mathematics, 10(3): 277–282, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883946.

Odaka:1959:SFR

[193] Kunio Odaka and Sigeki Nisihira. Some factors related to social mobility in Japan. Annals of the Institute of Statistical Mathematics, 10(3): 283–288, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02883947.

Anonymous:1959:HCc

[194] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 10(3):??, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akaike:1959:STP

[195] Hirotugu Akaike. On a successive transformation of probability distribution and its application to the analysis of the optimum gradient method. Annals of the Institute of Statistical Mathematics, 11(1):1–16, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01831719.

Ishii:1959:EPR

[196] Goro Ishii. On the exact probabilities of Renyi’s tests. Annals of the Institute of Statistical Mathematics, 11(1):17–24, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01831720.

Taga:1959:SMC

[197] Yasushi Taga and Keiiti Isii. On a stochastic model concerning the pattern of communication. Annals of the Institute of Statistical Mathematics, 11(1):25–43, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01831721.

Jilek:1959:CDO

[198] Milos Jílek and Otakar Líkar. Coefficients for the determination of onesided tolerance limits of normal distribution. Annals of the Institute of Statistical Mathematics, 11(1):45–48, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01831722.

Motoo:1959:SEC

[199] Minoru Motoo. Some evaluations for continuous Monte Carlo method by using Brownian hitting process. Annals of the Institute of Statistical Mathematics, 11(1):49–54, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01831723.

Aggarwal:1959:TCD

[200] Om P. Aggarwal and Irwin Guttman. Tables of the cumulative distribution functions of samples from symmetrically truncated normal distributions. Annals of the Institute of Statistical Mathematics, 11(1):55–68, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01831724.

Savage:1959:N

[201] I. Richard Savage. Notice. Annals of the Institute of Statistical Mathematics, 11(1):69, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01831725.

Walsh:1959:E

[202] John E. Walsh, Hirotugu Akaike, Sigeki Sakino, and Chikio Hayashi. Errata. Annals of the Institute of Statistical Mathematics, 11(1):70, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01831726.

Suzuki:1959:SIP

[203] Yukio Suzuki. On sampling inspection plans. Annals of the Institute of Statistical Mathematics, 11 (1):71–79, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01737394.

Walsh:1959:NPS

[204] John E. Walsh. Nonparametric properties of some maximum likelihood estimates for median of symmetrical population. Annals of the Institute of Statistical Mathematics, 11 (1):81–88, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01737395.

Isii:1959:BPN

[205] Keiiti Isii. Bounds on probability for non-negative random variables. Annals of the Institute of Statistical Mathematics, 11(1):89–99, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01737396.

Haight:1959:GPD

[206] Frank A. Haight. The generalized Poisson distribution. Annals of the Institute of Statistical Mathematics, 11(1): 101–105, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01737397.

Okamoto:1959:CTD

[207] Masashi Okamoto. A convergence theorem for discrete probability distributions. Annals of the Institute of Statistical Mathematics, 11(1): 107–112, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01737398.

Hudimoto:1959:TSN

[208] Hirosi Hudimoto. On a two-sample non-parametric test in the case that ties are present. Annals of the Institute of Statistical Mathematics, 11(1): 113–120, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01737399.

Ikeda:1959:NNA

[209] Sadao Ikeda. A note on the normal approximation to the sum of independent random variables. Annals of the Institute of Statistical Mathematics, 11(1):121–130, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01737400.

Ikeda:1959:CCS

[210] Sadao Ikeda. Continuity and characterization of Shannon–Wiener information measure for continuous probability distributions. Annals of the Institute of Statistical Mathematics, 11 (1):131–144, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01737401.

Anonymous:1959:HCd

[211] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 11(1):??, ???? 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Anonymous:1959:HCe

[212] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 11(1):??, September 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akaike:1959:ETE

[213] Hirotugu Akaike. Effect of timing-error on the power spectrum of sampleddata. Annals of the Institute of Statistical Mathematics, 11(2):145–165, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01682325.

Siotani:1959:NMC

[214] Minoru Siotani. Notes on multivariate confidence bounds. Annals of the Institute of Statistical Mathematics, 11(2):167–182, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01682326.

Walsh:1959:NTM

[215] John E. Walsh. Nonparametric tests for median by interpolation from sign tests. Annals of the Institute of Statistical Mathematics, 11(2):183–188, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01682327.

Cohen:1959:EPD

[216] A. Clifford Cohen, Jr. Estimation in the Poisson distribution when sample values of c + 1 are sometimes erroneously reported as c. Annals of the Institute of Statistical Mathematics, 11(2):189–193, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01682328.

Sibuya:1959:BES

[217] Masaaki Sibuya. Bivariate extreme statistics, I. Annals of the Institute of Statistical Mathematics, 11(2):195–210, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01682329.

Haga:1959:TSR

[218] Toshiro Haga. A two-sample rank test on location. Annals of the Institute of Statistical Mathematics, 11(2):211–219, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01682330.

Siotani:1959:E

[219] Minoru Siotani, Masaru Ozawa, and Masashi Okamoto. Errata. Annals of the Institute of Statistical Mathematics, 11(2):220, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01682331.

Anonymous:1959:HCf

[220] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 11(2):??, June 1959. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayashi:1960:SJN

[221] C. Hayashi, H. Aoyama, M. Isida, S. Nisihira, Y. Taga, M. Tutumi, H. Akaike, T. Uematu, T. Taguti, T. Suzuki, and K. Ôisi. A study of Japanese national character. Annals of the Institute of Statistical Mathematics, 11(3):1–30, ???? 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01907886.

Nisihira:1960:QCI

[222] Sigeki Nisihira. Quelques comparaisons internationales des attitudes sociales. (French) [Some international comparisons of social attitudes]. Annals of the Institute of Statistical Mathematics, 11(3):31–38, ???? 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01907887.

Anonymous:1960:HCa

[223] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 11(3):??, ???? 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akaike:1960:MMT

[224] Hirotugu Akaike and Yaeko Saigusa. On a min-max theorem and some of its applications. Annals of the Institute of Statistical Mathematics, 12 (1):1–5, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01577660.

Akaike:1960:LPW

[225] Hirotugu Akaike. On a limiting process which asymptotically produces f^{−2} spectral density. Annals of the Institute of Statistical Mathematics, 12(1):7–11, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01577661.

Crawford:1960:EEE

[226] J. R. Crawford and John E. Walsh. Empirical examination of Edgeworth series. Annals of the Institute of Statistical Mathematics, 12(1):13–26, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01577662.

Hudimoto:1960:CUO

[227] Hirosi Hudimoto. On a coefficient of unidimensional ordering for the individuals’ attitudes. Annals of the Institute of Statistical Mathematics, 12 (1):27–35, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01577663.

Motoo:1960:DPC

[228] Minoru Motoo. Diffusion process corresponding to \frac{1}{2} ∑ \frac{∂^{2}}{∂x^{i2}} + ∑ b^{i}(x)\frac{∂}{∂x^{i}}. Annals of the Institute of Statistical Mathematics, 12(1):37–61, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01577664.

Higuti:1960:RRF

[229] Zyunsirô Higuti. Remarque sur la répartition faible dans un espace localement convexe (I). (French) [Remark on the weak repartition in a locally-convex space (I)]. Annals of the Institute of Statistical Mathematics, 12(1):63–67, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01577665.

Ishii:1960:CBD

[230] Goro Ishii and Reiko Hayakawa. On the compound binomial distribution. Annals of the Institute of Statistical Mathematics, 12(1):69–80, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01577666.

Ikeda:1960:RCK

[231] Sadao Ikeda. A remark on the convergence of Kullback–Leibler’s mean information. Annals of the Institute of Statistical Mathematics, 12 (1):81–88, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01577667.

Anonymous:1960:E

[232] Anonymous. Errata. Annals of the Institute of Statistical Mathematics, 12(1):89, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01577668.

Anonymous:1960:HCb

[233] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 12(1):??, February 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Barankin:1960:SPS

[234] Edward W. Barankin. Sufficient parameters: Solution of the minimal dimensionality problem. Annals of the Institute of Statistical Mathematics, 12(2):91–118, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01733119.

Isii:1960:EPD

[235] Keiiti Isii. The extrema of probability determined by generalized moments (I) bounded random variables. Annals of the Institute of Statistical Mathematics, 12(2):119–134, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01733120.

Ogawa:1960:DOS

[236] Junjiro Ogawa. Determination of optimum spacings for the estimation of the scale parameter of an exponential distribution based on sample quantiles. Annals of the Institute of Statistical Mathematics, 12(2):135–141, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01733121.

Walsh:1960:PCM

[237] John E. Walsh. Probabilities for Cramér-von Mises–Smirnov test using grouped data. Annals of the Institute of Statistical Mathematics, 12(2):143–145, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01733122.

Siotani:1960:NIE

[238] Minoru Siotani. A note on the interval estimation related to the regression matrix. Annals of the Institute of Statistical Mathematics, 12(2):147–149, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01733123.

Sibuya:1960:CPM

[239] Masaaki Sibuya. Cutting out procedures for material with Poisson defects. Annals of the Institute of Statistical Mathematics, 12(2):151–159, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01733124.

Ishii:1960:ICT

[240] Goro Ishii. Intraclass contingency tables. Annals of the Institute of Statistical Mathematics, 12(2):161–207, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01733125. See

corrections [249].

Ishii:1960:E

[241] Goro Ishii and Reiko Hayakawa. Errata. Annals of the Institute of Statistical Mathematics, 12(2):208, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01733126.

Anonymous:1960:HCc

[242] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 12(2):??, June 1960. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Clark:1961:ECS

[243] Charles E. Clark and G. Trevor Williams. Estimates from censored samples. Annals of the Institute of Statistical Mathematics, 12(3):209–226, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728931.

Birnbaum:1961:MDP

[244] Allan Birnbaum. A multi-decision procedure related to the analysis of single degrees of freedom. Annals of the Institute of Statistical Mathematics, 12(3):227–236, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728932.

Takahasi:1961:MES

[245] Koiti Takahasi. Model for the estimation of the size of a population by using capture-recapture method. Annals of the Institute of Statistical Mathematics, 12(3):237–248, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728933.

Gumbel:1961:RPO

[246] E. J. Gumbel. The return period of order statistics. Annals of the Institute of Statistical Mathematics, 12(3):249–256, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728934.

Higuti:1961:SSR

[247] Isao Higuti. A statistical study of random packing of unequal spheres. Annals of the Institute of Statistical Mathematics, 12(3):257–271, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728935.

Ishii:1961:NTH

[248] Goro Ishii and Mitsuru Yamasaki. A note on the testing of homogeneity of k binomial experiments based on the range. Annals of the Institute of Statistical Mathematics, 12(3):273–278, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01728936.

Ishii:1961:CIC

[249] Goro Ishii. Corrections to “Intraclass contingency tables.” The same Annals Vol., XII, No. 2. Annals of the Institute of Statistical Mathematics, 12(3):279, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01728937. See

[240].

Isii:1961:E

[250] Keiiti Isii. Errata. Annals of the Institute of Statistical Mathematics, 12(3):280, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01728938.

Anonymous:1961:HCa

[251] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 12(3):??, October 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hoel:1961:SPO

[252] Paul G. Hoel. Some properties of optimal spacing in polynomial estimation. Annals of the Institute of Statistical Mathematics, 13(1):1–8, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868855.

Guttman:1961:BPT

[253] Irwin Guttman. Best populations and tolerance regions. Annals of the Institute of Statistical Mathematics, 13 (1):9–26, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868856.

Ikeda:1961:ADIa

[254] Sadao Ikeda. An application of the discrimination information measure to the theory of testing hypotheses. Part I. Annals of the Institute of Statistical Mathematics, 13(1):27–46, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868857.

Khatri:1961:DOV

[255] C. G. Khatri. On the distributions obtained by varying the number of trials in a binomial distribution. Annals of the Institute of Statistical Mathematics, 13(1):47–51, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868858.

Shimizu:1961:CND

[256] Ryoichi Shimizu. A characterization of the normal distribution. Annals of the Institute of Statistical Mathematics, 13(1):53–56, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868859.

Rao:1961:EVU

[257] J. N. K. Rao. On the estimate of the variance in unequal probability sampling. Annals of the Institute of Statistical Mathematics, 13(1):57–60, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868860.

Ikeda:1961:ADIb

[258] Sadao Ikeda. An application of the discrimination information measure to the theory of testing hypotheses. Part II. Annals of the Institute of Statistical Mathematics, 13(1):61–89, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868861.

Whittlesey:1961:CDE

[259] John R. B. Whittlesey and Frank A. Haight. Counting distributions for an Erlang process. Annals of the Institute of Statistical Mathematics, 13 (1):91–103, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868862.

Ogawa:1961:ERA

[260] Junjiro Ogawa. The effect of randomization on the analysis of randomized block design. Annals of the Institute of Statistical Mathematics, 13 (1):105–117, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868863.

Rustagi:1961:BVM

[261] J. S. Rustagi. Bounds for the variance of Mann–Whitney statistic. Annals of the Institute of Statistical Mathematics, 13(1):119–126, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868864.

Akaike:1961:UOS

[262] Hieotugu Akaike. Undamped oscillation of the sample autocovariance function and the effect of prewhitening operation. Annals of the Institute of Statistical Mathematics, 13(1): 127–143, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868865.

Khatri:1961:NIE

[263] C. G. Khatri. A note on the interval estimation related to the regression matrix. Annals of the Institute of Statistical Mathematics, 13 (1):145–146, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02868866.

Sakino:1961:AEM

[264] Sigeki Sakino. On the analysis of epidemic model II (theory and application). Annals of the Institute of Statistical Mathematics, 13(1):147–163, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868867.

Uematu:1961:SMC

[265] Tosio Uematu. Some models concerning statistical treatment of a certain congestion phenomenon. Annals of the Institute of Statistical Mathematics, 13(1):165–185, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868868.

Taga:1961:SSL

[266] Yasushi Taga. On some sequential life tests. Annals of the Institute of Statistical Mathematics, 13(1):187–199, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868869.

Tanaka:1961:CIG

[267] Masao Tanaka. On a confidence interval of given length for the parameter of the binomial and the Poisson distributions. Annals of the Institute of Statistical Mathematics, 13(1):201–215, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868870.

Berman:1961:CBL

[268] Simeon M. Berman. Convergence to bivariate limiting extreme value distributions. Annals of the Institute of Statistical Mathematics, 13(1):217–223, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868871.

LikeS:1961:DCL

[269] Jirí LikeS. On the distribution of certain linear functions of ordered sample from exponential population. Annals of the Institute of Statistical Mathematics, 13(1):225–230, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868872.

Sibuya:1961:EOR

[270] Masaaki Sibuya. On exponential and other random variable generators. Annals of the Institute of Statistical Mathematics, 13(1):231–237, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868873.

Khatri:1961:SCB

[271] C. G. Khatri. Simultaneous confidence bounds on the departures from a particular kind of multicollinearity. Annals of the Institute of Statistical Mathematics, 13(1):239–242, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868874.

Aoyama:1961:NOR

[272] Hirojiro Aoyama. Note on ordered random intervals and its application. Annals of the Institute of Statistical Mathematics, 13(1):243–250, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868875.

Sibuya:1961:MPA

[273] Masaaki Sibuya. On a model in probit analysis. Annals of the Institute of Statistical Mathematics, 13(1): 251–257, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868876.

Ikeda:1961:NCS

[274] Sadao Ikeda. A note on the characterization of Shannon–Wiener’s measure of information. Annals of the Institute of Statistical Mathematics, 13 (1):259–266, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868877.

Anonymous:1961:HCb

[275] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 13(1):??, December 1961. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akaike:1962:DLW

[276] Hirotugu Akaike. On the design of lag window for the estimation of spectra. Annals of the Institute of Statistical Mathematics, 14(1):1–21, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868621.

Akaike:1962:SEF

[277] Hirotugu Akaike and Yasufumi Yamanouchi. On the statistical estimation of frequency response function. Annals of the Institute of Statistical Mathematics, 14(1):23–56, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868622.

Khatri:1962:MEA

[278] C. G. Khatri. A method for estimating approximately the parameters of a certain class of distributions from grouped observations. Annals of the Institute of Statistical Mathematics, 14(1):57–62, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868623.

Kapadia:1962:MSS

[279] C. H. Kapadia. Minimal sufficient statistics for the partially balanced incomplete block (PBIB) design with two associate classes under an Eisenhart model II. Annals of the Institute of Statistical Mathematics, 14(1):63–71, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868624.

Ikeda:1962:CKL

[280] Sadao Ikeda. On characterization of the Kullback–Leibler mean information for continuous probability distributions. Annals of the Institute of Statistical Mathematics, 14(1):73–79, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868625.

Sibuya:1962:MGU

[281] Masaaki Sibuya. A method for generating uniformly distributed points on N-dimensional spheres. Annals of the Institute of Statistical Mathematics, 14 (1):81–85, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868626.

Sibuya:1962:NUM

[282] Masaaki Sibuya. A note on the use of median ranges. Annals of the Institute of Statistical Mathematics, 14 (1):87–89, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868627.

Pakshirajan:1962:NTP

[283] R. P. Pakshirajan and S. S. Chitgopekar. A note on a test procedure with a sample from a normal population when an upper bound to the standard deviation is known. Annals of the Institute of Statistical Mathematics, 14(1):91–93, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868628.

Khatri:1962:IBI

[284] C. G. Khatri and S. M. Shah. An inequality for balanced incomplete block design. Annals of the Institute of Statistical Mathematics, 14(1):95–96, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02868629.

Taga:1962:OLT

[285] Yasushi Taga. On the optimal life test procedures based on a cost model. Annals of the Institute of Statistical Mathematics, 14(1):97–106, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868630.

Ikeda:1962:NCC

[286] Sadao Ikeda. Necessary conditions for the convergence of Kullback–Leibler’s mean information. Annals of the Institute of Statistical Mathematics, 14 (1):107–118, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868631.

Sen:1962:SNP

[287] P. K. Sen. On Studentized nonparametric multi-sample location tests. Annals of the Institute of Statistical Mathematics, 14(1):119–131, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868632.

Khatri:1962:FPG

[288] C. G. Khatri. A fitting procedure for a generalised binomial distribution. Annals of the Institute of Statistical Mathematics, 14(1):133–141, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868633.

Rao:1962:ERE

[289] J. N. K. Rao. On the estimation of the relative efficiency of sampling procedures. Annals of the Institute of Statistical Mathematics, 14(1):143–150, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868634.

Bennett:1962:HTI

[290] B. M. Bennett. On a heuristic treatment of the ‘indices of dispersion’. Annals of the Institute of Statistical Mathematics, 14(1):151–157, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868635.

Sibuya:1962:FCN

[291] Masaaki Sibuya. Further consideration on normal random variable generator. Annals of the Institute of Statistical Mathematics, 14(1):159–165, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868636.

Khatri:1962:DOS

[292] C. G. Khatri. Distributions of order statistics for discrete case. Annals of the Institute of Statistical Mathematics, 14(1):167–171, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868637.

Shimizu:1962:CND

[293] Ryoichi Shimizu. Characterization of the normal distribution II. Annals of the Institute of Statistical Mathematics, 14(1):173–178, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868638.

Patil:1962:CPG

[294] G. P. Patil. Certain properties of the generalized power series distribution II. Annals of the Institute of Statistical Mathematics, 14(1):179–182, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868639.

Menon:1962:ISC

[295] M. V. Menon. An implication of stochastic convergence. Annals of the Institute of Statistical Mathematics, 14 (1):183–184, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02868640.

Isii:1962:STT

[296] Keiiti Isii. On sharpness of Tchebychefftype inequalities. Annals of the Institute of Statistical Mathematics, 14 (1):185–197, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868641.

Gupta:1962:SRP

[297] Shanti S. Gupta. On a selection and ranking procedure for gamma populations. Annals of the Institute of Statistical Mathematics, 14(1):199–212, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868642.

Kshirsagar:1962:ENC

[298] A. M. Kshirsagar. Effect of noncentrality on the Bartlett decomposition of a Wishart matrix. Annals of the Institute of Statistical Mathematics, 14(1):217–228, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868643.

Abbott:1962:TSE

[299] J. H. Abbott and J. I. Rosenblatt. Two stage estimation with one observation on the first stage. Annals of the Institute of Statistical Mathematics, 14(1):229–235, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868644.

John:1962:CS

[300] S. John. On classification by the statistics R and Z. Annals of the Institute of Statistical Mathematics, 14(1): 237–246, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868645.

Jaiswal:1962:AVC

[301] M. C. Jaiswal. Asymptotic variances and covariances of the moment estimates of parameters of a truncated Pearson type III distribution. Annals of the Institute of Statistical Mathematics, 14(1):247–250, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868646.

Aoyama:1962:SRS

[302] Hirojiro Aoyama. Stratified random sampling with optimum allocation for multivariate population. Annals of the Institute of Statistical Mathematics, 14(1):251–258, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868647.

Huzii:1962:SME

[303] Mituaki Huzii. On a simplified method of the estimation of the correlogram for a stationary Gaussian process. Annals of the Institute of Statistical Mathematics, 14(1):259–268, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868648.

Nisihira:1962:MSA

[304] Sigeki Nisihira. La mobilité sociale au Japon. (French) [Social mobility in Japan]. Annals of the Institute of Statistical Mathematics, 14(1): 269–278, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868649.

Anonymous:1962:HC

[305] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 14(1):??, December 1962. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Taga:1963:LDM

[306] Yasushi Taga. On the limiting distributions in Markov renewal processes with finitely many states. Annals of the Institute of Statistical Mathematics, 15(1):1–10, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865897.

Rao:1963:DA

[307] M. M. Rao. Discriminant analysis. Annals of the Institute of Statistical Mathematics, 15(1):11–24, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865898.

Rao:1963:TIA

[308] P. V. Rao. The F -test in the intrablock analysis of a class of PBIB designs. Annals of the Institute of Statistical Mathematics, 15(1):25–36, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865899.

Bhattacharya:1963:BLT

[309] Samir Kumar Bhattacharya and Mahabaleshwara Holla. Bivariate lifetesting models for two component systems. Annals of the Institute of Statistical Mathematics, 15(1):37–43, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865900.

Blum:1963:EQ

[310] J. R. Blum and Judah Rosenblatt. On estimating quantiles. Annals of the Institute of Statistical Mathematics, 15 (1):45–50, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02865901.

Kabe:1963:ESF

[311] D. G. Kabe. Estimation of a set of fixed variates for observed values of dependent variates with normal multivariate regression models subjected to linear restrictions. Annals of the Institute of Statistical Mathematics, 15(1):51–59, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865902.

Sibuya:1963:RUE

[312] Masaaki Sibuya. Randomized unbiased estimation of restricted parameters. Annals of the Institute of Statistical Mathematics, 15(1):61–66, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865903.

Rao:1963:TSU

[313] J. N. K. Rao. On two systems of unequal probability sampling without replacement. Annals of the Institute of Statistical Mathematics, 15 (1):67–72, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02865904.

Takamatsu:1963:CSIa

[314] Shunro Takamatsu. On the comeand-stay interarrival time in a modified queueing system M/G/1. Annals of the Institute of Statistical Mathematics, 15(1):73–78, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865905.

Konijn:1963:MIE

[315] H. S. Konijn. Minimax interval estimates with a shortness criterion: A new formulation. Annals of the Institute of Statistical Mathematics, 15 (1):79–81, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02865906.

Govindarajulu:1963:NII

[316] Zakkula Govindarajulu and Yukio Suzuki. A note on an identity involving binomial coefficients. Annals of the Institute of Statistical Mathematics, 15 (1):83–85, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02865907.

Ikeda:1963:AEP

[317] Sadao Ikeda. Asymptotic equivalence of probability distributions with applications to some problems of asymptotic independence. Annals of the Institute of Statistical Mathematics, 15 (1):87–116, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02865908.

Sen:1963:WRS

[318] P. K. Sen. On weighted rank-sum tests for dispersion. Annals of the Institute of Statistical Mathematics, 15 (1):117–135, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02865909.

Kendall:1963:ITL

[319] David G. Kendall. Information theory and the limit-theorem for Markov chains and processes with a countable infinity of states. Annals of the Institute of Statistical Mathematics, 15 (1):137–143, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02865910.

Shenton:1963:NBA

[320] L. R. Shenton. A note on bounds for the asymptotic sampling variance of the maximum likelihood estimator of a parameter in the negative binomial distribution. Annals of the Institute of Statistical Mathematics, 15(1):145–151, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865911.

Walsh:1963:BPP

[321] John E. Walsh. Bounded probability properties of Kolmogorov–Smirnov and similar statistics for discrete data. Annals of the Institute of Statistical Mathematics, 15(1):153–158, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865912.

Suzuki:1963:FT

[322] Giitiro Suzuki. On a functional transform. Annals of the Institute of Statistical Mathematics, 15(1):159–165, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865913.

Isii:1963:LTS

[323] Keiiti Isii. On a limit theorem for a stochastic process related to quantum biophysics of vision. Annals of the Institute of Statistical Mathematics, 15(1):167–175, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865914.

Weiss:1963:SBP

[324] Lionel Weiss. Sequential Bayes procedures which never observe more than a bounded number of observations. Annals of the Institute of Statistical Mathematics, 15(1):177–185, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865915.

Taga:1963:HOM

[325] Yasushi Taga. On high order moments of the number of renewals. Annals of the Institute of Statistical Mathematics, 15(1):187–196, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865916.

Clark:1963:SEM

[326] Charles E. Clark. Sampling efficiency in Monte Carlo analyses. Annals of the Institute of Statistical Mathematics, 15(1):197–206, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865917.

Takamatsu:1963:CSIb

[327] Shunro Takamatsu. On the comeand-stay interarrival time in a modified queueing system GI/M/1. Annals of the Institute of Statistical Mathematics, 15(1):207–213, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02865918.

Seshadri:1963:CBD

[328] V. Seshadri and G. P. Patil. A characterization of a bivariate distribution by the marginal and the conditional distributions of the same component. Annals of the Institute of Statistical Mathematics, 15(1):215–221, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02865919.

Anonymous:1963:HC

[329] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 15(1):??, December 1963. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Anscombe:1964:NLF

[330] F. J. Anscombe. Normal likelihood functions. Annals of the Institute of Statistical Mathematics, 16 (1):1–19, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868558.

Arrow:1964:OCP

[331] Kenneth J. Arrow. Optimal capital policy, the cost of capital, and myopic decision rules. Annals of the Institute of Statistical Mathematics, 16 (1):21–30, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868559.

Chernoff:1964:EM

[332] Herman Chernoff. Estimation of the mode. Annals of the Institute of Statistical Mathematics, 16(1):31–41, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868560.

Cochran:1964:CTM

[333] William G. Cochran. Comparison of two methods of handling covariates in discriminatory analysis. Annals of the Institute of Statistical Mathematics, 16(1):43–53, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868561.

Connor:1964:MER

[334] W. S. Connor and Gertrude M. Cox. Methodology for estimating reliability. Annals of the Institute of Statistical Mathematics, 16(1):55–67, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868562.

Guttman:1964:DTD

[335] Louis Guttman. Deviation theory for dichotomies. Annals of the Institute of Statistical Mathematics, 16 (1):69–78, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868563.

Klefer:1964:OEIa

[336] J. Klefer and J. Wolfowitz. Optimum extrapolation and interpolation designs, I. Annals of the Institute of Statistical Mathematics, 16(1):79–108, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868564.

Moran:1964:RCS

[337] P. A. P. Moran. On the range of cumulative sums. Annals of the Institute of Statistical Mathematics, 16(1): 109–112, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868565.

Nisihira:1964:LPJa

[338] Sigeki Nisihira. L’opinion publique des japonais au milieu du vingtiéme siécle. (French) [Japanese public opinion in the Twentieth Century]. Annals of the Institute of Statistical Mathematics, 16(1):113–128, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868566.

Renyi:1964:EPP

[339] Alfréd Rényi. On an extremal property of the Poisson process. Annals of the Institute of Statistical Mathematics, 16(1):129–133, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868567.

Slotani:1964:TRM

[340] Mlnoru Slotani. Tolerance regions for a multivariate normal population. Annals of the Institute of Statistical Mathematics, 16(1):135–153, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868568.

Stein:1964:IUE

[341] Charles Stein. Inadmissibility of the usual estimator for the variance of a normal distribution with unknown mean. Annals of the Institute of Statistical Mathematics, 16(1):155–160, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868569.

Suzuki:1964:USE

[342] Yukio Suzuki. On the use of some extraneous information in the estimation of the coefficients of regression. Annals of the Institute of Statistical Mathematics, 16(1):161–173, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868570.

Aoyama:1964:AND

[343] Hirojiro Aoyama. On an analysis of natural disaster on the railway. Annals of the Institute of Statistical Mathematics, 16(1):175–184, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868571.

Barankin:1964:PE

[344] Edward W. Barankin. Probability and the east. Annals of the Institute of Statistical Mathematics, 16(1): 185–230, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868572.

Hayashi:1964:MQD

[345] Chikio Hayashi. Multidimensional quantification of the data obtained by the method of paired comparison. Annals of the Institute of Statistical Mathematics, 16(1):231–245, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868573.

See corrections [481].

Hudimoto:1964:DFT

[346] Hirosi Hudimoto. On a distributionfree two-way classification. Annals of the Institute of Statistical Mathematics, 16(1):247–253, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868574.

Isida:1964:SFS

[347] Masatugu Isida. 10,000 spots forest survey. Annals of the Institute of Statistical Mathematics, 16(1): 255–276, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868575.

Isii:1964:ITC

[348] Keiiti Isii. Inequalities of the types of Chebyshev and Cramér–Rao and mathematical programming. Annals of the Institute of Statistical Mathematics, 16(1):277–293, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868576.

Klefer:1964:OEIb

[349] J. Klefer and J. Wolfowitz. Optimum extrapolation and interpolation designs II. Annals of the Institute of Statistical Mathematics, 16(1):295–303, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868577.

Matusita:1964:DDR

[350] Kameo Matusita. Distance and decision rules. Annals of the Institute of Statistical Mathematics, 16(1): 305–315, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868578.

Motoo:1964:SAF

[351] Mlnoru Motoo. The sweeping-out of additive functionals and processes on the boundary. Annals of the Institute of Statistical Mathematics, 16(1): 317–345, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868579.

Nisihira:1964:LPJb

[352] Sigeki Nisihira. L’opinion publique des japonais, II au milieu du vingtième siècle. (French) [Japanese public opinion in the Twentieth Century, II]. Annals of the Institute of Statistical Mathematics, 16(1):347–367, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868580.

Ogawa:1964:ADL

[353] Junjiro Ogawa and Sadao Ikeda. On the asymptotic distribution of the likelihood ratio under the regularity conditions due to Doob. Annals of the Institute of Statistical Mathematics, 16(1):369–385, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868581.

Skimizu:1964:DID

[354] Ryoichi Skimizu. On the decomposition of infinitely divisible characteristic functions with a continuous Poisson spectrum. Annals of the Institute of Statistical Mathematics, 16(1): 387–407, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868582.

Sibuya:1964:NMD

[355] Masaaki Sibuya, Isao Yoshimura, and Ryoichi Shimizu. Negative multinomial distribution. Annals of the Institute of Statistical Mathematics, 16 (1):409–426, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868583.

Taga:1964:NDN

[356] Yasushi Taga. A note on the degree of normal approximation to the distribution function of the mean of samples from finite populations. Annals of the Institute of Statistical Mathematics, 16(1):427–430, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868584.

Uematu:1964:MLD

[357] Tosio Uematu. On a multidimensional linear discriminant function. Annals of the Institute of Statistical Mathematics, 16(1):431–437, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868585.

Anonymous:1964:HC

[358] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 16(1):??, December 1964. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Asano:1965:EMP

[359] Chooichiro Asano. On estimating multinomial probabilities by pooling incomplete samples. Annals of the Institute of Statistical Mathematics, 17 (1):1–13, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868148.

Sukhatme:1965:CST

[360] B. V. Sukhatme and M. S. Avadhani. Controlled selection a technique in random sampling. Annals of the Institute of Statistical Mathematics, 17(1):15–28, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868149.

Patel:1965:EPT

[361] R. C. Patel. Estimates of parameters of truncated inverse Gaussian distribution. Annals of the Institute of Statistical Mathematics, 17(1):29–33, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868150.

Krishnaiah:1965:SAM

[362] P. R. Krishnaiah. On the simultaneous ANOVA and MANOVA tests. Annals of the Institute of Statistical Mathematics, 17(1):35–53, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868151.

Bennett:1965:MSR

[363] B. M. Bennett. On multivariate signed rank tests. Annals of the Institute of Statistical Mathematics, 17 (1):55–61, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868152.

Srivastava:1965:MEG

[364] J. N. Srivastava. A multivariate extension of the Gauss–Markov theorem. Annals of the Institute of Statistical Mathematics, 17(1):63–66, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868153.

Hayakawa:1965:THS

[365] Takesi Hayakawa and D. G. Kabe. On testing the hypothesis that submatrices of the multivariate regression matrices of k populations are equal. Annals of the Institute of Statistical Mathematics, 17(1):67–73, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868154.

Kabe:1965:NDR

[366] D. G. Kabe. On the noncentral distribution of Rao’s U statistic. Annals of the Institute of Statistical Mathematics, 17(1):75–80, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868155.

Swain:1965:LBP

[367] A. K. P. C. Swain. A lower bound to the probability of variance ratio. Annals of the Institute of Statistical Mathematics, 17(1):81–84, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868156.

Mardia:1965:TFO

[368] K. V. Mardia. Tippett’s formulas and other results on sample range and extremes. Annals of the Institute of Statistical Mathematics, 17(1):85–91, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868157.

Basu:1965:CED

[369] A. P. Basu. On characterizing the exponential distribution by order statistics. Annals of the Institute of Statistical Mathematics, 17(1):93–96, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868158.

Bhattacharya:1965:LTD

[370] S. K. Bhattacharya and M. S. Holla. On a life test distribution with stochastic deviations in the mean. Annals of the Institute of Statistical Mathematics, 17(1):97–104, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868159.

Taga:1965:OSP

[371] Yasushi Taga. The optimal sampling procedure for estimating the mean of stationary Markov processes. Annals of the Institute of Statistical Mathematics, 17(1):105–112, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868160.

Shimizu:1965:CCI

[372] Ryoichi Shimizu. Certain class of infinitely divisible characteristic functions. Annals of the Institute of Statistical Mathematics, 17(1):115–132, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868161.

Rao:1965:EDO

[373] M. M. Rao. Existence and determination of optimal estimators relative to convex loss. Annals of the Institute of Statistical Mathematics, 17(1): 133–147, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868162.

Kusumoto:1965:NCE

[374] Kumaichi Kusumoto. A necessary condition for the existence of regular and symmetrical PBIB designs of T_{3} type. Annals of the Institute of Statistical Mathematics, 17(1):149–165, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868163.

Krishnaiah:1965:MGS

[375] P. R. Krishnaiah. On a multivariate generalization of the simultaneous analysis of variance test. Annals of the Institute of Statistical Mathematics, 17(1):167–173, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868164.

Khatri:1965:NCB

[376] C. G. Khatri. A note on the confidence bounds for the characteristic roots of dispersion matrices of normal variates. Annals of the Institute of Statistical Mathematics, 17(1):175–183, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868165.

Akaike:1965:SEF

[377] Hirotugu Akaike. On the statistical estimation of the frequency response function of a system having multiple input. Annals of the Institute of Statistical Mathematics, 17(1):185–210, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868166.

Krishnan:1965:TQA

[378] T. Krishnan. Truncation in quantal assay. Annals of the Institute of Statistical Mathematics, 17(1):211–231, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868167.

Sen:1965:SAP

[379] P. K. Sen. On some asymptotic properties of a class of non-parametric tests based on the number of rare exceedances. Annals of the Institute of Statistical Mathematics, 17(1): 233–255, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868168.

Takahasi:1965:NMB

[380] Koiti Takahasi. Note on the multivariate Burr’s distribution. Annals of the Institute of Statistical Mathematics, 17(1):257–260, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868169.

Sugiura:1965:ETS

[381] Nariaki Sugiura. An example of the two-sided Wilcoxon test which is not unbiased. Annals of the Institute of Statistical Mathematics, 17(1): 261–263, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868170.

Gould:1965:IIS

[382] H. W. Gould. An identity involving Stirling numbers. Annals of the Institute of Statistical Mathematics, 17 (1):265–269, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868171.

Suzuki:1965:CEM

[383] Giitiro Suzuki. A consistent estimator for the mean deviation of the Pearson type distribution. Annals of the Institute of Statistical Mathematics, 17(1):271–285, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868172.

Kamat:1965:PMD

[384] A. R. Kamat. A property of the mean deviation (of some discrete distributions). Annals of the Institute of Statistical Mathematics, 17(1):287–293, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868173.

Ikeda:1965:BVY

[385] Sadao Ikeda. On Bouman–Velden– Yamamoto’s asymptotic evaluation formula for the probability of visual response in a certain experimental research in quantum-biophysics of vision. Annals of the Institute of Statistical Mathematics, 17(1):295–310, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868174.

Hanson:1965:ICM

[386] M. A. Hanson. Inequality constrained maximum likelihood estimation. Annals of the Institute of Statistical Mathematics, 17(1):311–321, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868175.

Puri:1965:STH

[387] Madan L. Puri. On some tests of homogeneity of variances. Annals of the Institute of Statistical Mathematics, 17(1):323–330, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868176.

Asano:1965:RTC

[388] Chooichiro Asano. Runs test for a circular distribution and a table of probabilities. Annals of the Institute of Statistical Mathematics, 17(1): 331–346, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02868177.

Kshirsagar:1965:GFT

[389] A. M. Kshirsagar and R. P. Gupta. The goodness of fit of two (or more) hypothetical principal components. Annals of the Institute of Statistical Mathematics, 17(1):347–356, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02868178.

Hocking:1965:DPL

[390] R. R. Hocking. The distribution of a projected least squares estimator. Annals of the Institute of Statistical Mathematics, 17(1):357–362, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868179.

Samanta:1965:NPO

[391] M. Samanta. A note on the problem of optimum truncation of a bivariate population in stratified random sampling. Annals of the Institute of Statistical Mathematics, 17(1):363–375, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868180.

Holla:1965:DCD

[392] M. S. Holla and S. K. Bhattacharya. On a discrete compound distribution. Annals of the Institute of Statistical Mathematics, 17(1):377–384, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02868181.

Ninomija:1965:NMA

[393] Satoki Ninomija. Über eine numerische Methode zur Auflösung der komplexen Gleichungen. (German) [On a numerical method for solution of complex equations]. Annals of the Institute of Statistical Mathematics, 17(1): 385–398, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02868182.

Anonymous:1965:HC

[394] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 17(1):??, December 1965. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kusama:1966:CBS

[395] Tokitake Kusama. On classes of Bayes solutions. Annals of the Institute of Statistical Mathematics, 18 (1):1–11, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869512.

DeGroot:1966:OAO

[396] M. H. DeGroot. Optimal allocation of observations. Annals of the Institute of Statistical Mathematics, 18 (1):13–28, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869513.

Suzuki:1966:GCT

[397] Giitiro Suzuki. On the Glivenko– Cantelli theorem. Annals of the Institute of Statistical Mathematics, 18 (1):29–37, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869514. See corrections

[429].

Patil:1966:MVU

[398] G. P. Patil and J. K. Wani. Minimum variance unbiased estimation of the distribution function admitting a sufficient statistic. Annals of the Institute of Statistical Mathematics, 18 (1):39–47, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869515.

Shimizu:1966:RSS

[399] Ryoichi Shimizu. Remarks on sufficient statistics. Annals of the Institute of Statistical Mathematics, 18 (1):49–55, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869516.

Srivastava:1966:SEP

[400] J. N. Srivastava and R. C. Bose. Some economic partially balanced 2^{m} factorial fractions. Annals of the Institute of Statistical Mathematics, 18 (1):57–73, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869517.

Khatri:1966:NMM

[401] C. G. Khatri. A note on a MANOVA model applied to problems in growth curve. Annals of the Institute of Statistical Mathematics, 18(1):75–86, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869518.

Sen:1966:CSW

[402] P. K. Sen and Z. Govindarajulu. On a class of c-sample weighted ranksum tests for location and scale. Annals of the Institute of Statistical Mathematics, 18(1):87–105, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869519.

Shah:1966:EPD

[403] S. M. Shah and M. C. Jaiswal. Estimation of parameters of doubly truncated normal distribution from first four sample moments. Annals of the Institute of Statistical Mathematics, 18(1):107–111, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02869520.

Bland:1966:NSN

[404] R. P. Bland and D. B. Owen. A note on singular normal distributions. Annals of the Institute of Statistical Mathematics, 18(1):113–116, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02869521.

Rao:1966:CUR

[405] T. J. Rao. On certain unbiased ratio estimators. Annals of the Institute of Statistical Mathematics, 18(1): 117–121, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869522.

Akaike:1966:NHO

[406] Hirotugu Akaike. Note on higher order spectra. Annals of the Institute of Statistical Mathematics, 18(1): 123–126, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869523.

Suzuki:1966:DCD

[407] Giitiro Suzuki. Discrete compound decision problem. Annals of the Institute of Statistical Mathematics, 18 (1):127–139, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869524.

Kusama:1966:RAD

[408] Tokitake Kusama. Remarks on admissibility of decision functions. Annals of the Institute of Statistical Mathematics, 18(1):141–148, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869525.

Weiss:1966:APC

[409] Lionel Weiss. On the asymptotic power of Cramér–von Mises tests of fit. Annals of the Institute of Statistical Mathematics, 18(1):149–153, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869526.

Kaufman:1966:AEM

[410] S. Kaufman. Asymptotic efficiency of the maximum likelihood estimator. Annals of the Institute of Statistical Mathematics, 18(1):155–178, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869527.

Cacoullos:1966:EMD

[411] Theophilos Cacoullos. Estimation of a multivariate density. Annals of the Institute of Statistical Mathematics, 18 (1):179–189, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869528.

Hayakawa:1966:DQF

[412] Takesi Hayakawa. On the distribution of a quadratic form in a multivariate normal sample. Annals of the Institute of Statistical Mathematics, 18(1):191–201, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869529.

Ruben:1966:SSV

[413] Harold Ruben. On the simultaneous stabilization of variances and covariances. Annals of the Institute of Statistical Mathematics, 18(1):203–210, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869530.

Bhat:1966:QGM

[414] U. N. Bhat. The queue gi/m/ 2 with service rate depending on the number of busy servers. Annals of the Institute of Statistical Mathematics, 18(1):211–221, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869531.

Shanbhag:1966:CSN

[415] D. N. Shanbhag. On congestion systems with negative exponential desired service time distributions. Annals of the Institute of Statistical Mathematics, 18(1):223–228, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869532.

Suzuki:1966:SDP

[416] Yukio Suzuki. On sequential decision problems with delayed observations. Annals of the Institute of Statistical Mathematics, 18(1):229–267, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869533.

See corrections [455].

Akaike:1966:UNG

[417] Hlrotugu Akaike. On the use of non–Gaussian process in the identification of a linear dynamic system. Annals of the Institute of Statistical Mathematics, 18(1):269–276, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869534.

Kambo:1966:EBB

[418] N. Singh Kambo and Samuel Kotz. On exponential bounds for binomial probabilities. Annals of the Institute of Statistical Mathematics, 18(1): 277–287, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869535.

Cobb:1966:ALB

[419] E. B. Cobb and Bernard Harris. An asymptotic lower bound for the entropy of discrete populations with application to the estimation of entropy for approximately uniform populations. Annals of the Institute of Statistical Mathematics, 18(1):289–297, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02869536.

Srivastava:1966:MSP

[420] M. S. Srivastava. On a multivariate slippage problem. I. Annals of the Institute of Statistical Mathematics, 18(1):299–305, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02869537.

Alam:1966:SMN

[421] Khursheed Alam and M. Haseeb Rizvi. Selection from multivariate normal populations. Annals of the Institute of Statistical Mathematics, 18(1): 307–318, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02869538.

Sen:1966:NSC

[422] Pranab Kumar Sen. On nonparametric simultaneous confidence regions and tests for the one criterion analysis of variance problem. Annals of the Institute of Statistical Mathematics, 18(1):319–336, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02869539.

Chu:1966:QOE

[423] J. T. Chu and K. Ya’Coub. Quadratic order estimates and moments of normal order statistics. Annals of the Institute of Statistical Mathematics, 18 (1):337–341, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869540.

Konijn:1966:NEC

[424] H. S. Konijn. Non-existence of consistent estimator sequences and unbiased estimators: a practical example. Annals of the Institute of Statistical Mathematics, 18(1):343–350, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869541.

Blum:1966:SSP

[425] J. R. Blum and Judah Rosenblatt. On some statistical problems requiring purely sequential sampling schemes. Annals of the Institute of Statistical Mathematics, 18(1):351–355, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869542.

Dubey:1966:CPD

[426] Satya D. Dubey. Compound Pascal distributions. Annals of the Institute of Statistical Mathematics, 18 (1):357–365, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02869543.

Kabe:1966:DTD

[427] D. G. Kabe. Dirichlet’s transformation and distributions of linear functions of ordered gamma variates. Annals of the Institute of Statistical Mathematics, 18(1):367–374, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869544.

Khatri:1966:NLS

[428] C. G. Khatri. A note on a large sample distribution of a transformed multiple correlation coefficient. Annals of the Institute of Statistical Mathematics, 18(1):375–380, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02869545.

Suzuki:1966:CGC

[429] Giitiro Suzuki. Corrections to “On the Glivenko–Cantelli theorem”. Annals of the Institute of Statistical Mathematics, 18(1):381, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02869546. See

[397].

Anonymous:1966:HC

[430] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 18(1):??, December 1966. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayakawa:1967:DML

[431] Takesi Hayakawa. On the distribution of the maximum latent root of a positive definite symmetric random matrix. Annals of the Institute of Statistical Mathematics, 19(1):1–17, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911665.

Naik:1967:BRC

[432] Umesh D. Naik. Bayes rules for comparisons of scale parameters of certain distributions. Annals of the Institute of Statistical Mathematics, 19 (1):19–37, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911666.

Kori:1967:RBM

[433] Tosiaki Kori. On the resolvent of a Brownian motion with drift. Annals of the Institute of Statistical Mathematics, 19(1):39–53, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911667.

Rao:1967:QMB

[434] S. Subba Rao. Queueing models with balking and reneging. Annals of the Institute of Statistical Mathematics, 19(1):55–71, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911668.

Kusumoto:1967:ASN

[435] Kumaichi Kusumoto. Association schemes of new types and necessary conditions for existence for regular and symmetrical PBIB designs with those association schemes. Annals of the Institute of Statistical Mathematics, 19(1):73–100, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911669.

Taga:1967:OSO

[436] Yasushi Taga. On optimum stratification for the objective variable based on concomitant variables using prior information. Annals of the Institute of Statistical Mathematics, 19(1): 101–129, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911670. See corrections

[456].

Chikkagoudar:1967:TPS

[437] M. S. Chikkagoudar. Two-phase sampling for PPS estimation. Annals of the Institute of Statistical Mathematics, 19(1):131–142, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911671.

Khatri:1967:MTT

[438] C. G. Khatri and K. C. S. Pillai. On the moments of traces of two matrices in multivariate analysis. Annals of the Institute of Statistical Mathematics, 19(1):143–156, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911672.

Gupta:1967:LRV

[439] R. P. Gupta. Latent roots and vectors of a Wishart matrix. Annals of the Institute of Statistical Mathematics, 19(1):157–165, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911673.

Pillai:1967:DSE

[440] K. C. Sreedharan Pillai and Arjun K. Gupta. On the distribution of the second elementary symmetric function of the roots of a matrix. Annals of the Institute of Statistical Mathematics, 19(1):167–179, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911674.

Matusita:1967:NAS

[441] Kameo Matusita. On the notion of affinity of several distributions and some of its applications. Annals of the Institute of Statistical Mathematics, 19(1):181–192, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911675.

Weiss:1967:MPE

[442] L. Weiss and J. Wolfowitz. Maximum probability estimators. Annals of the Institute of Statistical Mathematics, 19(1):193–206, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911676.

Stigum:1967:DTA

[443] Bernt P. Stigum. A decision theoretic approach to time series analysis. Annals of the Institute of Statistical Mathematics, 19(1):207–243, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911677.

Siotani:1967:SAL

[444] Minoru Siotani. Some applications of Loewner’s ordering on symmetric matrices. Annals of the Institute of Statistical Mathematics, 19(1): 245–259, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

1007/BF02911678. See cancellation [482].

Alam:1967:TSEa

[445] Khursheed Alam. A two-sample estimate of a common mean. Annals of the Institute of Statistical Mathematics, 19(1):261–270, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911679.

Alam:1967:TSEb

[446] Khursheed Alam. A two-sample estimate of the largest mean. Annals of the Institute of Statistical Mathematics, 19(1):271–283, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911680.

Puri:1967:CIO

[447] Madan L. Puri. Combining independent one-sample tests of significance. Annals of the Institute of Statistical Mathematics, 19(1):285–300, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911681.

Naik:1967:BUS

[448] Umesh D. Naik. On bilateral and unilateral statistics for tests concerning means of normal populations. Annals of the Institute of Statistical Mathematics, 19(1):301–312, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911682.

Ogawa:1967:NDS

[449] Junjiro Ogawa, Sadao Ikeda, and Motoyasu Ogasawara. On the nulldistribution of the F -statistics for testing a ‘partial’ null-hypothesis in a randomized partially balanced incomplete block design with m associate classes under the Neyman model. Annals of the Institute of Statistical Mathematics, 19(1):313–330, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911683.

Das:1967:GDR

[450] M. N. Das and A. Dey. Group-divisible rotatable designs. Annals of the Institute of Statistical Mathematics, 19 (1):331–347, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911684. See corrections

[497].

Rao:1967:ETT

[451] P. V. Rao. The effect of truncation on the F -test for a class of PBIB designs. Annals of the Institute of Statistical Mathematics, 19(1):349–354, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911685.

Press:1967:SCB

[452] S. James Press. On the sample covariance from a bivariate normal distribution. Annals of the Institute of Statistical Mathematics, 19(1):355–361, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911686.

Hayashi:1967:NMQ

[453] Chikio Hayashi. Note on multidimensional quantification of data obtained by paired comparison. Annals of the Institute of Statistical Mathematics, 19(1):363–365, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911687.

Dickey:1967:BHD

[454] James M. Dickey. A Bayesian hypothesis-decision procedure. Annals of the Institute of Statistical Mathematics, 19(1):367–369, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911688.

Suzuki:1967:CSD

[455] Yukio Suzuki. Corrections to “On sequential decision problems with delayed observations”. Annals of the Institute of Statistical Mathematics, 19(1):371, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02911689. See

[416].

Taga:1967:COS

[456] Yasushi Taga. Corrections to “On optimum stratification for the objective variable based on concomitant variables using prior information”. Annals of the Institute of Statistical Mathematics, 19(1):372, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02911690. See [436].

Suzuki:1967:EPS

[457] Giitiro Suzuki. On exact probabilities of some generalized Kolmogorov’s D-statistics. Annals of the Institute of Statistical Mathematics, 19(1): 373–388, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911691.

Weiss:1967:SPC

[458] Lionel Weiss. Some properties of a class of tests of fit. Annals of the Institute of Statistical Mathematics, 19 (1):389–399, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911692.

Ali:1967:BLE

[459] M. M. Ali and L. K. Chan. Ban linear estimates of the parameters of the normal distribution from censored samples. Annals of the Institute of Statistical Mathematics, 19(1):401–411, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911693.

Gupta:1967:UE

[460] Milan K. Gupta. Unbiased estimate for 1/p. Annals of the Institute of Statistical Mathematics, 19(1):413–416, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911694.

Srivastava:1967:EGM

[461] J. N. Srivastava. On the extension of Gauss–Markov theorem to complex multivariate linear models. Annals of the Institute of Statistical Mathematics, 19(1):417–437, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911695.

Nakajima:1967:MTR

[462] Nobuyuki Nakajima and Keiiti Isii. Multidimensional tolerance regions based on a large sample. Annals of the Institute of Statistical Mathematics, 19 (1):439–449, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911696.

Sen:1967:SNG

[463] Pranab Kumar Sen. On some nonparametric generalizations of Wilks’ tests for H_{M}, H_{V C} and H_{MV C}, I.

Annals of the Institute of Statistical Mathematics, 19(1):451–471, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911697.

Srivastava:1967:CMN

[464] M. S. Srivastava. Classification into multivariate normal populations when the population means are linearly restricted. Annals of the Institute of Statistical Mathematics, 19(1): 473–478, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911698.

Kazi:1967:KES

[465] Kimio Kazi. Kolmogorov’s ϵ-entropy of some Gaussian processes. Annals of the Institute of Statistical Mathematics, 19(1):479–503, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911699.

Smith:1967:CTB

[466] Armand V. Smith. Comparison of two Bernoulli processes by multiple stage sampling using Bayesian decision theory. Annals of the Institute of Statistical Mathematics, 19(1):505–518, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911700.

Gibbons:1967:CCB

[467] Jean D. Gibbons. Correlation coefficients between nonparametric tests for location and scale. Annals of the Institute of Statistical Mathematics, 19 (1):519–526, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911701.

Neuts:1967:MDW

[468] M. F. Neuts and S. Zacks. On mixtures of χ^{2}- and F -distributions which yield distributions of the same family. Annals of the Institute of Statistical Mathematics, 19(1):527–536, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911702.

Meyer:1967:NSP

[469] Paul L. Meyer. A note on the sum of Poisson probabilities and an application. Annals of the Institute of Statistical Mathematics, 19(1):537–542, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911703.

Anonymous:1967:HC

[470] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 19(1):??, December 1967. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takahasi:1968:UEP

[471] Koiti Takahasi and Kazumasa Wakimoto. On unbiased estimates of the population mean based on the sample stratified by means of ordering. Annals of the Institute of Statistical Mathematics, 20(1):1–31, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911622.

Ishii:1968:MIP

[472] Goro Ishii. Minimax invariant prediction regions. Annals of the Institute of Statistical Mathematics, 20 (1):33–53, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911623.

Akaike:1968:UIB

[473] Hirotugu Akaike. On the use of an index of bias in the estimation of power spectra. Annals of the Institute of Statistical Mathematics, 20(1):55–69, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911624.

Suzuki:1968:TCH

[474] Yukio Suzuki. On testing certain hypotheses. Annals of the Institute of Statistical Mathematics, 20(1):71–78, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911625.

Bartlett:1968:SDF

[475] Noel S. Bartlett and Zakkula Govindarajulu. Some distribution-free statistics and their application to the selection problem. Annals of the Institute of Statistical Mathematics, 20 (1):79–97, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911626.

Puri:1968:MSS

[476] Madan L. Puri. Multi-sample scale problem: Unknown location parameters. Annals of the Institute of Statistical Mathematics, 20(1):99–106, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911627.

Taguchi:1968:CCM

[477] Tokio Taguchi. Concentration-curve methods and structures of skew populations. Annals of the Institute of Statistical Mathematics, 20(1): 107–141, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911628.

Khatri:1968:NEM

[478] C. G. Khatri. A note on exact moments of ARC sine correlation coefficient with the help of characteristic function. Annals of the Institute of Statistical Mathematics, 20(1):143–149, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911629.

Holla:1968:DDP

[479] M. S. Holla. Discrete distributions with prior information. Annals of the Institute of Statistical Mathematics, 20(1):151–157, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911630.

Rao:1968:ASS

[480] T. J. Rao. On the allocation of sample size in stratified sampling. Annals of the Institute of Statistical Mathematics, 20(1):159–166, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911631.

Hayashi:1968:CMQ

[481] Chikio Hayashi. Corrections to “Multidimensional quantification of the data obtained by the method of paired comparison”. Annals of the Institute of Statistical Mathematics, 20(1):167, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02911632. See [345].

Siotani:1968:CSS

[482] Minoru Siotani. Cancellation of section 5 of “Some applications of Loewner’s ordering on symmetric matrices”. Annals of the Institute of Statistical Mathematics, 20(1):168, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02911633. See [444].

Hudimoto:1968:EBP

[483] Hirosi Hudimoto. On the empirical Bayes procedure. Annals of the Institute of Statistical Mathematics, 20 (1):169–185, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

                                        1007/BF02911634. See corrections

[536].

Shimizu:1968:CFS

[484] Ryoichi Shimizu. Characteristic functions satisfying a functional equation (I). Annals of the Institute of Statistical Mathematics, 20(1):187–209, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911635.

Hayashi:1968:REB

[485] Chikio Hayashi. Response errors and biased information. Annals of the Institute of Statistical Mathematics, 20 (1):211–228, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911636.

Govindarajulu:1968:DFC

[486] Zakkula Govindarajulu. Distributionfree confidence bounds for P (X < Y). Annals of the Institute of Statistical Mathematics, 20(1):229–238, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911637.

Bhattacharjee:1968:NNH

[487] G. P. Bhattacharjee. Non-normality and heterogeneity in two sample ttest. Annals of the Institute of Statistical Mathematics, 20(1):239–254, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911638.

Gokhale:1968:ARE

[488] D. V. Gokhale. On asymptotic relative efficiencies of a class of rank tests for independence of two variables. Annals of the Institute of Statistical Mathematics, 20(1):255–261, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911639.

Broemeling:1968:IOC

[489] L. D. Broemeling and H. O. Hartley. Investigations of the optimality of a confidence region for the parameters of a non-linear regression model. Annals of the Institute of Statistical Mathematics, 20(1):263–269, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911640.

Akaike:1968:LPF

[490] Hirotugu Akaike. Low pass filter design. Annals of the Institute of Statistical Mathematics, 20(1):271–297, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911641.

Bhattacharyya:1968:REL

[491] G. K. Bhattacharyya. Robust estimates of linear trend in multivariate time series. Annals of the Institute of Statistical Mathematics, 20(1): 299–310, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911642.

Rao:1968:PLL

[492] B. L. S. Prakasa Rao and Herman Rubin. A property of the log-likelihood-ratio process for Gaussian processes. Annals of the Institute of Statistical Mathematics, 20(1):311–314, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911643.

Heyde:1968:GRW

[493] C. C. Heyde. On the growth of a random walk. Annals of the Institute of Statistical Mathematics, 20(1): 315–321, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911644.

Griego:1968:NPC

[494] R. J. Griego. A note on projections of continuous additive functionals. Annals of the Institute of Statistical Mathematics, 20(1):323–326, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911645.

Dharmadhikari:1968:URC

[495] S. W. Dharmadhikari. A uniqueness result for compound normal and compound exponential distributions. Annals of the Institute of Statistical Mathematics, 20(1):327–329, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911646.

Holla:1968:CGD

[496] M. S. Holla and S. K. Bhattacharya. On a compound Gaussian distribution. Annals of the Institute of Statistical Mathematics, 20(1):331–336, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911647.

Das:1968:CGD

[497] M. N. Das and A. Dey. Corrections to “Group divisible rotatable designs”. Annals of the Institute of Statistical Mathematics, 20(1):337, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02911648. See [450].

Ikeda:1968:AER

[498] Sadao Ikeda. Asymptotic equivalence of real probability distributions. Annals of the Institute of Statistical Mathematics, 20(1):339–362, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911649.

Godambe:1968:BSS

[499] V. P. Godambe. Bayesian sufficiency in survey-sampling. Annals of the Institute of Statistical Mathematics, 20(1):363–373, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911650.

Bhattacharya:1968:BAC

[500] S. K. Bhattacharya. Bayes approach to compound distributions arising from truncated mixing densities. Annals of the Institute of Statistical Mathematics, 20(1):375–381, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911651.

Laha:1968:PWP

[501] R. G. Laha and E. Lukacs. On a property of the Wiener process. Annals of the Institute of Statistical Mathematics, 20(1):383–389, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911652.

Imai:1968:NLL

[502] Haruo Imai. Notes on a local limit theorem for discrete time Galton– Watson branching processes. Annals of the Institute of Statistical Mathematics, 20(1):391–410, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911653.

Moore:1968:ETD

[503] Erin H. Moore and Ronald Pyke. Estimation of the transition distributions of a Markov renewal process. Annals of the Institute of Statistical Mathematics, 20(1):411–424, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911654.

Akaike:1968:ULM

[504] Hirotugu Akaike. On the use of a linear model for the identification of feedback systems. Annals of the Institute of Statistical Mathematics, 20 (1):425–439, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911655.

Talukder:1968:MRE

[505] Anwar Hossain Talukder. Multivariate regression estimates for finite populations. Annals of the Institute of Statistical Mathematics, 20(1):441–455, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911656.

Roy:1968:EGL

[506] A. R. Roy and V. K. Srivastava. On the estimation of generalized linear probability model involving discrete random variables. Annals of the Institute of Statistical Mathematics, 20(1): 457–467, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911657.

Kshirsagar:1968:ABI

[507] A. M. Kshirsagar and P. S. Simha. Analysis of a balanced incomplete twoway design. Annals of the Institute of Statistical Mathematics, 20(1): 469–476, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02911658.

Dey:1968:GDR

[508] A. Dey and A. K. Nigam. Group divisible rotatable designs-some further considerations. Annals of the Institute of Statistical Mathematics, 20(1): 477–481, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02911659.

Rao:1968:NSM

[509] P. V. Rao. A note on a ksample model of Conover. Annals of the Institute of Statistical Mathematics, 20(1):483–487, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911660.

Chu:1968:SSM

[510] J. T. Chu. Some statistical methods for large scale and preliminary data analyses. Annals of the Institute of Statistical Mathematics, 20(1):489–499, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911661.

Rohatgi:1968:RCR

[511] V. K. Rohatgi. On the rate of convergence of the range of cumulative sums. Annals of the Institute of Statistical Mathematics, 20(1):501–503, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02911662.

Sugiura:1968:NCI

[512] Nariaki Sugiura and Masanori Ôtake. Numerical comparison of improved methods of testing in contingency tables with small frequencies. Annals of the Institute of Statistical Mathematics, 20(1):505–517, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911663.

Thomasson:1968:EPT

[513] R. L. Thomasson and C. H. Kapadia. On estimating the parameter of a truncated geometric distribution. Annals of the Institute of Statistical Mathematics, 20(1):519–523, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02911664.

Anonymous:1968:HC

[514] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 20(1):??, December 1968. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayakawa:1969:DLR

[515] Takesi Hayakawa. On the distribution of the latent roots of a positive definite random symmetric matrix. I. Annals of the Institute of Statistical Mathematics, 21(1):1–21, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532228.

Khatri:1969:NCD

[516] C. G. Khatri. Non-central distributions of i-th largest characteristic roots of three matrices concerning complex multivariate normal populations. Annals of the Institute of Statistical Mathematics, 21(1):23–32, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532229.

Gleser:1969:TEM

[517] Leon Jay Gleser and Ingram Olkin. Testing for equality of means, equality of variances, and equality of covariances under restrictions upon the parameter space. Annals of the Institute of Statistical Mathematics, 21 (1):33–48, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532230.

Pillai:1969:PCT

[518] K. C. Sreeharan Pillai and Charles O. Dotson. Power comparisons of tests of two multivariate hypotheses based on individual characteristic roots. Annals of the Institute of Statistical Mathematics, 21(1):49–66, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532231.

Asoh:1969:NNN

[519] Y. H. Asoh and Masashi Okamoto. A note on the non-null distribution of the Wilks statistic in MANOVA. Annals of the Institute of Statistical Mathematics, 21(1):67–71, ????

1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532232.

Roussas:1969:NEM

[520] George G. Roussas. Nonparametric estimation in Markov processes. Annals of the Institute of Statistical Mathematics, 21(1):73–87, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532233.

Uematu:1969:SMQ

[521] Tosio Uematu. On some model of queueing system with state-dependent service time distributions. Annals of the Institute of Statistical Mathematics, 21(1):89–106, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532234.

Choi:1969:EPF

[522] Keewhan Choi. Estimators for the parameters of a finite mixture of distributions. Annals of the Institute of Statistical Mathematics, 21 (1):107–116, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532235. See corrections

[606].

Choi:1969:EBP

[523] Keewhan Choi. Empirical Bayes procedure for (pattern) classification with stochastic learning. Annals of the Institute of Statistical Mathematics, 21 (1):117–125, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532236. See [607].

Jaiswal:1969:PFL

[524] M. C. Jaiswal and C. G. Khatri. Power function of the likelihood ratio test when range depends upon the parameter. Annals of the Institute of Statistical Mathematics, 21(1):127–136, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532237.

Khatri:1969:TEP

[525] C. G. Khatri and M. C. Jaiswal. On testing the equality of parameters in k triangular populations with unequal observations. Annals of the Institute of Statistical Mathematics, 21(1):137–148, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532238.

Guttman:1969:PBP

[526] Irwin Guttman and Roy C. Milton. Procedures for a best population problem when the criterion of bestness involves a fixed tolerance region. Annals of the Institute of Statistical Mathematics, 21(1):149–161, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532239.

Puri:1969:ATR

[527] Madan Lal Puri and Pranab Kumar Sen. On the asymptotic theory of rank order tests for experiments involving paired comparisons. Annals of the Institute of Statistical Mathematics, 21(1):163–173, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532240.

Matubara:1969:EPM

[528] Nozomu Matubara. On ergodic probability measures. Annals of the Institute of Statistical Mathematics, 21 (1):175–183, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532241.

Kubokawa:1969:FII

[529] Yoshihiro Kubokawa. Finite and infinite invariant measures for a measurable transformation. Annals of the Institute of Statistical Mathematics, 21 (1):185–193, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532242.

Kubokawa:1969:RFI

[530] Yoshihiro Kubokawa. Remarks on finite invariant measures for oneparameter group of measurable transformations. Annals of the Institute of Statistical Mathematics, 21(1):195–200, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532243.

Kotz:1969:DPQ

[531] Samuel Kotz and R. Srinivasan. Distribution of product and quotient of Bessel function variates. Annals of the Institute of Statistical Mathematics, 21(1):201–210, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532244.

Blum:1969:FPE

[532] J. R. Blum and Judah Rosenblatt. Fixed precision estimation in the class of IFR distribution. Annals of the Institute of Statistical Mathematics, 21 (1):211–213, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532245.

Brillinger:1969:CCC

[533] David R. Brillinger. The calculation of cumulants via conditioning. Annals of the Institute of Statistical Mathematics, 21(1):215–218, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532246.

Krafft:1969:NEB

[534] Olaf Krafft. A note on exponential bounds for binomial probabilities. Annals of the Institute of Statistical Mathematics, 21(1):219–220, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02532247.

Hayakawa:1969:DML

[535] Takesi Hayakawa. On the distribution of the maximum latent root of a positive definite symmetric random matrix. Annals of the Institute of Statistical Mathematics, 21(1):221, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02532248.

Hudimoto:1969:CEB

[536] Hirosi Hudimoto. Corrections to “On the empirical Bayes procedure (1)”. Annals of the Institute of Statistical Mathematics, 21(1):223, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02532249. See [483].

Akaike:1969:MSI

[537] Hirotugu Akaike. A method of statistical identification of discrete time parameter linear systems. Annals of the Institute of Statistical Mathematics, 21(1):225–242, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532250.

Akaike:1969:FAM

[538] Hirotugu Akaike. Fitting autoregressive models for prediction. Annals of the Institute of Statistical Mathematics, 21(1):243–247, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532251.

Takahasi:1969:EPM

[539] Koiti Takahasi. On the estimation of the population mean based on ordered samples from an equicorrelated multivariate distribution. Annals of the Institute of Statistical Mathematics, 21(1):249–255, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532252.

Weiss:1969:AJD

[540] Lionel Weiss. The asymptotic joint distribution of an increasing number of sample quantiles. Annals of the Institute of Statistical Mathematics, 21 (1):257–263, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532253.

Saxena:1969:MRO

[541] K. M. Lal Saxena and I. R. Savage. Monotonicity of rank order likelihood ratio. Annals of the Institute of Statistical Mathematics, 21(1):265–275, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532254.

Mudholkar:1969:GMC

[542] Govind S. Mudholkar. A generalized monotone character of d.f.’s and moments of statistics from some wellknown populations. Annals of the Institute of Statistical Mathematics, 21 (1):277–285, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532255.

Walsh:1969:AIB

[543] John E. Walsh. Asymptotic independence between largest and smallest of a set of independent observations. Annals of the Institute of Statistical Mathematics, 21(1):287–289, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532256.

Zidek:1969:RBI

[544] James V. Zidek. A representation of Bayes invariant procedures in terms of Haar measure. Annals of the Institute of Statistical Mathematics, 21 (1):291–308, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532257.

Pillai:1969:MES

[545] K. C. Sreedharan Pillai and Gary M. Jouris. On the moments of elementary symmetric functions of the roots of two matrices. Annals of the Institute of Statistical Mathematics, 21(1):309–320, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532258.

Pillai:1969:NCD

[546] K. C. S. Pillai and T. Sugiyama. Noncentral distributions of the largest latent roots of three matrices in multivariate analysis. Annals of the Institute of Statistical Mathematics, 21 (1):321–327, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532259.

Sen:1969:NMM

[547] Pranab Kumar Sen. On nonparametric T -method of multiple, comparisons for randomized blocks. Annals of the Institute of Statistical Mathematics, 21 (1):329–333, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532260.

Raktoe:1969:SNO

[548] B. L. Raktoe and W. T. Federer. Some non-orthogonal unsaturated main effect and resolution V plans derived from a one-restrictional lattice. Annals of the Institute of Statistical Mathematics, 21(1):335–342, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532261.

Dey:1969:NWD

[549] A. Dey. A note on weighing designs. Annals of the Institute of Statistical Mathematics, 21(1):343–346, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532262.

Kazi:1969:EDP

[550] Kimio Kazi. On the ϵ-entropy of diffusion processes. Annals of the Institute of Statistical Mathematics, 21 (1):347–356, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532263.

Makino:1969:IMW

[551] Toji Makino. Investigation of the mean waiting time for queueing system with many servers. Annals of the Institute of Statistical Mathematics, 21(1):357–366, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532264.

Shanbhag:1969:QSS

[552] D. N. Shanbhag. A queueing system with several types of customers. Annals of the Institute of Statistical Mathematics, 21(1):367–371, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532265.

Gun:1969:SLP

[553] A. M. Gun. On the significance level of preliminary tests in some “TE” procedures. Annals of the Institute of Statistical Mathematics, 21(1):373–376, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532266.

Inada:1969:SMG

[554] Ken ichi Inada. On the stability of multisectoral growth equilibrium. Annals of the Institute of Statistical Mathematics, 21(1):377–390, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532267.

Shimizu:1969:CFS

[555] Ryoichi Shimizu. Characteristic functions satisfying a functional equation (II). Annals of the Institute of Statistical Mathematics, 21(1):391–405, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL http://link.springer.com/
article/10.1007/BF02532268. See

corrections [588].

Akaike:1969:PSE

[556] Hirotugu Akaike. Power spectrum estimation through autoregressive model fitting. Annals of the Institute of Statistical Mathematics, 21(1):407–419, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532269.

Barankin:1969:TMG

[557] Edward W. Barankin. Toward the mathematics of a general theory of behavior, 1. The lattice 0[Hebrew letter tsade]. Annals of the Institute of Statistical Mathematics, 21 (1):421–456, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532270.

Alam:1969:LAR

[558] Khursheed Alam and James R. Thompson. Locally averaged risk. Annals of the Institute of Statistical Mathematics, 21(1):457–469, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532271.

Ishii:1969:OUP

[559] Goro Ishii. Optimality of unbiased predictors. Annals of the Institute of Statistical Mathematics, 21(1):471–488, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532272.

Yanagimoto:1969:POP

[560] Takemi Yanagimoto and Masashi Okamoto. Partial orderings of permutations and monotonicity of a rank correlation statistic. Annals of the Institute of Statistical Mathematics, 21 (1):489–506, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532273.

Srivastava:1969:COH

[561] J. N. Srivastava and L. L. McDonald. On the costwise optimality of hierarchical multiresponse randomized block designs under the trace criterion. Annals of the Institute of Statistical Mathematics, 21(1):507–514, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532274.

Singh:1969:OS

[562] Ravindra Singh and B. V. Sukhatme. Optimum stratification. Annals of the Institute of Statistical Mathematics, 21 (1):515–528, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532275.

Pakshirajan:1969:CNL

[563] R. P. Pakshirajan and N. R. Mohan. A characterization of the normal law. Annals of the Institute of Statistical Mathematics, 21(1):529–532, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532276.

Ikeda:1969:RIT

[564] Sadao Ikeda. A remark on the incomparability of two criteria for a uniform convergence of probability measures. Annals of the Institute of Statistical Mathematics, 21(1):533–536, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532277.

Kubokawa:1969:BMT

[565] Yoshihiro Kubokawa. Boundedness of a measurable transformation and a weakly wandering set. Annals of the Institute of Statistical Mathematics, 21(1):537–540, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532278.

Skibinsky:1969:SKR

[566] Morris Skibinsky. Some known results concerning zero-one sets. Annals of the Institute of Statistical Mathematics, 21(1):541–545, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02532279.

Lientz:1969:NLD

[567] B. P. Lientz. A note on limiting distributions of some Rényi-type statistics. Annals of the Institute of Statistical Mathematics, 21(1):547–550, ????

1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532280.

Kabe:1969:SDP

[568] D. G. Kabe. Some distribution problems of order statistics from discrete populations. Annals of the Institute of Statistical Mathematics, 21(1):551–556, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532281.

Anonymous:1969:HC

[569] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 21(1):??, ???? 1969. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matusita:1970:ZS

[570] Kameo Matusita. Zyoiti Suetuna, 1898–1970. Annals of the Institute of Statistical Mathematics, 22(1):v–vi, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02506317.

Inagaki:1970:LDS

[571] Nobuo Inagaki. On the limiting distribution of a sequence of estimators with uniformity property. Annals of the Institute of Statistical Mathematics, 22(1):1–13, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506318.

Wakimoto:1970:UEPa

[572] Kazumasa Wakimoto. On unbiased estimation of the population variance based on the stratified random sample. Annals of the Institute of Statistical Mathematics, 22(1):15–26, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL http://link.springer.com/
article/10.1007/BF02506319. See

correction [609].

Zacks:1970:BEE

[573] S. Zacks. Bayes equivariant estimators of variance components. Annals of the Institute of Statistical Mathematics, 22(1):27–40, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506320.

Moore:1970:ANE

[574] D. S. Moore. Asymptotically nearly efficient procedures for bivariate location parameters. Annals of the Institute of Statistical Mathematics, 22(1):41–49, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506321.

Ghosh:1970:ERS

[575] J. K. Ghosh and Rajinder Singh. Estimation of the reciprocal of scale parameter of a gamma density. Annals of the Institute of Statistical Mathematics, 22(1):51–55, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506322.

Yanagimoto:1970:MAR

[576] Takemi Yanagimoto. On measures of association and a related problem. Annals of the Institute of Statistical Mathematics, 22(1):57–63, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506323.

Hoel:1970:SMA

[577] David G. Hoel. Some modifications and applications of Wald’s OC formula. Annals of the Institute of Statistical Mathematics, 22(1):65–75, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506324.

Davis:1970:FAD

[578] A. W. Davis. Further applications of a differential equation for Hotelling’s generalized T_{0}^{2}. Annals of the Institute of Statistical Mathematics, 22 (1):77–87, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506325.

Mathai:1970:EDV

[579] A. M. Mathai and P. N. Rathie. The exact distribution of Votaw’s criteria. Annals of the Institute of Statistical Mathematics, 22(1):89–116, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506326.

Han:1970:DDF

[580] Chien-Pai Han. Distribution of discriminant function in circular models. Annals of the Institute of Statistical Mathematics, 22(1):117–125, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506327.

Alam:1970:TSP

[581] Khursheed Alam. A two-sample procedure for selecting the population with the largest mean from k normal populations. Annals of the Institute of Statistical Mathematics, 22(1):127–136, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506328.

Smith:1970:CMT

[582] Armand V. Smith, Jr. Comparison of the means of two normal processes by multiple stage sampling using Bayesian decision theory. Annals of the Institute of Statistical Mathematics, 22 (1):137–143, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506329.

Kulkarni:1970:LAM

[583] S. R. Kulkarni. Locally asymptotically most powerful tests about the effects of K treatments. Annals of the Institute of Statistical Mathematics, 22(1):145–158, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

1007/BF02506330. See comment [608].

Saha:1970:NOM

[584] G. M. Saha and S. Mohanty. On nonorthogonal main effect plans for asymmetrical factorials. Annals of the Institute of Statistical Mathematics, 22 (1):159–169, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506331.

Sclove:1970:AML

[585] Stanley L. Sclove. Admissibility of the maximum likelihood estimator in the regression of two predictands on one predictor. Annals of the Institute of Statistical Mathematics, 22(1):171–174, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506332.

Asoo:1970:NES

[586] Yasuhiro Asoo. Note on the estimation of the standardized covariance matrix. Annals of the Institute of Statistical Mathematics, 22(1):175–179, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506333.

Shantaram:1970:LTC

[587] R. Shantaram and William L. Harkness. A limit theorem for a certain transform of sums of independent random variables. Annals of the Institute of Statistical Mathematics, 22 (1):181–184, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506334.

Shimizu:1970:CCF

[588] Ryoichi Shimizu. Corrections to “Characteristic functions satisfying a functional equation (II)”. Annals of the Institute of Statistical Mathematics, 22 (1):185–186, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02506335. See

[555].

Barankin:1970:TMG

[589] Edward W. Barankin. Toward the mathematics of a general theory of behavior, II. Annals of the Institute of Statistical Mathematics, 22(1):187–202, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506336.

Akaike:1970:SPI

[590] Hirotugu Akaike. Statistical predictor identification. Annals of the Institute of Statistical Mathematics, 22(1):203–217, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506337.

Akaike:1970:FRB

[591] Hirotugu Akaike. A fundamental relation between predictor identification and power spectrum estimation. Annals of the Institute of Statistical Mathematics, 22(1):219–223, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506338.

Weiss:1970:MPE

[592] L. Weiss and J. Wolfowitz. Maximum probability estimators and asymptotic sufficiency. Annals of the Institute of Statistical Mathematics, 22(1):225–244, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506339.

Shimizu:1970:DPA

[593] Ryoichi Shimizu. On the domain of partial attraction of semi-stable distributions. Annals of the Institute of Statistical Mathematics, 22(1):245–255, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506340.

Csorgo:1970:DRP

[594] Miklós Csörgö and Mayer Alvo. Distribution results and power functions for Kac statistics. Annals of the Institute of Statistical Mathematics, 22 (1):257–260, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506341.

Blumenthal:1970:TFB

[595] Saul Blumenthal. Tests of fit based on partial sums of the ordered spacings. Annals of the Institute of Statistical Mathematics, 22(1):261–276, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506342.

Sen:1970:REC

[596] Pranab Kumar Sen. On the robustefficiency of the combination of independent nonparametric tests. Annals of the Institute of Statistical Mathematics, 22(1):277–280, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506343.

Sen:1970:NIR

[597] Pranab Kumar Sen. Nonparametric inference in n replicated 2^{m} factorial experiments. Annals of the Institute of Statistical Mathematics, 22(1):281–294, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506344.

Kshirsagar:1970:GFA

[598] A. M. Kshirsagar. Goodness of fit of an assigned set of scores for the analysis of association in a contingency table. Annals of the Institute of Statistical Mathematics, 22(1):295–306, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506345.

Pillai:1970:MPF

[599] K. C. S. Pillai and Hung C. Li. Monotonicity of the power functions of some tests of hypotheses concerning multivariate complex normal distributions. Annals of the Institute of Statistical Mathematics, 22(1):307–318, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506346.

Sclove:1970:SRN

[600] Stanley L. Sclove. Some remarks on normal multivariate regression. Annals of the Institute of Statistical Mathematics, 22(1):319–326, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506347.

Serfling:1970:VFE

[601] R. J. Serfling. The variance function of the Erlang process. Annals of the Institute of Statistical Mathematics, 22(1):327–337, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506348.

Takacs:1970:FIT

[602] Lajos Takács. A fundamental identity in the theory of queues. Annals of the Institute of Statistical Mathematics, 22(1):339–348, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506349.

Takamatsu:1970:QPA

[603] Shunro Takamatsu. Queueing processes with accumulated service. Annals of the Institute of Statistical Mathematics, 22(1):349–379, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506350.

Dey:1970:MEP

[604] A. Dey and G. M. Saha. Main effect plans for k^{n} factorials with blocks.

Annals of the Institute of Statistical Mathematics, 22(1):381–388, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506351.

Dey:1970:CBA

[605] A. Dey. On construction of balanced n-ary block designs. Annals of the Institute of Statistical Mathematics, 22 (1):389–393, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506352.

Choi:1970:CEP

[606] Keewhan Choi. Corrections for “Estimators for the parameters of a finite mixture of distributions”. Annals of the Institute of Statistical Mathematics, 22(1):395–396, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02506353. See

[522].

Choi:1970:CEB

[607] Keewhan Choi. Corrections for “Empirical Bayes procedure for (pattern) classification with stochastic learning”. Annals of the Institute of Statistical Mathematics, 22(1):397–398, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02506354. See [523].

Kulkarni:1970:CPL

[608] S. R. Kulkarni. A comment on the paper “Locally asymptotically most powerful tests about the effects of K treatments”. Annals of the Institute of Statistical Mathematics, 22(1):399, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02506355. See [583].

Wakimoto:1970:CUE

[609] Kazumasa Wakimoto. Correction to “On unbiased estimation of the population variance based on the stratified random sample”. Annals of the Institute of Statistical Mathematics, 22(1):400, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02506356. See

[572].

Talukder:1970:COA

[610] M. A. H. Talukder. Corrections to “On the order of approximation of the variance of multivariate regression estimates (MRE) for finite populations”. Annals of the Institute of Statistical Mathematics, 22(1):401, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02506357.

Takahasi:1970:ESC

[611] Koiti Takahasi. Estimation of several characteristics of distributions of order statistics. Annals of the Institute of Statistical Mathematics, 22(1):403–412, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506358.

Takahasi:1970:SNC

[612] Koiti Takahasi. Some nonparametric consistent estimates from censored samples. Annals of the Institute of Statistical Mathematics, 22(1):413–419, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506359.

Takahasi:1970:PNE

[613] Koiti Takahasi. Practical note on estimation of population means based on samples stratified by means of ordering. Annals of the Institute of Statistical Mathematics, 22(1):421–428, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506360.

Wakimoto:1970:UEPb

[614] Kazumasa Wakimoto. On unbiased estimation of the population variance based on the stratified random sample (II). Annals of the Institute of Statistical Mathematics, 22(1):429–433, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506361.

Ikeda:1970:AIO

[615] Sadao Ikeda and Tadashi Matsunawa. On asymptotic independence of order statistics. Annals of the Institute of Statistical Mathematics, 22(1):435–449, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506362.

Khatri:1970:FCS

[616] C. G. Khatri. Further contributions to some inequalities for normal distributions and their applications to simultaneous confidence bounds. Annals of the Institute of Statistical Mathematics, 22(1):451–458, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506363.

Raghavachari:1970:MST

[617] M. Raghavachari. Multi-sample tests for scale. Annals of the Institute of Statistical Mathematics, 22(1):459–464, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506364.

Milch:1970:JPS

[618] Paul R. Milch. The Jacobi polynomial and some hypergeometric type distributions. Annals of the Institute of Statistical Mathematics, 22(1):465–474, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506365.

Kabe:1970:SDO

[619] D. G. Kabe. Some distributions of ordered random intervals with applications. Annals of the Institute of Statistical Mathematics, 22(1):475–481, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506366.

Srivastava:1970:EES

[620] V. K. Srivastava. The efficiency of estimating seemingly unrelated regression equations. Annals of the Institute of Statistical Mathematics, 22(1):483–493, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506367.

Sharma:1970:ECL

[621] S. S. Sharma. On an estimator in T_{3}- class of linear estimators in sampling with varying probabilities from a finite population. Annals of the Institute of Statistical Mathematics, 22(1):495–500, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506368.

Raghavarao:1970:SRT

[622] Damaraju Raghavarao. Some results on tactical configurations and non-existence of difference set solutions for certain symmetrical PBIB designs. Annals of the Institute of Statistical Mathematics, 22(1):501–506, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506369.

Srivastava:1970:HTR

[623] J. N. Srivastava and L. L. Mc-Donald. On the hierarchical tworesponse (cyclic PBIB) designs, costwise optimal under the trace criterion. Annals of the Institute of Statistical Mathematics, 22(1):507–518, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506370.

Raktoe:1970:LBN

[624] B. L. Raktoe and W. T. Federer. A lower bound for the number of singular staturated main effect plans of an S^{m} factorial. Annals of the Institute of Statistical Mathematics, 22(1):519–525, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506371.

Huzii:1970:VSE

[625] Mituaki Huzii. On the variance of a simplified estimate of correlogram. Annals of the Institute of Statistical Mathematics, 22(1):527–534, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506372.

Patel:1970:COS

[626] H. I. Patel. The choice of optimum scores in a Markov chain of order one. Annals of the Institute of Statistical Mathematics, 22(1):535–538, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506373.

Rohatgi:1970:SL

[627] V. K. Rohatgi. On the strong law. Annals of the Institute of Statistical Mathematics, 22(1):539–541, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506374.

Sibuya:1970:SGI

[628] Masaaki Sibuya. Subclasses of generalized inverses of matrices. Annals of the Institute of Statistical Mathematics, 22(1):543–556, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02506375.

Anonymous:1970:HC

[629] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 22(1):??, ???? 1970. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Noda:1971:MEM

[630] Kazuo Noda and Yasushi Taga. Minimax estimation method for the optimum decomposition of a sample space based on prior information. Annals of the Institute of Statistical Mathematics, 23(1):1–29, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479208.

Hayashi:1971:RRA

[631] Chikio Hayashi. Response reliability and attitude change-supplement to response errors and biased information. Annals of the Institute of Statistical Mathematics, 23(1):31–34, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479209.

Barankin:1971:TMG

[632] Edward W. Barankin. Toward the mathematics of a general theory of behavior, III the flanking heredity theorem. Annals of the Institute of Statistical Mathematics, 23(1):35–65, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479210.

Guttman:1971:BAP

[633] Irwin Guttman. A Bayesian analogue of Paulson’s lemma and its use in tolerance region construction when sampling from the multi-variate normal. Annals of the Institute of Statistical Mathematics, 23(1):67–76, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479211.

Gupta:1971:DWL

[634] A. K. Gupta. Distribution of Wilks’ likelihood-ratio criterion in the complex case. Annals of the Institute of Statistical Mathematics, 23(1):77–87, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479212.

Pillai:1971:ADL

[635] K. C. S. Pillai and D. L. Young. An approximation to the distribution of the largest root of a complex Wishart matrix. Annals of the Institute of Statistical Mathematics, 23(1):89–96, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479213.

Gupta:1971:DCF

[636] R. P. Gupta and D. G. Kabe. Distribution of certain factors useful in discriminant analysis. Annals of the Institute of Statistical Mathematics, 23(1):97–103, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479214.

Church:1971:NEM

[637] J. D. Church and E. Benton Cobb. Nonparametric estimation of the mean using quantal response data. Annals of the Institute of Statistical Mathematics, 23(1):105–117, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479215.

Dixon:1971:BSP

[638] D. O. Dixon and R. P. Bland. A Bayes solution for the problem of ranking Poisson parameters. Annals of the Institute of Statistical Mathematics, 23(1):119–124, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479216.

Harris:1971:LDS

[639] B. Harris and C. J. Park. The limiting distribution of the sample occupancy numbers from the multinomial distribution with equal cell probabilities. Annals of the Institute of Statistical Mathematics, 23(1):125–133, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479217.

Patel:1971:ACO

[640] H. I. Patel. Acknowledgement to “The choice of optimum scores in a Markov chain of order one”. Annals of the Institute of Statistical Mathematics, 23(1):135, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02479218.

Matusita:1971:SPA

[641] Kameo Matusita. Some properties of affinity and applications. Annals of the Institute of Statistical Mathematics, 23(1):137–155, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479219.

Kirmani:1971:SLP

[642] S. N. U. A. Kirmani. Some limiting properties of Matusita’s measure of distance. Annals of the Institute of Statistical Mathematics, 23(1):157–162, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479220.

Akaike:1971:AMF

[643] Hirotugu Akaike. Autoregressive model fitting for control. Annals of the Institute of Statistical Mathematics, 23 (1):163–180, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479221.  See correction

[674].

Mathai:1971:DLR

[644] A. M. Mathai. On the distribution of the likelihood ratio criterion for testing linear hypotheses on regression coefficients. Annals of the Institute of Statistical Mathematics, 23(1):181–197, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479222.

Jaiswal:1971:CTM

[645] M. C. Jaiswal and C. G. Khatri. On certain tests and monotonicity of their power for the parameters involved in the non-regular density functions. Annals of the Institute of Statistical Mathematics, 23(1):199–210, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479223.

Mehta:1971:PDE

[646] J. S. Mehta and R. Srinivasan. On pooling data i: Estimation of the mean. Annals of the Institute of Statistical Mathematics, 23(1):211–224, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479224.

Chan:1971:SAP

[647] Lai K. Chan. Some asymptotic properties of the linearized maximum likelihood estimate and best linear unbiased estimate. Annals of the Institute of Statistical Mathematics, 23(1):225–232, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479225.

Wakimoto:1971:SRSa

[648] Kazumasa Wakimoto. Stratified random sampling (1) estimation of the population variance. Annals of the Institute of Statistical Mathematics, 23 (1):233–252, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479226.

Alam:1971:SPB

[649] Khursheed Alam and James R. Thompson. A selection procedure based on ranks. Annals of the Institute of Statistical Mathematics, 23(1):253–262, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479227.

Kshirsager:1971:RIR

[650] A. M. Kshirsager. Recovery of interrow and inter-column information in two-way designs. Annals of the Institute of Statistical Mathematics, 23 (1):263–278, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479228.

Gulati:1971:MS

[651] Bodh Raj Gulati. On maximal (k, t)- sets. Annals of the Institute of Statistical Mathematics, 23(1):279–292, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL http://link.springer.com/
article/10.1007/BF02479229. See

correction [673].

Krishnaiah:1971:EDS

[652] P. R. Krishnaiah and T. C. Chang. On the exact distribution of the smallest root of the Wishart matrix using zonal polynomials. Annals of the Institute of Statistical Mathematics, 23(1):293–295, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479230.

Abe:1971:SAR

[653] Shun ichi Abe. Statistical analysis of reliability data in renewal processes. Annals of the Institute of Statistical Mathematics, 23(1):297–320, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479231.

Holgate:1971:CPE

[654] P. Holgate. On a class of psychological experiments. Annals of the Institute of Statistical Mathematics, 23(1):321–325, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479232.

Wakimoto:1971:SRSb

[655] Kazumasa Wakimoto. Stratified random sampling (II) estimation of the population covariance. Annals of the Institute of Statistical Mathematics, 23 (1):327–337, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479233.

Wakimoto:1971:SRSc

[656] Kazumasa Wakimoto. Stratified random sampling (III) estimation of the correlation coefficient. Annals of the Institute of Statistical Mathematics, 23 (1):339–353, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479234.

Taga:1971:COS

[657] Yasushi Taga. On the convergence of optimum stratifications for empiric distribution function in univariate case. Annals of the Institute of Statistical Mathematics, 23(1):355–363, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479235.

Alam:1971:SSR

[658] Khursheed Alam, Kenzo Seo, and James R. Thompson. A sequential sampling rule for selecting the most probable multinomial event. Annals of the Institute of Statistical Mathematics, 23(1):365–374, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479236.

Bhapkar:1971:SSP

[659] V. P. Bhapkar and A. P. Gore. Some selection procedures based on U-statistics for the location and scale problems. Annals of the Institute of Statistical Mathematics, 23(1):375–386, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479237.

Jones:1971:SEM

[660] Richard H. Jones. Spectrum estimation with missing observations. Annals of the Institute of Statistical Mathematics, 23(1):387–398, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479238.

Huzii:1971:NEC

[661] Mituaki Huzii. Note on the estimation of correlogram by using transformed variables. Annals of the Institute of Statistical Mathematics, 23(1):399–410, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479239.

Alam:1971:SPP

[662] Khursheed Alam. Selection from Poisson processes. Annals of the Institute of Statistical Mathematics, 23(1):411–418, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479240.

Sedransk:1971:EPM

[663] N. Sedransk and Masashi Okamoto. Estimation of the probabilities of misclassification for a linear discriminant function in the univariate normal case. Annals of the Institute of Statistical Mathematics, 23(1):419–435, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479241.

Srivastava:1971:DVE

[664] V. K. Srivastava. Disturbance variance estimation in simultaneous equations by k-class method. Annals of the Institute of Statistical Mathematics, 23(1):437–449, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479242.

Krishnaiah:1971:STE

[665] P. R. Krishnaiah and V. B. Waikar. Simultaneous tests for equality of latent roots against certain alternativesi. Annals of the Institute of Statistical Mathematics, 23(1):451–468, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479243.

Sugiura:1971:AED

[666] Nariaki Sugiura and Hisao Nagao. Asymptotic expansion of the distribution of the generalized variance for noncentral Wishart matrix, when Ω = O(n). Annals of the Institute of Statistical Mathematics, 23(1):469–475, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479244.

Fujikoshi:1971:AEN

[667] Yasunori Fujikoshi. Asymptotic expansions of the non-null distributions of two criteria for the linear hypothesis concerning complex multivariate normal populations. Annals of the Institute of Statistical Mathematics, 23 (1):477–490, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479245.

Kulshreshtha:1971:CD

[668] A. C. Kulshreshtha, G. M. Saha, and A. Dey. On circular designs. Annals of the Institute of Statistical Mathematics, 23(1):491–497, ????

1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479246.

Saha:1971:CTR

[669] G. M. Saha and A. K. Mishra. A class of three-replicate three-associate p.b.i.b. designs. Annals of the Institute of Statistical Mathematics, 23 (1):499–505, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479247.

Harris:1971:NAN

[670] B. Harris and C. J. Park. A note on the asymptotic normality of the distribution of the number of empty cells in occupancy problems. Annals of the Institute of Statistical Mathematics, 23(1):507–513, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479248.

Nagasaka:1971:HDN

[671] Kenji Nagasaka. On Hausdorff dimension of non-normal sets. Annals of the Institute of Statistical Mathematics, 23(1):515–521, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479249.

Alam:1971:CN

[672] Khursheed Alam. A characterization of normality. Annals of the Institute of Statistical Mathematics, 23(1):523–525, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479250.

Gulati:1971:CNM

[673] Bodh Raj Gulati. Correction note to “On maximal (k, t)-sets”. Annals of the Institute of Statistical Mathematics, 23(1):527–529, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479251. See [651].

Akaike:1971:CAM

[674] H. Akaike. Correction to “Autoregressive model fitting for control”. Annals of the Institute of Statistical Mathematics, 23(1):531, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02479252. See

[643].

Anonymous:1971:HC

[675] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 23(1):??, ???? 1971. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayakawa:1972:DLR

[676] Takesi Hayakawa. On the distribution of the latent roots of a complex Wishart matrix (non-central case). Annals of the Institute of Statistical Mathematics, 24(1):1–17, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479734.

Hayakawa:1972:DAD

[677] Takesi Hayakawa. On the derivation of the asymptotic distribution of the generalized Hotelling’s T_{0}^{2}. Annals of the Institute of Statistical Mathematics, 24(1):19–32, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479735.

Ikeda:1972:UAJ

[678] Sadao Ikeda and Tadashi Matsunawa. On the uniform asymptotic joint normality of sample quantiles. Annals of the Institute of Statistical Mathematics, 24(1):33–52, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479736.

Mathai:1972:ENC

[679] A. M. Mathai. The exact non-central distribution of the generalized variance. Annals of the Institute of Statistical Mathematics, 24(1):53–65, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479737.

Nagao:1972:NND

[680] Hisao Nagao. Non-null distributions of the likelihood ratio criteria for independence and equality of mean vectors and covariance matrices. Annals of the Institute of Statistical Mathematics, 24(1):67–79, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

1007/BF02479738. See corrections

[842].

Krishnaiah:1972:STE

[681] P. R. Krishnaiah and V. B. Waikar. Simultaneous tests for equality of latent roots against certain alternatives-II. Annals of the Institute of Statistical Mathematics, 24(1):81–85, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479739.

Sugiyama:1972:DLL

[682] T. Sugiyama. Distributions of the largest latent root of the multivariate complex Gaussian distribution. Annals of the Institute of Statistical Mathematics, 24(1):87–94, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479740.

Kabe:1972:GRS

[683] D. G. Kabe. On a generalization of Rao’s u statistic. Annals of the Institute of Statistical Mathematics, 24 (1):95–100, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479741.

Csorgo:1972:DRD

[684] Miklós Csörgő. Distribution results for distance functions based on the modified empirical distribution function of M. Kac. Annals of the Institute of Statistical Mathematics, 24(1):101–110, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479742.

Davies:1972:EMG

[685] M. G. Davies. The expectation of Mahalanobis’ generalized distance. Annals of the Institute of Statistical Mathematics, 24(1):111–125, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479743.

Mudholkar:1972:GPC

[686] Govind S. Mudholkar. G-peakedness comparisons for random vectors. Annals of the Institute of Statistical Mathematics, 24(1):127–135, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479744.

Yanagawa:1972:ESF

[687] Takashi Yanagawa and Kazumasa Wakimoto. Estimation of some functional of the population distribution based on a stratified random sample. Annals of the Institute of Statistical Mathematics, 24(1):137–151, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479745.

Avadhani:1972:CMS

[688] M. S. Avadhani and A. K. Srivastava. A comparison of Midzuno–Sen scheme with P.P.S. sampling without replacement and its application to successive sampling. Annals of the Institute of Statistical Mathematics, 24(1):153–164, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479746.

Govindarajulu:1972:REM

[689] Z. Govindarajulu and Jayant V. Deshpandé. Random effects model: Nonparametric case. Annals of the Institute of Statistical Mathematics, 24 (1):165–170, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479747.

Kumar:1972:TGT

[690] S. Kumar. Trinomial group-testing with an unknown proportion of units in the three categories. Annals of the Institute of Statistical Mathematics, 24(1):171–181, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479748.

Coberly:1972:NOS

[691] W. A. Coberly and T. O. Lewis. A note on a one-sided Kolmogorov–Smirnov test of fit for discrete distribution functions. Annals of the Institute of Statistical Mathematics, 24(1):183–187, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479749.

Davenport:1972:NAI

[692] James M. Davenport. A note on an application of the identity given by Govindarajulu and Suzuki. Annals of the Institute of Statistical Mathematics, 24(1):189–192, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479750.

Shinozaki:1972:NAMa

[693] Nobuo Shinozaki, Masaaki Sibuya, and Kunio Tanabe. Numerical algorithms for the Moore–Penrose inverse of a matrix: Direct methods. Annals of the Institute of Statistical Mathematics, 24(1):193–203, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479751.

Hayakawa:1972:DMQ

[694] Takesi Hayakawa. On the distribution of the multivariate quadratic form in multivariate normal samples. Annals of the Institute of Statistical Mathematics, 24(1):205–230, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479752.

Hayakawa:1972:ADS

[695] Takesi Hayakawa. The asymptotic distributions of the statistics based on the complex Gaussian distribution. Annals of the Institute of Statistical Mathematics, 24(1):231–244, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479753.

Giri:1972:TPC

[696] N. Giri. On testing problems concerning mean of multivariate complex Gaussian distribution. Annals of the Institute of Statistical Mathematics, 24 (1):245–250, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479754.

Hayashi:1972:TDQ

[697] Chikio Hayashi. Two dimensional quantification based on the measure of dissimilarity among three elements. Annals of the Institute of Statistical Mathematics, 24(1):251–257, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479755.

Yanagimoto:1972:SLC

[698] Takemi Yanagimoto and Masaaki Sibuya. Stochastically larger component of a random vector. Annals of the Institute of Statistical Mathematics, 24(1):259–269, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479756.

Tong:1972:CSS

[699] Yung Liang Tong. On the consistency of single-stage ranking procedures. Annals of the Institute of Statistical Mathematics, 24(1):271–284, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479757.

Srivastava:1972:AMP

[700] M. S. Srivastava. Asymptotically most powerful rank tests for regression parameters in MANOVA. Annals of the Institute of Statistical Mathematics, 24 (1):285–297, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479758. See corrections

[1058].

Moore:1972:AEE

[701] D. S. Moore. Asymptotically efficient estimation by local location-parameter approximations. Annals of the Institute of Statistical Mathematics, 24 (1):299–308, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479759.

Yamazoe:1972:ROP

[702] Siro Yamazoe. A random observation process for stochastic approximation. Annals of the Institute of Statistical Mathematics, 24(1):309–317, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479760.

Lin:1972:RCE

[703] Pi-Erh Lin. Rates of convergence in empirical Bayes estimation problems: Discrete case. Annals of the Institute of Statistical Mathematics, 24(1):319–325, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479761.

Muddapur:1972:BEP

[704] M. V. Muddapur. Bayesian estimates of parameters in some queueing models. Annals of the Institute of Statistical Mathematics, 24(1):327–331, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479762.

Rao:1972:MLE

[705] B. L. S. Prakasa Rao. Maximum likelihood estimation for Markov processes. Annals of the Institute of Statistical Mathematics, 24(1):333–345, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479763.

Shimizu:1972:DSC

[706] Ryoichi Shimizu. On the decomposition of stable characteristic functions. Annals of the Institute of Statistical Mathematics, 24(1):347–353, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479764.

Taguchi:1972:TDCa

[707] Tokio Taguchi. On the two-dimensional concentration surface and extensions of concentration coefficient and Pareto distribution to the two dimensional case — I. Annals of the Institute of Statistical Mathematics, 24(1):355–381, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479765.

Kounias:1972:TLS

[708] E. G. Kounias and Bodh Raj Gulati. On two level symmetrical factorial designs. Annals of the Institute of Statistical Mathematics, 24(1):383–403, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479766.

Alam:1972:NMP

[709] Khursheed Alam. A note on a monotonicity property of a rank order probability ratio. Annals of the Institute of Statistical Mathematics, 24(1):405–408, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479767.

Wani:1972:NML

[710] J. K. Wani and D. G. Kabe. Note on a multidimensional linear discriminant function. Annals of the Institute of Statistical Mathematics, 24(1):409–412, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479768.

Raktoe:1972:ALB

[711] B. L. Raktoe and W. T. Federer. Addendum to “A lower bound for the number of singular saturated main effect plans of an S^{M} factorial”. Annals of the Institute of Statistical Mathematics, 24(1):413, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02479769.

Suzuki:1972:DKS

[712] Giitiro Suzuki. Distributions of Kacstatistics. Annals of the Institute of Statistical Mathematics, 24(1):415–421, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479770.

Yanagimoto:1972:TSO

[713] Takemi Yanagimoto and Masaaki Sibuya. Test of symmetry of a onedimensional distribution against positive biasedness. Annals of the Institute of Statistical Mathematics, 24 (1):423–434, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479771.

Bell:1972:CTC

[714] C. B. Bell and Paul J. Smith. Completeness theorems for characterizing distribution-free statistics. Annals of the Institute of Statistical Mathematics, 24(1):435–453, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479772.

Srivastava:1972:SSP

[715] M. S. Srivastava and V. S. Taneja. Some sequential procedures for ranking multivariate normal populations. Annals of the Institute of Statistical Mathematics, 24(1):455–464, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479773.

Sobel:1972:ACS

[716] Milton Sobel and S. P. Yen. An asymptotic comparison of subset selection procedures. Annals of the Institute of Statistical Mathematics, 24(1):465–468, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479774.

Bennett:1972:NRS

[717] B. M. Bennett. Note on the rank sum test of Wilcoxon under trend alternatives. Annals of the Institute of Statistical Mathematics, 24(1):469–472, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479775.

Mathai:1972:CMM

[718] A. M. Mathai and P. N. Rathie. Characterization of Matusita’s measure of affinity. Annals of the Institute of Statistical Mathematics, 24(1):473–483, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479776.

Singh:1972:OSS

[719] Ravindra Singh and B. V. Sukhatme. Optimum stratification in sampling with varying probabilities. Annals of the Institute of Statistical Mathematics, 24(1):485–494, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479777.

Roy:1972:BGD

[720] A. R. Roy and V. K. Srivastava. The bias of generalized double k-class estimators. Annals of the Institute of Statistical Mathematics, 24(1):495–508, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479778.

Hayakawa:1972:NAD

[721] Takesi Hayakawa. Note on the asymptotic distributions of the functions of a multivariate quadratic form in normal sample. Annals of the Institute of Statistical Mathematics, 24(1):509–515, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479779.

Sugiura:1972:ASH

[722] Nariaki Sugiura. Asymptotic solutions of the hypergeometric function_{1} F_{1} of matrix argument, useful in multivariate analysis. Annals of the Institute of Statistical Mathematics, 24(1):517–524, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479780.

Dey:1972:TSP

[723] A. Dey, A. C. Kulshreshtha, and G. M. Saha. Three symbol partially balanced arrays. Annals of the Institute of Statistical Mathematics, 24(1):525–528, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479781.

Barankin:1972:PFB

[724] Edward W. Barankin and Joaquin Curiel. Properties of flanking, beflanking and enflanking. Annals of the Institute of Statistical Mathematics, 24 (1):529–558, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479782.

Yanagimoto:1972:FPD

[725] Takemi Yanagimoto. Families of positively dependent random variables. Annals of the Institute of Statistical Mathematics, 24(1):559–573, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479783.

Imai:1972:CFI

[726] Haruo Imai. Convergence of functional iterates for branching processes. Annals of the Institute of Statistical Mathematics, 24(1):575–587, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479784.

Pakes:1972:GMQ

[727] A. G. Pakes. A GI/ M/ 1 queue with a modified service mechanism. Annals of the Institute of Statistical Mathematics, 24(1):589–597, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479785.

Taguchi:1972:TDCb

[728] Tokio Taguchi. On the two-dimensional concentration surface and extentions of concentration coefficient and Pareto distribution to the two dimensional case — II. Annals of the Institute of Statistical Mathematics, 24(1):599–619, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479786.

Shinozaki:1972:NAMb

[729] Nobuo Shinozaki, Masaaki Sibuya, and Kunio Tanabe. Numerical algorithms for the Moore–Penrose inverse of a matrix: Iterative methods. Annals of the Institute of Statistical Mathematics, 24(1):621–629, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479787.

Anonymous:1972:HC

[730] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 24(1):??, ???? 1972. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Inagaki:1973:ARB

[731] Nobuo Inagaki. Asymptotic relations between the likelihood estimating function and the maximum likelihood estimator. Annals of the Institute of Statistical Mathematics, 25(1):1–26, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479355.

Roussas:1973:EAF

[732] G. G. Roussas and A. Soms. On the exponential approximation of a family of probability measures and a representation theorem of Hájek– Inagaki. Annals of the Institute of Statistical Mathematics, 25(1):27–39, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479356.

Khan:1973:NCA

[733] A. H. Khan and S. M. Ali. A new coefficient of association. Annals of the Institute of Statistical Mathematics, 25(1):41–50, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479357.

Singh:1973:UKC

[734] J. Singh, B. N. Pandey, and K. Hirano. On the utilization of a known coefficient of kurtosis in the estimation procedure of variance. Annals of the Institute of Statistical Mathematics, 25(1):51–55, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479358.

Alam:1973:SBE

[735] Khursheed Alam and James R. Thompson. Some biased estimates of the mean of the normal distribution. Annals of the Institute of Statistical Mathematics, 25(1):57–64, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479359.

Mike:1973:RPT

[736] Valerie Miké. Robust Pitman-type estimators of location. Annals of the Institute of Statistical Mathematics, 25(1):65–86, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479360.

Walsh:1973:NEM

[737] John E. Walsh and Grace J. Kelleher. Nonparametric estimation of mean and variance when a few “sample” values possibly outliers. Annals of the Institute of Statistical Mathematics, 25(1):87–90, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479361.

Ghosh:1973:CAOa

[738] Malay Ghosh. On a class of asymptotically optimal nonparametric tests for grouped data. I. Annals of the Institute of Statistical Mathematics, 25(1):91–108, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479362.

Ghosh:1973:CAOb

[739] Malay Ghosh. On a class of asymptotically optimal nonparametric tests for grouped data II. Annals of the Institute of Statistical Mathematics, 25 (1):109–122, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479363.

Ghosh:1973:SSS

[740] Malay Ghosh and Pranab Kumar Sen. On some sequential simultaneous confidence intervals procedures. Annals of the Institute of Statistical Mathematics, 25(1):123–133, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479364.

Tan:1973:CAB

[741] W. Y. Tan. On the complex analogue of Bayesian estimation of a multivariate regression model. Annals of the Institute of Statistical Mathematics, 25(1):135–152, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479365.

Sugiura:1973:FAF

[742] Nariaki Sugiura. Further asymptotic formulas for the non-null distributions of three statistics for multivariate linear hypothesis. Annals of the Institute of Statistical Mathematics, 25(1):153–163, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479366.

Hirakawa:1973:NDM

[743] Fumiko Hirakawa. Note on the distribution of the minimum latent root. Annals of the Institute of Statistical Mathematics, 25(1):165–172, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479367.

Naik:1973:SPD

[744] Umesh D. Naik. Some posterior distributions concerning normal samples with applications to analysis of variance model I problems. Annals of the Institute of Statistical Mathematics, 25(1):173–186, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479368. See corrections

[843].

Kulkarni:1973:THA

[745] S. R. Kulkarni. On tests of hypotheses about treatment effects and treatment X places interactions, in two heteroscedastic experiments. Annals of the Institute of Statistical Mathematics, 25(1):187–203, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479369.

Rathie:1973:IFT

[746] P. N. Rathie and Pl. Kannappan. An inaccuracy function of type β. Annals of the Institute of Statistical Mathematics, 25(1):205–214, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479370.

Taguchi:1973:TDC

[747] Tokio Taguchi. On the two-dimensional concentration surface and extentions of concentration coefficient and Pareto distribution to the two dimensional case-III. Annals of the Institute of Statistical Mathematics, 25(1):215–237, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479371.

Ogawa:1973:ANN

[748] Junjiro Ogawa and Sadao Ikeda. The asymptotic non-null distribution of the F -static for testing a partial nullhypothesis in a randomized PBIB design with m associate classes under the Neyman model. Annals of the Institute of Statistical Mathematics, 25 (1):239–259, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479372.

Matsunawa:1973:UAJ

[749] Tadashi Matsunawa. Uniform asymptotic joint normality of sample quantities in censored cases. Annals of the Institute of Statistical Mathematics, 25(1):261–278, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479373.

Hettmansperger:1973:JLA

[750] Thomas P. Hettmansperger. On the Jodges–Lehmann approximate efficiency. Annals of the Institute of Statistical Mathematics, 25(1):279–286, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479374.

Chatterjee:1973:KST

[751] Shoutir Kishore Chatterjee and Pranab Kumar Sen. On Kolmogorov–Smirnovtype tests for symmetry. Annals of the Institute of Statistical Mathematics, 25(1):287–299, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479375.

Samanta:1973:EEL

[752] M. Samanta. Efficient estimation of the location parameters in the bivariate two sample problem. Annals of the Institute of Statistical Mathematics, 25(1):301–319, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479376.

Rohatgi:1973:SEM

[753] V. K. Rohatgi and R. T. O’Neill. On sequential estimation of the mean vector of a multinormal population. Annals of the Institute of Statistical Mathematics, 25(1):321–325, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479377.

Yamazoe:1973:ROP

[754] Siro Yamazoe. On random observation processes for stochastic approximation. Annals of the Institute of Statistical Mathematics, 25(1):327–334, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479378.

Han:1973:REB

[755] Chien-Pai Han. Regression estimation for bivariate normal distributions. Annals of the Institute of Statistical Mathematics, 25(1):335–344, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479379.

Khatri:1973:ENN

[756] C. G. Khatri and M. S. Srivastava. On the exact non-null distribution of likelihood ratio criteria for covariance matrices. Annals of the Institute of Statistical Mathematics, 25(1):345–354, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479380.

Patel:1973:TIM

[757] H. I. Patel and Rashid Ahmad. On tests of independence for r × c Markovian contingency tables. Annals of the Institute of Statistical Mathematics, 25(1):355–361, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479381.

John:1973:IPM

[758] S. John. On inferring the probability of misclassification by the linear discriminant function. Annals of the Institute of Statistical Mathematics, 25 (1):363–372, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479382.

Glick:1973:SPC

[759] Ned Glick. Separation and probability of correct classification among two or more distributions. Annals of the Institute of Statistical Mathematics, 25(1):373–382, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479383.

Srivastava:1973:EGM

[760] J. N. Srivastava and Lyman L. Mc-Donald. On the extensions of Gauss– Markov theorem to subsets of the parameter space under complex multivariate linear models. Annals of the Institute of Statistical Mathematics, 25 (1):383–393, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479384.

Hayakawa:1973:AED

[761] Takesi Hayakawa. An asymptotic expansion for the distribution of the determinant of a multivariate quadratic form in a normal sample. Annals of the Institute of Statistical Mathematics, 25(1):395–406, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479385.

Nagao:1973:AED

[762] Hisao Nagao. Asymptotic expansions of the distributions of Bartlett’s test and sphericity test under the local alternatives. Annals of the Institute of Statistical Mathematics, 25(1):407–422, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479386.

Fujikoshi:1973:AFD

[763] Yasunori Fujikoshi. Asymptotic formulas for the distributions of three statistics for multivariate linear hypothesis. Annals of the Institute of Statistical Mathematics, 25(1):423–437, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479387.

Saha:1973:CUB

[764] G. M. Saha and A. Dey. On construction and uses of balanced n-ary designs. Annals of the Institute of Statistical Mathematics, 25(1):439–445, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479388.

Chang:1973:ADC

[765] Tseng C. Chang. On an asymptotic distribution of the characteristic roots of S_{1} S_{2}^{−1} when roots are not all distinct. Annals of the Institute of Statistical Mathematics, 25(1):447–451, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479389.

Imai:1973:RLL

[766] Haruo Imai. Remarks to a local limit theorem for Galton–Watson processes. Annals of the Institute of Statistical Mathematics, 25(1):453–455, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479390.

Inagaki:1973:ARH

[767] Nobuo Inagaki. The asymptotic representation of the Hodges–Lehmann estimator based on Wilcoxon two-sample statistic. Annals of the Institute of Statistical Mathematics, 25(1):457–466, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479391.

Dudewicz:1973:MPE

[768] Edward J. Dudewicz. Maximum probability estimators for ranked means. Annals of the Institute of Statistical Mathematics, 25(1):467–477, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479392.

Hirano:1973:SPE

[769] Katuomi Hirano. Some properties of an estimator for the variance of a normal distribution. Annals of the Institute of Statistical Mathematics, 25(1):479–492, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479393.

Kirmani:1973:GFT

[770] S. N. U. A. Kirmani. On a goodness of fit test based on Matusita’s measure of distance. Annals of the Institute of Statistical Mathematics, 25(1):493–500, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479394.

Sinha:1973:CSE

[771] Bikas Kumar Sinha. Comparison of some experiments from sufficiency consideration. Annals of the Institute of Statistical Mathematics, 25(1):501–520, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479395.

Choi:1973:CEP

[772] Keewhan Choi. A consistent estimator of the parameters of continuous compound distribution functions. Annals of the Institute of Statistical Mathematics, 25(1):521–532, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479396.

Harville:1973:DTL

[773] David A. Harville. The distribution of a truncated linear difference between independent chi-square variates. Annals of the Institute of Statistical Mathematics, 25(1):533–548, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479397.

Guttman:1973:UDW

[774] Irwin Guttman and W. Y. Tan. The use of the disguised Wishart distribution in a Bayesian approach to tolerance region construction. Annals of the Institute of Statistical Mathematics, 25(1):549–556, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479398.

Mathai:1973:FRA

[775] A. M. Mathai. A few remarks about some recent articles on the exact distributions of multivariate test criteria: I. Annals of the Institute of Statistical Mathematics, 25(1):557–566, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479399.

Paik:1973:CFR

[776] U. B. Paik and W. T. Federer. On construction of fractional replicates and on aliasing schemes. Annals of the Institute of Statistical Mathematics, 25 (1):567–585, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479400.

Chopra:1973:OBF

[777] D. V. Chopra and J. N. Srivastava. Optimal balanced 2^{7} fractional factorial designs of resolution V, with N ≤ 42. Annals of the Institute of Statistical Mathematics, 25(1):587–604, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479401.

Saha:1973:CTP

[778] G. M. Saha. On construction of T_{m}- type PBIB designs. Annals of the Institute of Statistical Mathematics, 25 (1):605–616, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479402.

Gupta:1973:SPG

[779] B. C. Gupta and D. S. Tracy. Some properties and generating function of ordered partitions. Annals of the Institute of Statistical Mathematics, 25 (1):617–626, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479403.

Singh:1973:OSR

[780] Ravindra Singh and B. V. Sukhatme. Optimum stratification with ratio and regression methods of estimation. Annals of the Institute of Statistical Mathematics, 25(1):627–633, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479404.

Itoh:1973:RPI

[781] Yoshiaki Itoh. On a ruin problem with interaction. Annals of the Institute of Statistical Mathematics, 25(1):635–641, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479405.

Brock:1973:GFT

[782] Dwight B. Brock and A. M. Kshirsagar. A χ^{2} goodness-of-fit test for Markov renewal processes goodnessof-fit test for Markov renewal processes. Annals of the Institute of Statistical Mathematics, 25(1):643–654, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479406.

Anonymous:1973:HC

[783] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 25(1):??, ???? 1973. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Suzuki:1974:RBC

[784] Giitiro Suzuki. Robustness of Bayes classification region. Annals of the Institute of Statistical Mathematics, 26(1):1–13, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479800.

Britney:1974:BPE

[785] Robert R. Britney and Robert L. Winkler. Bayesian point estimation and prediction. Annals of the Institute of Statistical Mathematics, 26(1):15–34, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479801.

Govindarajulu:1974:BPR

[786] Z. Govindarajulu and Charles Harvey. Bayesian procedures for ranking and selection problems. Annals of the Institute of Statistical Mathematics, 26(1):35–53, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479802.

Sen:1974:CS

[787] Pranab Kumar Sen. On L^{p}- convergence of U-statistics. Annals of the Institute of Statistical Mathematics, 26(1):55–60, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479803.

Venkataraman:1974:CTL

[788] K. N. Venkataraman. Convergence theorems on the least square estimators of the structural parameters of a linear explosive model. Annals of the Institute of Statistical Mathematics, 26(1):61–85, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479804.

Susarla:1974:RCS

[789] V. Susarla. Rates of convergence in the sequence-compound squareddistance loss estimation and linearloss two-action problems for a family of scale parameter exponential distributions. Annals of the Institute of Statistical Mathematics, 26(1):87–102, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479805.

McCabe:1974:SER

[790] George P. McCabe, Jr. Sequential estimation of a restricted mean parameter of an exponential family. Annals of the Institute of Statistical Mathematics, 26(1):103–115, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479806.

Sugiura:1974:AFH

[791] Nariaki Sugiura. Asymptotic formulas for the hypergeometric function _{2}F_{1} of matrix argument, useful in multivariate analysis. Annals of the Institute of Statistical Mathematics, 26(1):117–125, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479807.

Han:1974:ADD

[792] Chien-Pai Han. Asymptotic distribution of discriminant function when covariance matrices are proportional and unknown. Annals of the Institute of Statistical Mathematics, 26(1):127–133, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479808.

Sen:1974:CSR

[793] P. K. Sen and P. R. Krishnaiah. On a class of simultaneous rank order tests in MANOCOVA. Annals of the Institute of Statistical Mathematics, 26 (1):135–145, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479809.

Walsh:1974:EIA

[794] John E. Walsh. Exact investigation of all effects for extensions of oneway ANOVA model with random effects. Annals of the Institute of Statistical Mathematics, 26(1):147–152, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479810.

Raghavarao:1974:SNS

[795] D. Raghavarao and K. R. Aggarwal. Some new series of PBIB designs and their applications. Annals of the Institute of Statistical Mathematics, 26 (1):153–161, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479811.

Kageyama:1974:NRA

[796] Sanpei Kageyama. Note on the reduction of associate classes for PBIB designs. Annals of the Institute of Statistical Mathematics, 26(1):163–170, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479812.

Dey:1974:ITC

[797] A. Dey and G. M. Saha. An inequality for tactical configurations. Annals of the Institute of Statistical Mathematics, 26(1):171–173, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479813.

Taguchi:1974:FTS

[798] Tokio Taguchi. On Fechner’s thesis and statistics with norm p. Annals of the Institute of Statistical Mathematics, 26(1):175–193, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479814.

Shimizu:1974:RTC

[799] Ryoichi Shimizu. On the remained term for the central limit theorem. Annals of the Institute of Statistical Mathematics, 26(1):195–201, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479815.

Anderson:1974:TCD

[800] Donald A. Anderson, Lyman L. Mc-Donald, and Kim D. Weaver. Tests on categorical data from the unionintersection principle. Annals of the Institute of Statistical Mathematics, 26 (1):203–213, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479816.

Khan:1974:SDE

[801] Rasul A. Khan. On sequential distinguishability for the exponential family. Annals of the Institute of Statistical Mathematics, 26(1):215–221, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479817.

Adichie:1974:SRP

[802] J. N. Adichie. On some robust properties of estimates of regression based on rank tests. Annals of the Institute of Statistical Mathematics, 26(1):223–231, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479818.

Ahmad:1974:SSS

[803] Rashid Ahmad. On the structure of symmetric sample testing: A distribution-free approach. Annals of the Institute of Statistical Mathematics, 26(1):233–245, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479819.

Waller:1974:BRS

[804] Ray A. Waller and David B. Duncan. A Bayes rule for the symmetric multiple comparisons problem II. Annals of the Institute of Statistical Mathematics, 26(1):247–264, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479820.

Ahuja:1974:CIL

[805] J. C. Ahuja and E. A. Enneking. Convolution of independent left-truncated negative binomial variables and limiting distributions. Annals of the Institute of Statistical Mathematics, 26 (1):265–270, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479821.

Janardan:1974:MMP

[806] K. G. Janardan and G. P. Patil. On multivariate modified Polya and inverse Polya distributions and their properties. Annals of the Institute of Statistical Mathematics, 26(1):271–276, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479822.

Chou:1974:AEN

[807] Charissa Chou and Minoru Siotani. Asymtotic expansion of the non-null distribution of the ratio of two conditionally independent Hotelling’s T_{0}^{2}- statistics. Annals of the Institute of Statistical Mathematics, 26(1):277–288, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479823.

Fujikoshi:1974:AEN

[808] Yasunori Fujikoshi. Asymptotic expansions of the non-null distributions of three statistics in GMANOVA. Annals of the Institute of Statistical Mathematics, 26(1):289–297, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479824.

Srivastava:1974:IMD

[809] J. N. Srivastava and M. K. Zaatar. Incomplete multivariate designs, optimal with respect to Fisher’s information matrix. Annals of the Institute of Statistical Mathematics, 26(1):299–313, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479825.

Aggarwal:1974:SHC

[810] K. R. Aggarwal. Some higher class PBIB designs and their application as confounded factorial experiments. Annals of the Institute of Statistical Mathematics, 26(1):315–323, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479826.

Nigam:1974:VFC

[811] A. K. Nigam. Variance functions for comparing mixture designs. Annals of the Institute of Statistical Mathematics, 26(1):325–329, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479827.

Hedayat:1974:PVB

[812] A. Hedayat and W. T. Federer. Pairwise and variance balanced incomplete block designs. Annals of the Institute of Statistical Mathematics, 26(1):331–338, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479828.

DeGraft-Johnson:1974:CRE

[813] K. T. DeGraft-Johnson and J. Sedransk. Comparison of ratio estimators in two-phase sampling. Annals of the Institute of Statistical Mathematics, 26(1):339–350, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479829.

Flaspohler:1974:QSD

[814] David C. Flaspohler. Quasi-stationary distributions for absorbing continuoustime denumerable Markov chains. Annals of the Institute of Statistical Mathematics, 26(1):351–356, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479830.

Wang:1974:NHP

[815] Y. H. Wang. A note on homogeneous processes with independent increments. Annals of the Institute of Statistical Mathematics, 26(1):357–360, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479831.

Kakwani:1974:NEZ

[816] N. C. Kakwani. A note on the efficiency of the Zelliner’s seemingly unrelated regressions estimator. Annals of the Institute of Statistical Mathematics, 26(1):361–362, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02479832.

Akaike:1974:MRS

[817] Hirotugu Akaike. Markovian representation of stochastic processes and its application to the analysis of autoregressive moving average processes. Annals of the Institute of Statistical Mathematics, 26(1):363–387, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479833.

Toussaint:1974:SPM

[818] Godfried T. Toussaint. Some properties of Matusita’s measure of affinity of several distributions. Annals of the Institute of Statistical Mathematics, 26(1):389–394, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479834.

Nagao:1974:ANN

[819] Hisao Nagao. Asymptotic non-null distributions of two test criteria for equality of covariance matrices under local alternatives. Annals of the Institute of Statistical Mathematics, 26(1):395–402, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479835.

Patil:1974:DMU

[820] S. A. Patil, J. L. Kovner, and D. C. Patel. The distribution of the MLE of the uniform correlation coefficient in the multivariate normal population. Annals of the Institute of Statistical Mathematics, 26(1):403–411, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479836.

Haq:1974:MMI

[821] M. Safiul Haq. A multivariate model with intra-class covariance structure. Annals of the Institute of Statistical Mathematics, 26(1):413–420, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02479837.

Puri:1974:ADR

[822] Madan L. Puri and Norman L. Wykoff. Asymptotic distribution of rank statistics for experiments involving incomplete block designs. Annals of the Institute of Statistical Mathematics, 26 (1):421–446, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02479838.

Banerjee:1974:SOR

[823] K. S. Banerjee. Some observations on repeated spring balance weighing designs. Annals of the Institute of Statistical Mathematics, 26(1):447–454, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02479839.

Hayashi:1974:QAC

[824] Chikio Hayashi and Tatsuzo Suzuki. Quantitative approach to a crosssocietal research; a comparative study of Japanese character. Annals of the Institute of Statistical Mathematics, 26 (1):455–516, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02479840.

Anonymous:1974:HC

[825] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 26(1):??, ???? 1974. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hayashi:1975:QAC

[826] Chikio Hayashi and Tatsuzo Suzuki. Quantitative approach to a crosssocietal research; a comparative study of Japanese character. Part II. Annals of the Institute of Statistical Mathematics, 27(1):1–32, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504621.

Yanagawa:1975:SRS

[827] Takashi Yanagawa. Stratified random sampling; gain in precision due to stratification in the case of proportional allocation. Annals of the Institute of Statistical Mathematics, 27(1):33–44, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504622.

Philippou:1975:ANM

[828] A. N. Philippou and G. G. Roussas. Asymptotic normality of the maximum likelihood estimate in the independent not identically distributed case. Annals of the Institute of Statistical Mathematics, 27(1):45–55, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504623.

Ghosh:1975:SPC

[829] Malay Ghosh. On some properties of a class of Spearman rank statistics with applications. Annals of the Institute of Statistical Mathematics, 27(1):57–68, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504624.

Srivastava:1975:CRS

[830] M. S. Srivastava. On a class of rank scores tests for censored data. Annals of the Institute of Statistical Mathematics, 27(1):69–78, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504625.

Lwin:1975:MD

[831] Thaung Lwin. On von Mises directions. Annals of the Institute of Statistical Mathematics, 27(1):79–86, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504626.

Huang:1975:CDE

[832] J. S. Huang. Characterization of distributions by the expected values of the order statistics. Annals of the Institute of Statistical Mathematics, 27(1):87–93, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504627.

Joshi:1975:IET

[833] S. W. Joshi. Integral expressions for tail probabilities of the negative multinomial distribution. Annals of the Institute of Statistical Mathematics, 27(1):95–97, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504628.

Fujikoshi:1975:AFN

[834] Yasunori Fujikoshi. Asymptotic formulas for the non-null distributions of three statistics for multivariate linear hypothesis. Annals of the Institute of Statistical Mathematics, 27(1):99–108, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504629.

Bhargava:1975:SRB

[835] R. P. Bhargava. Some results on beta distributions with application to multivariate problems. Annals of the Institute of Statistical Mathematics, 27 (1):109–116, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504630.

Pesotan:1975:CAG

[836] H. Pesotan, B. L. Raktoe, and W. T. Federer. On complexes of abelian groups with applications to fractional factorial designs. Annals of the Institute of Statistical Mathematics, 27 (1):117–142, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504631.

Yamamoto:1975:BAS

[837] Sumiyasu Yamamoto, Teruhiro Shirakura, and Masahide Kuwada. Balanced arrays of strength 2^{l} and balanced fractional 2^{m} factorial designs. Annals of the Institute of Statistical Mathematics, 27(1):143–157, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504632.

Banerjee:1975:SME

[838] A. K. Banerjee, A. Dey, and G. M. Saha. Some main effect plans for 3^{n} factorials. Annals of the Institute of Statistical Mathematics, 27(1):159–165, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504633.

Gupta:1975:SNS

[839] T. K. Gupta and A. Dey. On some new second order rotatable designs. Annals of the Institute of Statistical Mathematics, 27(1):167–175, ????

1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504634.

Kageyama:1975:NCP

[840] Sanpei Kageyama. Note on the construction of partially balanced arrays. Annals of the Institute of Statistical Mathematics, 27(1):177–180, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504635.

Sidak:1975:NCG

[841] Zbynek Sidák. A note on C. G. Khatri’s and A. Scott’s papers on multivariate normal distributions. Annals of the Institute of Statistical Mathematics, 27(1):181–184, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504636.

Nagao:1975:CNN

[842] Hisao Nagao. Corrections to “Nonnull distributions of the likelihood ratio criteria for independence and equality of mean vectors and covariance matrices”. Annals of the Institute of Statistical Mathematics, 27(1):185, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02504637. See [680].

Naik:1975:CSP

[843] Umesh D. Naik. Corrections to “Some posterior distributions concerning normal samples with applications to analysis of variance model I problems”. Annals of the Institute of Statistical Mathematics, 27(1):187, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02504638. See

[744].

Matsunawa:1975:EEJ

[844] T. Matsunawa. On the error evaluation of the joint normal approximation for sample quantiles. Annals of the Institute of Statistical Mathematics, 27(1):189–199, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504639.

Yamazaki:1975:PDT

[845] Genji Yamazaki and Hirotaka Sakasegawa. Properties of duality in tandem queueing systems. Annals of the Institute of Statistical Mathematics, 27 (1):201–212, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504640.  See correction

[962].

Fu:1975:CAP

[846] James C. Fu and Leon Jay Gleser. Classical asymptotic properties of a certain estimator related to the maximum likelihood estimator. Annals of the Institute of Statistical Mathematics, 27(1):213–233, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504641.

Cacoullos:1975:MVU

[847] T. Cacoullos and Ch. Charalambides. On minimum variance unbiased estimation for truncated binomial and negative binomial distributions. Annals of the Institute of Statistical Mathematics, 27(1):235–244, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504642.

Sinha:1975:SPU

[848] Bikas Kumar Sinha and Bimal Kumar Sinha. Some problems of unbiased sequential binomial estimation. Annals of the Institute of Statistical Mathematics, 27(1):245–258, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504643.

Sharma:1975:BAI

[849] Ran S. Sharma. Bayes approach to interval estimation of a binomial parameter. Annals of the Institute of Statistical Mathematics, 27(1):259–267, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504644.

Kapadia:1975:EPT

[850] C. H. Kapadia and R. L. Thomasson. On estimating the parameter of a truncated geometric distribution by the method of moments. Annals of the Institute of Statistical Mathematics, 27(1):269–272, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504645.

Singh:1975:OSE

[851] Ravindra Singh and Dev Parkash. Optimum stratification for equal allocation. Annals of the Institute of Statistical Mathematics, 27(1):273–280, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504646.

Ali:1975:CCC

[852] Mir M. Ali and Winston A. Richards. On a conjecture concerning the common content of an N-cube and a diagonal cylinder. Annals of the Institute of Statistical Mathematics, 27(1):281–287, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504647.

Mase:1975:DID

[853] Shigeru Mase. Decomposition of infinitely divisible characteristic functions with absolutely continuous Poisson spectral measure. Annals of the Institute of Statistical Mathematics, 27 (1):289–298, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504648.

Adichie:1975:NPS

[854] J. N. Adichie. Non-parametric c-sample tests with regression. Annals of the Institute of Statistical Mathematics, 27 (1):299–307, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504649.

Joshi:1975:SDT

[855] Prakash C. Joshi. Some distribution theory results for a regression model. Annals of the Institute of Statistical Mathematics, 27(1):309–317, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504650.

Gleser:1975:NBG

[856] Leon Jay Gleser and Ingram Olkin. A note on Box’s general method of approximation for the null distributions of likelihood criteria. Annals of the Institute of Statistical Mathematics, 27(1):319–326, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504651.

Bhargava:1975:SOS

[857] R. P. Bhargava. Some one-sample hypothesis testing problems when there is a monotone sample from a multivariate normal population. Annals of the Institute of Statistical Mathematics, 27(1):327–339, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504652.

Gupta:1975:SIW

[858] A. K. Gupta. On a stochastic inequality for the Wilks statistic. Annals of the Institute of Statistical Mathematics, 27(1):341–348, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504653.  See correction

[888].

Han:1975:TEC

[859] Chien-Pai Han. Testing the equality of covariance matrices under intraclass correlation models. Annals of the Institute of Statistical Mathematics, 27(1):349–356, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504654.

Hirakawa:1975:SDL

[860] Fumiko Hirakawa. Some distributions of the latent roots of a complex Wishart matrix variate. Annals of the Institute of Statistical Mathematics, 27 (1):357–363, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504655.

Bhargava:1975:ASM

[861] R. P. Bhargava and K. R. Shah. Analysis of some mixed-models for block and split-plot designs. Annals of the Institute of Statistical Mathematics, 27 (1):365–375, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504656.

Shirakura:1975:NBF

[862] Teruhiro Shirakura and Masahide Kuwada. Note on balanced fractional 2^{m} factorial designs of resolution 2l + 1. Annals of the Institute of Statistical Mathematics, 27(1):377–386, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504657.

Saha:1975:NRB

[863] G. M. Saha. A note on relations between incomplete block and weighing designs. Annals of the Institute of Statistical Mathematics, 27(1):387–390, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504658.

Inagaki:1975:WCL

[864] Nobuo Inagaki and Yosihiko Ogata. The weak convergence of likelihood ratio random fields and its applications. Annals of the Institute of Statistical Mathematics, 27(1):391–419, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504659.

Ikeda:1975:SCU

[865] Sadao Ikeda. Some criteria for uniform asymptotic equivalence of real probability distributions. Annals of the Institute of Statistical Mathematics, 27(1):421–428, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504660.

Koul:1975:ANE

[866] Hira Lal Koul. Asymptotic normality of H–L estimators based on dependent data. Annals of the Institute of Statistical Mathematics, 27(1):429–441, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504661.

Guttman:1975:EAC

[867] Irwin Guttman and Charles D. Palit. Effect of auto-correlations on the optimum allocations in two phase stratified sampling — a Bayesian approach. Annals of the Institute of Statistical Mathematics, 27(1):443–462, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504662.

Brown:1975:SDA

[868] George F. Brown, Jr. The smalldisturbance asymptotic moment matrix of k-class estimates of parameters of different equations in a complete system of simultaneous linear equations. Annals of the Institute of Statistical Mathematics, 27(1):463–472, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504663.

Muirhead:1975:ADE

[869] Robb J. Muirhead and Yasuko Chikuse. Approximations for the distributions of the extreme latent roots of three matrices. Annals of the Institute of Statistical Mathematics, 27(1):473–478, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504664.

Sen:1975:TDC

[870] Ashish Sen and S. Srivastava. On tests for detecting change in mean when variance is unknown. Annals of the Institute of Statistical Mathematics, 27 (1):479–486, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504665.

Gore:1975:SNT

[871] A. P. Gore. Some nonparametric tests and selection procedures for main effects in two-way layouts. Annals of the Institute of Statistical Mathematics, 27(1):487–500, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504666.

Sahai:1975:BEE

[872] Hardeo Sahai. Bayes equivariant estimators in a crossed classification random effects model. Annals of the Institute of Statistical Mathematics, 27 (1):501–505, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504667.

Maejima:1975:LLT

[873] Makoto Maejima. On local limit theorems and Blackwell’s renewal theorem for independent random variables. Annals of the Institute of Statistical Mathematics, 27(1):507–520, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504668.

Pirie:1975:NTT

[874] Walter R. Pirie and Myles Hollander. Note on a Tukey test for ordered alternatives. Annals of the Institute of Statistical Mathematics, 27(1):521–523, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504669.

Kelleher:1975:ITW

[875] Grace J. Kelleher and John E. Walsh. Intervals and tests for weighted sum of percentiles θ(p) and θ(1 − p) for symmetrical populations. Annals of the Institute of Statistical Mathematics, 27(1):525–527, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504670.

Kageyama:1975:NIT

[876] Sanpei Kageyama. Note on an inequality for tactical configurations. Annals of the Institute of Statistical Mathematics, 27(1):529–530, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02504671.

Anonymous:1975:HC

[877] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 27(1):??, ???? 1975. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Sakasegawa:1976:OMS

[878] Hirotaka Sakasegawa. Orthogonal mesh sampling method. Annals of the Institute of Statistical Mathematics, 28(1):1–7, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504725.

Dharmadhikari:1976:CUD

[879] S. W. Dharmadhikari and Kumar Jogdeo. On characterizations of unimodality of discrete distributions. Annals of the Institute of Statistical Mathematics, 28(1):9–18, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504726.

deOliveira:1976:ABM

[880] J. Tiago de Oliveira. Asymptotic behaviour of maxima with periodic disturbances. Annals of the Institute of Statistical Mathematics, 28(1):19–23, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504727.

Srinivasan:1976:FRS

[881] R. Srinivasan and R. M. Wharton. Further results on simultaneous confidence intervals for the normal distribution. Annals of the Institute of Statistical Mathematics, 28(1):25–33, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504728.

Akahira:1976:AEE

[882] Masafumi Akahira. On the asymptotic efficiency of estimators in an autoregressive process. Annals of the Institute of Statistical Mathematics, 28(1):35–48, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504729.

Stigum:1976:APD

[883] Bernt P. Stigum. Asymptotic properties of dynamic stochastic parameter estimates. Annals of the Institute of Statistical Mathematics, 28(1):49–75, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504730.

Beran:1976:AEA

[884] Rudolf Beran. Adaptive estimates for autoregressive processes. Annals of the Institute of Statistical Mathematics, 28(1):77–89, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504731.

Roy:1976:SAR

[885] Roch Roy. Spectral analysis for a random process on the sphere. Annals of the Institute of Statistical Mathematics, 28(1):91–97, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504732.

Rao:1976:ERS

[886] A. N. V. Rao and Chris P. Tsokos. On the existence of a random solution to a nonlinear perturbed stochastic integral equation. Annals of the Institute of Statistical Mathematics, 28(1):99–109, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504733.

Arvesen:1976:NTH

[887] James N. Arvesen. A note on the Tukey–Hooke variance component results. Annals of the Institute of Statistical Mathematics, 28(1):111–121, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504734.

Gupta:1976:CSI

[888] A. K. Gupta. Correction to “On a stochastic inequality for the Wilks statistic”. Annals of the Institute of Statistical Mathematics, 28(1):123, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02504735. See [858].

Matsubara:1976:BTI

[889] Nozomu Matsubara. Bayes theorem, information number and behavior of posterior distributions. Annals of the Institute of Statistical Mathematics, 28 (1):125–144, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504736. See corrections

[919].

Murakami:1976:RCC

[890] Masakatsu Murakami. On the reduction to a complete class in multiple decision problems. Annals of the Institute of Statistical Mathematics, 28 (1):145–165, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

                                        1007/BF02504737.  See correction

[920].

Tiao:1976:SBC

[891] G. C. Tiao and B. Afonja. Some Bayesian considerations of the choice of design for ranking, selection and estimation. Annals of the Institute of Statistical Mathematics, 28(1):167–185, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504738.

Chikuse:1976:PDE

[892] Yasuko Chikuse. Partial differential equations for hypergeometric functions of complex argument matrices and their applications. Annals of the Institute of Statistical Mathematics, 28 (1):187–199, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504739.

Gordon:1976:CWD

[893] F. S. Gordon and S. P. Gordon. Characterizations of the Wishart distribution using regression properties. Annals of the Institute of Statistical Mathematics, 28(1):201–220, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504740.

Noda:1976:ERF

[894] Kazuo Noda. Estimation of a regression function by the Parzen kerneltype density estimators. Annals of the Institute of Statistical Mathematics, 28(1):221–234, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504741.

Saleh:1976:HLE

[895] A. K. Md. Ehsanes Saleh. Hodges– Lehmann estimate of the location parameter in censored samples. Annals of the Institute of Statistical Mathematics, 28(1):235–247, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504742.

Sinha:1976:AMT

[896] Bimal Kumar Sinha. Approximately minimax tests for testing hypotheses about a random parameter with unknown distribution. Annals of the Institute of Statistical Mathematics, 28 (1):249–258, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504743.

Ahmad:1976:MSP

[897] R. Ahmad. On the multivariate ksample problem and the generalization of the Kolmogorov–Smirnov-test. Annals of the Institute of Statistical Mathematics, 28(1):259–265, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504744.

Kim:1976:SD

[898] P. J. Kim. The Smirnov distribution. Annals of the Institute of Statistical Mathematics, 28(1):267–275, ????

1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504745.

Subrahmanya:1976:EQL

[899] M. T. Subrahmanya. Equitable quality level and error-areas under the operating characteristic curves of normal single sampling inspection plans (with σ known). Annals of the Institute of Statistical Mathematics, 28(1):277–290, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504746.

Matsunawa:1976:SIB

[900] T. Matsunawa. Some inequalities based on inverse factorial series. Annals of the Institute of Statistical Mathematics, 28(1):291–305, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504747.

Govindarajulu:1976:NDF

[901] Z. Govindarajulu. A note on distribution-free confidence bounds for P (X < Y) when X and Y are dependent. Annals of the Institute of Statistical Mathematics, 28(1):307–308, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02504748.

Hayakawa:1976:NTC

[902] Takesi Hayakawa. The new test criterion for the homogeneity of parameters of several populations. Annals of the Institute of Statistical Mathematics, 28(1):309–328, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504749.

Yanagimoto:1976:ITS

[903] Takemi Yanagimoto and Masaaki Sibuya. Isotonic tests for spread and tail. Annals of the Institute of Statistical Mathematics, 28(1):329–342, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504750.

Wegner:1976:EMP

[904] Horst Wegner. On the existence of maximum probability estimators. Annals of the Institute of Statistical Mathematics, 28(1):343–347, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504751.

Chakravorti:1976:APM

[905] S. R. Chakravorti. On asymptotic properties of the maximum likelihood estimates of the general growth curve model. Annals of the Institute of Statistical Mathematics, 28(1):349–357, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504752.

Wolfowitz:1976:AEE

[906] J. Wolfowitz. Asymptotically efficient estimators when the densities of the observations have discontinuities. Annals of the Institute of Statistical Mathematics, 28(1):359–370, ????

1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504753.

Trenkler:1976:NMT

[907] G. Trenkler. On a new method of testing statistical hypotheses. Annals of the Institute of Statistical Mathematics, 28(1):371–384, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504754.

Kirmani:1976:LBB

[908] S. N. U. A. Kirmani. A lower bound on Bayes risk in classification problems. Annals of the Institute of Statistical Mathematics, 28(1):385–387, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504755.

OBryan:1976:REB

[909] Thomas E. O’Bryan and V. Susarla. Rates in the empirical Bayes estimation problem with non-identical components. Annals of the Institute of Statistical Mathematics, 28(1):389–397, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504756.

Greenwood:1976:MTG

[910] J. Arthur Greenwood. Moments of the time to generate random variables by rejection. Annals of the Institute of Statistical Mathematics, 28(1):399–401, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504757.

Nagao:1976:PST

[911] Hisao Nagao. Properties of some test criteria for covariance matrix. Annals of the Institute of Statistical Mathematics, 28(1):403–409, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504758.

Srivastava:1976:TST

[912] V. K. Srivastava and Ramji Tiwari. Two-stage and three-stage least squares estimation of dispersion matrix of disturbances in simultaneous equations. Annals of the Institute of Statistical Mathematics, 28(1):411–428, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02504759.

Janardan:1976:CEP

[913] K. G. Janardan. Certain estimation problems for multivariate hypergeometric models. Annals of the Institute of Statistical Mathematics, 28 (1):429–444, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504760.

Reddy:1976:SSR

[914] V. Nagi Reddy. Stratified simple random sampling and prior distributions. Annals of the Institute of Statistical Mathematics, 28(1):445–459, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504761.

Koop:1976:BRE

[915] J. C. Koop. Bias reduction and efficiency of reconstructed ratio estimators for a finite universe. Annals of the Institute of Statistical Mathematics, 28(1):461–467, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504762.

Davies:1976:IDL

[916] Laurie Davies and Ryoichi Shimizu. On identically distributed linear statistics. Annals of the Institute of Statistical Mathematics, 28(1):469–489, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504763.

Dallas:1976:CPP

[917] A. C. Dallas. Characterizing the Pareto and power distributions. Annals of the Institute of Statistical Mathematics, 28(1):491–497, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02504764.

Charalambides:1976:ANC

[918] Ch. A. Charalambides. The asymptotic normality of certain combinatorial distributions. Annals of the Institute of Statistical Mathematics, 28(1):499–506, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02504765.

Matsubara:1976:CBT

[919] Nozomu Matsubara. Corrections to “Bayes theorem, information number and behavior of posterior distributions”. Annals of the Institute of Statistical Mathematics, 28(1):507, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02504766. See [889].

Murakami:1976:CRC

[920] Masakatsu Murakami. Correction to “On the reduction to a complete class in multiple decision problems”. Annals of the Institute of Statistical Mathematics, 28(1):508, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02504767. See

[890].

Hasegawa:1976:PSD

[921] Masami Hasegawa and Masaharu Tanemura. On the pattern of space division by territories. Annals of the Institute of Statistical Mathematics, 28 (1):509–519, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02504768.

Anonymous:1976:HC

[922] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 28(1):??, ???? 1976. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takahasi:1977:RRT

[923] Koiti Takahasi and Hirotaka Sakasegawa. A randomized response technique without making use of any randomizing device. Annals of the Institute of Statistical Mathematics, 29(1):1–8, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532769.

Akaike:1977:OUB

[924] Hirotugu Akaike. An objective use of Bayesian models. Annals of the Institute of Statistical Mathematics, 29(1):9–20, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532770.

Hirano:1977:EPB

[925] Katuomi Hirano. Estimation procedures based on preliminary test, shrinkage technique and information criterion. Annals of the Institute of Statistical Mathematics, 29(1):21–34, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL http://link.springer.com/
article/10.1007/BF02532771. See

corrections [948].

Gupta:1977:TPM

[926] Somesh Das Gupta. Two problems in multivariate analysis: BLUS residuals and testability of linear hypothesis. Annals of the Institute of Statistical Mathematics, 29(1):35–41, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532772.

Stone:1977:UAC

[927] M. Stone. A unified approach to coordinate-free multivariate analysis. Annals of the Institute of Statistical Mathematics, 29(1):43–57, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532773.

Chiu:1977:CSR

[928] W. K. Chiu. On correct selection for a ranking problem. Annals of the Institute of Statistical Mathematics, 29(1):59–66, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532774.

Sakasegawa:1977:AF

[929] Hirotaka Sakasegawa. An approximation formula L_{q} ≃ α · ρ^{β}/(1 − ρ). Annals of the Institute of Statistical Mathematics, 29(1):67–75, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532775.

Puri:1977:ADM

[930] Prem S. Puri. On the asymptotic distribution of the maximum of sums of a random number of I.I.D. random variables. Annals of the Institute of Statistical Mathematics, 29(1):77–87, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532776.

Mudholkar:1977:SBD

[931] Govind S. Mudholkar and Siddhartha R. Dalal. Some bounds on the distribution functions of linear combinations and applications. Annals of the Institute of Statistical Mathematics, 29(1):89–100, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532777.

Venkataraman:1977:SLT

[932] K. N. Venkataraman. A spectral limit theorem on a non-linear stochastic process with non-additive, independent, linear components. Annals of the Institute of Statistical Mathematics, 29 (1):101–118, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532778.

Patel:1977:HMM

[933] Y. C. Patel. Higher moments of moment estimators and even point estimators for the parameters of the Hermite distribution. Annals of the Institute of Statistical Mathematics, 29(1):119–130, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532779.

Inagaki:1977:TES

[934] Nobuo Inagaki. Two errors in statistical model fitting. Annals of the Institute of Statistical Mathematics, 29 (1):131–152, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532780.

Akaike:1977:EMM

[935] Hirotugu Akaike. An extension of the method of maximum likelihood and the Stein’s problem. Annals of the Institute of Statistical Mathematics, 29 (1):153–164, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532781.

Ogata:1977:WCL

[936] Yosihiko Ogata and Nobuo Inagaki. The weak convergence of the likelihood ratio random fields for Markov observations. Annals of the Institute of Statistical Mathematics, 29(1):165–187, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532782.

Lind:1977:CTC

[937] Bruce Lind and George Roussas. Cramér-type conditions and quadratic mean differentiability. Annals of the Institute of Statistical Mathematics, 29 (1):189–201, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532783.

Roussas:1977:APM

[938] George G. Roussas. Asymptotic properties of the maximum probability estimates in Markov processes. Annals of the Institute of Statistical Mathematics, 29(1):203–219, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532784.

Chikuse:1977:AEJ

[939] Yasuko Chikuse. Asymptotic expansions for the joint and marginal distributions of the latent roots of S_{1}S_{2}^{−1}. Annals of the Institute of Statistical Mathematics, 29(1):221–233, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532785.

Isogai:1977:AED

[940] Takafumi Isogai. Asymptotic expansions for the distributions of latent roots of S_{h}S_{e}^{−1} and of certain test statistics in MANOVA. Annals of the Institute of Statistical Mathematics, 29(1):235–246, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532786.

Lachenbruch:1977:CAD

[941] Peter A. Lachenbruch. Covariance adjusted discriminant functions. Annals of the Institute of Statistical Mathematics, 29(1):247–257, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532787.

Nath:1977:ICP

[942] G. Baikunth Nath. Inference concerning the population correlation coefficient from bivariate normal samples based on minimal observations. Annals of the Institute of Statistical Mathematics, 29(1):259–273, ????

1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532788.

Mase:1977:DID

[943] Shigeru Mase. Decomposition of infinitely divisible characteristic functions without Gaussian component. Annals of the Institute of Statistical Mathematics, 29(1):275–286, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532789.

Stephens:1977:WRC

[944] M. A. Stephens. Whitworth runs on a circle. Annals of the Institute of Statistical Mathematics, 29(1):287–293, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532790.

Tashiro:1977:MGU

[945] Yoshihiro Tashiro. On methods for generating uniform random points on the surface of a sphere. Annals of the Institute of Statistical Mathematics, 29(1):295–300, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532791.

Srivastava:1977:ESD

[946] J. N. Srivastava and S. Ghosh. On the existence of search designs with continuous factors. Annals of the Institute of Statistical Mathematics, 29(1):301–306, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532792.

Makino:1977:III

[947] Toji Makino. On the independence of interdeparture intervals from single server queueing systems. Annals of the Institute of Statistical Mathematics, 29(1):307–315, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532793. See corrections

[1328].

Hirano:1977:CEP

[948] Katuomi Hirano. Corrections to “Estimation procedures based on preliminary test, shrinkage technique and information criterion”. Annals of the Institute of Statistical Mathematics, 29(1):317, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02532794. See

[925].

Kitagawa:1977:SPO

[949] Genshiro Kitagawa. On a search procedure for the optimal AR–MA order. Annals of the Institute of Statistical Mathematics, 29(1):319–332, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532795.

Matsunawa:1977:APB

[950] T. Matsunawa. Approximations to the probabilities of binomial and multinomial random variables and chi-square type statistics. Annals of the Institute of Statistical Mathematics, 29 (1):333–358, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532796.

Hayakawa:1977:LRC

[951] Takesi Hayakawa. The likelihood ratio criterion and the asymptotic expansion of its distribution. Annals of the Institute of Statistical Mathematics, 29(1):359–378, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532797. See corrections

[1456].

Fujikoshi:1977:AED

[952] Y. Fujikoshi. An asymptotic expansion for the distributions of the latent roots of the Wishart matrix with multiple population roots. Annals of the Institute of Statistical Mathematics, 29(1):379–387, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532798.

Konishi:1977:AED

[953] Sadanori Konishi. Asymptotic expansion for the distribution of a function of latent roots of the covariance matrix. Annals of the Institute of Statistical Mathematics, 29(1):389–396, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532799.

Takeuchi:1977:EET

[954] Kei Takeuchi and Masafumi Akahira. Extension of Edgeworth type expansion of the distribution of the sums of I.I.D. random variables in nonregular cases. Annals of the Institute of Statistical Mathematics, 29(1):397–406, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532800.

Nakatsuka:1977:RAC

[955] Toshinao Nakatsuka. Regions of autocorrelation coefficients and of their estimators in a stationary time series. Annals of the Institute of Statistical Mathematics, 29(1):407–414, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532801.

Huzii:1977:SEO

[956] Mituaki Huzii. On a spectral estimate obtained by an autoregressive model fitting. Annals of the Institute of Statistical Mathematics, 29 (1):415–431, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532802. See corrections

[957, 996].

Huzii:1977:CSE

[957] Mituaki Huzii. Corrections to “On a spectral estimate obtained by an autoregressive model fitting”. Annals of the Institute of Statistical Mathematics, 29(1):432, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02532803. See

[956].

Nishi:1977:LCP

[958] Akihiro Nishi. On linear classification procedures between two categories with known mean vectors and covariance matrices. Annals of the Institute of Statistical Mathematics, 29(1):433–444, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532804.

Sakasegawa:1977:IAF

[959] Hirotaka Sakasegawa and Genji Yamazaki. Inequalities and an approximation formula for the mean delay time in tandem queueing systems. Annals of the Institute of Statistical Mathematics, 29(1):445–466, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532805.

Finney:1977:ERC

[960] D. J. Finney. Estimation of the response curve in radioligand assays. Annals of the Institute of Statistical Mathematics, 29(1):467–477, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02532806.

Barankin:1977:BNF

[961] Edward W. Barankin. Build-up of the notion of flanking. Annals of the Institute of Statistical Mathematics, 29 (1):479–507, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02532807.

Yamazaki:1977:CPD

[962] Genji Yamazaki and Hirotaka Sakasegawa. Correction to “Properties of duality in tandem queueing systems”. Annals of the Institute of Statistical Mathematics, 29(1):509, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02532808. See

[845].

Suzuki:1977:MPI

[963] Tatsuzo Suzuki and Ted T. Jitodai. Migration and prefectural identification in four Japanese prefectures. Annals of the Institute of Statistical Mathematics, 29(1):511–525, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02532809.

Anonymous:1977:HC

[964] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 29(1):??, ???? 1977. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hirano:1978:LSP

[965] Katuomi Hirano. On level of significance of the preliminary test in pooling means. Annals of the Institute of Statistical Mathematics, 30(1):1–8, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480193.

Akaike:1978:BAM

[966] Hirotugu Akaike. A Bayesian analysis of the minimum AIC procedure. Annals of the Institute of Statistical Mathematics, 30(1):9–14, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480194.

Suzuki:1978:DOT

[967] Giitiro Suzuki. Detecting optimum time of control action for a manufacturing system. Annals of the Institute of Statistical Mathematics, 30(1):15–26, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480195.

Murakami:1978:RCC

[968] Masakatsu Murakami. On the reduction to a complete class in multiple decision problems (2). Annals of the Institute of Statistical Mathematics, 30(1):27–34, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480196.

Nogami:1978:SCO

[969] Yoshiko Nogami. The set-compound one-stage estimation in the nonregular family of distributions over the interval (0, θ). Annals of the Institute of Statistical Mathematics, 30(1):35–43, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480197.

Choi:1978:APL

[970] Keewhan Choi. Asymptotic properties of least squares estimators for discrete compound distributions. Annals of the Institute of Statistical Mathematics, 30(1):45–50, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480198.

Hayakawa:1978:AED

[971] Takesi Hayakawa. The asymptotic expansion of the distribution of Anderson’s statistic for testing a latent vector of a covariance matrix. Annals of the Institute of Statistical Mathematics, 30(1):51–55, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480199.

Chikuse:1978:ADL

[972] Yasuko Chikuse. Asymptotic distributions of the latent roots with multiple population roots in multiple discriminant analysis. Annals of the Institute of Statistical Mathematics, 30(1):57–62, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480200.

Anderson:1978:UFP

[973] T. W. Anderson and S. G. Ghurye. Unique factorization of products of bivariate normal cumulative distribution functions. Annals of the Institute of Statistical Mathematics, 30(1):63–69, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480201.

Davis:1978:ADG

[974] A. W. Davis. On the asymptotic distribution of Gower’s m^{2} goodnessof-fit criterion in a particular case. Annals of the Institute of Statistical Mathematics, 30(1):71–79, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480202.

Szatrowski:1978:ESO

[975] Ted H. Szatrowski. Explicit solutions, one iteration convergence and averaging in the multivariate normal estimation problem for patterned means and covariances. Annals of the Institute of Statistical Mathematics, 30(1):81–88, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480203.

Srivastava:1978:LSA

[976] V. K. Srivastava and Sushama Upadhyaya. Large-sample approximations in seemingly unrelated regression equations. Annals of the Institute of Statistical Mathematics, 30(1):89–96, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480204.

Wakimoto:1978:CGM

[977] Kazumasa Wakimoto and Masaaki Taguri. Constellation graphical method for representing multi-dimensional data. Annals of the Institute of Statistical Mathematics, 30(1):97–104, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480205.

Gyorfi:1978:DGA

[978] L. Györfi and T. Nemetz. f-Dissimilarity: A generalization of the affinity of several distributions. Annals of the Institute of Statistical Mathematics, 30(1):105–113, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480206.

Kageyama:1978:RNA

[979] S. Kageyama, G. M. Saha, and A. D. Das. Reduction of the number of associate classes of hypercubic association schemes. Annals of the Institute of Statistical Mathematics, 30(1):115–123, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480207.

Barankin:1978:BRVa

[980] Edward W. Barankin and Koiti Takashasi. Betweenness for real vectors and lines, I. Basic generalities. Annals of the Institute of Statistical Mathematics, 30(1):125–162, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480208.

Ahsanullah:1978:CED

[981] M. Ahsanullah. On a characterization of the exponential distribution by spacings. Annals of the Institute of Statistical Mathematics, 30(1):163–166, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480209.

Hwang:1978:NBM

[982] J. S. Hwang. A note on Bernstein and Müntz–Szasz theorems with applications to the order statistics. Annals of the Institute of Statistical Mathematics, 30(1):167–176, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480210.

Takahashi:1978:NPS

[983] Rinya Takahashi. Note on k-point separation measurement. Annals of the Institute of Statistical Mathematics, 30(1):177–179, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480211.

Eaton:1978:NGM

[984] Morris L. Eaton. A note on the Gauss– Markov theorem. Annals of the Institute of Statistical Mathematics, 30 (1):181–184, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480212.

Sakamoto:1978:ACC

[985] Yosiyuki Sakamoto and Hirotugu Akaike. Analysis of cross classified data by AIC. Annals of the Institute of Statistical Mathematics, 30(1):185–197, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480213.

Hayashi:1978:CTT

[986] Chikio Hayashi and Fumi Hayashi. Comparison of two types of multidimensional scaling methods. Annals of the Institute of Statistical Mathematics, 30(1):199–209, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480214.

Taguchi:1978:GGD

[987] Tokio Taguchi. On a generalization of Gaussian distribution. Annals of the Institute of Statistical Mathematics, 30(1):211–242, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480215. See corrections

[1009].

Ogata:1978:ABM

[988] Yoshiko Ogata. The asymptotic behaviour of maximum likelihood estimators for stationary point processes. Annals of the Institute of Statistical Mathematics, 30(1):243–261, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480216.

Shimizu:1978:EMP

[989] Ryoichi Shimizu. Entropy maximization principle and selection of the order of an autoregressive Gaussian process. Annals of the Institute of Statistical Mathematics, 30(1):263–270, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480217.

Sakasegawa:1978:GNP

[990] Hirotaka Sakasegawa. On a generation of normal pseudo-random numbers. Annals of the Institute of Statistical Mathematics, 30(1):271–279, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480218.

Tolley:1978:NPT

[991] H. Dennis Tolley. A non-parametric test for composite hypotheses in survival analysis. Annals of the Institute of Statistical Mathematics, 30(1):281–295, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480219.

Davies:1978:LSE

[992] Laurie Davies and Ludwig Baringhaus. Linear statistics and exponential families. Annals of the Institute of Statistical Mathematics, 30(1):297–314, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480220.

Nakatsuka:1978:RAC

[993] Toshinao Nakatsuka. Regions of autocorrelation coefficients in AR(p) and EX(p) processes. Annals of the Institute of Statistical Mathematics, 30 (1):315–319, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480221.

Rao:1978:DEM

[994] B. L. S. Prakasa Rao. Density estimation for Markov processes using deltasequences. Annals of the Institute of Statistical Mathematics, 30(1):321–328, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480222.

Tanaka:1978:SGM

[995] Yutaka Tanaka. Some generalized methods of optimal scaling and their asymptotic theories: The case of multiple responses-multiple factors. Annals of the Institute of Statistical Mathematics, 30(1):329–348, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480223.

Anonymous:1978:CSE

[996] Anonymous. Corrections to “On a spectral estimate obtained by an autoregressive model fitting”. Annals of the Institute of Statistical Mathematics, 30(1):349, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02480224. See

[956].

Kitagawa:1978:PMN

[997] Genshiro Kitagawa and Hirotugu Akaike. A procedure for the modeling of non-stationary time series. Annals of the Institute of Statistical Mathematics, 30(1):351–363, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480225.

Siotani:1978:STP

[998] Minoru Siotani, Dhanwant S. Gill, and Christian Löschcke. Stepwise test procedures and approximate chi-square analysis. Annals of the Institute of Statistical Mathematics, 30(1):365–375, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480226.

Nagao:1978:AED

[999] Hisao Nagao. An asymptotic expansion for the distribution of a function of latent roots of the noncentral Wishart matrix, when ω = O(n). Annals of the Institute of Statistical Mathematics, 30(1):377–383, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480227.

vanderMerwe:1978:AES

[1000] A. J. van der Merwe and D. J.

de Waal. The asymptotic expansion of the Stein estimators for the vector case. Annals of the Institute of Statistical Mathematics, 30(1):385–395, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480228.

Taylor:1978:UCC

[1001] R. L. Taylor and K. F. Cheng.

On the uniform complete convergence of density function estimates. Annals of the Institute of Statistical Mathematics, 30(1):397–406, ????

1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480229.

Bhattacharya:1978:YTE

[1002] C. G. Bhattacharya. Yates type estimators of a common mean. Annals of the Institute of Statistical Mathematics, 30(1):407–414, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480230.

Campos:1978:ECR

[1003] Antonio Dorival Campos. An extension of the Cramér-Rao inequality for a sequential procedure without assuming regularity conditions. Annals of the Institute of Statistical Mathematics, 30(1):415–419, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480231.

Cressie:1978:TET

[1004] Noel Cressie. Testing for the equality of two binomial proportions. Annals of the Institute of Statistical Mathematics, 30(1):421–427, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480232.

Ahsanullah:1978:RVE

[1005] M. Ahsanullah. Record values and the exponential distribution. Annals of the Institute of Statistical Mathematics, 30(1):429–433, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480233.

Cressie:1978:FTC

[1006] Noel Cressie. A finely tuned continuity correction. Annals of the Institute of Statistical Mathematics, 30(1):435–442, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480234.

Barankin:1978:BRVb

[1007] Edward W. Barankin and Koiti Takahasi. Betweenness for real vectors and lines, II. Relatedness of betweennesses. Annals of the Institute of Statistical Mathematics, 30(1):443–464, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480235.

Morettin:1978:HSP

[1008] P. A. Morettin. On homogeneous stochastic processes on compact Abelian groups. Annals of the Institute of Statistical Mathematics, 30 (1):465–472, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480236.

Taguchi:1978:SCS

[1009] Tokio Taguchi. Some corrections and supplements “On a generalization of Gaussian distribution”. Annals of the Institute of Statistical Mathematics, 30 (1):473–475, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480237. See [987].

Chiu:1978:AP

[1010] W. K. Chiu. Acknowledgement of priority. Annals of the Institute of Statistical Mathematics, 30(1):477, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02480238.

Ishiguro:1978:SAC

[1011] Makio Ishiguro. A scheme of adaptive control. Annals of the Institute of Statistical Mathematics, 30(1):479–498, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480239.

Akaike:1978:CMC

[1012] Hirotugu Akaike. Covariance matrix computation of the state variable of a stationary Gaussian process. Annals of the Institute of Statistical Mathematics, 30(1):499–504, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480240.

Anonymous:1978:HC

[1013] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 30(1):??, ???? 1978. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hudimoto:1979:EBA

[1014] Hirosi Hudimoto. On the empirical Bayes approach to classification in the case of discrete multivariate distribution having only finite mass points. Annals of the Institute of Statistical Mathematics, 31(1):1–7, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480261.

Bhapkar:1979:NTS

[1015] V. P. Bhapkar. Nonparametric tests for scalar profile analysis of several multivariate samples. Annals of the Institute of Statistical Mathematics, 31(1):9–20, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480262.

Akritas:1979:AEL

[1016] M. G. Akritas and G. G. Roussas. Asymptotic expansion of the log-likelihood function based on stopping times defined on a Markov process. Annals of the Institute of Statistical Mathematics, 31(1):21–38, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480263.

Akahira:1979:DLM

[1017] Masafumi Akahira and Kei Takeuchi.

Discretized likelihood methods — asymptotic properties of discretized likelihood estimators (DLE’s). Annals of the Institute of Statistical Mathematics, 31(1):39–56, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480264.

Shirakura:1979:ABA

[1018] Teruhiro Shirakura. Alias balanced and alias partially balanced fractional 2^{m} factorial designs of resolution 2l + 1. Annals of the Institute of Statistical Mathematics, 31(1):57–65, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480265.

Shaked:1979:SCP

[1019] Moshe Shaked. Some concepts of positive dependence for bivariate interchangeable distributions. Annals of the Institute of Statistical Mathematics, 31(1):67–84, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480266.

Pillai:1979:ERS

[1020] K. C. S. Pillai and Yu-Sheng Hsu. Exact robustness studies of the test of independence based on four multivariate criteria and their distribution problems under violations. Annals of the Institute of Statistical Mathematics, 31(1):85–101, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480267.

Aickin:1979:EMD

[1021] M. Aickin. Existence of MLEs for discrete linear exponential models. Annals of the Institute of Statistical Mathematics, 31(1):103–113, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480268.

Tanaka:1979:OSA

[1022] Yutaka Tanaka. Optimal scaling for arbitrarily ordered categories. Annals of the Institute of Statistical Mathematics, 31(1):115–124, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480269.

Niki:1979:MFN

[1023] Naoto Niki. Multi-folding the normal distribution and mutual transformation between uniform and normal random variables. Annals of the Institute of Statistical Mathematics, 31 (1):125–140, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480270. See corrections

[1045].

Ray:1979:NRP

[1024] S. K. Ray, Ashok Sahai, and Ajit Sahai. A note on ratio and product type estimators. Annals of the Institute of Statistical Mathematics, 31(1):141–144, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480271.

Ozaki:1979:MLE

[1025] T. Ozaki. Maximum likelihood estimation of Hawkes’ self-exciting point processes. Annals of the Institute of Statistical Mathematics, 31(1):145–155, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480272.

Itoh:1979:RPM

[1026] Yoshiaki Itoh and Sumie Ueda. A random packing model for elections. Annals of the Institute of Statistical Mathematics, 31(1):157–167, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480273.

Nogami:1979:ESC

[1027] Yoshiko Nogami. The k-extended setcompound estimation problem in a nonregular family of distributions over [θ, θ + 1). Annals of the Institute of Statistical Mathematics, 31(1):169–176, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480274.

Takada:1979:SPP

[1028] Yoshikazu Takada. Stein’s positive part estimator and Bayes estimator. Annals of the Institute of Statistical Mathematics, 31(1):177–183, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480275.

Chu:1979:PCT

[1029] S. Sylvia Chu and K. C. S. Pillai. Power comparisons of two-sided tests of equality of two covariance matrices based on six criteria. Annals of the Institute of Statistical Mathematics, 31(1):185–205, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480276.

Gupta:1979:CZP

[1030] Rameshwar D. Gupta and Donald Richards. Calculation of zonal polynomials of 3 × 3 positive definite symmetric matrices. Annals of the Institute of Statistical Mathematics, 31(1):207–213, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480277.

Mathai:1979:DEP

[1031] A. M. Mathai and R. S. Katiyar.

The distribution and the exact percentage points for Wilks’ L_{mvc} criterion. Annals of the Institute of Statistical Mathematics, 31(1):215–224, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480278.

Pandey:1979:DSE

[1032] B. N. Pandey. Double stage estimation of population variance. Annals of the Institute of Statistical Mathematics, 31(1):225–233, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480279.

Gupta:1979:SSP

[1033] Shanti S. Gupta and Ming-Wei Lu.

Subset selection procedures for restricted families of probability distributions. Annals of the Institute of Statistical Mathematics, 31(1):235–252, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480280.

Koop:1979:SIS

[1034] J. C. Koop. On statistical inference in sample surveys and the underlying role of randomization. Annals of the Institute of Statistical Mathematics, 31(1):253–269, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480281.

Brillinger:1979:CFP

[1035] David R. Brillinger. A continuous form of post-stratification. Annals of the Institute of Statistical Mathematics, 31(1):271–277, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480282.

Ahmad:1979:SCD

[1036] Ibrahim A. Ahmad. Strong consistency of density estimation by orthogonal series methods for dependent variables with applications. Annals of the Institute of Statistical Mathematics, 31 (1):279–288, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480283.

Kirmani:1979:RBM

[1037] S. N. U. A. Kirmani. On the relation between Matusita’s and Kolmogorov’s measures of distance. Annals of the Institute of Statistical Mathematics, 31 (1):289–291, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480284.

Kageyama:1979:MEI

[1038] Sanpei Kageyama. Mathematical expression of an inequality for a block design. Annals of the Institute of Statistical Mathematics, 31(1):293–298, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480285.

deHaan:1979:BRS

[1039] Laurens de Haan and Elselien Taconis-Haantjes. On Bahadur’s representation of sample quantiles. Annals of the Institute of Statistical Mathematics, 31(1):299–308, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480286.

Shimizu:1979:LMP

[1040] R. Shimizu. On a lack of memory property of the exponential distribution. Annals of the Institute of Statistical Mathematics, 31(1):309–313, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480287.

Killeen:1979:LDD

[1041] Timothy J. Killeen. On large deviations and density functions. Annals of the Institute of Statistical Mathematics, 31(1):315–317, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480288.

Nakatsuka:1979:NEW

[1042] Toshinao Nakatsuka. Nonexistence of estimates which minimize x′V^{−1}x in an exponential type of stationary time series. Annals of the Institute of Statistical Mathematics, 31(1):319–320, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02480289.

Joshi:1979:NMO

[1043] P. C. Joshi. A note on the moments of order statistics from doubly truncated exponential distribution. Annals of the Institute of Statistical Mathematics, 31(1):321–324, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480290.

Hayashi:1979:NSM

[1044] Chikio Hayashi, Tsutomu Komazawa, and Fumi Hayashi. A new statistical method to estimate the size of animal population. Annals of the Institute of Statistical Mathematics, 31(1):325–348, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480291.

Niki:1979:CMF

[1045] Naoto Niki. Corrections to “Multifolding the normal distribution and mutual transformation between uniform and normal random variables”. Annals of the Institute of Statistical Mathematics, 31(1):349, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/

BF02480292. See [1023].

Tanemura:1979:RCP

[1046] Masaharu Tanemura. On random complete packing by discs. Annals of the Institute of Statistical Mathematics, 31 (1):351–365, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480293.

Shimizu:1979:CED

[1047] R. Shimizu. A characterization of the exponential distribution. Annals of the Institute of Statistical Mathematics, 31(1):367–372, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480294.

Sibuya:1979:GHD

[1048] Masaaki Sibuya. Generalized hypergeometric, digamma and trigamma distributions. Annals of the Institute of Statistical Mathematics, 31(1):373–390, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480295.

Lingappaiah:1979:BAP

[1049] G. S. Lingappaiah. Bayesian approach to prediction and the spacings in the exponential distribution. Annals of the Institute of Statistical Mathematics, 31(1):391–401, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480296.

Takeuchi:1979:AOG

[1050] Kei Takeuchi and Masafumi Akahira. Asymptotic optimality of the generalized Bayes estimator in multiparameter cases. Annals of the Institute of Statistical Mathematics, 31(1):403–415, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480297.

Campos:1979:EIV

[1051] Antonio Dorival Campos. Extension of the inequality for the variance of an estimator by Bayesian process. Annals of the Institute of Statistical Mathematics, 31(1):417–421, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480298.

Tong:1979:SNP

[1052] Y. L. Tong. Some new properties of the Bechhofer–Kiefer–Sobel stopping rule. Annals of the Institute of Statistical Mathematics, 31(1):423–433, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480299.

McDunnough:1979:EIP

[1053] Philip McDunnough. Estimating an interaction parameter of an infinite particle system. Annals of the Institute of Statistical Mathematics, 31(1):435–443, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480300.

Pillai:1979:DCR

[1054] K. C. S. Pillai and Yu-Sheng Hsu. The distribution of the characteristic roots of S_{1}S_{2}^{−1} under violations in the complex case and power comparisons of four tests. Annals of the Institute of Statistical Mathematics, 31(1):445–463, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480301.

Davis:1979:IPT

[1055] A. W. Davis. Invariant polynomials with two matrix arguments extending the zonal polynomials: Applications to multivariate distribution theory. Annals of the Institute of Statistical Mathematics, 31(1):465–485, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480302.

McDunnough:1979:SSS

[1056] Philip McDunnough and David B.

Wolfson. On some sampling schemes for estimating the parameters of a continuous time series. Annals of the Institute of Statistical Mathematics, 31 (1):487–497, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480303.

Rao:1979:EBM

[1057] B. L. S. Prakasa Rao. The equivalence between (modified) Bayes estimator and maximum likelihood estimator for Markov processes. Annals of the Institute of Statistical Mathematics, 31(1):499–513, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480304.

Srivastava:1979:CAM

[1058] M. S. Srivastava. Correction to “Asymptotically most powerful rank tests for regression parameters in MANOVA”. Annals of the Institute of Statistical Mathematics, 31(1):515–516, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02480305. See [700].

Anonymous:1979:HC

[1059] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 31(1):??, ???? 1979. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shimizu:1980:FEE

[1060] Ryoichi Shimizu. Functional equation with an error term and the stability of some characterizations of the exponential distribution. Annals of the Institute of Statistical Mathematics, 32(1):1–16, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480306.

Brandhofe:1980:FET

[1061] Thomas Brandhofe and Laurie Davies. On a functional equation in the theory of linear statistics. Annals of the Institute of Statistical Mathematics, 32(1):17–23, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480307.

Sibuya:1980:MDD

[1062] Masaaki Sibuya. Multivariate digamma distribution. Annals of the Institute of Statistical Mathematics, 32(1):25–36, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480308.

Jensen:1980:BJD

[1063] D. R. Jensen. Bounds on the joint distribution of T_{0}^{2} statistics. Annals of the Institute of Statistical Mathematics, 32(1):37–42, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480309.

Mudholkar:1980:TSM

[1064] Govind S. Mudholkar and Perla Subbaiah. Testing significance of a mean vector — a possible alternative to Hotelling’s T^{2}. Annals of the Institute of Statistical Mathematics, 32 (1):43–52, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480310.

Suzuki:1980:SGO

[1065] Giitiro Suzuki. On a stochastic game with one-chance recovery. Annals of the Institute of Statistical Mathematics, 32(1):53–64, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480311.

Walter:1980:CSD

[1066] Gilbert G. Walter. A class of spectral density estimators. Annals of the Institute of Statistical Mathematics, 32(1):65–80, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480312.

Yajima:1980:PIM

[1067] Yoshihiro Yajima. On prediction of integrated moving average processes. Annals of the Institute of Statistical Mathematics, 32(1):81–94, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480313.

Wright:1980:NCS

[1068] A. L. Wright. Nonexistence of complete sufficient statistics for stationary k-state Markov chains, k ≥ 3. Annals of the Institute of Statistical Mathematics, 32(1):95–97, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480314.

Duthie:1980:SHA

[1069] Audrey I. Duthie. The split hypercubic association scheme. Annals of the Institute of Statistical Mathematics, 32(1):99–106, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480315.

Kageyama:1980:CCB

[1070] Sanpei Kageyama. Characterization of certain balanced n-ary block designs. Annals of the Institute of Statistical Mathematics, 32(1):107–110, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480316.

Srivenkataramana:1980:ARM

[1071] T. Srivenkataramana and D. S. Tracy.

An alternative to ratio method in sample surveys. Annals of the Institute of Statistical Mathematics, 32(1):111–120, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480317.

Drossos:1980:NMD

[1072] Constantine A. Drossos and Andreas N. Philippou. A note on minimum distance estimates. Annals of the Institute of Statistical Mathematics, 32(1):121–123, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480318.

Otake:1980:CTR

[1073] Masanori Otake. Comparison of time risks based on a multinomial logistic response model in longitudinal studies. Annals of the Institute of Statistical Mathematics, 32(1):125–142, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480319.

Suzuki:1980:FMF

[1074] Giitiro Suzuki. Further modified forms of binomial and Poisson distributions. Annals of the Institute of Statistical Mathematics, 32(1):143–159, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480320.

George:1980:GLD

[1075] E. Olusegun George and M. O. Ojo. On a generalization of the logistic distribution. Annals of the Institute of Statistical Mathematics, 32(1):161–169, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480321.

Akaike:1980:IPD

[1076] Hirotugu Akaike. Ignorance prior distribution of a hyperparameter and Stein’s estimator. Annals of the Institute of Statistical Mathematics, 32 (1):171–178, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480322.

Rehder:1980:WPB

[1077] Wulf Rehder. When is the pseudo-best estimator blue? Annals of the Institute of Statistical Mathematics, 32 (1):179–185, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480323.

Cheng:1980:UCC

[1078] K. F. Cheng and R. L. Taylor.

On the uniform complete convergence of estimates for multivariate density functions and regression curves. Annals of the Institute of Statistical Mathematics, 32(1):187–199, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480324.

Nagao:1980:SPE

[1079] Hisao Nagao and Michio Takada.

On sequential point estimation of the mean of a normal distribution. Annals of the Institute of Statistical Mathematics, 32(1):201–210, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480325.

Ahmad:1980:CST

[1080] Ibrahim A. Ahmad and Pi-Erh Lin. On the Chernoff–Savage theorem for dependent sequences. Annals of the Institute of Statistical Mathematics, 32 (1):211–222, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480326.

Ahmad:1980:NEA

[1081] Ibrahim A. Ahmad. Nonparametric estimation of an affinity measure between two absolutely continuous distributions with hypotheses testing applications. Annals of the Institute of Statistical Mathematics, 32(1):223–240, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480327.

Ahmad:1980:NEM

[1082] Ibrahim A. Ahmad. Nonparametric estimation of Matusita’s measure of affinity between absolutely continuous distributions. Annals of the Institute of Statistical Mathematics, 32(1):241–245, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480328.

Khan:1980:DDF

[1083] A. H. Khan and Mohd Yaqub. Distribution of a distance function. Annals of the Institute of Statistical Mathematics, 32(1):247–253, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480329.

Kageyama:1980:RCB

[1084] Sanpei Kageyama. Robustness of connected balanced block designs. Annals of the Institute of Statistical Mathematics, 32(1):255–261, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480330.

Kageyama:1980:CEV

[1085] Sanpei Kageyama and Takumi Tsuji. Characterization of equireplicated variance-balanced block designs. Annals of the Institute of Statistical Mathematics, 32(1):263–273, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480331.

Chang:1980:DOC

[1086] Der-Shin Chang and Yuang-Chin Chiang. Designs of φ-optimal control for second-order processes. Annals of the Institute of Statistical Mathematics, 32 (1):275–281, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480332.

Tweedie:1980:PCM

[1087] R. L. Tweedie. Perturbations of countable Markov chains and processes. Annals of the Institute of Statistical Mathematics, 32(1):283–290, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480333.

Yamazoe:1980:DRP

[1088] Shiro Yamazoe, Kazuo Fukutomi, and Takemi Yanagimoto. Dose-response problems in drug-induced diseases after consecutive intake of long duration. Annals of the Institute of Statistical Mathematics, 32(1):291–301, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480334.

Wakimoto:1980:SCM

[1089] Kazumasa Wakimoto. Sun chart method for looking multivariate data. Annals of the Institute of Statistical Mathematics, 32(1):303–310, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480335.

Akaike:1980:UPL

[1090] Hirotugu Akaike. On the use of the predictive likelihood of a Gaussian model. Annals of the Institute of Statistical Mathematics, 32(1):311–324, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480336.

Yanagimoto:1980:CTD

[1091] Takemi Yanagimoto and Masaaki Sibuya. Comparison of tails of distributions in models for estimating safe doses. Annals of the Institute of Statistical Mathematics, 32(1):325–340, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480337.

Ghorai:1980:ANQ

[1092] Jugal Ghorai. Asymptotic normality of a quadratic measure of orthogonal series type density estimate. Annals of the Institute of Statistical Mathematics, 32(1):341–350, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480338.

Hall:1980:EDP

[1093] Peter Hall. Estimating a density on the positive half line by the method of orthogonal series. Annals of the Institute of Statistical Mathematics, 32 (1):351–362, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480339.

Ahsanullah:1980:LPR

[1094] M. Ahsanullah. Linear prediction of record values for the two parameter exponential distribution. Annals of the Institute of Statistical Mathematics, 32(1):363–368, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480340.

Nagao:1980:STS

[1095] Hisao Nagao. On stopping times of sequential estimations of the mean of a log-normal distribution. Annals of the Institute of Statistical Mathematics, 32(1):369–375, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480341.

Skibinsky:1980:MPI

[1096] Morris Skibinsky. The minimum probability on an interval when the mean and variance are known. Annals of the Institute of Statistical Mathematics, 32(1):377–385, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480342.

Srivastava:1980:DDC

[1097] V. K. Srivastava, B. S. Agnihotri, and T. D. Dwivedi. Dominance of double k-class estimators in simultaneous equations. Annals of the Institute of Statistical Mathematics, 32(1):387–392, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480343.

Sakai:1980:FAR

[1098] Hideaki Sakai. Fitting autoregression with regularly missed observations. Annals of the Institute of Statistical Mathematics, 32(1):393–400, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480344.

Taniguchi:1980:SOS

[1099] Masanobu Taniguchi. On selection of the order of the spectral density model for a stationary process. Annals of the Institute of Statistical Mathematics, 32(1):401–419, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480345.

Gulati:1980:SCD

[1100] Chandra M. Gulati. Selection of certain dichotomous experiments. Annals of the Institute of Statistical Mathematics, 32(1):421–431, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480346.

Dayal:1980:ASU

[1101] Shambhu Dayal. On allocation of sample using estimates of both proportions of stratum sizes and standard deviations. Annals of the Institute of Statistical Mathematics, 32(1):433–444, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480347.

Pandian:1980:CPB

[1102] V. S. Soundara Pandian. Construction of partially balanced n-ary designs using difference sets. Annals of the Institute of Statistical Mathematics, 32(1):445–464, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480348.

Yanagimoto:1980:CME

[1103] Takemi Yanagimoto and David G.

Hoel. Comparisons of models for estimation of safe doses using measures of the heaviness of tail of a distribution. Annals of the Institute of Statistical Mathematics, 32(1):465–480, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480349.

Akaike:1980:TEM

[1104] Hirotugu Akaike and Makio Ishiguro.

Trend estimation with missing observations. Annals of the Institute of Statistical Mathematics, 32(1):481–488, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480350.

Anonymous:1980:HC

[1105] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 32(1):??, ???? 1980. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aki:1981:ADC

[1106] Sigeo Aki. Asymptotic distribution of a Cramér–von Mises type statistic for testing symmetry when the center is estimated. Annals of the Institute of Statistical Mathematics, 33(1):1–14, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480914.

Hayakawa:1981:ADG

[1107] Takesi Hayakawa. Asymptotic distribution of a generalized Hotelling’s T_{0}^{2} in the doubly noncentral case. Annals of the Institute of Statistical Mathematics, 33(1):15–25, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480915.

Konishi:1981:IAD

[1108] Sadanori Konishi and Takakazu Sugiyama. Improved approximations to distributions of the largest and the smallest latent roots of a Wishart matrix. Annals of the Institute of Statistical Mathematics, 33(1):27–33, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480916.

Mathai:1981:DCC

[1109] A. M. Mathai. Distribution of the canonical correlation matrix. Annals of the Institute of Statistical Mathematics, 33(1):35–43, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480917.

Pillai:1981:ENN

[1110] K. C. S. Pillai and Antia Singh.

On the exact non-null distribution of Wilks’ L_{V C} criterion and power studies. Annals of the Institute of Statistical Mathematics, 33(1):45–55, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480918.

Al-Hussaini:1981:BIG

[1111] Essam K. Al-Hussaini and Nagi S.

Abd-El-Hakim. Bivariate inverse Gaussian distribution. Annals of the Institute of Statistical Mathematics, 33 (1):57–66, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480919.

Nogami:1981:SCO

[1112] Yoshiko Nogami. The set-compound one-stage estimation in the nonregular family of distributions over the interval [θ, θ + 1). Annals of the Institute of Statistical Mathematics, 33(1):67–80, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480920.

Umbach:1981:BES

[1113] Dale Umbach. Bayes estimation with spherically symmetric, convex loss. Annals of the Institute of Statistical Mathematics, 33(1):81–90, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480921.

Takada:1981:IPR

[1114] Yoshikazu Takada. Invariant prediction rules and an adequate statistic. Annals of the Institute of Statistical Mathematics, 33(1):91–100, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480922.

Arimoto:1981:ABD

[1115] Akio Arimoto. Asymptotic behavior of difference between a finite predictor and an infinite predictor for a weakly stationary stochastic process. Annals of the Institute of Statistical Mathematics, 33(1):101–113, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480923.

Yamazaki:1981:ORB

[1116] Genji Yamazaki. An ordering relation of the blocking two-stage tandem queueing system to the reduced single server queueing system. Annals of the Institute of Statistical Mathematics, 33(1):115–123, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480924.

George:1981:CLD

[1117] E. Olusegun George and Govind S. Mudholkar. A characterization of the logistic distribution by a sample median. Annals of the Institute of Statistical Mathematics, 33(1):125–129, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480925.

Huang:1981:LMP

[1118] J. S. Huang. On a “lack of memory”

property. Annals of the Institute of Statistical Mathematics, 33(1):131–134, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480926.

Umbach:1981:NMD

[1119] Dale Umbach. A note on the median of a distribution. Annals of the Institute of Statistical Mathematics, 33 (1):135–140, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480927.

Kageyama:1981:SBP

[1120] Sanpei Kageyama. Some bounds for partially balanced block designs. Annals of the Institute of Statistical Mathematics, 33(1):141–153, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480928.

Cheng:1981:CPD

[1121] Ching-Shui Cheng. On the comparison of PBIB designs with two associate classes. Annals of the Institute of Statistical Mathematics, 33(1):155–164, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480929.

Dayal:1981:SEW

[1122] Shambhu Dayal. Sampling for estimating weighted totals and averages. Annals of the Institute of Statistical Mathematics, 33(1):165–176, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480930.

Sibuya:1981:GHF

[1123] Masaaki Sibuya and Ryoichi Shimizu. The generalized hypergeometric family of distributions. Annals of the Institute of Statistical Mathematics, 33 (1):177–190, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480931.

Panaretos:1981:JDT

[1124] John Panaretos. On the joint distribution of two discrete random variables. Annals of the Institute of Statistical Mathematics, 33(1):191–198, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480932.

Richards:1981:STS

[1125] Donald St. P. Richards. Stability theorems for some characterizations of the exponential distribution. Annals of the Institute of Statistical Mathematics, 33(1):199–204, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480933.

Grosswald:1981:ILM

[1126] E. Grosswald and Samuel Kotz. An integrated lack of memory characterization of the exponential distribution. Annals of the Institute of Statistical Mathematics, 33(1):205–214, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480934.

Tsui:1981:SES

[1127] Kam-Wah Tsui. Simultaneous estimation of several Poisson parameters under squared error loss. Annals of the Institute of Statistical Mathematics, 33(1):215–223, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480935.

Rao:1981:CAU

[1128] T. J. Rao. On a class of almost unbiased ratio estimators. Annals of the Institute of Statistical Mathematics, 33(1):225–231, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480936.

Rust:1981:CKE

[1129] Albert E. Rust and Chris P. Tsokos.

On the convergence of kernel estimators of probability density functions. Annals of the Institute of Statistical Mathematics, 33(1):233–246, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480937.

Ahmad:1981:NND

[1130] Ibrahim A. Ahmad. A note on nonparametric density estimation for dependent variables using a delta sequence. Annals of the Institute of Statistical Mathematics, 33(1):247–254, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480938.

Aiyar:1981:AER

[1131] R. J. Aiyar. Asymptotic efficiency of rank tests of randomness against autocorrelation. Annals of the Institute of Statistical Mathematics, 33(1):255–262, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480939.

Venkataraman:1981:SLT

[1132] K. N. Venkataraman. Some limit theorems on an explosive model for time series, and their statistical applications. Annals of the Institute of Statistical Mathematics, 33(1):263–278, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480940.

Fujikoshi:1981:PLR

[1133] Yasunori Fujikoshi. The power of the likelihood ratio test for additional information in a multivariate linear model. Annals of the Institute of Statistical Mathematics, 33(1):279–285, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480941.

Bhargava:1981:DPI

[1134] R. P. Bhargava and C. G. Khatri.

The distribution of product of independent beta random variables with application to multivariate analysis. Annals of the Institute of Statistical Mathematics, 33(1):287–296, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480942.

Davis:1981:CCI

[1135] A. W. Davis. On the construction of a class of invariant polynomials in several matrices, extending the zonal polynomials. Annals of the Institute of Statistical Mathematics, 33(1):297–313, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480943.

Ogata:1981:EIP

[1136] Yosihiko Ogata and Masaharu Tanemura. Estimation of interaction potentials of spatial point patterns through the maximum likelihood procedure. Annals of the Institute of Statistical Mathematics, 33(1):315–338, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480944.

Shimizu:1981:SCE

[1137] Ryoichi Shimizu. On the stability of characterizations of the exponential distribution. Annals of the Institute of Statistical Mathematics, 33(1):339–346, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480945.

Ruschendorf:1981:CDC

[1138] Ludger Rüschendorf. Characterization of dependence concepts in normal distributions. Annals of the Institute of Statistical Mathematics, 33(1):347–359, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480946.

Lwin:1981:MPS

[1139] T. Lwin. A modified power series distribution. Annals of the Institute of Statistical Mathematics, 33(1):361–374, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480947.

Ahmad:1981:MNA

[1140] Ibrahim A. Ahmad. Mixing normal approximations of vectors of sums and maximum sums. Annals of the Institute of Statistical Mathematics, 33 (1):375–383, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480948.

Menzefricke:1981:PDQ

[1141] Ulrich Menzefricke. On positive definite quadratic forms in correlated t variables. Annals of the Institute of Statistical Mathematics, 33(1):385–390, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480949.

Aki:1981:ABF

[1142] Sigeo Aki. Asymptotic behavior of functionals of empirical distribution functions for the two-sample problem. Annals of the Institute of Statistical Mathematics, 33(1):391–403, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480950.

Chen:1981:SLL

[1143] Wen Chen Chen. Some local limit theorems in the symmetric Dirichlet– multinomial urn models. Annals of the Institute of Statistical Mathematics, 33(1):405–415, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480951.

Csorgo:1981:SAE

[1144] Sándor Csörgő. Strong approximation of empirical Kac processes. Annals of the Institute of Statistical Mathematics, 33(1):417–423, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480952.

Blum:1981:MDT

[1145] J. Blum and V. Susarla. Maximal deviation theory of some estimators of prior distribution functions. Annals of the Institute of Statistical Mathematics, 33(1):425–436, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480953.

Nagaraja:1981:SFS

[1146] H. N. Nagaraja. Some finite sample results for the selection differential. Annals of the Institute of Statistical Mathematics, 33(1):437–448, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480954.

Hall:1981:LSP

[1147] Peter Hall. Large sample properties of Jaeckel’s adaptive trimmed mean. Annals of the Institute of Statistical Mathematics, 33(1):449–462, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480955.

Yang:1981:LCC

[1148] Shie-Shien Yang. Linear combination of concomitants of order statistics with application to testing and estimation. Annals of the Institute of Statistical Mathematics, 33(1):463–470, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480956.

Feigin:1981:IZA

[1149] Paul D. Feigin and Ishay Weissman. On the indifference zone approach to selection — a consistency result. Annals of the Institute of Statistical Mathematics, 33(1):471–474, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480957.

Otake:1981:CRR

[1150] Masanori Otake. Comparison of relative risk, attributable risk and logistic response procedures for 2 × 2 × 2 and c 2 × 2 × 2 contingency tables. Annals of the Institute of Statistical Mathematics, 33(1):475–486, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480958.

Anonymous:1981:HC

[1151] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 33(1):??, ???? 1981. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matsunawa:1982:UEP

[1152] T. Matsunawa. Uniform θ-equivalence of probability distributions based on information and related measures of discrepancy. Annals of the Institute of Statistical Mathematics, 34(1):1–17, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481004.

Kuboki:1982:UESa

[1153] Hisataka Kuboki. Unbiased estimators in the sense of Lehmann and their discrimination rates. Annals of the Institute of Statistical Mathematics, 34(1):19–37, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481005.

Ahmad:1982:NEL

[1154] Ibrahim A. Ahmad. Nonparametric estimation of the location and scale parameters based on density estimation. Annals of the Institute of Statistical Mathematics, 34(1):39–53, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481006.

Withers:1982:DQF

[1155] C. S. Withers. The distribution and quantiles of a function of parameter estimates. Annals of the Institute of Statistical Mathematics, 34(1):55–68, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481007.

Akahira:1982:AOE

[1156] Masafumi Akahira. Asymptotic optimality of estimators in non-regular cases. Annals of the Institute of Statistical Mathematics, 34(1):69–82, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481008.

Nakano:1982:PEA

[1157] Junji Nakano. Parameter estimation of an autoregressive moving average model. Annals of the Institute of Statistical Mathematics, 34(1):83–90, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481009.

Jain:1982:ESR

[1158] Aridaman K. Jain. Estimation in stratified random sampling: Adjustment for changes in strata composition. Annals of the Institute of Statistical Mathematics, 34(1):91–103, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481010.

Hudlet:1982:ESO

[1159] Raul Hudlet and Richard A. Johnson. An extension of some optimal properties of principal components. Annals of the Institute of Statistical Mathematics, 34(1):105–110, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481011.

Richards:1982:DOAa

[1160] Donald St. P. Richards. Differential operators associated with zonal polynomials. I. Annals of the Institute of Statistical Mathematics, 34(1):111–117, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481012.

Richards:1982:DOAb

[1161] Donald St. P. Richards. Differential operators associated with zonal polynomials. II. Annals of the Institute of Statistical Mathematics, 34(1):119–121, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481013.

Stone:1982:LAA

[1162] Charles J. Stone. Local asymptotic admissibility of a generalization of Akaike’s model selection rule. Annals of the Institute of Statistical Mathematics, 34(1):123–133, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481014.

Cooke:1982:SPC

[1163] P. J. Cooke and M. K. Vagholkar. Sequential procedures for a class of distributions related to the uniform. Annals of the Institute of Statistical Mathematics, 34(1):135–142, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481015.

Bansal:1982:SBD

[1164] Ashok K. Bansal. Sensitivity of Bayes decisions for the success probability for samples from a nonbinomial population. Annals of the Institute of Statistical Mathematics, 34(1):143–149, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481016.

Trenkler:1982:PSU

[1165] G. Trenkler. Partitions, sufficiency and undominated families of probability measures. Annals of the Institute of Statistical Mathematics, 34(1):151–160, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481017.

Bhattacharya:1982:UBP

[1166] Binay K. Bhattacharya and Godfried T. Toussaint. An upper bound on the probability of misclassification in terms of Matusita’s measure of affinity. Annals of the Institute of Statistical Mathematics, 34(1):161–165, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481018.

Ruben:1982:EDK

[1167] Harold Ruben and Jack Gambino.

The exact distribution of Kolmogorov’s statistic D_{n} for n ≤ 10. Annals of the Institute of Statistical Mathematics, 34(1):167–173, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481019.

Stepniak:1982:OAO

[1168] C. Stepniak. Optimal allocation of observations in one-way random normal model. Annals of the Institute of Statistical Mathematics, 34(1):175–179, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481020.

Roy:1982:AFB

[1169] F. Roy and Gurajada S. Murty. Application of FPE based AR model for signal detection and digital data compression. Annals of the Institute of Statistical Mathematics, 34(1):181–188, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481021.

Vere-Jones:1982:SES

[1170] D. Vere-Jones and T. Ozaki. Some examples of statistical estimation applied to earthquake data. Annals of the Institute of Statistical Mathematics, 34(1):189–207, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481022. See corrections

[1419].

Matsunawa:1982:SSE

[1171] T. Matsunawa. Some strong ϵequivalence of random variables. Annals of the Institute of Statistical Mathematics, 34(1):209–224, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481023.

Rivest:1982:SAD

[1172] Louis-Paul Rivest. Some asymptotic distributions in the location-scale model. Annals of the Institute of Statistical Mathematics, 34(1):225–239, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481024.

Nogami:1982:RCS

[1173] Yoshiko Nogami. A rate of convergence for the set compound estimation in a family of certain retracted distributions. Annals of the Institute of Statistical Mathematics, 34(1):241–257, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481025.

Akritas:1982:ATE

[1174] Michael G. Akritas. Asymptotic theory for estimating the parameters of a Lévy process. Annals of the Institute of Statistical Mathematics, 34(1):259–280, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481026.

Kariya:1982:MAP

[1175] Takeaki Kariya and Koichi Maekawa. A method for approximations to the pdf’s and cdf’s of GLSE’s and its application to the seemingly unrelated regression model. Annals of the Institute of Statistical Mathematics, 34(1):281–297, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481027.

Szatrowski:1982:REE

[1176] Ted H. Szatrowski. Relative efficiencies of estimates using patterned covariances or correlations in the multivariate normal estimation problem. Annals of the Institute of Statistical Mathematics, 34(1):299–307, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481028.

Esimal:1982:MAR

[1177] Grace Esimal and Chien-Pai Han.

Multi-auxiliary regression estimation based on conditional specification. Annals of the Institute of Statistical Mathematics, 34(1):309–317, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481029.

Guttman:1982:ULF

[1178] Irwin Guttman and Ulrich Menzefricke. On the use of loss functions in the changepoint problem. Annals of the Institute of Statistical Mathematics, 34(1):319–326, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481030.

Rao:1982:NTH

[1179] K. S. Madhava Rao. Nonparametric tests for homogeneity of scale against ordered alternatives. Annals of the Institute of Statistical Mathematics, 34 (1):327–334, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481031.

Preda:1982:SDP

[1180] Vasile C. Preda. The Student distribution and the principle of maximum entropy. Annals of the Institute of Statistical Mathematics, 34(1):335–338, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481032.

Ahmad:1982:IMS

[1181] Ibrahim A. Ahmad. Integrated mean square properties of density estimation by orthogonal series methods for dependent variables. Annals of the Institute of Statistical Mathematics, 34 (1):339–350, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481033.

Ahmad:1982:DBD

[1182] Ibrahim A. Ahmad. On the deviation between the distributions of sums and maximum sums of IID random vectors. Annals of the Institute of Statistical Mathematics, 34(1):351–358, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481034.

Lai:1982:CGM

[1183] C. D. Lai. A characterization of gamma, Meixner hypergeometric and negative binomial distributions based on canonical measures. Annals of the Institute of Statistical Mathematics, 34 (1):359–363, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481035.

Huda:1982:STO

[1184] S. Huda. Some third-order rotatable designs in three dimensions. Annals of the Institute of Statistical Mathematics, 34(1):365–371, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481036.

Ogata:1982:ALI

[1185] Y. Ogata, H. Akaike, and K. Katsura. The application of linear intensity models to the investigation of causal relations between a point process and another stochastic process.

Annals of the Institute of Statistical Mathematics, 34(1):373–387, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481037.

Kitagawa:1982:QBA

[1186] Genshiro Kitagawa and Hirotugu Akaike. A quasi Bayesian approach to outlier detection. Annals of the Institute of Statistical Mathematics, 34 (1):389–398, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481038.

Kuboki:1982:UESb

[1187] Hisataka Kuboki. Unbiased estimators in the sense of Lehmann and their discrimination rates (II): Multi-parameter cases. Annals of the Institute of Statistical Mathematics, 34(1):399–409, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481039.

deOliveira:1982:DEE

[1188] J. Tiago de Oliveira. A definition of estimator efficiency in k-parameter case. Annals of the Institute of Statistical Mathematics, 34(1):411–421, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481040.

Ferreira:1982:MEE

[1189] Pedro E. Ferreira. Multiparametric estimating equations. Annals of the Institute of Statistical Mathematics, 34 (1):423–431, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481041.

Hwang:1982:EEP

[1190] Jiunn Tzon Hwang. Examples of estimation problems. Annals of the Institute of Statistical Mathematics, 34 (1):433–439, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481042.

Lin:1982:PBM

[1191] Pi-Erh Lin and Amany Mousa. Proper Bayes minimax estimators for a multivariate normal mean with unknown common variance under a convex loss function. Annals of the Institute of Statistical Mathematics, 34(1):441–456, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481043.

Ahmad:1982:CMD

[1192] Ibrahim A. Ahmad. L_{p}-consistency of multivariate density estimates. Annals of the Institute of Statistical Mathematics, 34(1):457–466, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481044.

Ghorai:1982:NQM

[1193] J. K. Ghorai. A note on a quadratic measure of deviation of density estimates. Annals of the Institute of Statistical Mathematics, 34(1):467–477, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481045.

Cheng:1982:EDD

[1194] K. F. Cheng. On estimation of a density and its derivatives. Annals of the Institute of Statistical Mathematics, 34(1):479–489, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481046.

Takada:1982:AAS

[1195] Yoshikazu Takada. Application of an adequate statistic to the invariant prediction region. Annals of the Institute of Statistical Mathematics, 34 (1):491–503, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481047.

Konishi:1982:APE

[1196] Sadanori Konishi. Asymptotic properties of estimators of interclass correlation from familial data. Annals of the Institute of Statistical Mathematics, 34(1):505–515, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481048.

Davis:1982:RRG

[1197] A. W. Davis. On a result of Roy and Gnanadesikan concerning multivariete variance components. Annals of the Institute of Statistical Mathematics, 34(1):517–521, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481049.

Fujikoshi:1982:TAI

[1198] Yasunori Fujikoshi. A test for additional information in canonical correlation analysis. Annals of the Institute of Statistical Mathematics, 34(1):523–530, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481050.

Isogai:1982:MMS

[1199] Takafumi Isogai. On a measure of multivariate skewness and a test for multivariate normality. Annals of the Institute of Statistical Mathematics, 34 (1):531–541, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481051.

Ahmad:1982:RNI

[1200] Khalaf E. Ahmad and Essam K.

Al-Hussaini. Remarks on the nonidentifiability of mixtures of distributions. Annals of the Institute of Statistical Mathematics, 34(1):543–544, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02481052.

Obremski:1982:LMF

[1201] Thomas E. Obremski. Locating the minimum of a function when the errors of observation have unknown density. Annals of the Institute of Statistical Mathematics, 34(1):545–558, ????

1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481053.

Klein:1982:HTC

[1202] Steven W. Klein. Hypothesis testing for the common mean of two normal distributions in the presence of an indifference zone. Annals of the Institute of Statistical Mathematics, 34(1):559–577, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481054.

Akritas:1982:ETE

[1203] Michael G. Akritas and Richard A. Johnson. Efficiencies of tests and estimators for p-order autoregressive processes when the error distribution is nonnormal. Annals of the Institute of Statistical Mathematics, 34(1):579–589, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481055.

Mathai:1982:SCD

[1204] A. M. Mathai. Storage capacity of a dam with gamma type inputs. Annals of the Institute of Statistical Mathematics, 34(1):591–597, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481056.

Yamazaki:1982:GGQ

[1205] Genji Yamazaki. The GI/G/1 queue with last-come-first-served. Annals of the Institute of Statistical Mathematics, 34(1):599–604, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481057.

Sekkappan:1982:BSR

[1206] R. M. Sekkappan. Bayesian stratified random sampling using auxiliary information. Annals of the Institute of Statistical Mathematics, 34(1):605–609, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481058.

Anonymous:1982:HC

[1207] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 34(1):??, ???? 1982. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Amari:1983:DGE

[1208] Shun ichi Amari and Masayuki Kumon. Differential geometry of Edgeworth expansions in curved exponential family. Annals of the Institute of Statistical Mathematics, 35(1):1–24, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480959.

Tsui:1983:CAE

[1209] K. W. Tsui. A class of admissible estimators of a finite population total. Annals of the Institute of Statistical Mathematics, 35(1):25–30, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480960.

vanderMerwe:1983:AEW

[1210] A. J. van der Merwe. The asymptotic expansion as well as the exact moments of the Stein estimator when the population means are nearly equal. Annals of the Institute of Statistical Mathematics, 35(1):31–39, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480961.

Ogawa:1983:CBR

[1211] Junjiro Ogawa. On the ‘confidence bounds” of the ratio of the means of a bivariate normal distribution. Annals of the Institute of Statistical Mathematics, 35(1):41–48, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480962.

Withers:1983:ACI

[1212] C. S. Withers. Accurate confidence intervals for distributions with one parameter. Annals of the Institute of Statistical Mathematics, 35(1):49–61, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480963.

Sarkar:1983:STU

[1213] S. K. Sarkar, B. K. Sinha, and P. R. Krishnaiah. Some tests with unbalanced data from a bivariate normal population. Annals of the Institute of Statistical Mathematics, 35(1):63–75, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480964.

Hirotsu:1983:ADP

[1214] C. Hirotsu. An approach to defining the pattern of interaction effects in a twoway layout. Annals of the Institute of Statistical Mathematics, 35(1):77–90, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480965.

Shimizu:1983:CPU

[1215] R. Shimizu and J. S. Huang. On a characteristic property of the uniform distribution. Annals of the Institute of Statistical Mathematics, 35(1):91–94, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480966.

Cacoullos:1983:CDD

[1216] T. Cacoullos and H. Papageorgiou.

Characterizations of discrete distributions by a conditional distribution and a regression function. Annals of the Institute of Statistical Mathematics, 35(1):95–103, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480967.

Klebanov:1983:MAC

[1217] L. B. Klebanov and J. A. Melamed.

A method associated with characterizations of the exponential distribution. Annals of the Institute of Statistical Mathematics, 35(1):105–114, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480968.

Ishiguro:1983:BAB

[1218] Makio Ishiguro and Yosiyuki Sakamoto. A Bayesian approach to binary response curve estimation. Annals of the Institute of Statistical Mathematics, 35(1):115–137, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480969.

Akaike:1983:MIP

[1219] Hirotugu Akaike. On minimum information prior distributions. Annals of the Institute of Statistical Mathematics, 35(1):139–149, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480970.

Inagaki:1983:DFI

[1220] Nobuo Inagaki. The decomposition of the Fisher information. Annals of the Institute of Statistical Mathematics, 35(1):151–165, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480971.

Sono:1983:NPD

[1221] Shintaro Sono. On a noniformative prior distribution for Bayesian inference of multinomial distribution’s parameters. Annals of the Institute of Statistical Mathematics, 35(1):167–174, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480972.

Murakami:1983:SPR

[1222] Masakatsu Murakami. Some properties of the risk set in multiple decision problems. Annals of the Institute of Statistical Mathematics, 35(1):175–183, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480973.

Sono:1983:AMS

[1223] Shintaro Sono. On an approximation for a multi-stage decision problem. Annals of the Institute of Statistical Mathematics, 35(1):185–191, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480974.

Nagata:1983:AEO

[1224] Yasushi Nagata. An admissible estimator in the one-parameter exponential family with ambiguous information. Annals of the Institute of Statistical Mathematics, 35(1):193–199, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480975.

Isokawa:1983:EFR

[1225] Y. Isokawa. Estimation of frequency by random sampling. Annals of the Institute of Statistical Mathematics, 35 (1):201–213, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480976.

Greblicki:1983:OSE

[1226] Wlodzimierz Greblicki, Danuta Rutkowska, and Leszek Rutkowski. An orthogonal series estimate of time-varying regression. Annals of the Institute of Statistical Mathematics, 35(1):215–228, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480977.

Henna:1983:NCE

[1227] Jogi Henna. A note on a consistent estimator of a mixing distribution function. Annals of the Institute of Statistical Mathematics, 35(1):229–233, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480978.

Gutmann:1983:MRR

[1228] Sam Gutmann. A minimax result related to Stein’s effect. Annals of the Institute of Statistical Mathematics, 35 (1):235–241, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480979.

Keilson:1983:DBD

[1229] Julian Keilson and Ushio Sumita. A decomposition of the beta distribution, related order and asymptotic behavior. Annals of the Institute of Statistical Mathematics, 35(1):243–253, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480980.

Smith:1983:IFO

[1230] Norman L. Smith and Y. L. Tong.

Inequalities for functions of order statistics under an additive model. Annals of the Institute of Statistical Mathematics, 35(1):255–265, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480981.

Dunsmore:1983:FOR

[1231] Ian R. Dunsmore. The future occurrence of records. Annals of the Institute of Statistical Mathematics, 35 (1):267–277, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480982.

Xekalaki:1983:IDC

[1232] Evdokia Xekalaki. Infinite divisibility, completeness and regression properties of the univariate generalized Waring distribution. Annals of the Institute of Statistical Mathematics, 35(1):279–289, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480983.

Sakasegawa:1983:SRS

[1233] H. Sakasegawa. Stratified rejection and squeeze method for generating beta random numbers. Annals of the Institute of Statistical Mathematics, 35 (1):291–302, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480984.

Watson:1983:LST

[1234] Geoffrey S. Watson. Large sample theory for distributions on the hypersphere with rotational symmetries. Annals of the Institute of Statistical Mathematics, 35(1):303–319, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480985.

Yamamoto:1983:JMN

[1235] Eiji Yamamoto, Kazumasa Wakimoto, and Seiji Nabeya. Joint moments of the number of + runs and the number of + signs in a random sequence. Annals of the Institute of Statistical Mathematics, 35(1):321–328, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480986.

Ikeda:1983:UAJ

[1236] Sadao Ikeda and Yoshiyuki Nonaka. Uniform asymptotic joint normality of a set of increasing number of sample quantiles. Annals of the Institute of Statistical Mathematics, 35(1):329–341, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480987.

Kohne:1983:NUA

[1237] W. Kohne and R.-D. Reiss. A note on uniform approximation to distributions of extreme order statistics. Annals of the Institute of Statistical Mathematics, 35(1):343–345, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480988.

Anonymous:1983:CQF

[1238] Anonymous. Corrections to “Quadratic forms of a matric-t variate”. Annals of the Institute of Statistical Mathematics, 35(1):346, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02480989. See

[1239].

Marx:1983:QFM

[1239] D. G. Marx. Quadratic forms of a matric-t variate. Annals of the Institute of Statistical Mathematics, 35 (1):347–353, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480990. See corrections

[1238, 1348].

Alamatsaz:1983:CSD

[1240] M. H. Alamatsaz. Completeness and self-decomposability of mixtures. Annals of the Institute of Statistical Mathematics, 35(1):355–363, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480991.

Nagata:1983:ASP

[1241] Yasushi Nagata. Admissibility of some preliminary test estimators for the mean of normal distribution. Annals of the Institute of Statistical Mathematics, 35(1):365–373, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480992.

Steniak:1983:LBR

[1242] Czeslaw Steniak. Lower bound of risk in linear unbiased estimation and its application. Annals of the Institute of Statistical Mathematics, 35(1):375–378, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480993.

Ghosh:1983:SEP

[1243] Malay Ghosh and Richard Auer. Simultaneous estimation of parameters in exponential families. Annals of the Institute of Statistical Mathematics, 35 (1):379–387, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02480994.

Phadia:1983:NBE

[1244] E. G. Phadia and V. Susarla. Nonparametric Bayesian estimation of a survival curve with dependent censoring mechanism. Annals of the Institute of Statistical Mathematics, 35(1):389–400, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02480995.

Ahmad:1983:CRS

[1245] Ibrahim A. Ahmad. On L_{p}-convergence rates for statistical functions with application to L-estimates. Annals of the Institute of Statistical Mathematics, 35(1):401–406, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02480996.

Cheng:1983:MPQ

[1246] Smiley W. Cheng. On the most powerful quantile test of the scale parameter. Annals of the Institute of Statistical Mathematics, 35(1):407–414, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480997.

Shibata:1983:AME

[1247] Ritei Shibata. Asymptotic mean efficiency of a selection of regression variables. Annals of the Institute of Statistical Mathematics, 35(1):415–423, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480998.

Kishino:1983:LSE

[1248] Hirohisa Kishino. The least squares estimation of the transition probabilities of binary processes on the basis of sample paths. Annals of the Institute of Statistical Mathematics, 35(1):425–438, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02480999.

Chanda:1983:DEL

[1249] Kamal C. Chanda. Density estimation for linear processes. Annals of the Institute of Statistical Mathematics, 35(1):439–446, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481000.

Kageyama:1983:NCO

[1250] S. Kageyama and G. M. Saha. Note on the construction of optimum chemical balance weighing designs. Annals of the Institute of Statistical Mathematics, 35(1):447–452, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481001.

OShaughnessy:1983:PAS

[1251] C. D. O’Shaughnessy, Abdul Hoque, D. C. Frank, and Hee Tang Ooi. The polytopal association scheme. Annals of the Institute of Statistical Mathematics, 35(1):453–459, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481002.

Stepniak:1983:OAU

[1252] Czeslaw Stepniak. Optimal allocation of units in experimental designs with hierarchical and cross classification. Annals of the Institute of Statistical Mathematics, 35(1):461–473, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481003.

Anonymous:1983:HC

[1253] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 35(1):??, ???? 1983. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hirano:1984:PTP

[1254] Katuomi Hirano. A preliminary test procedure for the scale parameter of exponential distribution when the selection parameter is unknown. Annals of the Institute of Statistical Mathematics, 36(1):1–9, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481948.

Leonard:1984:SDA

[1255] Tom Leonard. Some data-analytic modifications to Bayes–Stein estimation. Annals of the Institute of Statistical Mathematics, 36(1):11–21, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481949.

Matsubara:1984:GDE

[1256] Nozomu Matsubara. General derivation of exact OC and ASN of SPRT when log of likelihood ratio takes only two integral multiples of a constant. Annals of the Institute of Statistical Mathematics, 36(1):23–33, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481950.

Chiang:1984:RPT

[1257] Ching-Yuan Chiang and Madan L.

Puri. Rank procedures for testing subhypotheses in linear regression. Annals of the Institute of Statistical Mathematics, 36(1):35–50, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481951.

Alam:1984:BPS

[1258] Khursheed Alam. A Bayes procedure for selecting the population with the largest p-th quantile. Annals of the Institute of Statistical Mathematics, 36(1):51–58, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481952.

Read:1984:CAA

[1259] Timothy R. C. Read. Closer asymptotic approximations for the distributions of the power divergence goodnessof-fit statistics. Annals of the Institute of Statistical Mathematics, 36 (1):59–69, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481953.

Gomes:1984:PLF

[1260] M. Ivette Gomes. Penultimate limiting forms in extreme value theory. Annals of the Institute of Statistical Mathematics, 36(1):71–85, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481954.

Basu:1984:COH

[1261] A. P. Basu and Nader Ebrahimi. Corrections to “On K-order harmonic new better than used in expectation distributions”. Annals of the Institute of Statistical Mathematics, 36(1):86, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/

BF02481955. See [1262].

Basu:1984:OHN

[1262] A. P. Basu and Nader Ebrahimi. On k-order harmonic new better than used in expectation distributions. Annals of the Institute of Statistical Mathematics, 36(1):87–100, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481956. See corrections

[1327].

Venkataraman:1984:LTL

[1263] K. N. Venkataraman and K. Suresh Chandra. Limit theorems on a linear explosive stochastic model for time series with moving average error. Annals of the Institute of Statistical Mathematics, 36(1):101–118, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481957.

Fujikoshi:1984:APM

[1264] Yasunori Fujikoshi and Yoshimichi Ochi. Asymptotic properties of the maximum likelihood estimate in the first order autoregressive process. Annals of the Institute of Statistical Mathematics, 36(1):119–128, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481958.

Bhattacharya:1984:TIA

[1265] C. G. Bhattacharya. Two inequalities with an application. Annals of the Institute of Statistical Mathematics, 36(1):129–134, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481959.

Hayashi:1984:CBS

[1266] Chikio Hayashi and Tatsuzo Suzuki. Changes in belief systems, quality of life issues and social conditions over 25 years in post-war Japan. Annals of the Institute of Statistical Mathematics, 36(1):135–161, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481960.

Bozdogan:1984:CMS

[1267] Hamparsum Bozdogan and Stanley L. Sclove. Correction to “Multi-sample cluster analysis using Akaike’s information criterion”. Annals of the Institute of Statistical Mathematics, 36(1):162, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/

BF02481961. See [1268].

Bozdogan:1984:MSC

[1268] Hamparsum Bozdogan and Stanley L. Sclove. Multi-sample cluster analysis using Akaike’s information criterion. Annals of the Institute of Statistical Mathematics, 36(1):163–180, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02481962. See

corrections [1267].

Kuboki:1984:GRC

[1269] Hisataka Kuboki. A generalization of the relative conditional expectation — further properties of Pitman’s T^{∗} and their applications to statistics. Annals of the Institute of Statistical Mathematics, 36(1):181–197, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481963.

Eguchi:1984:CSO

[1270] Shinto Eguchi. A characterization of second order efficiency in a curved exponential family. Annals of the Institute of Statistical Mathematics, 36 (1):199–206, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481964.

Inada:1984:MRE

[1271] Kôichi Inada. A minimax regret estimator of a normal mean after preliminary test. Annals of the Institute of Statistical Mathematics, 36(1):207–215, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481965.

Bar-Lev:1984:LSP

[1272] Shaul K. Bar-Lev. Large sample properties of the MLE and MCLE for the natural parameter of a truncated exponential family. Annals of the Institute of Statistical Mathematics, 36 (1):217–222, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481966.

Shiraishi:1984:RAL

[1273] Taka-Aki Shiraishi. Rank analogues of the likelihood ratio test for an ordered alternative in a two-way layout. Annals of the Institute of Statistical Mathematics, 36(1):223–237, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481967.

Crowder:1984:CML

[1274] Martin Crowder. On constrained maximum likelihood estimation with noni.i.d. observations. Annals of the Institute of Statistical Mathematics, 36 (1):239–249, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481968.

Kimura:1984:RST

[1275] Miyoshi Kimura. Robust slippage tests. Annals of the Institute of Statistical Mathematics, 36(1):251–270, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481969.

Mathal:1984:EWI

[1276] A. M. Mathal. Extensions of Wilks’ integral equations and distributions of test statistics. Annals of the Institute of Statistical Mathematics, 36(1):271–288, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481970.

Nagaraja:1984:ALP

[1277] H. N. Nagaraja. Asymptotic linear prediction of extreme order statistics. Annals of the Institute of Statistical Mathematics, 36(1):289–299, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481971.

Berman:1984:LDS

[1278] Simeon M. Berman. Limiting distribution of sums of nonnegative stationary random variables. Annals of the Institute of Statistical Mathematics, 36(1):301–321, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481972.

Agrawal:1984:RSS

[1279] M. C. Agrawal. A robustness study in sampling on successive occasions. Annals of the Institute of Statistical Mathematics, 36(1):323–335, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481973.

Nagasaka:1984:BL

[1280] Kenji Nagasaka. On Benford’s law.

Annals of the Institute of Statistical Mathematics, 36(1):337–352, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481974.

Itoh:1984:RCM

[1281] Yoshiaki Itoh. Random collision model for random genetic drift and stochastic difference equation. Annals of the Institute of Statistical Mathematics, 36 (1):353–362, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481975.

Yanagimoto:1984:MFP

[1282] Takemi Yanagimoto and Toshinari Kamakura. The maximum full and partial likelihood estimators in the proportional hazard model. Annals of the Institute of Statistical Mathematics, 36(1):363–373, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481976.

Nakamura:1984:ETM

[1283] Tadashi Nakamura. Existence theorems of a maximum likelihood estimate from a generalized censored data sample. Annals of the Institute of Statistical Mathematics, 36(1):375–393, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481977.

Srivastava:1984:ERC

[1284] V. K. Srivastava, Baldev Raj, and Kuldeep Kumar. Estimation of a random coefficient model under linear stochastic constraints. Annals of the Institute of Statistical Mathematics, 36(1):395–401, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481978.

Robinson:1984:KEI

[1285] P. M. Robinson. Kernel estimation and interpolation for time series containing missing observations. Annals of the Institute of Statistical Mathematics, 36(1):403–417, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481979.

Yamamoto:1984:ABL

[1286] Taku Yamamoto and Naoto Kunitomo.

Asymptotic bias of the least squares estimator for multivariate autoregressive models. Annals of the Institute of Statistical Mathematics, 36(1):419–430, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481980.

Aki:1984:DDO

[1287] Sigeo Aki, Hisataka Kuboki, and Katuomi Hirano. On discrete distributions of order k. Annals of the Institute of Statistical Mathematics, 36 (1):431–440, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481981.

Loh:1984:SUS

[1288] Wei-Yin Loh. Strong unimodality and scale mixtures. Annals of the Institute of Statistical Mathematics, 36 (1):441–449, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481982.

Lin:1984:NED

[1289] Gwo Dong Lin. A note on equal distributions. Annals of the Institute of Statistical Mathematics, 36(1):451–453, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481983.

Georgiev:1984:SCN

[1290] Alexander A. Georgiev. Speed of convergence in nonparametric kernel estimation of a regression function and its derivatives. Annals of the Institute of Statistical Mathematics, 36 (1):455–462, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481984. See corrections

[1755].

Shiraishi:1984:SAR

[1291] Taka-Aki Shiraishi. Semi-aligned rank tests. Annals of the Institute of Statistical Mathematics, 36(1):463–473, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481985.

Azuma:1984:MSC

[1292] Shinji Azuma, Kenji Hayashi, and Akio Kudô. Moments of a statistic caused by random combinations or random matings. Annals of the Institute of Statistical Mathematics, 36(1):475–479, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481986.

Huda:1984:VFG

[1293] S. Huda. Variance functions for mgrouped cylindrically rotatable designs of type 3. Annals of the Institute of Statistical Mathematics, 36(1):481–483, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481987.

Menzefricke:1984:DTA

[1294] Ulrich Menzefricke. A decisiontheoretic approach to some screening problems. Annals of the Institute of Statistical Mathematics, 36(1):485–497, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481988.

Barankin:1984:BRV

[1295] Edward W. Barankin and Koiti Takahasi. Betweeness for real vectors and lines, III alternative characterizations of betweennesses. Annals of the Institute of Statistical Mathematics, 36 (1):499–520, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481989.

Ishiguro:1984:BAP

[1296] Makio Ishiguro and Yosiyuki Sakamoto.

A Bayesian approach to the probability density estimation. Annals of the Institute of Statistical Mathematics, 36(1):523–538, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481990.

Anonymous:1984:HC

[1297] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 36(1):??, ???? 1984. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matsunawa:1985:EAD

[1298] T. Matsunawa. The exact and approximate distributions of linear combinations of selected order statistics from a uniform distribution. Annals of the Institute of Statistical Mathematics, 37(1):1–16, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481076.

Akahira:1985:ECP

[1299] Masafumi Akahira and Kei Takeuchi. Estimation of a common parameter for pooled samples from the uniform distributions. Annals of the Institute of Statistical Mathematics, 37(1):17–26, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481077.

Hall:1985:LTM

[1300] Peter Hall and A. H. Welsh. Limit theorems for the median deviation. Annals of the Institute of Statistical Mathematics, 37(1):27–36, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481078.

Yanagimoto:1985:SLA

[1301] Takemi Yanagimoto and Eiji Yamamoto. Simple linear approximations to the likelihood equation for combining evidence in multiple 2 × 2 tables: A critique of conventional procedures. Annals of the Institute of Statistical Mathematics, 37(1):37–49, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481079.

Puri:1985:RCN

[1302] Madan L. Puri and Munsup Seoh. On the rate of convergence to normality for generalized linear rank statistics. Annals of the Institute of Statistical Mathematics, 37(1):51–69, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481080.

Singh:1985:EBE

[1303] R. S. Singh. Empirical Bayes estimation in a multiple linear regression model. Annals of the Institute of Statistical Mathematics, 37(1):71–86, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481081.

Konishi:1985:NVS

[1304] Sadanori Konishi. Normalizing and variance stabilizing transformations for intraclass correlations. Annals of the Institute of Statistical Mathematics, 37(1):87–94, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481082.

Hayakawa:1985:AED

[1305] Takesi Hayakawa and Madan L. Puri. Asymptotic expansions of the distributions of some test statistics. Annals of the Institute of Statistical Mathematics, 37(1):95–108, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481083.

Esty:1985:ADT

[1306] W. Esty, R. Gillette, M. Hamilton, and D. Taylor. Asymptotic distribution theory of statistical functionals: The compact derivative approach for robust estimators. Annals of the Institute of Statistical Mathematics, 37(1):109–129, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481084.

Taylor:1985:ADT

[1307] Donald C. Taylor. Asymptotic distribution theory for general statistical functionals. Annals of the Institute of Statistical Mathematics, 37(1):131–138, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481085.

Kageyama:1985:CPD

[1308] Sanpei Kageyama. Connectedness of PBIB designs having asymmetrical association schemes. Annals of the Institute of Statistical Mathematics, 37 (1):139–143, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481086.

Banerjee:1985:SCP

[1309] Shakti Banerjee, Bhagwandas, and S. Kageyama. Some constructions of PBIB designs. Annals of the Institute of Statistical Mathematics, 37 (1):145–150, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481087.

Lingappaiah:1985:SSM

[1310] G. S. Lingappaiah. A study of shifting models in life tests via Bayesian approach using semi-or-used priors (SOUPS). Annals of the Institute of Statistical Mathematics, 37(1):151–163, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481088.

Joe:1985:CLD

[1311] Harry Joe. Characterizations of life distributions from percentile residual lifetimes. Annals of the Institute of Statistical Mathematics, 37(1):165–172, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481089.

Tanabe:1985:CGM

[1312] Kunio Tanabe. The conjugate gradient method for computing all the extremal stationary probability vectors of a stochastic matrix. Annals of the Institute of Statistical Mathematics, 37 (1):173–187, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481090.

Tanabe:1985:GAC

[1313] Kunio Tanabe. Global analysis of continuous analogues of the Levenberg– Marquardt and Newton–Raphson methods for solving nonlinear equations. Annals of the Institute of Statistical Mathematics, 37(1):189–203, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481091.

Aki:1985:DDO

[1314] Sigeo Aki. Discrete distributions of order k on a binary sequence. Annals of the Institute of Statistical Mathematics, 37(1):205–224, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481092.

Ruschendorf:1985:CMD

[1315] Ludger Rüschendorf. Construction of multivariate distributions with given marginals. Annals of the Institute of Statistical Mathematics, 37(1):225–233, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481093.

Henna:1985:ENC

[1316] Jõgi Henna. On estimating of the number of constituents of a finite mixture of continuous distributions. Annals of the Institute of Statistical Mathematics, 37(1):235–240, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481094.

Kyriakoussis:1985:AMV

[1317] A. Kyriakoussis. Asymptotically minimum variance unbiased estimation for a class of power series distributions. Annals of the Institute of Statistical Mathematics, 37(1):241–250, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481095.

Inagaki:1985:PTM

[1318] Nobuo Inagaki and Pranab Kumar Sen.

On progressively truncated maximum likelihood estimators. Annals of the Institute of Statistical Mathematics, 37 (1):251–269, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481096.

Cox:1985:PMN

[1319] Dennis D. Cox. A penalty method for nonparametric estimation of the logarithmic derivative of a density function. Annals of the Institute of Statistical Mathematics, 37(1):271–288, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481097.

Isogai:1985:SEH

[1320] Takafumi Isogai. Some extension of Haldane’s multivariate median and its application. Annals of the Institute of Statistical Mathematics, 37(1):289–301, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481098.

Chen:1985:SSL

[1321] Pinyuen Chen. Subset selection for the least probable multinomial cell. Annals of the Institute of Statistical Mathematics, 37(1):303–314, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481099.

Laippala:1985:EBR

[1322] Pekka Laippala. The empirical Bayes rules with floating optimal sample size for exponential conditional distributions. Annals of the Institute of Statistical Mathematics, 37(1):315–327, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481100.

Raktoe:1985:LSA

[1323] B. L. Raktoe and W. T. Federer. Lattice square approach to construction of mutually orthogonal F -squares. Annals of the Institute of Statistical Mathematics, 37(1):329–336, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481101.

Saha:1985:CBA

[1324] G. M. Saha and B. K. Samanta. A construction of balanced arrays of strength t and some related incomplete block designs. Annals of the Institute of Statistical Mathematics, 37(1):337–345, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481102.

Basu:1985:TWS

[1325] A. P. Basu and Nader Ebrahimi. Testing whether survival function is harmonic new better than used in expectation. Annals of the Institute of Statistical Mathematics, 37(1):347–359, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481103.

Klefsjo:1985:SCP

[1326] Bengt Klefsjö. Some comments on a paper on k-HNBUE life distributions. Annals of the Institute of Statistical Mathematics, 37(1):361–364, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481104.

Basu:1985:COH

[1327] A. P. Basu and Nader Ebrahimi. Corrections to “On k-order harmonic new better than used in expectation distributions”. Annals of the Institute of Statistical Mathematics, 37(1):365–366, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/

BF02481105. See [1262].

Makino:1985:CII

[1328] Toji Makino. Corrections to “On the independence of interdeparture intervals from single server queueing systems”. Annals of the Institute of Statistical Mathematics, 37(1):367, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1007/
                                        BF02481106. See [947].

Sakamoto:1985:BBR

[1329] Yosiyuki Sakamoto and Makio Ishiguro. Bayesian binary regression involving two explanatory variables. Annals of the Institute of Statistical Mathematics, 37(1):369–387, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481107. See corrections

[1420].

Yajima:1985:EDD

[1330] Yoshihiro Yajima. Estimation of the degree of differencing of an ARIMA process. Annals of the Institute of Statistical Mathematics, 37(1):389–408, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481108.

Koteeswaran:1985:SCT

[1331] P. Koteeswaran, K. Nanthi, and K. Suresh Chandra. Some convergence theorems on a supercritical Galton– Watson process. Annals of the Institute of Statistical Mathematics, 37 (1):409–414, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481109.

Noda:1985:OCS

[1332] Kazuo Noda. Optimal construction of a selection of a subpopulation. Annals of the Institute of Statistical Mathematics, 37(1):415–435, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481110.

Nagata:1985:MPT

[1333] Yasushi Nagata and Taichi Inaba. Minimaxity of a preliminary test estimator for the mean of normal distribution. Annals of the Institute of Statistical Mathematics, 37(1):437–442, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481111.

Greblicki:1985:FHS

[1334] Wlodzimierz Greblicki and Miroslaw Pawlak. Fourier and Hermite series estimates of regression functions. Annals of the Institute of Statistical Mathematics, 37(1):443–454, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481112.

Sinha:1985:USE

[1335] Bikas Kumar Sinha and Arup Bose.

Unbiased sequential estimation of 1/ p: Settlement of a conjecture. Annals of the Institute of Statistical Mathematics, 37(1):455–460, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481113.

vanEeden:1985:MIS

[1336] Constance van Eeden. Mean integrated squared error of kernel estimators when the density and its derivative are not necessarily continuous. Annals of the Institute of Statistical Mathematics, 37(1):461–472, ????

1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481114.

Moothathu:1985:DML

[1337] T. S. K. Moothathu. Distributions of maximum likelihood estimators of Lorenz curve and Gini index of exponential distribution. Annals of the Institute of Statistical Mathematics, 37 (1):473–479, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481115.

deWaal:1985:RSM

[1338] D. J. de Waal, P. C. N. Groenewald, J. M. van Zyl, and J. V. Zidek. A randomized solution for multi-Bayes estimates of the multinormal mean. Annals of the Institute of Statistical Mathematics, 37(1):481–486, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481116.

Ki:1985:ICS

[1339] Fanny Ki and Kam-Wah Tsui. Improved confidence set estimators of a multivariate normal mean and generalizations. Annals of the Institute of Statistical Mathematics, 37(1):487–498, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481117.

Chiou:1985:ECM

[1340] Wen-Jau Chiou and Arthur Cohen. On estimating a common multivariate normal mean vector. Annals of the Institute of Statistical Mathematics, 37(1):499–506, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481118.

Kaminsky:1985:MLP

[1341] Kenneth S. Kaminsky and Lennart S.

Rhodin. Maximum likelihood prediction. Annals of the Institute of Statistical Mathematics, 37(1):507–517, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481119.

Shiraishi:1985:LPT

[1342] Taka-Aki Shiraishi. Local powers of two-sample and multi-sample rank tests for Lehmann’s contaminated alternative. Annals of the Institute of Statistical Mathematics, 37(1):519–527, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481120.

Reiss:1985:ADO

[1343] R.-D. Reiss. Approximations to the distributions of ordered distance random variables. Annals of the Institute of Statistical Mathematics, 37(1):529–533, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481121.

Choi:1985:CLC

[1344] B. S. Choi. A conditional limit construction of the normal probability density. Annals of the Institute of Statistical Mathematics, 37(1):535–539, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481122.

Moschopoulos:1985:DSI

[1345] P. G. Moschopoulos. The distribution of the sum of independent gamma random variables. Annals of the Institute of Statistical Mathematics, 37 (1):541–544, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481123.

Blough:1985:MLP

[1346] David K. Blough. Measures of location in the plane. Annals of the Institute of Statistical Mathematics, 37(1):545–555, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481124.

Jacroux:1985:SOD

[1347] Mike Jacroux. Some E and MV - optimal designs for the two-way elimination of heterogeneity. Annals of the Institute of Statistical Mathematics, 37(1):557–566, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481125.

Marx:1985:CQF

[1348] D. G. Marx. Corrections to “Quadratic forms of a matric-t variate”. Annals of the Institute of Statistical Mathematics, 37(1):567, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02481126. See

[1239].

Anonymous:1985:HC

[1349] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 37(1):??, ???? 1985. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aki:1986:STS

[1350] Sigeo Aki. Some test statistics based on the martingale term of the empirical distribution function. Annals of the Institute of Statistical Mathematics, 38(1):1–21, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482496.

Evans:1986:MCR

[1351] Michael Evans. Moments of coverage of a random ellipsoid. Annals of the Institute of Statistical Mathematics, 38(1):23–33, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482497.

Akahira:1986:BBV

[1352] Masafumi Akahira, Madan L. Puri, and Kei Takeuchi. Bhattacharyya bound of variances of unbiased estimators in nonregular cases. Annals of the Institute of Statistical Mathematics, 38(1):35–44, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482498.

Tsui:1986:MES

[1353] Kam-Wah Tsui. Multiparameter estimation for some multivariate discrete distributions with possibly dependent components. Annals of the Institute of Statistical Mathematics, 38(1):45–56, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482499.

Kano:1986:CCL

[1354] Yutaka Kano. Consistency conditions on the least squares estimator in single common factor analysis model. Annals of the Institute of Statistical Mathematics, 38(1):57–68, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482500.

Isogai:1986:ACF

[1355] Eiichi Isogai. Asymptotic consistency of fixed-width sequential confidence intervals for a multiple regression function. Annals of the Institute of Statistical Mathematics, 38(1):69–83, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482501.

Akai:1986:SEL

[1356] Toyoaki Akai. Simultaneous estimation of location parameters of the distribution with finite support. Annals of the Institute of Statistical Mathematics, 38(1):85–99, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482502.

Dey:1986:TME

[1357] Dipak K. Dey and C. Srinivasan.

Trimmed minimax estimator of a covariance matrix. Annals of the Institute of Statistical Mathematics, 38 (1):101–108, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482503.

Chikuse:1986:SPI

[1358] Yasuko Chikuse and A. W. Davis.

Some properties of invariant polynomials with matrix arguments and their applications in econometrics. Annals of the Institute of Statistical Mathematics, 38(1):109–122, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482504.

Gates:1986:CES

[1359] David J. Gates and Mark Westcott.

Clustering estimates for spatial point distributions with unstable potentials. Annals of the Institute of Statistical Mathematics, 38(1):123–135, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482505.

Huang:1986:CLD

[1360] Wei-Min Huang. A characterization of limiting distributions of estimators in an autoregressive process. Annals of the Institute of Statistical Mathematics, 38(1):137–144, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482506.

Shirakura:1986:BPF

[1361] Teruhiro Shirakura. Block plan for a fractional 2^{m} factorial design derived from a 2^{r} factorial design. Annals of the Institute of Statistical Mathematics, 38(1):145–159, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482507.

Constantine:1986:OBD

[1362] Gregory M. Constantine. On the optimality of block designs. Annals of the Institute of Statistical Mathematics, 38(1):161–174, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482508.

Barankin:1986:TSI

[1363] Edward W. Barankin. Time, space and incidence in general, with a consequent proof of the law of homeostasis. Annals of the Institute of Statistical Mathematics, 38(1):175–194, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482509.

Shimizu:1986:IDM

[1364] Ryoichi Shimizu. Inequalities for a distribution with monotone hazard rate. Annals of the Institute of Statistical Mathematics, 38(1):195–204, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482510.

Matsunawa:1986:MIC

[1365] T. Matsunawa. Modified information criteria for a uniform approximate equivalence of probability distributions. Annals of the Institute of Statistical Mathematics, 38(1):205–222, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482511.

Panaretos:1986:GBM

[1366] John Panaretos and Evdokia Xekalaki.

On generalized binomial and multinomial distributions and their relation to generalized Poisson distributions. Annals of the Institute of Statistical Mathematics, 38(1):223–231, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482512.

Bowman:1986:RDU

[1367] K. O. Bowman and L. R. Shenton.

A reinforcement-depletion urn model: A contiguity case. Annals of the Institute of Statistical Mathematics, 38 (1):233–243, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482513.

Falk:1986:RUC

[1368] M. Falk. Rates of uniform convergence of extreme order statistics. Annals of the Institute of Statistical Mathematics, 38(1):245–262, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482514.

Watson:1986:SET

[1369] G. S. Watson. Some estimation theory on the sphere. Annals of the Institute of Statistical Mathematics, 38(1):263–275, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482515.

Bar-lev:1986:LRT

[1370] Shaul K. Bar-lev and Benzion Boukai. Likelihood ratio tests for comparing k populations — the two-parameter nonregular models. Annals of the Institute of Statistical Mathematics, 38 (1):277–283, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02482516.

Provost:1986:STS

[1371] Serge B. Provost. Some test statistics for the structural coefficients of the multivariate linear functional relationship model. Annals of the Institute of Statistical Mathematics, 38(1):285–296, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482517.

Kwoun:1986:AMT

[1372] Gea Hwa Kwoun and Yoshihiro Yajima. On an autoregressive model with time-dependent coefficients. Annals of the Institute of Statistical Mathematics, 38(1):297–309, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02482518.

Chandra:1986:ISS

[1373] K. Suresh Chandra and P. Koteeswaran. Inference on superimposed subcritical Galton–Watson processes with immigration. Annals of the Institute of Statistical Mathematics, 38 (1):311–318, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482519.

Nishii:1986:CSR

[1374] Ryuei Nishii. Criteria for selection of response variables and the asymptotic properties in a multivariate calibration. Annals of the Institute of Statistical Mathematics, 38(1):319–329, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482520.

Mukerjee:1986:FOP

[1375] Rahul Mukerjee and Haruo Yanai. Factorial orthogonality in the presence of covariates. Annals of the Institute of Statistical Mathematics, 38(1):331–341, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482521.

Kuwada:1986:OPB

[1376] Masahide Kuwada. Optimal partially balanced fractional 2^{m_{1}+m_{2}} factorial designs of resolution IV. Annals of the Institute of Statistical Mathematics, 38(1):343–351, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482522.

Nakamura:1986:BCM

[1377] Takashi Nakamura. Bayesian cohort models for general cohort table analyses. Annals of the Institute of Statistical Mathematics, 38(1):353–370, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482523.

Niki:1986:ETH

[1378] Naoto Niki and Sadanori Konishi. Effects of transformations in higher order asymptotic expansions. Annals of the Institute of Statistical Mathematics, 38(1):371–383, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02482524.

Eguchi:1986:PME

[1379] Shinto Eguchi. A projection method of estimation for a subfamily of exponential families. Annals of the Institute of Statistical Mathematics, 38(1):385–398, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482525.

Alam:1986:CRM

[1380] Khursheed Alam and Amitava Mitra. Component risk in multiparameter estimation. Annals of the Institute of Statistical Mathematics, 38(1):399–410, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482526.

Sen:1986:SLE

[1381] Pranab Kumar Sen. On stable laws for estimating functions and derived estimators. Annals of the Institute of Statistical Mathematics, 38(1):411–417, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482527.

Ghosh:1986:IPT

[1382] Malay Ghosh and Dipak K. Dey. On the inadmissibility of preliminary-test estimators when the loss involves a complexity cost. Annals of the Institute of Statistical Mathematics, 38 (1):419–427, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482528.

Barabas:1986:BCB

[1383] Béla Barabás, Miklós Csörgő, Lajos Horváth, and Brian S. Yandell. Bootstrapped confidence bands for percentile lifetime. Annals of the Institute of Statistical Mathematics, 38 (1):429–438, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482529.

Babu:1986:NBV

[1384] Gutti Jogesh Babu. A note on bootstrapping the variance of sample quantile. Annals of the Institute of Statistical Mathematics, 38(1):439–443, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482530.

Bofinger:1986:CDS

[1385] Eve Bofinger. ψ-correct decision for selection and elimination. Annals of the Institute of Statistical Mathematics, 38(1):445–450, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02482531.

Dumitrescu:1986:APM

[1386] Monica E. Bad Dumitrescu. The application of the principle of minimum cross-entropy to the characterization of the exponential-type probability distributions. Annals of the Institute of Statistical Mathematics, 38(1):451–457, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482532.

Shibata:1986:SNR

[1387] Ritei Shibata. Selection of the number of regression variables; a minimax choice of generalized FPE. Annals of the Institute of Statistical Mathematics, 38(1):459–474, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482533.

Rao:1986:WCL

[1388] B. L. S. Prakasa Rao. The weak convergence of least squares random fields and its application. Annals of the Institute of Statistical Mathematics, 38 (1):475–483, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02482534.

Koziol:1986:REG

[1389] James A. Koziol. Relative efficiencies of goodness of fit procedures for assessing univariate normality. Annals of the Institute of Statistical Mathematics, 38(1):485–493, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482535.

Chou:1986:LMT

[1390] Rouh-Jane Chou and Wen da Lo.

On the local minimaxity of a test of independence in incomplete samples. Annals of the Institute of Statistical Mathematics, 38(1):495–502, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482536.

Sutrick:1986:APC

[1391] Kenneth H. Sutrick. Asymptotic power comparison of the chi-square and likelihood ratio tests. Annals of the Institute of Statistical Mathematics, 38 (1):503–511, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482537.

Shiraishi:1986:APR

[1392] Taka-Aki Shiraishi. The asymptotic power of rank tests under scalealternatives including contaminated distributions. Annals of the Institute of Statistical Mathematics, 38(1):513–522, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482538.

Hyakutake:1986:CMC

[1393] Hiroto Hyakutake. A construction method of certain matrices required in the multivariate heteroscedastic method. Annals of the Institute of Statistical Mathematics, 38(1):523–528, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482539.

Balakrishnan:1986:EMB

[1394] N. Balakrishnan, S. Kocherlakota, and K. Kocherlakota. On the errors of misclassification based on dichotomous and normal variables. Annals of the Institute of Statistical Mathematics, 38(1):529–538, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482540.

Robinson:1986:CFS

[1395] P. M. Robinson. On the consistency and finite-sample properties of nonparametric kernel time series regression, autoregression and density estimators. Annals of the Institute of Statistical Mathematics, 38(1):539–549, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482541.

David:1986:IOS

[1396] H. A. David. Inequalities for ordered sums. Annals of the Institute of Statistical Mathematics, 38(1):551–555, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02482542.

Charalambides:1986:DDO

[1397] Ch. A. Charalambides. On discrete distributions of order k. Annals of the Institute of Statistical Mathematics, 38(1):557–568, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02482543.

Jacroux:1986:URL

[1398] Mike Jacroux. On the usage of refined linear models for determining N-way classification designs which are optimal for comparing test treatments with a standard treatment. Annals of the Institute of Statistical Mathematics, 38(1):569–581, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482544.

Itoh:1986:GCR

[1399] Yoshiaki Itoh. Golay code and random packing. Annals of the Institute of Statistical Mathematics, 38(1):583–588, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02482545.

Anonymous:1986:HC

[1400] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 38(1):??, ???? 1986. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kuboki:1987:AMC

[1401] Hisataka Kuboki. Analysis of marginal and conditional density functions for separate inference. Annals of the Institute of Statistical Mathematics, 39(1):1–23, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491445.

Akahira:1987:SOA

[1402] Masafumi Akahira. Second order asymptotic comparison of estimators of a common parameter in the double exponential case. Annals of the Institute of Statistical Mathematics, 39(1):25–36, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491446.

Dorea:1987:EEV

[1403] Chang C. Y. Dorea. Estimation of the extreme value and the extreme points. Annals of the Institute of Statistical Mathematics, 39(1):37–48, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491447.

Inaba:1987:MNP

[1404] Taichi Inaba and Yasushi Nagata. Minimaxity and nonminimaxity of a preliminary test estimator for the multivariate normal mean. Annals of the Institute of Statistical Mathematics, 39(1):49–54, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491448.

Chaturvedi:1987:SPE

[1405] Ajit Chaturvedi. Sequential point estimation of regression parameters in a linear model. Annals of the Institute of Statistical Mathematics, 39(1):55–67, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491449.

Sibuya:1987:RSB

[1406] Masaaki Sibuya and Yoshiaki Itoh.

Random sequential bisection and its associated binary tree. Annals of the Institute of Statistical Mathematics, 39(1):69–84, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491450.

Takeuchi:1987:SRVa

[1407] Kei Takeuchi and Akimichi Takemura.

On sum of 0–1 random variables I. Univariate case. Annals of the Institute of Statistical Mathematics, 39(1):85–102, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491451.

Cowan:1987:BED

[1408] Richard Cowan. A bivariate exponential distribution arising in random geometry. Annals of the Institute of Statistical Mathematics, 39(1):103–111, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491452.

Kimeldorf:1987:PDO

[1409] George Kimeldorf and Allan R. Sampson. Positive dependence orderings. Annals of the Institute of Statistical Mathematics, 39(1):113–128, ????

1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491453.

Lin:1987:RBT

[1410] Gwo Dong Lin. Relationships between two extensions of Farlie–Gumbel– Morgenstern distribution. Annals of the Institute of Statistical Mathematics, 39(1):129–140, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02491454.

Davis:1987:SDT

[1411] A. W. Davis and T. Hayakawa. Some distribution theory relating to confidence regions in multivariate calibration. Annals of the Institute of Statistical Mathematics, 39(1):141–152, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491455.

Fujikoshi:1987:EBA

[1412] Yasunori Fujikoshi. Error bounds for asymptotic expansions of the distribution of the MLE in a GMANOVA model. Annals of the Institute of Statistical Mathematics, 39(1):153–161, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491456.

Sarkar:1987:TSU

[1413] Shakuntala Sarkar and P. R. Krishnaiah. Tests for sphericity under correlated multivariate regression equations model. Annals of the Institute of Statistical Mathematics, 39(1):163–175, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491457.

Dudewicz:1987:HMM

[1414] Edward J. Dudewicz and Vidya S.

Taneja. The heteroscedastic method: Multivariate implementation. Annals of the Institute of Statistical Mathematics, 39(1):177–190, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491458.

Uesaka:1987:LSL

[1415] Hiroyuki Uesaka and Chooichiro Asano. Latent scale linear models for multivariate ordinal responses and analysis by the method of weighted least squares. Annals of the Institute of Statistical Mathematics, 39(1):191–210, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491459.

Dumitrescu:1987:NPE

[1416] Monica E. Bad Dumitrescu. On the normality a posteriori for exponential distributions, using the Bayesian estimation. Annals of the Institute of Statistical Mathematics, 39(1):211–218, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491460.

Draper:1987:BAH

[1417] Norman Draper and Irwin Guttman.

Bayesian analysis of hybrid life tests with exponential failure times. Annals of the Institute of Statistical Mathematics, 39(1):219–225, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491461.

Tamura:1987:ANP

[1418] Yoshiyasu-Hamada Tamura. An approach to the nonstationary process analysis. Annals of the Institute of Statistical Mathematics, 39(1):227–241, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491462.

Vere-Jones:1987:CSE

[1419] D. Vere-Jones and T. Ozaki. Corrections to “Some examples of statistical estimation applied to earthquake data”. Annals of the Institute of Statistical Mathematics, 39(1):243, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/

BF02491463. See [1170].

Sakamoto:1987:CBB

[1420] Yosiyuki Sakamoto and Makio Ishiguro. Corrections to “Bayesian binary regression involving two explanatory variables”. Annals of the Institute of Statistical Mathematics, 39(1):245, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/

BF02491464. See [1329].

Yanagimoto:1987:NOR

[1421] Takemi Yanagimoto. A notion of an obstructive residual likelihood. Annals of the Institute of Statistical Mathematics, 39(1):247–261, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491465.

Mussmann:1987:CML

[1422] Dieter Mussmann. On a characterization of monotone likelihood ratio experiments. Annals of the Institute of Statistical Mathematics, 39(1):263–274, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491466.

Kishino:1987:VSS

[1423] Hirohisa Kishino. Variance of sightings in the survey of patchily distributed objects. Annals of the Institute of Statistical Mathematics, 39(1):275–287, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491467.

Tong:1987:IEC

[1424] Y. L. Tong. Interval estimation of the critical value in a general linear model. Annals of the Institute of Statistical Mathematics, 39(1):289–297, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491468.

Hayakawa:1987:NVS

[1425] Takesi Hayakawa. Normalizing and variance stabilizing transformations of multivariate statistics under an elliptical population. Annals of the Institute of Statistical Mathematics, 39 (1):299–306, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491469.

Takeuchi:1987:SRVb

[1426] Kei Takeuchi and Akimichi Takemura. On sum of 0–1 random variables II. Multivariate case. Annals of the Institute of Statistical Mathematics, 39 (1):307–324, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491470.

Chen:1987:RPS

[1427] Pinyuen Chen. The k-in-a-row procedure in selection theory. Annals of the Institute of Statistical Mathematics, 39(1):325–330, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491471.

Cheng:1987:NID

[1428] K. F. Cheng. Nonparametric inference on the difference of location parameters of correlated variables from fragmentary samples. Annals of the Institute of Statistical Mathematics, 39(1):331–347, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491472.

Govindarajulu:1987:AES

[1429] Z. Govindarajulu and Bo. H. Lindqvist. Asymptotic efficiency of the Spearman estimator and characterizations of distributions. Annals of the Institute of Statistical Mathematics, 39(1):349–361, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491473.

Maesono:1987:CWS

[1430] Yoshihiko Maesono. Competitors of the Wilcoxon signed rank test. Annals of the Institute of Statistical Mathematics, 39(1):363–375, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491474.

Yao:1987:NTC

[1431] Yi-Ching Yao. A note on testing for constant hazard against a change-point alternative. Annals of the Institute of Statistical Mathematics, 39(1):377–383, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491475.

Lee:1987:SSC

[1432] K. Y. Lee and Mike Jacroux. Some sufficient conditions for the E- and MV -optimality of block designs having blocks of unequal size. Annals of the Institute of Statistical Mathematics, 39(1):385–397, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491476.

Cacoullos:1987:CPP

[1433] Theophilos Cacoullos. Characterizing priors by posterior expectations in multiparameter exponential families. Annals of the Institute of Statistical Mathematics, 39(1):399–405, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491477.

Sato:1987:MMU

[1434] Ken iti Sato. Modes and moments of unimodal distributions. Annals of the Institute of Statistical Mathematics, 39(1):407–415, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491478.

Masry:1987:LNF

[1435] Elias Masry and Bernard Picinbono. Linear/ nonlinear forms and the normal law: Characterization by high order correlations. Annals of the Institute of Statistical Mathematics, 39 (1):417–428, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491479.

Muliere:1987:CMO

[1436] Pietro Muliere and Marco Scarsini. Characterization of a Marshall–Olkin type class of distributions. Annals of the Institute of Statistical Mathematics, 39(1):429–441, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491480.

Sato:1987:PTI

[1437] Manabu Sato. Pragmatic treatment of improper solutions in factor analysis. Annals of the Institute of Statistical Mathematics, 39(1):443–455, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491481.

Aki:1987:NTS

[1438] Sigeo Aki. On nonparametric tests for symmetry. Annals of the Institute of Statistical Mathematics, 39(1):457–472, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491482.

Nabeya:1987:ANT

[1439] Seiji Nabeya. On Aki’s nonparametric test for symmetry. Annals of the Institute of Statistical Mathematics, 39 (1):473–482, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491483.

Bergman:1987:NST

[1440] Bo Bergman and Bengt Klefsjö. A note on some test statistics against HNBUE. Annals of the Institute of Statistical Mathematics, 39(1):483–488, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491484.

Nomakuchi:1987:NTT

[1441] Kentaro Nomakuchi and Toshio Sakata. A note on testing twodimensional normal mean. Annals of the Institute of Statistical Mathematics, 39(1):489–495, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02491485.

Sarma:1987:APR

[1442] Y. Rama Krishna Sarma. Asymptotic properties of Rao’s test for testing hypotheses in discrete parameter stochastic processes. Annals of the Institute of Statistical Mathematics, 39(1):497–512, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491486.

Uesaka:1987:AOC

[1443] Hiroyuki Uesaka and Chooichiro Asano. Analysis of ordered categorical data from repeated measurements assuming a quantitative latent variable. Annals of the Institute of Statistical Mathematics, 39(1):513–531, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491487.

Hardle:1987:ESR

[1444] Wolfgang Härdle. An effective selection of regression variables when the error distribution is incorrectly specified. Annals of the Institute of Statistical Mathematics, 39(1):533–548, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491488.

Singh:1987:MNK

[1445] Radhey S. Singh and Manzoor Ahmad. Modified nonparametric kernel estimates of a regression function and their consistencies with rates.

Annals of the Institute of Statistical Mathematics, 39(1):549–562, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491489.

Stepniak:1987:CCL

[1446] C. Stepniak. A complete class for linear estimation in a general linear model. Annals of the Institute of Statistical Mathematics, 39(1):563–573, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491490.

Kunitomo:1987:TOO

[1447] Naoto Kunitomo. A third order optimum property of the ML estimator in a linear functional relationship model and simultaneous equation system in econometrics. Annals of the Institute of Statistical Mathematics, 39(1):575–591, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491491.

Akahira:1987:LBV

[1448] Masafumi Akahira and Kei Takeuchi.

The lower bound for the variance of unbiased estimators for one-directional family of distributions. Annals of the Institute of Statistical Mathematics, 39 (1):593–610, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02491492.

Shimizu:1987:EBA

[1449] Ryoichi Shimizu. Error bounds for asymptotic expansion of the scale mixtures of the normal distribution. Annals of the Institute of Statistical Mathematics, 39(1):611–622, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491493.

Bhattacharya:1987:BNA

[1450] Samir K. Bhattacharya. Bayesian normal analysis with an inverse Gaussian prior. Annals of the Institute of Statistical Mathematics, 39(1):623–626, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491494.

George:1987:LM

[1451] E. Olusegun George and Cecil C. Rousseau. On the logistic midrange. Annals of the Institute of Statistical Mathematics, 39(1):627–635, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491495.

Takahashi:1987:SPM

[1452] Rinya Takahashi. Some properties of multivariate extreme value distributions and multivariate tail equivalence. Annals of the Institute of Statistical Mathematics, 39(1):637–647, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02491496.

Kuwada:1987:RBF

[1453] Masahide Kuwada. On the robustness of balanced fractional 2^{m} factorial designs of resolution 2l+1 in the presence of outliers. Annals of the Institute of Statistical Mathematics, 39(1):649–659, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491497.

Kageyama:1987:SCL

[1454] Sanpei Kageyama. Some characterization of locally resistant BIB designs of degree one. Annals of the Institute of Statistical Mathematics, 39(1):661–669, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491498.

Banerjee:1987:SCT

[1455] Snigdha Banerjee, Sanpei Kageyama, and Bhagwandas. Some constructions of two-associate class PBIB designs. Annals of the Institute of Statistical Mathematics, 39(1):671–679, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02491499.

Hayakawa:1987:CLR

[1456] Takesi Hayakawa. Correction to “The likelihood ratio criterion and the asymptotic expansion of its distribution”. Annals of the Institute of Statistical Mathematics, 39(1):681, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02491500. See [951].

Anonymous:1987:HC

[1457] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 39(1):??, ???? 1987. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nakamura:1988:SIO

[1458] H. Nakamura and Y. Toyota. Statistical identification and optimal control of thermal power plants. Annals of the Institute of Statistical Mathematics, 40(1):1–28, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053952.

Ogata:1988:LAS

[1459] Yosihiko Ogata and Koichi Katsura. Likelihood analysis of spatial inhomogeneity for marked point patterns. Annals of the Institute of Statistical Mathematics, 40(1):29–39, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053953.

Yasue:1988:SN

[1460] Kunio Yasue, Mari Jibu, Tetsuya Misawa, and Jean-Claude Zambrini. Stochastic neurodynamics. Annals of the Institute of Statistical Mathematics, 40(1):41–59, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053954.

Kocherlakota:1988:CBP

[1461] S. Kocherlakota. On the compounded bivariate Poisson distribution: A unified treatment. Annals of the Institute of Statistical Mathematics, 40 (1):61–76, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053955.

Nishimura:1988:OTT

[1462] Kazuo Nishimura and Masaaki Sibuya.

Occupancy with two types of balls. Annals of the Institute of Statistical Mathematics, 40(1):77–91, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053956.

Nomakuchi:1988:CCC

[1463] Kentaro Nomakuchi and Toshio Sakata. Characterization of conditional covariance and unified theory in the problem of ordering random variables. Annals of the Institute of Statistical Mathematics, 40(1):93–99, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053957.

Wei:1988:EAL

[1464] William W. S. Wei and Daniel O.

Stram. An eigenvalue approach to the limiting behavior of time series aggregates. Annals of the Institute of Statistical Mathematics, 40(1):101–110, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053958.

Mukhopadhyay:1988:CRT

[1465] N. Mukhopadhyay and G. Vik. Convergence rates for two-stage confidence intervals based on U-statistics. Annals of the Institute of Statistical Mathematics, 40(1):111–117, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053959.

Sugiura:1988:ECP

[1466] Nariaki Sugiura and Tatsuya Kubokawa. Estimating common parameters of growth curve models. Annals of the Institute of Statistical Mathematics, 40 (1):119–135, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053960.

Dey:1988:SEE

[1467] Dipak K. Dey. Simultaneous estimation of eigenvalues. Annals of the Institute of Statistical Mathematics, 40 (1):137–147, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053961.

Anraku:1988:TMP

[1468] Kazuo Anraku, Akihiro Nishi, and Takashi Yanagawa. Tests for the marginal probabilities in the two-way contingency table under restricted alternatives. Annals of the Institute of Statistical Mathematics, 40(1):149–163, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053962.

Mielke:1988:CMS

[1469] P. W. Mielke, Jr. and Y. C. Yao. A class of multiple sample tests based on empirical coverages. Annals of the Institute of Statistical Mathematics, 40(1):165–178, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053963.

Tsai:1988:SWT

[1470] Kao-Tai Tsai and James A. Koziol.

Score and Wald tests for the multivariate growth curve model with missing data. Annals of the Institute of Statistical Mathematics, 40(1):179–186, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053964.

Gilliland:1988:EBS

[1471] Dennis C. Gilliland and Rohana Karunamuni. On empirical Bayes with sequential component. Annals of the Institute of Statistical Mathematics, 40 (1):187–193, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053965.

Mukerjee:1988:KFD

[1472] Rahul Mukerjee and Mausumi Sen.

Kronecker factorial designs for multiway elimination of heterogeneity. Annals of the Institute of Statistical Mathematics, 40(1):195–210, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053966.

Anonymous:1988:HCa

[1473] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 40(1):??, March 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Wada:1988:AAM

[1474] Takao Wada, Makoto Jinnouchi, and Yasuo Matsumura. Application of autoregressive modelling for the analysis of clinical and other biological data. Annals of the Institute of Statistical Mathematics, 40(2):211–227, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00052339.

Suzuki:1988:MDT

[1475] Naomichi Suzuki and Minoru Biyajima. Multiplicity distributions in a two-component branching process. Annals of the Institute of Statistical Mathematics, 40(2):229–246, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00052340.

Tran:1988:ROS

[1476] Lanh Tat Tran. Rank order statistics for time series models. Annals of the Institute of Statistical Mathematics, 40(2):247–260, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00052341.

Ahmad:1988:IFM

[1477] Khalaf E. Ahmad. Identifiability of finite mixtures using a new transform. Annals of the Institute of Statistical Mathematics, 40(2):261–265, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052342.

Nair:1988:CBE

[1478] K. R. Muraleedharan Nair and N. Unnikrishnan Nair. On characterizing the bivariate exponential and geometric distributions. Annals of the Institute of Statistical Mathematics, 40(2):267–271, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052343.

Salakbishnan:1988:RRO

[1479] N. Salakbishnan. Recurrence relations for order statistics from n independent and non-identically distributed random variables. Annals of the Institute of Statistical Mathematics, 40(2):273–277, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052344.

Konishi:1988:AED

[1480] Sadanori Konishi, Naoto Niki, and Arjun K. Gupta. Asymptotic expansions for the distribution of quadratic forms in normal variables. Annals of the Institute of Statistical Mathematics, 40(2):279–296, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052345.

Crowder:1988:AEP

[1481] Martin Crowder. Asymptotic expansions of posterior expectations, distributions and densities for stochastic processes. Annals of the Institute of Statistical Mathematics, 40(2):297–309, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00052346.

Akahira:1988:SOA

[1482] Masaflimi Akahira. Second order asymptotic optimality of estimators for a density with finite cusps. Annals of the Institute of Statistical Mathematics, 40(2):311–328, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00052347.

Sugiura:1988:ELR

[1483] Nariaki Sugiura and Yoshihiko Konno. Entropy loss and risk of improved estimators for the generalized variance and precision. Annals of the Institute of Statistical Mathematics, 40(2):329–341, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00052348.

Blough:1988:CEL

[1484] David K. Blough. Consistent estimation of location region. Annals of the Institute of Statistical Mathematics, 40(2):343–352, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00052349.

Fraser:1988:GAV

[1485] D. A. S. Fraser and P. McDunnough. On generalization of the analysis of variance. Annals of the Institute of Statistical Mathematics, 40(2):353–366, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052350.

Kikuchi:1988:IHI

[1486] Yasuki Kikuchi and Takashi Yanagawa. Incorporating historical information in testing for a trend in Poisson means. Annals of the Institute of Statistical Mathematics, 40(2):367–379, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052351.

Giri:1988:LMT

[1487] Narayan C. Giri. Locally minimax tests in symmetrical distributions. Annals of the Institute of Statistical Mathematics, 40(2):381–394, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052352.

Hara:1988:DMO

[1488] Takahiko Hara. Detection of multivariate outliers with location slippage or scale inflation in left orthogonally invariant or elliptically contoured distributions. Annals of the Institute of Statistical Mathematics, 40(2):395–406, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052353.

Jacroux:1988:MOB

[1489] Mike Jacroux and Dexter C. Whittinghill III. On the E- and MVoptimality of block designs having k ≥ v. Annals of the Institute of Statistical Mathematics, 40(2):407–418, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00052354.

Anonymous:1988:HCb

[1490] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 40(2):??, June 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hagimura:1988:ATS

[1491] S. Hagimura, T. Saitoh, and Y. Yagihara. Application of time series analysis and modern control theory to the cement plant. Annals of the Institute of Statistical Mathematics, 40(3):419–438, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053056.

Janardan:1988:RBM

[1492] K. G. Janardan. Relationship between Morisita’s model for estimating the environmental density and the generalized Eulerian numbers. Annals of the Institute of Statistical Mathematics, 40(3):439–450, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053057.

Hirotsu:1988:CEC

[1493] C. Hirotsu. A class of estimable contrasts in an age-period-cohort model. Annals of the Institute of Statistical Mathematics, 40(3):451–465, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053058.

Philippou:1988:MDO

[1494] Andreas N. Philippou. On multiparameter distributions of order k. Annals of the Institute of Statistical Mathematics, 40(3):467–475, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053059.

Nagao:1988:JSE

[1495] Hisao Nagao. On the jackknife statistics for eigenvalues and eigenvectors of a correlation matrix. Annals of the Institute of Statistical Mathematics, 40(3):477–489, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053060.

Bai:1988:LPO

[1496] Z. D. Bai, P. R. Krishnaiah, and Y. G. Yin. Limiting properties of the occurrence/exposure rate and simple risk rate. Annals of the Institute of Statistical Mathematics, 40(3): 491–505, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053061.

Kreiss:1988:SE

[1497] Jens Peter Kreiss. On stochastic estimation. Annals of the Institute of Statistical Mathematics, 40(3): 507–520, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053062.

Inagaki:1988:PTE

[1498] Nobuo Inagaki. The progressively truncated estimating functions and estimators. Annals of the Institute of Statistical Mathematics, 40(3): 521–540, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053063.

Huang:1988:ATE

[1499] Wei-Min Huang and Wei-Yann Tsai. Asymptotic theorems for estimating the distribution function under random truncation. Annals of the Institute of Statistical Mathematics, 40(3): 541–553, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053064.

Kubokawa:1988:IUT

[1500] Tatsuya Kubokawa. Inadmissibility of the uncombined two-stage estimator when additional samples are available. Annals of the Institute of Statistical Mathematics, 40(3):555–563, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053065.

Romano:1988:BM

[1501] Joseph P. Romano. Bootstrapping the mode. Annals of the Institute of Statistical Mathematics, 40(3): 565–586, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053066.

Sakamoto:1988:BAN

[1502] Yosiyuki Sakamoto and Makio Ishiguro. A Bayesian approach to nonparametric test problems. Annals of the Institute of Statistical Mathematics, 40 (3):587–602, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053067.

Eubank:1988:RTT

[1503] R. L. Eubank and V. N. LaRiccia.

Regression type tests for parametric hypotheses based on optimally selected subsets of the order statistics. Annals of the Institute of Statistical Mathematics, 40(3):603–613, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053068.

Chen:1988:IFS

[1504] Pinyuen Chen. An integrated formulation for selecting the most probable multinomial cell. Annals of the Institute of Statistical Mathematics, 40(3): 615–625, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053069.

Anonymous:1988:HCc

[1505] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 40(3):??, September 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Gil:1988:LID

[1506] Maria Angeles Gil. On the loss of information due to fuzziness in experimental observations. Annals of the Institute of Statistical Mathematics, 40 (4):627–639, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049422.

Bohning:1988:MQA

[1507] Dankmar Böhning and Bruce G. Lindsay. Monotonicity of quadraticapproximation algorithms. Annals of the Institute of Statistical Mathematics, 40(4):641–663, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049423.

Maehara:1988:TSR

[1508] Hiroshi Maehara. A threshold for the size of random caps to cover a sphere. Annals of the Institute of Statistical Mathematics, 40(4):665–670, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049424.

Deheuvels:1988:RBU

[1509] P. Deheuvels and D. Pfeifer. On a relationship between Uspensky’s theorem and Poisson approximations. Annals of the Institute of Statistical Mathematics, 40(4):671–681, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049425.

Fu:1988:ERC

[1510] James C. Fu and Robert E. Kass. The exponential rates of convergence of posterior distributions. Annals of the Institute of Statistical Mathematics, 40 (4):683–691, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049426.

Sibuya:1988:LCS

[1511] Masaaki Sibuya. Log-concavity of Stirling numbers and unimodality of Stirling distributions. Annals of the Institute of Statistical Mathematics, 40 (4):693–714, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049427.

Mussmann:1988:SJI

[1512] Dieter Mussmann. Sufficiency and Jensen’s inequality for conditional expectations. Annals of the Institute of Statistical Mathematics, 40(4): 715–726, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049428.

Withers:1988:NCI

[1513] C. S. Withers. Nonparametric confidence intervals for functions of several distributions. Annals of the Institute of Statistical Mathematics, 40 (4):727–746, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049429.

vanderMerwe:1988:ABE

[1514] A. J. van der Merwe, P. C. N. Groenewald, and C. A. van der Merwe. Approximated Bayes and empirical Bayes confidence intervals — the known variance case. Annals of the Institute of Statistical Mathematics, 40(4):747–767, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049430.

Chaturvedi:1988:SPP

[1515] Ajit Chaturvedi. On sequential procedures for the point estimation of the mean of a normal population. Annals of the Institute of Statistical Mathematics, 40(4):769–783, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049431.

Hamdy:1988:TSP

[1516] H. I. Hamdy, N. Mukhopadhyay, M. C. Costanza, and M. S. Son. Triple stage point estimation for the exponential location parameter. Annals of the Institute of Statistical Mathematics, 40(4):785–797, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049432.

Yoshida:1988:RED

[1517] Nakahiro Yoshida. Robust Mestimators in diffusion processes. Annals of the Institute of Statistical Mathematics, 40(4):799–820, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049433.

Rafajlowicz:1988:OED

[1518] Ewaryst Rafajlowicz and Wojciech Myszka. Optimum experimental design for a regression on a hypercubegeneralization of Hoel’s result. Annals of the Institute of Statistical Mathematics, 40(4):821–827, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049434.

Anonymous:1988:HCd

[1519] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 40(4):??, December 1988. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mallows:1989:ZCP

[1520] Colin L. Mallows and Vijayan N.

Nair. On a zero-crossing probability. Annals of the Institute of Statistical Mathematics, 41(1):1–8, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049104.

Itoh:1989:ACI

[1521] Yoshiaki Itoh. An application of the convolution inequality for the Fisher information. Annals of the Institute of Statistical Mathematics, 41(1):9–12, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049105.

Johnson:1989:CBC

[1522] Norman L. Johnson and Samuel Kotz. Characterization based on conditional distributions. Annals of the Institute of Statistical Mathematics, 41(1):13–17, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049106.

Falk:1989:NUA

[1523] Michael Falk. A note on uniform asymptotic normality of intermediate order statistics. Annals of the Institute of Statistical Mathematics, 41(1):19–29, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049107.

Kimeldorf:1989:FPD

[1524] George Kimeldorf and Allan R. Sampson. A framework for positive dependence. Annals of the Institute of Statistical Mathematics, 41(1):31–45, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049108.

Aki:1989:EPD

[1525] S. Aki and K. Hirano. Estimation of parameters in the discrete distributions of order k. Annals of the Institute of Statistical Mathematics, 41(1):47–61, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049109.

Dabrowska:1989:REC

[1526] Dorota M. Dabrowska, Kjell A. Doksum, and Ryozo Miura. Rank estimates in a class of semiparametric two-sample models. Annals of the Institute of Statistical Mathematics, 41(1):63–79, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049110.

Lwin:1989:EBA

[1527] T. Lwin and J. S. Maritz. Empirical Bayes approach to multiparameter estimation: with special reference to multinomial distribution. Annals of the Institute of Statistical Mathematics, 41(1):81–99, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049111.

Walter:1989:BEB

[1528] G. G. Walter and G. G. Hamedani.

Bayes empirical Bayes estimation for discrete exponential families. Annals of the Institute of Statistical Mathematics, 41(1):101–119, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049112.

Mukhopadhyay:1989:SRP

[1529] Nitis Mukhopadhyay, Pranab Kumar Sen, and Bikas Kumar Sinha. Stopping rules, permutation invariance and sufficiency principle. Annals of the Institute of Statistical Mathematics, 41(1):121–138, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049113.

Dharmadhikari:1989:ESP

[1530] A. D. Dharmadhikari, U. V. Naik-Nimbalkar, and S. Bhyri. Estimation of the scale parameter of a power law process using power law counts. Annals of the Institute of Statistical Mathematics, 41(1):139–148, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049114.

Lee:1989:TMH

[1531] Chu-In Charles Lee, Tim Robertson, and F. T. Wright. On the testing of marginal homogeneity with a onesided alternative in the analysis of variance. Annals of the Institute of Statistical Mathematics, 41(1):149–167, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049115.

Isogai:1989:UIF

[1532] Takafumi Isogai. On using influence functions for testing multivariate normality. Annals of the Institute of Statistical Mathematics, 41(1):169–186, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049116.

Taniguchi:1989:NTS

[1533] M. Taniguchi, P. R. Krishnaiah, and R. Chao. Normalizing transformations of some statistics of Gaussian ARMA processes. Annals of the Institute of Statistical Mathematics, 41(1):187–197, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049117.

Brzeskwiniewicz:1989:CPB

[1534] Henryk Brzeskwiniewicz. On the connectedness of partially balanced block designs. Annals of the Institute of Statistical Mathematics, 41(1):199–204, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049118.

Anonymous:1989:HCa

[1535] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 41(1):??, March 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Taniguchi:1989:SAD

[1536] M. Taniguchi, L. C. Zhao, P. R. Krishnaiah, and Z. D. Bai. Statistical analysis of dyadic stationary processes. Annals of the Institute of Statistical Mathematics, 41(2):205–225, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049392.

Lo:1989:CBN

[1537] Albert Y. Lo and Chung-Sing Weng.

On a class of Bayesian nonparametric estimates: II. hazard rate estimates. Annals of the Institute of Statistical Mathematics, 41(2):227–245, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049393.

Barndorff-Nielsen:1989:AEM

[1538] O. E. Barndorff-Nielsen and P. E. Jupp. Approximating exponential models. Annals of the Institute of Statistical Mathematics, 41(2):247–267, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049394.

Yanagimoto:1989:PSC

[1539] Takemi Yanagimoto and Kazuo Anraku. Possible superiority of the conditional MLE over the unconditional MLE. Annals of the Institute of Statistical Mathematics, 41(2):269–278, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049395.

Hirano:1989:CII

[1540] Katuomi Hirano and Kosei Iwase. Conditional information for an inverse Gaussian distribution with known coefficient of variation. Annals of the Institute of Statistical Mathematics, 41(2):279–287, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049396.

Mitchell:1989:IMS

[1541] Ann E. S. Mitchell. The information matrix, skewness tensor and aconnections for the general multivariate elliptic distribution. Annals of the Institute of Statistical Mathematics, 41(2):289–304, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049397.

Korwar:1989:NSD

[1542] R. M. Korwar. On a new system of discrete distributions and characterizations of several discrete distributions by equality of distributions. Annals of the Institute of Statistical Mathematics, 41(2):305–321, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049398.

Balakrishnan:1989:RRA

[1543] N. Balakrishnan. Recurrence relations among moments of order statistics from two related sets of independent and non-identically distributed random variables. Annals of the Institute of Statistical Mathematics, 41(2):323–329, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049399.

Shinozaki:1989:ICS

[1544] Nobuo Shinozaki. Improved confidence sets for the mean of a multivariate normal distribution. Annals of the Institute of Statistical Mathematics, 41(2):331–346, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049400.

Hawkins:1989:MAR

[1545] D. L. Hawkins and Chien-Pai Han.

A minimum average risk approach to shrinkage estimators of the normal mean. Annals of the Institute of Statistical Mathematics, 41(2):347–363, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049401.

Shao:1989:ADW

[1546] Jun Shao. Asymptotic distribution of the weighted least squares estimator. Annals of the Institute of Statistical Mathematics, 41(2):365–382, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049402.

Cox:1989:SSB

[1547] Dennis Cox and Eunmee Koh. A smoothing spline based test of model adequacy in polynomial regression. Annals of the Institute of Statistical Mathematics, 41(2):383–400, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049403.

Milton:1989:CIS

[1548] Roy C. Milton and M. Haseeb Rizvi. On computation of integrals for selection from multivariate normal populations on the basis of distances. Annals of the Institute of Statistical Mathematics, 41(2):401–408, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049404.

Sinha:1989:CCG

[1549] Kishore Sinha and Sanpei Kageyama. Composite construction of group divisible designs. Annals of the Institute of Statistical Mathematics, 41(2):409–414, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049405.

Anonymous:1989:HCb

[1550] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 41(2):??, June 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Pham:1989:AND

[1551] Dinh Tuan Pham, Joachim Möcks, and Lothar Sroka. Asymptotic normality of double-indexed linear permutation statistics. Annals of the Institute of Statistical Mathematics, 41(3): 415–427, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050659.

Vos:1989:FES

[1552] Paul W. Vos. Fundamental equations for statistical submanifolds with applications to the Bartlett correction. Annals of the Institute of Statistical Mathematics, 41(3):429–450, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050660.

Zografos:1989:LPS

[1553] K. Zografos, K. Ferentinos, and T. Papaioannou. Limiting properties of some measures of information. Annals of the Institute of Statistical Mathematics, 41(3):451–460, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050661.

Blough:1989:MSP

[1554] David K. Blough. Multivariate symmetry via projection pursuit. Annals of the Institute of Statistical Mathematics, 41(3):461–475, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050662.

Kubokawa:1989:CEC

[1555] Tatsuya Kubokawa. Closer estimators of a common mean in the sense of Pitman. Annals of the Institute of Statistical Mathematics, 41(3): 477–484, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050663.

Akai:1989:SEM

[1556] Toyoaki Akai. Simultaneous estimation of means of classified normal observations. Annals of the Institute of Statistical Mathematics, 41(3): 485–502, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050664.

Yu:1989:MIE

[1557] Qiquing Yu. Methodology for the invariant estimation of a continuous distribution function. Annals of the Institute of Statistical Mathematics, 41 (3):503–520, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050665.

Martinsek:1989:SER

[1558] Adam T. Martinsek. Sequential estimation in regression models using analogues of trimmed means. Annals of the Institute of Statistical Mathematics, 41(3):521–540, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050666.

Zhou:1989:BES

[1559] Xian Zhou and S. Rao Jammalamadaka. Bahadur efficiencies of spacings tests for goodness of fit. Annals of the Institute of Statistical Mathematics, 41(3):541–553, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050667.

LeBreton:1989:BLS

[1560] Alain Le Breton and Dinh Tuan Pham. On the bias of the least squares estimator for the first order autoregressive process. Annals of the Institute of Statistical Mathematics, 41(3):555–563, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050668.

Moller:1989:RCS

[1561] Jesper Møller. On the rate of convergence of spatial birth-and-death processes. Annals of the Institute of Statistical Mathematics, 41(3):

565–581, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050669.

Ogata:1989:LES

[1562] Yosihiko Ogata and Masaharu Tanemura. Likelihood estimation of softcore interaction potentials for Gibbsian point patterns. Annals of the Institute of Statistical Mathematics, 41 (3):583–600, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050670.

Kalbakam:1989:EIS

[1563] S. Kalbakam and G. Abivarignan. On exhibiting inventory systems with Erlangian lifetimes under renewal demands. Annals of the Institute of Statistical Mathematics, 41(3):601– 616, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050671.

Al-Hussaini:1989:FRI

[1564] Essam K. Al-Hussaini and Nagi S. Abd-El-Hakim. Failure rate of the inverse Gaussian–Weibull mixture model. Annals of the Institute of Statistical Mathematics, 41(3): 617–622, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050672.

Anonymous:1989:HCc

[1565] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 41(3):??, September 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Amari:1989:FIU

[1566] Shun ichi Amari. Fisher information under restriction of Shannon information in multi-terminal situations. Annals of the Institute of Statistical Mathematics, 41(4):623–648, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057730.

Wu:1989:CAW

[1567] Tiee-Jian Wu. Contiguous alternatives which preserve Cramér-type large deviations for a general class of statistics. Annals of the Institute of Statistical Mathematics, 41(4):649–660, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057731.

Dallas:1989:SPR

[1568] A. C. Dallas. Some properties of record values coming from the geometric distribution. Annals of the Institute of Statistical Mathematics, 41 (4):661–669, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00057732.

Kyriakoussis:1989:CPS

[1569] A. Kyriakoussis and H. Papageorgiou. On characterization of power series distributions by a marginal distribution and a regression function. Annals of the Institute of Statistical Mathematics, 41(4):671–676, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057733.

Ramallingam:1989:SCE

[1570] T. Ramallingam. Symbolic computing the exact distributions of Lstatistics from a uniform distribution. Annals of the Institute of Statistical Mathematics, 41(4):677–681, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057734.

Joe:1989:EEO

[1571] Harry Joe. Estimation of entropy and other functionals of a multivariate density. Annals of the Institute of Statistical Mathematics, 41(4):683–697, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057735.

Skibinsky:1989:AEB

[1572] Morris Skibinsky and Andrew L. Rukhin. Admissible estimators of binomial probability and the inverse Bayes rule map. Annals of the Institute of Statistical Mathematics, 41 (4):699–716, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00057736.

Sarkar:1989:CAS

[1573] Nityananda Sarkar. Comparisons among some estimators in misspecified linear models with multicollinearity. Annals of the Institute of Statistical Mathematics, 41(4):717–724, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057737.

Akahira:1989:HOA

[1574] Masafumi Akahira and Kei Takeuchi.

Higher order asymptotics in estimation for two-sided Weibull type distributions. Annals of the Institute of Statistical Mathematics, 41(4):725–752, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057738.

Aki:1989:APS

[1575] Sigeo Aki and Nobuhisa Kashiwagi. Asymptotic properties of some goodness-of-fit tests based on the L_{1}- norm. Annals of the Institute of Statistical Mathematics, 41(4):753–764, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057739.

Nakano:1989:TPP

[1576] Junji Nakano and Shigemi Tagami. A test for the presence of pure feedback in multivariate dynamic stochastic systems. Annals of the Institute of Statistical Mathematics, 41(4):765–779, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00057740.

Antoniadis:1989:PMN

[1577] Anestis Antoniadis. A penalty method for nonparametric estimation of the intensity function of a counting process. Annals of the Institute of Statistical Mathematics, 41(4):781–807, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00057741.

Chao:1989:LTC

[1578] M. T. Chao and James C. Fu. A limit theorem of certain repairable systems. Annals of the Institute of Statistical Mathematics, 41(4):809–818, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00057742.

Anonymous:1989:HCd

[1579] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 41(4):??, December 1989. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shaked:1990:BAQ

[1580] Moshe Shaked and Nozer D. Singpurwalla. A Bayesian approach for quantile and response probability estimation with applications to reliability. Annals of the Institute of Statistical Mathematics, 42(1):1–19, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050775.

Chung:1990:CBQ

[1581] Chang-Jo F. Chung, Miklós Csörgő, and Lajos Horváth. Confidence bands for quantile function under random censorship. Annals of the Institute of Statistical Mathematics, 42(1):21–36, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050776.

Hosoya:1990:IAH

[1582] Yuzo Hosoya. Information amount and higher-order efficiency in estimation. Annals of the Institute of Statistical Mathematics, 42(1):37–49, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050777.

Venter:1990:ETN

[1583] J. H. Venter and S. J. Steel. Estimation of two normal means which may be common. Annals of the Institute of Statistical Mathematics, 42(1):51–64, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050778.

Hwang:1990:MCM

[1584] Tea-Yuan Hwang and Chin-Yuan Hu. More comparisons of MLE with UMVUE for exponential families. Annals of the Institute of Statistical Mathematics, 42(1):65–75, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050779.

Madi:1990:ERS

[1585] Mohamed Madi and Kam-Wah Tsui.

Estimation of the ratio of the scale parameters of two exponential distributions with unknown location parameters. Annals of the Institute of Statistical Mathematics, 42(1):77–87, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050780.

Inagaki:1990:PES

[1586] Nobuo Inagaki and Toshihabu Hayashi. Parameter estimation for the simple self-correcting point process. Annals of the Institute of Statistical Mathematics, 42(1):89–98, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050781.

Jerdack:1990:NTR

[1587] George R. Jerdack and Pranab Kumar Sen. Nonparametric test of restricted interchangeability. Annals of the Institute of Statistical Mathematics, 42(1):99–114, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050782.

Wang:1990:SAR

[1588] Soujin Wang. Saddlepoint approximations in resampling analysis. Annals of the Institute of Statistical Mathematics, 42(1):115–131, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050783.

Hosmane:1990:SLR

[1589] Balakrishna S. Hosmane. Smoothing of likelihood ratio statistic for equiprobable multinomial goodness-of-fit. Annals of the Institute of Statistical Mathematics, 42(1):133–147, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050784.

Chen:1990:NDS

[1590] Shande Chen and Govind S. Mudholkar. Null distribution of the sum of squared z-transforms in testing complete independence. Annals of the Institute of Statistical Mathematics, 42(1):149–155, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050785.

Pillai:1990:MLF

[1591] R. N. Pillai. On Mittag-Leffler functions and related distributions. Annals of the Institute of Statistical Mathematics, 42(1):157–161, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050786.

Baksalary:1990:PBV

[1592] Jerzy K. Baksalary and P. D. Puri.

Pairwise-balanced, variance-balanced and resistant incomplete block designs revisited. Annals of the Institute of Statistical Mathematics, 42(1):163–171, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050787.

Jacroux:1990:SOD

[1593] Mike Jacroux. Some optimal designs for comparing a set of test treatments with a set of controls. Annals of the Institute of Statistical Mathematics, 42(1):173–185, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050788.

Abaffy:1990:CSD

[1594] Jozsef Abaffy and Emilio Spedicato. A class of scaled direct methods for linear systems. Annals of the Institute of Statistical Mathematics, 42(1):187–201, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050789.

Anonymous:1990:HCa

[1595] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 42(1):??, March 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mase:1990:MCG

[1596] Shigeru Mase. Mean characteristics of Gibbsian point processes. Annals of the Institute of Statistical Mathematics, 42(2):203–220, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050833.

Yoshida:1990:ABE

[1597] Nakahiro Yoshida. Asymptotic behavior of M-estimator and related random field for diffusion process. Annals of the Institute of Statistical Mathematics, 42(2):221–251, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050834.

Rajarshi:1990:BMS

[1598] M. B. Rajarshi. Bootstrap in Markovsequences based on estimates of transition density. Annals of the Institute of Statistical Mathematics, 42(2):253–268, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050835.

Guiasu:1990:CMP

[1599] Silviu Guiasu. A classification of the main probability distributions by minimizing the weighted logarithmic measure of deviation. Annals of the Institute of Statistical Mathematics, 42 (2):269–279, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050836.

Alzaid:1990:MAS

[1600] Abdulhamid A. Alzaid. A moment’s approach to some characterization problems. Annals of the Institute of Statistical Mathematics, 42(2):281–285, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050837.

Cline:1990:OKE

[1601] Daren B. H. Cline. Optimal kernel estimation of densities. Annals of the Institute of Statistical Mathematics, 42(2):287–303, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050838.

Tran:1990:RKD

[1602] Lanh Tat Tran. Recursive kernel density estimators under a weak dependence condition. Annals of the Institute of Statistical Mathematics, 42 (2):305–329, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050839.

Kubokawa:1990:ECM

[1603] Tatsuya Kubokawa and Yoshihiko Konno. Estimating the covariance matrix and the generalized variance under a symmetric loss. Annals of the Institute of Statistical Mathematics, 42(2):331–343, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050840.

Jureckova:1990:EIE

[1604] Jana Jurecková and Pranab Kumar Sen. Effect of the initial estimator on the asymptotic behavior of one-step Mestimator. Annals of the Institute of Statistical Mathematics, 42(2):345–357, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050841.

Hayakawa:1990:TMD

[1605] T. Hayakawa. On tests for the mean direction of the Langevin distribution. Annals of the Institute of Statistical Mathematics, 42(2):359–373, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050842.

Tsai:1990:AER

[1606] Ming-Tan M. Tsai and Pranas Kumar Sen. Asymptotically efficient rank MANOVA tests for restricted alternatives in randomized block designs. Annals of the Institute of Statistical Mathematics, 42(2):375–385, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050843.

Chen:1990:CBL

[1607] X. Chen. On the convergence of Broyden-like methods for nonlinear equations with nondifferentiable terms. Annals of the Institute of Statistical Mathematics, 42(2):387–401, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050844.

Anonymous:1990:HCb

[1608] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 42(2):??, June 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Ogata:1990:MCM

[1609] Yosihiko Ogata. A Monte Carlo method for an objective Bayesian procedure. Annals of the Institute of Statistical Mathematics, 42(3): 403–433, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049299.

Bolfarine:1990:BLP

[1610] Heleno Bolfarine. Bayesian linear prediction in finite populations. Annals of the Institute of Statistical Mathematics, 42(3):435–444, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049300.

Jensen:1990:SER

[1611] E. B. Jensen, K. Kiêu, and H. J. G. Gundersen. On the stereological estimation of reduced moment measures. Annals of the Institute of Statistical Mathematics, 42(3):445–461, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049301.

Frankl:1990:SGA

[1612] Peter Frankl and Hiroshi Maehara. Some geometric applications of the beta distribution. Annals of the Institute of Statistical Mathematics, 42 (3):463–474, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049302.

Yamazaki:1990:IRS

[1613] Genji Yamazaki. Invariance relations in single server queues with LCFS service discipline. Annals of the Institute of Statistical Mathematics, 42(3): 475–488, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049303.

Yoshida:1990:REP

[1614] Nakahiro Yoshida and Toshiharu Hayashi. On the robust estimation in Poisson processes with periodic intensities. Annals of the Institute of Statistical Mathematics, 42(3): 489–507, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049304.

Shaked:1990:PSC

[1615] Moshe Shaked and J. George Shanthikumar. Parametric stochastic convexity and concavity of stochastic processes. Annals of the Institute of Statistical Mathematics, 42(3): 509–531, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00049305.

Fujikoshi:1990:GCM

[1616] Y. Fujikoshi, T. Kanda, and N. Tanimura. The growth curve model with an autoregressive covariance structure. Annals of the Institute of Statistical Mathematics, 42(3):533–542, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049306.

Lu:1990:SNC

[1617] Jye-Chyl Lu and Gouri K. Bhattacharyya. Some new constructions of bivariate Weibull models. Annals of the Institute of Statistical Mathematics, 42(3):543–559, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049307.

Konishi:1990:III

[1618] Sadanori Konishi and C. G. Khatri. Inferences on interclass and intraclass correlations in multivariate familial data. Annals of the Institute of Statistical Mathematics, 42(3): 561–580, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049308.

Bansal:1990:TLH

[1619] Naveen K. Bansal. Testing linear hypotheses in errors in variables model. Annals of the Institute of Statistical Mathematics, 42(3):581–596, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049309.

Kozlowska:1990:OSR

[1620] Maria Kozlowska and Ryszard Walkowiak. E-optimality of some row and column designs. Annals of the Institute of Statistical Mathematics, 42(3): 597–602, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049310.

Anonymous:1990:HCc

[1621] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 42(3):??, September 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Buhler:1990:HSL

[1622] Wolfgang J. Bühler, Prem S. Puri, and Hans-J. Schuh. Hitting straight lines by compound Poisson process paths. Annals of the Institute of Statistical Mathematics, 42(4):603–621, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481140.

Cohen:1990:AEP

[1623] Arthur Cohen and Harold B. Sackrowitz. Admissibility of estimators of the probability of unobserved outcomes. Annals of the Institute of Statistical Mathematics, 42(4):623–636, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481141.

See corrections [1807].

Karunamuni:1990:EBA

[1624] R. J. Karunamuni. On the empirical Bayes approach to multiple decision problems with sequential components. Annals of the Institute of Statistical Mathematics, 42(4):637–655, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02481142.

Parsian:1990:AEN

[1625] Ahmad Parsian. On the admissibility of an estimator of a normal mean vector under a linex loss function. Annals of the Institute of Statistical Mathematics, 42(4):657–669, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481143.

Jureckova:1990:ARB

[1626] Jana Jurecková and A. H. Welsh. Asymptotic relations between L- and M-estimators in the linear model. Annals of the Institute of Statistical Mathematics, 42(4):671–698, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481144.

Isogai:1990:NER

[1627] Eiichi Isogai. Nonparametric estimation of a regression function by delta sequences. Annals of the Institute of Statistical Mathematics, 42 (4):699–708, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02481145.

Leger:1990:BCT

[1628] Christian Léger and Joseph P. Romano. Bootstrap choice of tuning parameters. Annals of the Institute of Statistical Mathematics, 42(4):709–735, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481146.

Shao:1990:BEA

[1629] Jun Shao. Bootstrap estimation of the asymptotic variances of statistical functionals. Annals of the Institute of Statistical Mathematics, 42(4): 737–752, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481147.

Bose:1990:BMA

[1630] Arup Bose. Bootstrap in moving average models. Annals of the Institute of Statistical Mathematics, 42(4): 753–768, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481148.

Fujikoshi:1990:SRS

[1631] Yasunori Fujikoshi and Chinubal G.

Khatri. A study of redundancy of some variables in covariate discriminant analysis. Annals of the Institute of Statistical Mathematics, 42(4): 769–782, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481149.

Chatterjee:1990:SDS

[1632] Kashinath Chatterjee. Search designs for searching for one among the twoand three-factor interaction effects in the general symmetric and asymmetric factorials. Annals of the Institute of Statistical Mathematics, 42(4): 783–803, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02481150.

Zhang:1990:NSC

[1633] Shao Liang Zhang and Yoshio Oyanagi.

A necessary and sufficient convergence condition of orthomin(k) methods for least squares problem with weight.

Annals of the Institute of Statistical Mathematics, 42(4):805–811, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02481151.

Anonymous:1990:HCd

[1634] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 42(4):??, December 1990. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Besag:1991:BIR

[1635] Julian Besag, Jeremy York, and Annie Mollié. Bayesian image restoration, with two applications in spatial statistics. Annals of the Institute of Statistical Mathematics, 43(1):1–20, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00116466.

Anonymous:1991:D

[1636] Anonymous. Discussion. Annals of the Institute of Statistical Mathematics, 43(1):21–45, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00116467.

Besag:1991:R

[1637] Julian Besag. Rejoinder. Annals of the Institute of Statistical Mathematics, 43(1):45–59, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00116468.

Stein:1991:KAK

[1638] Michael L. Stein. A kernel approximation to the kriging predictor of a spatial process. Annals of the Institute of Statistical Mathematics, 43(1):61–75, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116469.

Kashiwagi:1991:BDS

[1639] Nobuhisa Kashiwagi. Bayesian detection of structural changes. Annals of the Institute of Statistical Mathematics, 43(1):77–93, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00116470.

Blaesild:1991:YTD

[1640] P. Blæsild. Yokes and tensors derived from yokes. Annals of the Institute of Statistical Mathematics, 43(1):95–113, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116471.

Rutherford:1991:EIN

[1641] B. Rutherford and S. Yakowitz. Error inference for nonparametric regression. Annals of the Institute of Statistical Mathematics, 43(1):115–129, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116472.

Morgenthaler:1991:ITF

[1642] Stephan Morgenthaler and Clifford Hurvich. An information-theoretic framework for robustness. Annals of the Institute of Statistical Mathematics, 43(1):131–146, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00116473.

Withers:1991:CMS

[1643] C. S. Withers. A class of multiple shrinkage estimators. Annals of the Institute of Statistical Mathematics, 43 (1):147–156, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00116474.

Konno:1991:NEE

[1644] Yoshihiko Konno. A note on estimating eigenvalues of scale matrix of the multivariate F -distribution. Annals of the Institute of Statistical Mathematics, 43(1):157–165, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116475.

From:1991:EMN

[1645] Steven G. From. Estimating means from a non-I.I.D. mixture of Poisson samples. Annals of the Institute of Statistical Mathematics, 43(1):167–179, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116476.

Akahira:1991:OAE

[1646] Masafumi Akahira. The 3/2th and 2nd order asymptotic efficiency of maximum probability estimators in nonregular cases. Annals of the Institute of Statistical Mathematics, 43(1):181–195, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116477.

Charalambides:1991:GED

[1647] Ch. A. Charalambides. On a generalized Eulerian distribution. Annals of the Institute of Statistical Mathematics, 43(1):197–206, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00116478.

Anonymous:1991:HCa

[1648] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 43(1):??, March 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Imoto:1991:ABA

[1649] Masajiro Imoto. An application of Bayesian (ABIC) smoothing methods to estimating space and time variations in the magnitude distributions of earthquakes. Annals of the Institute of Statistical Mathematics, 43(2):207–225, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00118632.

Bansal:1991:BEN

[1650] N. K. Bansal and M. Bhandary.

Bayes estimation of number of signals. Annals of the Institute of Statistical Mathematics, 43(2):227–243, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00118633.

Eguchi:1991:GLN

[1651] Shinto Eguchi. A geometric look at nuisance parameter effect of local powers in testing hypothesis. Annals of the Institute of Statistical Mathematics, 43(2):245–260, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00118634.

Sibuya:1991:BTI

[1652] Masaaki Sibuya. Bonferroni-type inequalities; Chebyshev-type inequalities for the distributions on [0, n ]. Annals of the Institute of Statistical Mathematics, 43(2):261–285, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00118635.

Korwar:1991:CDM

[1653] R. M. Korwar. On characterizations of distributions by mean absolute deviation and variance bounds. Annals of the Institute of Statistical Mathematics, 43(2):287–295, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00118636.

Akahira:1991:BME

[1654] Masafumi Akahira and Kei Takeuchi. Bootstrap method and empirical process. Annals of the Institute of Statistical Mathematics, 43(2):297–310, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00118637.

Pawlak:1991:AEP

[1655] Miroslaw Pawlak. On the almost everywhere properties of the kernel regression estimate. Annals of the Institute of Statistical Mathematics, 43(2):311–326, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00118638.

Karunamuni:1991:OCP

[1656] R. J. Karunamuni and K. L. Mehra.

Optimal convergence properties of kernel density estimators without differentiability conditions. Annals of the Institute of Statistical Mathematics, 43(2):327–346, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00118639.

Kaur:1991:EOM

[1657] Amarjot Kaur and Harshinder Singh.

On the estimation of ordered means of two exponential populations. Annals of the Institute of Statistical Mathematics, 43(2):347–356, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00118640.

Ahmad:1991:ECM

[1658] Manzoor Ahmad, Y. P. Chaubey, and B. K. Sinha. Estimation of a common mean of several univariate inverse Gaussian populations. Annals of the Institute of Statistical Mathematics, 43(2):357–367, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00118641.

Sarkar:1991:STI

[1659] Sanat K. Sarkar. Stein-type improvements of confidence intervals for the generalized variance. Annals of the Institute of Statistical Mathematics, 43(2):369–375, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00118642.

Jacroux:1991:DCO

[1660] Mike Jacroux and Rita Saha Ray. On the determination and construction of optimal row-column designs having unequal row and column sizes. Annals of the Institute of Statistical Mathematics, 43(2):377–390, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00118643.

Habibullah:1991:MSD

[1661] Mohamed Habibullah and S. K. Katti. A modified steepest descent method with applications to maximizing likelihood functions. Annals of the Institute of Statistical Mathematics, 43 (2):391–404, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00118644.

Anonymous:1991:HCb

[1662] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 43(2):??, June 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Anonymous:1991:A

[1663] Anonymous. Announcement. Annals of the Institute of Statistical Mathematics, 43(3):405, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00053363.

Takanami:1991:EAT

[1664] Tetsuo Takanami and Genshiro Kitagawa. Estimation of the arrival times of seismic waves by multivariate time series model. Annals of the Institute of Statistical Mathematics, 43(3):407–433, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053364.

Kishino:1991:DHE

[1665] Hirohisa Kishino, Hidehiro Kato, Fujio Kasamatsu, and Yoshihiro Fujise. Detection of heterogeneity and estimation of population characteristics from the field survey data: 1987/ 88 Japanese feasibility study of the southern hemisphere Minke whales. Annals of the Institute of Statistical Mathematics, 43(3):435–453, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053365.

Jensen:1991:RDS

[1666] Eva B. Vedel Jensen. Recent developments in the stereological analysis of particles. Annals of the Institute of Statistical Mathematics, 43(3): 455–468, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053366.

Higuchi:1991:FDC

[1667] T. Higuchi. Frequency domain characteristics of linear operator to decompose a time series into the multicomponents. Annals of the Institute of Statistical Mathematics, 43(3): 469–492, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053367.

Abraham:1991:NTS

[1668] Bovas Abraham and A. Thavaneswaran. A nonlinear time series model and estimation of missing observations. Annals of the Institute of Statistical Mathematics, 43(3):493–504, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053368.

Findley:1991:CPB

[1669] David F. Findley. Counterexamples to parsimony and BIC. Annals of the Institute of Statistical Mathematics, 43(3):505–514, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053369.

Vos:1991:GD

[1670] Paul W. Vos. Geometry of fdivergence. Annals of the Institute of Statistical Mathematics, 43(3): 515–537, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053370.

Yanagimoto:1991:CEP

[1671] Takemi Yanagimoto and Eiji Yamamoto. Constructing elementary procedures for inference of the gamma distribution. Annals of the Institute of Statistical Mathematics, 43(3): 539–550, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053371.

Baringhaus:1991:CCT

[1672] Ludwig Baringhaus and Norbert Henze. A class of consistent tests for exponentiality based on the empirical Laplace transform. Annals of the Institute of Statistical Mathematics, 43 (3):551–564, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053372.

Hoglund:1991:BSS

[1673] Thomas Höglund. Bounds for the sample size to justify normal approximation of the confidence level. Annals of the Institute of Statistical Mathematics, 43(3):565–578, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053373.

Phadia:1991:MAP

[1674] E. G. Phadia and Qiqing Yu. Minimaxity and admissibility of the product limit estimator. Annals of the Institute of Statistical Mathematics, 43 (3):579–596, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053374.

Galantai:1991:AEP

[1675] A. Galantai. Analysis of error propagation in the ABS class for linear systems. Annals of the Institute of Statistical Mathematics, 43(3):597–603, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053375.

Anonymous:1991:ADS

[1676] Anonymous. AISM data 43-3-01 seismograms of foreshocks of 1982 Urakawa-oki earthquake. Annals of the Institute of Statistical Mathematics, 43 (3):605–606, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00053376.

Anonymous:1991:HCc

[1677] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 43(3):??, September 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shohoji:1991:PIG

[1678] Takao Shohoji, Kouji Kanefuji, Takahiro Sumiya, and Tao Qin. A prediction of individual growth of height according to an empirical Bayesian approach. Annals of the Institute of Statistical Mathematics, 43(4):607–619, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121642.

Xu:1991:DGS

[1679] Daming Xu. Differential geometrical structures related to forecasting error variance ratios. Annals of the Institute of Statistical Mathematics, 43(4): 621–646, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00121643.

Wang:1991:DSS

[1680] P. C. Wang. Diagnostics and score statistics in regression. Annals of the Institute of Statistical Mathematics, 43 (4):647–656, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00121644.

Iwasa:1991:AUT

[1681] Manabu Iwasa. Admissibility of unbiased tests for a composite hypothesis with a restricted alternative. Annals of the Institute of Statistical Mathematics, 43(4):657–665, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121645.

Shi:1991:LB

[1682] Sheng G. Shi. Local bootstrap.

Annals of the Institute of Statistical Mathematics, 43(4):667–676, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121646.

Kubokawa:1991:REC

[1683] T. Kubokawa, C. Robert, and A. K. Md. E. Saleh. Robust estimation of common regression coefficients under spherical symmetry. Annals of the Institute of Statistical Mathematics, 43(4):677–688, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00121647.

Bhandapy:1991:RED

[1684] Madhusudan Bhandapy. Robust Mestimation of a dispersion matrix with a structure. Annals of the Institute of Statistical Mathematics, 43(4): 689–705, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00121648.

Lee:1991:JVE

[1685] Youngjo Lee. Jackknife variance estimators of the location estimator in the one-way random-effects model. Annals of the Institute of Statistical Mathematics, 43(4):707–714, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00121649.

Shiraishi:1991:SIB

[1686] Taka aki Shiraishi. Statistical inference based on aligned ranks for two-way MANOVA with interaction. Annals of the Institute of Statistical Mathematics, 43(4):715–734, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121650.

Yanagimoto:1991:EMT

[1687] Takemi Yanagimoto. Estimating a model through the conditional MLE. Annals of the Institute of Statistical Mathematics, 43(4):735–746, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121651.

Ghorai:1991:ESQ

[1688] J. K. Ghorai. Estimation of a smooth quantile function under the proportional hazards model. Annals of the Institute of Statistical Mathematics, 43(4):747–760, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121652.

Krishnamoorthy:1991:ECM

[1689] K. Krishnamoorthy. Estimation of a common multivariate normal mean vector. Annals of the Institute of Statistical Mathematics, 43(4):761–771, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121653.

Saran:1991:SJL

[1690] Jagdish Saran. On some joint laws in fluctuations of sums of random variables. Annals of the Institute of Statistical Mathematics, 43(4):773–791, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00121654.

Nguyen:1991:NCM

[1691] Truc T. Nguyen and Allan R. Sampson. A note on characterizations of multivariate stable distributions. Annals of the Institute of Statistical Mathematics, 43(4):793–801, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00121655.

Anonymous:1991:HCd

[1692] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 43(4):??, December 1991. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Musmeci:1992:STC

[1693] F. Musmeci and D. Vere-Jones. A space-time clustering model for historical earthquakes. Annals of the Institute of Statistical Mathematics, 44(1):1–11, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048666.

Thomson:1992:SEU

[1694] P. J. Thomson. Signal estimation using stochastic velocity models and irregular arrays. Annals of the Institute of Statistical Mathematics, 44(1):13–25, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048667.

Zimmerman:1992:MSP

[1695] Dale L. Zimmerman and Noel Cressie. Mean squared prediction error in the spatial linear model with estimated covariance parameters. Annals of the Institute of Statistical Mathematics, 44(1):27–43, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00048668.

Picard:1992:SMR

[1696] Dominique B. Picard. Statistical morphisms and related invariance properties. Annals of the Institute of Statistical Mathematics, 44(1):45–61, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048669.

Mora:1992:GED

[1697] Marianne Mora. Geometrical expansions for the distributions of the score vector and the maximum likelihood estimator. Annals of the Institute of Statistical Mathematics, 44(1):63–83, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048670.

Tolley:1992:LSP

[1698] H. Dennis Tolley and Kenneth G. Manton. Large sample properties of estimates of a discrete grade of membership model. Annals of the Institute of Statistical Mathematics, 44(1):85–95, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048671.

Mathal:1992:FMG

[1699] A. M. Mathal and P. G. Moschopoulos. A form of multivariate gamma distribution. Annals of the Institute of Statistical Mathematics, 44(1):97–106, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048672.

Lee:1992:LCF

[1700] Chu-In Charles Lee. On Laplace continued fraction for the normal integral. Annals of the Institute of Statistical Mathematics, 44(1):107–120, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048673.

Kanagawa:1992:LTM

[1701] S. Kanagawa, Y. Mochizuki, and H. Tanaka. Limit theorems for the minimum interpoint distance between any pair of i.i.d. random points in R^{d}. Annals of the Institute of Statistical Mathematics, 44(1):121–131, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048674.

Liang:1992:MEB

[1702] TaChen Liang and S. Panchapakesan. On a monotone empirical Bayes test procedure in geometric model. Annals of the Institute of Statistical Mathematics, 44(1):133–140, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048675.

Nomakuchi:1992:NUM

[1703] Kentaro Nomakuchi. A note on the uniformly most powerful tests in the presence of nuisance parameters.

Annals of the Institute of Statistical Mathematics, 44(1):141–145, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048676.

Watamori:1992:TGL

[1704] Yoko Watamori. Tests for a given linear structure of the mean direction of the Langevin distribution. Annals of the Institute of Statistical Mathematics, 44(1):147–156, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048677.

Chaturvedi:1992:FWC

[1705] Ajit Chaturvedi, N. D. Shukla, and Pramod S. Shukla. Fixed-width confidence intervals for contrasts in the means. Annals of the Institute of Statistical Mathematics, 44(1):157–167, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048678.

English:1992:ICE

[1706] Brian J. English, Raphael Gillett, and Michael J. Phillips. Inequalities concerning the expected selection differentials. Annals of the Institute of Statistical Mathematics, 44(1):169–175, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00048679.

Balakrishnan:1992:GRI

[1707] N. Balakrishnan, S. M. Bendre, and H. J. Malik. General relations and identities for order statistics from nonindependent non-identical variables. Annals of the Institute of Statistical Mathematics, 44(1):177–183, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048680.

Shirakura:1992:SSD

[1708] Teruhiro Shirakura and Shinsei Tazawa. A series of search designs for 2^{m} factorial designs of resolution V which permit search of one or two unknown extra three-factor interactions. Annals of the Institute of Statistical Mathematics, 44(1):185–196, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00048681.

Bohning:1992:MLR

[1709] Dankmar Böhning. Multinomial logistic regression algorithm. Annals of the Institute of Statistical Mathematics, 44(1):197–200, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00048682.

Anonymous:1992:HCa

[1710] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 44(1):??, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

vanderMerwe:1992:BIN

[1711] A. J. van der Merwe, C. A. van der Merwe, and P. C. N. Groenewald. Bayesian inferences on nonlinear functions of the parameters in linear regression. Annals of the Institute of Statistical Mathematics, 44(2):201–211, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00058636.

Ramanathan:1992:OER

[1712] T. V. Ramanathan and M. B. Rajarshi.

Optimal estimation in random coefficient regression models. Annals of the Institute of Statistical Mathematics, 44(2):213–227, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00058637.

Bischoff:1992:EOD

[1713] Wolfgang Bischoff. On exact D-optimal designs for regression models with correlated observations. Annals of the Institute of Statistical Mathematics, 44(2):229–238, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00058638.

Kumon:1992:INM

[1714] Masayuki Kumon. Identification of non-minimum phase transfer function using higher-order spectrum. Annals of the Institute of Statistical Mathematics, 44(2):239–260, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00058639.

Vos:1992:MDE

[1715] Paul W. Vos. Minimum f-divergence estimators and quasi-likelihood functions. Annals of the Institute of Statistical Mathematics, 44(2):261–279, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00058640.

Venter:1992:SCS

[1716] J. H. Venter and S. J. Steel. Some contributions to selection and estimation in the normal linear model. Annals of the Institute of Statistical Mathematics, 44(2):281–297, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00058641.

Rukhin:1992:ARB

[1717] Andrew L. Rukhin. Asymptotic risk behavior of mean vector and variance estimators and the problem of positive normal mean. Annals of the Institute of Statistical Mathematics, 44(2):299–311, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00058642.

Tsukahara:1992:RET

[1718] Hideatsu Tsukahara. A rank estimator in the two-sample transformation model with randomly censored data. Annals of the Institute of Statistical Mathematics, 44(2):313–333, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00058643.

Albers:1992:AET

[1719] Willem Albers. Asymptotic expansions for two-stage rank tests.

Annals of the Institute of Statistical Mathematics, 44(2):335–356, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00058644.

Giri:1992:OTE

[1720] N. Giri. On an optimum test of the equality of two covariance matrices. Annals of the Institute of Statistical Mathematics, 44(2):357–362, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00058645.

Aki:1992:WTP

[1721] Sigeo Aki. Waiting time problems for a sequence of discrete random variables. Annals of the Institute of Statistical Mathematics, 44(2):363–378, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00058646.

Deng:1992:SCU

[1722] Lih-Yuan Deng and E. Olusegun George. Some characterizations of the uniform distribution with applications to random number generation. Annals of the Institute of Statistical Mathematics, 44(2):379–385, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00058647.

Babu:1992:ESI

[1723] Gutti Jogesh Babu and C. Radhakrishna Rao. Expansions for statistics involving the mean absolute deviations. Annals of the Institute of Statistical Mathematics, 44(2):387–403, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00058648.

Anonymous:1992:HCb

[1724] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 44(2):??, June 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Crowder:1992:BPB

[1725] Martin Crowder. Bayesian priors based on a parameter transformation using the distribution function. Annals of the Institute of Statistical Mathematics, 44(3):405–416, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00050695.

Eaves:1992:PME

[1726] D. M. Eaves and T. Chang. Posterior mode estimation for the generalized linear model. Annals of the Institute of Statistical Mathematics, 44 (3):417–434, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00050696.

Singh:1992:EBR

[1727] R. S. Singh and Laisheng Wei. Empirical Bayes with rates and best rates of convergence in u(x)C(θ) exp(−x/θ)- family: Estimation case. Annals of the Institute of Statistical Mathematics, 44 (3):435–449, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050697.

Gil:1992:SFR

[1728] María Angeles Gil. Sufficiency and fuzziness in random experiments. Annals of the Institute of Statistical Mathematics, 44(3):451–462, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050698.

Ebrahimi:1992:ITF

[1729] Nader Ebrahimi. Information theory and the failure time of a system. Annals of the Institute of Statistical Mathematics, 44(3):463–477, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050699.

Mase:1992:ABP

[1730] Shigeru Mase. Approximations to the birthday problem with unequal occurrence probabilities and their application to the surname problem in Japan. Annals of the Institute of Statistical Mathematics, 44(3):479–499, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050700.

Fang:1992:FDS

[1731] Z. Fang and H. Joe. Further developments on some dependence orderings for continuous bivariate distributions. Annals of the Institute of Statistical Mathematics, 44(3):501–517, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050701.

Kanda:1992:MPG

[1732] T. Kanda. MSE’s of prediction in growth curve model with covariance structures. Annals of the Institute of Statistical Mathematics, 44(3): 519–528, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050702.

Sato:1992:ICR

[1733] Tosiya Sato. On inconsistency of the common rate difference estimators from sparse follow-up data. Annals of the Institute of Statistical Mathematics, 44(3):529–535, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050703.

Tan:1992:MEL

[1734] M. Tan and L. J. Gleser. Minimax estimators for location vectors in elliptical distributions with unknown scale parameter and its application to variance reduction in simulation. Annals of the Institute of Statistical Mathematics, 44(3):537–550, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050704.

Vellaisamy:1992:AWS

[1735] P. Vellaisamy. Average worth and simultaneous estimation of the selected subset. Annals of the Institute of Statistical Mathematics, 44(3):

551–562, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050705.

Ghosh:1992:OBK

[1736] B. K. Ghosh and Wei-Min Huang. Optimum bandwidths and kernels for estimating certain discontinuous densities. Annals of the Institute of Statistical Mathematics, 44(3):563–577, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050706.

Shirahata:1992:ISE

[1737] Shingo Shirahata and In-Sun Chu.

Integrated squared error of kerneltype estimator of distribution function. Annals of the Institute of Statistical Mathematics, 44(3):579–591, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050707.

Ghosh:1992:ECB

[1738] Subir Ghosh, Sanpei Kageyama, and Rahul Mukerjee. Efficiency of connected binary block designs when a single observation is unavailable. Annals of the Institute of Statistical Mathematics, 44(3):593–603, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050708.

Anonymous:1992:HCc

[1739] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 44(3):??, September 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Papastavridis:1992:CSM

[1740] Stavros G. Papastavridis and Markos V. Koutras. Consecutive k-out-of-n systems with maintenance. Annals of the Institute of Statistical Mathematics, 44 (4):605–612, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053392.

Chen:1992:TSP

[1741] Pinyuen Chen. Truncated selection procedures for the most probable event and the least probable event. Annals of the Institute of Statistical Mathematics, 44(4):613–622, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053393.

Bar-Lev:1992:BIP

[1742] Shaul K. Bar-Lev, Idit Lavi, and Benjamin Reiser. Bayesian inference for the power law process. Annals of the Institute of Statistical Mathematics, 44(4):623–639, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053394.

Maritz:1992:APE

[1743] J. S. Maritz and T. Lwin. Assessing the performance of empirical Bayes estimators. Annals of the Institute of Statistical Mathematics, 44(4):641–657, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053395.

Draper:1992:TBV

[1744] Norman R. Draper and Irwin Guttman.

Treating bias as variance for experimental design purposes. Annals of the Institute of Statistical Mathematics, 44(4):659–671, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053396.

Shao:1992:JGL

[1745] Jun Shao. Jackknifing in generalized linear models. Annals of the Institute of Statistical Mathematics, 44(4): 673–686, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053397.

Shao:1992:OSJ

[1746] Jun Shao. One-step jackknife for M-estimators computed using Newton’s method. Annals of the Institute of Statistical Mathematics, 44(4): 687–701, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053398.

Babu:1992:SHS

[1747] Gutti Jogesh Babu. Subsample and half-sample methods. Annals of the Institute of Statistical Mathematics, 44 (4):703–720, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053399.

Jones:1992:EDQ

[1748] M. C. Jones. Estimating densities, quantiles, quantile densities and density quantiles. Annals of the Institute of Statistical Mathematics, 44(4): 721–727, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00053400.

Yu:1992:MIE

[1749] Qiqing Yu. Minimax invariant estimator of a continuous distribution function. Annals of the Institute of Statistical Mathematics, 44(4):729–735, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053401.

Gajek:1992:UBR

[1750] L. Gajek and M. Kaluszka. Upper bounds for the L_{1}-risk of the minimum L_{1}-distance regression estimator. Annals of the Institute of Statistical Mathematics, 44(4):737–744, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053402.

Akritas:1992:SAS

[1751] Michael G. Akritas and Richard A. Johnson. Symmetrized approximate score rank tests for the two-sample case. Annals of the Institute of Statistical Mathematics, 44(4):745–753, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00053403.

Allaire:1992:PAV

[1752] Jérôme Allaire and Yves Lepage. A procedure for assessing vector correlations. Annals of the Institute of Statistical Mathematics, 44(4):755–768, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053404.

Mathai:1992:BFN

[1753] A. M. Mathai. On bilinear forms in normal variables. Annals of the Institute of Statistical Mathematics, 44 (4):769–779, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00053405.

Bowman:1992:SEE

[1754] K. O. Bowman and L. R. Shenton. Some exact expressions for the mean and higher moments of functions of sample moments. Annals of the Institute of Statistical Mathematics, 44(4):781–798, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00053406.

Anonymous:1992:CSC

[1755] Anonymous. Corrections to “Speed of convergence in nonparametric kernel estimation of a regression function and its derivatives”. Annals of the Institute of Statistical Mathematics, 44(4):799, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/

                                        article/10.1007/BF00053407. See

[1290].

Anonymous:1992:HCd

[1756] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 44(4):??, December 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Anonymous:1993:Aa

[1757] Anonymous. Acknowledgment. Annals of the Institute of Statistical Mathematics, 45(1):i–iii, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00773679.

Sigman:1993:HLT

[1758] Karl Sigman and Genji Yamazaki. Heavy and light traffic in fluid models with burst arrivals. Annals of the Institute of Statistical Mathematics, 45(1):1–7, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773664.

Ebrahimi:1993:ESR

[1759] Nader Ebrahimi and T. Ramallingam. Estimation of system reliability in Brownian stress-strength models based on sample paths. Annals of the Institute of Statistical Mathematics, 45(1):9–19, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773665.

Kashiwagi:1993:UKF

[1760] Nobuhisa Kashiwagi. On use of the Kalman filter for spatial smoothing. Annals of the Institute of Statistical Mathematics, 45(1):21–34, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773666.

Speed:1993:MSP

[1761] T. P. Speed and Bin Yu. Model selection and prediction: Normal regression. Annals of the Institute of Statistical Mathematics, 45(1):35–54, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773667.

Mantel:1993:EFC

[1762] H. J. Mantel and V. P. Godambe.

Estimating functions for conditional inference: Many nuisance parameter case. Annals of the Institute of Statistical Mathematics, 45(1):55–67, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773668.

Hall:1993:EE

[1763] Peter Hall and Sally C. Morton. On the estimation of entropy. Annals of the Institute of Statistical Mathematics, 45(1):69–88, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773669.

Magdalinos:1993:AML

[1764] Michael A. Magdalinos. Approximate maximum likelihood estimation in linear regression. Annals of the Institute of Statistical Mathematics, 45(1):89–104, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773670.

Zhang:1993:EPE

[1765] Ping Zhang. On the estimation of prediction errors in linear regression models. Annals of the Institute of Statistical Mathematics, 45(1):105–111, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773671.

Ouyang:1993:GRE

[1766] Z. Ouyang, J. N. Srivastava, and H. T. Schreuder. A general ratio estimator and its application in model based inference. Annals of the Institute of Statistical Mathematics, 45(1):113–127, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773672.

Gokhale:1993:MDI

[1767] D. V. Gokhale, Koichi Inada, and Hea-Jung Kim. A minimum discrimination information estimator of preliminary conjectured normal variance. Annals of the Institute of Statistical Mathematics, 45(1):129–136, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773673.

Sengupta:1993:OTN

[1768] Ashis Sengupta and Chandranath Pal.

Optimal tests for no contamination in symmetric multivariate normal mixtures. Annals of the Institute of Statistical Mathematics, 45(1):137–146, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773674.

Fujioka:1993:ATC

[1769] Teruo Fujioka. An approximate test for common principal component subspaces in two groups. Annals of the Institute of Statistical Mathematics, 45 (1):147–158, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773675.

Silvapulle:1993:RTG

[1770] Mervyn J. Silvapulle and Pranab K.

Sen. Robust tests in group sequential analysis: One- and twosided hypotheses in the linear model. Annals of the Institute of Statistical Mathematics, 45(1):159–171, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773676.

Nishii:1993:CGC

[1771] Ryuei Nishii. Convergence of the Gram–Charier expansion after the normalizing Box–Cox transformation. Annals of the Institute of Statistical Mathematics, 45(1):173–186, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773677.

Sivaganesan:1993:RPP

[1772] S. Sivaganesan. Range of the posterior probability of an interval for priors with unimodality preserving contaminations. Annals of the Institute of Statistical Mathematics, 45(1):187–199, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773678.

Anonymous:1993:HCa

[1773] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 45(1):??, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Charalambides:1993:GMM

[1774] Ch. A. Charalambides and M. V. Koutras. On a generalization of Morisita’s model for estimating the habitat preference. Annals of the Institute of Statistical Mathematics, 45 (2):201–210, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00775807.

Stoyan:1993:EVE

[1775] D. Stoyan, U. Bertram, and H. Wendrock. Estimation variances for estimators of product densities and pair correlation functions of planar point processes. Annals of the Institute of Statistical Mathematics, 45(2):211–221, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775808.

Alzaid:1993:SAM

[1776] A. A. Alzaid and M. A. Al-Osh.

Some autoregressive moving average processes with generalized Poisson marginal distributions. Annals of the Institute of Statistical Mathematics, 45 (2):223–232, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00775809.

Swanepoel:1993:ANI

[1777] J. W. H. Swanepoel. The asymptotic normality of an intermediate order statistic of the ranges of subsamples. Annals of the Institute of Statistical Mathematics, 45(2):233–242, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775810.

Balakrishnan:1993:RBM

[1778] N. Balakrishnan, Z. Govindarajulu, and K. Balasubramanian. Relationships between moments of two related sets of order statistics and some extensions. Annals of the Institute of Statistical Mathematics, 45(2):243–247, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775811.

Gurler:1993:NEH

[1779] Ülkü Gürler and Jane-Ling Wang. Nonparametric estimation of hazard functions and their derivatives under truncation model. Annals of the Institute of Statistical Mathematics, 45(2):249–264, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00775812.

Shiraishi:1993:SPB

[1780] Taka aki Shiraishi. Statistical procedures based on signed ranks in k samples with unequal variances. Annals of the Institute of Statistical Mathematics, 45(2):265–278, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00775813.

Singh:1993:LPS

[1781] Bahadur Singh and F. T. Wright. The level probabilities for a simple loop ordering. Annals of the Institute of Statistical Mathematics, 45(2):279–292, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00775814.

Ghosh:1993:FVH

[1782] J. K. Ghosh and Rahul Mukerjee. Frequentist validity of highest posterior density regions in the multiparameter case. Annals of the Institute of Statistical Mathematics, 45 (2):293–302, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00775815. See corrections

[1808].

Consonni:1993:UBE

[1783] Guido Consonni and Piero Veronese. Unbiased Bayes estimates and improper priors. Annals of the Institute of Statistical Mathematics, 45(2):303–315, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775816.

Janas:1993:SBE

[1784] Daniel Janas. A smoothed bootstrap estimator for a Studentized sample quantile. Annals of the Institute of Statistical Mathematics, 45(2):317–329, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775817.

Pham:1993:BCP

[1785] Tuan D. Pham and Hung T. Nguyen.

Bootstrapping the change-point of a hazard rate. Annals of the Institute of Statistical Mathematics, 45(2):331–340, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775818.

Kallenberg:1993:IME

[1786] W. C. M. Kallenberg. Interpretation and manipulation of Edgeworth expansions. Annals of the Institute of Statistical Mathematics, 45(2):341–351, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00775819.

Jensen:1993:NAE

[1787] J. L. Jensen. A note on asymptotic expansions for sums over a weakly dependent random field with application to the Poisson and Strauss processes. Annals of the Institute of Statistical Mathematics, 45(2):353–360, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00775820.

Fujita:1993:GRP

[1788] Yasuhiro Fujita. A generalization of the results of Pillai. Annals of the Institute of Statistical Mathematics, 45 (2):361–365, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00775821.

Milne:1993:GMH

[1789] R. K. Milne and M. Westcott. Generalized multivariate Hermite distributions and related point processes. Annals of the Institute of Statistical Mathematics, 45(2):367–381, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00775822.

Alam:1993:RDD

[1790] Khursheed Alam and Calvin L. Williams. Relative difference in diversity between populations. Annals of the Institute of Statistical Mathematics, 45(2):383–399, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00775823.

Anonymous:1993:HCb

[1791] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 45(2):??, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Yafune:1993:BAL

[1792] Akifumi Yafune, Toshiki Matsubara, and Makio Ishiguro. Bayesian analysis of lymphatic spreading patterns in cancer of the thoracic esophagus. Annals of the Institute of Statistical Mathematics, 45(3):401–418, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773343.

Saito:1993:SSD

[1793] Yoshihiro Saito and Taketomo Mitsui. Simulation of stochastic differential equations. Annals of the Institute of Statistical Mathematics, 45(3):419–432, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773344.

Kaufmann:1993:SCM

[1794] E. Kaufmann and R.-D. Reiss. Strong convergence of multivariate point processes of exceedances. Annals of the Institute of Statistical Mathematics, 45 (3):433–444, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773345.

Luschgy:1993:SOS

[1795] Harald Luschgy. On a singularity occurring in a self-correcting point process model. Annals of the Institute of Statistical Mathematics, 45(3):445–452, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773346.

Yamato:1993:PUM

[1796] Hajime Yamato. A pólya urn model with a continuum of colors. Annals of the Institute of Statistical Mathematics, 45(3):453–458, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773347.

Sibuya:1993:RCP

[1797] Masaaki Sibuya. A random clustering process. Annals of the Institute of Statistical Mathematics, 45(3):459–465, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773348.

Wei:1993:AES

[1798] Gaoyuan Wei and B. E. Eichinger. Asymptotic expansions of some matrix argument hypergeometric functions, with applications to macromolecules. Annals of the Institute of Statistical Mathematics, 45(3):467–475, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773349.

Fu:1993:LDE

[1799] J. C. Fu, Gang Li, and D. L. C. Zhao. On large deviation expansion of distribution of maximum likelihood estimator and its application in large sample estimation. Annals of the Institute of Statistical Mathematics, 45(3):477–498, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773350.

Takagi:1993:EFA

[1800] Yoshiji Takagi and Nobuo Inagaki. Estimating function with asymptotic bias and its estimator. Annals of the Institute of Statistical Mathematics, 45 (3):499–510, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773351.

Joseph:1993:MLE

[1801] Lawrence Joseph and David B. Wolfson. Maximum likelihood estimation in the multi-path change-point problem. Annals of the Institute of Statistical Mathematics, 45(3):511–530, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773352.

Bolfarine:1993:ESL

[1802] Heleno Bolfarine and Lisbeth K. Cordani. Estimation of a structural linear regression model with a known reliability ratio. Annals of the Institute of Statistical Mathematics, 45(3):531–540, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773353.

Bose:1993:DMM

[1803] Arup Bose and Probal Chaudhuri. On the dispersion of multivariate median. Annals of the Institute of Statistical Mathematics, 45(3):541–550, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773354.

Brandwein:1993:SEL

[1804] Ann Cohen Brandwein, Stefan Ralescu, and William E. Strawderman. Shrinkage estimators of the location parameter for certain spherically symmetric distributions. Annals of the Institute of Statistical Mathematics, 45(3):551–565, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773355.

Kuboki:1993:IDN

[1805] Hisataka Kuboki. Inferential distributions for non-Bayesian predictive fit. Annals of the Institute of Statistical Mathematics, 45(3):567–578, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773356.

Huang:1993:OIT

[1806] Wen-Tao Huang and Bimal K. Sinha. On optimum invariant tests of equality of intraclass correlation coefficients. Annals of the Institute of Statistical Mathematics, 45(3):579–597, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773357.

Cohen:1993:CAE

[1807] Arthur Cohen and H. B. Sackrowitz. Corrections to “Admissibility of estimators of the probability of unobserved outcomes”. Annals of the Institute of Statistical Mathematics, 45 (3):599–601, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773358. See [1623].

Ghosh:1993:CFV

[1808] J. K. Ghosh and Rahul Mukerjee.

Corrections to “Frequentist validity of highest posterior density regions in the multiparameter case”. Annals of the Institute of Statistical Mathematics, 45(3):602, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00773359. See

[1782].

Anonymous:1993:HCc

[1809] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 45(3):??, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Anonymous:1993:Ab

[1810] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 45(4):I–II, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00774789.

Navard:1993:CDU

[1811] Sharon E. Navard, John W. Seaman, Jr., and Dean M. Young. A characterization of discrete unimodality with applications to variance upper bounds. Annals of the Institute of Statistical Mathematics, 45(4):603–614, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774775.

Wu:1993:RCI

[1812] Y. Wu. On rates of convergence of information theoretic criterion in rank determination of one-way random effects models. Annals of the Institute of Statistical Mathematics, 45(4):615–620, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00774776.

Chen:1993:AEL

[1813] Song Xi Chen. On the accuracy of empirical likelihood confidence regions for linear regression model. Annals of the Institute of Statistical Mathematics, 45(4):621–637, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00774777.

Beran:1993:SRC

[1814] Rudolf Beran. Semiparametric random coefficient regression models. Annals of the Institute of Statistical Mathematics, 45(4):639–654, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00774778.

Park:1993:MRS

[1815] Sung H. Park, Jun H. Lim, and Yasumasa Baba. A measure of rotatability for second order response surface designs. Annals of the Institute of Statistical Mathematics, 45(4):655–664, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00774779.

Tran:1993:OSN

[1816] Lanh Tat Tran and Berlin Wu. Order statistics for nonstationary time series. Annals of the Institute of Statistical Mathematics, 45(4):665–686, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774780.

Puri:1993:ABS

[1817] Madan L. Puri and Frits H. Ruymgaart. Asymptotic behavior of Lstatistics for a large class of time series. Annals of the Institute of Statistical Mathematics, 45(4):687–701, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774781.

Pfanzagl:1993:CCM

[1818] J. Pfanzagl. On the consistency of conditional maximum likelihood estimators. Annals of the Institute of Statistical Mathematics, 45(4):703–719, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774782.

Thavaneswaran:1993:NSE

[1819] A. Thavaneswaran and Jagbir Singh.

A note on smoothed estimating functions. Annals of the Institute of Statistical Mathematics, 45(4):721–729, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774783.

Kuriki:1993:OIE

[1820] Satoshi Kuriki. Orthogonally invariant estimation of the skew-symmetric normal mean matrix. Annals of the Institute of Statistical Mathematics, 45 (4):731–739, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00774784.

Nakamura:1993:EMC

[1821] Tadashi Nakamura and Chae-Shin Lee. On the existence of minimum contrast estimates in binary response model. Annals of the Institute of Statistical Mathematics, 45(4):741–758, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774785.

Mukerjee:1993:ECL

[1822] Rahul Mukerjee. An extension of the conditional likelihood ratio test to the general multiparameter case. Annals of the Institute of Statistical Mathematics, 45(4):759–771, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774786.

Tripathi:1993:STI

[1823] Ram C. Tripathi, Ramesh C. Gupta, and Robert K. Pair. Statistical tests involving several independent gamma distributions. Annals of the Institute of Statistical Mathematics, 45(4):773–786, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00774787.

Aki:1993:NTS

[1824] Sigeo Aki. On nonparametric tests for symmetry in R^{m}. Annals of the Institute of Statistical Mathematics, 45 (4):787–800, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00774788.

Anonymous:1993:HCd

[1825] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 45(4):??, ???? 1993. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kunsch:1994:RPS

[1826] Hans R. Künsch. Robust priors for smoothing and image restoration. Annals of the Institute of Statistical Mathematics, 46(1):1–19, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773588.

Nicolaou:1994:CPB

[1827] Anna Nicolaou. Conditional properties of Bayesian interval estimates. Annals of the Institute of Statistical Mathematics, 46(1):21–28, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773589.

Yanagimoto:1994:KLR

[1828] Takemi Yanagimoto. The Kullback– Leibler risk of the Stein estimator and the conditional MLE. Annals of the Institute of Statistical Mathematics, 46(1):29–41, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773590.

Blaesild:1994:MLE

[1829] Preben Blæsild. Maximum likelihood estimation in exponential orthogeodesic models. Annals of the Institute of Statistical Mathematics, 46(1):43–55, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773591.

delCastillo:1994:STN

[1830] Joan del Castillo. The singly truncated normal distribution: A nonsteep exponential family. Annals of the Institute of Statistical Mathematics, 46(1):57–66, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773592.

Grubel:1994:EDF

[1831] Rudolf Grübel. Estimation of density functionals. Annals of the Institute of Statistical Mathematics, 46(1):67–75, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773593.

Isogal:1994:SEP

[1832] Eiichi Isogal and Chikara Uno. Sequential estimation of a parameter of an exponential distribution. Annals of the Institute of Statistical Mathematics, 46(1):77–82, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773594.

Clarke:1994:REC

[1833] B. R. Clarke and C. R. Heathcote. Robust estimation of k-component univariate normal mixtures. Annals of the Institute of Statistical Mathematics, 46(1):83–93, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773595.

Kubokawa:1994:DSE

[1834] Tatsuya Kubokawa. Double shrinkage estimation of ratio of scale parameters. Annals of the Institute of Statistical Mathematics, 46(1):95–116, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773596.

Qin:1994:SEL

[1835] Jing Qin. Semi-empirical likelihood ratio confidence intervals for the difference of two sample means. Annals of the Institute of Statistical Mathematics, 46(1):117–126, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773597.

Bilodeau:1994:LTI

[1836] Martin Bilodeau and Takeaki Kariya.

LBI tests of independence in bivariate exponential distributions. Annals of the Institute of Statistical Mathematics, 46(1):127–136, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773598.

Hu:1994:LRT

[1837] Xiaomi Hu and F. T. Wright. Likelihood ratio tests for a class of nonoblique hypotheses. Annals of the Institute of Statistical Mathematics, 46 (1):137–145, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773599.

Loh:1994:DEE

[1838] Wei-Liem Loh. On m-dependence and Edgeworth expansions. Annals of the Institute of Statistical Mathematics, 46 (1):147–164, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773600.

Hwang:1994:JDS

[1839] Tea-Yuan Hwang and Chin-Yuan Hu. On the joint distribution of Studentized order statistics. Annals of the Institute of Statistical Mathematics, 46 (1):165–177, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773601.

Fu:1994:PAD

[1840] James C. Fu and Markos V. Koutras. Poisson approximations for 2-dimensional patterns. Annals of the Institute of Statistical Mathematics, 46(1):179–192, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773602.

Aki:1994:DNF

[1841] Sigeo Aki and Katuomi Hirano. Distributions of numbers of failures and successes until the first consecutive k successes. Annals of the Institute of Statistical Mathematics, 46(1):193–202, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773603.

Anonymous:1994:HCa

[1842] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 46(1):??, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Leonard:1994:BLI

[1843] Tom Leonard, John S. J. Hsu, Kam-Wah Tsui, and James F. Murray. Bayesian and likelihood inference from equally weighted mixtures. Annals of the Institute of Statistical Mathematics, 46(2):203–220, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01720581.

Sun:1994:BSR

[1844] Dongchu Sun and James O. Berger.

Bayesian sequential reliability for Weibull and related distributions. Annals of the Institute of Statistical Mathematics, 46(2):221–249, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01720582.

Chen:1994:TNE

[1845] Juei-Chao Chen. Testing for no effect in nonparametric regression via spline smoothing techniques. Annals of the Institute of Statistical Mathematics, 46(2):251–265, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01720583.

Wei:1994:SMR

[1846] Bo-Cheng Wei and Jian-Qing Shih. On statistical models for regression diagnostics. Annals of the Institute of Statistical Mathematics, 46(2):267–278, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01720584.

Truong:1994:NTS

[1847] Young K. Truong. Nonparametric time series regression. Annals of the Institute of Statistical Mathematics, 46 (2):279–293, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01720585.

Ferguson:1994:MEV

[1848] Thomas S. Ferguson and Lynn Kuo. Minimax estimation of a variance. Annals of the Institute of Statistical Mathematics, 46(2):295–308, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01720586.

Wefelmeyer:1994:EEE

[1849] W. Wefelmeyer. An efficient estimator for the expectation of a bounded function under the residual distribution of an autoregressive process. Annals of the Institute of Statistical Mathematics, 46(2):309–315, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01720587.

Tripathi:1994:EPB

[1850] Ram C. Tripathi, Ramesh C. Gupta, and John Gurland. Estimation of parameters in the beta binomial model. Annals of the Institute of Statistical Mathematics, 46(2):317–331, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01720588.

Bassan:1994:PDO

[1851] Bruno Bassan and Marco Scarsini. Positive dependence orderings and stopping times. Annals of the Institute of Statistical Mathematics, 46(2):333–342, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01720589.

Finner:1994:LBJ

[1852] H. Finner and M. Roters. On the limit behaviour of the joint distribution function of order statistics. Annals of the Institute of Statistical Mathematics, 46(2):343–349, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01720590.

Li:1994:CPP

[1853] Shun-Hwa Li, Wen-Jang Huang, and Mong-Na Lo Huang. Characterizations of the Poisson process as a renewal process via two conditional moments. Annals of the Institute of Statistical Mathematics, 46(2):351–360, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01720591.

Royen:1994:SMG

[1854] T. Royen. On some multivariate gamma-distributions connected with spanning trees. Annals of the Institute of Statistical Mathematics, 46 (2):361–371, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01720592.

Bohning:1994:DLR

[1855] Dankmar Böhning, Ekkehart Dietz, Rainer Schaub, Peter Schlattmann, and Bruce G. Lindsay. The distribution of the likelihood ratio for mixtures of densities from the one-parameter exponential family. Annals of the Institute of Statistical Mathematics, 46 (2):373–388, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01720593.

Dette:1994:ODR

[1856] Holger Dette and William J. Studden. Optimal designs with respect to Elfving’s partial minimax criterion in polynomial regression. Annals of the Institute of Statistical Mathematics, 46 (2):389–403, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01720594.

Anonymous:1994:HCb

[1857] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 46(2):??, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Higuchi:1994:SSS

[1858] T. Higuchi, G. K. Crawford, R. J.

Strangeway, and C. T. Russell. Separation of spin synchronized signals. Annals of the Institute of Statistical Mathematics, 46(3):405–428, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773508.

Liu:1994:MBF

[1859] Shu-Ing Liu. Multiperiod Bayesian forecasts for AR models. Annals of the Institute of Statistical Mathematics, 46(3):429–452, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773509.

vonSachs:1994:ENL

[1860] Rainer von Sachs. Estimating nonlinear functions of the spectral density, using a data-taper. Annals of the Institute of Statistical Mathematics, 46(3):453–474, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773510.

Jensen:1994:ANP

[1861] Jens Ledet Jensen and Hans R. Künsch. On asymptotic normality of pseudo likelihood estimates for pairwise interaction processes. Annals of the Institute of Statistical Mathematics, 46(3):475–486, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773511.

Wells:1994:BBE

[1862] Martin T. Wells and Ram C. Tiwari. Bootstrapping a Bayes estimator of a survival function with censored data. Annals of the Institute of Statistical Mathematics, 46(3):487–495, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773512.

Lu:1994:ABS

[1863] Wang-Shu Lu. Approximate Bayesian shrinkage estimation. Annals of the Institute of Statistical Mathematics, 46 (3):497–507, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773513.

Mukhopadhyay:1994:ISE

[1864] N. Mukhopadhyay. Improved sequential estimation of means of exponential distributions. Annals of the Institute of Statistical Mathematics, 46(3):509–519, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773514.

Jones:1994:VLS

[1865] M. C. Jones, I. J. McKay, and T.- C. Hu. Variable location and scale kernel density estimation. Annals of the Institute of Statistical Mathematics, 46(3):521–535, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773515.

Pitts:1994:NEC

[1866] S. M. Pitts. Nonparametric estimation of compound distributions with applications in insurance. Annals of the Institute of Statistical Mathematics, 46(3):537–555, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773516.

Provost:1994:EDF

[1867] Serge B. Provost and Edmund M.

Rudiuk. The exact density function of the ratio of two dependent linear combinations of chi-square variables. Annals of the Institute of Statistical Mathematics, 46(3):557–571, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773517.

Boulerice:1994:DDD

[1868] Bernard Boulerice and Gilles R.

Ducharme. Decentered directional data. Annals of the Institute of Statistical Mathematics, 46(3):573–586, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773518.

Lin:1994:EWC

[1869] Gwo Dong Lin. On equivalence of weak convergence and moment convergence of life distributions. Annals of the Institute of Statistical Mathematics, 46(3):587–592, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773519.

Bhapkar:1994:FII

[1870] Vasant P. Bhapkar and Cidambi Srinivasan. On Fisher information inequalities in the presence of nuisance parameters. Annals of the Institute of Statistical Mathematics, 46(3):593–604, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773520.

Anonymous:1994:HCc

[1871] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 46(3):??, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kitagawa:1994:TFF

[1872] Genshiro Kitagawa. The two-filter formula for smoothing and an implementation of the Gaussian-sum smoother. Annals of the Institute of Statistical Mathematics, 46(4):605–623, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773470.

Pourahmadi:1994:CCR

[1873] Mohsen Pourahmadi. Canonical correlation and reduction of multiple time series. Annals of the Institute of Statistical Mathematics, 46(4):625–631, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773471.

Inagaki:1994:ADM

[1874] Nobuo Inagaki. Asymptotic distribution of maximal autoregressive process with weight tending to 1. Annals of the Institute of Statistical Mathematics, 46(4):633–640, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773472.

Yannaros:1994:WRP

[1875] Nikos Yannaros. Weibull renewal processes. Annals of the Institute of Statistical Mathematics, 46(4):641–648, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773473.

Pace:1994:GSE

[1876] Luigi Pace and Alessandra Salvan.

The geometric structure of the expected/observed likelihood expansions. Annals of the Institute of Statistical Mathematics, 46(4):649–666, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773474.

Janas:1994:EES

[1877] Daniel Janas. Edgeworth expansions for spectral mean estimates with applications to Whittle estimates. Annals of the Institute of Statistical Mathematics, 46(4):667–682, ????

1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773475.

Basu:1994:MDE

[1878] Ayanendranath Basu and Bruce G. Lindsay. Minimum disparity estimation for continuous models: Efficiency, distributions and robustness. Annals of the Institute of Statistical Mathematics, 46(4):683–705, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773476.

Su:1994:SDR

[1879] Yingcai Su and Stamatis Cambanis. Sampling designs for regression coefficient estimation with correlated errors. Annals of the Institute of Statistical Mathematics, 46(4):707–722, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773477.

Lam:1994:EPT

[1880] Kin Lam, Bimal K. Sinha, and Zhong Wu. Estimation of parameters in a two-parameter exponential distribution using ranked set sample. Annals of the Institute of Statistical Mathematics, 46(4):723–736, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773478.

Jureckova:1994:ACT

[1881] Jana Jurecková, Roger Koenker, and A. H. Welsh. Adaptive choice of trimming proportions. Annals of the Institute of Statistical Mathematics, 46 (4):737–755, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773479.

Nagaraja:1994:TLS

[1882] H. N. Nagaraja. Tukey’s linear sensitivity and order statistics. Annals of the Institute of Statistical Mathematics, 46(4):757–768, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773480.

Hwang:1994:JDG

[1883] Tea-Yuan Hwang and Chin-Yuan Hu.

On the joint distribution of Grubbs’ statistics. Annals of the Institute of Statistical Mathematics, 46(4):769–775, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773481.

Mohanty:1994:SRL

[1884] S. G. Mohanty. Success runs of length k in Markov dependent trials. Annals of the Institute of Statistical Mathematics, 46(4):777–796, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773482.

Pakes:1994:NCC

[1885] Anthony G. Pakes. Necessary conditions for characterization of laws via mixed sums. Annals of the Institute of Statistical Mathematics, 46(4):797–802, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773483.

Anonymous:1994:A

[1886] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 46(4):803–804, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00773484.

Anonymous:1994:HCd

[1887] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 46(4):??, ???? 1994. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Brockwell:1995:ACT

[1888] P. J. Brockwell and O. Stramer. On the approximation of continuous time threshold ARMA processes. Annals of the Institute of Statistical Mathematics, 47(1):1–20, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773408.

Nagaev:1995:SPC

[1889] A. V. Nagaev. Some properties of convex hulls generated by homogeneous Poisson point processes in an unbounded convex domain. Annals of the Institute of Statistical Mathematics, 47(1):21–29, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773409.

Yamazaki:1995:OOT

[1890] Genji Yamazaki and Hiroshi Ito. Optimal order for two servers in tandem. Annals of the Institute of Statistical Mathematics, 47(1):31–48, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773410.

Vos:1995:QLE

[1891] Paul W. Vos. Quasi-likelihood or extended quasi-likelihood? an information-geometric approach. Annals of the Institute of Statistical Mathematics, 47(1):49–64, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773411.

Wang:1995:OSB

[1892] Suojin Wang. Optimizing the smoothed bootstrap. Annals of the Institute of Statistical Mathematics, 47(1):65–80, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773412.

Wei:1995:CRE

[1893] Laisheng Wei and Shunpu Zhang. The convergence rates of empirical Bayes estimation in a multiple linear regression model. Annals of the Institute of Statistical Mathematics, 47(1):81–97, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773413.

Bose:1995:NAS

[1894] Arup Bose and Nitis Mukhopadhyay. A note on accelerated sequential estimation of the mean of NEF–PVF distributions. Annals of the Institute of Statistical Mathematics, 47(1):99–104, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773414.

Xiang:1995:EQS

[1895] Xiaojing Xiang. Estimation of a quantile in some nonstandard cases. Annals of the Institute of Statistical Mathematics, 47(1):105–117, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773415.

Takada:1995:API

[1896] Yoshikazu Takada. Admissibility of prediction intervals. Annals of the Institute of Statistical Mathematics, 47 (1):119–128, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773416.

vonRosen:1995:RGC

[1897] Dietrich von Rosen. Residuals in the growth curve model. Annals of the Institute of Statistical Mathematics, 47 (1):129–136, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773417.

Pan:1995:MOD

[1898] Jian-Xin Pan and Kai-Tai Fang. Multiple outlier detection in growth curve model with unstructured covariance matrix. Annals of the Institute of Statistical Mathematics, 47(1):137–153, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773418.

Dumbgen:1995:MTC

[1899] Lutz Dümbgen. Minimax tests for convex cones. Annals of the Institute of Statistical Mathematics, 47(1):155–165, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773419.

Amrhein:1995:ETS

[1900] Peter Amrhein. An example of a twosided Wilcoxon signed rank test which is not unbiased. Annals of the Institute of Statistical Mathematics, 47 (1):167–170, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773420.

Papathanasiou:1995:CPS

[1901] V. Papathanasiou. A characterization of the Pearson system of distributions and the associated orthogonal polynomials. Annals of the Institute of Statistical Mathematics, 47(1):171–176, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773421.

Wesolowski:1995:BDP

[1902] Jacek Wesolowski. Bivariate distributions via a Pareto conditional distribution and a regression function. Annals of the Institute of Statistical Mathematics, 47(1):177–183, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773422.

Papadatos:1995:MVO

[1903] Nickos Papadatos. Maximum variance of order statistics. Annals of the Institute of Statistical Mathematics, 47 (1):185–193, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773423.

Anonymous:1995:HCa

[1904] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 47(1):??, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nishiyama:1995:LAN

[1905] Yoichi Nishiyama. Local asymptotic normality of a sequential model for marked point processes and its applications. Annals of the Institute of Statistical Mathematics, 47(2):195–209, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773457.

Liu:1995:BMF

[1906] Shu-Ing Liu. Bayesian multiperiod forecasts for ARX models. Annals of the Institute of Statistical Mathematics, 47(2):211–224, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773458.

Aki:1995:JDN

[1907] Sigeo Aki and Katuomi Hirano. Joint distributions of numbers of successruns and failures until the first consecutive k successes. Annals of the Institute of Statistical Mathematics, 47 (2):225–235, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773459.

Xiang:1995:BET

[1908] Xiaojing Xiang. A Berry–Esseen theorem for the kernel quantile estimator with application to studying the deficiency of quantile estimators. Annals of the Institute of Statistical Mathematics, 47(2):237–251, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773460.

Bagai:1995:KTD

[1909] Isha Bagai and B. L. S. Prakasa Rao. Kernel-type density and failure rate estimation for associated sequences. Annals of the Institute of Statistical Mathematics, 47(2):253–266, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773461.

Bose:1995:EAD

[1910] Arup Bose. Estimating the asymptotic dispersion of the L_{1} median. Annals of the Institute of Statistical Mathematics, 47(2):267–271, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773462.

Shinozaki:1995:SMI

[1911] Nobuo Shinozaki. Some modifications of improved estimators of a normal variance. Annals of the Institute of Statistical Mathematics, 47(2):273–286, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773463.

Vijayasree:1995:CEO

[1912] G. Vijayasree, Neeraj Misra, and Harshinder Singh. Componentwise estimation of ordered parameters of k(≥ 2) exponential populations. Annals of the Institute of Statistical Mathematics, 47(2):287–307, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773464.

Yokoyama:1995:LTR

[1913] Takahisa Yokoyama. LR test for random-effects covariance structure in a parallel profile model. Annals of the Institute of Statistical Mathematics, 47(2):309–320, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773465.

Kowalski:1995:CCT

[1914] Jacek P. Kowalski. Complete classes of tests for regularly varying distributions. Annals of the Institute of Statistical Mathematics, 47(2):321–350, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773466.

Chang:1995:SRM

[1915] Der-Shin Chang and Guan-Chyun Lin.

Stochastic regression model with heteroscedastic disturbance. Annals of the Institute of Statistical Mathematics, 47(2):351–369, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773467.

Abdous:1995:RQE

[1916] Belkacem Abdous and Bruno Remillard. Relating quantiles and expectiles under weighted-symmetry. Annals of the Institute of Statistical Mathematics, 47(2):371–384, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773468.

Bischoff:1995:DFA

[1917] Wolfgang Bischoff. Determinant formulas with applications to designing when the observations are correlated. Annals of the Institute of Statistical Mathematics, 47(2):385–399, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773469.

Anonymous:1995:HCb

[1918] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 47(2):??, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Cain:1995:REP

[1919] Michael Cain and Christian Janssen. Real estate price prediction under asymmetric loss. Annals of the Institute of Statistical Mathematics, 47 (3):401–414, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773391.

Uchida:1995:SLW

[1920] Masayuki Uchida and Sigeo Aki. Sooner and later waiting time problems in a two-state Markov chain. Annals of the Institute of Statistical Mathematics, 47(3):415–433, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773392.

Fu:1995:ELD

[1921] James C. Fu. Exact and limiting distributions of the number of successions in a random permutation. Annals of the Institute of Statistical Mathematics, 47(3):435–446, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773393.

Sapatinas:1995:IMP

[1922] Theofanis Sapatinas. Identifiability of mixtures of power-series distributions and related characterizations. Annals of the Institute of Statistical Mathematics, 47(3):447–459, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773394.

Barndorff-Nielsen:1995:QPD

[1923] O. E. Barndorff-Nielsen. Quasi profile and directed likelihoods from estimating functions. Annals of the Institute of Statistical Mathematics, 47(3):461–464, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773395.

Stokes:1995:PRS

[1924] Lynne Stokes. Parametric ranked set sampling. Annals of the Institute of Statistical Mathematics, 47(3):465–482, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773396.

Mitra:1995:CPE

[1925] Murari Mitra and Sujit K. Basu.

Change point estimation in nonmonotonic aging models. Annals of the Institute of Statistical Mathematics, 47(3):483–491, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00773397.

Gupta:1995:BEM

[1926] Ramesh C. Gupta and H. Olcay Akman. Bayes estimation in a mixture inverse Gaussian model. Annals of the Institute of Statistical Mathematics, 47(3):493–503, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773398.

Moonesinghe:1995:THO

[1927] Ramal Moonesinghe and F. T. Wright. Testing homogeneity with an ordered alternative in a two-factor layout by combining p-values. Annals of the Institute of Statistical Mathematics, 47 (3):505–523, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773399.

Ren:1995:GCM

[1928] Jian-Jian Ren. Generalized Cramérvon Mises tests of goodness of fit for doubly censored data. Annals of the Institute of Statistical Mathematics, 47 (3):525–549, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00773400.

Garel:1995:LAN

[1929] Bernard Garel and Marc Hallin. Local asymptotic normality of multivariate ARMA processes with a linear trend. Annals of the Institute of Statistical Mathematics, 47(3):551–579, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00773401.

Taniguchi:1995:HOA

[1930] Masanobu Taniguchi and Madan L. Puri. Higher order asymptotic theory for normalizing transformations of maximum likelihood estimators. Annals of the Institute of Statistical Mathematics, 47(3):581–600, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00773402.

Anonymous:1995:HCc

[1931] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 47(3):??, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Baddeley:1995:AIP

[1932] A. J. Baddeley and M. N. M. van Lieshout. Area-interaction point processes. Annals of the Institute of Statistical Mathematics, 47(4):601–619, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01856536.

Patil:1995:FPC

[1933] G. P. Patil, A. K. Sinha, and C. Taillie. Finite population corrections for ranked set sampling. Annals of the Institute of Statistical Mathematics, 47 (4):621–636, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01856537.

He:1995:EDT

[1934] Kun He. On estimating domain totals over a subpopulation. Annals of the Institute of Statistical Mathematics, 47(4):637–644, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01856538.

Hesse:1995:DDC

[1935] Christian H. Hesse. Deconvolving a density from contaminated dependent observations. Annals of the Institute of Statistical Mathematics, 47(4):645–663, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01856539.

Shiraishi:1995:CIE

[1936] T. Shiraishi and Y. Konno. On construction of improved estimators in multiple-design multivariate linear models under general restriction. Annals of the Institute of Statistical Mathematics, 47(4):665–674, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01856540.

Albers:1995:TSR

[1937] Willem Albers. A two-stage rank test using density estimation. Annals of the Institute of Statistical Mathematics, 47(4):675–691, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01856541.

Falk:1995:LEO

[1938] Michael Falk. LAN of extreme order statistics. Annals of the Institute of Statistical Mathematics, 47(4):693–717, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01856542.

Dykstra:1995:LRT

[1939] Richard Dykstra, Subhash Kochar, and Tim Robertson. Likelihood ratio tests for symmetry against one-sided alternatives. Annals of the Institute of Statistical Mathematics, 47(4):719–730, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01856543.

Ananda:1995:GTU

[1940] Malwane M. A. Ananda. Generalized F -tests for unbalanced nested designs under heteroscedasticity. Annals of the Institute of Statistical Mathematics, 47(4):731–742, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF01856544.

Koutras:1995:RSU

[1941] M. V. Koutras and V. A. Alexandrou.

Runs, scans and URN model distributions: A unified Markov chain approach. Annals of the Institute of Statistical Mathematics, 47(4):743–766, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01856545.

Kollo:1995:AWD

[1942] Tönu Kollo and Dietrich von Rosen.

Approximating by the Wishart distribution. Annals of the Institute of Statistical Mathematics, 47(4):767–783, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01856546.

Letac:1995:CST

[1943] G. Letac and H. Massam. Craig– Sakamoto’s theorem for the Wishart distributions on symmetric cones. Annals of the Institute of Statistical Mathematics, 47(4):785–799, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01856547.

Anonymous:1995:HCd

[1944] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 47(4):??, ???? 1995. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Beran:1996:CSC

[1945] Rudolf Beran. Confidence sets centered at C_{p}-estimators. Annals of the Institute of Statistical Mathematics, 48(1):1–15, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049285.

duPlessis:1996:BCE

[1946] J. L. du Plessis and A. J. van der Merwe. Bayesian calibration in the estimation of the age of rhinoceros. Annals of the Institute of Statistical Mathematics, 48(1):17–28, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049286.

Thomson:1996:ESO

[1947] Peter J. Thomson and Peter M. Robinson. Estimation of second-order properties from jittered time series.

Annals of the Institute of Statistical Mathematics, 48(1):29–48, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049287.

Boshnakov:1996:BFC

[1948] Georgi N. Boshnakov. Bartlett’s formulae — closed forms and recurrent equations. Annals of the Institute of Statistical Mathematics, 48(1):49–59, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049288.

Kuchler:1996:CEF

[1949] Uwe Küchler and Michael Sørensen.

Curved exponential families of stochastic processes and their envelope families. Annals of the Institute of Statistical Mathematics, 48(1):61–74, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049289.

Babu:1996:ABI

[1950] Gutti Jogesh Babu and Yogendra P.

Chaubey. Asymptotics and bootstrap for inverse Gaussian regression. Annals of the Institute of Statistical Mathematics, 48(1):75–88, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049290.

Jimichi:1996:CER

[1951] Masayuki Jimichi and Nobuo Inagaki. r-k Class estimation in regression model with concomitant variables. Annals of the Institute of Statistical Mathematics, 48(1):89–95, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049291.

Wu:1996:FWS

[1952] Jong-Wuu Wu, Jiahn-Bang Jang, and Tzong-Ru Tsai. Fuzzy weighted scaled coefficients in semi-parametric model. Annals of the Institute of Statistical Mathematics, 48(1):97–110, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049292.

Arellano-Valle:1996:NSS

[1953] R. B. Arellano-Valle and H. Bolfarine. A note on the simple structural regression model. Annals of the Institute of Statistical Mathematics, 48(1):111–125, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00049293.

Takahashi:1996:MSP

[1954] Rinya Takahashi and Masaaki Sibuya. The maximum size of the planar sections of random spheres and its application to metallurgy. Annals of the Institute of Statistical Mathematics, 48 (1):127–144, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00049294.

Rauhut:1996:IPD

[1955] B. Rauhut. Iterated probability distributions and extremes with random sample size. Annals of the Institute of Statistical Mathematics, 48(1):145–155, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049295.

Dimaki:1996:TUC

[1956] Caterina Dimaki and Evdokia Xekalaki.

Towards a unification of certain characterizations by conditional expectations. Annals of the Institute of Statistical Mathematics, 48(1):157–168, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049296.

Shenton:1996:MSU

[1957] L. R. Shenton and K. O. Bowman.

Moment solution to an urn model. Annals of the Institute of Statistical Mathematics, 48(1):169–184, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049297.

Aki:1996:LDE

[1958] Sigeo Aki and Katuomi Hirano. Lifetime distribution and estimation problems of consecutive-k-out-of-n:F systems. Annals of the Institute of Statistical Mathematics, 48(1):185–199, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00049298.

Anonymous:1996:HCa

[1959] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 48(1):??, March 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mase:1996:TME

[1960] Shigeru Mase. The threshold method for estimating total rainfall. Annals of the Institute of Statistical Mathematics, 48(2):201–213, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00054785.

Kashiwagi:1996:SSA

[1961] Nobuhisa Kashiwagi. A state-space approach to polygonal line regression. Annals of the Institute of Statistical Mathematics, 48(2):215–228, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054786.

Karunamuni:1996:EBD

[1962] Rohana J. Karunamuni and Shunpu Zhang. Empirical Bayes detection of a change in distribution. Annals of the Institute of Statistical Mathematics, 48(2):229–246, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00054787.

Rojo:1996:RBP

[1963] Javier Rojo. Relationships between pure tail orderings of lifetime distributions and some concepts of residual life. Annals of the Institute of Statistical Mathematics, 48(2):247–255, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054788.

Ebrahimi:1996:MDB

[1964] Nader Ebrahimi and S. N. U. A.

Kirmani. A measure of discrimination between two residual lifetime distributions and its applications. Annals of the Institute of Statistical Mathematics, 48(2):257–265, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00054789.

Roussas:1996:MDR

[1965] George G. Roussas and Yannis G. Yatracos. Minimum distance regressiontype estimates with rates under weak dependence. Annals of the Institute of Statistical Mathematics, 48(2):267–281, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00054790.

Wu:1996:QLE

[1966] Jong-Wuu Wu. The quasi-likelihood estimation in regression. Annals of the Institute of Statistical Mathematics, 48(2):283–294, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00054791.

He:1996:BER

[1967] Xuming He and Qi man Shao. Bahadur efficiency and robustness of Studentized score tests. Annals of the Institute of Statistical Mathematics, 48(2):295–314, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00054792.

Heckman:1996:NTB

[1968] Nancy E. Heckman and Bing Li. Nonparametric tests for bounds on the derivative of a regression function. Annals of the Institute of Statistical Mathematics, 48(2):315–336, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054793.

Fueda:1996:LNT

[1969] Kaoru Fueda. The limiting normality of the test statistic for the twosample problem induced by a convex sum distance. Annals of the Institute of Statistical Mathematics, 48(2):337–347, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054794.

Akahira:1996:LIS

[1970] Masafumi Akahira. Loss of information of a statistic for a family of non-regular distributions. Annals of the Institute of Statistical Mathematics, 48(2):349–364, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054795.

Akahira:1996:TOE

[1971] Masafumi Akahira. Third order efficiency implies fourth order efficiency: A resolution of the conjecture of J. K. Ghosh. Annals of the Institute of Statistical Mathematics, 48(2):365–380, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00054796.

Provost:1996:EDI

[1972] Serge B. Provost and Edmund M.

Rudiuk. The exact distribution of indefinite quadratic forms in noncentral normal vectors. Annals of the Institute of Statistical Mathematics, 48(2):381–394, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00054797.

Draper:1996:OTO

[1973] Norman R. Draper, Berthold Heiligers, and Friedrich Pukelsheim. On optimal third order rotatable designs. Annals of the Institute of Statistical Mathematics, 48(2):395–402, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00054798.

Anonymous:1996:HCb

[1974] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 48(2):??, June 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hall:1996:BED

[1975] Peter Hall and Raoul LePage. On bootstrap estimation of the distribution of the Studentized mean. Annals of the Institute of Statistical Mathematics, 48(3):403–421, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050845.

Fujisawa:1996:MLE

[1976] Hironori Fujisawa. The maximum likelihood estimators in a multivariate normal distribution with AR(1) covariance structure for monotone data. Annals of the Institute of Statistical Mathematics, 48(3):423–428, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050846.

Hallin:1996:KDE

[1977] Marc Hallin and Lanh Tat Tran. Kernel density estimation for linear processes: Asymptotic normality and optimal bandwidth derivation. Annals of the Institute of Statistical Mathematics, 48(3):429–449, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050847.

Wu:1996:MKD

[1978] Colin O. Wu and Andrew Q. Mao. Minimax kernels for density estimation with biased data. Annals of the Institute of Statistical Mathematics, 48 (3):451–467, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050848.

Visek:1996:SAE

[1979] Jan Ámos Vísek. Sensitivity analysis of M-estimates. Annals of the Institute of Statistical Mathematics, 48(3): 469–495, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050849.

Mukhopadhyay:1996:SFW

[1980] Nitis Mukhopadhyay and Sujay Datta.

On sequential fixed-width confidence intervals for the mean and secondorder expansions of the associated coverage probabilities. Annals of the Institute of Statistical Mathematics, 48 (3):497–507, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050850.

Kourouklis:1996:IEU

[1981] Stavros Kourouklis. Improved estimation under Pitman’s measure of closeness. Annals of the Institute of Statistical Mathematics, 48(3): 509–518, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050851.

Balakrishnan:1996:RMO

[1982] N. Balakrishnan and Rita Aggarwala.

Relationships for moments of order statistics from the right-truncated generalized half logistic distribution. Annals of the Institute of Statistical Mathematics, 48(3):519–534, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050852.

Cramer:1996:SOS

[1983] Erhard Cramer and Udo Kamps. Sequential order statistics and k-out-of-n systems with sequentially adjusted failure rates. Annals of the Institute of Statistical Mathematics, 48(3): 535–549, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050853.

Antzoulakos:1996:DPD

[1984] Demetrios L. Antzoulakos and Andreas N. Philippou. Derivation of the probability distribution functions for succession quota random variables. Annals of the Institute of Statistical Mathematics, 48(3):551–561, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050854.

Ruiz:1996:CBC

[1985] J. M. Ruiz and J. Navarro. Characterizations based on conditional expectations of the doubled truncated distribution. Annals of the Institute of Statistical Mathematics, 48(3): 563–572, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050855.

Nguyen:1996:CCD

[1986] T. T. Nguyen, A. K. Gupta, and Y. Wang. A characterization of certain discrete exponential families. Annals of the Institute of Statistical Mathematics, 48(3):573–576, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00050856.

Bhansali:1996:AEA

[1987] R. J. Bhansali. Asymptotically efficient autoregressive model selection for multistep prediction. Annals of the Institute of Statistical Mathematics, 48 (3):577–602, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00050857.

Anonymous:1996:HCc

[1988] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 48(3):??, September 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takemura:1996:PIB

[1989] Akimichi Takemura and Satoshi Kuriki. A proof of independent Bartlett correctability of nested likelihood ratio tests. Annals of the Institute of Statistical Mathematics, 48(4):603–620, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052322.

Ku:1996:QTT

[1990] Simon Ku and Eugene Seneta.

Quenouille-type theorem on autocorrelations. Annals of the Institute of Statistical Mathematics, 48(4): 621–630, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052323.

Biscay:1996:LLM

[1991] R. Biscay, J. C. Jimenez, J. J. Riera, and P. A. Valdes. Local linearization method for the numerical solution of stochastic differential equations. Annals of the Institute of Statistical Mathematics, 48(4):631–644, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052324.

Bhattachariee:1996:SCB

[1992] M. C. Bhattachariee. Stochastic comparisons and bounds for aging renewal process shock models and their applications. Annals of the Institute of Statistical Mathematics, 48(4):645–662, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052325.

Muliere:1996:BNP

[1993] P. Muliere and P. Secchi. Bayesian nonparametric predictive inference and bootstrap techniques. Annals of the Institute of Statistical Mathematics, 48 (4):663–673, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052326.

Sturgeon:1996:MSR

[1994] Michael Sturgeon. Mass shifting roles of negative kernel mass in density estimation. Annals of the Institute of Statistical Mathematics, 48(4): 675–686, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052327.

Quiroz:1996:EMB

[1995] Adolfo J. Quiroz, Miguel Nakamura, and Francisco J. Pérez. Estimation of a multivariate Box–Cox transformation to elliptical symmetry via the empirical characteristic function. Annals of the Institute of Statistical Mathematics, 48(4):687–709, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052328.

Karunamuni:1996:EBS

[1996] Rohana J. Karunamuni. Empirical Bayes sequential estimation for exponential families: The untruncated component. Annals of the Institute of Statistical Mathematics, 48(4): 711–730, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052329.

Kong:1996:LTM

[1997] Fanhui Kong and Heliang Fei. Limit theorems for the maximum likelihood estimate under general multiply type II censoring. Annals of the Institute of Statistical Mathematics, 48(4): 731–755, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00052330.

Aggarwala:1996:RRS

[1998] Rita Aggarwala and N. Balakrishnan. Recurrence relations for single and product moments of progressive Type-II right censored order statistics from exponential and truncated exponential distributions. Annals of the Institute of Statistical Mathematics, 48(4):757–771, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052331.

Aki:1996:SLW

[1999] S. Aki, N. Balakrishnan, and S. G. Mohanty. Sooner and later waiting time problems for success and failure runs in higher order Markov dependent trials. Annals of the Institute of Statistical Mathematics, 48(4):773–787, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052332.

Koutras:1996:WTD

[2000] M. V. Koutras. On a waiting time distribution in a sequence of Bernoulli trials. Annals of the Institute of Statistical Mathematics, 48(4):789–806, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052333.

Anonymous:1996:A

[2001] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 48(4):807–809, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00052334.

Anonymous:1996:HCd

[2002] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 48(4):??, December 1996. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Beran:1997:DBS

[2003] Rudolf Beran. Diagnosing bootstrap success. Annals of the Institute of Statistical Mathematics, 49(1): 1–24, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003114420352.

Lo:1997:MSG

[2004] Albert Y. Lo and V. V. Sazonov. Von Mises ω^{2}-statistic and the generalized Bayesian bootstraps. Annals of the Institute of Statistical Mathematics, 49(1):25–34, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003106504422.

Zhang:1997:QPP

[2005] Biao Zhang. Quantile processes in the presence of auxiliary information. Annals of the Institute of Statistical Mathematics, 49(1):35–55, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003158521261.

Kundu:1997:BDN

[2006] Subrata Kundu and Adam T. Martinsek. Bounding the L_{1} distance in nonparametric density estimation. Annals of the Institute of Statistical Mathematics, 49(1):57–78, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003110605331.

Fan:1997:LPR

[2007] Jianqing Fan, Theo Gasser, Irène Gijbels, Michael Brockmann, and Joachim Engel. Local polynomial regression: Optimal kernels and asymptotic minimax efficiency. Annals of the Institute of Statistical Mathematics, 49 (1):79–99, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003162622169.

Nakashima:1997:SME

[2008] Eiji Nakashima. Some methods for estimation in a negative-binomial model. Annals of the Institute of Statistical Mathematics, 49(1):101– 115, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003114706239.

Papadatos:1997:NMV

[2009] Nickos Papadatos. A note on maximum variance of order statistics from symmetric populations. Annals of the Institute of Statistical Mathematics, 49 (1):117–121, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003166723078.

Koutras:1997:WTD

[2010] M. V. Koutras. Waiting time distributions associated with runs of fixed length in two-state Markov chains. Annals of the Institute of Statistical Mathematics, 49(1):123– 139, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003118807148.

Godbole:1997:FRJ

[2011] Anant P. Godbole, Stavros G. Papastavridis, and Robert S. Weishaar. Formulae and recursions for the joint distribution of success runs of several lengths. Annals of the Institute of Statistical Mathematics, 49 (1):141–153, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003170823986.

Balakrishnan:1997:SDT

[2012] N. Balakrishnan, S. G. Mohanty, and S. Aki. Start-up demonstration tests under Markov dependence model with corrective actions. Annals of the Institute of Statistical Mathematics, 49 (1):155–169, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003122908057.

Gupta:1997:UMP

[2013] Arjun K. Gupta and Jacek Wesolowski.

Uniform mixtures via posterior means. Annals of the Institute of Statistical Mathematics, 49(1):171–180, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003175024895.

Ghosal:1997:AEP

[2014] Subhashis Ghosal and Tapas Samanta. Asymptotic expansions of posterior distributions in nonregular cases. Annals of the Institute of Statistical Mathematics, 49(1):181–197, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003127108965.

Anonymous:1997:HCa

[2015] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 49(1):??, March 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hirose:1997:CRF

[2016] Keiichi Hirose, Eiichi Isogai, and Chikara Uno. The convergence rate of fixed-width sequential confidence intervals for a parameter of an exponential distribution. Annals of the Institute of Statistical Mathematics, 49(2):199–209, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003121726848.

Sen:1997:EDR

[2017] Ananda Sen and Arthur Fries. Estimation in a discrete reliability growth model under an inverse sampling Scheme. Annals of the Institute of Statistical Mathematics, 49(2):211– 229, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003102610918.

Chen:1997:NBI

[2018] Zhiqiang Chen and Eswar Phadia. A note on the best invariant estimator of a distribution function under the Kolmogorov–Smirnov loss. Annals of the Institute of Statistical Mathematics, 49(2):231–235, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003154627756.

Bhattacharya:1997:TSA

[2019] Bhaskar Bhattacharya. On tests of symmetry against one-sided alternatives. Annals of the Institute of Statistical Mathematics, 49(2):237– 254, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003106711827.

DelCastillo:1997:TDG

[2020] Joan Del Castillo and Pedro Puig.

Testing departures from gamma, Rayleigh and truncated normal distributions. Annals of the Institute of Statistical Mathematics, 49(2):255– 269, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003158828665.

Horvath:1997:DCL

[2021] Lajos Horváth. Detection of changes in linear sequences. Annals of the Institute of Statistical Mathematics, 49(2):271–283, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003110912735.

Shimizu:1997:SEB

[2022] Ryoichi Shimizu and Yasunori Fujikoshi. Sharp error bounds for asymptotic expansions of the distribution functions for scale mixtures. Annals of the Institute of Statistical Mathematics, 49(2):285–297, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003162929574.

Ramachandran:1997:GSL

[2023] B. Ramachandran. On geometricstable laws, a related property of stable processes, and stable densities of exponent one. Annals of the Institute of Statistical Mathematics, 49 (2):299–313, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003115013644.

Pusz:1997:RCG

[2024] J. Pusz. Regressional characterization of the generalized inverse Gaussian population. Annals of the Institute of Statistical Mathematics, 49 (2):315–319, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003167030482.

Venter:1997:LMS

[2025] J. H. Venter and J. L. J. Snyman. Linear model selection based on risk estimation. Annals of the Institute of Statistical Mathematics, 49 (2):321–340, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003119114553.

Li:1997:AES

[2026] Ta-Hsin Li and Gerald R. North. Aliasing effects and sampling theorems of spherical random fields when sampled on a finite Grid. Annals of the Institute of Statistical Mathematics, 49(2):341–354, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003171131391.

Cuevas:1997:DFS

[2027] Antonio Cuevas and Juan Romo. Differentiable functionals and smoothed bootstrap. Annals of the Institute of Statistical Mathematics, 49(2):355– 370, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003123215461.

Woodbury:1997:CMH

[2028] Max A. Woodbury, Kenneth G. Manton, and H. Dennis Tolley. Convex models of high dimensional discrete data. Annals of the Institute of Statistical Mathematics, 49 (2):371–393, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003175232300.

Anonymous:1997:HCb

[2029] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 49(2):??, June 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shimodaira:1997:AEP

[2030] Hidetoshi Shimodaira. Assessing the error probability of the model selection test. Annals of the Institute of Statistical Mathematics, 49(3): 395–410, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003140609666.

Ishiguro:1997:BLL

[2031] Makio Ishiguro, Yosiyuki Sakamoto, and Genshiro Kitagawa. Bootstrapping log likelihood and EIC, an extension of AIC. Annals of the Institute of Statistical Mathematics, 49(3): 411–434, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003158526504.

Bhattacharya:1997:FDA

[2032] Bhaskar Bhattacharya and Richard L. Dykstra. A Fenchel duality aspect of iterative I-projection procedures. Annals of the Institute of Statistical Mathematics, 49(3):435–446, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003110627413.

Adimari:1997:ELT

[2033] Gianfranco Adimari. Empirical likelihood type confidence intervals under random censorship. Annals of the Institute of Statistical Mathematics, 49 (3):447–466, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003114711483.

VanKeilegom:1997:EBC

[2034] Ingrid Van Keilegom and Noël Veraverbeke. Estimation and bootstrap with censored data in fixed design nonparametric regression. Annals of the Institute of Statistical Mathematics, 49 (3):467–491, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003166728321.

Gao:1997:SIS

[2035] Jiti Gao and Hua Liang. Statistical inference in single-index and partially nonlinear models. Annals of the Institute of Statistical Mathematics, 49 (3):493–517, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003118812392.

Balakrishnan:1997:JDN

[2036] N. Balakrishnan. Joint distributions of numbers of success-runs and failures until the first consecutive k successes in a binary sequence. Annals of the Institute of Statistical Mathematics, 49 (3):519–529, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003170829230.

Antzoulakos:1997:PDF

[2037] Demetrios L. Antzoulakos and Andreas N. Philippou. Probability distribution functions of succession quotas in the case of Markov dependent trials. Annals of the Institute of Statistical Mathematics, 49(3): 531–539, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003123013300.

Wu:1997:DMF

[2038] Eden K. H. Wu and P. S. Chan. Distributions minimizing Fisher information for location in Kolmogorov neighbourhoods. Annals of the Institute of Statistical Mathematics, 49(3): 541–554, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003175030138.

Baringhaus:1997:CCP

[2039] Ludwig Baringhaus and Rudolf Grübel. On a class of characterization problems for random convex combinations. Annals of the Institute of Statistical Mathematics, 49(3):555–567, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003127114209.

Stepniak:1997:CNL

[2040] Czeslaw Stepniak. Comparison of normal linear experiments by quadratic forms. Annals of the Institute of Statistical Mathematics, 49(3):569–584, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003179131047.

Larsson:1997:AES

[2041] Rolf Larsson. On the asymptotic expectations of some unit root tests in a first order autoregressive process in the presence of trend. Annals of the Institute of Statistical Mathematics, 49 (3):585–599, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003131215117.

Anonymous:1997:HCc

[2042] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 49(3):??, September 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Copas:1997:FND

[2043] J. B. Copas and C. B. Stride. Fitting a normal distribution when the model is wrong. Annals of the Institute of Statistical Mathematics, 49 (4):601–614, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003220407410.

Falk:1997:MC

[2044] Michael Falk. On mad and comedians. Annals of the Institute of Statistical Mathematics, 49(4):615–644, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003258024248.

Marohn:1997:LAN

[2045] Frank Marohn. Local asymptotic normality in extreme value index estimation. Annals of the Institute of Statistical Mathematics, 49(4):645–666, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003210125157.

Bingham:1997:EDC

[2046] N. H. Bingham and Bruce Dunham. Estimating diffusion coefficients from count data: Einstein–Smoluchowski theory revisited. Annals of the Institute of Statistical Mathematics, 49 (4):667–679, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003214209227.

Son:1997:CTI

[2047] M. S. Son, L. D. Haugh, H. I. Hamdy, and M. C. Costanza. Controlling type II error while constructing triple sampling fixed precision confidence intervals for the normal mean. Annals of the Institute of Statistical Mathematics, 49 (4):681–692, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003266326065.

Lingham:1997:THA

[2048] Rama T. Lingham and S. Sivaganesan. Testing hypotheses about the power law process under failure truncation using intrinsic Bayes factors. Annals of the Institute of Statistical Mathematics, 49(4):693–710, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003218410136.

Goutis:1997:RBP

[2049] Constantinos Goutis and George Casella. Relationships between postdata accuracy measures. Annals of the Institute of Statistical Mathematics, 49 (4):711–726, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003270426974.

Papadatos:1997:EBE

[2050] Nickos Papadatos. Exact bounds for the expectations of order statistics from non-negative populations. Annals of the Institute of Statistical Mathematics, 49(4):727–736, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003222527882.

Maejima:1997:MLL

[2051] Makoto Maejima. Moments of limits of lightly trimmed sums of random vectors in the generalized domain of normal attraction of non–Gaussian operator-stable laws. Annals of the Institute of Statistical Mathematics, 49 (4):737–747, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003226611953.

Shenton:1997:RDU

[2052] L. R. Shenton and K. O. Bowman.

Replenishment-depletion urn in equilibrium. Annals of the Institute of Statistical Mathematics, 49(4):749–760, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003278628791.

Caussinus:1997:CLM

[2053] Henri Caussinus and Faouzi Lyazrhi. Choosing a linear model with a random number of change-points and outliers. Annals of the Institute of Statistical Mathematics, 49(4):761–775, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003230713770.

Hallin:1997:BET

[2054] Marc Hallin and Khalid Rifi. A Berry-Esséen theorem for serial rank statistics. Annals of the Institute of Statistical Mathematics, 49(4):777–799, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003286814679.

Anonymous:1997:HCd

[2055] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 49(4):??, December 1997. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shimodaira:1998:AMC

[2056] Hidetoshi Shimodaira. An application of multiple comparison techniques to model selection. Annals of the Institute of Statistical Mathematics, 50(1):1–13, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003483128844.

Ebrahimi:1998:EFP

[2057] Nader Ebrahimi. Estimating the finite population versions of mean residual life-time function and hazard function using Bayes method. Annals of the Institute of Statistical Mathematics, 50(1):15–27, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003493112915.

Larsson:1998:OET

[2058] Rolf Larsson. The order of the error term for moments of the log likelihood ratio unit root test in an autoregressive process. Annals of the Institute of Statistical Mathematics, 50(1):29–48, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003445229753.

Takahasi:1998:DBO

[2059] Koiti Takahasi and Masao Futatsuya. Dependence between order statistics in samples from finite population and its application to ranked set sampling. Annals of the Institute of Statistical Mathematics, 50 (1):49–70, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003497213823.

Arcones:1998:BKR

[2060] Miguel A. Arcones. The Bahadur– Kiefer representation of the two dimensional spatial medians. Annals of the Institute of Statistical Mathematics, 50(1):71–86, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003449330662.

Arcones:1998:SOR

[2061] Miguel A. Arcones. Second order representations of the least absolute deviation regression estimator. Annals of the Institute of Statistical Mathematics, 50(1):87–117, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003401414732.

Zheng:1998:SFW

[2062] Shen Zheng, T. N. Sriram, and Andrew F. Seila. Sequential fixed-width confidence interval for the product of two means. Annals of the Institute of Statistical Mathematics, 50 (1):119–145, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003453431570.

Tsimikas:1998:AMU

[2063] John V. Tsimikas and Johannes Ledolter. Analysis of multi-unit variance components models with state space profiles. Annals of the Institute of Statistical Mathematics, 50 (1):147–164, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003405615641.

Mohapl:1998:MLE

[2064] Jaroslav Mohapl. On maximum likelihood estimation for Gaussian spatial autoregression models. Annals of the Institute of Statistical Mathematics, 50 (1):165–186, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003457632479.

Benn:1998:RFB

[2065] A. G. Benn and R. J. Kulperger. A remark on a Fourier bounding method of proof for convergence of sums of periodograms. Annals of the Institute of Statistical Mathematics, 50 (1):187–202, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003409716549.

Anonymous:1998:HCa

[2066] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 50(1):??, March 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Uchida:1998:JDN

[2067] Masayuki Uchida. Joint distributions of numbers of success-runs until the first consecutive k successes in a higherorder two-state Markov chain. Annals of the Institute of Statistical Mathematics, 50(2):203–222, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003489713112.

Zhang:1998:DSS

[2068] Jian Zhang. Data sphering: Some properties and applications. Annals of the Institute of Statistical Mathematics, 50(2):223–240, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003482914021.

Nel:1998:DBF

[2069] D. G. Nel and I. Pienaar. The decomposition of the Behrens–Fisher statistic in q-dimensional common principal component submodels. Annals of the Institute of Statistical Mathematics, 50(2):241–252, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003487030859.

Dette:1998:SGF

[2070] Holger Dette and Axel Munk. A simple goodness-of-fit test for linear models under a random design assumption. Annals of the Institute of Statistical Mathematics, 50 (2):253–275, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003439114929.

Wei:1998:TVD

[2071] Bo-Cheng Wei, Jian-Qing Shi, Wing-Kam Fung, and Yue-Qing Hu. Testing for varying dispersion in exponential family nonlinear models. Annals of the Institute of Statistical Mathematics, 50(2):277–294, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003491131768.

Zografos:1998:DSD

[2072] K. Zografos. f-Dissimilarity of several distributions in testing statistical hypotheses. Annals of the Institute of Statistical Mathematics, 50 (2):295–310, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003443215838.

Chen:1998:STW

[2073] Yuh-Ing Chen. Simple-tree weighted logrank tests for right-censored data. Annals of the Institute of Statistical Mathematics, 50(2):311–324, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003495232676.

Takada:1998:NPB

[2074] Yoshikazu Takada. The nonexistence of procedures with bounded performance characteristics in certain parametric inference problems. Annals of the Institute of Statistical Mathematics, 50(2):325–335, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003447316747.

Datta:1998:ITP

[2075] Somnath Datta and William P. Mc-Cormick. Inference for the tail parameters of a linear process with heavy tail innovations. Annals of the Institute of Statistical Mathematics, 50 (2):337–359, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003499300817.

Takahashi:1998:PMS

[2076] Rinya Takahashi and Masaaki Sibuya. Prediction of the maximum size in Wicksell’s corpuscle problem. Annals of the Institute of Statistical Mathematics, 50(2):361–377, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003451417655.

Ogata:1998:STP

[2077] Yosihiko Ogata. Space-time pointprocess models for earthquake occurrences. Annals of the Institute of Statistical Mathematics, 50(2):379– 402, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003403601725.

Anonymous:1998:HCb

[2078] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 50(2):??, June 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Achcar:1998:UMC

[2079] Jorge A. Achcar and Roseli A. Leandro. Use of Markov chain Monte Carlo methods in a Bayesian analysis of the block and Basu bivariate exponential distribution. Annals of the Institute of Statistical Mathematics, 50 (3):403–416, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003582409664.

Leung:1998:CAL

[2080] Chi-Ying Leung. The covariance adjusted location linear discriminant function for classifying data with unequal dispersion matrices in different locations. Annals of the Institute of Statistical Mathematics, 50(3): 417–431, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003517226502.

Corcuera:1998:CMR

[2081] J. M. Corcuera and F. Giummolè. A characterization of monotone and regular divergences. Annals of the Institute of Statistical Mathematics, 50(3): 433–450, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003569210573.

Maesono:1998:ACS

[2082] Yoshihiko Maesono. Asymptotic comparisons of several variance estimators and their effects for Studentizations. Annals of the Institute of Statistical Mathematics, 50(3):451–470, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003521327411.

Dumbgen:1998:TFS

[2083] Lutz Dümbgen. On Tyler’s Mfunctional of scatter in high dimension. Annals of the Institute of Statistical Mathematics, 50(3):471–491, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003573311481.

Hopfner:1998:MDE

[2084] R. Höpfner and Yu. A. Kutoyants. On minimum distance estimation in recurrent Markov step processes II. Annals of the Institute of Statistical Mathematics, 50(3):493–502, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003525428320.

Pak:1998:MDE

[2085] Ro Jin Pak and Ayanendranath Basu. Minimum disparity estimation in linear regression models: Distribution and efficiency. Annals of the Institute of Statistical Mathematics, 50(3): 503–521, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003577412390.

Leung:1998:ESM

[2086] Pui Lam Leung and Wai Yin Chan. Estimation of the scale matrix and its eigenvalues in the Wishart and the multivariate F distributions. Annals of the Institute of Statistical Mathematics, 50 (3):523–530, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003529529228.

Krishnamoorthy:1998:SST

[2087] K. Krishnamoorthy and Maruthy K.

Pannala. Some simple test procedures for normal mean vector with incomplete data. Annals of the Institute of Statistical Mathematics, 50(3): 531–542, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003581513299.

Wesolowski:1998:DPE

[2088] Jacek Wesolowski and Mohammad Ahsanullah. Distributional properties of exceedance statistics. Annals of the Institute of Statistical Mathematics, 50 (3):543–565, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003533730137.

DelCastillo:1998:WPD

[2089] Joan Del Castillo and Marta Pérez-Casany. Weighted Poisson distributions for overdispersion and underdispersion situations. Annals of the Institute of Statistical Mathematics, 50 (3):567–585, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003585714207.

Uchida:1998:NOS

[2090] Masayuki Uchida. On number of occurrences of success runs of specified length in a higher-order two-state Markov chain. Annals of the Institute of Statistical Mathematics, 50(3): 587–601, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003537831046.

Anonymous:1998:HCc

[2091] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 50(3):??, September 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Matsunawa:1998:PSU

[2092] T. Matsunawa. Parametric statistical uncertainty relations and parametric statistical fundamental equations. Annals of the Institute of Statistical Mathematics, 50(4):603–626, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003715611735.

Vu:1998:APC

[2093] H. T. V. Vu, R. A. Maller, and X. Zhou. Asymptotic properties of a class of mixture models for failure data: The interior and boundary cases. Annals of the Institute of Statistical Mathematics, 50 (4):627–653, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003704728573.

Uchida:1998:GFW

[2094] Masayuki Uchida. On generating functions of waiting time problems for sequence patterns of discrete random variables. Annals of the Institute of Statistical Mathematics, 50(4): 655–671, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003756712643.

Jensen:1998:LDO

[2095] Jens Ledet Jensen and Andrew T. A.

Wood. Large deviation and other results for minimum contrast estimators. Annals of the Institute of Statistical Mathematics, 50(4):673–695, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003708829482.

Brown:1998:CLS

[2096] Bruce M. Brown and Song Xi Chen.

Combined and least squares empirical likelihood. Annals of the Institute of Statistical Mathematics, 50(4): 697–714, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003760813552.

Sengupta:1998:BEA

[2097] Ashis Sengupta and Ranjan Maitra.

On best equivariance and admissibility of simultaneous MLE for mean direction vectors of several Langevin distributions. Annals of the Institute of Statistical Mathematics, 50(4): 715–727, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003712930390.

Vilar-Fernandez:1998:RER

[2098] J. A. Vilar-Fernández and J. M. Vilar-Fernández. Recursive estimation of regression functions by local polynomial fitting. Annals of the Institute of Statistical Mathematics, 50(4): 729–754, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003764914460.

Gu:1998:ALI

[2099] Hong Gu and Wing K. Fung. Assessing local influence in canonical correlation analysis. Annals of the Institute of Statistical Mathematics, 50 (4):755–772, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003717031299.

Pal:1998:ECM

[2100] Nabendu Pal and Wooi K. Lim. Estimation of the coefficient of multiple determination. Annals of the Institute of Statistical Mathematics, 50 (4):773–788, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003769115369.

Chang:1998:SSM

[2101] Yen-Chang Chang and Lii-Yuh Leu. A state space model for software reliability. Annals of the Institute of Statistical Mathematics, 50(4):789–799, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003721232207.

Anonymous:1998:HCd

[2102] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 50(4):??, December 1998. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aki:1999:DRC

[2103] Sigeo Aki. Distributions of runs and consecutive systems on directed trees. Annals of the Institute of Statistical Mathematics, 51(1):1–15, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003877016437.

Aki:1999:SLW

[2104] Sigeo Aki and Katuomi Hirano. Sooner and later waiting time problems for runs in Markov dependent bivariate trials. Annals of the Institute of Statistical Mathematics, 51(1):17– 29, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003874900507.

Shao:1999:SCM

[2105] Yongzhao Shao and Marjorie G. Hahn.

Strong consistency of the maximum product of spacings estimates with applications in nonparametrics and in estimation of unimodal densities. Annals of the Institute of Statistical Mathematics, 51(1):31–49, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003827017345.

Sanchez-Sellero:1999:BSD

[2106] C. Sánchez-Sellero, W. González-Manteiga, and R. Cao. Bandwidth selection in density estimation with truncated and censored data. Annals of the Institute of Statistical Mathematics, 51(1):51–70, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003879001416.

Bingham:1999:NPE

[2107] N. H. Bingham and Susan M. Pitts. Non-parametric estimation for the M/G/∞ queue. Annals of the Institute of Statistical Mathematics, 51 (1):71–97, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003831118254.

Takagi:1999:PIR

[2108] Yoshiji Takagi. Parametrization invariance with respect to second order admissibility under mean squared error. Annals of the Institute of Statistical Mathematics, 51(1):99–110, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003883202324.

Vidakovic:1999:LVN

[2109] Brani Vidakovic. Linear versus nonlinear rules for mixture normal priors. Annals of the Institute of Statistical Mathematics, 51(1):111–124, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003835319163.

Delicado:1999:GFT

[2110] Pedro Delicado and Juan Romo. Goodness of fit tests in random coefficient regression models. Annals of the Institute of Statistical Mathematics, 51 (1):125–148, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003887303233.

Karlis:1999:TNC

[2111] Dimitris Karlis and Evdokia Xekalaki.

On testing for the number of components in a mixed Poisson model. Annals of the Institute of Statistical Mathematics, 51(1):149–162, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003839420071.

Clarke:1999:AEP

[2112] Bertrand Clarke and Dongchu Sun.

Asymptotics of the expected posterior. Annals of the Institute of Statistical Mathematics, 51(1):163–185, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003891404142.

Wu:1999:SOE

[2113] Yanhong Wu. Second order expansions for the moments of minimum point of an unbalanced two-sided normal random walk. Annals of the Institute of Statistical Mathematics, 51 (1):187–200, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003843520980.

Anonymous:1999:HCa

[2114] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 51(1):??, March 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Godambe:1999:LBO

[2115] V. P. Godambe. Linear Bayes and optimal estimation. Annals of the Institute of Statistical Mathematics, 51(2):201–215, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003893706155.

Basu:1999:CCI

[2116] Sanjib Basu. Conservatism of the z confidence interval under symmetric and asymmetric departures from normality. Annals of the Institute of Statistical Mathematics, 51(2):217– 230, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1003849922993.

Gijbels:1999:EJP

[2117] Irene Gijbels, Peter Hall, and Aloïs Kneip. On the estimation of jump points in smooth curves. Annals of the Institute of Statistical Mathematics, 51(2):231–251, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003802007064.

Shi:1999:MEM

[2118] Daoji Shi and Shengsheng Zhou. Moment estimation for multivariate extreme value distribution in a nested logistic model. Annals of the Institute of Statistical Mathematics, 51 (2):253–264, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003854023902.

Helmers:1999:EPI

[2119] Roelof Helmers and Ricardas Zitikis. On estimation of Poisson intensity functions. Annals of the Institute of Statistical Mathematics, 51 (2):265–280, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003806107972.

Huang:1999:EBP

[2120] Wen-Tao Huang and Yao-Tsung Lai.

Empirical Bayes procedures for selecting the best population with multiple criteria. Annals of the Institute of Statistical Mathematics, 51 (2):281–299, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003858124810.

Saha:1999:DEU

[2121] Krishna K. Saha and B. C. Sutradhar. On the distribution of the extremes of unequally correlated normal variables with applications to antedependent cluster data. Annals of the Institute of Statistical Mathematics, 51(2):301–322, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003810208881.

Antzoulakos:1999:WTP

[2122] Demetrios L. Antzoulakos. On waiting time problems associated with runs in Markov dependent trials. Annals of the Institute of Statistical Mathematics, 51(2):323–330, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003862225719.

Kottas:1999:BDP

[2123] Athanasios Kottas, Konstantinos Adamidis, and Sotirios Loukas. Bivariate distributions with Pearson type VII conditionals. Annals of the Institute of Statistical Mathematics, 51 (2):331–344, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003814309789.

Huang:1999:EDC

[2124] J. S. Huang and G. D. Lin. Equality in distribution in a convex ordering family. Annals of the Institute of Statistical Mathematics, 51 (2):345–349, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003866326628.

Barakat:1999:MBO

[2125] H. M. Barakat. On moments of bivariate order statistics. Annals of the Institute of Statistical Mathematics, 51(2):351–358, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003818410698.

Aly:1999:FET

[2126] Emad-Eldin A. A. Aly and Nadjib Bouzar. On the first entry time of a Z_{+}-valued AR(1) process. Annals of the Institute of Statistical Mathematics, 51(2):359–367, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003870427536.

Alsmeyer:1999:LTS

[2127] Gerold Alsmeyer and Allan Gut. Limit theorems for stopped functionals of Markov renewal processes. Annals of the Institute of Statistical Mathematics, 51(2):369–382, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003822611607.

Lee:1999:CTD

[2128] Andy H. Lee and John S. Yick. Covariate transformation diagnostics for generalized linear models. Annals of the Institute of Statistical Mathematics, 51(2):383–398, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003874628445.

Anonymous:1999:HCb

[2129] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 51(2):??, June 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kleibergen:1999:OPS

[2130] Frank Kleibergen, Herman K. van Dijk, and Jean-Pierre Urbain. Oil price shocks and long run price and import demand behavior. Annals of the Institute of Statistical Mathematics, 51 (3):399–417, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003978303477.

Han:1999:JDR

[2131] Qing Han and Sigeo Aki. Joint distributions of runs in a sequence of multi-state trials. Annals of the Institute of Statistical Mathematics, 51 (3):419–447, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003941920316.

Maejima:1999:CPA

[2132] Makoto Maejima and Gennady Samorodnitsky. Certain probabilistic aspects of semistable laws. Annals of the Institute of Statistical Mathematics, 51 (3):449–462, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003993904386.

Pazman:1999:SPI

[2133] Andrej Pázman. Some properties and improvements of the saddlepoint approximation in nonlinear regression. Annals of the Institute of Statistical Mathematics, 51(3):463–478, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003946021224.

Ghosal:1999:APD

[2134] Subhashis Ghosal, Jayanta K. Ghosh, and Tapas Samanta. Approximation of the posterior distribution in a changepoint problem. Annals of the Institute of Statistical Mathematics, 51(3): 479–497, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003998005295.

Basu:1999:PSS

[2135] Sanjib Basu. Posterior sensitivity to the sampling distribution and the prior: More than one observation. Annals of the Institute of Statistical Mathematics, 51(3):499–513, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003950122133.

Aerts:1999:BPM

[2136] Marc Aerts and Gerda Claeskens.

Bootstrapping pseudolikelihood models for clustered binary data. Annals of the Institute of Statistical Mathematics, 51(3):515–530, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1003902206203.

Ardanuy:1999:WEW

[2137] Ramón Ardanuy, J. López-Fidalgo, Patrick J. Laycock, and Weng Kee Wong. When is an equally-weighted design D-optimal? Annals of the Institute of Statistical Mathematics, 51 (3):531–540, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003954207112.

Fraiman:1999:CB

[2138] Ricardo Fraiman and Jean Meloche. Counting bumps. Annals of the Institute of Statistical Mathematics, 51 (3):541–569, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003958323950.

Gupta:1999:PIE

[2139] Ramesh C. Gupta, S. Ramakrishnan, and Xingwang Zhou. Point and interval estimation of P (X < Y): The normal case with common coefficient of variation. Annals of the Institute of Statistical Mathematics, 51(3): 571–584, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1003910408020.

Kim:1999:MTN

[2140] Hea-Jung Kim. A method for testing nested point null hypotheses using multiple Bayes factor. Annals of the Institute of Statistical Mathematics, 51 (3):585–602, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1003962424859.

Anonymous:1999:HCc

[2141] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 51(3):??, September 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Wang:1999:NEF

[2142] Jinpang Wang. Nonconservative estimating functions and approximate quasi-likelihoods. Annals of the Institute of Statistical Mathematics, 51 (4):603–619, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004096728035.

Mukhopadhyay:1999:TSP

[2143] Nitis Mukhopadhyay and William Duggan. On a two-stage procedure having second-order properties with applications. Annals of the Institute of Statistical Mathematics, 51(4): 621–636, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004074912105.

Chen:1999:EBE

[2144] Ming-Hui Chen and Qi-Man Shao. Existence of Bayesian estimates for the polychotomous quantal response models. Annals of the Institute of Statistical Mathematics, 51(4):637–656, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004027028943.

Assuncao:1999:RIP

[2145] Renato Assunção and Peter Guttorp. Robustness for inhomogeneous Poisson point processes. Annals of the Institute of Statistical Mathematics, 51 (4):657–678, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004079013014.

Blaker:1999:ACR

[2146] Helge Blaker. On adaptive combination of regression estimators. Annals of the Institute of Statistical Mathematics, 51 (4):679–689, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004031129852.

Lu:1999:MLP

[2147] Zhan-Qian Lu. Multivariate local polynomial fitting for martingale nonlinear regression models. Annals of the Institute of Statistical Mathematics, 51 (4):691–706, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004083213922.

Fujikoshi:1999:GCM

[2148] Yasunori Fujikoshi, Takashi Kanda, and Megu Ohtaki. Growth curve model with hierarchical within-individuals design matrices. Annals of the Institute of Statistical Mathematics, 51(4): 707–721, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004035330761.

Schwabe:1999:EBP

[2149] Rainer Schwabe and Weng Kee Wong.

Efficiency bounds for product designs in linear models. Annals of the Institute of Statistical Mathematics, 51 (4):723–730, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004087314831.

Fotopoulos:1999:EBA

[2150] Stergios B. Fotopoulos and Lijian He.

Error bounds for asymptotic expansion of the conditional variance of the scale mixtures of the multivariate normal distribution. Annals of the Institute of Statistical Mathematics, 51 (4):731–747, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004039431669.

Hwang:1999:CGD

[2151] Tea-Yuan Hwang and Chin-Yuan Hu.

On a characterization of the gamma distribution: The independence of the sample mean and the sample coefficient of variation. Annals of the Institute of Statistical Mathematics, 51 (4):749–753, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004091415740.

Pfanzagl:1999:RLD

[2152] J. Pfanzagl. On rates and limit distributions. Annals of the Institute of Statistical Mathematics, 51(4): 755–778, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004043532578.

Misawa:1999:CQS

[2153] Tetsuya Misawa. Conserved quantities and symmetries related to stochastic dynamical systems. Annals of the Institute of Statistical Mathematics, 51 (4):779–802, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004095516648.

Anonymous:1999:HCd

[2154] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 51(4):??, December 1999. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kuriki:2000:SGC

[2155] Satoshi Kuriki and Akimichi Takemura. Some geometry of the cone of nonnegative definite matrices and weights of associated X^{2} distribution. Annals of the Institute of Statistical Mathematics, 52(1):1–14, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004191713375.

Huimei:2000:UMP

[2156] Liu Huimei. Uniformly more powerful, two-sided tests for hypotheses about linear inequalities. Annals of the Institute of Statistical Mathematics, 52(1):15–27, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004176730214.

Rousseau:2000:CPO

[2157] Judith Rousseau. Coverage properties of one-sided intervals in the discrete case and application to matching priors. Annals of the Institute of Statistical Mathematics, 52(1):28– 42, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004128814284.

Bar-Hen:2000:CLF

[2158] A. Bar-Hen and H. Kishino. Comparing the likelihood functions of phylogenetic trees. Annals of the Institute of Statistical Mathematics, 52 (1):43–56, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004180831122.

Vidoni:2000:MSU

[2159] Paolo Vidoni. Model selection using the estimative and the approximate p∗ predictive densities. Annals of the Institute of Statistical Mathematics, 52(1):57–70, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004132915192.

Bhattacharyya:2000:TUR

[2160] B. B. Bhattacharyya, X. Li, M. Pensky, and G. D. Richardson. Testing for unit roots in a nearly nonstationary spatial autoregressive process. Annals of the Institute of Statistical Mathematics, 52(1):71–83, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004184932031.

Mohapl:2000:SAD

[2161] Jaroslav Mohapl. A stochastic advection–diffusion model for the rocky flats soil plutonium data. Annals of the Institute of Statistical Mathematics, 52(1):84–107, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004137016101.

Grillenzoni:2000:TVP

[2162] Carlo Grillenzoni. Time-varying parameters prediction. Annals of the Institute of Statistical Mathematics, 52 (1):108–122, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004189000171.

Feuerverger:2000:PET

[2163] Andrey Feuerverger and Yehuda Vardi. Positron emission tomography and random coefficients regression. Annals of the Institute of Statistical Mathematics, 52(1):123–138, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004141117010.

Paparoditis:2000:LBK

[2164] Efstathios Paparoditis and Dimitris N. Politis. The local bootstrap for kernel estimators under general dependence conditions. Annals of the Institute of Statistical Mathematics, 52 (1):139–159, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004193117918.

Datta:2000:CAN

[2165] Somnath Datta, Glen A. Satten, and John M. Williamson. Consistency and asymptotic normality of estimators in a proportional hazards model with interval censoring and left truncation. Annals of the Institute of Statistical Mathematics, 52 (1):160–172, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004197201989.

Bhapkar:2000:OEB

[2166] Vasant P. Bhapkar. On the optimality of estimators based on P - sufficient statistics. Annals of the Institute of Statistical Mathematics, 52 (1):173–183, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004149318827.

Balakrishnan:2000:SDT

[2167] N. Balakrishnan and P. S. Chan. Startup demonstration tests with rejection of units upon observing d failures. Annals of the Institute of Statistical Mathematics, 52(1):184–196, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004101402897.

Anonymous:2000:HCa

[2168] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 52(1):??, March 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Yamada:2000:QAO

[2169] Tomoya Yamada and Tadashi Matsunawa. Quantitative approximation to the ordered Dirichlet distribution under varying basic probability spaces. Annals of the Institute of Statistical Mathematics, 52 (2):197–214, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004153419736.

Lindsay:2000:MBA

[2170] Bruce G. Lindsay, Ramani S. Pilla, and Prasanta Basak. Moment-based approximations of distributions using mixtures: Theory and applications. Annals of the Institute of Statistical Mathematics, 52(2):215– 230, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004105603806.

Kozubowski:2000:EMR

[2171] Tomasz J. Kozubowski. Exponential mixture representation of geometric stable distributions. Annals of the Institute of Statistical Mathematics, 52(2):231–238, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004157620644.

Kaur:2000:OAS

[2172] Amarjot Kaur, G. P. Patil, and C. Taillie. Optimal allocation for symmetric distributions in ranked set sampling. Annals of the Institute of Statistical Mathematics, 52(2):239– 254, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004109704714.

Stoimenova:2000:RTB

[2173] Eugenia Stoimenova. Rank tests based on exceeding observations. Annals of the Institute of Statistical Mathematics, 52(2):255–266, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004161721553.

Gurtler:2000:GFT

[2174] Nora Gürtler and Norbert Henze.

Goodness-of-fit tests for the Cauchy distribution based on the empirical characteristic function. Annals of the Institute of Statistical Mathematics, 52(2):267–286, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004113805623.

Buhlmann:2000:MSV

[2175] Peter Bühlmann. Model selection for variable length Markov chains and tuning the context algorithm. Annals of the Institute of Statistical Mathematics, 52(2):287–315, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004165822461.

Yao:2000:LSE

[2176] Jian-Feng Yao. On least squares estimation for stable nonlinear AR processes. Annals of the Institute of Statistical Mathematics, 52(2):316– 331, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004117906532.

Wan:2000:OVM

[2177] Alan T. K. Wan and Anoop Chaturvedi. Operational variants of the minimum mean squared error estimator in linear regression models with non-spherical disturbances. Annals of the Institute of Statistical Mathematics, 52 (2):332–342, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004169923370.

Yuan:2000:EEN

[2178] Ke-Hai Yuan and Robert I. Jennrich. Estimating equations with nuisance parameters: Theory and applications. Annals of the Institute of Statistical Mathematics, 52(2):343– 350, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004122007440.

Collins:2000:RCS

[2179] John R. Collins. Robustness comparisons of some classes of location parameter estimators. Annals of the Institute of Statistical Mathematics, 52(2):351–366, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004174024279.

Zhong:2000:IAL

[2180] Xu-Ping Zhong, Bo-Cheng Wei, and Wing-Kam Fung. Influence analysis for linear measurement error models. Annals of the Institute of Statistical Mathematics, 52(2):367– 379, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004126108349.

Zou:2000:SES

[2181] Guohua Zou and Alan T. K. Wan. Simultaneous estimation of several stratum means under error-in-variables superpopulation models. Annals of the Institute of Statistical Mathematics, 52(2):380–396, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004178125187.

Lin:2000:SOS

[2182] Chien-Tai Lin and C. J. Shiau. Some optimal strategies for bandit problems with beta prior distributions. Annals of the Institute of Statistical Mathematics, 52(2):397–405, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004130209258.

Anonymous:2000:HCb

[2183] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 52(2):??, June 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Han:2000:SLW

[2184] Qing Han and Sigeo Aki. Sooner and later waiting time problems based on a dependent sequence. Annals of the Institute of Statistical Mathematics, 52 (3):407–414, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004182030590.

Fu:2000:JDR

[2185] James C. Fu and W. Y. Wendy Lou. Joint distribution of rises and falls. Annals of the Institute of Statistical Mathematics, 52(3):415–425, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004148814661.

Garren:2000:ADE

[2186] Steven T. Garren. Asymptotic distribution of estimated affinity between multiparameter exponential families. Annals of the Institute of Statistical Mathematics, 52(3):426–437, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004100931499.

Yuan:2000:BDR

[2187] Lin Yuan and John D. Kalbfleisch. On the Bessel distribution and related problems. Annals of the Institute of Statistical Mathematics, 52(3): 438–447, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004152916478.

Bairamov:2000:CPE

[2188] Ismihan G. Bairamov. On the characteristic properties of exponential distribution. Annals of the Institute of Statistical Mathematics, 52(3): 448–458, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004109117386.

Honda:2000:NEC

[2189] Toshio Honda. Nonparametric estimation of a conditional quantile for αmixing processes. Annals of the Institute of Statistical Mathematics, 52 (3):459–470, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004113201457.

Chen:2000:PDF

[2190] Song Xi Chen. Probability density function estimation using gamma kernels. Annals of the Institute of Statistical Mathematics, 52(3):471–480, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004165218295.

Seidel:2000:CNL

[2191] Wilfried Seidel, Karl Mosler, and Manfred Alker. A cautionary note on likelihood ratio tests in mixture models. Annals of the Institute of Statistical Mathematics, 52(3):481–487, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004117419204.

Larocque:2000:BST

[2192] Denis Larocque, Serge Tardif, and Constance van Eeden. Bivariate sign tests based on the Sup, L_{1} and L_{2} norms. Annals of the Institute of Statistical Mathematics, 52(3):488–506, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004121503274.

Hall:2000:RBP

[2193] Peter Hall, Brett Presnell, and Berwin A. Turlach. Reducing bias without prejudicing sign. Annals of the Institute of Statistical Mathematics, 52 (3):507–518, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1004173520112.

Wu:2000:TSS

[2194] Colin O. Wu, Kai Fun Yu, and Chin-Tsang Chiang. A two-step smoothing method for varying-coefficient models with repeated measurements. Annals of the Institute of Statistical Mathematics, 52(3):519–543, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1004125621021.

Chung:2000:DAW

[2195] Hie-Choon Chung and Chien-Pai Han.

Discriminant analysis when a block of observations is missing. Annals of the Institute of Statistical Mathematics, 52 (3):544–556, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004129706000.

Dette:2000:COD

[2196] Holger Dette and Mong-Na Lo Huang. Convex optimal designs for compound polynomial extrapolation. Annals of the Institute of Statistical Mathematics, 52(3):557–573, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1004185822838.

Dette:2000:UAS

[2197] Holger Dette and Yuri Grigoriev. A unified approach to second order optimality criteria in nonlinear design of experiments. Annals of the Institute of Statistical Mathematics, 52(3): 574–597, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1004137906908.

Anonymous:2000:HCc

[2198] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 52(3):??, September 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Honda:2000:NDE

[2199] Toshio Honda. Nonparametric density estimation for a long-range dependent linear process. Annals of the Institute of Statistical Mathematics, 52 (4):599–611, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017504723799.

Zhang:2000:BBC

[2200] Shunpu Zhang and Rohana J. Karunamuni. Boundary bias correction for nonparametric deconvolution. Annals of the Institute of Statistical Mathematics, 52(4):612–629, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017564907869.

Mehra:2000:LIL

[2201] K. L. Mehra, Y. S. Ramakrishnaiah, and P. Sashikala. Laws of iterated logarithm and related asymptotics for estimators of conditional density and mode. Annals of the Institute of Statistical Mathematics, 52(4): 630–645, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017517124707.

Martinsek:2000:SEM

[2202] Adam T. Martinsek. Sequential estimation of the maximum in a model for corrosion data. Annals of the Institute of Statistical Mathematics, 52 (4):646–657, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017569108778.

Bischoff:2000:AOT

[2203] Wolfgang Bischoff and Frank Miller.

Asymptotically optimal tests and optimal designs for testing the mean in regression models with applications to change-point problems. Annals of the Institute of Statistical Mathematics, 52 (4):658–679, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017521225616.

Pan:2000:NMC

[2204] Guohua Pan. Nonparametric methods for checking the validity of prior order information. Annals of the Institute of Statistical Mathematics, 52 (4):680–697, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017573209686.

Gimenez:2000:HTE

[2205] Patricia Gimenez, Heleno Bolfarine, and Enrico A. Colosimo. Hypotheses testing for error-in-variables models. Annals of the Institute of Statistical Mathematics, 52(4):698–711, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017525326525.

Xiao:2000:LUP

[2206] Yushan Xiao. Linex unbiasedness in a prediction problem. Annals of the Institute of Statistical Mathematics, 52 (4):712–721, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017577310595.

Lee:2000:JBA

[2207] Sik-Yum Lee and Jian-Qing Shi. Joint Bayesian analysis of factor scores and structural parameters in the factor analysis model. Annals of the Institute of Statistical Mathematics, 52(4): 722–736, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017529427433.

Pan:2000:BIA

[2208] Jian-Xin Pan and Wing-Kam Fung. Bayesian influence assessment in the growth curve model with unstructured covariance. Annals of the Institute of Statistical Mathematics, 52 (4):737–752, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017581411504.

Gu:2000:IDC

[2209] Hong Gu and Wing K. Fung. Influence diagnostics in the common canonical variates model. Annals of the Institute of Statistical Mathematics, 52 (4):753–766, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017533528342.

Aki:2000:NSR

[2210] Sigeo Aki and Katuomi Hirano. Numbers of success-runs of specified length until certain stopping time rules and generalized binomial distributions of order k. Annals of the Institute of Statistical Mathematics, 52(4):767–777, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017585512412.

Han:2000:WTP

[2211] Qing Han and Sigeo Aki. Waiting time problems in a two-state Markov chain. Annals of the Institute of Statistical Mathematics, 52(4): 778–789, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017537629251.

Aly:2000:GID

[2212] Emad-Eldin A. A. Aly and Nadjib Bouzar. On geometric infinite divisibility and stability. Annals of the Institute of Statistical Mathematics, 52 (4):790–799, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017589613321.

Anonymous:2000:HCd

[2213] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 52(4):??, December 2000. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Professor:2001:GSM

[2214] H. Akaike Emeritus Professor. Golf swing motion analysis: An experiment on the use of verbal analysis in statistical reasoning. Annals of the Institute of Statistical Mathematics, 53(1):1–10, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017952001329.

Kitagawa:2001:SIN

[2215] Genshiro Kitagawa and Tomoyuki Higuchi. Special issue on nonlinear non-Gaussian models and related filtering methods. Annals of the Institute of Statistical Mathematics, 53(1):3, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1017947900421.

Fahrmeir:2001:BSR

[2216] Ludwig Fahrmeir and Stefan Lang. Bayesian semiparametric regression analysis of multicategorical time–space data. Annals of the Institute of Statistical Mathematics, 53(1):11–30, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017904118167.

Fruhwirth-Schnatter:2001:FBA

[2217] Sylvia Frühwirth-Schnatter. Fully Bayesian analysis of switching Gaussian state space models. Annals of the Institute of Statistical Mathematics, 53(1):31–49, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017908219076.

Takahashi:2001:MCF

[2218] Akihiko Takahashi and Seisho Sato.

A Monte Carlo filtering approach for estimating the term structure of interest rates. Annals of the Institute of Statistical Mathematics, 53 (1):50–62, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017964304055.

Tanizaki:2001:NNG

[2219] Hisashi Tanizaki. Nonlinear and non– Gaussian state space modeling using sampling techniques. Annals of the Institute of Statistical Mathematics, 53(1):63–81, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017916420893.

Godsill:2001:MPS

[2220] Simon Godsill, Arnaud Doucet, and Mike West. Maximum a posteriori sequence estimation using Monte Carlo particle filters. Annals of the Institute of Statistical Mathematics, 53(1):82–96, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017968404964.

Bergman:2001:OEC

[2221] Niclas Bergman, Arnaud Doucet, and Neil Gordon. Optimal estimation and Cramér–Rao bounds for partial non– Gaussian state space models. Annals of the Institute of Statistical Mathematics, 53(1):97–112, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017920621802.

Brockwell:2001:LDC

[2222] P. J. Brockwell. Lévy-driven Carma processes. Annals of the Institute of Statistical Mathematics, 53(1):113– 124, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1017972605872.

Chandra:2001:EFN

[2223] S. Ajay Chandra and Masanobu Taniguchi. Estimating functions for nonlinear time series models. Annals of the Institute of Statistical Mathematics, 53(1):125–141, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017924722711.

Shiohama:2001:SEF

[2224] Takayuki Shiohama and Masanobu Taniguchi. Sequential estimation for a functional of the spectral density of a Gaussian stationary process. Annals of the Institute of Statistical Mathematics, 53(1):142–158, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1017976706781.

Truong:2001:AWB

[2225] Young K. Truong and Prakash N. Patil. Asymptotics for wavelet based estimates of piecewise smooth regression for stationary time series. Annals of the Institute of Statistical Mathematics, 53(1):159–178, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017928823619.

Hardle:2001:WQT

[2226] Wolfgang Härdle, Torsten Kleinow, and Rolf Tschernig. Web quantlets for time series analysis. Annals of the Institute of Statistical Mathematics, 53 (1):179–188, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1017980807689.

Anonymous:2001:HCa

[2227] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 53(1):??, March 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Sibuya:2001:OTM

[2228] Masaaki Sibuya and Kazuyuki Suzuki.

Optimal threshold for the k-out-of-n monitor with dual failure modes. Annals of the Institute of Statistical Mathematics, 53(2):189–202, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1012451319428.

Kirmani:2001:POM

[2229] S. N. U. A. Kirmani and Ramesh C.

Gupta. On the proportional odds model in survival analysis. Annals of the Institute of Statistical Mathematics, 53(2):203–216, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012458303498.

Jiang:2001:EBP

[2230] Jiming Jiang and P. Lahiri. Empirical best prediction for small area inference with binary data. Annals of the Institute of Statistical Mathematics, 53(2):217–243, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012410420337.

Kusnier:2001:TBB

[2231] Jozef Kusnier and Ivan Mizera. Tail behavior and breakdown properties of equivariant estimators of location. Annals of the Institute of Statistical Mathematics, 53(2):244– 261, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1012462404407.

Suzukawa:2001:KLI

[2232] Akio Suzukawa, Hideyuki Imai, and Yoshiharu Sato. Kullback–Leibler information consistent estimation for censored data. Annals of the Institute of Statistical Mathematics, 53 (2):262–276, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012414621245.

Menendez:2001:MDE

[2233] M. Menéndez, D. Morales, L. Pardo, and I. Vajda. Minimum divergence estimators based on grouped data. Annals of the Institute of Statistical Mathematics, 53(2):277–288, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012466605316.

Belitser:2001:ALM

[2234] Eduard Belitser and Boris Levit.

Asymptotically local minimax estimation of infinitely smooth density with censored data. Annals of the Institute of Statistical Mathematics, 53(2):289–306, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012418722154.

Cramer:2001:ESO

[2235] Erhard Cramer and Udo Kamps. Estimation with sequential order statistics from exponential distributions. Annals of the Institute of Statistical Mathematics, 53(2):307–324, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012470706224.

Ebrahimi:2001:TUR

[2236] Nader Ebrahimi. Testing for uniformity of the residual life time based on dynamic Kullback–Leibler information. Annals of the Institute of Statistical Mathematics, 53(2):325– 337, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1012085320762.

Klar:2001:GFT

[2237] Bernhard Klar. Goodness-of-fit tests for the exponential and the normal distribution based on the integrated distribution function. Annals of the Institute of Statistical Mathematics, 53(2):338–353, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012422823063.

Ahmed:2001:SES

[2238] S. E. Ahmed, A. K. Gupta, S. M. Khan, and C. J. Nicol. Simultaneous estimation of several intraclass correlation coefficients. Annals of the Institute of Statistical Mathematics, 53(2):354–369, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012474807133.

Gnot:2001:SIL

[2239] Stanislaw Gnot, Ewaryst Rafajlowicz, and Agnieszka Urbańska-Motyka. Statistical inference in a linear model for spatially located sensors and random input. Annals of the Institute of Statistical Mathematics, 53 (2):370–379, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012426923971.

Chakraborty:2001:AEM

[2240] Biman Chakraborty. On affine equivariant multivariate quantiles. Annals of the Institute of Statistical Mathematics, 53(2):380–403, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012478908041.

Singh:2001:GCA

[2241] Sarjinder Singh. Generalized calibration approach for estimating variance in survey sampling. Annals of the Institute of Statistical Mathematics, 53(2):404–417, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012431008950.

Montepiedra:2001:NDC

[2242] Grace Montepiedra and Weng Kee Wong. A new design criterion when heteroscedasticity is ignored. Annals of the Institute of Statistical Mathematics, 53(2):418–426, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1012435125788.

Anonymous:2001:HCb

[2243] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 53(2):??, June 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Akahira:2001:IIF

[2244] Masafumi Akahira and Kei Takeuchi.

Information inequalities in a family of uniform distributions. Annals of the Institute of Statistical Mathematics, 53(3):427–435, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014685808327.

Iliopoulos:2001:DTE

[2245] George Iliopoulos. Decision theoretic estimation of the ratio of variances in a bivariate normal distribution. Annals of the Institute of Statistical Mathematics, 53(3):436–446, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014600625165.

Lu:2001:ANK

[2246] Zudi Lu. Asymptotic normality of kernel density estimators under dependence. Annals of the Institute of Statistical Mathematics, 53(3):447– 468, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1014652626073.

Yu:2001:APS

[2247] Qiqing Yu, George Y. C. Wong, and Linxiong Li. Asymptotic properties of self-consistent estimators with mixed interval-censored data. Annals of the Institute of Statistical Mathematics, 53 (3):469–486, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014656726982.

Barakat:2001:ADT

[2248] H. M. Barakat. The asymptotic distribution theory of bivariate order statistics. Annals of the Institute of Statistical Mathematics, 53 (3):487–497, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014660811052.

Ren:2001:WEL

[2249] Jian-Jian Ren. Weighted empirical likelihood ratio confidence intervals for the mean with censored data. Annals of the Institute of Statistical Mathematics, 53(3):498–516, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014612911961.

Wang:2001:ELC

[2250] Qi-Hua Wang and Bing-Yi Jing. Empirical likelihood for a class of functionals of survival distribution with censored data. Annals of the Institute of Statistical Mathematics, 53 (3):517–527, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014617112870.

Tsairidis:2001:IQR

[2251] Ch. Tsairidis, K. Zografos, K. Ferentinos, and T. Papaioannou. Information in quantal response data and random censoring. Annals of the Institute of Statistical Mathematics, 53 (3):528–542, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014621229708.

Hill:2001:EVM

[2252] Theodore P. Hill and Victor Perez-Abreu. Extreme-value moment goodness-of-fit tests. Annals of the Institute of Statistical Mathematics, 53 (3):543–551, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014673230617.

Grabovsky:2001:CPD

[2253] Irina Grabovsky and Lajos Horváth. Change-point detection in angular data. Annals of the Institute of Statistical Mathematics, 53(3):552–566, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014677314687.

Muller:2001:SOM

[2254] Alfred Müller. Stochastic ordering of multivariate normal distributions. Annals of the Institute of Statistical Mathematics, 53(3):567–575, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014629416504.

Chadjiconstantinidis:2001:DNF

[2255] Stathis Chadjiconstantinidis and Markos V. Koutras. Distributions of the numbers of failures and successes in a waiting time problem. Annals of the Institute of Statistical Mathematics, 53 (3):576–598, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014637600574.

Antzoulakos:2001:DNS

[2256] Demetrios L. Antzoulakos and Stathis Chadjiconstantinidis. Distributions of numbers of success runs of fixed length in Markov dependent trials. Annals of the Institute of Statistical Mathematics, 53(3):599–619, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014689617413.

Khaledi:2001:DPM

[2257] Baha-Eldin Khaledi and Subhash Kochar. Dependence properties of multivariate mixture distributions and their applications. Annals of the Institute of Statistical Mathematics, 53 (3):620–630, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014641701483.

Kochar:2001:RSR

[2258] Subhash C. Kochar and Ramesh Korwar. On random sampling without replacement from a finite population. Annals of the Institute of Statistical Mathematics, 53(3):631– 646, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1014693702392.

Takahashi:2001:PMS

[2259] Rinya Takahashi and Masaaki Sibuya.

Prediction of the maximum size in Wicksell’s corpuscle problem, II. Annals of the Institute of Statistical Mathematics, 53(3):647–660, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1014697919230.

Anonymous:2001:HCc

[2260] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 53(3):??, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mase:2001:PDS

[2261] S. Mase, J. Møller, D. Stoyan, R. P. Waagepetersen, and G. Döge. Packing densities and simulated tempering for hard core Gibbs point processes. Annals of the Institute of Statistical Mathematics, 53(4):661– 680, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014662415827.

Pensky:2001:NES

[2262] Marianna Pensky and Brani Vidakovic. On non-equally spaced wavelet regression. Annals of the Institute of Statistical Mathematics, 53(4):681– 690, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014640632666.

Walk:2001:SUP

[2263] Harro Walk. Strong universal pointwise consistency of recursive regression estimates. Annals of the Institute of Statistical Mathematics, 53 (4):691–707, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014692616736.

Li:2001:BAN

[2264] Gang Li and Somnath Datta. A bootstrap approach to nonparametric regression for right censored data. Annals of the Institute of Statistical Mathematics, 53(4):708–729, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014644700806.

VanKeilegom:2001:HRE

[2265] Ingrid Van Keilegom and Noël Veraverbeke. Hazard rate estimation in nonparametric regression with censored data. Annals of the Institute of Statistical Mathematics, 53 (4):730–745, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014696717644.

Petropoulos:2001:EEQ

[2266] Constantinos Petropoulos and Stavros Kourouklis. Estimation of an exponential quantile under a general loss and an alternative estimator under quadratic loss. Annals of the Institute of Statistical Mathematics, 53 (4):746–759, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014648819462.

Zhou:2001:EMN

[2267] Xian Zhou, Xiaoqian Sun, and Jinglong Wang. Estimation of the multivariate normal precision matrix under the entropy loss. Annals of the Institute of Statistical Mathematics, 53 (4):760–768, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014657020370.

Leung:2001:IEP

[2268] Pui Lam Leung and Foon Yip Ng. Improved estimation of parameter matrices in a one-sample and twosample problems. Annals of the Institute of Statistical Mathematics, 53 (4):769–780, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014661221279.

vanRooij:2001:AIE

[2269] Arnoud C. M. van Rooij and Frits H. Ruymgaart. Abstract inverse estimation with application to deconvolution on locally compact Abelian groups. Annals of the Institute of Statistical Mathematics, 53(4):781– 798, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014665305349.

Novikov:2001:UAE

[2270] Andrei Novikov. Uniform asymptotic expansion of likelihood ratio for Markov dependent observations. Annals of the Institute of Statistical Mathematics, 53(4):799–809, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1014617422188.

Inglot:2001:IAC

[2271] Tadeusz Inglot and Teresa Ledwina. Intermediate approach to comparison of some goodness-of-fit tests. Annals of the Institute of Statistical Mathematics, 53(4):810–834, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014669423096.

Lee:2001:WKM

[2272] Wei-Chun Lee and Yuh-Ing Chen.

Weighted Kaplan–Meier tests for umbrella alternatives. Annals of the Institute of Statistical Mathematics, 53 (4):835–852, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014673524005.

Nanda:2001:HRR

[2273] Asok K. Nanda and Moshe Shaked.

The hazard rate and the reversed hazard rate orders, with applications to order statistics. Annals of the Institute of Statistical Mathematics, 53 (4):853–864, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1014677608075.

Gelfand:2001:NBM

[2274] Alan E. Gelfand and Athanasios Kottas. Nonparametric Bayesian modeling for stochastic order. Annals of the Institute of Statistical Mathematics, 53(4):865–876, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014629724913.

Rychlik:2001:SOS

[2275] Tomasz Rychlik. Stability of order statistics under dependence. Annals of the Institute of Statistical Mathematics, 53(4):877–894, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014681708984.

Robin:2001:EDD

[2276] S. Robin and J.-J. Daudin. Exact distribution of the distances between any occurrences of a set of words. Annals of the Institute of Statistical Mathematics, 53(4):895–905, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1014633825822.

Anonymous:2001:A

[2277] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 53(4):907–909, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1017313413010.

Anonymous:2001:HCd

[2278] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 53(4):??, ???? 2001. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Huang:2002:GPL

[2279] Fuchun Huang and Yosihiko Ogata. Generalized pseudo-likelihood estimates for Markov random fields on lattice. Annals of the Institute of Statistical Mathematics, 54(1):1–18, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1016170102988.

Liebscher:2002:KDH

[2280] Eckhard Liebscher. Kernel density and hazard rate estimation for censored data under α-mixing condition. Annals of the Institute of Statistical Mathematics, 54(1):19–28, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016157519826.

Bogdan:2002:DDS

[2281] M. Bogdan, K. Bogdan, and A. Futschik.

A data driven smooth test for circular uniformity. Annals of the Institute of Statistical Mathematics, 54 (1):29–44, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016109603897.

Wong:2002:DDC

[2282] H. Wong and W. K. Li. Detecting and diagnostic checking multivariate conditional heteroscedastic time series models. Annals of the Institute of Statistical Mathematics, 54 (1):45–59, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016161620735.

Derbort:2002:TAN

[2283] Stephan Derbort, Holger Dette, and Axel Munk. A test for additivity in nonparametric regression. Annals of the Institute of Statistical Mathematics, 54(1):60–82, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1016113704805.

Pensky:2002:LAW

[2284] Marianna Pensky. Locally adaptive wavelet empirical Bayes estimation of a location parameter. Annals of the Institute of Statistical Mathematics, 54(1):83–99, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1016165721644.

Lin:2002:BSP

[2285] Yu-Pin Lin, TaChen Liang, and Wen-Tao Huang. Bayesian sampling plans for exponential distribution based on Type I censoring data. Annals of the Institute of Statistical Mathematics, 54(1):100–113, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1016117805714.

Wang:2002:EDS

[2286] Nan Wang and Wei Liu. Expansions for the distributions of some normalized summations of random numbers of I.I.D. random variables. Annals of the Institute of Statistical Mathematics, 54(1):114–124, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016169822552.

Aly:2002:NMG

[2287] Emad-Eldin A. A. Aly and Nadjib Bouzar. A notion of αmonotonicity with generalized multiplications. Annals of the Institute of Statistical Mathematics, 54 (1):125–137, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016121906623.

Mudholkar:2002:IGM

[2288] Govind S. Mudholkar and Rajeshwari Natarajan. The inverse Gaussian models: Analogues of symmetry, skewness and kurtosis. Annals of the Institute of Statistical Mathematics, 54 (1):138–154, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016173923461.

Hunter:2002:CEP

[2289] David R. Hunter and Kenneth Lange.

Computing estimates in the proportional odds model. Annals of the Institute of Statistical Mathematics, 54 (1):155–168, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016126007531.

Higuchi:2002:SSN

[2290] Tomoyuki Higuchi and Genshiro Kitagawa. Special section on nonparametric approach to time series analysis. Annals of the Institute of Statistical Mathematics, 54(1):169, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1016178024370.

Ombao:2002:SMN

[2291] Hernando Ombao, Jonathan Raz, Rainer von Sachs, and Wensheng Guo. The SLEX model of a non-stationary random process. Annals of the Institute of Statistical Mathematics, 54 (1):171–200, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016130108440.

Stoffer:2002:LSE

[2292] David S. Stoffer, Hernando C. Ombao, and David E. Tyler. Local spectral envelope: An approach using dyadic treebased adaptive segmentation. Annals of the Institute of Statistical Mathematics, 54(1):201–223, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016182125278.

Kilminster:2002:EIP

[2293] Devin Kilminster, David Allingham, and Alistair Mees. Estimating invariant probability densities for dynamical systems. Annals of the Institute of Statistical Mathematics, 54 (1):224–233, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1016134209348.

Anonymous:2002:HCa

[2294] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 54(1):??, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Adimari:2002:QPL

[2295] G. Adimari and L. Ventura. Quasiprofile log likelihoods for unbiased estimating functions. Annals of the Institute of Statistical Mathematics, 54 (2):235–244, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022425816250.

Schick:2002:EID

[2296] Anton Schick and Wolfgang Wefelmeyer. Estimating the innovation distribution in nonlinear autoregressive models. Annals of the Institute of Statistical Mathematics, 54 (2):245–260, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022413700321.

Visek:2002:SAE

[2297] Jan Ámos Vísek. Sensitivity analysis of M-estimates of nonlinear regression model: Influence of data subsets. Annals of the Institute of Statistical Mathematics, 54(2):261–290, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022465701229.

Beran:2002:LPF

[2298] Jan Beran and Yuanhua Feng. Local polynomial fitting with longmemory, short-memory and antipersistent errors. Annals of the Institute of Statistical Mathematics, 54 (2):291–311, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022469818068.

Chen:2002:LLS

[2299] Song Xi Chen. Local linear smoothers using asymmetric kernels. Annals of the Institute of Statistical Mathematics, 54(2):312–323, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022422002138.

Shieh:2002:BPA

[2300] Gwowen Shieh and Jack C. Lee. Bayesian prediction analysis for growth curve model using noninformative priors. Annals of the Institute of Statistical Mathematics, 54(2):324–337, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022474018976.

Chen:2002:PWM

[2301] Ming-Hui Chen and Qi-Man Shao. Partition-weighted Monte Carlo estimation. Annals of the Institute of Statistical Mathematics, 54(2):338– 354, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022426103047.

Mukerjee:2002:PDI

[2302] Rahul Mukerjee and Brajendra C. Sutradhar. On the positive definiteness of the information matrix under the binary and Poisson mixed models. Annals of the Institute of Statistical Mathematics, 54(2):355– 366, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022478119885.

Chatterjee:2002:DAG

[2303] Snigdhansu Chatterjee and Arup Bose.

Dimension asymptotics for generalised bootstrap in linear regression. Annals of the Institute of Statistical Mathematics, 54(2):367–381, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022430203955.

Nikitin:2002:AEM

[2304] Ya. Yu. Nikitin and E. V. Ponikarov.

Asymptotic efficiency of Maesono statistics for testing of symmetry. Annals of the Institute of Statistical Mathematics, 54(2):382–390, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022482220794.

Kadri:2002:ABC

[2305] M’hammed Kadri and Khalid Rifi.

Asymptotic bound on the characteristic function of signed linear serial rank statistics. Annals of the Institute of Statistical Mathematics, 54 (2):391–403, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022434304864.

Roy:2002:BLM

[2306] Dilip Roy. On bivariate lack of memory property and a new definition. Annals of the Institute of Statistical Mathematics, 54(2):404–410, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022486321702.

Chryssaphinou:2002:CPA

[2307] Ourania Chryssaphinou and Eutichia Vaggelatou. Compound Poisson approximation for multiple runs in a Markov chain. Annals of the Institute of Statistical Mathematics, 54 (2):411–424, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022438422611.

Henze:2002:GFT

[2308] Norbert Henze and Bernhard Klar. Goodness-of-fit tests for the inverse Gaussian distribution based on the empirical Laplace transform. Annals of the Institute of Statistical Mathematics, 54(2):425–444, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022442506681.

Schoenberg:2002:RPP

[2309] Frederic Paik Schoenberg. On rescaled Poisson processes and the Brownian bridge. Annals of the Institute of Statistical Mathematics, 54 (2):445–457, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022494523519.

Anonymous:2002:HCb

[2310] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 54(2):??, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kanamori:2002:SAT

[2311] Takafumi Kanamori. Statistical asymptotic theory of active learning. Annals of the Institute of Statistical Mathematics, 54(3):459–475, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022446624428.

Kosorok:2002:ATG

[2312] Michael R. Kosorok, Jason P. Fine, Hongyu Jiang, and Rick Chappell. Asymptotic theory for the gamma frailty model with dependent censoring. Annals of the Institute of Statistical Mathematics, 54(3):476–499, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022450708498.

Sivaganesan:2002:ASI

[2313] S. Sivaganesan and Rama T. Lingham. On the asymptotic stability of the intrinsic and fractional Bayes factors for testing some diffusion models. Annals of the Institute of Statistical Mathematics, 54(3):500–516, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022402925337.

Patra:2002:GCC

[2314] Kaushik Patra and Dipak K. Dey. A general class of change point and change curve modeling for life time data. Annals of the Institute of Statistical Mathematics, 54(3):517–530, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022454909407.

Swanepoel:2002:GFT

[2315] Jan W. H. Swanepoel and François C. Van Graan. Goodness-of-fit tests based on estimated expectations of probability integral transformed order statistics. Annals of the Institute of Statistical Mathematics, 54 (3):531–542, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022407026245.

Fokianos:2002:PDF

[2316] Konstantinos Fokianos. Power divergence family of tests for categorical time series models. Annals of the Institute of Statistical Mathematics, 54 (3):543–564, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022459010316.

Bhattacharya:2002:TPS

[2317] Bhaskar Bhattacharya. Tests of parameters of several gamma distributions with inequality restrictions. Annals of the Institute of Statistical Mathematics, 54(3):565–576, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022411127154.

Zhang:2002:BAD

[2318] Jin Zhang and Yuehua Wu. Beta approximation to the distribution of Kolmogorov–Smirnov statistic. Annals of the Institute of Statistical Mathematics, 54(3):577–584, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022463111224.

Ciuperca:2002:LRS

[2319] Gabriela Ciuperca. Likelihood ratio statistic for exponential mixtures. Annals of the Institute of Statistical Mathematics, 54(3):585–594, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022415228062.

Zhong:2002:ELM

[2320] Xu-Ping Zhong, Wing-Kam Fung, and Bo-Cheng Wei. Estimation in linear models with random effects and errors-in-variables. Annals of the Institute of Statistical Mathematics, 54 (3):595–606, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022467212133.

Babu:2002:LPI

[2321] Gutti Jogesh Babu and Eugenijus Manstavicius. Limit processes with independent increments for the Ewens sampling formula. Annals of the Institute of Statistical Mathematics, 54 (3):607–620, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022419328971.

Nomakuchi:2002:MMC

[2322] Kentaro Nomakuchi. A monotonicity of moments concerned with order restricted statistical inference. Annals of the Institute of Statistical Mathematics, 54(3):621–625, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022471313041.

Zhao:2002:RSS

[2323] Xiaoyue Zhao and Zehua Chen. On the ranked-set sampling M-estimates for symmetric location families. Annals of the Institute of Statistical Mathematics, 54(3):626–640, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022423429880.

Zheng:2002:MML

[2324] Gang Zheng and Mohammad F. Al-Saleh. Modified maximum likelihood estimators based on ranked set samples. Annals of the Institute of Statistical Mathematics, 54 (3):641–658, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022475413950.

Arnab:2002:ESM

[2325] Raghunath Arnab and Sarjinder Singh.

Estimation of the size and mean value of a stigmatized characteristic of a hidden gang in a finite population: A unified approach. Annals of the Institute of Statistical Mathematics, 54 (3):659–666, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022427530788.

Abo-Eleneen:2002:FIO

[2326] Z. A. Abo-Eleneen and H. N. Nagaraja.

Fisher information in an order statistic and its concomitant. Annals of the Institute of Statistical Mathematics, 54(3):667–680, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022479514859.

Inoue:2002:GWT

[2327] Kiyoshi Inoue and Sigeo Aki. Generalized waiting time problems associated with pattern in Polya’s urn Scheme. Annals of the Institute of Statistical Mathematics, 54(3):681– 688, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022431631697.

Bradley:2002:DSN

[2328] David M. Bradley and Ramesh C. Gupta. On the distribution of the sum of n non-identically distributed uniform random variables. Annals of the Institute of Statistical Mathematics, 54(3):689–700, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022483715767.

Pensky:2002:DDD

[2329] Marianna Pensky and Ahmed I. Zayed. Density deconvolution of different conditional distributions. Annals of the Institute of Statistical Mathematics, 54(3):701–712, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022435832605.

Anonymous:2002:HCc

[2330] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 54(3):??, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aki:2002:WTR

[2331] Sigeo Aki and Katuomi Hirano. On waiting time for reversed patterns in random sequences. Annals of the Institute of Statistical Mathematics, 54 (4):713–718, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022472731832.

Fu:2002:ELD

[2332] James C. Fu, W. Y. Wendy Lou, Zhi-Dong Bai, and Gang Li. The exact and limiting distributions for the number of successes in success runs within a sequence of Markov-dependent two-state trials. Annals of the Institute of Statistical Mathematics, 54 (4):719–730, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022451015903.

Lefevre:2002:GPD

[2333] Claude Lefèvre, Vasilis Papathanasiou, and Sergey Utev. Generalized Pearson distributions and related characterization problems. Annals of the Institute of Statistical Mathematics, 54 (4):731–742, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022403132741.

Yang:2002:SAD

[2334] Bo Yang and John E. Kolassa. Saddlepoint approximation for the distribution function near the mean. Annals of the Institute of Statistical Mathematics, 54(4):743–747, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022455116811.

Bacro:2002:BLS

[2335] J.-N. Bacro, J.-J. Daudin, S. Mercier, and S. Robin. Back to the local score in the logarithmic case: A direct and simple proof. Annals of the Institute of Statistical Mathematics, 54 (4):748–757, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022407200882.

Ellis:2002:BDW

[2336] Steven P. Ellis. Blind deconvolution when noise is symmetric: Existence and examples of solutions. Annals of the Institute of Statistical Mathematics, 54(4):758–767, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022459217720.

Chiaromonte:2002:SDR

[2337] Francesca Chiaromonte and R. Dennis Cook. Sufficient dimension reduction and graphics in regression. Annals of the Institute of Statistical Mathematics, 54(4):768–795, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022411301790.

Chandrasekar:2002:MVT

[2338] B. Chandrasekar and N. Balakrishnan. On a multiparameter version of Tukey’s linear sensitivity measure and its properties. Annals of the Institute of Statistical Mathematics, 54 (4):796–805, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022463318629.

Akahira:2002:IIB

[2339] Masafumi Akahira and Nao Ohyauchi. Information inequalities for the Bayes risk for a family of non-regular distributions. Annals of the Institute of Statistical Mathematics, 54(4):806– 815, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022415402699.

Kotz:2002:MLE

[2340] Samuel Kotz, Tomasz J. Kozubowski, and Krzysztof Podgórski. Maximum likelihood estimation of asymmetric Laplace parameters. Annals of the Institute of Statistical Mathematics, 54 (4):816–826, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022467519537.

Bohning:2002:ENI

[2341] Dankmar Böhning, Uwe Malzahn, Jesus Sarol, Jr., Sasivimol Rattanasiri, and Annibale Biggeri. Efficient noniterative and nonparametric estimation of heterogeneity variance for the standardized mortality ratio. Annals of the Institute of Statistical Mathematics, 54(4):827–839, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022419603608.

Hwang:2002:NME

[2342] Tea-Yuan Hwang and Ping-Huang Huang. On new moment estimation of parameters of the gamma distribution using its characterization. Annals of the Institute of Statistical Mathematics, 54(4):840–847, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022471620446.

Chang:2002:CRU

[2343] Yuan-Tsung Chang and Nobuo Shinozaki. A comparison of restricted and unrestricted estimators in estimating linear functions of ordered scale parameters of two gamma distributions. Annals of the Institute of Statistical Mathematics, 54(4):848– 860, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1022423704516.

Yu:2002:ECM

[2344] Philip L. H. Yu, Yijun Sun, and Bimal K. Sinha. Estimation of the common mean of a bivariate normal population. Annals of the Institute of Statistical Mathematics, 54 (4):861–878, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022475721354.

Kohler:2002:UCL

[2345] Michael Kohler. Universal consistency of local polynomial kernel regression estimates. Annals of the Institute of Statistical Mathematics, 54 (4):879–899, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1022427805425.

Beran:2002:IPL

[2346] Rudolf Beran. Improving penalized least squares through adaptive selection of penalty and shrinkage. Annals of the Institute of Statistical Mathematics, 54(4):900–917, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022479822263.

Hayakawa:2002:ILR

[2347] Takesi Hayakawa. Independence of likelihood ratio criteria for homogeneity of several populations. Annals of the Institute of Statistical Mathematics, 54(4):918–933, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022431906333.

Park:2002:MPT

[2348] Hyo-Il Park. Multivariate percentile tests for incomplete data. Annals of the Institute of Statistical Mathematics, 54(4):934–944, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022483923172.

Dette:2002:ODT

[2349] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. D-optimal designs for trigonometric regression models on a partial circle. Annals of the Institute of Statistical Mathematics, 54 (4):945–959, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022436007242.

Anonymous:2002:HCd

[2350] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 54(4):??, ???? 2002. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Datta:2003:PMP

[2351] Gauri Sankar Datta and Rahul Mukerjee. Probability matching priors for predicting a dependent variable with application to regression models. Annals of the Institute of Statistical Mathematics, 55(1):1–6, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530481.

Basu:2003:BES

[2352] Sanjib Basu and Rama T. Lingham. Bayesian estimation of system reliability in Brownian stress-strength models. Annals of the Institute of Statistical Mathematics, 55(1):7–19, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530482.

Wang:2003:EPL

[2353] Qi-Hua Wang. Estimation of partial linear error-in-response models with validation data. Annals of the Institute of Statistical Mathematics, 55(1):21–39, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530483.

Bibi:2003:CAN

[2354] Abdelouahab Bibi and Christian Francq. Consistent and asymptotically normal estimators for cyclically timedependent linear models. Annals of the Institute of Statistical Mathematics, 55(1):41–68, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530484.

Chanda:2003:DEC

[2355] Kamal C. Chanda. Density estimation for a class of stationary nonlinear processes. Annals of the Institute of Statistical Mathematics, 55(1):69–82, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530485.

Babu:2003:EEC

[2356] Gutti Jogesh Babu, Kesar Singh, and Yaning Yang. Edgeworth expansions for compound Poisson processes and the bootstrap. Annals of the Institute of Statistical Mathematics, 55 (1):83–94, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530486.

Pfanzagl:2003:ABE

[2357] J. Pfanzagl. Asymptotic bounds for estimators without limit distribution. Annals of the Institute of Statistical Mathematics, 55(1):95–110, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530487.

Dewan:2003:MWT

[2358] Isha Dewan and B. L. S. Prakasa Rao.

Mann–Whitney test for associated sequences. Annals of the Institute of Statistical Mathematics, 55(1):111–119, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530488.

Liu:2003:TIC

[2359] Xinsheng Liu and Jinde Wang. Testing for increasing convex order in several populations. Annals of the Institute of Statistical Mathematics, 55(1):121–136, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530489.

Meintanis:2003:TFR

[2360] Simos Meintanis and George Iliopoulos. Tests of fit for the Rayleigh distribution based on the empirical Laplace transform. Annals of the Institute of Statistical Mathematics, 55(1):137–151, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530490.

Inoue:2003:GBN

[2361] Kiyoshi Inoue and Sigeo Aki. Generalized binomial and negative binomial distributions of order k by the l-overlapping enumeration scheme. Annals of the Institute of Statistical Mathematics, 55(1):153–167, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530491.

Vidoni:2003:PCG

[2362] Paolo Vidoni. Prediction and calibration in generalized linear models. Annals of the Institute of Statistical Mathematics, 55(1):169–185, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530492.

Sartori:2003:NLA

[2363] N. Sartori. A note on likelihood asymptotics in normal linear regression. Annals of the Institute of Statistical Mathematics, 55(1):187–195, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530493.

Lu:2003:LIM

[2364] Zudi Lu and Y. V. Hui. L_{1} linear interpolator for missing values in time series. Annals of the Institute of Statistical Mathematics, 55(1):197–216, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530494.

Bradley:2003:LBM

[2365] David M. Bradley and Ramesh C.

Gupta. Limiting behaviour of the mean residual life. Annals of the Institute of Statistical Mathematics, 55(1):217–226, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530495.

Anonymous:2003:HCa

[2366] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 55(1):??, March 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hobolth:2003:CPS

[2367] Asger Hobolth, Jan Pedersen, and Eva B. Vedel Jensen. A continuous parametric shape model. Annals of the Institute of Statistical Mathematics, 55(2):227–242, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530496.

Pan:2003:EOS

[2368] Wei Pan, Walter W. Piegorsch, and R. Webster West. Exact one-sided simultaneous confidence bands via Uusipaikka’s method. Annals of the Institute of Statistical Mathematics, 55 (2):243–250, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530497.

Gao:2003:PIC

[2369] Wei Gao and Ning-Zhong Shi. Iprojection onto isotonic cones and its applications to maximum likelihood estimation for log-linear models. Annals of the Institute of Statistical Mathematics, 55(2):251–263, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530498.

Croux:2003:BIR

[2370] Christophe Croux, Stefan Van Aelst, and Catherine Dehon. Bounded influence regression using high breakdown scatter matrices. Annals of the Institute of Statistical Mathematics, 55(2):265–285, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530499.

Kohler:2003:SCA

[2371] Michael Kohler, Adam Krzyzak, and Harro Walk. Strong consistency of automatic kernel regression estimates. Annals of the Institute of Statistical Mathematics, 55(2):287–308, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530500.

Tanaka:2003:FDA

[2372] Hidekazu Tanaka and Masafumi Akahira. On a family of distributions attaining the Bhattacharyya bound. Annals of the Institute of Statistical Mathematics, 55(2):309–317, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530501.

Childs:2003:ELI

[2373] A. Childs, B. Chandrasekar, N. Balakrishnan, and D. Kundu. Exact likelihood inference based on Type– I and Type–II hybrid censored samples from the exponential distribution. Annals of the Institute of Statistical Mathematics, 55(2):319–330, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530502.

Perez:2003:BCD

[2374] M. C. Iglesias Pérez and W. González Manteiga. Bootstrap for the conditional distribution function with truncated and censored data. Annals of the Institute of Statistical Mathematics, 55(2):331–357, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530503.

Kim:2003:BCS

[2375] Choongrak Kim, Byeong U. Park, Woochul Kim, and Chiyon Lim. Bézier curve smoothing of the Kaplan–Meier estimator. Annals of the Institute of Statistical Mathematics, 55(2):359–367, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530504.

Fukumizu:2003:P

[2376] Kenji Fukumizu, Yukito Iba, Takashi Tsuchiya, and Koji Tsuda. Preface. Annals of the Institute of Statistical Mathematics, 55(2):369–370, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02530505.

Bousquet:2003:NAS

[2377] Olivier Bousquet. New approaches to statistical learning theory. Annals of the Institute of Statistical Mathematics, 55(2):371–389, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530506.

Weston:2003:DLD

[2378] Jason Weston, Bernhard Schölkopf, Eleazar Eskin, Christina Leslie, and William Stafford Noble. Dealing with large diagonals in kernel matrices.

Annals of the Institute of Statistical Mathematics, 55(2):391–408, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530507.

Doucet:2003:PEG

[2379] Arnaud Doucet and Vladislav B. Tadić.

Parameter estimation in general statespace models using particle methods. Annals of the Institute of Statistical Mathematics, 55(2):409–422, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530508.

Guo:2003:NAD

[2380] Dong Guo, Xiaodong Wang, and Rong Chen. Nonparametric adaptive detection in fading channels based on sequential Monte Carlo and Bayesian model averaging. Annals of the Institute of Statistical Mathematics, 55 (2):423–436, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530509.

Tanaka-Yamawaki:2003:SMS

[2381] Mieko Tanaka-Yamawaki. Stability of Markovian structure observed in high frequency foreign exchange data. Annals of the Institute of Statistical Mathematics, 55(2):437–446, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530510.

Anonymous:2003:HCb

[2382] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 55(2):??, June 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Tsukuma:2003:EML

[2383] Hisayuki Tsukuma. On estimation in multivariate linear calibration with elliptical errors. Annals of the Institute of Statistical Mathematics, 55(3):447–466, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02517800.

Lee:2003:CST

[2384] Sangyeol Lee, Okyoung Na, and Seongryong Na. On the cusum of squares test for variance change in nonstationary and nonparametric time series models. Annals of the Institute of Statistical Mathematics, 55(3): 467–485, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02517801.

Carolan:2003:CLC

[2385] Chris Carolan and Richard Dykstra. Characterization of the least concave majorant of Brownian motion, conditional on a vertex point, with application to construction. Annals of the Institute of Statistical Mathematics, 55(3):487–497, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02517802.

Hu:2003:CED

[2386] Chin-Yuan Hu and Gwo Dong Lin.

Characterizations of the exponential distribution by stochastic ordering properties of the geometric compound. Annals of the Institute of Statistical Mathematics, 55(3):499–506, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02517803.

Hashorva:2003:MGT

[2387] Enkelejd Hashorva and Jürg Hüsler.

On multivariate Gaussian tails. Annals of the Institute of Statistical Mathematics, 55(3):507–522, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02517804.

Aly:2003:DU

[2388] Emad-Eldin A. A. Aly and Nadjib Bouzar. On discrete α-unimodality. Annals of the Institute of Statistical Mathematics, 55(3):523–535, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02517805.

Fujikoshi:2003:CVC

[2389] Yasunori Fujikoshi, Takafumi Noguchi, Megu Ohtaki, and Hirokazu Yanagihara. Corrected versions of crossvalidation criteria for selecting multivariate regression and growth curve models. Annals of the Institute of Statistical Mathematics, 55(3):537– 553, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02517806.

Gottlieb:2003:AEJ

[2390] Alex D. Gottlieb. Asymptotic equivalence of the jackknife and infinitesimal jackknife variance estimators for some smooth statistics. Annals of the Institute of Statistical Mathematics, 55 (3):555–561, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02517807.

Arcones:2003:AAB

[2391] Miguel A. Arcones. On the asymptotic accuracy of the bootstrap under arbitrary resampling size. Annals of the Institute of Statistical Mathematics, 55(3):563–583, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02517808.

Wang:2003:ELP

[2392] Qi-Hua Wang and Bing-Yi Jing. Empirical likelihood for partial linear models. Annals of the Institute of Statistical Mathematics, 55(3):585–595, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02517809.

Qin:2003:EEC

[2393] Gensheng Qin and Bing-Yi Jing. Edgeworth expansion in censored linear regression model. Annals of the Institute of Statistical Mathematics, 55(3): 597–617, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02517810.

Ghorai:2003:CLT

[2394] J. K. Ghorai. A central limit theorem for the L_{2} error of positive wavelet density estimator. Annals of the Institute of Statistical Mathematics, 55(3):619–637, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02517811.

Osterreicher:2003:NCM

[2395] Ferdinand Österreicher and Igor Vajda. A new class of metric divergences on probability spaces and its applicability in statistics. Annals of the Institute of Statistical Mathematics, 55(3):639–653, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02517812.

Huang:2003:OVC

[2396] Su-Yun Huang, Chuhsing Kate Hsiao, and Ching-Wei Chang. Optimal volume-corrected Laplace–Metropolis method. Annals of the Institute of Statistical Mathematics, 55(3):655– 670, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02517813.

Anonymous:2003:HCc

[2397] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 55(3):??, September 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Imoto:2003:SSP

[2398] Seiya Imoto and Sadanori Konishi. Selection of smoothing parameters in Bspline nonparametric regression models using information criteria. Annals of the Institute of Statistical Mathematics, 55(4):671–687, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02523388.

Zhou:2003:AEL

[2399] Wang Zhou and Bing-Yi Jing. Adjusted empirical likelihood method for quantiles. Annals of the Institute of Statistical Mathematics, 55(4): 689–703, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02523389.

Hidalgo:2003:SEL

[2400] J. Hidalgo and Y. Yajima. Semiparametric estimation of the longrange parameter. Annals of the Institute of Statistical Mathematics, 55 (4):705–736, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02523390.

Fryzlewicz:2003:FNS

[2401] Piotr Fryzlewicz, Sébastien Van Bellegem, and Rainer von Sachs. Forecasting non-stationary time series by wavelet process modelling. Annals of the Institute of Statistical Mathematics, 55(4):737–764, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523391.

Alonso:2003:RTS

[2402] Andrés M. Alonso, Daniel Peña, and Juan Romo. Resampling time series using missing values techniques. Annals of the Institute of Statistical Mathematics, 55(4):765–796, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523392.

Ristic:2003:BUA

[2403] Miroslav M. Ristić and Biljana C.

Popović. A bivariate uniform autoregressive process. Annals of the Institute of Statistical Mathematics, 55 (4):797–802, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02523393.

Fourdrinier:2003:BUE

[2404] Dominique Fourdrinier and William E.

Strawderman. On Bayes and unbiased estimators of loss. Annals of the Institute of Statistical Mathematics, 55 (4):803–816, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02523394.

Karlis:2003:MEM

[2405] Dimitris Karlis. ML estimation for multivariate shock models via an EM algorithm. Annals of the Institute of Statistical Mathematics, 55(4):817–830, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523395.

Taneichi:2003:IGF

[2406] Nobuhiro Taneichi, Yuri Sekiya, and Hideyuki Imai. Improvements of goodness-of-fit statistics for sparse multinomials based on normalizing transformations. Annals of the Institute of Statistical Mathematics, 55 (4):831–848, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02523396.

Bischoff:2003:EAB

[2407] Wolfgang Bischoff, Enkelejd Hashorva, Jürg Hüsler, and Frank Miller. Exact asymptotics for boundary crossings of the Brownian bridge with trend with application to the Kolmogorov test. Annals of the Institute of Statistical Mathematics, 55(4):849–864, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523397.

Antzoulakos:2003:DTN

[2408] D. L. Antzoulakos, S. Bersimis, and M. V. Koutras. On the distribution of the total number of run lengths. Annals of the Institute of Statistical Mathematics, 55(4):865–884, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523398.

Mahmoud:2003:OSV

[2409] Hosam M. Mahmoud. One-sided variations on binary search trees. Annals of the Institute of Statistical Mathematics, 55(4):885–900, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523399.

Anonymous:2003:A

[2410] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 55(4):901–903, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02523400.

Anonymous:2003:HCd

[2411] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 55(4):??, December 2003. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takemura:2004:SCM

[2412] Akimichi Takemura and Satoshi Aoki.

Some characterizations of minimal Markov basis for sampling from discrete conditional distributions. Annals of the Institute of Statistical Mathematics, 56(1):1–17, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530522.

Delaigle:2004:BBS

[2413] A. Delaigle and I. Gijbels. Bootstrap bandwidth selection in kernel density estimation from a contaminated sample. Annals of the Institute of Statistical Mathematics, 56(1):19–47, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530523.

Honda:2004:NRC

[2414] Toshio Honda. Nonparametric regression with current status data. Annals of the Institute of Statistical Mathematics, 56(1):49–72, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530524.

Peng:2004:NRU

[2415] Liang Peng and Qiwei Yao. Nonparametric regression under dependent errors with infinite variance. Annals of the Institute of Statistical Mathematics, 56(1):73–86, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530525.

Chiang:2004:SER

[2416] Chin-Tsang Chiang and Mei-Cheng Wang. Smoothing estimation of rate function for recurrent event data with informative censoring. Annals of the Institute of Statistical Mathematics, 56(1):87–100, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530526.

Sheena:2004:EEN

[2417] Yo Sheena, A. K. Gupta, and Y. Fujikoshi. Estimation of the eigenvalues of noncentrality parameter in matrix variate noncentral beta distribution. Annals of the Institute of Statistical Mathematics, 56(1):101–125, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530527.

Wu:2004:BEC

[2418] Yanhong Wu. Bias of estimator of change point detected by a CUSUM procedure. Annals of the Institute of Statistical Mathematics, 56(1):127–142, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530528.

Inoue:2004:JDA

[2419] Kiyoshi Inoue. Joint distributions associated with patterns, successes and failures in a sequence of multi-state trials. Annals of the Institute of Statistical Mathematics, 56(1):143–168, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530529.

Aki:2004:WTP

[2420] Sigeo Aki and Katuomi Hirano.

Waiting time problems for a twodimensional pattern. Annals of the Institute of Statistical Mathematics, 56 (1):169–182, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530530.

Cramer:2004:UUG

[2421] Erhard Cramer, Udo Kamps, and Tomasz Rychlik. Unimodality of uniform generalized order statistics, with applications to mean bounds. Annals of the Institute of Statistical Mathematics, 56(1):183–192, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530531.

Chang:2004:SSR

[2422] Yen-Chang Chang. A sequential software release policy. Annals of the Institute of Statistical Mathematics, 56 (1):193–204, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530532.

Anonymous:2004:HCa

[2423] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 56(1):??, March 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Efromovich:2004:ABS

[2424] Sam Efromovich. Analysis of blockwise shrinkage wavelet estimates via lower bounds for no-signal setting. Annals of the Institute of Statistical Mathematics, 56(2):205–223, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530542.

Lahiri:2004:ADE

[2425] S. N. Lahiri and Kanchan Mukherjee. Asymptotic distributions of M-estimators in a spatial regression model under some fixed and stochastic spatial sampling designs.

Annals of the Institute of Statistical Mathematics, 56(2):225–250, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530543.

Wang:2004:AEC

[2426] Lihong Wang. Asymptotics of estimates in constrained nonlinear regression with long-range dependent innovations. Annals of the Institute of Statistical Mathematics, 56(2):251–264, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530544.

Yatracos:2004:DDR

[2427] Yannis Yatracos. Dependence and the dimensionality reduction principle. Annals of the Institute of Statistical Mathematics, 56(2):265–277, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530545.

Dauxois:2004:LRC

[2428] Jacques Dauxois, Guy Martial Nkiet, and Yves Romain. Linear relative canonical analysis of Euclidean random variables, asymptotic study and some applications. Annals of the Institute of Statistical Mathematics, 56(2):279–304, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530546.

Gupta:2004:CMS

[2429] Arjun K. Gupta and John T. Chen.

A class of multivariate skew-normal models. Annals of the Institute of Statistical Mathematics, 56(2):305–315, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530547.

Sarkar:2004:WTD

[2430] Anish Sarkar, Kanwar Sen, and Anuradha. Waiting time distributions of runs in higher order Markov chains. Annals of the Institute of Statistical Mathematics, 56(2):317–349, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530548.

Gupta:2004:CSN

[2431] Arjun K. Gupta, Truc T. Nguyen, and Jose Almer T. Sanqui. Characterization of the skew-normal distribution. Annals of the Institute of Statistical Mathematics, 56(2):351–360, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530549.

Navarro:2004:CMN

[2432] Jorge Navarro and Jose M. Ruiz. A characterization of the multivariate normal distribution by using the hazard gradient. Annals of the Institute of Statistical Mathematics, 56(2):361–367, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530550.

Gibilisco:2004:CPM

[2433] Paolo Gibilisco and Tommaso Isola. On the characterisation of paired monotone metrics. Annals of the Institute of Statistical Mathematics, 56(2):369–381, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530551.

Hofmann:2004:FIR

[2434] Glenn Hofmann and N. Balakrishnan. Fisher information in k-records. Annals of the Institute of Statistical Mathematics, 56(2):383–396, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530552.

Maehara:2004:WDU

[2435] H. Maehara. When does the union of random spherical caps become connected? Annals of the Institute of Statistical Mathematics, 56(2):397–402, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530553.

Anonymous:2004:HCb

[2436] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 56(2):??, June 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Wang:2004:LBI

[2437] Qi-Hua Wang. Likelihood-based imputation inference for mean functionals in the presence of missing responses. Annals of the Institute of Statistical Mathematics, 56(3):403–414, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530533.

Peruggia:2004:DSW

[2438] Mario Peruggia, Thomas J. Santner, and Yu-Yun Ho. Detecting stage-wise outliers in hierarchical Bayesian linear models of repeated measures data. Annals of the Institute of Statistical Mathematics, 56(3):415–433, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530534.

Ebrahimi:2004:IAB

[2439] Nader Ebrahimi. Indirect assessment of the bivariate survival function. Annals of the Institute of Statistical Mathematics, 56(3):435–448, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02530535.

Comte:2004:NAF

[2440] F. Comte and Y. Rozenholc. A new algorithm for fixed design regression and denoising. Annals of the Institute of Statistical Mathematics, 56(3): 449–473, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530536.

Cristobal:2004:CBN

[2441] J. A. Cristóbal, J. L. Ojeda, and J. T. Alcalá. Confidence bands in nonparametric regression with length biased data. Annals of the Institute of Statistical Mathematics, 56(3): 475–496, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530537.

Bouzar:2004:DSS

[2442] Nadjib Bouzar. Discrete semi-stable distributions. Annals of the Institute of Statistical Mathematics, 56(3): 497–510, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02530538.

Pace:2004:TLE

[2443] Luigi Pace and Alessandra Salvan.

Tensors and likelihood expansions in the presence of nuisance parameters. Annals of the Institute of Statistical Mathematics, 56(3):511–528, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530539.

Nandi:2004:APL

[2444] Swagata Nandi and Debasis Kundu.

Asymptotic properties of the least squares estimators of the parameters of the chirp signals. Annals of the Institute of Statistical Mathematics, 56(3):529–544, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02530540.

Sakamoto:2004:AEF

[2445] Yuji Sakamoto and Nakahiro Yoshida.

Asymptotic expansion formulas for functionals of ϵ-Markov processes with a mixing property. Annals of the Institute of Statistical Mathematics, 56 (3):545–597, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02530541.

Anonymous:2004:HCc

[2446] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 56(3):??, September 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Fan:2004:BAS

[2447] Tsai-Hung Fan and Eng-Nan Kuo. A Bayesian analysis for the seismic data on Taiwan. Annals of the Institute of Statistical Mathematics, 56 (4):599–609, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02506478.

Jiang:2004:SDE

[2448] Jiancheng Jiang and Y. V. Hui. Spectral density estimation with amplitude modulation and outlier detection. Annals of the Institute of Statistical Mathematics, 56(4):611–630, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506479.

Seidel:2004:TLM

[2449] Wilfried Seidel and Hana Sevcíková. Types of likelihood maxima in mixture models and their implication on the performance of tests. Annals of the Institute of Statistical Mathematics, 56(4):631–654, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506480.

Adam:2004:CCM

[2450] Maria Adam and John Maroulas.

Canonical correlations in multi-way layout. Annals of the Institute of Statistical Mathematics, 56(4):655–666, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506481.

DeUna-alvarez:2004:NEU

[2451] Jacobo De Uña-Álvarez. Nonparametric estimation under length-biased sampling and Type I censoring: A moment based approach. Annals of the Institute of Statistical Mathematics, 56 (4):667–681, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506482.

Gomes:2004:AWL

[2452] Antonio Eduardo Gomes. Asymptotics for a weighted least squares estimator of the disease onset distribution function for a survival-sacrifice model. Annals of the Institute of Statistical Mathematics, 56(4):683–700, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506483.

Ozturk:2004:ORR

[2453] Omer Ozturk and Steven N. MacEachern. Order restricted randomized designs for control versus treatment comparison. Annals of the Institute of Statistical Mathematics, 56(4):

701–720, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506484.

Gupta:2004:SCR

[2454] Ramesh C. Gupta and Mohammad Ahsanullah. Some characterization results based on the conditional expectation of a function of non-adjacent order statistic (record value). Annals of the Institute of Statistical Mathematics, 56(4):721–732, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506485.

Jones:2004:MDD

[2455] M. C. Jones. The Möbius distribution on the disc. Annals of the Institute of Statistical Mathematics, 56 (4):733–742, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506486.

Buchmann:2004:DPR

[2456] Boris Buchmann and Rudolf Grübel. Decompounding Poisson random sums: Recursively truncated estimates in the discrete case. Annals of the Institute of Statistical Mathematics, 56(4): 743–756, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506487.

Sheena:2004:NED

[2457] Yo Sheena and Arjun K. Gupta. New estimators of discriminant coefficients as the gradient of log-odds. Annals of the Institute of Statistical Mathematics, 56(4):757–770, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506488.

Chen:2004:AAA

[2458] Zhiqiang Chen and Evarist Giné. Another approach to asymptotics and bootstrap of randomly trimmed means. Annals of the Institute of Statistical Mathematics, 56(4):771–790, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506489.

Vidal-Sanz:2004:UCD

[2459] Jose M. Vidal-Sanz and Miguel A.

Delgado. Universal consistency of delta estimators. Annals of the Institute of Statistical Mathematics, 56 (4):791–818, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506490.

Harel:2004:UCC

[2460] Michel Harel and Madan L. Puri.

Universally consistent conditional Ustatistics for absolutely regular processes and its applications for hidden Markov models. Annals of the Institute of Statistical Mathematics, 56 (4):819–832, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506491.

Anonymous:2004:A

[2461] Anonymous. Acknowledgement. Annals of the Institute of Statistical Mathematics, 56(4):833–835, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506492.

Anonymous:2004:HCd

[2462] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 56(4):??, December 2004. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Tanaka:2005:SCM

[2463] Kentaro Tanaka and Akimichi Takemura. Strong consistency of MLE for finite uniform mixtures when the scale parameters are exponentially small. Annals of the Institute of Statistical Mathematics, 57(1):1–19, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506875.

Pozdnyakov:2005:MAS

[2464] Vladimir Pozdnyakov, Joseph Glaz, Martin Kulldorff, and J. Michael Steele. A martingale approach to scan statistics. Annals of the Institute of Statistical Mathematics, 57(1):21–37, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02506876.

Kirmani:2005:RSS

[2465] S. Kirmani and J. Wesolowski. Regressions for sums of squares of spacings. Annals of the Institute of Statistical Mathematics, 57(1):39–47, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506877.

Inoue:2005:GPU

[2466] Kiyoshi Inoue and Sigeo Aki. A generalized Pólya urn model and related multivariate distributions. Annals of the Institute of Statistical Mathematics, 57(1):49–59, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506878.

Tse:2005:QPL

[2467] Szeman Tse. Quantile process for left truncated and right censored data. Annals of the Institute of Statistical Mathematics, 57(1):61–69, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506879.

Mrkvicka:2005:OLU

[2468] Tomás Mrkvicka and Ilya Molchanov.

Optimisation of linear unbiased intensity estimators for point processes. Annals of the Institute of Statistical Mathematics, 57(1):71–81, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506880.

Wong:2005:JMC

[2469] Heung Wong, W. K. Li, and Shiqing Ling. Joint modeling of cointegration and conditional heteroscedasticity with applications. Annals of the Institute of Statistical Mathematics, 57(1):83–103, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506881.

Kim:2005:TIT

[2470] Eunhee Kim and Sangyeol Lee. A test for independence of two stationary infinite order autoregressive processes. Annals of the Institute of Statistical Mathematics, 57(1):105–127, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506882.

Marchand:2005:IMR

[2471] Éric Marchand and William E. Strawderman. Improving on the minimum risk equivariant estimator of a location parameter which is constrained to an interval or a halfinterval. Annals of the Institute of Statistical Mathematics, 57(1):129–143, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506883.

Maruyama:2005:SMA

[2472] Yuzo Maruyama and Katsunori Iwasaki. Sensitivity of minimaxity and admissibility in the estimation of a positive normal mean. Annals of the Institute of Statistical Mathematics, 57 (1):145–156, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02506884.

Maruyama:2005:NCD

[2473] Yuzo Maruyama and William E. Strawderman. Necessary conditions for dominating the James–Stein estimator. Annals of the Institute of Statistical Mathematics, 57(1):157–165, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506885.

Xiao:2005:MCB

[2474] Yushan Xiao, Yoshikazu Takada, and Ningzhong Shi. Minimax confidence bound of the normal mean under an asymmetric loss function. Annals of the Institute of Statistical Mathematics, 57(1):167–182, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506886.

Matsui:2005:ECF

[2475] Muneya Matsui and Akimichi Takemura. Empirical characteristic function approach to goodness-of-fit tests for the Cauchy distribution with parameters estimated by MLE or EISE. Annals of the Institute of Statistical Mathematics, 57(1):183–199, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02506887.

Anonymous:2005:HCa

[2476] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 57(1):??, March 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Beran:2005:AFM

[2477] Rodolf Beran. ASP fits to multi-way layouts. Annals of the Institute of Statistical Mathematics, 57(2):201–220, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507022.

Ma:2005:SST

[2478] Chunsheng Ma. Semiparametric spatio-temporal covariance models with the ARMA temporal margin. Annals of the Institute of Statistical Mathematics, 57(2):221–233, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507023.

Delouille:2005:ENA

[2479] Véronique Delouille and Rainer von Sachs. Estimation of nonlinear autoregressive models using design-adapted wavelets. Annals of the Institute of Statistical Mathematics, 57(2):235–253, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507024.

Dixon:2005:FIS

[2480] John R. Dixon, Michael R. Kosorok, and Bee Leng Lee. Functional inference in semiparametric models using the piggyback bootstrap. Annals of the Institute of Statistical Mathematics, 57(2):255–277, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02507025.

Cheung:2005:VES

[2481] K. Y. Cheung and Stephen M. S.

Lee. Variance estimation for sample quantiles using the m out of n bootstrap. Annals of the Institute of Statistical Mathematics, 57(2):279–290, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02507026.

Lachout:2005:SCE

[2482] Petr Lachout, Eckhard Liebscher, and Silvia Vogel. Strong convergence of estimators as ϵ_{n}-minimisers of optimisation problems of optimisation problems. Annals of the Institute of Statistical Mathematics, 57(2):291–313, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02507027.

Bochkina:2005:PME

[2483] Natalia Bochkina and Theofanis Sapatinas. On the posterior median estimators of possibly sparse sequences. Annals of the Institute of Statistical Mathematics, 57(2):315–351, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02507028.

Inoue:2005:JDN

[2484] Kiyoshi Inoue and Sigeo Aki. Joint distributions of numbers of success runs of specified lengths in linear and circular sequences. Annals of the Institute of Statistical Mathematics, 57(2):353–368, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507029.

Hoshino:2005:EEN

[2485] Nobuaki Hoshino. Engen’s extended negative binomial model revisited. Annals of the Institute of Statistical Mathematics, 57(2):369–387, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507030.

Genton:2005:GSE

[2486] Marc G. Genton and Nicola M. R. Loperfido. Generalized skew-elliptical distributions and their quadratic forms. Annals of the Institute of Statistical Mathematics, 57(2):389–401, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02507031.

Anonymous:2005:HCb

[2487] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 57(2):??, June 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Honda:2005:EAC

[2488] Toshio Honda. Estimation in additive Cox models by marginal integration. Annals of the Institute of Statistical Mathematics, 57(3):403–423, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02509232.

Fernandes:2005:CLT

[2489] Marcelo Fernandes and Paulo Klinger Monteiro. Central limit theorem for asymmetric kernel functionals. Annals of the Institute of Statistical Mathematics, 57(3):425–442, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02509233.

Campos:2005:KES

[2490] Viviane S. M. Campos and Chang C. Y.

Dorea. Kernel estimation for stationary density of Markov chains with general state space. Annals of the Institute of Statistical Mathematics, 57(3): 443–453, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02509234.

Sun:2005:EMN

[2491] Dongchu Sun and Xiaoqian Sun. Estimation of the multivariate normal precision and covariance matrices in a star-shape model. Annals of the Institute of Statistical Mathematics, 57 (3):455–484, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02509235.

Lin:2005:PEL

[2492] Lu Lin, Lixing Zhu, and K. C. Yuen.

Profile empirical likelihood for parametric and semiparametric models. Annals of the Institute of Statistical Mathematics, 57(3):485–505, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02509236.

Janssen:2005:RST

[2493] Arnold Janssen. Resampling Student’s t-type statistics. Annals of the Institute of Statistical Mathematics, 57(3): 507–529, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02509237.

Han:2005:CCG

[2494] Dong Han and Fugee Tsung. Comparison of the cuscore, GLRT and cusum control charts for detecting a dynamic mean change. Annals of the Institute of Statistical Mathematics, 57(3): 531–552, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02509238.

Lee:2005:TPC

[2495] Sangyeol Lee and Okyoung Na. Test for parameter change based on the estimator minimizing density-based divergence measures. Annals of the Institute of Statistical Mathematics, 57 (3):553–573, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02509239.

Duchesne:2005:TSC

[2496] Pierre Duchesne. Testing for serial correlation of unknown form in cointegrated time series models. Annals of the Institute of Statistical Mathematics, 57(3):575–595, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02509240.

Huillet:2005:UOS

[2497] Thierry Huillet. Unordered and ordered sample from Dirichlet distribution. Annals of the Institute of Statistical Mathematics, 57(3):597– 616, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02509241.

Anonymous:2005:HCc

[2498] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 57(3):??, September 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nonaka:2005:NRM

[2499] Yoshisuke Nonaka and Sadanori Konishi. Nonlinear regression modeling using regularized local likelihood method. Annals of the Institute of Statistical Mathematics, 57(4):617–635, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915429.

Chiang:2005:CBS

[2500] Chin-Tsang Chiang. Comparisons between simultaneous and componentwise splines for varying coefficient models. Annals of the Institute of Statistical Mathematics, 57(4):637–653, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915430.

Henna:2005:ENC

[2501] Jogi Henna. Estimation of the number of components of finite mixtures of multivariate distributions. Annals of the Institute of Statistical Mathematics, 57(4):655–664, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915431.

Walk:2005:SUC

[2502] Harro Walk. Strong universal consistency of smooth kernel regression estimates. Annals of the Institute of Statistical Mathematics, 57(4):665–685, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915432.

Wang:2005:DMD

[2503] Xiaogang Wang and James V. Zidek. Derivation of mixture distributions and weighted likelihood function as minimizers of KL-divergence subject to constraints. Annals of the Institute of Statistical Mathematics, 57(4): 687–701, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915433.

Bandyopadhyay:2005:NSA

[2504] Nibedita Bandyopadhyay and Ananda Sen. Non-standard asymptotics in an inhomogeneous gamma process. Annals of the Institute of Statistical Mathematics, 57(4):703–732, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915434.

So:2005:NIV

[2505] Beong Soo So. A new instrumental variable estimation for diffusion processes. Annals of the Institute of Statistical Mathematics, 57(4):733–745, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915435.

Kawczak:2005:EDF

[2506] Janusz Kawczak, Reg Kulperger, and Hao Yu. The empirical distribution function and partial sum process of residuals from a stationary arch with drift process. Annals of the Institute of Statistical Mathematics, 57(4):747–765, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915436.

Pelletier:2005:IBS

[2507] Bruno Pelletier. Informative barycentres in statistics. Annals of the Institute of Statistical Mathematics, 57 (4):767–780, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915437.

Ubeda-Flores:2005:MVB

[2508] Manuel Úbeda-Flores. Multivariate versions of Blomqvist’s beta and Spearman’s footrule. Annals of the Institute of Statistical Mathematics, 57(4): 781–788, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915438.

Shieh:2005:IBB

[2509] Grace S. Shieh and Richard A. Johnson. Inferences based on a bivariate distribution with von Mises marginals. Annals of the Institute of Statistical Mathematics, 57(4):789–802, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915439.

Belzunce:2005:TDO

[2510] F. Belzunce, J. F. Pinar, and J. M. Ruiz. On testing the dilation order and HNBUE alternatives. Annals of the Institute of Statistical Mathematics, 57(4):803–815, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915440.

Alaya:2005:FEG

[2511] Mohamed Ben Alaya and Thierry Huillet. On a functional equation generalizing the class of semistable distributions. Annals of the Institute of Statistical Mathematics, 57(4):817–831, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915441.

Chang:2005:ODW

[2512] Fu-Chuen Chang. D-optimal designs for weighted polynomial regression — a functional approach. Annals of the Institute of Statistical Mathematics, 57(4):833–844, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915442.

Anonymous:2005:HCd

[2513] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 57(4):??, December 2005. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Chakrabarti:2006:OAIa

[2514] Arijit Chakrabarti and Jayanta K.

Ghosh. Optimality of AIC in inference about Brownian motion. Annals of the Institute of Statistical Mathematics, 58(1):1–20, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0007-7.

Arcones:2006:LDE

[2515] Miguel A. Arcones. Large deviations for M-estimators. Annals of the Institute of Statistical Mathematics, 58(1):21–52, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0017-5.

Ghosh:2006:ECG

[2516] Sucharita Ghosh and Jan Beran. On estimating the cumulant generating function of linear processes. Annals of the Institute of Statistical Mathematics, 58(1):53–71, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0009-5.

Fourdrinier:2006:ELP

[2517] Dominique Fourdrinier, William E. Strawderman, and Martin T. Wells. Estimation of a location parameter with restrictions or “vague information” for spherically symmetric distributions. Annals of the Institute of Statistical Mathematics, 58(1):73–92, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0001-0.

Sugiura:2006:BUT

[2518] Nariaki Sugiura, Hidetoshi Murakami, Seong Keon Lee, and Yasutomo Maeda. Biased and unbiased twosided Wilcoxon tests for equal sample sizes. Annals of the Institute of Statistical Mathematics, 58(1):93–100, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0019-3.

Gupta:2006:STC

[2519] Arjun K. Gupta and Jin Xu. On some tests of the covariance matrix under general conditions. Annals of the Institute of Statistical Mathematics, 58 (1):101–114, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0010-z.

Valenca:2006:THWa

[2520] Dione M. Valença and Heleno Bolfarine. Testing homogeneity in Weibull error in variables models. Annals of the Institute of Statistical Mathematics, 58(1):115–129, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0006-8.

Hamers:2006:NBE

[2521] Michael Hamers and Michael Kohler.

Nonasymptotic bounds on the L_{2} error of neural network regression estimates. Annals of the Institute of Statistical Mathematics, 58(1):131–151, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0005-9.

Roy:2006:FMG

[2522] Surupa Roy and Tathagata Banerjee.

A flexible model for generalized linear regression with measurement error. Annals of the Institute of Statistical Mathematics, 58(1):153–169, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0002-z.

Balaji:2006:ELDa

[2523] Srinivasan Balaji and Hosam M. Mahmoud. Exact and limiting distributions in diagonal Pólya processes. Annals of the Institute of Statistical Mathematics, 58(1):171–185, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0012-x.

Cekanavicius:2006:CBAa

[2524] Vydas Cekanavicius and Bero Roos.

Compound binomial approximations.

Annals of the Institute of Statistical Mathematics, 58(1):187–210, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-005-
                                        0018-4.

Anonymous:2006:HCa

[2525] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 58(1):??, March 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Lee:2006:TPC

[2526] Sangyeol Lee, Yoichi Nishiyama, and Nakahiro Yoshida. Test for parameter change in diffusion processes by Cusum statistics based on one-step estimators. Annals of the Institute of Statistical Mathematics, 58(2):211–222, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0037-9.

Schoenberg:2006:NSM

[2527] Frederic Paik Schoenberg. On nonsimple marked point processes. Annals of the Institute of Statistical Mathematics, 58(2):223–233, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-005-0003-y.

vanLieshout:2006:FMP

[2528] M. N. M. van Lieshout. A J-function for marked point patterns. Annals of the Institute of Statistical Mathematics, 58(2):235–259, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0015-7.

Falk:2006:TTI

[2529] Michael Falk and René Michel. Testing for tail independence in extreme value models. Annals of the Institute of Statistical Mathematics, 58(2):261–290, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0016-6.

Fu:2006:WTD

[2530] James C. Fu and W. Y. Wendy Lou.

Waiting time distributions of simple and compound patterns in a sequence of r-th order Markov dependent multistate trials. Annals of the Institute of Statistical Mathematics, 58(2):291–310, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0038-8.

Baba:2006:MC

[2531] Kunihiro Baba and Ritei Shibata. Multiplicative correlations. Annals of the Institute of Statistical Mathematics, 58(2):311–326, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0036-x.

Lee:2006:EKL

[2532] Young Kyung Lee and Byeong U. Park.

Estimation of Kullback–Leibler divergence by local likelihood. Annals of the Institute of Statistical Mathematics, 58(2):327–340, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0014-8.

Basu:2006:REP

[2533] Srabashi Basu, Ayanendranath Basu, and M. C. Jones. Robust and efficient parametric estimation for censored survival data. Annals of the Institute of Statistical Mathematics, 58(2):341–355, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0004-x.

Ould-Said:2006:APN

[2534] Elias Ould-Saïd and Mohamed Lemdani. Asymptotic properties of a nonparametric regression function estimator with randomly truncated data. Annals of the Institute of Statistical Mathematics, 58(2):357–378, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0011-y.

Yatracos:2006:CSP

[2535] Yannis G. Yatracos. On consistent statistical procedures in regression. Annals of the Institute of Statistical Mathematics, 58(2):379–387, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0021-9.

Martins-Filho:2006:NUS

[2536] Carlos Martins-Filho and Feng Yao. A note on the use of V and U statistics in nonparametric models of regression. Annals of the Institute of Statistical Mathematics, 58(2):389–406, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0034-z.

Dette:2006:LOD

[2537] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Local cand E-optimal designs for exponential regression models. Annals of the Institute of Statistical Mathematics, 58(2):407–426, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0031-2.

Anonymous:2006:HCb

[2538] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 58(2):??, June 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Chang:2006:ARG

[2539] In Hong Chang and Rahul Mukerjee.

Asymptotic results on a general class of empirical statistics: Power and confidence interval properties. Annals of the Institute of Statistical Mathematics, 58 (3):427–440, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0040-1.

Tsao:2006:CET

[2540] C. Andy. Tsao and Yu-Ling Tseng. Confidence estimation for tolerance intervals. Annals of the Institute of Statistical Mathematics, 58(3):441– 456, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0008-6.

Ghosh:2006:LBI

[2541] Malay Ghosh, Gauri Sankar Datta, Dalho Kim, and Trevor J. Sweeting. Likelihood-based inference for the ratios of regression coefficients in linear models. Annals of the Institute of Statistical Mathematics, 58(3): 457–473, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0027-3.

Fokianos:2006:EMD

[2542] Konstantinos Fokianos and Irene Kaimi. On the effect of misspecifying the density ratio model. Annals of the Institute of Statistical Mathematics, 58 (3):475–497, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0022-8.

Wu:2006:ITA

[2543] Changchun Wu and Runchu Zhang. An information-theoretic approach to the effective usage of auxiliary information from survey data. Annals of the Institute of Statistical Mathematics, 58 (3):499–509, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0013-9.

Ma:2006:APE

[2544] Shuangge Ma and Michael R. Kosorok.

Adaptive penalized M-estimation with current status data. Annals of the Institute of Statistical Mathematics, 58 (3):511–526, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-005-0026-4.

Al-Jarallah:2006:CPS

[2545] R. A. Al-Jarallah, A. R. Soltani, and N. A. Al-Kandari. On continuity of the Pearson statistic and sample quantiles. Annals of the Institute of Statistical Mathematics, 58(3):527–535, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0028-2.

Kotwal:2006:JDR

[2546] K. S. Kotwal and R. L. Shinde. Joint distributions of runs in a sequence of higher-order two-state Markov trials. Annals of the Institute of Statistical Mathematics, 58(3):537–554, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0024-6.

Kozubowski:2006:SLD

[2547] Tomasz J. Kozubowski and Seidu Inusah. A skew Laplace distribution on integers. Annals of the Institute of Statistical Mathematics, 58 (3):555–571, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-005-0029-1.

Kamae:2006:SPR

[2548] Teturo Kamae and Hayato Takahashi. Statistical problems related to irrational rotations. Annals of the Institute of Statistical Mathematics, 58(3): 573–593, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-005-0023-7.

Balakrishnan:2006:CTW

[2549] N. Balakrishnan and Po Yang. Classification of three-word indicator functions of two-level factorial designs. Annals of the Institute of Statistical Mathematics, 58(3):595–608, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0033-0.

Balakrishnan:2006:CBR

[2550] N. Balakrishnan and Po Yang. Connections between the resolutions of general two-level factorial designs. Annals of the Institute of Statistical Mathematics, 58(3):609–618, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-005-
                                        0020-x.

Abraham:2006:KRF

[2551] C. Abraham, G. Biau, and B. Cadre. On the kernel rule for function classification. Annals of the Institute of Statistical Mathematics, 58(3): 619–633, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0032-1.

Chanda:2006:SPS

[2552] Kamal C. Chanda. Sampling properties of U-statistics for a class of stationary nonlinear processes. Annals of the Institute of Statistical Mathematics, 58 (3):635–646, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0030-3.

Anonymous:2006:HCc

[2553] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 58(3):??, September 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Sugimoto:2006:PEB

[2554] Tomoyuki Sugimoto and Toshimitsu Hamasaki. Properties of estimators of baseline hazard functions in a semiparametric cure model. Annals of the Institute of Statistical Mathematics, 58 (4):647–674, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0041-0.

Tse:2006:LCT

[2555] Sze-Man Tse. Lorenz curve for truncated and censored data. Annals of the Institute of Statistical Mathematics, 58 (4):675–686, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0039-7.

Zhang:2006:SML

[2556] Zhiwei Zhang and Howard E. Rockette. Semiparametric maximum likelihood for missing covariates in parametric regression. Annals of the Institute of Statistical Mathematics, 58 (4):687–706, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0047-7.

Zhang:2006:PEL

[2557] Biao Zhang. A partial empirical likelihood based score test under a semiparametric finite mixture model. Annals of the Institute of Statistical Mathematics, 58(4):707–719, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0043-y.

Betro:2006:GMT

[2558] Bruno Betrò, Antonella Bodini, and Alessandra Guglielmi. Generalized moment theory and Bayesian robustness analysis for hierarchical mixture models. Annals of the Institute of Statistical Mathematics, 58(4):721–738, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0046-8.

Oono:2006:EEV

[2559] Youhei Oono and Nobuo Shinozaki. Estimation of error variance in ANOVA model and order restricted scale parameters. Annals of the Institute of Statistical Mathematics, 58(4):739–756, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-005-
0025-5.

Balakrishnan:2006:CIQ

[2560] N. Balakrishnan and T. Li. Confidence intervals for quantiles and tolerance intervals based on ordered ranked set samples. Annals of the Institute of Statistical Mathematics, 58(4): 757–777, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0035-y.

Antoniadis:2006:WBE

[2561] Anestis Antoniadis, Andrey Feuerverger, and Paulo Gonçalves. Wavelet-based estimation for univariate stable laws. Annals of the Institute of Statistical Mathematics, 58(4):779–807, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0042-z.

Valenca:2006:THWb

[2562] Dione M. Valença and Heleno Bolfarine. Testing homogeneity in Weibull error in variables models. Annals of the Institute of Statistical Mathematics, 58(4):809, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-006-0091-
                                        3.

Chakrabarti:2006:OAIb

[2563] Arijit Chakrabarti and Jayanta K. Ghosh. Optimality of AIC in inference about Brownian motion. Annals of the Institute of Statistical Mathematics, 58(4):811, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-006-0092-
                                        2.

Balaji:2006:ELDb

[2564] Srinivasan Balaji and Hosam M. Mahmoud. Exact and limiting distributions in diagonal Pólya processes. Annals of the Institute of Statistical Mathematics, 58(4):813, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-006-0093-
                                        1.

Cekanavicius:2006:CBAb

[2565] Vydas Cekanavicius and Bero Roos. Compound binomial approximations. Annals of the Institute of Statistical Mathematics, 58(4):815, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10463-006-0094-0.

Anonymous:2006:HCd

[2566] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 58(4):??, December 2006. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Amari:2007:P

[2567] Shun ichi Amari and Shiro Ikeda. Preface. Annals of the Institute of Statistical Mathematics, 59(1):1–2, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-006-0104-2.

Fujimoto:2007:MEA

[2568] Yu Fujimoto and Noboru Murata.

A modified EM algorithm for mixture models based on Bregman divergence. Annals of the Institute of Statistical Mathematics, 59(1):3–25, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0097-x.

Cena:2007:ESM

[2569] Alberto Cena and Giovanni Pistone. Exponential statistical manifold. Annals of the Institute of Statistical Mathematics, 59(1):27–56, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0096-y.

Kawanabe:2007:NAN

[2570] Motoaki Kawanabe, Masashi Sugiyama, Gilles Blanchard, and Klaus-Robert Müller. A new algorithm of non-Gaussian component analysis with radial kernel functions. Annals of the Institute of Statistical Mathematics, 59(1):57–75, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0098-9.

Dawid:2007:GPS

[2571] A. P. Dawid. The geometry of proper scoring rules. Annals of the Institute of Statistical Mathematics, 59(1):77–93, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0099-8.

Marriott:2007:ELM

[2572] Paul Marriott. Extending local mixture models. Annals of the Institute of Statistical Mathematics, 59(1):95–110, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0100-6.

Anaya-Izquierdo:2007:LME

[2573] K. A. Anaya-Izquierdo and P. K. Marriott. Local mixtures of the exponential distribution. Annals of the Institute of Statistical Mathematics, 59 (1):111–134, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0095-z.

Komaki:2007:BPB

[2574] Fumiyasu Komaki. Bayesian prediction based on a class of shrinkage priors for location-scale models. Annals of the Institute of Statistical Mathematics, 59(1):135–146, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0102-4.

Gibilisco:2007:UPQ

[2575] Paolo Gibilisco and Tommaso Isola.

Uncertainty principle and quantum Fisher information. Annals of the Institute of Statistical Mathematics, 59 (1):147–159, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0103-3.

Zhang:2007:NCC

[2576] Jun Zhang. A note on curvature of α-connections of a statistical manifold. Annals of the Institute of Statistical Mathematics, 59(1):161–170, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0105-1.

Anonymous:2007:HCa

[2577] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 59(1):??, March 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Beran:2007:MPR

[2578] Rudolf Beran. Multiple penalty regression: Fitting and extrapolating a discrete incomplete multi-way layout. Annals of the Institute of Statistical Mathematics, 59(2):171–195, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0050-z.

Banerjee:2007:TAR

[2579] Tathagata Banerjee and Atanu Biswas. Testing for the absence of random effects in a two-way nested design with mixed effects model: A nonparametric approach. Annals of the Institute of Statistical Mathematics, 59(2):197–210, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0049-5.

Sun:2007:EMN

[2580] Xiaoqian Sun and Dongchu Sun. Estimation of a multivariate normal covariance matrix with staircase pattern data. Annals of the Institute of Statistical Mathematics, 59(2):211–233, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0044-x.

Gijbels:2007:JPR

[2581] Irène Gijbels, Alexandre Lambert, and Peihua Qiu. Jump-preserving regression and smoothing using local linear fitting: A compromise. Annals of the Institute of Statistical Mathematics, 59(2):235–272, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0045-9.

Heuchenne:2007:PRC

[2582] Cédric Heuchenne and Ingrid Van Keilegom. Polynomial regression with censored data based on preliminary nonparametric estimation. Annals of the Institute of Statistical Mathematics, 59(2):273–297, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0066-4.

Li:2007:MIS

[2583] Linyuan Li and Yimin Xiao. Mean integrated squared error of nonlinear wavelet-based estimators with long memory data. Annals of the Institute of Statistical Mathematics, 59(2):299–324, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0048-6.

Mena:2007:SVG

[2584] Ramsés H. Mena and Stephen G.

Walker. On the stationary version of the generalized hyperbolic ARCH model. Annals of the Institute of Statistical Mathematics, 59(2):325–348, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0052-x.

Chopin:2007:DDC

[2585] Nicolas Chopin. Dynamic detection of change points in long time series. Annals of the Institute of Statistical Mathematics, 59(2):349–366, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0053-9.

Miao:2007:ANR

[2586] Baiqi Miao, Yuehua Wu, Donghai Liu, and Qian Tong. Asymptotic normality of the recursive M-estimators of the scale parameters. Annals of the Institute of Statistical Mathematics, 59(2):367–384, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0051-y.

Omelka:2007:SOL

[2587] Marek Omelka. Second-order linearity of Wilcoxon statistics. Annals of the Institute of Statistical Mathematics, 59(2):385–402, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0054-8.

Anonymous:2007:HCb

[2588] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 59(2):??, June 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Cappuccio:2007:AND

[2589] Nunzio Cappuccio and Diego Lubian. Asymptotic null distributions of stationarity and nonstationarity tests under local-to-finite variance errors. Annals of the Institute of Statistical Mathematics, 59(3):403–423, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0060-x.

Abramovich:2007:POB

[2590] Felix Abramovich, Claudia Angelini, and Daniela De Canditiis. Pointwise optimality of Bayesian wavelet estimators. Annals of the Institute of Statistical Mathematics, 59(3): 425–434, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0071-7.

Fourdrinier:2007:IMS

[2591] D. Fourdrinier and S. Pergamenshchikov. Improved model selection method for a regression function with dependent noise. Annals of the Institute of Statistical Mathematics, 59(3): 435–464, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0063-7.

Kluszczynski:2007:ISP

[2592] R. Kluszczyński, M. N. M. van Lieshout, and T. Schreiber. Image segmentation by polygonal Markov fields. Annals of the Institute of Statistical Mathematics, 59(3):465–486, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0062-8.

Durante:2007:GAC

[2593] Fabrizio Durante, José Juan Quesada-Molina, and Carlo Sempi. A generalization of the Archimedean class of bivariate copulas. Annals of the Institute of Statistical Mathematics, 59 (3):487–498, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0061-9.

Royen:2007:IRA

[2594] T. Royen. Integral representations and approximations for multivariate gamma distributions. Annals of the Institute of Statistical Mathematics, 59 (3):499–513, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0057-5.

Zavala:2007:CET

[2595] Arturo A. Z. Zavala, Heleno Bolfarine, and Mário de Castro. Consistent estimation and testing in heteroscedastic polynomial errors-in-variables models. Annals of the Institute of Statistical Mathematics, 59(3):515–530, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0069-1.

Harrar:2007:AEN

[2596] Solomon W. Harrar and Arjun K. Gupta. Asymptotic expansion for the null distribution of the F -statistic in one-way ANOVA under non-normality. Annals of the Institute of Statistical Mathematics, 59(3):531–556, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0055-7.

Vellaisamy:2007:SCR

[2597] P. Vellaisamy and V. Vijay. Some collapsibility results for n-dimensional contingency tables. Annals of the Institute of Statistical Mathematics, 59 (3):557–576, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0058-4.

Inoue:2007:JDN

[2598] Kiyoshi Inoue and Sigeo Aki. Joint distributions of numbers of runs of specified lengths in a sequence of Markov dependent multistate trials. Annals of the Institute of Statistical Mathematics, 59 (3):577–595, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0056-6.

Aki:2007:WTF

[2599] Sigeo Aki and Katuomi Hirano. On the waiting time for the first success run. Annals of the Institute of Statistical Mathematics, 59(3):597–602, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0059-3.

da-Silva:2007:APS

[2600] Cibele Queiroz da Silva. Asymptotics for a population size estimator of a partially uncatchable population. Annals of the Institute of Statistical Mathematics, 59(3):603–615, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0065-5.

Anonymous:2007:HCc

[2601] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 59(3):??, September 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Efromovich:2007:ONE

[2602] Sam Efromovich. Optimal nonparametric estimation of the density of regression errors with finite support. Annals of the Institute of Statistical Mathematics, 59(4):617–654, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0067-3.

Dette:2007:ODE

[2603] Holger Dette, Viatcheslav B. Melas, and Piter Shpilev. Optimal designs for estimating the coefficients of the lower frequencies in trigonometric regression models. Annals of the Institute of Statistical Mathematics, 59 (4):655–673, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0068-2.

Yu:2007:BJT

[2604] Qiqing Yu, George Y. C. Wong, and Menggang Yu. Buckley–James-type of estimators under the classical case cohort design. Annals of the Institute of Statistical Mathematics, 59(4): 675–695, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0086-0.

Poskitt:2007:AAN

[2605] D. S. Poskitt. Autoregressive approximation in nonstandard situations: the fractionally integrated and noninvertible cases. Annals of the Institute of Statistical Mathematics, 59 (4):697–725, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0074-4.

Rapallo:2007:TSM

[2606] Fabio Rapallo. Toric statistical models:

parametric and binomial representations. Annals of the Institute of Statistical Mathematics, 59(4):727–740, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0079-z.

Hassenforder:2007:EDL

[2607] Claudie Hassenforder and Sabine Mercier. Exact distribution of the local score for Markovian sequences. Annals of the Institute of Statistical Mathematics, 59(4):741–755, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0064-6.

Ghosh:2007:EPA

[2608] Kaushik Ghosh and Ram C. Tiwari.

Empirical process approach to some two-sample problems based on ranked set samples. Annals of the Institute of Statistical Mathematics, 59(4): 757–787, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0073-5.

Taylor:2007:MMC

[2609] M. D. Taylor. Multivariate measures of concordance. Annals of the Institute of Statistical Mathematics, 59(4): 789–806, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0076-2.

Takane:2007:CGI

[2610] Yoshio Takane, Yongge Tian, and Haruo Yanai. On constrained generalized inverses of matrices and their properties. Annals of the Institute of Statistical Mathematics, 59(4):807–820, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0075-3.

Anonymous:2007:HCd

[2611] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 59(4):??, December 2007. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aoyama:2008:FPT

[2612] Kazuki Aoyama, Kunio Shimizu, and S. H. Ong. A first-passage time random walk distribution with five transition probabilities: a generalization of the shifted inverse trinomial. Annals of the Institute of Statistical Mathematics, 60(1):1–20, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0084-2.

Zychaluk:2008:CVM

[2613] Kamila Zychaluk and Prakash N. Patil.

A cross-validation method for data with ties in kernel density estimation. Annals of the Institute of Statistical Mathematics, 60(1):21–44, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0077-1.

Antognini:2008:RGE

[2614] Alessandro Baldi Antognini, Paola Bortot, and Alessandra Giovagnoli. Randomized group up and down experiments. Annals of the Institute of Statistical Mathematics, 60(1):45–59, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0081-5.

Boik:2008:ACI

[2615] Robert J. Boik. Accurate confidence intervals in regression analyses of nonnormal data. Annals of the Institute of Statistical Mathematics, 60(1):61–83, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0085-1.

Fourdrinier:2008:EGF

[2616] D. Fourdrinier and P. Lepelletier.

Estimating a general function of a quadratic function. Annals of the Institute of Statistical Mathematics, 60 (1):85–119, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0072-6.

Zhang:2008:EMP

[2617] Zhengjun Zhang. The estimation of M4 processes with geometric moving patterns. Annals of the Institute of Statistical Mathematics, 60(1):121–150, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0078-0.

Balakrishnan:2008:PCH

[2618] N. Balakrishnan and Erhard Cramer. Progressive censoring from heterogeneous distributions with applications to robustness. Annals of the Institute of Statistical Mathematics, 60(1):151–171, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0070-8.

Koutras:2008:BMC

[2619] M. V. Koutras, S. Bersimis, and D. L. Antzoulakos. Bivariate Markov chain embeddable variables of polynomial type. Annals of the Institute of Statistical Mathematics, 60(1):173–191, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0080-6.

Pozdnyakov:2008:OSM

[2620] Vladimir Pozdnyakov. On occurrence of subpattern and method of gambling teams. Annals of the Institute of Statistical Mathematics, 60(1):193–203, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0082-4.

Field:2008:SAMa

[2621] Chris Field, John Robinson, and Elvezio Ronchetti. Saddlepoint approximations for multivariate M-estimates with applications to bootstrap accuracy. Annals of the Institute of Statistical Mathematics, 60(1):205–224, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0083-3.

Field:2008:SAMb

[2622] Chris Field, John Robinson, and Elvezio Ronchetti. Saddlepoint approximations for multivariate M-estimates with applications to bootstrap accuracy. Annals of the Institute of Statistical Mathematics, 60(1):225–227, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0101-5.

Anonymous:2008:HCa

[2623] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 60(1):??, March 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Aoki:2008:MIM

[2624] Satoshi Aoki and Akimichi Takemura.

Minimal invariant Markov basis for sampling contingency tables with fixed marginals. Annals of the Institute of Statistical Mathematics, 60(2):229–256, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0089-x.

Nkiet:2008:CED

[2625] Guy Martial Nkiet. Consistent estimation of the dimensionality in sliced inverse regression. Annals of the Institute of Statistical Mathematics, 60 (2):257–271, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0106-0.

Molanes-Lopez:2008:PBS

[2626] Elisa María Molanes-López and Ricardo Cao. Plug-in bandwidth selector for the kernel relative density estimator. Annals of the Institute of Statistical Mathematics, 60(2):273–300, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0108-y.

Chacko:2008:EPM

[2627] Manoj Chacko and P. Yageen Thomas. Estimation of a parameter of Morgenstern type bivariate exponential distribution by ranked set sampling. Annals of the Institute of Statistical Mathematics, 60(2):301–318, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0088-y.

Feng:2008:LLR

[2628] Qunqiang Feng, Hosam M. Mahmoud, and Alois Panholzer. Limit laws for the Randić index of random binary tree models. Annals of the Institute of Statistical Mathematics, 60(2):319–343, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0107-z.

Davydov:2008:WCR

[2629] Youri Davydov and Ricardas Zitikis.

On weak convergence of random fields. Annals of the Institute of Statistical Mathematics, 60(2):345–365, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-006-
0090-4.

Hayashi:2008:ANC

[2630] Takaki Hayashi and Nakahiro Yoshida.

Asymptotic normality of a covariance estimator for nonsynchronously observed diffusion processes. Annals of the Institute of Statistical Mathematics, 60(2):367–406, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0138-0.

Hyndman:2008:APS

[2631] Rob J. Hyndman, Muhammad Akram, and Blyth C. Archibald. The admissible parameter space for exponential smoothing models. Annals of the Institute of Statistical Mathematics, 60(2):407–426, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-006-0109-x.

Wang:2008:SSS

[2632] Jiantian Wang. Small-sample studies on right censored data with discrete failure times. Annals of the Institute of Statistical Mathematics, 60(2):427–440, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-006-
                                        0087-z.

Zhao:2008:ELI

[2633] Yichuan Zhao and Song Yang. Empirical likelihood inference for censored median regression with weighted empirical hazard functions. Annals of the Institute of Statistical Mathematics, 60(2):441–457, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0110-4.

Bouzar:2008:SSDa

[2634] Nadjib Bouzar. The semi-Sibuya distribution. Annals of the Institute of Statistical Mathematics, 60(2):459–464, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0118-4.

Anonymous:2008:HCb

[2635] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 60(2):??, June 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nakashima:2008:CMO

[2636] Eiji Nakashima, Kazuo Neriishi, and Atsushi Minamoto. Comparison of methods for ordinal lens opacity data from atomic-bomb survivors: univariate worse-eye method and bivariate GEE method using global odds ratio. Annals of the Institute of Statistical Mathematics, 60(3):465–482, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0113-9.

Solari:2008:SRB

[2637] Fabrizio Solari, Brunero Liseo, and Dongchu Sun. Some remarks on Bayesian inference for one-way ANOVA models. Annals of the Institute of Statistical Mathematics, 60 (3):483–498, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0117-5.

Sun:2008:BHL

[2638] Dongchu Sun and Paul L. Speckman.

Bayesian hierarchical linear mixed models for additive smoothing splines. Annals of the Institute of Statistical Mathematics, 60(3):499–517, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0127-3.

Ojeda:2008:BAM

[2639] J. L. Ojeda, J. A. Cristóbal, and J. T. Alcalá. A bootstrap approach to model checking for linear models under length-biased data. Annals of the Institute of Statistical Mathematics, 60 (3):519–543, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-006-0111-3.

Lu:2008:MLE

[2640] Wenbin Lu. Maximum likelihood estimation in the proportional hazards cure model. Annals of the Institute of Statistical Mathematics, 60(3): 545–574, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0120-x.

Shen:2008:ELC

[2641] Junshan Shen and Shuyuan He. Empirical likelihood confidence intervals for hazard and density functions under right censorship. Annals of the Institute of Statistical Mathematics, 60 (3):575–589, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0114-8.

Nair:2008:SRL

[2642] N. Unnikrishnan Nair and K. K. Sudheesh. Some results on lower variance bounds useful in reliability modeling and estimation. Annals of the Institute of Statistical Mathematics, 60(3): 591–603, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0119-3.

Beutner:2008:NIS

[2643] Eric Beutner. Nonparametric inference for sequential k-out-of-n systems. Annals of the Institute of Statistical Mathematics, 60(3):605–626, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0115-7.

Baddeley:2008:PRS

[2644] A. Baddeley, J. Møller, and A. G.

Pakes. Properties of residuals for spatial point processes. Annals of the Institute of Statistical Mathematics, 60 (3):627–649, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0116-6.

Beirlant:2008:HOE

[2645] J. Beirlant, A. Berlinet, and G. Biau.

Higher order estimation at Lebesgue points. Annals of the Institute of Statistical Mathematics, 60(3):651– 677, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0112-x.

Liang:2008:SNU

[2646] Jiajuan Liang, Kai-Tai Fang, and Fred J. Hickernell. Some necessary uniform tests for spherical symmetry. Annals of the Institute of Statistical Mathematics, 60(3):679–696, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0121-9.

Anonymous:2008:HCc

[2647] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 60(3):??, September 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Higuchi:2008:PFS

[2648] Tomoyuki Higuchi and Takashi Washio. Preface: Featured section on datamining and statistical science. Annals of the Institute of Statistical Mathematics, 60(4):697–698, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10463-008-0208-y.

Sugiyama:2008:DIE

[2649] Masashi Sugiyama, Taiji Suzuki, Shinichi Nakajima, Hisashi Kashima, Paul von Bünau, and Motoaki Kawanabe. Direct importance estimation for covariate shift adaptation. Annals of the Institute of Statistical Mathematics, 60(4):699–746, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0197-x.

Sheridan:2008:PAM

[2650] Paul Sheridan, Yuichi Yagahara, and Hidetoshi Shimodaira. A preferential attachment model with Poisson growth for scale-free networks. Annals of the Institute of Statistical Mathematics, 60 (4):747–761, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0181-5.

Ando:2008:MBS

[2651] Tomohiro Ando. Measuring the baseline sales and the promotion effect for incense products: a Bayesian statespace modeling approach. Annals of the Institute of Statistical Mathematics, 60 (4):763–780, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0194-0.

Hokimoto:2008:ALT

[2652] Tsukasa Hokimoto and Kunio Shimizu.

An angular-linear time series model for wave-height prediction. Annals of the Institute of Statistical Mathematics, 60 (4):781–800, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0207-z.

Kumon:2008:SSW

[2653] Masayuki Kumon and Akimichi Takemura. On a simple strategy weakly forcing the strong law of large numbers in the bounded forecasting game. Annals of the Institute of Statistical Mathematics, 60(4):801–812, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0125-5.

Khmaladze:2008:LEP

[2654] Estate Khmaladze and Wolfgang Weil.

Local empirical processes near boundaries of convex bodies. Annals of the Institute of Statistical Mathematics, 60 (4):813–842, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0123-7.

Beran:2008:EMM

[2655] Rudolf Beran. Estimating a mean matrix: boosting efficiency by multiple affine shrinkage. Annals of the Institute of Statistical Mathematics, 60 (4):843–864, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0128-2.

Gregori:2008:PNS

[2656] P. Gregori, E. Porcu, J. Mateu, and Z. Sasvári. On potentially negative space time covariances obtained as sum of products of marginal ones. Annals of the Institute of Statistical Mathematics, 60(4):865–882, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0122-8.

Wang:2008:SON

[2657] Liqun Wang and Alexandre Leblanc. Second-order nonlinear least squares estimation. Annals of the Institute of Statistical Mathematics, 60(4): 883–900, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0139-z.

Bouzar:2008:SSDb

[2658] Nadjib Bouzar. Semi-self-decomposable distributions on Z_{+}. Annals of the Institute of Statistical Mathematics, 60 (4):901–917, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0124-6.

Anonymous:2008:HCd

[2659] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 60(4):??, December 2008. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kakizawa:2009:MCS

[2660] Yoshihide Kakizawa. Multiple comparisons of several homoscedastic multivariate populations. Annals of the Institute of Statistical Mathematics, 61(1):1–26, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0134-4.

Liang:2009:GPL

[2661] Hua Liang. Generalized partially linear mixed-effects models incorporating mismeasured covariates. Annals of the Institute of Statistical Mathematics, 61(1):27–46, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0146-0.

Wang:2009:SEP

[2662] Qi-Hua Wang. Statistical estimation in partial linear models with covariate data missing at random. Annals of the Institute of Statistical Mathematics, 61(1):47–84, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0137-1.

Perez-Gonzalez:2009:APL

[2663] A. Pérez-González, J. M. Vilar-Fernández, and W. González-Manteiga. Asymptotic properties of local polynomial regression with missing data and correlated errors. Annals of the Institute of Statistical Mathematics, 61(1):85–109, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0136-2.

Dette:2009:EMC

[2664] Holger Dette and Kay Pilz. On the estimation of a monotone conditional variance in nonparametric regression. Annals of the Institute of Statistical Mathematics, 61(1):111–141, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0126-4.

Biedermann:2009:COD

[2665] Stefanie Biedermann, Holger Dette, and Philipp Hoffmann. Constrained optimal discrimination designs for Fourier regression models. Annals of the Institute of Statistical Mathematics, 61(1):143–157, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0133-5.

Fruhwirth-Schnatter:2009:BES

[2666] Sylvia Frühwirth-Schnatter and Leopold Sögner. Bayesian estimation of stochastic volatility models based on OU processes with marginal gamma law. Annals of the Institute of Statistical Mathematics, 61(1):159–179, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0130-8.

Masuda:2009:NEI

[2667] Hiroki Masuda. Notes on estimating inverse-Gaussian and gamma subordinators under high-frequency sampling. Annals of the Institute of Statistical Mathematics, 61(1):181–195, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0131-7.

Chiang:2009:VCM

[2668] Chin-Tsang Chiang and Mei-Cheng Wang. Varying-coefficient model for the occurrence rate function of recurrent events. Annals of the Institute of Statistical Mathematics, 61(1):197–213, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0129-1.

Marchand:2009:EBP

[2669] Éric Marchand and François Perron.

Estimating a bounded parameter for symmetric distributions. Annals of the Institute of Statistical Mathematics, 61 (1):215–234, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0132-6.

Ozturk:2009:ETS

[2670] Omer Ozturk and N. Balakrishnan. Exact two-sample nonparametric test for quantile difference between two populations based on ranked set samples. Annals of the Institute of Statistical Mathematics, 61(1):235–249, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0141-5.

Balakrishnan:2009:EIS

[2671] N. Balakrishnan, Qihao Xie, and D. Kundu. Exact inference for a simple step-stress model from the exponential distribution under time constraint. Annals of the Institute of Statistical Mathematics, 61(1):251–274, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0135-3.

Anonymous:2009:HCa

[2672] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 61(1):??, March 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Chia:2009:DSU

[2673] Nicholas Chia and Junji Nakano. Mdecomposability and symmetric unimodal densities in one dimension. Annals of the Institute of Statistical Mathematics, 61(2):275–289, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10463-007-
0144-2.

Wei:2009:RMS

[2674] Wen Hsiang Wei. On regression model selection for the data with correlated errors. Annals of the Institute of Statistical Mathematics, 61(2):291–308, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0145-1.

Vellaisamy:2009:LLM

[2675] P. Vellaisamy and V. Vijay. Log-linear modeling using conditional log-linear structures. Annals of the Institute of Statistical Mathematics, 61(2):309–329, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0153-1.

Ando:2009:NLD

[2676] Tomohiro Ando and Sadanori Konishi. Nonlinear logistic discrimination via regularized radial basis functions for classifying high-dimensional data. Annals of the Institute of Statistical Mathematics, 61(2):331–353, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0143-3.

Alvarez:2009:BIC

[2677] Enrique E. Alvarez and Dipak K. Dey.

Bayesian isotonic changepoint analysis. Annals of the Institute of Statistical Mathematics, 61(2):355–370, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0148-y.

Shcherbakov:2009:MSP

[2678] V. Shcherbakov. On a model of sequential point patterns. Annals of the Institute of Statistical Mathematics, 61(2):371–390, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0147-z.

Chan:2009:MEN

[2679] Ngai Hang Chan and Rongmao Zhang. M-estimation in nonparametric regression under strong dependence and infinite variance. Annals of the Institute of Statistical Mathematics, 61(2):391–411, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0142-4.

Honda:2009:NDE

[2680] Toshio Honda. Nonparametric density estimation for linear processes with infinite variance. Annals of the Institute of Statistical Mathematics, 61(2):413–439, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0149-x.

Rao:2009:CIC

[2681] B. L. S. Prakasa Rao. Conditional independence, conditional mixing and conditional association. Annals of the Institute of Statistical Mathematics, 61(2):441–460, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0152-2.

Shen:2009:CRB

[2682] Pao sheng Shen. A class of rankbased test for left-truncated and rightcensored data. Annals of the Institute of Statistical Mathematics, 61 (2):461–476, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0151-3.

Berg:2009:HOA

[2683] Arthur Berg and Dimitris N. Politis. Higher-order accurate polyspectral estimation with flat-top lag-windows. Annals of the Institute of Statistical Mathematics, 61(2):477–498, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0154-0.

Inoue:2009:WTD

[2684] Kiyoshi Inoue and Sigeo Aki. On waiting time distributions associated with compound patterns in a sequence of multi-state trials. Annals of the Institute of Statistical Mathematics, 61(2):499–516, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0140-6.

Tian:2009:OPA

[2685] Yongge Tian and Yoshio Takane.

On V -orthogonal projectors associated with a semi-norm. Annals of the Institute of Statistical Mathematics, 61(2):517–530, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0150-4.

Anonymous:2009:HCb

[2686] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 61(2):??, June 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kato:2009:IPM

[2687] Kengo Kato. Improved prediction for a multivariate normal distribution with unknown mean and variance. Annals of the Institute of Statistical Mathematics, 61(3):531–542, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0163-z.

Visek:2009:CIW

[2688] Jan Ámos Vísek. Consistency of the instrumental weighted variables. Annals of the Institute of Statistical Mathematics, 61(3):543–578, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0159-8.

Jureckova:2009:TTI

[2689] Jana Jurecková, Hira L. Koul, and Jan Picek. Testing the tail index in autoregressive models. Annals of the Institute of Statistical Mathematics, 61 (3):579–598, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0155-z.

Helmers:2009:EIC

[2690] Roelof Helmers and I. Wayan Mangku.

Estimating the intensity of a cyclic Poisson process in the presence of linear trend. Annals of the Institute of Statistical Mathematics, 61(3): 599–628, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0160-2.

Sakamoto:2009:TOA

[2691] Yuji Sakamoto and Nakahiro Yoshida.

Third-order asymptotic expansion of M-estimators for diffusion processes. Annals of the Institute of Statistical Mathematics, 61(3):629–661, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0190-4.

Wang:2009:EFS

[2692] Jing Wang and Lijian Yang. Efficient and fast spline-backfitted kernel smoothing of additive models. Annals of the Institute of Statistical Mathematics, 61(3):663–690, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0157-x.

Abramovich:2009:OTA

[2693] Felix Abramovich, Italia De Feis, and Theofanis Sapatinas. Optimal testing for additivity in multiple nonparametric regression. Annals of the Institute of Statistical Mathematics, 61 (3):691–714, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0164-y.

Qiu:2009:JPS

[2694] Peihua Qiu. Jump-preserving surface reconstruction from noisy data. Annals of the Institute of Statistical Mathematics, 61(3):715–751, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-007-
                                        0166-9.

Balakrishnan:2009:SMM

[2695] N. Balakrishnan and G. Iliopoulos. Stochastic monotonicity of the MLE of exponential mean under different censoring schemes. Annals of the Institute of Statistical Mathematics, 61(3): 753–772, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-007-0156-y.

Anonymous:2009:HCc

[2696] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 61(3):??, September 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Otsu:2009:GNP

[2697] Taisuke Otsu. Generalized Neyman– Pearson optimality of empirical likelihood for testing parameter hypotheses. Annals of the Institute of Statistical Mathematics, 61(4):773–787, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0172-6.

Huwang:2009:URT

[2698] Longcheen Huwang, Y. H. Steve Huang, and Yi-Hua Tina Wang. Uniformly robust tests in errors-in-variables models. Annals of the Institute of Statistical Mathematics, 61 (4):789–810, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0167-8.

Araki:2009:FRM

[2699] Yuko Araki, Sadanori Konishi, Shuichi Kawano, and Hidetoshi Matsui. Functional regression modeling via regularized Gaussian basis expansions. Annals of the Institute of Statistical Mathematics, 61(4):811–833, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0161-1.

Choi:2009:APP

[2700] Taeryon Choi. Asymptotic properties of posterior distributions in nonparametric regression with non-Gaussian errors. Annals of the Institute of Statistical Mathematics, 61(4):835–859, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0168-2.

Dette:2009:STP

[2701] Holger Dette and Benjamin Hetzler. A simple test for the parametric form of the variance function in nonparametric regression. Annals of the Institute of Statistical Mathematics, 61 (4):861–886, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0169-1.

Alvarez-Andrade:2009:PHR

[2702] Sergio Alvarez-Andrade, N. Balakrishnan, and Laurent Bordes. Proportional hazards regression under progressive Type-II censoring. Annals of the Institute of Statistical Mathematics, 61 (4):887–903, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0170-8.

Zou:2009:LIA

[2703] Qingming Zou, Zhongyi Zhu, and Jinglong Wang. Local influence analysis for penalized Gaussian likelihood estimation in partially linear singleindex models. Annals of the Institute of Statistical Mathematics, 61(4): 905–918, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-007-0158-9.

Negri:2009:GFT

[2704] Ilia Negri and Yoichi Nishiyama. Goodness of fit test for ergodic diffusion processes. Annals of the Institute of Statistical Mathematics, 61(4):919–928, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-007-
0162-0.

Adelfio:2009:PPD

[2705] Giada Adelfio and Frederic Paik Schoenberg. Point process diagnostics based on weighted second-order statistics and their asymptotic properties. Annals of the Institute of Statistical Mathematics, 61(4):929–948, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0177-1.

Kamatani:2009:MHA

[2706] Kengo Kamatani. Metropolis–Hastings algorithms with acceptance ratios of nearly 1. Annals of the Institute of Statistical Mathematics, 61(4): 949–967, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0180-6.

Vovk:2009:MOG

[2707] Vladimir Vovk. Merging of opinions in game-theoretic probability. Annals of the Institute of Statistical Mathematics, 61(4):969–993, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-007-
                                        0165-x.

Ogasawara:2009:AES

[2708] Haruhiko Ogasawara. Asymptotic expansions in the singular value decomposition for cross covariance and correlation under nonnormality. Annals of the Institute of Statistical Mathematics, 61(4):995–1017, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0174-4.

Anonymous:2009:HCd

[2709] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 61(4):??, December 2009. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Konishi:2010:PSI

[2710] Sadanori Konishi and Genshiro Kitagawa. Preface: Special issue in honor of Dr. hirotugu Akaike. Annals of the Institute of Statistical Mathematics, 62(1):1–2, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-009-0269-
                                        6.

Akaike:2010:MST

[2711] Hirotugu Akaike. Making statistical thinking more productive. Annals of the Institute of Statistical Mathematics, 62(1):3–9, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0238-0.

Dean:2010:LCA

[2712] Nema Dean and Adrian E. Raftery.

Latent class analysis variable selection. Annals of the Institute of Statistical Mathematics, 62(1):11–35, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0258-9.

Koyama:2010:BDN

[2713] Shinsuke Koyama, Uri T. Eden, Emery N. Brown, and Robert E. Kass. Bayesian decoding of neural spike trains. Annals of the Institute of Statistical Mathematics, 62(1):37–59, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0249-x.

Briers:2010:SAS

[2714] Mark Briers, Arnaud Doucet, and Simon Maskell. Smoothing algorithms for state-space models. Annals of the Institute of Statistical Mathematics, 62 (1):61–89, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0236-2.

Terui:2010:FMS

[2715] Nobuhiko Terui, Masataka Ban, and Toshihiko Maki. Finding market structure by sales count dynamics — multivariate structural time series models with hierarchical structure for count data. Annals of the Institute of Statistical Mathematics, 62(1):91–107, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0244-2.

Tsai:2010:DBA

[2716] Chih-Ling Tsai, Hansheng Wang, and Ning Zhu. Does a Bayesian approach generate robust forecasts? evidence from applications in portfolio investment decisions. Annals of the Institute of Statistical Mathematics, 62 (1):109–116, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0250-4.

Lukacs:2010:MSB

[2717] Paul M. Lukacs, Kenneth P. Burnham, and David R. Anderson. Model selection bias and Freedman’s paradox. Annals of the Institute of Statistical Mathematics, 62(1):117–125, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0234-4.

Adelfio:2010:HKE

[2718] Giada Adelfio and Yosihiko Ogata. Hybrid kernel estimates of spacetime earthquake occurrence rates using the epidemic-type aftershock sequence model. Annals of the Institute of Statistical Mathematics, 62(1): 127–143, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0268-7.

Martins:2010:DDB

[2719] Leonardo de Oliveira Martins and Hirohisa Kishino. Distribution of distances between topologies and its effect on detection of phylogenetic recombination. Annals of the Institute of Statistical Mathematics, 62(1):145–159, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0259-8.

Uchida:2010:CBI

[2720] Masayuki Uchida. Contrast-based information criterion for ergodic diffusion processes from discrete observations. Annals of the Institute of Statistical Mathematics, 62(1):161–187, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0245-1.

Shimodaira:2010:FBM

[2721] Hidetoshi Shimodaira. Frequentist and Bayesian measures of confidence via multiscale bootstrap for testing three regions. Annals of the Institute of Statistical Mathematics, 62(1): 189–208, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0247-z.

Kitagawa:2010:BVR

[2722] Genshiro Kitagawa and Sadanori Konishi. Bias and variance reduction techniques for bootstrap information criteria. Annals of the Institute of Statistical Mathematics, 62(1):209–234, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-009-
                                        0237-1.

Anonymous:2010:HCa

[2723] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(1):??, February 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Ueki:2010:BLQ

[2724] Masao Ueki and Kaoru Fueda. Boosting local quasi-likelihood estimators. Annals of the Institute of Statistical Mathematics, 62(2):235–248, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0173-5.

Efromovich:2010:OIC

[2725] Sam Efromovich. Oracle inequality for conditional density estimation and an actuarial example. Annals of the Institute of Statistical Mathematics, 62(2):249–275, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0185-1.

Qi:2010:TIH

[2726] Yongcheng Qi. On the tail index of a heavy tailed distribution. Annals of the Institute of Statistical Mathematics, 62(2):277–298, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0176-2.

Aoki:2010:MBG

[2727] Satoshi Aoki, Takayuki Hibi, Hidefumi Ohsugi, and Akimichi Takemura. Markov basis and Gröbner basis of Segrè–Veronese configuration for testing independence in group-wise selections. Annals of the Institute of Statistical Mathematics, 62(2):299–321, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0171-7.

deUna-Alvarez:2010:NEC

[2728] Jacobo de Uña-Álvarez and M. Carmen Iglesias-Pérez. Nonparametric estimation of a conditional distribution from length-biased data. Annals of the Institute of Statistical Mathematics, 62(2):323–341, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0178-0.

Sharia:2010:RPE

[2729] Teo Sharia. Recursive parameter estimation: asymptotic expansion. Annals of the Institute of Statistical Mathematics, 62(2):343–362, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0179-z.

Arellano-Valle:2010:IPQ

[2730] Reinaldo B. Arellano-Valle and Marc G.

Genton. An invariance property of quadratic forms in random vectors with a selection distribution, with application to sample variogram and covariogram estimators. Annals of the Institute of Statistical Mathematics, 62(2):363–381, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0175-3.

Mougeot:2010:PTC

[2731] Mathilde Mougeot and Karine Tribouley. Procedure of test to compare the tail indices. Annals of the Institute of Statistical Mathematics, 62(2):383–412, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0198-9.

Anonymous:2010:HCb

[2732] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(2):??, April 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Ueki:2010:OTP

[2733] Masao Ueki and Kaoru Fueda. Optimal tuning parameter estimation in maximum penalized likelihood method. Annals of the Institute of Statistical Mathematics, 62(3):413–438, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0186-0.

Heuchenne:2010:ENL

[2734] Cédric Heuchenne and Ingrid Van Keilegom. Estimation in nonparametric location-scale regression models with censored data. Annals of the Institute of Statistical Mathematics, 62(3):439–463, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0219-3.

Teodorescu:2010:GTD

[2735] Bianca Teodorescu, Ingrid Van Keilegom, and Ricardo Cao. Generalized time-dependent conditional linear models under left truncation and right censoring. Annals of the Institute of Statistical Mathematics, 62(3):465–485, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0187-z.

Xu:2010:SEV

[2736] Jinfeng Xu and Zhiliang Ying. Simultaneous estimation and variable selection in median regression using lassotype penalty. Annals of the Institute of Statistical Mathematics, 62(3):487–514, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0184-2.

Sancetta:2010:BMS

[2737] Alessio Sancetta. Bootstrap model selection for possibly dependent and heterogeneous data. Annals of the Institute of Statistical Mathematics, 62(3):515–546, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0183-3.

Hayashi:2010:LBR

[2738] Masahito Hayashi. Limiting behavior of relative Rényi entropy in a nonregular location shift family. Annals of the Institute of Statistical Mathematics, 62(3):547–569, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0182-4.

Aoshima:2010:ASO

[2739] Makoto Aoshima and Kazuyoshi Yata. Asymptotic second-order consistency for two-stage estimation methodologies and its applications. Annals of the Institute of Statistical Mathematics, 62(3):571–600, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0188-y.

Anonymous:2010:HCc

[2740] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(3):??, June 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takemura:2010:P

[2741] Akimichi Takemura. Preface. Annals of the Institute of Statistical Mathematics, 62(4):601–602, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-010-0297-
                                        2.

Sturmfels:2010:MGS

[2742] Bernd Sturmfels and Caroline Uhler. Multivariate Gaussians, semidefinite matrix completion, and convex algebraic geometry. Annals of the Institute of Statistical Mathematics, 62 (4):603–638, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0295-4.

Ohsugi:2010:NVA

[2743] Hidefumi Ohsugi and Takayuki Hibi.

Non-very ample configurations arising from contingency tables. Annals of the Institute of Statistical Mathematics, 62 (4):639–644, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0288-3.

Kuriki:2010:GPM

[2744] Satoshi Kuriki and Yasuhide Numata.

Graph presentations for moments of noncentral Wishart distributions and their applications. Annals of the Institute of Statistical Mathematics, 62 (4):645–672, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0279-4.

Berstein:2010:MAD

[2745] Yael Berstein, Hugo Maruri-Aguilar, Shmuel Onn, Eva Riccomagno, and Henry Wynn. Minimal average degree aberration and the state polytope for experimental designs. Annals of the Institute of Statistical Mathematics, 62 (4):673–698, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0291-8.

Aoki:2010:SOC

[2746] Satoshi Aoki. Some optimal criteria of model-robustness for two-level non-regular fractional factorial designs. Annals of the Institute of Statistical Mathematics, 62(4):699–716, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0292-7.

Sullivant:2010:NBG

[2747] Seth Sullivant. Normal binary graph models. Annals of the Institute of Statistical Mathematics, 62(4):717–726, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0296-3.

Kahle:2010:DBI

[2748] Thomas Kahle. Decompositions of binomial ideals. Annals of the Institute of Statistical Mathematics, 62 (4):727–745, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0290-9.

Wang:2010:UAA

[2749] Jinfang Wang. A universal algebraic approach for conditional independence. Annals of the Institute of Statistical Mathematics, 62(4):747–773, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0278-5.

Drton:2010:FSF

[2750] Mathias Drton and Han Xiao. Finiteness of small factor analysis models. Annals of the Institute of Statistical Mathematics, 62(4):775–783, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0293-6.

Rapallo:2010:MBS

[2751] Fabio Rapallo and Ruriko Yoshida.

Markov bases and subbases for bounded contingency tables. Annals of the Institute of Statistical Mathematics, 62(4):785–805, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0289-2.

Sumi:2010:AMR

[2752] Toshio Sumi, Mitsuhiro Miyazaki, and Toshio Sakata. About the maximal rank of 3-tensors over the real and the complex number field. Annals of the Institute of Statistical Mathematics, 62 (4):807–822, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0294-5.

Anonymous:2010:HCd

[2753] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(4):??, August 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Nishiyama:2010:NIM

[2754] Yoichi Nishiyama. Nonparametric inference in multiplicative intensity model by discrete time observation. Annals of the Institute of Statistical Mathematics, 62(5):823–833, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0196-y.

Shen:2010:NAD

[2755] Pao sheng Shen. Nonparametric analysis of doubly truncated data. Annals of the Institute of Statistical Mathematics, 62(5):835–853, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0192-2.

Bogsted:2010:DRS

[2756] Martin Bøgsted and Susan M. Pitts. Decompounding random sums: a nonparametric approach. Annals of the Institute of Statistical Mathematics, 62 (5):855–872, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0200-6.

Grasselli:2010:DCN

[2757] M. R. Grasselli. Dual connections in nonparametric classical information geometry. Annals of the Institute of Statistical Mathematics, 62(5):873–896, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0191-3.

Omelka:2010:UAR

[2758] Marek Omelka and Matías Salibián-Barrera. Uniform asymptotics for S- and MM-regression estimators. Annals of the Institute of Statistical Mathematics, 62(5):897–927, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0189-x.

Tian:2010:WLS

[2759] Yongge Tian. Weighted least-squares estimators of parametric functions of the regression coefficients under a general linear model. Annals of the Institute of Statistical Mathematics, 62 (5):929–941, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0199-8.

Mondal:2010:WVA

[2760] Debashis Mondal and Donald B. Percival. Wavelet variance analysis for gappy time series. Annals of the Institute of Statistical Mathematics, 62 (5):943–966, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0195-z.

Dryden:2010:SIF

[2761] Ian L. Dryden, Alfred Kume, Huiling Le, and Andrew T. A. Wood. Statistical inference for functions of the covariance matrix in the stationary Gaussian time-orthogonal principal components model. Annals of the Institute of Statistical Mathematics, 62 (5):967–994, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0202-4.

Anonymous:2010:HCe

[2762] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(5):??, October 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Honda:2010:NEC

[2763] Toshio Honda. Nonparametric estimation of conditional medians for linear and related processes. Annals of the Institute of Statistical Mathematics, 62 (6):995–1021, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0203-3.

Lee:2010:SPE

[2764] Alan Lee and Yuichi Hirose. Semiparametric efficiency bounds for regression models under response-selective sampling: the profile likelihood approach. Annals of the Institute of Statistical Mathematics, 62(6):1023– 1052, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0205-1.

Cribari-Neto:2010:SBA

[2765] Francisco Cribari-Neto and Maria da Glória A. Lima. Sequences of biasadjusted covariance matrix estimators under heteroskedasticity of unknown form. Annals of the Institute of Statistical Mathematics, 62(6):1053–1082, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0201-5.

Konev:2010:GMS

[2766] Victor Konev and Serguei Pergamenchtchikov. General model selection estimation of a periodic regression with a Gaussian noise. Annals of the Institute of Statistical Mathematics, 62(6): 1083–1111, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0193-1.

Withers:2010:TEE

[2767] Christopher S. Withers and Saralees Nadarajah. Tilted Edgeworth expansions for asymptotically normal vectors. Annals of the Institute of Statistical Mathematics, 62(6):1113–1142, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0206-0.

Zhang:2010:ARD

[2768] Baoxue Zhang, Tianqing Liu, and Z. D.

Bai. Analysis of rounded data from dependent sequences. Annals of the Institute of Statistical Mathematics, 62(6): 1143–1173, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0224-6.

Anonymous:2010:HCf

[2769] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 62(6):??, December 2010. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Carbonell:2011:GWR

[2770] F. Carbonell, K. J. Worsley, and L. Galan. The geometry of the Wilks’s Λ random field. Annals of the Institute of Statistical Mathematics, 63 (1):1–27, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0204-2.

Ohlson:2011:EEU

[2771] Martin Ohlson, Zhanna Andrushchenko, and Dietrich von Rosen. Explicit estimators under m-dependence for a multivariate normal distribution. Annals of the Institute of Statistical Mathematics, 63(1):29–42, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0213-1.

Ghosh:2011:GDC

[2772] Malay Ghosh, Victor Mergel, and Ruitao Liu. A general divergence criterion for prior selection. Annals of the Institute of Statistical Mathematics, 63(1):43–58, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0226-4.

Chung:2011:LDP

[2773] Yeonseung Chung and David B. Dunson. The local Dirichlet process. Annals of the Institute of Statistical Mathematics, 63(1):59–80, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0218-9.

McIntyre:2011:DER

[2774] Julie McIntyre and Leonard A. Stefanski. Density estimation with replicate heteroscedastic measurements. Annals of the Institute of Statistical Mathematics, 63(1):81–99, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0220-x.

Eryilmaz:2011:NPS

[2775] Serkan Eryilmaz. A new perspective to stress-strength models. Annals of the Institute of Statistical Mathematics, 63 (1):101–115, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0211-3.

Ma:2011:ARM

[2776] Shuangge Ma. Additive risk model for current status data with a cured subgroup. Annals of the Institute of Statistical Mathematics, 63(1): 117–134, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0212-2.

Balakrishnan:2011:CMS

[2777] N. Balakrishnan and Xingqiu Zhao. A class of multi-sample nonparametric tests for panel count data. Annals of the Institute of Statistical Mathematics, 63(1):135–156, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0209-x.

Chatterjee:2011:APS

[2778] Arindam Chatterjee. Asymptotic properties of sample quantiles from a finite population. Annals of the Institute of Statistical Mathematics, 63 (1):157–179, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-008-0210-4.

Lengyel:2011:GRW

[2779] Tamás Lengyel. Gambler’s ruin and winning a series by m games. Annals of the Institute of Statistical Mathematics, 63(1):181–195, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-008-
                                        0214-0.

Inoue:2011:BFP

[2780] Kiyoshi Inoue and Sigeo Aki. Bivariate Fibonacci polynomials of order k with statistical applications. Annals of the Institute of Statistical Mathematics, 63 (1):197–210, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-008-0217-x.

Anonymous:2011:HCa

[2781] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(1):??, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Negri:2011:GFT

[2782] Ilia Negri and Yoichi Nishiyama. Goodness of fit test for small diffusions by discrete time observations. Annals of the Institute of Statistical Mathematics, 63(2):211–225, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0228-2.

Ma:2011:APS

[2783] Yanyuan Ma, Marc G. Genton, and Emanuel Parzen. Asymptotic properties of sample quantiles of discrete distributions. Annals of the Institute of Statistical Mathematics, 63(2):227–243, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-008-
0215-z.

Dette:2011:EAQ

[2784] Holger Dette and Regine Scheder. Estimation of additive quantile regression. Annals of the Institute of Statistical Mathematics, 63(2):245–265, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0225-5.

Liang:2011:APC

[2785] Han-Ying Liang and Jacobo de Uña-Álvarez. Asymptotic properties of conditional quantile estimator for censored dependent observations. Annals of the Institute of Statistical Mathematics, 63(2):267–289, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0230-8.

Stoyanov:2011:MPS

[2786] Jordan Stoyanov and Gwo Dong Lin. Mixtures of power series distributions: identifiability via uniqueness in problems of moments. Annals of the Institute of Statistical Mathematics, 63 (2):291–303, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0221-9.

Hernandez-Quintero:2011:ASM

[2787] Angelica Hernandez-Quintero, Jean-François Dupuy, and Gabriel Escarela. Analysis of a semiparametric mixture model for competing risks. Annals of the Institute of Statistical Mathematics, 63(2):305–329, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0229-1.

Yu:2011:ELM

[2788] Wen Yu, Yunting Sun, and Ming Zheng. Empirical likelihood method for linear transformation models. Annals of the Institute of Statistical Mathematics, 63(2):331–346, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0223-7.

Kojadinovic:2011:TSI

[2789] Ivan Kojadinovic and Jun Yan. Tests of serial independence for continuous multivariate time series based on a Möbius decomposition of the independence empirical copula process. Annals of the Institute of Statistical Mathematics, 63(2):347–373, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0257-x.

Balakrishnan:2011:FFW

[2790] N. Balakrishnan and Po Yang. Forms of four-word indicator functions with implications to two-level factorial designs. Annals of the Institute of Statistical Mathematics, 63(2):375–386, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0222-8.

Fujikoshi:2011:PEC

[2791] Yasunori Fujikoshi, Tamio Kan, Shin Takahashi, and Tetsuro Sakurai. Prediction error criterion for selecting variables in a linear regression model. Annals of the Institute of Statistical Mathematics, 63(2):387–403, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0233-5.

Chesneau:2011:AWE

[2792] Christophe Chesneau. On adaptive wavelet estimation of a quadratic functional from a deconvolution problem. Annals of the Institute of Statistical Mathematics, 63(2):405–429, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-009-
                                        0232-6.

Anonymous:2011:HCb

[2793] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(2):??, April 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Yoshida:2011:PTL

[2794] Nakahiro Yoshida. Polynomial type large deviation inequalities and quasilikelihood analysis for stochastic differential equations. Annals of the Institute of Statistical Mathematics, 63(3):431–479, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0263-z.

Trippa:2011:EBP

[2795] Lorenzo Trippa, Paolo Bulla, and Sonia Petrone. Extended Bernstein prior via reinforced urn processes. Annals of the Institute of Statistical Mathematics, 63(3):481–496, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0227-3.

Chen:2011:RES

[2796] Hua Yun Chen. Representations of efficient score for coarse data problems based on Neumann series expansion. Annals of the Institute of Statistical Mathematics, 63(3):497–509, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0231-7.

Yi:2011:SMA

[2797] Grace Y. Yi, Wenqing He, and Hua Liang. Semiparametric marginal and association regression methods for clustered binary data. Annals of the Institute of Statistical Mathematics, 63(3):511–533, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0239-z.

Crowder:2011:EFR

[2798] Martin Crowder. Estimating functions for repeated measures with incidental parameters. Annals of the Institute of Statistical Mathematics, 63(3):535–558, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0243-3.

Godbole:2011:BCC

[2799] A. P. Godbole, M. V. Koutras, and F. S. Milienos. Binary consecutive covering arrays. Annals of the Institute of Statistical Mathematics, 63(3):559–584, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-009-
                                        0240-6.

Antoniadis:2011:PLR

[2800] Anestis Antoniadis, Irène Gijbels, and Mila Nikolova. Penalized likelihood regression for generalized linear models with non-quadratic penalties. Annals of the Institute of Statistical Mathematics, 63(3):585–615, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0242-4.

Maesono:2011:EEK

[2801] Yoshihiko Maesono and Spiridon Penev. Edgeworth expansion for the kernel quantile estimator. Annals of the Institute of Statistical Mathematics, 63(3):617–644, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-009-0241-5.

Anonymous:2011:HCc

[2802] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(3):??, June 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Sugimoto:2011:WTV

[2803] Tomoyuki Sugimoto. A Wald-type variance estimation for the nonparametric distribution estimators for doubly censored data. Annals of the Institute of Statistical Mathematics, 63 (4):645–670, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.

1007/s10463-009-0251-3. See erra-

tum [2804].

Sugimoto:2011:EWT

[2804] Tomoyuki Sugimoto. Erratum to:

“A Wald-type variance estimation for the nonparametric distribution estimators for doubly censored data”. Annals of the Institute of Statistical Mathematics, 63(4):671–674, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10463-
011-0331-z.pdf. See [2803].

Sei:2011:GMM

[2805] Tomonari Sei. Gradient modeling for multivariate quantitative data. Annals of the Institute of Statistical Mathematics, 63(4):675–688, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0261-1.

Hayashi:2011:LDT

[2806] Masahito Hayashi. Large deviation theory for non-regular location shift family. Annals of the Institute of Statistical Mathematics, 63(4):689–716, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0254-0.

Ciuperca:2011:ENR

[2807] Gabriela Ciuperca. Estimating nonlinear regression with and without change-points by the LAD method. Annals of the Institute of Statistical Mathematics, 63(4):717–743, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0256-y.

Chida:2011:LSI

[2808] Satoshi Chida and Naoto Miyoshi. Limiting size index distributions for ball-bin models with Zipf-type frequencies. Annals of the Institute of Statistical Mathematics, 63(4):745–768, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0276-7.

Frey:2011:CEU

[2809] Jesse Frey and Omer Ozturk. Constrained estimation using judgment post-stratification. Annals of the Institute of Statistical Mathematics, 63 (4):769–789, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0255-z.

Zhao:2011:FIW

[2810] Yanxing Zhao and H. N. Nagaraja. Fisher information in window censored renewal process data and its applications. Annals of the Institute of Statistical Mathematics, 63(4):791–825, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0252-2.

Ho:2011:BIB

[2811] Man-Wai Ho. On Bayes inference for a bathtub failure rate via S-paths. Annals of the Institute of Statistical Mathematics, 63(4):827–850, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0253-1.

Luati:2011:EWL

[2812] Alessandra Luati and Tommaso Proietti. On the equivalence of the weighted least squares and the generalised least squares estimators, with applications to kernel smoothing. Annals of the Institute of Statistical Mathematics, 63 (4):851–871, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0267-8.

Anonymous:2011:HCd

[2813] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(4):??, August 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Takemura:2011:GZO

[2814] Akimichi Takemura, Vladimir Vovk, and Glenn Shafer. The generality of the zero-one laws. Annals of the Institute of Statistical Mathematics, 63 (5):873–885, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0262-0.

Hsieh:2011:CLE

[2815] S. H. Hsieh, S. M. Lee, P. S. Shen, and M. F. Liu. Conditional likelihood estimation and efficiency comparisons in proportional odds model with missing covariates. Annals of the Institute of Statistical Mathematics, 63 (5):887–921, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0266-9.

Pace:2011:APL

[2816] Luigi Pace, Alessandra Salvan, and Laura Ventura. Adjustments of profile likelihood through predictive densities. Annals of the Institute of Statistical Mathematics, 63(5):923–937, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0274-9.

Munk:2011:NSG

[2817] Axel Munk, Jean-Pierre Stockis, Janis Valeinis, and Götz Giese. Neyman smooth goodness-of-fit tests for the marginal distribution of dependent data. Annals of the Institute of Statistical Mathematics, 63(5):939–959, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0260-2.

Komori:2011:BMM

[2818] Osamu Komori. A boosting method for maximization of the area under the ROC curve. Annals of the Institute of Statistical Mathematics, 63 (5):961–979, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-009-0264-y.

Dette:2011:ODS

[2819] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal design for smoothing splines. Annals of the Institute of Statistical Mathematics, 63(5):981–1003, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-009-
0265-x.

Ren:2011:FLI

[2820] Jian-Jian Ren and Mai Zhou. Full likelihood inferences in the Cox model: an empirical likelihood approach. Annals of the Institute of Statistical Mathematics, 63(5):1005–1018, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0272-y.

Bagkavos:2011:LLH

[2821] Dimitrios Bagkavos. Local linear hazard rate estimation and bandwidth selection. Annals of the Institute of Statistical Mathematics, 63(5):1019–1046, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0277-6.

Berlinet:2011:LLR

[2822] A. Berlinet, A. Elamine, and A. Mas.

Local linear regression for functional data. Annals of the Institute of Statistical Mathematics, 63(5):1047–1075, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0275-8.

Anonymous:2011:HCe

[2823] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(5):??, October 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Varron:2011:UBE

[2824] Davit Varron and Ingrid Van Keilegom. Uniform in bandwidth exact rates for a class of kernel estimators. Annals of the Institute of Statistical Mathematics, 63(6):1077–1102, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0286-5.

Chang:2011:TSM

[2825] Chien-Hsun Chang and Frederic Paik Schoenberg. Testing separability in marked multidimensional point processes with covariates. Annals of the Institute of Statistical Mathematics, 63(6):1103–1122, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0284-7.

Jiang:2011:BEF

[2826] Jiming Jiang and En-Tzu Tang. The best EBLUP in the Fay–Herriot model. Annals of the Institute of Statistical Mathematics, 63(6):1123–1140, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0281-x.

Vexler:2011:OAH

[2827] Albert Vexler and Sergey Tarima. An optimal approach for hypothesis testing in the presence of incomplete data. Annals of the Institute of Statistical Mathematics, 63(6):1141–1163, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0270-0.

Pinheiro:2011:CAN

[2828] Aluísio Pinheiro, Pranab Kumar Sen, and Hildete P. Pinheiro. A class of asymptotically normal degenerate quasi U-statistics. Annals of the Institute of Statistical Mathematics, 63(6): 1165–1182, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0271-z.

Detais:2011:MLE

[2829] Amélie Detais and Jean-François Dupuy. Maximum likelihood estimation in a partially observed stratified regression model with censored data. Annals of the Institute of Statistical Mathematics, 63(6):1183–1206, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0273-x.

Shen:2011:NES

[2830] Pao sheng Shen. Nonparametric estimators of the survival function with twice censored data. Annals of the Institute of Statistical Mathematics, 63(6):1207–1219, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0285-6.

deMiranda:2011:EIN

[2831] José Carlos Simon de Miranda and Pedro A. Morettin. Estimation of the intensity of non-homogeneous point processes via wavelets. Annals of the Institute of Statistical Mathematics, 63(6): 1221–1246, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0283-8.

Hirose:2011:EPL

[2832] Yuichi Hirose. Efficiency of profile likelihood in semi-parametric models. Annals of the Institute of Statistical Mathematics, 63(6):1247–1275, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0280-y.

Zhu:2011:VSC

[2833] Li-Ping Zhu, Lin-Yi Qian, and Jin-Guan Lin. Variable selection in a class of single-index models. Annals of the Institute of Statistical Mathematics, 63(6):1277–1293, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0287-4.

Anonymous:2011:HCf

[2834] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 63(6):??, December 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Fushing:2012:SEI

[2835] Hsieh Fushing. Semiparametric efficient inferences for lifetime regression model with time-dependent covariates. Annals of the Institute of Statistical Mathematics, 64(1):1–25, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0301-x.

Mondal:2012:EWV

[2836] Debashis Mondal and Donald B. Percival. M-estimation of wavelet variance. Annals of the Institute of Statistical Mathematics, 64(1):27–53, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0282-9.

Chang:2012:DDG

[2837] Yung-Ming Chang, James C. Fu, and Han-Ying Lin. Distribution and double generating function of number of patterns in a sequence of Markov dependent multistate trials. Annals of the Institute of Statistical Mathematics, 64(1):55–68, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0300-y.

Yu:2012:RPM

[2838] Qiqing Yu, G. Y. C. Wong, Hao Qin, and Jiaping Wang. Random partition masking model for censored and masked competing risks data. Annals of the Institute of Statistical Mathematics, 64(1):69–85, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0303-8.

Lu:2012:MMP

[2839] Shaochuan Lu. Markov modulated Poisson process associated with statedependent marks and its applications to the deep earthquakes. Annals of the Institute of Statistical Mathematics, 64 (1):87–106, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0302-9.

DeOliveira:2012:BAC

[2840] Victor De Oliveira. Bayesian analysis of conditional autoregressive models. Annals of the Institute of Statistical Mathematics, 64(1):107–133, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0298-1.

Harrar:2012:MTF

[2841] Solomon W. Harrar and Arne C. Bathke. A modified two-factor multivariate analysis of variance: asymptotics and small sample approximations. Annals of the Institute of Statistical Mathematics, 64(1): 135–165, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0299-0. See erra-

tum [2888].

Buch-Kromann:2012:MDE

[2842] Tine Buch-Kromann and Jens Perch Nielsen. Multivariate density estimation using dimension reducing information and tail flattening transformations for truncated or censored data. Annals of the Institute of Statistical Mathematics, 64(1):167–192, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0313-6.

Shimizu:2012:LAM

[2843] Yasutaka Shimizu. Local asymptotic mixed normality for discretely observed non-recurrent Ornstein– Uhlenbeck processes. Annals of the Institute of Statistical Mathematics, 64 (1):193–211, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0307-4.

Zhao:2012:VSS

[2844] Peixin Zhao and Liugen Xue. Variable selection in semiparametric regression analysis for longitudinal data. Annals of the Institute of Statistical Mathematics, 64(1):213–231, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0312-7.

Anonymous:2012:HCa

[2845] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(1):??, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Liu:2012:CML

[2846] Song Liu and Yuhong Yang. Combining models in longitudinal data analysis. Annals of the Institute of Statistical Mathematics, 64(2):233–254, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0306-5.

Kato:2012:ANP

[2847] Kengo Kato. Asymptotic normality of Powell’s kernel estimator. Annals of the Institute of Statistical Mathematics, 64(2):255–273, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0310-9.

Wang:2012:MTT

[2848] Jing Wang. Modelling time trend via spline confidence band. Annals of the Institute of Statistical Mathematics, 64(2):275–301, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0311-8.

Balakrishnan:2012:SOS

[2849] N. Balakrishnan, U. Kamps, and M. Kateri. A sequential order statistics approach to step-stress testing. Annals of the Institute of Statistical Mathematics, 64(2):303–318, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0309-2.

Hsu:2012:BEC

[2850] Chih-Wen Hsu, Marick S. Sinay, and John S. J. Hsu. Bayesian estimation of a covariance matrix with flexible prior specification. Annals of the Institute of Statistical Mathematics, 64(2):319–342, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0314-5.

Bobotas:2012:ERT

[2851] Panayiotis Bobotas, George Iliopoulos, and Stavros Kourouklis. Estimating the ratio of two scale parameters: a simple approach. Annals of the Institute of Statistical Mathematics, 64(2):343–357, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0308-3.

Vos:2012:EBD

[2852] Paul Vos and Qiang Wu. Estimators for the binomial distribution that dominate the MLE in terms of Kullback– Leibler risk. Annals of the Institute of Statistical Mathematics, 64(2):359–371, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0316-3.

Lemonte:2012:LPG

[2853] Artur J. Lemonte and Silvia L. P. Ferrari. The local power of the gradient test. Annals of the Institute of Statistical Mathematics, 64(2):373–381, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0315-4.

Kutoyants:2012:ITD

[2854] Yury A. Kutoyants. On identification of the threshold diffusion processes. Annals of the Institute of Statistical Mathematics, 64(2):383–413, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0318-1.

Wang:2012:HFE

[2855] Qihua Wang, Gregg E. Dinse, and Chunling Liu. Hazard function estimation with cause-of-death data missing at random. Annals of the Institute of Statistical Mathematics, 64(2):415–438, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0317-2.

Frey:2012:CNE

[2856] Jesse Frey. Constrained nonparametric estimation of the mean and the CDF using ranked-set sampling with a covariate. Annals of the Institute of Statistical Mathematics, 64(2):439–456, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0326-9.

Anonymous:2012:HCb

[2857] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(2):??, April 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Hoshino:2012:RPS

[2858] Nobuaki Hoshino. Random partitioning over a sparse contingency table. Annals of the Institute of Statistical Mathematics, 64(3):457–474, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0327-8.

Abarin:2012:IVA

[2859] Taraneh Abarin and Liqun Wang. Instrumental variable approach to covariate measurement error in generalized linear models. Annals of the Institute of Statistical Mathematics, 64(3):475–493, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-010-0319-0.

Smith:2012:IRP

[2860] Jim Q. Smith and Fabio Rigat. Isoseparation and robustness in parametric Bayesian inference. Annals of the Institute of Statistical Mathematics, 64(3):495–519, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0334-9.

Wang:2012:MMH

[2861] Ting Wang, Mark Bebbington, and David Harte. Markov-modulated Hawkes process with stepwise decay. Annals of the Institute of Statistical Mathematics, 64(3):521–544, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0320-7.

Shimizu:2012:EPD

[2862] Yasutaka Shimizu. Estimation of parameters for discretely observed diffusion processes with a variety of rates for information. Annals of the Institute of Statistical Mathematics, 64(3):545–575, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-010-
                                        0323-4.

Yue:2012:PBA

[2863] Yu Ryan Yue, Paul L. Speckman, and Dongchu Sun. Priors for Bayesian adaptive spline smoothing. Annals of the Institute of Statistical Mathematics, 64(3):577–613, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-010-0321-6.

Takazawa:2012:EIL

[2864] Shin ichiro Takazawa. Exponential inequalities and the law of the iterated logarithm in the unbounded forecasting game. Annals of the Institute of Statistical Mathematics, 64(3):615–632, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-010-
0322-5.

Aki:2012:SMD

[2865] Sigeo Aki. Statistical modeling for discrete patterns in a sequence of exchangeable trials. Annals of the Institute of Statistical Mathematics, 64 (3):633–655, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0325-x.

Dutta:2012:RCU

[2866] Subhajit Dutta and Anil K. Ghosh.

On robust classification using projection depth. Annals of the Institute of Statistical Mathematics, 64(3):657–676, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0324-y.

Nelsen:2012:DDM

[2867] Roger B. Nelsen and Manuel Úbeda-Flores. Directional dependence in multivariate distributions. Annals of the Institute of Statistical Mathematics, 64(3):677–685, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0329-6.

Anonymous:2012:HCc

[2868] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(3):??, June 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Bhattacharya:2012:SCN

[2869] Abhishek Bhattacharya and David B. Dunson. Strong consistency of nonparametric Bayes density estimation on compact metric spaces with applications to specific manifolds. Annals of the Institute of Statistical Mathematics, 64(4):687–714, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0341-x.

Fan:2012:OIP

[2870] Chunpeng Fan, Jason P. Fine, and Jong-Hyeon Jeong. Optimal inferences for proportional hazards model with parametric covariate transformations. Annals of the Institute of Statistical Mathematics, 64(4):715–736, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-011-
                                        0330-0.

Azzalini:2012:SMU

[2871] Adelchi Azzalini. Selection models under generalized symmetry settings. Annals of the Institute of Statistical Mathematics, 64(4):737–750, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-011-
                                        0328-7.

Kim:2012:ESO

[2872] Mijeong Kim and Yanyuan Ma. The efficiency of the second-order nonlinear least squares estimator and its extension. Annals of the Institute of Statistical Mathematics, 64(4):751–764, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0332-y.

Liang:2012:ELC

[2873] Han-Ying Liang and Jacobo de Uña-Álvarez. Empirical likelihood for conditional quantile with left-truncated and dependent data. Annals of the Institute of Statistical Mathematics, 64 (4):765–790, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0335-8.

Chretien:2012:SAP

[2874] Stéphane Chrétien, Alfred Hero, and Hervé Perdry. Space alternating penalized Kullback proximal point algorithms for maximizing likelihood with nondifferentiable penalty. Annals of the Institute of Statistical Mathematics, 64 (4):791–809, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0333-x.

Genest:2012:TSB

[2875] Christian Genest, Johanna Neslehová, and Jean-François Quessy. Tests of symmetry for bivariate copulas. Annals of the Institute of Statistical Mathematics, 64(4):811–834, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-011-
                                        0337-6.

Sun:2012:EMS

[2876] Yan Sun, Jialiang Li, and Wenyang Zhang. Estimation and model selection in a class of semiparametric models for cluster data. Annals of the Institute of Statistical Mathematics, 64 (4):835–856, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0342-9.

Azzalini:2012:SPS

[2877] Adelchi Azzalini and Giuliana Regoli. Some properties of skew-symmetric distributions. Annals of the Institute of Statistical Mathematics, 64(4):857–879, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-011-
                                        0338-5.

Anonymous:2012:HCd

[2878] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(4):??, August 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Kunitomo:2012:OML

[2879] Naoto Kunitomo. An optimal modification of the LIML estimation for many instruments and persistent heteroscedasticity. Annals of the Institute of Statistical Mathematics, 64 (5):881–910, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0336-7.

Negri:2012:ADF

[2880] Ilia Negri and Yoichi Nishiyama.

Asymptotically distribution free test for parameter change in a diffusion process model. Annals of the Institute of Statistical Mathematics, 64 (5):911–918, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0345-6.

Leblanc:2012:EDF

[2881] Alexandre Leblanc. On estimating distribution functions using Bernstein polynomials. Annals of the Institute of Statistical Mathematics, 64(5):919–943, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0339-4.

Li:2012:OED

[2882] Gang Li. Optimal and efficient designs for Gompertz regression models. Annals of the Institute of Statistical Mathematics, 64(5):945–957, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-011-
0340-y.

Diao:2012:GTC

[2883] Guoqing Diao and Guosheng Yin. A general transformation class of semiparametric cure rate frailty models. Annals of the Institute of Statistical Mathematics, 64(5):959–989, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-012-
                                        0354-0.

Fujii:2012:SPN

[2884] Takayuki Fujii and Yoichi Nishiyama. Some problems in nonparametric inference for the stress release process related to the local time. Annals of the Institute of Statistical Mathematics, 64 (5):991–1007, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0344-7.

Sugiyama:2012:DRM

[2885] Masashi Sugiyama, Taiji Suzuki, and Takafumi Kanamori. Density-ratio matching under the Bregman divergence: a unified framework of densityratio estimation. Annals of the Institute of Statistical Mathematics, 64 (5):1009–1044, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-011-0343-8.

Dette:2012:TCC

[2886] Holger Dette, Mareen Marchlewski, and Jens Wagener. Testing for a constant coefficient of variation in nonparametric regression by empirical processes. Annals of the Institute of Statistical Mathematics, 64(5):1045–1070, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-011-
                                        0346-5.

Falk:2012:ACD

[2887] Michael Falk and Diana Tichy. Asymptotic conditional distribution of exceedance counts: fragility index with different margins. Annals of the Institute of Statistical Mathematics, 64 (5):1071–1085, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-011-0348-3.

Harrar:2012:EMT

[2888] Solomon W. Harrar and Arne C.

Bathke. Erratum to: “A modified twofactor multivariate analysis of variance: asymptotics and small sample approximations”. Annals of the Institute of Statistical Mathematics, 64(5):1087, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10463-
012-0364-y.pdf. See [2841].

Anonymous:2012:HCe

[2889] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(5):??, October 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Brockwell:2012:SSS

[2890] Peter Brockwell, Alexander Lindner, and Bernd Vollenbröker. Strictly stationary solutions of multivariate ARMA equations with i.i.d. noise. Annals of the Institute of Statistical Mathematics, 64(6):1089–1119, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0357-x.

Akahira:2012:LIS

[2891] Masafumi Akahira, Hyo Gyeong Kim, and Nao Ohyauchi. Loss of information of a statistic for a family of nonregular distributions, II: more general case. Annals of the Institute of Statistical Mathematics, 64(6):1121–1138, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-011-
                                        0347-4.

Bissiri:2012:CIP

[2892] Pier Giovanni Bissiri and Stephen G. Walker. Converting information into probability measures with the Kullback–Leibler divergence. Annals of the Institute of Statistical Mathematics, 64(6):1139–1160, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0350-4.

Reiss:2012:RBI

[2893] Philip T. Reiss, Lei Huang, Joseph E. Cavanaugh, and Amy Krain Roy. Resampling-based information criteria for best-subset regression. Annals of the Institute of Statistical Mathematics, 64(6):1161–1186, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0353-1.

Grane:2012:EGF

[2894] Aurea Grané. Exact goodness-of-fit tests for censored data. Annals of the Institute of Statistical Mathematics, 64(6):1187–1203, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0356-y.

Fokianos:2012:NPA

[2895] Konstantinos Fokianos and Dag Tjøstheim. Nonlinear Poisson autoregression. Annals of the Institute of Statistical Mathematics, 64(6): 1205–1225, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0351-3.

Huckemann:2012:MMS

[2896] Stephan F. Huckemann. On the meaning of mean shape: manifold stability, locus and the two sample test. Annals of the Institute of Statistical Mathematics, 64(6):1227–1259, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0352-2.

Helmers:2012:PCP

[2897] Roelof Helmers and I. Wayan Mangku.

Predicting a cyclic Poisson process. Annals of the Institute of Statistical Mathematics, 64(6):1261–1279, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0349-x.

Anonymous:2012:HCf

[2898] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 64(6):??, December 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Qin:2013:ELB

[2899] Gengsheng Qin, Baoying Yang, and Nelly E. Belinga-Hall. Empirical likelihood-based inferences for the Lorenz curve. Annals of the Institute of Statistical Mathematics, 65 (1):1–21, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0355-z.

Honda:2013:NQR

[2900] Toshio Honda. Nonparametric quantile regression with heavy-tailed and strongly dependent errors. Annals of the Institute of Statistical Mathematics, 65(1):23–47, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0359-8.

Joutard:2013:SLD

[2901] Cyrille Joutard. Strong large deviations for arbitrary sequences of random variables. Annals of the Institute of Statistical Mathematics, 65 (1):49–67, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0361-1.

Yatracos:2013:EPB

[2902] Yannis G. Yatracos. Equal percent bias reduction and variance proportionate modifying properties with mean-covariance preserving matching.

Annals of the Institute of Statistical Mathematics, 65(1):69–87, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0358-9.

Zhang:2013:LSE

[2903] Shibin Zhang and Xinsheng Zhang.

A least squares estimator for discretely observed Ornstein–Uhlenbeck processes driven by symmetric αstable motions. Annals of the Institute of Statistical Mathematics, 65 (1):89–103, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0362-0.

Lai:2013:ECS

[2904] P. Y. Lai and Stephen M. S. Lee. Estimation of central shapes of error distributions in linear regression problems. Annals of the Institute of Statistical Mathematics, 65(1):105–124, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0360-2.

Landajo:2013:NPL

[2905] Manuel Landajo and María José Presno. Nonparametric pseudo-Lagrange multiplier stationarity testing. Annals of the Institute of Statistical Mathematics, 65(1):125–147, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0363-z.

Fan:2013:MWT

[2906] Jie Fan and Somnath Datta. On Mann– Whitney tests for comparing sojourn time distributions when the transition times are right censored. Annals of the Institute of Statistical Mathematics, 65 (1):149–166, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0378-5.

Maesono:2013:ICI

[2907] Yoshihiko Maesono and Spiridon Penev. Improved confidence intervals for quantiles. Annals of the Institute of Statistical Mathematics, 65 (1):167–189, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0369-6.

Ogawa:2013:GBU

[2908] Mitsunori Ogawa, Hisayuki Hara, and Akimichi Takemura. Graver basis for an undirected graph and its application to testing the beta model of random graphs. Annals of the Institute of Statistical Mathematics, 65 (1):191–212, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0367-8.

Anonymous:2013:HCa

[2909] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 65(1):??, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Lee:2013:MDP

[2910] Sangyeol Lee and Junmo Song. Minimum density power divergence estimator for diffusion processes. Annals of the Institute of Statistical Mathematics, 65(2):213–236, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0366-9.

Zhang:2013:PLS

[2911] Jun Zhang, Yao Yu, Li-Xing Zhu, and Hua Liang. Partial linear single index models with distortion measurement errors. Annals of the Institute of Statistical Mathematics, 65(2):237–267, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0371-z.

Pistone:2013:ARM

[2912] Giovanni Pistone and Maria Piera Rogantin. The algebra of reversible Markov chains. Annals of the Institute of Statistical Mathematics, 65 (2):269–293, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0368-7.

Qian:2013:MSS

[2913] Wei Qian and Yuhong Yang. Model selection via standard error adjusted adaptive lasso. Annals of the Institute of Statistical Mathematics, 65 (2):295–318, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0370-0.

Basu:2013:TSH

[2914] A. Basu, A. Mandal, N. Martin, and L. Pardo. Testing statistical hypotheses based on the density power divergence. Annals of the Institute of Statistical Mathematics, 65(2):319–348, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0372-y.

Leucht:2013:DSU

[2915] Anne Leucht and Michael H. Neumann. Degenerate U- and V -statistics under ergodicity: asymptotics, bootstrap and applications in statistics. Annals of the Institute of Statistical Mathematics, 65(2):349–386, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0374-9.

Prokesova:2013:APL

[2916] Michaela Prokesová and Eva B. Vedel Jensen. Asymptotic Palm likelihood theory for stationary point processes. Annals of the Institute of Statistical Mathematics, 65(2):387–412, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0376-7.

Anonymous:2013:HCb

[2917] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 65(2):??, April 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Martin:2013:ICP

[2918] James S. Martin, Ajay Jasra, and Emma McCoy. Inference for a class of partially observed point process models. Annals of the Institute of Statistical Mathematics, 65(3):413–437, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0375-8.

McElroy:2013:FCT

[2919] Tucker S. McElroy. Forecasting continuous-time processes with applications to signal extraction. Annals of the Institute of Statistical Mathematics, 65(3):439–456, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0373-x.

Ren:2013:OBA

[2920] Cuirong Ren and Dongchu Sun. Objective Bayesian analysis for CAR models. Annals of the Institute of Statistical Mathematics, 65(3):457–472, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0377-6.

Xu:2013:CAP

[2921] Wangli Xu and Xu Guo. Checking the adequacy of partial linear models with missing covariates at random. Annals of the Institute of Statistical Mathematics, 65(3):473–490, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0379-4.

Andrade:2013:MCI

[2922] J. A. A. Andrade and Edward Omey. Modelling conflicting information using subexponential distributions and related classes. Annals of the Institute of Statistical Mathematics, 65(3):491–511, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0380-y.

Hsieh:2013:REF

[2923] Yu-Fei Hsieh and Tung-Lung Wu. Recursive equations in finite Markov chain imbedding. Annals of the Institute of Statistical Mathematics, 65(3):513–527, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0381-x.

Liu:2013:CEL

[2924] Yukun Liu, Changliang Zou, and Zhaojun Wang. Calibration of the empirical likelihood for high-dimensional data. Annals of the Institute of Statistical Mathematics, 65(3):529–550, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0384-7.

Frey:2013:VEU

[2925] Jesse Frey and Timothy G. Feeman. Variance estimation using judgment post-stratification. Annals of the Institute of Statistical Mathematics, 65 (3):551–569, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0386-5.

Aki:2013:CCP

[2926] Sigeo Aki and Katuomi Hirano.

Coupon collector’s problems with statistical applications to rankings. Annals of the Institute of Statistical Mathematics, 65(3):571–587, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0382-9.

Yu:2013:NEE

[2927] Keming Yu, Bing Xing Wang, and Valentin Patilea. New estimating equation approaches with application in lifetime data analysis. Annals of the Institute of Statistical Mathematics, 65(3):589–615, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0385-6.

Anonymous:2013:HCc

[2928] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 65(3):??, June 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Zheng:2013:REJ

[2929] Xueying Zheng, Wing Kam Fung, and Zhongyi Zhu. Robust estimation in joint mean-covariance regression model for longitudinal data. Annals of the Institute of Statistical Mathematics, 65 (4):617–638, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0383-8.

Tang:2013:ELS

[2930] Nian-Sheng Tang and Pu-Ying Zhao. Empirical likelihood semiparametric nonlinear regression analysis for longitudinal data with responses missing at random. Annals of the Institute of Statistical Mathematics, 65 (4):639–665, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0387-4.

Maruri-Aguilar:2013:ADE

[2931] Hugo Maruri-Aguilar, Eduardo Sáenz de Cabezón, and Henry P. Wynn. Alexander duality in experimental designs. Annals of the Institute of Statistical Mathematics, 65(4):667–686, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0390-9.

Afendras:2013:UEV

[2932] Giorgos Afendras. Unified extension of variance bounds for integrated Pearson family. Annals of the Institute of Statistical Mathematics, 65(4):687–702, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0388-3.

Macci:2013:LDP

[2933] Claudio Macci and Stefano Trapani.

Large deviations for posterior distributions on the parameter of a multivariate AR(p) process. Annals of the Institute of Statistical Mathematics, 65 (4):703–719, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0389-2.

Bravo:2013:PLV

[2934] Francesco Bravo. Partially linear varying coefficient models with missing at random responses. Annals of the Institute of Statistical Mathematics, 65 (4):721–762, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0391-8.

Xi:2013:ENZ

[2935] Jing Xi, Ruriko Yoshida, and David Haws. Estimating the number of zeroone multi-way tables via sequential importance sampling. Annals of the Institute of Statistical Mathematics, 65 (4):763–783, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-012-0392-7.

Chen:2013:ESM

[2936] Song Xi Chen and Ingrid Van Keilegom. Estimation in semiparametric models with missing data. Annals of the Institute of Statistical Mathematics, 65(4):785–805, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-012-0393-6.

Anonymous:2013:HCd

[2937] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 65(4):??, August 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Shen:2013:CRH

[2938] Xiaotong Shen, Wei Pan, Yunzhang Zhu, and Hui Zhou. On constrained and regularized high-dimensional regression. Annals of the Institute of Statistical Mathematics, 65(5):807–832, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0396-3.

Abe:2013:ECD

[2939] Toshihiro Abe, Arthur Pewsey, and Kunio Shimizu. Extending circular distributions through transformation of argument. Annals of the Institute of Statistical Mathematics, 65(5):833–858, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-012-
                                        0394-5.

Sugimoto:2013:ADN

[2940] Tomoyuki Sugimoto. Asymptotic distribution of the nonparametric distribution estimator based on a martingale approach in doubly censored data. Annals of the Institute of Statistical Mathematics, 65(5):859–888, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-012-
0395-4.

Ahn:2013:CRU

[2941] Kwang Woo Ahn and Kung-Sik Chan.

On the convergence rate of the unscented transformation. Annals of the Institute of Statistical Mathematics, 65 (5):889–912, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0397-x.

Hirano:2013:CTP

[2942] Toshihiro Hirano and Yoshihiro Yajima. Covariance tapering for prediction of large spatial data sets in transformed random fields. Annals of the Institute of Statistical Mathematics, 65 (5):913–939, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0399-8.

Ma:2013:MLV

[2943] Chunsheng Ma. Mittag-Leffler vector random fields with Mittag-Leffler direct and cross covariance functions. Annals of the Institute of Statistical Mathematics, 65(5):941–958, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0398-9.

Chigansky:2013:ETA

[2944] P. Chigansky and Yu. A. Kutoyants.

Estimation in threshold autoregressive models with correlated innovations. Annals of the Institute of Statistical Mathematics, 65(5):959–992, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-013-
                                        0402-4.

Liang:2013:OAB

[2945] You Liang, Xikui Wang, and Yanqing Yi. One-armed bandit process with a covariate. Annals of the Institute of Statistical Mathematics, 65 (5):993–1006, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0401-5.

Anonymous:2013:HCe

[2946] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 65(5):??, October 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Pati:2014:BNR

[2947] Debdeep Pati and David B. Dunson. Bayesian nonparametric regression with varying residual density. Annals of the Institute of Statistical Mathematics, 66(1):1–31, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0415-z.

Ahmad:2014:SAH

[2948] M. Rauf Ahmad. A U-statistic approach for a high-dimensional twosample mean testing problem under non-normality and Behrens–Fisher setting. Annals of the Institute of Statistical Mathematics, 66(1):33–61, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0404-2.

Kamatani:2014:LCM

[2949] Kengo Kamatani. Local consistency of Markov chain Monte Carlo methods. Annals of the Institute of Statistical Mathematics, 66(1):63–74, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0403-3.

Brockwell:2014:BCT

[2950] Peter J. Brockwell, Jens-Peter Kreiss, and Tobias Niebuhr. Bootstrapping continuous-time autoregressive processes. Annals of the Institute of Statistical Mathematics, 66(1):75–92, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0406-0.

Fujii:2014:MSA

[2951] Masaaki Fujii. Momentum-space approach to asymptotic expansion for stochastic filtering. Annals of the Institute of Statistical Mathematics, 66 (1):93–120, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0405-1.

Feng:2014:BCS

[2952] Sanying Feng and Liugen Xue. Biascorrected statistical inference for partially linear varying coefficient errorsin-variables models with restricted condition. Annals of the Institute of Statistical Mathematics, 66(1): 121–140, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0407-z.

Vasiliev:2014:TEM

[2953] Vyacheslav A. Vasiliev. A truncated estimation method with guaranteed accuracy. Annals of the Institute of Statistical Mathematics, 66 (1):141–163, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0409-x.

Zhao:2014:REV

[2954] Weihua Zhao, Riquan Zhang, Jicai Liu, and Yazhao Lv. Robust and efficient variable selection for semiparametric partially linear varying coefficient model based on modal regression. Annals of the Institute of Statistical Mathematics, 66(1):165–191, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0410-4.

Beran:2014:HMT

[2955] Jan Beran, Dieter Schell, and Milan Stehlík. The harmonic moment tail index estimator: asymptotic distribution and robustness. Annals of the Institute of Statistical Mathematics, 66 (1):193–220, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0412-2.

Anonymous:2014:HCa

[2956] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 66(1):??, February 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Leng:2014:BAL

[2957] Chenlei Leng, Minh-Ngoc Tran, and David Nott. Bayesian adaptive lasso. Annals of the Institute of Statistical Mathematics, 66(2):221–244, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0429-6.

Xu:2014:OBA

[2958] Chang Xu, Dongchu Sun, and Chong He. Objective Bayesian analysis for a capture-recapture model. Annals of the Institute of Statistical Mathematics, 66(2):245–278, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0413-1.

Lan:2014:TCH

[2959] Wei Lan, Hansheng Wang, and Chih-Ling Tsai. Testing covariates in highdimensional regression. Annals of the Institute of Statistical Mathematics, 66(2):279–301, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0414-0.

Chakraborty:2014:DDI

[2960] Anirvan Chakraborty and Probal Chaudhuri. On data depth in infinite dimensional spaces. Annals of the Institute of Statistical Mathematics, 66(2):303–324, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0416-y.

Yang:2014:JDT

[2961] Yujiao Yang and Qiongxia Song. Jump detection in time series nonparametric regression models: a polynomial spline approach. Annals of the Institute of Statistical Mathematics, 66(2):325–344, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0411-3.

Chaudhuri:2014:QIS

[2962] Sanjay Chaudhuri. Qualitative inequalities for squared partial correlations of a Gaussian random vector. Annals of the Institute of Statistical Mathematics, 66(2):345–367, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0417-x.

Bharath:2014:AEC

[2963] Karthik Bharath, Vladimir Pozdnyakov, and Dipak K. Dey. Asymptotics of the empirical cross-over function. Annals of the Institute of Statistical Mathematics, 66(2):369–382, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10463-013-
0423-z.

Bravo:2014:VCP

[2964] Francesco Bravo. Varying coefficients partially linear models with randomly censored data. Annals of the Institute of Statistical Mathematics, 66 (2):383–412, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0420-2.

Galvao:2014:TLA

[2965] Antonio F. Galvao, Kengo Kato, Gabriel Montes-Rojas, and Jose Olmo. Testing linearity against threshold effects: uniform inference in quantile regression. Annals of the Institute of Statistical Mathematics, 66(2):413–439, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0418-9.

Anonymous:2014:HCb

[2966] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 66(2):??, April 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Yoshida:2014:P

[2967] Ryo Yoshida, Genta Ueno, and Arnaud Doucet. Preface. Annals of the Institute of Statistical Mathematics, 66(3):441–442, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10463-014-0459-8.pdf.

Kitagawa:2014:CAS

[2968] Genshiro Kitagawa. Computational aspects of sequential Monte Carlo filter and smoother. Annals of the Institute of Statistical Mathematics, 66 (3):443–471, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0446-0.

Mukherjee:2014:SVS

[2969] Chiranjit Mukherjee, Prasad S. Kasibhatla, and Mike West. Spatially varying SAR models and Bayesian inference for high-resolution lattice data. Annals of the Institute of Statistical Mathematics, 66(3):473–494, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0426-9.

Nguyen:2014:BNM

[2970] XuanLong Nguyen and Alan E. Gelfand. Bayesian nonparametric modeling for functional analysis of variance. Annals of the Institute of Statistical Mathematics, 66(3):495–526, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0436-7.

Pitt:2014:SLI

[2971] Michael K. Pitt, Sheheryar Malik, and Arnaud Doucet. Simulated likelihood inference for stochastic volatility models using continuous particle filtering. Annals of the Institute of Statistical Mathematics, 66(3):527–552, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0456-y.

Nam:2014:PSM

[2972] Christopher F. H. Nam, John A. D.

Aston, and Adam M. Johansen. Parallel sequential Monte Carlo samplers and estimation of the number of states in a hidden Markov model. Annals of the Institute of Statistical Mathematics, 66(3):553–575, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0450-4.

Finke:2014:SPE

[2973] Axel Finke, Adam M. Johansen, and Dario Spanò. Static-parameter estimation in piecewise deterministic processes using particle Gibbs samplers. Annals of the Institute of Statistical Mathematics, 66(3):577–609, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0455-z.

Iba:2014:MMS

[2974] Yukito Iba, Nen Saito, and Akimasa Kitajima. Multicanonical MCMC for sampling rare events: an illustrative review. Annals of the Institute of Statistical Mathematics, 66(3):611–645, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0460-2.

Anonymous:2014:HCc

[2975] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 66(3):??, June 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Brockwell:2014:RRT

[2976] P. J. Brockwell. Recent results in the theory and applications of CARMA processes. Annals of the Institute of Statistical Mathematics, 66(4):647–685, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0468-7.

Tanaka:2014:IES

[2977] Ushio Tanaka and Yosihiko Ogata. Identification and estimation of superposed Neyman–Scott spatial cluster processes. Annals of the Institute of Statistical Mathematics, 66(4):687–702, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-013-
                                        0431-z.

Bedford:2014:CMI

[2978] Tim Bedford and Kevin J. Wilson. On the construction of minimum information bivariate copula families. Annals of the Institute of Statistical Mathematics, 66(4):703–723, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-013-
                                        0422-0.

Kuriki:2014:ATP

[2979] Satoshi Kuriki, Yoshiaki Harushima, Hironori Fujisawa, and Nori Kurata. Approximate tail probabilities of the maximum of a chi-square field on multidimensional lattice points and their applications to detection of loci interactions. Annals of the Institute of Statistical Mathematics, 66(4):725–757, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0419-8.

Balakrishnan:2014:SBS

[2980] N. Balakrishnan, M. V. Koutras, and F. S. Milienos. Some binary start-up demonstration tests and associated inferential methods. Annals of the Institute of Statistical Mathematics, 66 (4):759–787, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0424-y.

Yanagimoto:2014:PBP

[2981] Takemi Yanagimoto and Toshio Ohnishi. Permissible boundary prior function as a virtually proper prior density. Annals of the Institute of Statistical Mathematics, 66(4):789–809, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0421-1.

Jozani:2014:ENN

[2982] Mohammad Jafari Jozani, Éric Marchand, and William E. Strawderman.

Estimation of a non-negative location parameter with unknown scale. Annals of the Institute of Statistical Mathematics, 66(4):811–832, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0425-x.

Anonymous:2014:HCd

[2983] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 66(4):??, August 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Sibuya:2014:PEP

[2984] Masaaki Sibuya. Prediction in Ewens– Pitman sampling formula and random samples from number partitions. Annals of the Institute of Statistical Mathematics, 66(5):833–864, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0427-8.

Wu:2014:REV

[2985] Tiee-Jian Wu, Chih-Yuan Hsu, Huang-Yu Chen, and Hui-Chun Yu. Root n estimates of vectors of integrated density partial derivative functionals. Annals of the Institute of Statistical Mathematics, 66(5):865–895, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0428-7.

Gupta:2014:POF

[2986] Ramesh C. Gupta and Cheng Peng.

Proportional odds frailty model and stochastic comparisons. Annals of the Institute of Statistical Mathematics, 66 (5):897–912, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0432-y.

Ren:2014:ELB

[2987] Jian-Jian Ren and Tonya Riddlesworth. Empirical likelihood bivariate nonparametric maximum likelihood estimator with right censored data. Annals of the Institute of Statistical Mathematics, 66(5):913–930, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0433-x.

Ma:2014:EIP

[2988] Yunbei Ma, Alan T. K. Wan, Xuerong Chen, and Yong Zhou. On estimation and inference in a partially linear hazard model with varying coefficients. Annals of the Institute of Statistical Mathematics, 66(5):931–960, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0430-0.

Paige:2014:ESB

[2989] Robert L. Paige, A. Alexandre Trindade, and R. Indika P. Wickramasinghe. Extensions of saddlepoint-based bootstrap inference. Annals of the Institute of Statistical Mathematics, 66 (5):961–981, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0434-9.

Aoshima:2014:DBM

[2990] Makoto Aoshima and Kazuyoshi Yata. A distance-based, misclassification rate adjusted classifier for multiclass, highdimensional data. Annals of the Institute of Statistical Mathematics, 66 (5):983–1010, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0435-8.

Anonymous:2014:HCe

[2991] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 66(5):??, October 2014. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Colling:2015:EED

[2992] Benjamin Colling, Cédric Heuchenne, Rawane Samb, and Ingrid Van Keilegom. Estimation of the error density in a semiparametric transformation model. Annals of the Institute of Statistical Mathematics, 67(1):1–18, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-013-
                                        0441-x.

Podgorski:2015:MLO

[2993] Krzysztof Podgórski and Jonas Wallin. Maximizing leave-one-out likelihood for the location parameter of unbounded densities. Annals of the Institute of Statistical Mathematics, 67 (1):19–38, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0437-6.

Akashi:2015:LIM

[2994] Kentaro Akashi and Naoto Kunitomo.

The limited information maximum likelihood approach to dynamic panel structural equation models. Annals of the Institute of Statistical Mathematics, 67(1):39–73, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0438-5.

Kuang:2015:MLE

[2995] Nenghui Kuang and Huantian Xie.

Maximum likelihood estimator for the sub-fractional Brownian motion approximated by a random walk. Annals of the Institute of Statistical Mathematics, 67(1):75–91, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-013-
0439-4.

Cheng:2015:SEE

[2996] Guang Cheng, Hao Helen Zhang, and Zuofeng Shang. Sparse and efficient estimation for partial spline models with increasing dimension. Annals of the Institute of Statistical Mathematics, 67 (1):93–127, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0440-y.

Thavaneswaran:2015:GDM

[2997] Aerambamoorthy Thavaneswaran, Nalini Ravishanker, and You Liang. Generalized duration models and optimal estimation using estimating functions. Annals of the Institute of Statistical Mathematics, 67(1):129–156, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-013-
                                        0442-9.

Chen:2015:CCB

[2998] Ting-Li Chen. On the convergence and consistency of the blurring meanshift process. Annals of the Institute of Statistical Mathematics, 67(1): 157–176, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0443-8.

Hotz:2015:IMC

[2999] T. Hotz and S. Huckemann. Intrinsic means on the circle: uniqueness, locus and asymptotics. Annals of the Institute of Statistical Mathematics, 67 (1):177–193, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-013-0444-7.

Elijio:2015:CPA

[3000] A. Elijio and V. Cekanavicius. Compound Poisson approximation to weighted sums of symmetric discrete variables. Annals of the Institute of Statistical Mathematics, 67 (1):195–210, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-013-0445-6.

Anonymous:2015:HC

[3001] Anonymous. Help & contacts. Annals of the Institute of Statistical Mathematics, 67(1):??, February 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Jiang:2015:EES

[3002] Binyan Jiang. An empirical estimator for the sparsity of a large covariance matrix under multivariate normal assumptions. Annals of the Institute of Statistical Mathematics, 67(2):211–227, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0447-z.

Guo:2015:MCP

[3003] Xu Guo, Wangli Xu, and Lixing Zhu.

Model checking for parametric regressions with response missing at random. Annals of the Institute of Statistical Mathematics, 67(2):229–259, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0451-3.

Tsukuma:2015:MER

[3004] Hisayuki Tsukuma and Tatsuya Kubokawa. Minimaxity in estimation of restricted and non-restricted scale parameter matrices. Annals of the Institute of Statistical Mathematics, 67(2):261–285, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0449-x.

Luo:2015:EBI

[3005] Shan Luo, Jinfeng Xu, and Zehua Chen. Extended Bayesian information criterion in the Cox model with a highdimensional feature space. Annals of the Institute of Statistical Mathematics, 67(2):287–311, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0448-y.

Sun:2015:NHM

[3006] Yan Sun and Dan Ralescu. A normal hierarchical model and minimum contrast estimation for random intervals. Annals of the Institute of Statistical Mathematics, 67(2):313–333, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0453-1.

Terada:2015:SCF

[3007] Yoshikazu Terada. Strong consistency of factorial K-means clustering. Annals of the Institute of Statistical Mathematics, 67(2):335–357, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0454-0.

Li:2015:GEB

[3008] Weiming Li and Jianfeng Yao. On generalized expectation-based estimation of a population spectral distribution from high-dimensional data. Annals of the Institute of Statistical Mathematics, 67(2):359–373, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0452-2.

Lv:2015:QRV

[3009] Yazhao Lv, Riquan Zhang, Weihua Zhao, and Jicai Liu. Quantile regression and variable selection of partial linear single-index model. Annals of the Institute of Statistical Mathematics, 67(2):375–409, April 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0457-x.

Ojeda:2015:TRM

[3010] J. L. Ojeda, W. González-Manteiga, and J. A. Cristóbal. Testing regression models with selection-biased data. Annals of the Institute of Statistical Mathematics, 67(3):411–436, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0463-z.

Dette:2015:TAN

[3011] Holger Dette, Matthias Guhlich, and Natalie Neumeyer. Testing for additivity in nonparametric quantile regression. Annals of the Institute of Statistical Mathematics, 67(3):437–477, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0461-1.

Goegebeur:2015:RCW

[3012] Yuri Goegebeur, Armelle Guillou, and Théo Rietsch. Robust conditional Weibull-type estimation. Annals of the Institute of Statistical Mathematics, 67(3):479–514, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0458-9.

Nagaraja:2015:SAO

[3013] H. N. Nagaraja, Karthik Bharath, and Fangyuan Zhang. Spacings around an order statistic. Annals of the Institute of Statistical Mathematics, 67 (3):515–540, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0466-9.

Zhou:2015:ECS

[3014] Haojin Zhou and Tapan K. Nayak. On the equivariance criterion in statistical prediction. Annals of the Institute of Statistical Mathematics, 67(3):541–555, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0464-y.

Kashikar:2015:PPS

[3015] Akanksha S. Kashikar and S. R. Deshmukh. Probabilistic properties of second order branching process. Annals of the Institute of Statistical Mathematics, 67(3):557–572, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0462-0.

Pewsey:2015:SAL

[3016] Arthur Pewsey and Toshihiro Abe.

The sinh-arcsinhed logistic family of distributions: properties and inference. Annals of the Institute of Statistical Mathematics, 67(3):573–594, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0465-x.

Doring:2015:SCP

[3017] Maik Döring and Uwe Jensen. Smooth change point estimation in regression models with random design. Annals of the Institute of Statistical Mathematics, 67(3):595–619, June 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0467-8.

Slavkovic:2015:FMW

[3018] Aleksandra Slavković, Xiaotian Zhu, and Sonja Petrović. Fibers of multiway contingency tables given conditionals: relation to marginals, cell bounds and Markov bases. Annals of the Institute of Statistical Mathematics, 67(4):621–648, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0471-z.

Fernandes:2015:TSC

[3019] Marcelo Fernandes, Eduardo F.

Mendes, and Olivier Scaillet. Testing for symmetry and conditional symmetry using asymmetric kernels. Annals of the Institute of Statistical Mathematics, 67(4):649–671, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0469-6.

Yin:2015:MDC

[3020] Yue Yin and Julie Zhou. Minimax design criterion for fractional factorial designs. Annals of the Institute of Statistical Mathematics, 67(4):673–685, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0470-0.

Dinwoodie:2015:ETS

[3021] Ian H. Dinwoodie and Kruti Pandya. Exact tests for singular network data. Annals of the Institute of Statistical Mathematics, 67(4):687–706, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0472-y.

Bibinger:2015:EQC

[3022] Markus Bibinger and Mathias Vetter. Estimating the quadratic covariation of an asynchronously observed semimartingale with jumps. Annals of the Institute of Statistical Mathematics, 67 (4):707–743, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0473-x.

Kim:2015:EIF

[3023] Daeyoung Kim and Bruce G. Lindsay. Empirical identifiability in finite mixture models. Annals of the Institute of Statistical Mathematics, 67 (4):745–772, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0474-9.

Liang:2015:EHM

[3024] Yuli Liang, Dietrich von Rosen, and Tatjana von Rosen. On estimation in hierarchical models with block circular covariance structures. Annals of the Institute of Statistical Mathematics, 67(4):773–791, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0475-8.

Xu:2015:NCP

[3025] Wangli Xu and Lixing Zhu. Nonparametric check for partial linear errorsin-covariables models with validation data. Annals of the Institute of Statistical Mathematics, 67(4):793–815, August 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0476-7.

Wang:2015:PVC

[3026] Xuan Wang, Qihua Wang, and Xiao-Hua Andrew Zhou. Partially varying coefficient single-index additive hazard models. Annals of the Institute of Statistical Mathematics, 67(5):817–841, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0484-7.

Jones:2015:CCC

[3027] M. C. Jones, Arthur Pewsey, and Shogo Kato. On a class of circulas: copulas for circular distributions. Annals of the Institute of Statistical Mathematics, 67 (5):843–862, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0493-6.

Chang:2015:ETO

[3028] Yuan-Tsung Chang and Nobuo Shinozaki. Estimation of two ordered normal means under modified Pitman nearness criterion. Annals of the Institute of Statistical Mathematics, 67 (5):863–883, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0479-4.

Lemonte:2015:LPP

[3029] Artur J. Lemonte. On local power properties of the LR, Wald, score and gradient tests in nonlinear mixedeffects models. Annals of the Institute of Statistical Mathematics, 67 (5):885–895, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0478-5.

Zhao:2015:ECB

[3030] Xiao Bing Zhao and Xian Zhou. Estimation of copula-based models for lifetime medical costs. Annals of the Institute of Statistical Mathematics, 67 (5):897–915, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0477-6.

Dyckerhoff:2015:DBR

[3031] Rainer Dyckerhoff, Christophe Ley, and Davy Paindaveine. Depth-based runs tests for bivariate central symmetry. Annals of the Institute of Statistical Mathematics, 67(5):917–941, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0480-y.

Ninomiya:2015:CPM

[3032] Yoshiyuki Ninomiya. Change-point model selection via AIC. Annals of the Institute of Statistical Mathematics, 67 (5):943–961, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0481-x.

Mukhopadhyay:2015:COB

[3033] Minerva Mukhopadhyay, Tapas Samanta, and Arijit Chakrabarti. On consistency and optimality of Bayesian variable selection based ong-prior in normal linear regression models. Annals of the Institute of Statistical Mathematics, 67 (5):963–997, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0483-8.

Liu:2015:QRL

[3034] Peng Liu, Yixin Wang, and Yong Zhou. Quantile residual lifetime with right-censored and length-biased data.

Annals of the Institute of Statistical Mathematics, 67(5):999–1028, October 2015. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0482-9.

Malinowski:2016:IWM

[3035] Alexander Malinowski, Martin Schlather, and Zhengjun Zhang. Intrinsically weighted means and non-ergodic marked point processes. Annals of the Institute of Statistical Mathematics, 68(1):1–24, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0485-6.

Efromovich:2016:MTN

[3036] Sam Efromovich. Minimax theory of nonparametric hazard rate estimation: efficiency and adaptation. Annals of the Institute of Statistical Mathematics, 68(1):25–75, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0487-4.

Ducharme:2016:CMD

[3037] Gilles R. Ducharme, Pierre Lafaye de Micheaux, and Bastien Marchina. The complex multinormal distribution, quadratic forms in complex random vectors and an omnibus goodness-of-fit test for the complex normal distribution. Annals of the Institute of Statistical Mathematics, 68(1):77–104, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10463-014-
0486-5.

Jimenez-Gamero:2016:FMM

[3038] M. D. Jiménez-Gamero, A. Batsidis, and M. V. Alba-Fernández. Fourier methods for model selection. Annals of the Institute of Statistical Mathematics, 68(1):105–133, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0491-8.

Zhu:2016:TPE

[3039] Xuehu Zhu, Xu Guo, Lu Lin, and Lixing Zhu. Testing for positive expectation dependence. Annals of the Institute of Statistical Mathematics, 68 (1):135–153, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0492-7.

He:2016:ESP

[3040] Yawei He and Zehua Chen. The EBIC and a sequential procedure for feature selection in interactive linear models with high-dimensional data. Annals of the Institute of Statistical Mathematics, 68(1):155–180, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0497-2.

Ma:2016:EIF

[3041] Shujie Ma. Estimation and inference in functional single-index models. Annals of the Institute of Statistical Mathematics, 68(1):181–208, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0488-3.

Proksch:2016:CBM

[3042] Katharina Proksch. On confidence bands for multivariate nonparametric regression. Annals of the Institute of Statistical Mathematics, 68(1): 209–236, February 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0494-5.

Kim:2016:TII

[3043] Moosup Kim and Sangyeol Lee. On the tail index inference for heavy-tailed GARCH-type innovations. Annals of the Institute of Statistical Mathematics, 68(2):237–267, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0495-4.

Huang:2016:PMM

[3044] Zhiyue Huang and Paul Marriott. Parameterizing mixture models with generalized moments. Annals of the Institute of Statistical Mathematics, 68 (2):269–297, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0490-9.

Huang:2016:EPM

[3045] Zhiyue Huang and Paul Marriott. Erratum to: Parameterizing mixture models with generalized moments.

Annals of the Institute of Statistical Mathematics, 68(2):299–300, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-015-
0521-1;   http://link.springer.
com/content/pdf/10.1007/s10463-
015-0521-1.pdf.

Hwang:2016:KEM

[3046] Eunju Hwang and Dong Wan Shin.

Kernel estimators of mode under ψweak dependence. Annals of the Institute of Statistical Mathematics, 68 (2):301–327, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-014-0489-2.

Yuan:2016:ERA

[3047] Ke-Hai Yuan, Wai Chan, and Yubin Tian. Expectation-robust algorithm and estimating equations for means and dispersion matrix with missing data. Annals of the Institute of Statistical Mathematics, 68(2):329–351, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0498-1.

Avendano:2016:SGP

[3048] M. L. Avendaño and M. C. Pardo.

A semiparametric generalized proportional hazards model for right-censored data. Annals of the Institute of Statistical Mathematics, 68(2):353–384, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-014-
                                        0496-3.

Drapatz:2016:SSS

[3049] Martin Drapatz. Strictly stationary solutions of spatial ARMA equations. Annals of the Institute of Statistical Mathematics, 68(2):385–412, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0500-y.

Ghosh:2016:RBE

[3050] Abhik Ghosh and Ayanendranath Basu. Robust Bayes estimation using the density power divergence. Annals of the Institute of Statistical Mathematics, 68(2):413–437, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-014-
                                        0499-0.

Girardin:2016:EDM

[3051] Valérie Girardin and Philippe Regnault. Escort distributions minimizing the Kullback–Leibler divergence for a large deviations principle and tests of entropy level. Annals of the Institute of Statistical Mathematics, 68(2):439–468, April 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-014-0501-x.

Akahira:2016:SOA

[3052] Masafumi Akahira. Second-order asymptotic comparison of the MLE and MCLE of a natural parameter for a truncated exponential family of distributions. Annals of the Institute of Statistical Mathematics, 68(3):469–490, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-014-
0502-9.

Jentsch:2016:BSQ

[3053] Carsten Jentsch and Anne Leucht.

Bootstrapping sample quantiles of discrete data. Annals of the Institute of Statistical Mathematics, 68(3):491–539, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-015-
0503-3.

Banerjee:2016:GSF

[3054] Swarnali Banerjee and Nitis Mukhopadhyay. A general sequential fixedaccuracy confidence interval estimation methodology for a positive parameter: illustrations using health and safety data. Annals of the Institute of Statistical Mathematics, 68(3):541–570, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-015-
0504-2.

Rukhin:2016:DTI

[3055] Andrew L. Rukhin. Decision-theoretic issues in heterogeneity variance estimation. Annals of the Institute of Statistical Mathematics, 68(3):571–588, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-015-
0505-1.

Osorio:2016:IDR

[3056] Felipe Osorio. Influence diagnostics for robust P -splines using scale mixture of normal distributions. Annals of the Institute of Statistical Mathematics, 68(3):589–619, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-015-0506-0.

Lee:2016:PCT

[3057] Sangyeol Lee and Haejune Oh. Parameter change test for autoregressive conditional duration models. Annals of the Institute of Statistical Mathematics, 68(3):621–637, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-015-0541-x.

Bedbur:2016:IMS

[3058] Stefan Bedbur, Marco Burkschat, and Udo Kamps. Inference in a model of successive failures with shape-adjusted hazard rates. Annals of the Institute of Statistical Mathematics, 68(3):639–657, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-015-
                                        0508-y.

Wang:2016:OPB

[3059] Yaping Wang, Mingyao Ai, and Kang Li. Optimality of pairwise blocked definitive screening designs. Annals of the Institute of Statistical Mathematics, 68(3):659–671, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-015-0510-4.

Lin:2016:ORQ

[3060] Liang-Ching Lin and Meihui Guo. Optimal restricted quadratic estimator of integrated volatility. Annals of the Institute of Statistical Mathematics, 68(3):673–703, June 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-015-0507-z.

Kurata:2016:BEE

[3061] Hiroshi Kurata and Shun Matsuura.

Best equivariant estimator of regression coefficients in a seemingly unrelated regression model with known correlation matrix. Annals of the Institute of Statistical Mathematics, 68 (4):705–723, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0512-
2.

Ma:2016:PEG

[3062] Shujie Ma, Zijian Huang, and Chih-Ling Tsai. Parameter estimation for a generalized semiparametric model with repeated measurements. Annals of the Institute of Statistical Mathematics, 68 (4):725–764, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0513-
1.

Li:2016:CES

[3063] Huiqin Li, Zhi Dong Bai, and Jiang Hu. Convergence of empirical spectral distributions of large dimensional quaternion sample covariance matrices. Annals of the Institute of Statistical Mathematics, 68(4):765–785, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0514-0.

Mukerjee:2016:ATA

[3064] Rahul Mukerjee and S. Huda. Approximate theory-aided robust efficient factorial fractions under baseline parametrization. Annals of the Institute of Statistical Mathematics, 68 (4):787–803, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0509-
x.

Ghosh:2016:RAB

[3065] Palash Ghosh and Anup Dewanji. Regression analysis of biased case-control data. Annals of the Institute of Statistical Mathematics, 68(4):805–825, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0511-3.

Bissantz:2016:SBA

[3066] Nicolai Bissantz, Holger Dette, Thimo Hildebrandt, and Kathrin Bissantz. Smooth backfitting in additive inverse regression. Annals of the Institute of Statistical Mathematics, 68(4):827–853, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

com/accesspage/article/10.1007/
s10463-015-0517-x.

Zheng:2016:CRD

[3067] Ming Zheng, Renxin Lin, and Wen Yu.

Competing risks data analysis under the accelerated failure time model with missing cause of failure. Annals of the Institute of Statistical Mathematics, 68 (4):855–876, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0516-
y.

Min:2016:BMS

[3068] Xiaoyi Min and Dongchu Sun. Bayesian model selection for a linear model with grouped covariates. Annals of the Institute of Statistical Mathematics, 68 (4):877–903, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0518-
9.

Cronie:2016:SSI

[3069] O. Cronie and M. N. M. van Lieshout.

Summary statistics for inhomogeneous marked point processes. Annals of the Institute of Statistical Mathematics, 68 (4):905–928, August 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0515-
z.

Kojadinovic:2016:TCS

[3070] Ivan Kojadinovic, Jean-François Quessy, and Tom Rohmer. Testing the constancy of Spearman’s rho in multivariate time series. Annals of the Institute of Statistical Mathematics, 68 (5):929–954, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0520-
2.

Barmi:2016:TUS

[3071] Hammou El Barmi and Ian W. McKeague. Testing for uniform stochastic ordering via empirical likelihood. Annals of the Institute of Statistical Mathematics, 68(5):955–976, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0523-z.

Qin:2016:REG

[3072] Guoyou Qin, Zhongyi Zhu, and Wing K. Fung. Robust estimation of generalized partially linear model for longitudinal data with dropouts. Annals of the Institute of Statistical Mathematics, 68(5):977–1000, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0519-8.

Zhao:2016:AFT

[3073] Mu Zhao, Yixin Wang, and Yong Zhou. Accelerated failure time model with quantile information. Annals of the Institute of Statistical Mathematics, 68 (5):1001–1024, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/

article/10.1007/s10463-015-0522-
0.

Bizjajeva:2016:ASS

[3074] Svetlana Bizjajeva and Jimmy Olsson. Antithetic sampling for sequential Monte Carlo methods with application to state-space models. Annals of the Institute of Statistical Mathematics, 68 (5):1025–1053, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0524-
y.

Aki:2016:MEV

[3075] Sigeo Aki and Katuomi Hirano. On monotonicity of expected values of some run-related distributions. Annals of the Institute of Statistical Mathematics, 68(5):1055–1072, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0525-x.

Emura:2016:SIA

[3076] Takeshi Emura and Weijing Wang.

Semiparametric inference for an accelerated failure time model with dependent truncation. Annals of the Institute of Statistical Mathematics, 68(5): 1073–1094, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0526-
9.

Yang:2016:JEL

[3077] Hanfang Yang, Shen Liu, and Yichuan Zhao. Jackknife empirical likelihood for linear transformation models with right censoring. Annals of the Institute of Statistical Mathematics, 68(5): 1095–1109, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0528-
7.

Lu:2016:SBS

[3078] Minggen Lu and Chin-Shang Li. Spline-based semiparametric estimation of a zero-inflated Poisson regression single-index model. Annals of the Institute of Statistical Mathematics, 68 (5):1111–1134, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0527-
8.

deWet:2016:KRW

[3079] Tertius de Wet, Yuri Goegebeur, Armelle Guillou, and Michael Osmann. Kernel regression with Weibull-type tails. Annals of the Institute of Statistical Mathematics, 68(5):1135–1162, October 2016. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0531-z.

Mano:2017:ESG

[3080] Shuhei Mano. Extreme sizes in Gibbs-type exchangeable random partitions. Annals of the Institute of Statistical Mathematics, 69(1): 1–37, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/

article/10.1007/s10463-015-0530-
0;   http://link.springer.com/article/
10.1007/s10463-015-0530-0.

Ley:2017:EAD

[3081] Christophe Ley, Yvik Swan, and Thomas Verdebout. Efficient ANOVA for directional data. Annals of the Institute of Statistical Mathematics, 69 (1):39–62, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0533-
x;   http://link.springer.com/article/
10.1007/s10463-015-0533-x.

Hu:2017:CSI

[3082] Qinqin Hu and Lu Lin. Conditional sure independence screening by conditional marginal empirical likelihood. Annals of the Institute of Statistical Mathematics, 69 (1):63–96, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0534-
9;   http://link.springer.com/article/
10.1007/s10463-015-0534-9.

Zhang:2017:GVC

[3083] Jun Zhang, Zhenghui Feng, Peirong Xu, and Hua Liang. Generalized varying coefficient partially linear measurement errors models. Annals of the Institute of Statistical Mathematics, 69 (1):97–120, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0532-
y;   http://link.springer.com/article/
10.1007/s10463-015-0532-y.

Guo:2017:SCW

[3084] Huijun Guo and Youming Liu. Strong consistency of wavelet estimators for errors-in-variables regression model. Annals of the Institute of Statistical Mathematics, 69(1): 121–144, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-015-0529-
                                        6;   http://link.springer.com/article/
                                        10.1007/s10463-015-0529-6.

Zhang:2017:CCP

[3085] Liwen Zhang, Huixia Judy Wang, and Zhongyi Zhu. Composite change point estimation for bent line quantile regression. Annals of the Institute of Statistical Mathematics, 69(1): 145–168, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-015-0538-
                                        5;   http://link.springer.com/article/
                                        10.1007/s10463-015-0538-5.

Li:2017:PEE

[3086] Yongjin Li, Qingzhao Zhang, and Qihua Wang. Penalized estimation equation for an extended singleindex model. Annals of the Institute of Statistical Mathematics, 69(1): 169–187, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-015-0544-
                                        7;   http://link.springer.com/article/
                                        10.1007/s10463-015-0544-7.

Bott:2017:NEC

[3087] Ann-Kathrin Bott and Michael Kohler.

Nonparametric estimation of a conditional density. Annals of the Institute of Statistical Mathematics, 69 (1):189–214, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0535-
8;   http://link.springer.com/article/
10.1007/s10463-015-0535-8.

Bodnar:2017:SIE

[3088] Taras Bodnar and Thorsten Dickhaus. On the Simes inequality in elliptical models. Annals of the Institute of Statistical Mathematics, 69 (1):215–230, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0539-
4;   http://link.springer.com/article/
10.1007/s10463-015-0539-4.

Martin:2017:FED

[3089] Donald E. K. Martin and Laurent Noé. Faster exact distributions of pattern statistics through sequential elimination of states. Annals of the Institute of Statistical Mathematics, 69 (1):231–248, February 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0540-
y;   http://link.springer.com/article/
10.1007/s10463-015-0540-y.

Liu:2017:CJE

[3090] Xiaohui Liu, Qihua Wang, and Yi Liu.

A consistent jackknife empirical likelihood test for distribution functions. Annals of the Institute of Statistical Mathematics, 69(2):249–269, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0550-9.

Yuan:2017:SCK

[3091] Ao Yuan, Mihai Giurcanu, George Luta, and Ming T. Tan. U-statistics with conditional kernels for incomplete data models. Annals of the Institute of Statistical Mathematics, 69(2):271–302, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0537-6.

Coeurjolly:2017:MBE

[3092] Jean-François Coeurjolly. Medianbased estimation of the intensity of a spatial point process. Annals of the Institute of Statistical Mathematics, 69(2):303–331, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0536-
7.

Raim:2017:AIM

[3093] Andrew M. Raim, Nagaraj K. Neerchal, and Jorge G. Morel. An approximation to the information matrix of exponential family finite mixtures. Annals of the Institute of Statistical Mathematics, 69(2):333–364, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

com/accesspage/article/10.1007/
s10463-015-0542-9.

Hu:2017:TEH

[3094] Jiang Hu, Zhidong Bai, Chen Wang, and Wei Wang. On testing the equality of high dimensional mean vectors with unequal covariance matrices. Annals of the Institute of Statistical Mathematics, 69(2):365–387, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0543-
8.

Aletti:2017:BBG

[3095] Giacomo Aletti and Matteo Ruffini.

Is the Brownian bridge a good noise model on the boundary of a circle? Annals of the Institute of Statistical Mathematics, 69(2):389–416, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10463-015-0546-5.

Kozyra:2017:LUB

[3096] Pawel Marcin Kozyra and Tomasz Rychlik. Lower and upper bounds on the variances of spacings. Annals of the Institute of Statistical Mathematics, 69(2):417–428, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10463-015-0545-
6.

Partlett:2017:MAT

[3097] Christopher Partlett and Prakash Patil. Measuring asymmetry and testing symmetry. Annals of the Institute of Statistical Mathematics, 69 (2):429–460, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-015-0547-
                                        4.

Paulauskas:2017:CNT

[3098] Vygantas Paulauskas and Marijus Vaiciulis. A class of new tail index estimators. Annals of the Institute of Statistical Mathematics, 69(2):461–487, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10463-015-0548-3.

Kong:2017:NAE

[3099] Yong Kong. Number of appearances of events in random sequences: a new generating function approach to Type II and Type III runs. Annals of the Institute of Statistical Mathematics, 69(2):489–495, April 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10463-015-0549-
                                        2.

Kong:2017:TLR

[3100] Yong Kong. The m-th longest runs of multivariate random sequences. Annals of the Institute of Statistical Mathematics, 69(3):497–512, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Prokesova:2017:TSE

[3101] Michaela Prokesová, Jirí Dvorák, and Eva B. Vedel Jensen. Two-step estimation procedures for inhomogeneous shot-noise Cox processes. Annals of the Institute of Statistical Mathematics, 69(3):513–542, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Fourdrinier:2017:BMR

[3102] Dominique Fourdrinier, Fatiha Mezoued, and William E. Strawderman. A Bayes minimax result for spherically symmetric unimodal distributions. Annals of the Institute of Statistical Mathematics, 69(3):543–570, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Yuan:2017:ICR

[3103] Ke-Hai Yuan and Peter M. Bentler. Improving the convergence rate and speed of Fisher-scoring algorithm: ridge and anti-ridge methods in structural equation modeling. Annals of the Institute of Statistical Mathematics, 69 (3):571–597, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Luo:2017:NNP

[3104] Shan Luo and Gengsheng Qin. New non-parametric inferences for lowincome proportions. Annals of the Institute of Statistical Mathematics, 69(3):599–626, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Wu:2017:NVS

[3105] Ping Wu, Xinchao Luo, Peirong Xu, and Lixing Zhu. New variable selection for linear mixed-effects models. Annals of the Institute of Statistical Mathematics, 69(3):627–646, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Xi:2017:DTT

[3106] Jing Xi, Jin Xie, and Ruriko Yoshida. Distributions of topological tree metrics between a species tree and a gene tree. Annals of the Institute of Statistical Mathematics, 69(3):647–671, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Gross:2017:GFL

[3107] Elizabeth Gross, Sonja Petrović, and Despina Stasi. Goodness of fit for log-linear network models: dynamic Markov bases using hypergraphs. Annals of the Institute of Statistical Mathematics, 69(3):673–704, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Mao:2017:EPS

[3108] Chang Xuan Mao, Cuiying Yang, Yitong Yang, and Wei Zhuang. Estimating population sizes with the Rasch model. Annals of the Institute of Statistical Mathematics, 69(3):705–716, June 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

Balakrishnan:2017:ISD

[3109] N. Balakrishnan, M. V. Koutras, and F. S. Milienos. On the identifiability of start-up demonstration mixture models. Annals of the Institute of Statistical Mathematics, 69(4):717–735, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0569-6.

Dentcheva:2017:SEC

[3110] Darinka Dentcheva, Spiridon Penev, and Andrzej Ruszczyński. Statistical estimation of composite risk functionals and risk optimization problems. Annals of the Institute of Statistical Mathematics, 69(4):737–760, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0559-8.

Zhao:2017:QRV

[3111] Weihua Zhao, Riquan Zhang, Yazhao Lv, and Jicai Liu. Quantile regression and variable selection of single-index coefficient model. Annals of the Institute of Statistical Mathematics, 69 (4):761–789, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0558-9.

Vaiter:2017:DFP

[3112] Samuel Vaiter, Charles Deledalle, Jalal Fadili, Gabriel Peyré, and Charles Dossal. The degrees of freedom of partly smooth regularizers. Annals of the Institute of Statistical Mathematics, 69 (4):791–832, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0563-z.

Tsukuda:2017:CDP

[3113] Koji Tsukuda. A change detection procedure for an ergodic diffusion process. Annals of the Institute of Statistical Mathematics, 69(4):833–864, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0564-y.

Donovan:2017:SAC

[3114] Diane Donovan, Benjamin Haaland, and David J. Nott. A simple approach to constructing quasi-Sudoku-based sliced space-filling designs. Annals of the Institute of Statistical Mathematics, 69(4):865–878, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-016-0565-x.

Le:2017:UFM

[3115] Hông Vân Lê. The uniqueness of the Fisher metric as information metric. Annals of the Institute of Statistical Mathematics, 69(4):879–896, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0562-0.

Lv:2017:UPM

[3116] Shaogao Lv, Xin He, and Junhui Wang. A unified penalized method for sparse additive quantile models: an RKHS approach. Annals of the Institute of Statistical Mathematics, 69(4):897–923, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0566-9.

Mies:2017:CPP

[3117] Fabian Mies and Stefan Bedbur. On the coverage probabilities of parametric confidence bands for continuous distribution and quantile functions constructed via confidence regions for a location-scale parameter. Annals of the Institute of Statistical Mathematics, 69 (4):925–944, August 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0570-0.

Kim:2017:ETE

[3118] Moosup Kim and Sangyeol Lee. Estimation of the tail exponent of multivariate regular variation. Annals of the Institute of Statistical Mathematics, 69 (5):945–968, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0574-9.

Baringhaus:2017:LDW

[3119] L. Baringhaus, B. Ebner, and N. Henze. The limit distribution of weighted L^{2}-goodness-of-fit statistics under fixed alternatives, with applications. Annals of the Institute of Statistical Mathematics, 69(5):969–995, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0567-8.

Kwon:2017:DSA

[3120] Sunghoon Kwon, Jeongyoun Ahn, Woncheol Jang, Sangin Lee, and Yongdai Kim. A doubly sparse approach for group variable selection. Annals of the Institute of Statistical Mathematics, 69 (5):997–1025, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.

                                        1007/s10463-016-0571-z. See erra-

tum [3121].

Kwon:2017:EDS

[3121] Sunghoon Kwon, Jeongyoun Ahn, Woncheol Jang, Sangin Lee, and Yongdai Kim. Erratum to: A doubly sparse approach for group variable selection. Annals of the Institute of Statistical Mathematics, 69 (5):1027, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10463-017-0612-2.pdf;
                                        http://link.springer.com/article/
                                        10.1007/s10463-017-0612-2.  See

[3120].

Ozturk:2017:SIE

[3122] Omer Ozturk. Statistical inference with empty strata in judgment post stratified samples. Annals of the Institute of Statistical Mathematics, 69 (5):1029–1057, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-016-0572-y.

Yang:2017:SJE

[3123] Hanfang Yang and Yichuan Zhao. Smoothed jackknife empirical likelihood for the difference of two quantiles. Annals of the Institute of Statistical Mathematics, 69(5):1059–1073, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0576-7.

Grosjean:2017:AAG

[3124] Nicolas Grosjean and Thierry Huillet. Additional aspects of the generalized linear-fractional branching process. Annals of the Institute of Statistical Mathematics, 69(5):1075–1097, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0573-x.

Lu:2017:EEQ

[3125] Minggen Lu. Efficient estimation of quasi-likelihood models using Bsplines. Annals of the Institute of Statistical Mathematics, 69(5):1099–1127, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0575-8.

Fu:2017:CCD

[3126] James C. Fu and Wan-Chen Lee. On coupon collector’s and Dixie cup problems under fixed and random sample size sampling schemes. Annals of the Institute of Statistical Mathematics, 69 (5):1129–1139, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0578-5.

Shimizu:2017:MCR

[3127] Yusuke Shimizu. Moment convergence of regularized least-squares estimator for linear regression model. Annals of the Institute of Statistical Mathematics, 69(5):1141–1154, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0577-6.

Vellaisamy:2017:CSA

[3128] P. Vellaisamy. Collapsibility of some association measures and survival models. Annals of the Institute of Statistical Mathematics, 69(5):1155–1176, October 2017. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-016-
                                        0580-y.

Khmaladze:2018:FDS

[3129] Estate Khmaladze and Wolfgang Weil. Fold-up derivatives of set-valued functions and the change-set problem: A survey. Annals of the Institute of Statistical Mathematics, 70 (1):1–38, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0628-7.

Chernoyarov:2018:PEC

[3130] O. V. Chernoyarov, S. Dachian, and Yu. A. Kutoyants. On parameter estimation for cusp-type signals. Annals of the Institute of Statistical Mathematics, 70(1):39–62, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-016-0581-x.

Bhuyan:2018:IID

[3131] Prajamitra Bhuyan, Murari Mitra, and Anup Dewanji. Identifiability issues in dynamic stress-strength modeling. Annals of the Institute of Statistical Mathematics, 70(1):63–81, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0579-4.

Ruggiero:2018:CDC

[3132] Matteo Ruggiero and Matteo Sordello.

Clustering dynamics in a class of normalised generalised gamma dependent priors. Annals of the Institute of Statistical Mathematics, 70(1):83–98, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0583-8.

Barden:2018:LBF

[3133] D. Barden, H. Le, and M. Owen. Limiting behaviour of Fréchet means in the space of phylogenetic trees. Annals of the Institute of Statistical Mathematics, 70(1):99–129, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0582-9.

Xiang:2018:SMN

[3134] Sijia Xiang and Weixin Yao. Semiparametric mixtures of nonparametric regressions. Annals of the Institute of Statistical Mathematics, 70 (1):131–154, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0584-7.

Liang:2018:WEC

[3135] Han-Ying Liang and Elias Ould Saïd.

A weighted estimator of conditional hazard rate with left-truncated and dependent data. Annals of the Institute of Statistical Mathematics, 70 (1):155–189, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0587-4.

Wang:2018:VSE

[3136] Yanxin Wang, Qibin Fan, and Li Zhu. Variable selection and estimation using a continuous approximation to the L_{0} penalty. Annals of the Institute of Statistical Mathematics, 70 (1):191–214, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0588-3.

Zheng:2018:ISP

[3137] Nan Zheng and Brajendra C. Sutradhar. Inferences in semi-parametric dynamic mixed models for longitudinal count data. Annals of the Institute of Statistical Mathematics, 70 (1):215–247, February 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0590-9.

Wu:2018:FLE

[3138] C. F. Jeff Wu. A fresh look at effect aliasing and interactions: some new wine in old bottles. Annals of the Institute of Statistical Mathematics, 70(2):249–268, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0646-0. See discus-

sion [3139, 3140] and rejoinder [3141].

Peng:2018:D

[3139] Chien-Yu Peng. Discussion. Annals of the Institute of Statistical Mathematics, 70(2):269–274, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0640-y. See [3138].

Yoshida:2018:DPP

[3140] Ryo Yoshida. Discussion on the paper by Professor Wu. Annals of the Institute of Statistical Mathematics, 70(2):275–278, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0641-x. See [3138].

Wu:2018:R

[3141] C. F. Jeff Wu. Rejoinder. Annals of the Institute of Statistical Mathematics, 70(2):279–281, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0639-4. See [3138,

3139, 3140].

Liu:2018:MFF

[3142] Yi Liu and Qihua Wang. Modelfree feature screening for ultrahighdimensional data conditional on some variables. Annals of the Institute of Statistical Mathematics, 70(2):283–301, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0597-2.

Chen:2018:CTT

[3143] Chen Chen and Hosam Mahmoud. The continuous-time triangular Pólya process. Annals of the Institute of Statistical Mathematics, 70(2):303–321, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0594-5.

Wang:2018:VSS

[3144] Kangning Wang. Variable selection for spatial semivarying coefficient models. Annals of the Institute of Statistical Mathematics, 70(2):323–351, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-016-
0589-2.

Bandyopadhyay:2018:FWC

[3145] Uttam Bandyopadhyay and Atanu Biswas. Fixed-width confidence interval for covariate-adjusted responseadaptive designs. Annals of the Institute of Statistical Mathematics, 70 (2):353–371, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0596-3.

Eyjolfsson:2018:SEJ

[3146] Heidar Eyjolfsson and Dag Tjøstheim. Self-exciting jump processes with applications to energy markets. Annals of the Institute of Statistical Mathematics, 70(2):373–393, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0591-8.

Hsieh:2018:QRB

[3147] Jin-Jian Hsieh and Hong-Rui Wang.

Quantile regression based on counting process approach under semicompeting risks data. Annals of the Institute of Statistical Mathematics, 70(2):395–419, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0593-6.

Shimodaira:2018:ICM

[3148] Hidetoshi Shimodaira and Haruyoshi Maeda. An information criterion for model selection with missing data via complete-data divergence. Annals of the Institute of Statistical Mathematics, 70(2):421–438, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0592-7.

Kohler:2018:NQE

[3149] Michael Kohler, Adam Krzyzak, Reinhard Tent, and Harro Walk. Nonparametric quantile estimation using importance sampling. Annals of the Institute of Statistical Mathematics, 70(2):439–465, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-016-0595-4.

Kevei:2018:AMA

[3150] Péter Kevei. Asymptotic moving average representation of high-frequency sampled multivariate CARMA processes. Annals of the Institute of Statistical Mathematics, 70(2):467–487, April 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0601-5.

Tang:2018:RVS

[3151] Qingguo Tang and R. J. Karunamuni. Robust variable selection for finite mixture regression models. Annals of the Institute of Statistical Mathematics, 70(3):489–521, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0602-4.

Chiang:2018:VEC

[3152] Chin-Tsang Chiang, Shao-Hsuan Wang, and Ming-Yueh Huang. Versatile estimation in censored singleindex hazards regression. Annals of the Institute of Statistical Mathematics, 70(3):523–551, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0600-6.

Zhao:2018:AVC

[3153] Weihua Zhao, Jianbo Li, and Heng Lian. Adaptive varying-coefficient linear quantile model: a profiled estimating equations approach. Annals of the Institute of Statistical Mathematics, 70(3):553–582, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0599-8.

Razmkhah:2018:SIB

[3154] M. Razmkhah and S. Simriz. Statistical inferences based on INID progressively type II censored order statistics. Annals of the Institute of Statistical Mathematics, 70(3):583–604, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0598-9.

Giurcanu:2018:BIM

[3155] Mihai Giurcanu and Brett Presnell.

Bootstrap inference for misspecified moment condition models. Annals of the Institute of Statistical Mathematics, 70(3):605–630, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0604-2.

Kong:2018:FSD

[3156] Xiangshun Kong, Mingyao Ai, and Kwok Leung Tsui. Flexible sliced designs for computer experiments. Annals of the Institute of Statistical Mathematics, 70(3):631–646, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0603-3.

Eckle:2018:MIM

[3157] Konstantin Eckle, Nicolai Bissantz, Holger Dette, Katharina Proksch, and Sabrina Einecke. Multiscale inference for a multivariate density with applications to X-ray astronomy. Annals of the Institute of Statistical Mathematics, 70(3):647–689, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0605-1.

Banerjee:2018:MPT

[3158] Buddhananda Banerjee and Satyaki Mazumder. A more powerful test identifying the change in mean of functional data. Annals of the Institute of Statistical Mathematics, 70(3):691–715, June 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-017-
                                        0606-0.

Thorpe:2018:PCP

[3159] Matthew Thorpe and Adam M. Johansen. Pointwise convergence in probability of general smoothing splines. Annals of the Institute of Statistical Mathematics, 70(4):717–744, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-017-
                                        0609-x.

Bandyopadhyay:2018:ATV

[3160] Soutir Bandyopadhyay and Arnab Maity. Asymptotic theory for varying coefficient regression models with dependent data. Annals of the Institute of Statistical Mathematics, 70 (4):745–759, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0607-z.

Wu:2018:ERT

[3161] Jingjing Wu and Rohana J. Karunamuni. Efficient and robust tests for semiparametric models. Annals of the Institute of Statistical Mathematics, 70 (4):761–788, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0608-y.

An:2018:JEL

[3162] Yueheng An and Yichuan Zhao. Jackknife empirical likelihood for the difference of two volumes under ROC surfaces. Annals of the Institute of Statistical Mathematics, 70(4):789–806, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0631-z.

Chen:2018:ICP

[3163] Fuqi Chen, Rogemar Mamon, and Sévérien Nkurunziza. Inference for a change-point problem under a generalised Ornstein–Uhlenbeck setting. Annals of the Institute of Statistical Mathematics, 70(4):807–853, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0610-4.

Kozubowski:2018:GSD

[3164] Tomasz J. Kozubowski and Krzysztof Podgórski. A generalized Sibuya distribution. Annals of the Institute of Statistical Mathematics, 70(4):855–887, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0611-3.

Efromovich:2018:HRE

[3165] Sam Efromovich and Jufen Chu. Hazard rate estimation for left truncated and right censored data. Annals of the Institute of Statistical Mathematics, 70 (4):889–917, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0617-x.

Boente:2018:TEB

[3166] Graciela Boente, Daniela Rodriguez, and Mariela Sued. Testing equality between several populations covariance operators. Annals of the Institute of Statistical Mathematics, 70(4):919–950, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-017-
                                        0613-1.

Koutras:2018:ARJ

[3167] Markos V. Koutras and Demetrios P. Lyberopoulos. Asymptotic results for jump probabilities associated to the multiple scan statistic. Annals of the Institute of Statistical Mathematics, 70 (5):951–968, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0621-1.

Maesono:2018:SNT

[3168] Yoshihiko Maesono, Taku Moriyama, and Mengxin Lu. Smoothed nonparametric tests and approximations of pvalues. Annals of the Institute of Statistical Mathematics, 70(5):969–982, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0614-0.

Kahle:2018:HSE

[3169] David Kahle, Ruriko Yoshida, and Luis Garcia-Puente. Hybrid schemes for exact conditional inference in discrete exponential families. Annals of the Institute of Statistical Mathematics, 70 (5):983–1011, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0615-z.

Niu:2018:RAM

[3170] Cuizhen Niu and Lixing Zhu. A robust adaptive-to-model enhancement test for parametric single-index models. Annals of the Institute of Statistical Mathematics, 70(5):1013–1045, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-

0626-9. See correction [3171].

Niu:2018:CRA

[3171] Cuizhen Niu and Lixing Zhu. Correction to: A robust adaptive-to-model enhancement test for parametric singleindex models. Annals of the Institute of Statistical Mathematics, 70(5):1047, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0633-x;   http://link.springer.
com/content/pdf/10.1007/s10463-
017-0633-x.pdf. See [3170].

Mukhopadhyay:2018:PSB

[3172] Nitis Mukhopadhyay and Sudeep R. Bapat. Purely sequential bounded-risk point estimation of the negative binomial mean under various loss functions: one-sample problem. Annals of the Institute of Statistical Mathematics, 70 (5):1049–1075, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0620-2.

Zhang:2018:CHT

[3173] Jun Zhang, Zhenghui Feng, and Xiaoguang Wang. A constructive hypothesis test for the single-index models with two groups. Annals of the Institute of Statistical Mathematics, 70 (5):1077–1114, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0616-y.

Bindele:2018:GRB

[3174] Huybrechts F. Bindele, Ash Abebe, and Karlene N. Meyer. General rankbased estimation for regression single index models. Annals of the Institute of Statistical Mathematics, 70(5): 1115–1146, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-017-0618-9.

Fang:2018:GPL

[3175] Yixin Fang, Heng Lian, and Hua Liang. A generalized partially linear framework for variance functions. Annals of the Institute of Statistical Mathematics, 70(5):1147–1175, October 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0619-8.

Kim:2019:SEE

[3176] Mijeong Kim and Yanyuan Ma. Semiparametric efficient estimators in heteroscedastic error models. Annals of the Institute of Statistical Mathematics, 71(1):1–28, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0622-0.

Belomestny:2019:SHV

[3177] Denis Belomestny, Fabienne Comte, and Valentine Genon-Catalot. Sobolev– Hermite versus Sobolev nonparametric density estimation on R. Annals of the Institute of Statistical Mathematics, 71(1):29–62, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0624-y.

Ozturk:2019:TSC

[3178] Omer Ozturk. Two-stage cluster samples with ranked set sampling designs. Annals of the Institute of Statistical Mathematics, 71(1):63–91, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0623-z.

Kuo:2019:PGS

[3179] Kun-Lin Kuo and Yuchung J. Wang.

Pseudo-Gibbs sampler for discrete conditional distributions. Annals of the Institute of Statistical Mathematics, 71 (1):93–105, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0625-x.

Bauer:2019:ESM

[3180] Benedikt Bauer, Felix Heimrich, Michael Kohler, and Adam Krzyzak. On estimation of surrogate models for multivariate computer experiments. Annals of the Institute of Statistical Mathematics, 71(1):107–136, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0627-8.

Liang:2019:CUT

[3181] Jiajuan Liang, Kai Wang Ng, and Guoliang Tian. A class of uniform tests for goodness-of-fit of the multivariate L_{p}-norm spherical distributions and the l_{p}-norm symmetric distributions. Annals of the Institute of Statistical Mathematics, 71(1):137–162, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0630-0.

Gaunt:2019:AEN

[3182] Robert E. Gaunt, Satish Iyengar, Adri B. Olde Daalhuis, and Burcin Simsek. An asymptotic expansion for the normalizing constant of the Conway– Maxwell–Poisson distribution. Annals of the Institute of Statistical Mathematics, 71(1):163–180, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0629-6.

Muller:2019:IAS

[3183] Ursula U. Müller, Hanxiang Peng, and Anton Schick. Inference about the slope in linear regression: an empirical likelihood approach. Annals of the Institute of Statistical Mathematics, 71 (1):181–211, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0632-y.

Dobler:2019:BKM

[3184] Dennis Dobler. Bootstrapping the Kaplan–Meier estimator on the whole line. Annals of the Institute of Statistical Mathematics, 71(1):213–246, February 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-017-
0634-9.

Umezu:2019:ANC

[3185] Yuta Umezu, Yusuke Shimizu, Hiroki Masuda, and Yoshiyuki Ninomiya. AIC for the non-concave penalized likelihood method. Annals of the Institute of Statistical Mathematics, 71(2):247–274, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0649-x.

Gao:2019:FMA

[3186] Yan Gao, Xinyu Zhang, Shouyang Wang, Terence Tai leung Chong, and Guohua Zou. Frequentist model averaging for threshold models. Annals of the Institute of Statistical Mathematics, 71(2):275–306, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0642-9.

Aki:2019:WTC

[3187] Sigeo Aki. Waiting time for consecutive repetitions of a pattern and related distributions. Annals of the Institute of Statistical Mathematics, 71 (2):307–325, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0644-2.

Chatterjee:2019:TNA

[3188] Debajit Chatterjee and Uttam Bandyopadhyay. Testing in nonparametric ANCOVA model based on ridit reliability functional. Annals of the Institute of Statistical Mathematics, 71 (2):327–364, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-017-0643-8.

Qi:2019:WEE

[3189] Lihong Qi, Xu Zhang, Yanqing Sun, Lu Wang, and Yichuan Zhao. Weighted estimating equations for additive hazards models with missing covariates. Annals of the Institute of Statistical Mathematics, 71(2):365–387, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0648-y.

Rafik:2019:GUM

[3190] Aguech Rafik, Lasmar Nabil, and Selmi Olfa. A generalized urn with multiple drawing and random addition. Annals of the Institute of Statistical Mathematics, 71(2):389–408, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0651-3.

Liao:2019:PER

[3191] Lina Liao, Cheolwoo Park, and Hosik Choi. Penalized expectile regression: an alternative to penalized quantile regression. Annals of the Institute of Statistical Mathematics, 71(2):409–438, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0645-1.

Graczyk:2019:WEF

[3192] Piotr Graczyk, Hideyuki Ishi, and Salha Mamane. Wishart exponential families on cones related to tridiagonal matrices. Annals of the Institute of Statistical Mathematics, 71(2):439–471, April 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0647-z.

Aoshima:2019:DBC

[3193] Makoto Aoshima and Kazuyoshi Yata.

Distance-based classifier by data transformation for high-dimension, strongly spiked eigenvalue models. Annals of the Institute of Statistical Mathematics, 71(3):473–503, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0655-z.

Park:2019:THP

[3194] Junyong Park. Testing homogeneity of proportions from sparse binomial data with a large number of groups. Annals of the Institute of Statistical Mathematics, 71(3):505–535, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0652-2.

Nezakati:2019:RAS

[3195] Ensiyeh Nezakati, Mostafa Razmkhah, and Firoozeh Haghighi. Reliability analysis of a k-out-of-n: F system under a linear degradation model with calibrations. Annals of the Institute of Statistical Mathematics, 71(3):537–552, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-018-
                                        0650-4.

Guo:2019:REU

[3196] Huijun Guo and Youming Liu. Regression estimation under strong mixing data. Annals of the Institute of Statistical Mathematics, 71(3):553–576, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-018-
                                        0653-1.

Lima:2019:MBS

[3197] Italo R. Lima, Guanqun Cao, and Nedret Billor. M-based simultaneous inference for the mean function of functional data. Annals of the Institute of Statistical Mathematics, 71(3):577–598, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0656-y.

Aulbach:2019:TNG

[3198] Stefan Aulbach, Michael Falk, and Timo Fuller. Testing for a δneighborhood of a generalized Pareto copula. Annals of the Institute of Statistical Mathematics, 71(3):599–626, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0657-x.

Kuchibhotla:2019:SIB

[3199] Arun Kumar Kuchibhotla, Somabha Mukherjee, and Ayanendranath Basu. Statistical inference based on bridge divergences. Annals of the Institute of Statistical Mathematics, 71(3):627–656, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0665-x.

Wang:2019:SEU

[3200] Zhaoliang Wang, Liugen Xue, Gaorong Li, and Fei Lu. Spline estimator for ultra-high dimensional partially linear varying coefficient models. Annals of the Institute of Statistical Mathematics, 71(3):657–677, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0654-0.

Barreto-Souza:2019:SGE

[3201] Wagner Barreto-Souza and Vinícius Diniz Mayrink. Semiparametric generalized exponential frailty model for clustered survival data. Annals of the Institute of Statistical Mathematics, 71(3):679–701, June 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0658-9.

Koike:2019:APR

[3202] Yuta Koike and Zhi Liu. Asymptotic properties of the realized skewness and related statistics. Annals of the Institute of Statistical Mathematics, 71 (4):703–741, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0659-8.

Levine:2019:RFE

[3203] Michael Levine. Robust functional estimation in the multivariate partial linear model. Annals of the Institute of Statistical Mathematics, 71(4):743–770, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-018-
                                        0661-1.

Miroshnikov:2019:APP

[3204] Alexey Miroshnikov and Evgeny Savelev. Asymptotic properties of parallel Bayesian kernel density estimators. Annals of the Institute of Statistical Mathematics, 71(4):771–810, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10463-018-
0662-0.

Gribkova:2019:WAT

[3205] Nadezhda Gribkova and Ricardas Zitikis. Weighted allocations, their concomitant-based estimators, and asymptotics. Annals of the Institute of Statistical Mathematics, 71(4):811–835, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0660-2.

Takagi:2019:BRU

[3206] Yoshiharu Takagi and Yutaka Kano.

Bias reduction using surrogate endpoints as auxiliary variables. Annals of the Institute of Statistical Mathematics, 71(4):837–852, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0667-8.

Mynbaev:2019:UED

[3207] Kairat Mynbaev and Carlos Martins-Filho. Unified estimation of densities on bounded and unbounded domains. Annals of the Institute of Statistical Mathematics, 71(4):853–887, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0663-z.

Wang:2019:DRK

[3208] Lei Wang. Dimension reduction for kernel-assisted M-estimators with missing response at random. Annals of the Institute of Statistical Mathematics, 71(4):889–910, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0664-y.

Qi:2019:LDL

[3209] Yongcheng Qi, Fang Wang, and Lin Zhang. Limiting distributions of likelihood ratio test for independence of components for high-dimensional normal vectors. Annals of the Institute of Statistical Mathematics, 71(4):911–946, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0666-9.

Kim:2019:TTI

[3210] Moosup Kim and Sangyeol Lee. Test for tail index constancy of GARCH innovations based on conditional volatility. Annals of the Institute of Statistical Mathematics, 71(4):947–981, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0669-6.

Gu:2019:SCB

[3211] Lijie Gu, Suojin Wang, and Lijian Yang. Simultaneous confidence bands for the distribution function of a finite population in stratified sampling. Annals of the Institute of Statistical Mathematics, 71(4):983–1005, August 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0668-7.

Chen:2019:JFS

[3212] Xiaolin Chen, Yi Liu, and Qihua Wang. Joint feature screening for ultrahigh-dimensional sparse additive hazards model by the sparsity-restricted pseudo-score estimator. Annals of the Institute of Statistical Mathematics, 71 (5):1007–1031, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0675-8.

Lee:2019:CTG

[3213] Youngmi Lee and Sangyeol Lee.

CUSUM test for general nonlinear integer-valued GARCH models: comparison study. Annals of the Institute of Statistical Mathematics, 71(5): 1033–1057, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0676-7.

Oh:2019:MRC

[3214] Haejune Oh and Sangyeol Lee. Modified residual CUSUM test for locationscale time series models with heteroscedasticity. Annals of the Institute of Statistical Mathematics, 71(5): 1059–1091, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0679-4.

Wu:2019:HSM

[3215] Xianyi Wu and Xian Zhou. On Hodges’ superefficiency and merits of oracle property in model selection. Annals of the Institute of Statistical Mathematics, 71(5):1093–1119, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0670-0.

Choi:2019:MLE

[3216] Jungjun Choi and In Choi. Maximum likelihood estimation of autoregressive models with a near unit root and Cauchy errors. Annals of the Institute of Statistical Mathematics, 71 (5):1121–1142, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0671-z.

Wang:2019:BEB

[3217] Xuejun Wang, Yi Wu, and Shuhe Hu. The Berry–Esseen bounds of the weighted estimator in a nonparametric regression model. Annals of the Institute of Statistical Mathematics, 71 (5):1143–1162, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0677-6.

Lv:2019:QEM

[3218] Jing Lv and Chaohui Guo. Quantile estimations via modified Cholesky decomposition for longitudinal singleindex models. Annals of the Institute of Statistical Mathematics, 71(5): 1163–1199, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0673-x.

Sun:2019:SER

[3219] Zhuoer Sun and Suojin Wang. Semiparametric estimation in regression with missing covariates using singleindex models. Annals of the Institute of Statistical Mathematics, 71(5): 1201–1232, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0672-y.

Hansmann:2019:SUC

[3220] Matthias Hansmann, Michael Kohler, and Harro Walk. On the strong universal consistency of local averaging regression estimates. Annals of the Institute of Statistical Mathematics, 71 (5):1233–1263, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0674-9. See correc-

tion [3221].

Hansmann:2019:CSU

[3221] Matthias Hansmann, Michael Kohler, and Harro Walk. Correction to: On the strong universal consistency of local averaging regression estimates. Annals of the Institute of Statistical Mathematics, 71(5):1265–1269, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0687-4;   http://link.springer.
com/content/pdf/10.1007/s10463-
018-0687-4.pdf. See [3220].

Schoenberg:2019:RPP

[3222] Frederic Paik Schoenberg, Marc Hoffmann, and Ryan J. Harrigan. A recursive point process model for infectious diseases. Annals of the Institute of Statistical Mathematics, 71(5): 1271–1287, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0690-9.

Maji:2019:RSI

[3223] Avijit Maji, Abhik Ghosh, Ayanendranath Basu, and Leandro Pardo. Robust statistical inference based on the C-divergence family. Annals of the Institute of Statistical Mathematics, 71 (5):1289–1322, October 2019. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0678-5.

West:2020:BFM

[3224] Mike West. Bayesian forecasting of multivariate time series: scalability, structure uncertainty and decisions. Annals of the Institute of Statistical Mathematics, 72(1):1–31, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-

00741-3. See discussion [3225, 3226] and reply [3227].

Nakajima:2020:DBF

[3225] Jouchi Nakajima. Discussion of “Bayesian forecasting of multivariate time series: scalability, structure uncertainty and decisions”. Annals of the Institute of Statistical Mathematics, 72(1):33–36, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00742-2. See [3224,

3227].

Glynn:2020:DBF

[3226] Chris Glynn. Discussion of “Bayesian forecasting of multivariate time series: scalability, structure uncertainty and decisions”. Annals of the Institute of Statistical Mathematics, 72 (1):37–39, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00743-1. See [3224,

3227].

West:2020:RDB

[3227] Mike West. Reply to discussion of “Bayesian forecasting of multivariate time series: scalability, structure uncertainty and decisions”. Annals of the Institute of Statistical Mathematics, 72(1):41–44, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00744-0. See [3224,

3225, 3226].

He:2020:DMS

[3228] Xin He and Junhui Wang. Discovering model structure for partially linear models. Annals of the Institute of Statistical Mathematics, 72 (1):45–63, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0682-9.

Chen:2020:TSS

[3229] Zehua Chen and Yiwei Jiang. A twostage sequential conditional selection approach to sparse high-dimensional multivariate regression models. Annals of the Institute of Statistical Mathematics, 72(1):65–90, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0686-5.

Berger:2020:ELA

[3230] Yves G. Berger. An empirical likelihood approach under cluster sampling with missing observations. Annals of the Institute of Statistical Mathematics, 72(1):91–121, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0681-x.

Kohler:2020:EQI

[3231] Michael Kohler and Adam Krzyzak. Estimating quantiles in imperfect simulation models using conditional density estimation. Annals of the Institute of Statistical Mathematics, 72 (1):123–155, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0683-8.

Dumbgen:2020:IDF

[3232] Lutz Dümbgen and Ehsan Zamanzade. Inference on a distribution function from ranked set samples. Annals of the Institute of Statistical Mathematics, 72 (1):157–185, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0680-y.

Moriyama:2020:NKE

[3233] Taku Moriyama and Yoshihiko Maesono.

New kernel estimators of the hazard ratio and their asymptotic properties. Annals of the Institute of Statistical Mathematics, 72(1):187–211, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0685-6.

Zhao:2020:MQR

[3234] Weihua Zhao, Weiping Zhang, and Heng Lian. Marginal quantile regression for varying coefficient models with longitudinal data. Annals of the Institute of Statistical Mathematics, 72 (1):213–234, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0684-7.

Konev:2020:SFA

[3235] Victor Konev and Bogdan Nazarenko.

Sequential fixed accuracy estimation for nonstationary autoregressive processes. Annals of the Institute of Statistical Mathematics, 72(1): 235–264, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0689-2.

Bravo:2020:SQR

[3236] Francesco Bravo. Semiparametric quantile regression with random censoring. Annals of the Institute of Statistical Mathematics, 72(1): 265–295, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0688-3.

Poignard:2020:ATA

[3237] Benjamin Poignard. Asymptotic theory of the adaptive Sparse Group Lasso. Annals of the Institute of Statistical Mathematics, 72(1): 297–328, February 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0692-7.

Copas:2020:SMD

[3238] John Copas and Shinto Eguchi. Strong model dependence in statistical analysis: goodness of fit is not enough for model choice. Annals of the Institute of Statistical Mathematics, 72(2):329–352, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0691-8.

Luu:2020:SOI

[3239] Tung Duy Luu, Jalal Fadili, and Christophe Chesneau. Sharp oracle inequalities for low-complexity priors. Annals of the Institute of Statistical Mathematics, 72(2):353–397, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0693-6.

Afendras:2020:LDS

[3240] Georgios Afendras, Nickos Papadatos, and Violetta E. Piperigou. On the limiting distribution of sample central moments. Annals of the Institute of Statistical Mathematics, 72(2):399–425, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0695-4.

Zou:2020:EDE

[3241] Feng Zou and Hengjian Cui. Error density estimation in high-dimensional sparse linear model. Annals of the Institute of Statistical Mathematics, 72(2):427–449, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0699-0.

Barmi:2020:TPS

[3242] Hammou El Barmi. A test for the presence of stochastic ordering under censoring: the k-sample case. Annals of the Institute of Statistical Mathematics, 72(2):451–470, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0694-5.

Chowdhury:2020:CRK

[3243] Joydeep Chowdhury and Probal Chaudhuri. Convergence rates for kernel regression in infinite-dimensional spaces. Annals of the Institute of Statistical Mathematics, 72(2):471–509, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10463-018-
                                        0697-2.

Tsuruta:2020:TPB

[3244] Yasuhito Tsuruta and Masahiko Sagae. Theoretical properties of bandwidth selectors for kernel density estimation on the circle. Annals of the Institute of Statistical Mathematics, 72(2):511–530, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-018-
                                        0701-x.

Wu:2020:CWT

[3245] Tung-Lung Wu. Conditional waiting time distributions of runs and patterns and their applications. Annals of the Institute of Statistical Mathematics, 72(2):531–543, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0696-3.

Olsson:2020:PBO

[3246] Jimmy Olsson and Johan Westerborn Alenlöv. Particle-based online estimation of tangent filters with application to parameter estimation in nonlinear state-space models. Annals of the Institute of Statistical Mathematics, 72(2):545–576, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-0698-1.

Delsol:2020:SEN

[3247] Laurent Delsol and Ingrid Van Keilegom. Semiparametric M-estimation with non-smooth criterion functions.

Annals of the Institute of Statistical Mathematics, 72(2):577–605, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0700-y.

Durante:2020:SHC

[3248] Fabrizio Durante, Juan Fernández Sánchez, and Wolfgang Trutschnig. Spatially homogeneous copulas. Annals of the Institute of Statistical Mathematics, 72(2):607–626, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-018-
0703-8.

Jones:2020:PCR

[3249] Ben Jones and Andreas Artemiou. On principal components regression with Hilbertian predictors. Annals of the Institute of Statistical Mathematics, 72(2):627–644, April 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-018-0702-9.

Jones:2020:USD

[3250] M. C. Jones. On univariate slash distributions, continuous and discrete. Annals of the Institute of Statistical Mathematics, 72(3):645–657, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00708-4.

Brehmer:2020:PCP

[3251] Jonas R. Brehmer and Tilmann Gneiting. Properization: constructing proper scoring rules via Bayes acts. Annals of the Institute of Statistical Mathematics, 72(3):659–673, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-019-
                                        00705-7.

Farinetto:2020:PSL

[3252] C. Farinetto, Yu. A. Kutoyants, and A. Top. Poisson source localization on the plane: change-point case. Annals of the Institute of Statistical Mathematics, 72(3):675–698, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10463-018-00704-0.

Zhang:2020:EHT

[3253] Jun Zhang, Xia Cui, and Heng Peng. Estimation and hypothesis test for partial linear single-index multiplicative models. Annals of the Institute of Statistical Mathematics, 72(3):699–740, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10463-019-
                                        00706-6.

Henze:2020:MGN

[3254] Norbert Henze and Celeste Mayer. More good news on the HKM test for multivariate reflected symmetry about an unknown centre. Annals of the Institute of Statistical Mathematics, 72(3):741–770, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00707-5.

Abrams:2020:NEC

[3255] Steven Abrams, Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Nonparametric estimation of the cross ratio function. Annals of the Institute of Statistical Mathematics, 72 (3):771–801, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00709-3.

Kato:2020:SBM

[3256] Ryo Kato and Takahiro Hoshino. Semiparametric Bayesian multiple imputation for regression models with missing mixed continuous-discrete covariates. Annals of the Institute of Statistical Mathematics, 72(3):803–825, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00710-w.

Feng:2020:NVS

[3257] Zhenghui Feng, Lu Lin, Ruoqing Zhu, and Lixing Zhu. Nonparametric variable selection and its application to additive models. Annals of the Institute of Statistical Mathematics, 72(3):827–854, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00711-9.

Agostinelli:2020:RES

[3258] Claudio Agostinelli, Ana M. Bianco, and Graciela Boente. Robust estimation in single-index models when the errors have a unimodal density with unknown nuisance parameter. Annals of the Institute of Statistical Mathematics, 72(3):855–893, June 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00712-8.

Martin:2020:DPS

[3259] Donald E. K. Martin. Distributions of pattern statistics in sparse Markov models. Annals of the Institute of Statistical Mathematics, 72(4):895–913, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00714-6.

Gardes:2020:EEC

[3260] Laurent Gardes, Armelle Guillou, and Claire Roman. Estimation of extreme conditional quantiles under a general tail-first-order condition. Annals of the Institute of Statistical Mathematics, 72 (4):915–943, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00713-7.

Li:2020:STI

[3261] Chenlong Li, Zhanjie Song, and Wenjun Wang. Space-time inhomogeneous background intensity estimators for semi-parametric space-time selfexciting point process models. Annals of the Institute of Statistical Mathematics, 72(4):945–967, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10463-019-
00715-5.

Eriksson:2020:LSR

[3262] Frank Eriksson, Torben Martinussen, and Søren Feodor Nielsen. Large sample results for frequentist multiple imputation for Cox regression with missing covariate data. Annals of the Institute of Statistical Mathematics, 72 (4):969–996, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00716-4.

Dobler:2020:NMM

[3263] Dennis Dobler, Sarah Friedrich, and Markus Pauly. Nonparametric MANOVA in meaningful effects. Annals of the Institute of Statistical Mathematics, 72(4):997–1022, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00717-3.

Comte:2020:RFE

[3264] F. Comte and V. Genon-Catalot. Regression function estimation as a partly inverse problem. Annals of the Institute of Statistical Mathematics, 72 (4):1023–1054, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http:

//link.springer.com/article/10.
1007/s10463-019-00718-2.

Bucher:2020:DDS

[3265] Axel Bücher, Holger Dette, and Florian Heinrichs. Detecting deviations from second-order stationarity in locally stationary functional time series. Annals of the Institute of Statistical Mathematics, 72(4):1055–1094, August 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10463-019-
00721-7.

Dearborn:2020:IEM

[3266] Krisztina Dearborn and Rafael Frongillo. On the indirect elicitability of the mode and modal interval. Annals of the Institute of Statistical Mathematics, 72 (5):1095–1108, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00719-1.

Henze:2020:TND

[3267] Norbert Henze and Jaco Visagie. Testing for normality in any dimension based on a partial differential equation involving the moment generating function. Annals of the Institute of Statistical Mathematics, 72(5): 1109–1136, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00720-8.

Chernoyarov:2020:PSL

[3268] O. V. Chernoyarov, S. Dachian, and Yu. A. Kutoyants. Poisson source localization on the plane: cusp case. Annals of the Institute of Statistical Mathematics, 72(5): 1137–1157, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00722-6.

Chen:2020:BCF

[3269] Kun Chen, Ngai Hang Chan, and Chun Yip Yau. Bartlett correction of frequency domain empirical likelihood for time series with unknown innovation variance. Annals of the Institute of Statistical Mathematics, 72 (5):1159–1173, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00723-5.

Fonseca:2020:WED

[3270] Rodney V. Fonseca and Aluísio Pinheiro. Wavelet estimation of the dimensionality of curve time series. Annals of the Institute of Statistical Mathematics, 72(5):1175– 1204, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00724-4.

Beltaief:2020:MSR

[3271] Slim Beltaief, Oleg Chernoyarov, and Serguei Pergamenchtchikov. Model selection for the robust efficient signal processing observed with small Lévy noise. Annals of the Institute of Statistical Mathematics, 72(5): 1205–1235, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00726-2.

Harsha:2020:SII

[3272] K. V. Harsha and Alladi Subramanyam. Some information inequalities for statistical inference.

Annals of the Institute of Statistical Mathematics, 72(5):1237– 1256, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00725-3.

Nakayama:2020:BCS

[3273] Yugo Nakayama, Kazuyoshi Yata, and Makoto Aoshima. Bias-corrected support vector machine with Gaussian kernel in high-dimension, low-sample-size settings. Annals of the Institute of Statistical Mathematics, 72(5): 1257–1286, October 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00727-1.

Neumeyer:2020:SPT

[3274] Natalie Neumeyer, Leonie Selk, and Charles Tillier. Semi-parametric transformation boundary regression models. Annals of the Institute of Statistical Mathematics, 72(6):1287– 1315, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00731-5.

Hadjicosta:2020:ITM

[3275] Elena Hadjicosta and Donald Richards. Integral transform methods in goodnessof-fit testing, II: the Wishart distributions. Annals of the Institute of Statistical Mathematics, 72(6): 1317–1370, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00737-z.

Kim:2020:REG

[3276] Byungsoo Kim and Sangyeol Lee. Robust estimation for general integervalued time series models. Annals of the Institute of Statistical Mathematics, 72(6):1371–1396, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL https://link.springer.com/
article/10.1007/s10463-019-00728-
0.

Mandal:2020:OTA

[3277] Abhijit Mandal. An optimal test for the additive model with discrete or categorical predictors. Annals of the Institute of Statistical Mathematics, 72(6): 1397–1417, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00729-z.

Doukhan:2020:CMD

[3278] Paul Doukhan, Ieva Grublyte, and Laurence Reboul. Comparing the marginal densities of two strictly stationary linear processes. Annals of the Institute of Statistical Mathematics, 72(6):1419–1447, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL https://link.springer.com/
article/10.1007/s10463-019-00730-
6.

Cui:2020:FBP

[3279] Yan Cui, Qi Li, and Fukang Zhu. Flexible bivariate Poisson integer-valued GARCH model. Annals of the Institute of Statistical Mathematics, 72(6): 1449–1477, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00732-4.

Benelmadani:2020:RKH

[3280] D. Benelmadani, K. Benhenni, and S. Louhichi. The reproducing kernel Hilbert space approach in nonparametric regression problems with correlated observations. Annals of the Institute of Statistical Mathematics, 72(6): 1479–1500, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00733-3.

Ohishi:2020:EBA

[3281] Mineaki Ohishi, Hirokazu Yanagihara, and Shuichi Kawano. Equivalence between adaptive Lasso and generalized ridge estimators in linear regression with orthogonal explanatory variables after optimizing regularization parameters. Annals of the Institute of Statistical Mathematics, 72(6): 1501–1516, December 2020. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00734-2.

Fujisawa:2021:RE

[3282] H. Fujisawa, Y. Ninomiya, and T. Sei. Report of the editors. Annals of the Institute of Statistical Mathematics, 73(1):1–2, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00778-9.

Honda:2021:BGL

[3283] Toshio Honda. The de-biased group Lasso estimation for varying coefficient models. Annals of the Institute of Statistical Mathematics, 73 (1):3–29, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00740-4.

Betsch:2021:FPC

[3284] Steffen Betsch and Bruno Ebner. Fixed point characterizations of continuous univariate probability distributions and their applications. Annals of the Institute of Statistical Mathematics, 73 (1):31–59, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00735-1.

Dette:2021:SES

[3285] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. Some explicit solutions of c-optimal design problems for polynomial regression with no intercept. Annals of the Institute of Statistical Mathematics, 73 (1):61–82, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-019-00736-0.

Jiang:2021:PER

[3286] Yingying Jiang, Fuming Lin, and Yong Zhou. The k-th power expectile regression. Annals of the Institute of Statistical Mathematics, 73 (1):83–113, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00738-y.

Cao:2021:PHM

[3287] Hongyuan Cao and Jason P. Fine. On the proportional hazards model with last observation carried forward covariates. Annals of the Institute of Statistical Mathematics, 73 (1):115–134, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00739-x.

Nong:2021:CSM

[3288] Quynh Van Nong and Chi Tim Ng. Clustering of subsample means based on pairwise L_{1} regularized empirical likelihood. Annals of the Institute of Statistical Mathematics, 73(1): 135–174, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-019-00745-z. See cor-

rection [3289].

Nong:2021:CCS

[3289] Quynh Van Nong and Chi Tim Ng. Correction to: Clustering of subsample means based on pairwise L_{1} regularized empirical likelihood. Annals of the Institute of Statistical Mathematics, 73(1):175, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00767-y.  See

[3288].

Nakakita:2021:QLA

[3290] Shogo H. Nakakita, Yusuke Kaino, and Masayuki Uchida. Quasilikelihood analysis and Bayes-type estimators of an ergodic diffusion plus noise. Annals of the Institute of Statistical Mathematics, 73(1): 177–225, February 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00746-3.

Vexler:2021:VVE

[3291] Albert Vexler. Valid p-values and expectations of p-values revisited. Annals of the Institute of Statistical Mathematics, 73(2): 227–248, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00747-2.

Gotz:2021:EIS

[3292] Benedict Götz, Sebastian Kersting, and Michael Kohler. Estimation of an improved surrogate model in uncertainty quantification by neural networks. Annals of the Institute of Statistical Mathematics, 73 (2):249–281, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00748-1.

Gibberd:2021:CMC

[3293] A. Gibberd and S. Roy. Consistent multiple changepoint estimation with fused Gaussian graphical models. Annals of the Institute of Statistical Mathematics, 73(2):283–309, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00749-0.

Chen:2021:PSS

[3294] Shiyun Chen and Ery Arias-Castro. On the power of some sequential multiple testing procedures. Annals of the Institute of Statistical Mathematics, 73(2):311–336, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00752-5.

Kampf:2021:NEK

[3295] Jürgen Kampf, Georgiy Shevchenko, and Evgeny Spodarev. Nonparametric estimation of the kernel function of symmetric stable moving average random functions. Annals of the Institute of Statistical Mathematics, 73 (2):337–367, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00751-6.

Albrecher:2021:MMM

[3296] Hansjörg Albrecher, Martin Bladt, and Mogens Bladt. Multivariate matrix Mittag-Leffler distributions. Annals of the Institute of Statistical Mathematics, 73(2):369–394, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00750-7.

Taleb:2021:MAP

[3297] Youssef Taleb and Edward A. K. Cohen. Multiresolution analysis of point processes and statistical thresholding for Haar wavelet-based intensity estimation. Annals of the Institute of Statistical Mathematics, 73 (2):395–423, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00753-4.

Wong:2021:PPC

[3298] Ka Yiu Wong and Dietrich Stoyan.

Poles of pair correlation functions: When they are real? Annals of the Institute of Statistical Mathematics, 73(2):425–440, April 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00754-3.

Wang:2021:CCQ

[3299] Kangning Wang and Wen Shan. Copula and composite quantile regressionbased estimating equations for longitudinal data. Annals of the Institute of Statistical Mathematics, 73 (3):441–455, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00756-1.

Lai:2021:MIS

[3300] Peng Lai, Fangjian Wang, and Qingzhao Zhang. Model identification and selection for single-index varyingcoefficient models. Annals of the Institute of Statistical Mathematics, 73 (3):457–480, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00757-0.

Chen:2021:SML

[3301] Li-Pang Chen and Grace Y. Yi. Semiparametric methods for left-truncated and right-censored survival data with covariate measurement error. Annals of the Institute of Statistical Mathematics, 73(3):481–517, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00755-2.

Chen:2021:ISP

[3302] Ji Chen, Jun Shao, and Fang Fang. Instrument search in pseudolikelihood approach for nonignorable nonresponse. Annals of the Institute of Statistical Mathematics, 73 (3):519–533, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00758-z.

Wei:2021:MAL

[3303] Yuting Wei, Qihua Wang, and Wei Liu. Model averaging for linear models with responses missing at random. Annals of the Institute of Statistical Mathematics, 73(3):535–553, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00759-y.

Inatsugu:2021:GJF

[3304] Haruhiko Inatsugu and Nakahiro Yoshida. Global jump filters and quasilikelihood analysis for volatility. Annals of the Institute of Statistical Mathematics, 73(3):555–598, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00768-x.

Ishii:2021:HTH

[3305] Aki Ishii, Kazuyoshi Yata, and Makoto Aoshima. Hypothesis tests for high-dimensional covariance structures. Annals of the Institute of Statistical Mathematics, 73(3): 599–622, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00760-5.

Wang:2021:IEL

[3306] Lei Wang and Wei Ma. Improved empirical likelihood inference and variable selection for generalized linear models with longitudinal nonignorable dropouts. Annals of the Institute of Statistical Mathematics, 73 (3):623–647, June 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00761-4.

Clinet:2021:EHF

[3307] Simon Clinet and Yoann Potiron. Estimation for high-frequency data under parametric market microstructure noise. Annals of the Institute of Statistical Mathematics, 73 (4):649–669, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00762-3.

Kutoyants:2021:LSH

[3308] Yury A. Kutoyants. On localization of source by hidden Gaussian processes with small noise. Annals of the Institute of Statistical Mathematics, 73 (4):671–702, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00763-2.

Ren:2021:RHD

[3309] Mingyang Ren, Sanguo Zhang, and Qingzhao Zhang. Robust highdimensional regression for data with anomalous responses. Annals of the Institute of Statistical Mathematics, 73 (4):703–736, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00764-1.

Zheng:2021:DNC

[3310] Jiasen Zheng and Lixing Zhu. Determining the number of canonical correlation pairs for high-dimensional vectors. Annals of the Institute of Statistical Mathematics, 73(4): 737–756, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00776-x.

Misra:2021:GGM

[3311] Pratik Misra and Seth Sullivant. Gaussian graphical models with toric vanishing ideals. Annals of the Institute of Statistical Mathematics, 73 (4):757–785, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00765-0.

Qin:2021:HDS

[3312] Shanshan Qin, Hao Ding, and Feng Liu. High-dimensional signconstrained feature selection and grouping. Annals of the Institute of Statistical Mathematics, 73 (4):787–819, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00766-z.

Kim:2021:RTS

[3313] Byungsoo Kim, Junmo Song, and Changryong Baek. Robust test for structural instability in dynamic factor models. Annals of the Institute of Statistical Mathematics, 73 (4):821–853, August 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00773-0.

Dette:2021:ISB

[3314] Holger Dette, Subhra Sankar Dhar, and Weichi Wu. Identifying shifts between two regression curves. Annals of the Institute of Statistical Mathematics, 73 (5):855–889, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00771-2.

Chandra:2021:ATD

[3315] Noirrit Kiran Chandra and Sourabh Bhattacharya. Asymptotic theory of dependent Bayesian multiple testing procedures under possible model misspecification. Annals of the Institute of Statistical Mathematics, 73 (5):891–920, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00770-3.

DeGregorio:2021:RBT

[3316] Alessandro De Gregorio and Francesco Iafrate. Regularized bridge-type estimation with multiple penalties. Annals of the Institute of Statistical Mathematics, 73(5):921–951, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10463-020-00769-
                                        w.

Geenens:2021:MMK

[3317] Gery Geenens. Mellin–Meijer kernel density estimation on R^{+}. Annals of the Institute of Statistical Mathematics, 73(5):953–977, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00772-1.

Piancastelli:2021:GIG

[3318] Luiza S. C. Piancastelli, Wagner Barreto-Souza, and Vinícius D. Mayrink. Generalized inverse-Gaussian frailty models with application to TARGET neuroblastoma data. Annals of the Institute of Statistical Mathematics, 73(5):979–1010, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10463-020-00774-
                                        z.

Camerlenghi:2021:ABM

[3319] Federico Camerlenghi, Claudio Macci, and Elena Villa. Asymptotic behavior of mean density estimators based on a single observation: the Boolean model case. Annals of the Institute of Statistical Mathematics, 73(5): 1011–1035, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00775-y.

Wylupek:2021:PTT

[3320] Grzegorz Wyłupek. A permutation test for the two-sample rightcensored model. Annals of the Institute of Statistical Mathematics, 73(5): 1037–1061, October 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00777-w.

Khmaladze:2021:DFT

[3321] Estate V. Khmaladze. Distributionfree testing in linear and parametric regression. Annals of the Institute of Statistical Mathematics, 73(6): 1063–1087, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00786-3.

Vetter:2021:UAE

[3322] Mathias Vetter. A universal approach to estimate the conditional variance in semimartingale limit theorems. Annals of the Institute of Statistical Mathematics, 73(6):1089– 1125, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00781-0.

Yuan:2021:FEM

[3323] Baichuan Yuan, Frederic P. Schoenberg, and Andrea L. Bertozzi. Fast estimation of multivariate spatiotemporal Hawkes processes and network reconstruction. Annals of the Institute of Statistical Mathematics, 73(6): 1127–1152, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00780-1.

Nagatsuka:2021:ELB

[3324] Hideki Nagatsuka and N. Balakrishnan. Efficient likelihood-based inference for the generalized Pareto distribution. Annals of the Institute of Statistical Mathematics, 73(6): 1153–1185, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00782-z.

Druilhet:2021:IVF

[3325] Pierre Druilhet and Erwan Saint Loubert Bié. Improper versus finitely additive distributions as limits of countably additive probabilities. Annals of the Institute of Statistical Mathematics, 73(6):1187– 1202, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-020-00779-8.

Gomez:2021:NRW

[3326] Luz M. Gómez, Rogério F. Porto, and Pedro A. Morettin. Nonparametric regression with warped wavelets and strong mixing processes. Annals of the Institute of Statistical Mathematics, 73(6):1203– 1228, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00789-0.

Hanebeck:2021:SDF

[3327] Ariane Hanebeck and Bernhard Klar. Smooth distribution function estimation for lifetime distributions using Szasz–Mirakyan operators. Annals of the Institute of Statistical Mathematics, 73(6):1229– 1247, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-020-00783-y.

Cai:2021:SCB

[3328] Li Cai, Lijie Gu, and Suojin Wang. Simultaneous confidence bands for nonparametric regression with missing covariate data. Annals of the Institute of Statistical Mathematics, 73(6): 1249–1279, December 2021. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00784-5.

Poignard:2022:FSP

[3329] Benjamin Poignard and Jean-David Fermanian. The finite sample properties of sparse M-estimators with pseudo-observations. Annals of the Institute of Statistical Mathematics, 74 (1):1–31, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00785-4.

Amari:2022:WSO

[3330] Shun ichi Amari and Takeru Matsuda. Wasserstein statistics in onedimensional location scale models. Annals of the Institute of Statistical Mathematics, 74(1):33–47, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10463-021-00788-
                                        1.

Yang:2022:TSL

[3331] Kai Yang and Peihua Qiu. A three-step local smoothing approach for estimating the mean and covariance functions of spatio-temporal data. Annals of the Institute of Statistical Mathematics, 74 (1):49–68, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00787-2.

Sun:2022:BAR

[3332] Zhihua Sun, Yi Liu, and Gang Li. Broken adaptive ridge regression for rightcensored survival data. Annals of the Institute of Statistical Mathematics, 74 (1):69–91, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00794-3.

Li:2022:ELM

[3333] Mengke Li, Yukun Liu, and Jing Qin. Empirical likelihood meta-analysis with publication bias correction under copas-like selection model. Annals of the Institute of Statistical Mathematics, 74(1):93–112, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00793-4.

Pchelintsev:2022:EEM

[3334] Evgeny Pchelintsev, Serguei Pergamenshchikov, and Maria Povzun. Efficient estimation methods for non-Gaussian regression models in continuous time. Annals of the Institute of Statistical Mathematics, 74(1): 113–142, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00790-7.

DeBlasi:2022:ABN

[3335] Pierpaolo De Blasi, Ramsés H. Mena, and Igor Prünster. Asymptotic behavior of the number of distinct values in a sample from the geometric stickbreaking process. Annals of the Institute of Statistical Mathematics, 74 (1):143–165, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00791-6.

Werner:2022:ALE

[3336] Tino Werner. Asymptotic linear expansion of regularized Mestimators. Annals of the Institute of Statistical Mathematics, 74(1): 167–194, February 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00792-5.

Dette:2022:DRD

[3337] Holger Dette and Kevin Kokot. Detecting relevant differences in the covariance operators of functional time series: a sup-norm approach. Annals of the Institute of Statistical Mathematics, 74(2):195–231, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00795-2.

Fasen-Hartmann:2022:WEC

[3338] Vicky Fasen-Hartmann and Celeste Mayer. Whittle estimation for continuous-time stationary state space models with finite second moments. Annals of the Institute of Statistical Mathematics, 74 (2):233–270, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00802-6.

Jiang:2022:SMA

[3339] Bochuan Jiang, Yaping Wang, and Mingyao Ai. Search for minimum aberration designs with uniformity. Annals of the Institute of Statistical Mathematics, 74(2): 271–287, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-021-00796-1.

Hoang:2022:URP

[3340] Anh-Tuan Hoang and Thorsten Dickhaus. On the usage of randomized p-values in the Schweder–Spjøtvoll estimator. Annals of the Institute of Statistical Mathematics, 74 (2):289–319, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00797-0.

Feng:2022:VSF

[3341] Sanying Feng, Menghan Zhang, and Tiejun Tong. Variable selection for functional linear models with strong heredity constraint. Annals of the Institute of Statistical Mathematics, 74(2):321–339, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00798-z.

Adam:2022:LPE

[3342] C. Adam and I. Gijbels. Local polynomial expectile regression. Annals of the Institute of Statistical Mathematics, 74(2):341–378, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00799-y.

Zhang:2022:SVB

[3343] Jing Zhang, Qihua Wang, and Xuan Wang. Surrogate-variable-based model-free feature screening for survival data under the general censoring mechanism. Annals of the Institute of Statistical Mathematics, 74 (2):379–397, April 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00801-7.

Nakashima:2022:WWE

[3344] Hideto Nakashima and Piotr Graczyk.

Wigner and Wishart ensembles for sparse Vinberg models. Annals of the Institute of Statistical Mathematics, 74(3):399–433, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00800-8.

Luan:2022:RDE

[3345] Jiaming Luan, Hongwei Wang, and Benle Zhang. Robust distributed estimation and variable selection for massive datasets via rank regression. Annals of the Institute of Statistical Mathematics, 74 (3):435–450, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00803-5.

Yuan:2022:SIG

[3346] Meng Yuan, Chunlin Wang, and Pengfei Li. Semiparametric inference on general functionals of two semicontinuous populations. Annals of the Institute of Statistical Mathematics, 74(3):451–472, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00804-4.

Hu:2022:CND

[3347] Chin-Yuan Hu and Gwo Dong Lin. Characterizations of the normal distribution via the independence of the sample mean and the feasible definite statistics with ordered arguments. Annals of the Institute of Statistical Mathematics, 74 (3):473–488, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00805-3.

Jordan:2022:COS

[3348] Alexander I. Jordan, Anja Mühlemann, and Johanna F. Ziegel. Characterizing the optimal solutions to the isotonic regression problem for identifiable functionals. Annals of the Institute of Statistical Mathematics, 74 (3):489–514, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00808-0.

Fujimori:2022:VSD

[3349] Kou Fujimori. The variable selection by the Dantzig selector for Cox’s proportional hazards model. Annals of the Institute of Statistical Mathematics, 74(3):515–537, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00807-1.

Liang:2022:RMS

[3350] Zhongqi Liang, Qihua Wang, and Yuting Wei. Robust model selection with covariables missing at random. Annals of the Institute of Statistical Mathematics, 74(3):539–557, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00806-2.

Luo:2022:HDE

[3351] Bin Luo and Xiaoli Gao. A highdimensional M-estimator framework for bi-level variable selection. Annals of the Institute of Statistical Mathematics, 74(3):559–579, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00809-z.

Bhattacharya:2022:BFA

[3352] Minerva Mukhopadhyay and Sourabh Bhattacharya. Bayes factor asymptotics for variable selection in the Gaussian process framework. Annals of the Institute of Statistical Mathematics, 74(3):581–613, June 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL thtps:

//link.springer.com/article/10.
1007/s10463-021-00810-6.

Copas:2022:AML

[3353] John Copas. Akaike Memorial Lecture 2020: Some of the challenges of statistical applications. Annals of the Institute of Statistical Mathematics, 74 (4):615–637, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00831-9. See dis-

cussion [3354, 3355] and rejoinder [3356].

Henmi:2022:DAM

[3354] Masayuki Henmi. Discussion of Akaike Memorial Lecture 2020: Some of the challenges of statistical applications. Annals of the Institute of Statistical Mathematics, 74 (4):639–641, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00833-7. See [3353,

3355].

Taguri:2022:DAM

[3355] Masataka Taguri. Discussion of “Akaike Memorial Lecture 2020: Some of the challenges of statistical applications”. Annals of the Institute of Statistical Mathematics, 74 (4):643–647, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00829-3. See [3353,

3355].

Copas:2022:ARD

[3356] John Copas. Author’s rejoinder to the discussion of the Akaike Memorial Lecture 2020. Annals of the Institute of Statistical Mathematics, 74 (4):649–652, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00830-w. See [3353]

and discussion [3354, 3355].

Cho:2022:TSD

[3357] Haeran Cho and Claudia Kirch. Twostage data segmentation permitting multiscale change points, heavy tails and dependence. Annals of the Institute of Statistical Mathematics, 74 (4):653–684, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00811-5.

Konev:2022:FAE

[3358] Victor V. Konev and Sergey E. Vorobeychikov. Fixed accuracy estimation of parameters in a threshold autoregressive model. Annals of the Institute of Statistical Mathematics, 74 (4):685–711, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00812-4.

Gribkova:2022:ETC

[3359] N. V. Gribkova, J. Su, and R. Zitikis. Empirical tail conditional allocation and its consistency under minimal assumptions. Annals of the Institute of Statistical Mathematics, 74 (4):713–735, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00813-3.

Bouzebda:2022:AFD

[3360] Salim Bouzebda, Mohamed Chaouch, and Sultana Didi Biha. Asymptotics for function derivatives estimators based on stationary and ergodic discrete time processes. Annals of the Institute of Statistical Mathematics, 74 (4):737–771, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00814-2.

Proksch:2022:SIB

[3361] Katharina Proksch, Nicolai Bissantz, and Hajo Holzmann. Simultaneous inference for Berkson errors-in-variables regression under fixed design. Annals of the Institute of Statistical Mathematics, 74(4):773–800, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00817-z.

Kengne:2022:INT

[3362] William Kengne and Isidore S. Ngongo.

Inference for nonstationary time series of counts with application to changepoint problems. Annals of the Institute of Statistical Mathematics, 74 (4):801–835, August 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00815-1.

Bakoyannis:2022:NTM

[3363] Giorgos Bakoyannis and Dipankar Bandyopadhyay. Nonparametric tests for multistate processes with clustered data. Annals of the Institute of Statistical Mathematics, 74 (5):837–867, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00819-x.

Di:2022:MRS

[3364] Fengrui Di and Lei Wang. Multi-round smoothed composite quantile regression for distributed data. Annals of the Institute of Statistical Mathematics, 74 (5):869–893, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00816-0.

Akaoka:2022:BEM

[3365] Yuichi Akaoka, Kazuki Okamura, and Yoshiki Otobe. Bahadur efficiency of the maximum likelihood estimator and one-step estimator for quasi-arithmetic means of the Cauchy distribution. Annals of the Institute of Statistical Mathematics, 74 (5):895–923, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-021-00818-y.

Barbu:2022:AEE

[3366] Vlad Stefan Barbu, Slim Beltaief, and Serguei Pergamenchtchikov. Adaptive efficient estimation for generalized semi-Markov big data models. Annals of the Institute of Statistical Mathematics, 74(5):925– 955, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00820-y.

Wang:2022:DHR

[3367] Tiandong Wang and Panpan Zhang. Directed hybrid random networks mixing preferential attachment with uniform attachment mechanisms. Annals of the Institute of Statistical Mathematics, 74(5):957–986, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL https://link.springer.com/
article/10.1007/s10463-022-00827-
5.

Li:2022:ORB

[3368] Lu Li, Niwen Zhou, and Lixing Zhu. Outcome regression-based estimation of conditional average treatment effect. Annals of the Institute of Statistical Mathematics, 74(5): 987–1041, October 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00821-x.

Xiao:2022:BNA

[3369] Bofei Xiao, Bo Lei, and Bin Guo.

A blockwise network autoregressive model with application for fraud detection. Annals of the Institute of Statistical Mathematics, 74(6):1043– 1065, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00822-w.

ElBarmi:2022:CCR

[3370] Hammou El Barmi. On comparing competing risks using the ratio of their cumulative incidence functions. Annals of the Institute of Statistical Mathematics, 74(6):1067– 1083, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00823-9.

Kohler:2022:RCI

[3371] Michael Kohler, Adam Krzyzak, and Benjamin Walter. On the rate of convergence of image classifiers based on convolutional neural networks. Annals of the Institute of Statistical Mathematics, 74(6):1085– 1108, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00828-4.

Yu:2022:SFS

[3372] Ke Yu and Shan Luo. A sequential feature selection procedure for highdimensional Cox proportional hazards model. Annals of the Institute of Statistical Mathematics, 74(6):

1109–1142, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00824-8.

Chang:2022:IRE

[3373] Chih-Hao Chang, Hsin-Cheng Huang, and Ching-Kang Ing. Inference of random effects for linear mixed-effects models with a fixed number of clusters. Annals of the Institute of Statistical Mathematics, 74(6):1143– 1161, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00825-7.

Dette:2022:AEN

[3374] Holger Dette and Martin Kroll. Asymptotic equivalence for nonparametric regression with dependent errors: Gauss–Markov processes. Annals of the Institute of Statistical Mathematics, 74(6):1163– 1196, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00826-6.

Tsukurimichi:2022:CSI

[3375] Toshiaki Tsukurimichi, Yu Inatsu, and Ichiro Takeuchi. Conditional selective inference for robust regression and outlier detection using piecewise-linear homotopy continuation. Annals of the Institute of Statistical Mathematics, 74(6):1197– 1228, December 2022. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00846-2.

Ni:2023:EMO

[3376] Lyu Ni and Jun Shao. Estimation with multivariate outcomes having nonignorable item nonresponse. Annals of the Institute of Statistical Mathematics, 75(1):1–15, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00836-4.

Guo:2023:JBP

[3377] Jinhui Guo and Yingyin Lu. Joint behavior of point processes of clusters and partial sums for stationary bivariate Gaussian triangular arrays. Annals of the Institute of Statistical Mathematics, 75(1):17–37, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00832-8.

Wu:2023:SMT

[3378] Jingjing Wu, Tasnima Abedin, and Qiang Zhao. Semiparametric modelling of two-component mixtures with stochastic dominance. Annals of the Institute of Statistical Mathematics, 75 (1):39–70, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00835-5.

Chatla:2023:NIA

[3379] Suneel Babu Chatla. Nonparametric inference for additive models estimated via simplified smooth backfitting. Annals of the Institute of Statistical Mathematics, 75 (1):71–97, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00840-8.

Terada:2023:SIA

[3380] Yoshikazu Terada and Hidetoshi Shimodaira. Selective inference after feature selection via multiscale bootstrap. Annals of the Institute of Statistical Mathematics, 75 (1):99–125, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00838-2.

Duy:2023:ESI

[3381] Vo Nguyen Le Duy and Ichiro Takeuchi. Exact statistical inference for the Wasserstein distance by selective inference. Annals of the Institute of Statistical Mathematics, 75 (1):127–157, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00837-3.

Baillien:2023:FAM

[3382] Jonas Baillien, Irène Gijbels, and Anneleen Verhasselt. Flexible asymmetric multivariate distributions based on two-piece univariate distributions. Annals of the Institute of Statistical Mathematics, 75(1): 159–200, February 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00842-6.

Goegebeur:2023:REC

[3383] Yuri Goegebeur, Armelle Guillou, and Jing Qin. Robust estimation of the conditional stable tail dependence function. Annals of the Institute of Statistical Mathematics, 75 (2):201–231, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00839-1.

Tsao:2023:GLS

[3384] Min Tsao. Group least squares regression for linear models with strongly correlated predictor variables. Annals of the Institute of Statistical Mathematics, 75(2):233–250, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00841-7.

Tsao:2023:CGL

[3385] Min Tsao. Correction to: Group least squares regression for linear models with strongly correlated predictor variables. Annals of the Institute of Statistical Mathematics, 75(2):251, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00861-3.

Shahzadi:2023:IHS

[3386] Amina Shahzadi, Ting Wang, Mark Bebbington, and Matthew Parry. Inhomogeneous hidden semi-Markov models for incompletely observed point processes. Annals of the Institute of Statistical Mathematics, 75 (2):253–280, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00843-5.

Hanada:2023:IUE

[3387] Keisuke Hanada and Tomoyuki Sugimoto. Inference using an exact distribution of test statistic for randomeffects meta-analysis. Annals of the Institute of Statistical Mathematics, 75(2):281–302, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00844-4.

Bieniek:2023:COS

[3388] Mariusz Bieniek and Luiza Pa’nczyk. On the choice of the optimal single order statistic in quantile estimation. Annals of the Institute of Statistical Mathematics, 75 (2):303–333, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00845-3.

Werner:2023:QRI

[3389] Tino Werner. Quantitative robustness of instance ranking problems. Annals of the Institute of Statistical Mathematics, 75(2):335–368, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00847-1.

Wang:2023:ATN

[3390] Qiuping Wang, Yuan Zhang, and Ting Yan. Asymptotic theory in network models with covariates and a growing number of node parameters. Annals of the Institute of Statistical Mathematics, 75(2):369–392, April 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00848-0.

Honda:2023:FVS

[3391] Toshio Honda and Chien-Tong Lin.

Forward variable selection for ultrahigh dimensional quantile regression models. Annals of the Institute of Statistical Mathematics, 75 (3):393–424, June 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00849-z.

Sutradhar:2023:RAE

[3392] Brajendra C. Sutradhar. Regression analysis for exponential family data in a finite population setup using twostage cluster sample. Annals of the Institute of Statistical Mathematics, 75(3):425–462, June 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00850-6.

Mao:2023:MCU

[3393] Xiaojun Mao, Zhonglei Wang, and Shu Yang. Matrix completion under complex survey sampling. Annals of the Institute of Statistical Mathematics, 75(3):463–492, June 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00851-5.

Li:2023:REN

[3394] Shaomin Li, Kangning Wang, and Yong Xu. Robust estimation for nonrandomly distributed data. Annals of the Institute of Statistical Mathematics, 75(3):493–509, June 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00852-4.

Goto:2023:TEG

[3395] Yuichi Goto, Kotone Suzuki, Xiaofei Xu, and Masanobu Taniguchi. Tests for the existence of group effects and interactions for two-way models with dependent errors. Annals of the Institute of Statistical Mathematics, 75 (3):511–532, June 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-022-00853-3.

Uehara:2023:BMM

[3396] Yuma Uehara. Bootstrap method for misspecified ergodic Lévy driven stochastic differential equation models. Annals of the Institute of Statistical Mathematics, 75(4):533–565, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10463-022-00854-
                                        2.

Kamila:2023:DDM

[3397] Kare Kamila. Data-driven model selection for same-realization predictions in autoregressive processes. Annals of the Institute of Statistical Mathematics, 75(4):567–592, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00855-1.

Arendarczyk:2023:SDG

[3398] M. Arendarczyk, T. J. Kozubowski, and A. K. Panorska. Slash distributions, generalized convolutions, and extremes. Annals of the Institute of Statistical Mathematics, 75 (4):593–617, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00858-y.

Wu:2023:UPM

[3399] Zeyu Wu, Cheng Wang, and Weidong Liu. A unified precision matrix estimation framework via sparse column-wise inverse operator under weak sparsity. Annals of the Institute of Statistical Mathematics, 75 (4):619–648, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00856-0.

Zhan:2023:MAS

[3400] Zishu Zhan, Yang Li, Yuhong Yang, and Cunjie Lin. Model averaging for semiparametric varying coefficient quantile regression models. Annals of the Institute of Statistical Mathematics, 75(4):649–681, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00857-z.

Pesce:2023:GAR

[3401] Elena Pesce, Fabio Rapallo, Eva Riccomagno, and Henry P. Wynn.

Generation of all randomizations using circuits. Annals of the Institute of Statistical Mathematics, 75 (4):683–704, August 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00860-4.

Zhang:2023:CST

[3402] Shibin Zhang. A copula spectral test for pairwise time reversibility. Annals of the Institute of Statistical Mathematics, 75 (5):705–729, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00859-x.

Dussap:2023:NMR

[3403] Florian Dussap. Nonparametric multiple regression by projection on noncompactly supported bases. Annals of the Institute of Statistical Mathematics, 75(5):731–771, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00863-1.

Mandal:2023:RDP

[3404] Abhijit Mandal, Beste Hamiye Beyaztas, and Soutir Bandyopadhyay. Robust density power divergence estimates for panel data models. Annals of the Institute of Statistical Mathematics, 75(5):773–798, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00862-2.

Ma:2023:LAD

[3405] Nannan Ma, Hailin Sang, and Guangyu Yang. Least absolute deviation estimation for AR(1) processes with roots close to unity. Annals of the Institute of Statistical Mathematics, 75 (5):799–832, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-022-00864-0.

Li:2023:MSB

[3406] Shaoting Li and Jiahua Chen. Mixture of shifted binomial distributions for rating data. Annals of the Institute of Statistical Mathematics, 75 (5):833–853, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00865-7.

Tsuruta:2023:ADB

[3407] Yasuhito Tsuruta and Masahiko Sagae.

Automatic data-based bin width selection for rose diagram. Annals of the Institute of Statistical Mathematics, 75 (5):855–886, October 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00868-4.

Kresin:2023:PES

[3408] Conor Kresin and Frederic Schoenberg. Parametric estimation of spatial-temporal point processes using the Stoyan–Grabarnik statistic. Annals of the Institute of Statistical Mathematics, 75(6):887– 909, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00866-6.

Wang:2023:SIU

[3409] Hengfang Wang and Jae Kwang Kim. Statistical inference using regularized m-estimation in the reproducing kernel Hilbert space for handling missing data. Annals of the Institute of Statistical Mathematics, 75(6): 911–929, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00872-8.

Fang:2023:GEI

[3410] Kuangnan Fang, Jingmao Li, Yaqing Xu, Shuangge Ma, and Qingzhao Zhang. Gene-environment interaction analysis under the Cox model. Annals of the Institute of Statistical Mathematics, 75(6):931–948, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic).

URL https://link.springer.com/
article/10.1007/s10463-023-00871-
9.

Wang:2023:RVS

[3411] Xiuli Wang, Jingchang Shao, Jingjing Wu, and Qiang Zhao. Robust variable selection with exponential squared loss for partially linear spatial autoregressive models. Annals of the Institute of Statistical Mathematics, 75 (6):949–977, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00870-w.

Watanabe:2023:GFT

[3412] Chihiro Watanabe and Taiji Suzuki.

A goodness-of-fit test on the number of biclusters in a relational data matrix. Annals of the Institute of Statistical Mathematics, 75(6):979– 1009, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00869-3.

Ebner:2023:GFT

[3413] Bruno Ebner, Adrian Fischer, Norbert Henze, and Celeste Mayer. Goodnessof-fit tests for the Weibull distribution based on the Laplace transform and Stein’s method. Annals of the Institute of Statistical Mathematics, 75(6): 1011–1038, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00873-7.

Ma:2023:ECC

[3414] Yuqing Ma, Peijie Wang, and Jianguo Sun. Estimation of complier causal treatment effects with informatively interval-censored failure time data. Annals of the Institute of Statistical Mathematics, 75(6):1039– 1062, December 2023. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00874-6.

Hyvarinen:2024:ILV

[3415] Aapo Hyvärinen, Ilyes Khemakhem, and Ricardo Monti. Identifiability of latent-variable and structuralequation models: from linear to nonlinear. Annals of the Institute of Statistical Mathematics, 76 (1):1–33, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00884-4. See dis-

cussion [3416, 3417, ?] and rejoinder [3418].

Morioka:2024:DIL

[3416] Hiroshi Morioka. Discussion of “Identifiability of latent-variable and structural-equation models: from linear to nonlinear”. Annals of the Institute of Statistical Mathematics, 76 (1):35–37, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00886-2. See [3415,

3418].

Matsuda:2024:DIL

[3417] Takeru Matsuda. Discussion of “Identifiability of latent-variable and structural-equation models: from linear to nonlinear”. Annals of the Institute of Statistical Mathematics, 76 (1):39–42, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00885-3. See [3415,

3418].

Hyvarinen:2024:RIL

[3418] Aapo Hyvärinen. Rejoinder of “Identifiability of latent-variable and structural-equation models: from linear to nonlinear”. Annals of the Institute of Statistical Mathematics, 76 (1):43–46, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00887-1.

Brehmer:2024:CEP

[3419] Jonas R. Brehmer, Tilmann Gneiting, Marcus Herrmann, Warner Marzocchi, Martin Schlather, and Kirstin Strokorb. Comparative evaluation of point process forecasts. Annals of the Institute of Statistical Mathematics, 76 (1):47–71, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00875-5.

Zhao:2024:MAE

[3420] Zhihao Zhao, Xinyu Zhang, Guohua Zou, Alan T. K. Wan, and Geoffrey K. F. Tso. Model averaging for estimating treatment effects. Annals of the Institute of Statistical Mathematics, 76(1):73–92, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00876-4.

Xu:2024:TFE

[3421] Kai Xu and Nan An. A tuningfree efficient test for marginal linear effects in high-dimensional quantile regression. Annals of the Institute of Statistical Mathematics, 76 (1):93–110, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00877-3.

Eguchi:2024:GQI

[3422] Shoichi Eguchi and Hiroki Masuda. Gaussian quasi-information criteria for ergodic Lévy driven SDE. Annals of the Institute of Statistical Mathematics, 76 (1):111–157, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-023-00878-2.

Bastian:2024:CRC

[3423] Patrick Bastian, Holger Dette, Lukas Koletzko, and Kathrin Möllenhoff. Comparing regression curves: an L^{1}- point of view. Annals of the Institute of Statistical Mathematics, 76(1): 159–183, February 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-023-00880-8.

Dumbgen:2024:ASE

[3424] Lutz Dümbgen and Klaus Nordhausen. Approximating symmetrized estimators of scatter via balanced incomplete U-statistics. Annals of the Institute of Statistical Mathematics, 76 (2):185–207, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-023-00879-1.

Halconruy:2024:PLS

[3425] Hélène Halconruy and Nicolas Marie. On a projection least squares estimator for jump diffusion processes. Annals of the Institute of Statistical Mathematics, 76(2):209–234, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00881-7.

Zheng:2024:ENR

[3426] Qi Zheng, Yunwei Cui, and Rongning Wu. On estimation of nonparametric regression models with autoregressive and moving average errors. Annals of the Institute of Statistical Mathematics, 76(2):235–262, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00882-6.

Carbon:2024:MFP

[3427] Michel Carbon and Thierry Duchesne.

Multivariate frequency polygon for stationary random fields. Annals of the Institute of Statistical Mathematics, 76(2):263–287, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00883-5.

Paindaveine:2024:UHT

[3428] Davy Paindaveine. On UMPS hypothesis testing. Annals of the Institute of Statistical Mathematics, 76 (2):289–312, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00888-0.

vanLieshout:2024:NPA

[3429] M. N. M. van Lieshout. Nonparametric adaptive bandwidth selection for kernel estimators of spatial intensity functions. Annals of the Institute of Statistical Mathematics, 76 (2):313–331, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00890-6.

Lee:2024:TCQ

[3430] Sangyeol Lee and Chang Kyeom Kim. Test for conditional quantile change in general conditional heteroscedastic time series models. Annals of the Institute of Statistical Mathematics, 76(2):333–359, April 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00889-z.

Drews:2024:UCP

[3431] Selina Drews and Michael Kohler. On the universal consistency of an over-parametrized deep neural network estimate learned by gradient descent. Annals of the Institute of Statistical Mathematics, 76 (3):361–391, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-024-00898-6.

Kirch:2024:DST

[3432] Claudia Kirch and Kerstin Reckruehm. Data segmentation for time series based on a general moving sum approach. Annals of the Institute of Statistical Mathematics, 76 (3):393–421, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00892-4.

Quessy:2024:GCP

[3433] Jean-François Quessy. Gradual change-point analysis based on Spearman matrices for multivariate time series. Annals of the Institute of Statistical Mathematics, 76(3): 423–446, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00891-5.

Adamczyk-Chauvat:2024:SIR

[3434] Katarzyna Adamczyk-Chauvat, Mouna Kassa, Julien Papaïx, Kiên Kiêu, and Radu S. Stoica. Statistical inference for random t-tessellations models. application to agricultural landscape modeling. Annals of the Institute of Statistical Mathematics, 76 (3):447–479, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00893-3.

You:2024:RNR

[3435] Hojun You, Kyubaek Yoon, Wei-Ying Wu, Jongeun Choi, and Chae Young Lim. Regularized nonlinear regression with dependent errors and its application to a biomechanical model. Annals of the Institute of Statistical Mathematics, 76(3):481–510, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

//link.springer.com/article/10.
1007/s10463-023-00895-1.

vonRosen:2024:UGC

[3436] Dietrich von Rosen and Martin Singull. Using the growth curve model in classification of repeated measurements. Annals of the Institute of Statistical Mathematics, 76 (3):511–534, June 2024. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10463-024-00900-1.